A Complete Bibliography of Publications in Marine Biotechnology

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

29 August 2024 Version 1.24

Title word cross-reference

17β [673]. 1β [392]. 3β [1713]. ^{+} [805, 327, 607]. ^{2+} [859, 957]. ^{31} [30]. ^{Cip1} [780]. ^{T} [945, 1363]. _{1} [713]. _{12} [1337]. _{2} [1098, 497, 1806, 1337, 646, 1615, 1529, 1144, 1199, 1663, 1356, 1196, 1649]. _{3} [913, 1196, 594]. _{4} [950]. _{n} [578, 104]. α [845, 34, 1860, 1288, 436, 27, 1030, 954, 1043, 230, 28, 95, 886, 1285, 1024, 620, 576, 978, 1638, 1792, 679]. α_{1} [6]. β [520, 359, 1384, 511, 1329, 1057, 1145, 668, 1339, 1827, 1941, 721, 1381, 954, 93, 1630, 840, 349, 879, 924, 1306, 1325, 1021, 109, 1044, 133, 1025]. d [1021, 1025]. Δ [611, 795]. Δ5 [905, 1058, 1665, 1132]. Δ6 [1132, 906]. Δ6Δ5 [1482]. Δ9 [1132]. γ [1482, 1591]. ι [945]. κ [911, 1306, 1510]. l [1150]. λ [1919, 217, 179]. μ [992]. ω [497, 1383]. π [1456]. ′ [1227, 94, 905, 863, 1752, 99, 1015, 577, 92, 1539, 1518]. × [852, 1274, 1493, 1492, 1632].

-/ [1306]. -1 [359, 1630, 879, 924, 1021, 1044, 1025]. -2 [863].

-2-Methyl-3-Phenylacrylaldehyde [1579]. -3 [497]. -6 [1461]. -7- [863]. -Actin [93, 349, 1325]. -Agarase [840, 1827]. -Ambrox(R) [1240]. -Amylase [1043, 230, 679]. -ATPase [859, 327, 291]. -Azido [1285]. -Bromoacetophenones [886]. -Carotene [511]. -Carrageenan [945, 1306]. -Carrageenase [945]. -Carrageenin [1919]. -Concentrating [1663]. -D-glucans [721]. -d-Glucosidase [978]. -Dependent [1098, 957]. -Desaturase [906]. -dihydroxyphenoxy [863]. -Discodermolide [1140]. -Enolase [1792]. -Estradiol [1381, 673]. -Fructofuranosidase [1339]. -Galactosidase [845]. -Glucan [133]. -glucan-encapsulated [1025]. -Glucosidase [576]. -Glucosidases [1329]. -glucoside [28]. -Glucosylation [576]. -GO [1806]. -Haloacid [1150]. -Hydroxy [1384]. -Hydroxybutyrate [1941]. -Hydroxychalcones [1518]. -Hydroxysteroid [109]. -Induced [950]. -K [327, 291]. -like [1058]. -Linolenic [436]. -monophosphate [1015]. -one-Containing [1259]. -Polyglucans [620]. -producing [713]. -Proteinase [6]. -Sclareol [1240]. -Sclareolide [1240]. -Tailed [577]. -Upstream [905]. -Xylose [1021]. -Xylulose [1021].

. [1715].

[?]C18Ohara:2013:TFA. [?]C18-Ohara:2013:TFA. [?]C20Ohara:2013:TFA. [?]C20-Ohara:2013:TFA. /NDP [19]. /srsf3 [1952].

0.2- [992]. 008041 [801]. 01 [826].

1 [498, 219, 1417, 452, 34, 1860, 1091, 346, 319, 95, 9, 220, 1199, 1712]. 1- [863, 1053]. 1-Like [9]. 10 [905, 483]. 10a [1836]. 11-1 [1487]. 11.5 [280]. 11.5-kDa [280]. 1186 [1579]. 122 [1651]. 1237 [1576]. 13a [1766]. 14 [841]. 14-3-3E1 [841]. 15 [1198]. 15/16 [1591]. 15041c [1419]. 16 [1591]. 16s [496, 528, 440, 462, 351, 490, 229, 1008, 246, 732]. 18S [528, 77, 370, 490]. 18W [1766]. 18W-13a [1766]. 19 [1835, 1638]. 1996 [1609]. 1A [94, 45, 43]. 1A1 [48, 47]. 1st [1007].

2 [911, 460, 879, 1433, 1717, 1044, 145]. 2-40 [1296]. 2-Cyano-3-Phenylpropanamide [1576]. 2-Hydroxyphenylacetic [1101]. 2-Nitroimidazole [258]. 2/4 [1694]. 20 [1799]. 208029 [1843]. 20K [975]. 214 [1104, 1021]. 221 [1298]. 222 [1298]. 2316 [1545]. 26185 [1132]. 263 [1601]. 28209 [1726]. 2D [1078, 1932]. 2D-DIGE [1078]. 2” [863].

3 [1289, 1739, 761, 1037, 1766]. 3- [1025]. 3-Desaturase [1383]. 3-Glucanase [359, 879, 1044, 924]. 3-Hydroxy-3-Methylglutaryl-CoA [1766]. 3-Methylcholanthrene [43]. 3-Methylcrotonyl [443]. 3-Xylan [1021]. 3/ [1461]. 30.7 [590]. 30.7-kDa [590]. 30543 [1654]. 310 [1725]. 32 [1321]. 33 [1522]. 36B4 [743]. 38 [1827]. 382 [1770]. 3D [1823]. 3Dp86E [1199]. 3Dp86E-1 [1199]. 3E1 [841]. 3p [1836]. 3T3 [770, 812, 863]. 3T3-L1 [770, 812, 863]. 3Y1 [42, 314].

4 [779]. 4-Deoxy-L-Etychro-Hexoseulose [1442]. 4-dioxin [863]. 4-Glucuronosyltransferase [1630]. 4-Hydroxylase [1258]. 4-Methoxyphenyl [1053]. 4-tert-Octylphenol [673]. 40 [1006]. 4635 [1896]. 4” [863]. 4S [1116]. 4SNc [527]. 4SNc-Tudor [527].

5 [611, 754, 795, 1572, 1808]. 5-Aminolevulinic [1479]. 5-Desaturase-Like [611]. 5-Diketopiperazines [1189]. 5-fatty [795]. 5.8S [381]. 520P1 [1354]. 5n [414]. 5n-3 [414].

6 [947]. 6- [721]. 6-Hour-Old [228]. 6-Pyruvoyl [382]. 6-trihydroxyphenoxy [863]. 60 [235]. 60-kDa [235]. 60S [745]. 67 [1292]. 6n [414]. 6n-3 [414].

7422 [612].

9 [1775, 1601]. 9-trihydroxydibenzo-1 [863]. 934 [1169].

A. [510]. a1 [291]. a3 [291]. A94 [945]. Aaptamine [567]. Aaptamines [1893]. Aaptos [1893]. Abalone [543, 35, 609, 816, 1519, 1299, 375, 357, 526, 570, 300, 555, 524, 587, 140, 577, 232, 1223, 1370, 736, 450, 1632, 960, 1446, 150, 1014, 496]. ABCG4 [1393]. Aberrant [554, 474]. Ability [1211, 736]. Abiotic [1213, 1373, 1459, 853]. Ablation [878]. Abnormal [1458]. Abnormality [240]. Absence [1703]. Absorbing [342]. Abundance [1175, 1084, 1054, 1500]. Abundant [1698, 1264]. Acanthopagrus [605, 276, 1759, 1903]. Acartia [985]. Accelerate [1196]. Accelerated [2]. Access [1126]. Accessibility [1769]. Acclimated [1676, 1632]. Acclimation [1680, 598, 1283, 943, 1684]. Accumulating [1924]. Accumulation [737, 1297, 1390, 314, 1950, 1139, 1911, 1924]. Accumulative [1824]. ACE [1583]. Acetate [1699]. Acetyl [1457]. Acetylcholine [816]. Acetylcholine-Binding [816]. Acetylcholinesterase [1230]. Acetylglucosamine [82]. Acetylome [1571]. Acid [414, 1227, 1358, 269, 934, 757, 1372, 941, 1295, 436, 370, 905, 1748, 457, 659, 955, 549, 1442, 227, 1457, 1094, 1603, 1171, 1195, 1858, 819, 562, 748, 1132, 1479, 892, 542, 795, 311, 983, 956, 791, 1356, 1537, 1747, 804, 1065, 851, 1911, 943, 1101, 497, 753]. Acid-Containing [542]. Acid-Directed [934]. Acid-Producing [370]. Acidic [523, 1954, 838, 1433, 220]. Acidification [1881, 1809, 1356, 1662, 1716, 1494, 1847]. Acids [414, 1374, 1760, 436, 1225, 1141, 1058, 687, 1207, 1231, 616, 983, 956, 1808, 1155]. ACL1 [1094]. Acotylean [1621]. Acquisition [1172]. Acrochaetium [1778]. Acropora [768, 1506, 1657, 765, 1617]. Acrosome [1249]. Across [1503]. Actimineral [950]. Actin [543, 35, 93, 349, 1325]. acting [1646]. Actiniaria [1569].

Actiniidae [1569]. Actinobacteria [1256]. Actinobacterial [729]. Actinomycetal [732]. Actinomycetes [1045, 1127]. Action [58, 391, 1953, 779, 1912]. Activate [1380]. Activated [36, 561]. Activating [947]. Activation [911, 273, 253]. Active [1421, 1465, 1232, 475, 798, 361, 782]. Activin [1958]. Activities [15, 809, 1530, 394, 711, 793, 1923, 10, 391, 687, 367, 772, 1006, 990, 522, 297]. Activity [1175, 520, 799, 1241, 1449, 764, 1491, 567, 41, 354, 1320, 1002, 876, 1475, 1333, 1307, 1313, 1143, 590, 1410, 1339, 1049, 1289, 1941, 949, 1700, 253, 424, 1631, 778, 1138, 1756, 1936, 657, 1520, 319, 917, 1665, 925, 1132, 670, 1273, 879, 839, 1229, 1054, 420, 674, 994, 1284, 1245, 1695, 634, 1217, 1386, 38, 26, 855, 1044, 1293, 482, 612, 948, 492]. aculeatus [1041]. Acute [1910, 1938, 1892, 1773, 1961, 1402, 1906, 1849, 1523]. Acyl [436, 1482, 1058, 1040]. Acylhomoserine [1354]. Adaptation [270, 1570, 1348, 1750, 1914, 1533, 1532, 1759, 1903, 1342, 505]. Adapted [291, 826, 986, 628, 1100, 738, 679, 1047]. Adaptive [1881, 1534, 912, 1414]. Addition [1322]. Adducts [1576]. Adenosine [1015]. Adhesion [660]. Adhesive [323, 119, 952, 683, 1686, 1216, 1484, 824]. Adipicola [672]. Adipocyte [812, 863, 1911]. Adipocytes [770]. Adipose [1110]. Adjuvant [1783]. Administration [1715, 1294, 1450, 698, 1102, 1388, 1917]. ADP [401]. Adriatic [434, 4, 266]. Adsorption [903, 1686]. Adult [1931, 251, 1485, 774, 95, 1484, 636, 225, 191]. Advanced [926]. Advances [1746]. Advantages [1719]. aeglefinius [596]. aeglefinus [295]. Aeluropus [1371]. Aequorea [265]. Aerial [1361]. Aerobic [440]. aerogenes [469]. Aeromonas [1715, 1608, 1012, 1185, 380, 1510, 1960, 1429, 1790, 998, 1795, 1879]. aerophoba [367]. aeruginosa [1454, 1755]. aff [1631, 241]. Affect [1925, 971]. Affected [699, 155, 1107, 802, 810, 714]. Affecting [517, 391, 1144, 1129, 1186, 1717, 1791]. Affects [1851, 1202, 1735, 1078]. Affinity [1333, 1307, 1313]. AFLP [709, 626, 433, 111, 263, 959, 390, 871, 664, 422, 447]. AFLP-Based [390]. AFMB [801]. AFMB-008041 [801]. African [555]. After [1820, 274, 1068, 1832, 1721, 1364, 24, 1283, 1758, 1077, 1880, 257, 1960, 1604, 895, 98]. Against [1292, 1886, 259, 1691, 1449, 757, 41, 1801, 1835, 1187, 590, 1413, 646, 1221, 1049, 1941, 1504, 1756, 1303, 1456, 1266, 319, 792, 1953, 1153, 1077, 773, 970, 1464, 1510, 1921, 862, 1950, 1488, 1524, 1350, 1139, 1025]. Agamaki [861]. Agar [980, 1451, 1810]. Agaran [1350]. Agarase [1406, 840, 1827]. Agarases [1692]. Agardh [1143]. Agarivorans [840, 883, 1264]. Agarolytic [1410, 1827, 1311]. Agassiz [944]. Age [1779, 227]. Age-Associated [1779]. Agent [313, 1488]. Agents [1691, 1191, 1087]. Agglutination [742]. Aggregate [682]. Aggregated [218]. Aggregation [1686, 363]. aggregatum [1726]. Aging [1779, 1543, 1347, 1842]. AHL [1505, 1217]. AHL-Degrading [1505]. AHPND [1705]. Aiptasia [401, 832]. Aiptasia-Symbiodinum [401]. Air [1288, 1341, 1634]. Air-Breathing [1341, 1634]. Air-Exposed [1288]. akaara [1734, 635]. Akoya [1156, 1554].

Akt [1713]. al [601]. Alanine [1301, 1024]. albacares [148]. albicostatus [1799]. albiflora [1550, 1769]. Albino [1825]. Albumin [337]. alburnus [1875]. albus [1732]. Alcivar [601]. Alcohol [980]. Alcyonarian [674]. Alcyonium [1137]. aldingensis [1350]. Alendronic [1748]. Alexandrium [1877, 195, 510, 448, 480, 730, 1436, 746]. ALG [754]. ALG-5 [754]. Alga [958, 1877, 1793, 1143, 234, 541, 1738, 584, 554, 143, 1239, 813, 342, 544, 977, 839, 1583, 1284, 849, 1109, 1177, 1599, 782, 1778]. Algae [269, 534, 394, 107, 1005, 1373, 501, 1410, 1409, 980, 1670, 620, 1090, 943]. Algal [893, 1641, 801, 1049, 1593, 326, 55]. Algal-Lytic [801]. Algicidal [1049]. Alginate [1294, 929, 579, 754, 1936, 818, 1122, 84, 1296, 883, 1388, 1524]. Alginolytic [1442]. alginolyticus [899]. Algoriphagus [1839]. Alive [1931]. Alkaline [1938, 1635, 1570, 1232, 1128, 632, 706, 761, 475, 1933, 656]. Alkaline-Tolerant [1635]. Alkalinity [1906, 1849]. Alkaloid [1463, 842]. Alkaloids [1191, 1259, 1656]. All-Female [1913, 1344, 1274]. All-Male [1119]. Allele [242, 592]. Alleles [526]. Allelic [360]. Allelically [1721]. Allergen [135]. Allergens [358, 748]. Allergy [358]. Alleviates [1713]. Allotetraploid [630, 1003]. Allotetraploids [1274]. Allozymes [37]. Along [20, 1012, 1072, 1022]. Alpha [1328, 1584]. Alpha-Amylases [1584]. AlphaMax^{T M} [1290]. alpinus [199]. Alterations [303, 1587, 736]. Altered [1445]. Alternaria [1321]. Alternations [1472]. Alternative [1820, 269, 395, 1832, 1348, 1402, 1956, 99, 1758, 1906, 1708, 1471]. Alteromonas [82]. Alters [1716]. altivelis [422]. Aluminum [50]. alvarezii [1333, 1307, 1313]. Alveolar [1434]. Alvinocaris [486]. alyA [1122]. AlyVG2 [84]. Amberjack [1695]. amblycephala [1274, 1493, 1492, 1567]. Ambrox(R) [1240]. Ameliorates [1763]. Amelioration [1431]. American [196, 1922, 256, 18, 885]. americana [771]. americanus [900, 1100]. amine [452]. Amino [1374, 1358, 748, 687, 1065]. Aminolevulinic [1479]. Aminopeptidase [574]. Aminotransferase [1301, 1024]. Ammonia [1680, 1421, 1084, 1137, 1523]. Ammonia-Oxidizing [1084, 1137]. Ammonia-Scavenging [1421]. Amoebic [714]. Among [605, 212, 412, 451, 1750, 290, 18, 1540, 1467, 663, 937, 1342]. Amorphous [816]. Ampelomyces [595]. AmphiCL [487]. Amphidromous [23]. Amphioxus [1756, 487]. Amphipod [1664]. Amphiprion [1873]. Amphitrite [211, 258, 489]. Amphiura [87]. Amplicon [1642]. Amplification [141, 91, 643, 1083, 684, 288, 448, 93, 1935, 1795, 1907]. Amplified [709, 196, 193, 608, 4]. Amur [1570]. amurensis [1915]. Amusium [633]. Amylase [986, 1043, 230, 360, 679]. Amylases [1584]. Amylopectin [620]. Amylose [620]. Amyotrophic [1485]. Anabaena [346]. Anabas [1890]. Anadromous [495, 1212]. Anaerobic [1050]. Analogs [567, 1937]. Analogue [281]. Analogues [58, 678, 670, 1800]. Analyses [966, 1768, 768, 219, 315, 263, 171, 955, 226, 1121, 1206, 1533, 1532, 1275, 1634, 1685, 172, 868, 333, 311, 1617, 1960, 714, 1719, 1618]. Analysis [1415, 527, 48, 430, 675, 1807, 1228, 658, 1312, 1331, 543, 1424, 1192, 196, 478, 1676, 1870, 1822, 1925, 1922, 1826, 662, 511, 783, 1282, 296, 1658, 716, 210, 335, 101, 1271, 440, 111, 861, 1038, 829, 1445, 1715, 434, 514, 1252, 1883, 1111, 1754, 1039, 1832, 107, 523, 617, 902, 1525, 1026, 370, 1115, 1348, 1341, 1889, 1469, 228, 1852, 1750, 256, 785, 1209, 1821, 1926, 517, 953, 1146, 599, 579, 163, 793, 1289, 96, 193, 309, 1841, 134, 1142, 584, 93, 1902, 1364, 1020, 1353, 1335, 1393, 1361, 1476, 1571, 1660, 1884, 1936, 508, 1311, 1163, 566, 630, 852, 1398, 1499, 208, 1863, 1266]. Analysis [1858, 745, 1122, 87, 202, 1815, 1120, 1338, 260, 854, 1133, 185, 992, 1597, 1640, 1092, 1309, 1397, 352, 697, 47, 1052, 853, 1258, 1546, 1159, 1758, 229, 704, 756, 1484, 1592, 1033, 906, 151, 577, 1324, 1428, 83, 1583, 1708, 1880, 1781, 676, 1296, 310, 1780, 1527, 1040, 1581, 1440, 4, 374, 1261, 746, 417, 277, 623, 776, 1201, 1426, 1510, 1586, 1668, 246, 1113, 1521, 1542, 1783, 1921, 1706, 1891, 405, 1298, 1412, 1515, 1897, 1908, 1034, 625, 1219, 732, 302, 447, 1377, 1446, 1846, 306, 1254, 1427, 907, 546, 593, 655, 942, 920, 1157, 1549, 1582, 1612, 1790, 1776, 1849, 1946, 1086, 1342]. Analysis [1648, 1782, 1879, 1895, 496, 494, 305, 1552]. Analyzed [455]. Anatomically [1443]. Anchored [656]. Anchovy [156, 302]. andersoni [1436]. andrewsi [241]. Androgenesis [418]. Androgenic [1661, 1344]. Anemia [927]. Anemone [1193, 235, 645, 1569, 1120, 1574]. Anemones [531]. Anemonia [235]. Angiotensin [1270, 1929, 1778]. Angiotensin-Converting [1929]. Angiotensin-I [1778]. Anguilla [1922, 1832, 1661, 1703, 45, 202, 209, 446, 1783]. anguillarum [1528, 1252, 1832, 1165, 1153, 1077, 1705, 1201, 1783, 747, 592, 998]. anguillicaudatus [1341]. angulata [228, 1821, 937]. Anhydrase [1001, 258, 1328, 1392, 674, 1356]. Animal [1625, 1670, 1920, 1244]. Animals [783, 934, 1209]. Anion [1417]. Annelida [250]. Annotation [785, 1787]. Announcements [112]. annularis [13]. anophagefferens [143]. Anoxia [1757]. Answers [922, 472]. Antagonism [792, 677]. Antagonistic [1705]. Antarctic [894, 1200, 1345, 413, 628, 512, 748, 600, 1517, 1656, 738, 679, 943]. Antennapedia [386]. Anthozoa [768, 458, 442, 500, 892, 416]. Anthozoan [707]. Anthropogenic [222]. Anti [1415, 1368, 1080, 1691, 1922, 1788, 1333, 1307, 1313, 808, 391, 1365, 839, 1899, 299, 1284, 990]. Anti- [1922]. Anti-Bacterial [1899]. Anti-cancer [1080]. Anti-coagulant [1284]. Anti-Fouling [1368]. Anti-HIV [1333, 1307, 1313, 299]. Anti-infective [1691]. Anti-Inflammatory [1415, 391]. Anti-microfouling [839]. Anti-obesity [808]. Anti-proliferative [1365]. Anti-protease [990]. Anti-Trichomonas [1788]. Antiangiogenic [308, 484]. Antibacterial [280, 1248, 1358, 764, 757, 1801, 876, 1901, 7, 827, 391, 1229, 885]. Antibacterial/Antibiofilm [1358]. Antibiofilm [1358, 1488]. Antibiotic [338, 1395, 1177, 782]. Antibiotic-Active [782]. Antibiotics [1722, 65, 899, 1575]. Antibodies [29, 513]. Antibody [300, 1257]. Anticancer [1491, 925]. Anticoagulant [1475, 391]. Antidiabetic [808]. Antifoulant [1700]. Antifoulants [58, 637, 769]. Antifouling [1016, 259, 1545, 567, 1463, 394, 595, 1631, 1138, 1430, 670, 638, 772, 114, 298, 1259, 1276, 1386, 297]. Antifreeze [894, 1200]. Antifungal [767, 1191, 1756, 391, 988]. Antigen [514, 29, 79]. Antigens [513]. Antihypertensive [213]. Antilipopolysaccharide [519]. Antimalarial [1191, 391]. Antimicrobial [809, 470, 1449, 1320, 793, 595, 453, 657, 1470, 367, 1159, 1650, 729, 830, 880, 967, 484, 1599, 289, 1917, 297, 788, 789]. Antinociceptive [1415]. Antioxidant [1415, 1653, 799, 1241, 1530, 354, 1929, 1173, 949, 1361, 1797, 1529, 760, 1666, 1679, 1859, 994, 76, 1006, 773, 1695, 1559, 1712, 1947]. Antioxidant-Related [1559]. Antioxidants [978]. Antioxidation [1904]. Antiplasmodial [575]. Antiplatelet [391]. Antipsychotic [1464]. Antisense [1161, 278]. Antiserum [143]. Antisettlement [58]. Antituberculosis [391]. Antitumor [1143, 1633]. Antitumour [949]. Antitumoural [1140]. Antiviral [15, 1104, 10, 391, 389, 1497, 961]. Aplidium [1631]. Aplysia [1358, 129, 576, 978]. Aplysina [367]. Apocyclops [993]. Apoptosis [1757, 1872, 1909, 1894, 961]. Apoptosis-Related [1909]. Apoptotic [41, 1478, 311, 957, 1911, 1529]. Apostichopus [599, 1338, 1719]. Aposymbiotic [532, 907]. Apparatuses [1828]. Apparent [668]. Applicability [1517]. Applicable [1005, 1439]. Application [272, 366, 1444, 799, 1481, 643, 75, 1793, 434, 1672, 784, 458, 1923, 260, 1150, 165, 399, 530, 151, 244, 753, 1325, 1208, 446, 1833, 306, 488, 1948]. Applications [614, 1545, 1772, 1866, 465, 1641, 1529]. Applied [400]. Apply [605]. Appraisal [922]. Appreciation [160]. Approach [1585, 928, 1788, 639, 1145, 1209, 1437, 398, 1745, 79, 1669, 695, 431, 691, 1865, 499]. Approaches [922, 1269, 1074, 1928]. Appropriate [362]. ApRab4 [832]. Aptamer [1877]. Aptamers [1088]. aqps [1684]. Aquaculture [1374, 392, 563, 1344, 1073, 1670, 71, 915, 1403, 1899, 881, 638, 542, 1650, 1067, 232, 753, 1178, 1695, 1905, 1946]. Aquaporin [598]. Aquaporins [1684]. Aquaria [181, 167]. Aquarium [159, 175]. Aquatic [1292, 752]. Aquimarina [1827]. AR06 [1122]. Arachidonic [1231]. Aragonite [322]. Arbitrarily [153]. Archaeal [1175, 1084]. Archaeol [70]. Archaeon [643, 574]. Architecture [1648]. Architectures [1282]. Arctic [1215, 784, 701, 199]. Arctogadus [701]. Area [52, 715]. Area-Efficient [715]. Areal [1624]. arenaria [1406]. arenysensis [1744, 1263]. areolata [1445, 1946]. Aresch [871]. Arginine [1141, 1047, 1124, 507]. Arginine-Rich [1141]. Argopecten [77, 1020, 608, 548, 1861, 1874]. Argopecten- [77]. argus [1511]. ariakensis [1533, 1532, 518]. Ark [621]. Arm [146]. ARNT [225]. ARNT2 [121]. ARNT2-like [121]. Arranged [764]. Arrayed [123, 267]. Arrays [1816]. Arrested [53]. Arrived [178]. Arsenal [1953]. Arsenate [541]. Arsenic [509, 1858]. Artemia [207, 1902, 1437, 1942, 718]. Arthrobacter [1189, 1248, 1265]. Arthrospira [28, 425, 1803, 1786, 488]. ARTICLE [1255]. Articular [1444]. Artificial [15, 1886, 777, 583, 858, 142, 1221, 1671, 1499, 564, 691, 710, 848]. Artificially [1097]. ARTP [1810]. Aryl [177, 225, 1293]. Ascidia [1326, 1033, 215, 405]. Ascidian [330, 268, 153, 1631, 1033, 1256, 215, 405].

Ascorbate [1184]. Asian [976, 538, 1303, 633, 669, 1461, 1375, 1399, 1537, 1662, 1155, 1794, 145, 306, 1865]. asiatica [1583]. asinina [375, 140, 450]. Asparaginase [1755]. Aspartic [955, 900, 1100]. Aspects [531]. Aspergillus [1080, 1449, 1134, 1756, 935, 886, 1169]. Aspidochirotida [462]. Aspolin [955]. Assay [46, 335, 354, 147, 510, 1413, 14, 1796, 300, 239, 741, 363, 114, 983, 956, 38, 1293, 1066]. Assays [1135, 667, 686]. assembling [1799]. Assembly [1875, 1759, 1771, 1905]. Assess [263, 1440]. Assessed [393, 404, 854, 518]. Assessing [1585, 1913, 139, 222, 1916]. Assessment [37, 328, 1192, 384, 993, 1171, 892, 1087, 1803, 1786]. Assignment [504, 353, 232]. Assimilation [1296]. Assisted [367, 1306]. Associated [338, 1746, 809, 1098, 1270, 1468, 1545, 652, 51, 1241, 1701, 719, 1491, 973, 1404, 1830, 1111, 1418, 767, 523, 951, 590, 1503, 1779, 161, 1673, 403, 1660, 1251, 1093, 664, 1862, 396, 1302, 1963, 1205, 1216, 725, 199, 183, 1179, 1447, 1809, 749, 1208, 1256, 1461, 990, 672, 1780, 1581, 484, 1253, 1127, 1082, 1201, 1375, 782, 1513, 1735, 1874, 1379, 1387, 872]. Association [1798, 1404, 1889, 717, 1170, 1945, 1605, 1636, 24, 857, 1800, 1172, 1920, 1767, 1781, 1375, 1598, 1542, 1377, 592, 1379, 1648, 1387, 1458, 1538]. Associations [1393]. Astaxanthin [1367, 1797, 1529, 237, 1679]. Astaxanthin-Rich [237]. Asterias [952, 1915]. Asteromonas [1372]. Astrocoeniina [731]. ATCC [1726, 1132]. ATGL [1110]. Atlantic [1910, 975, 928, 1820, 1331, 1594, 1681, 373, 419, 1760, 335, 506, 1035, 315, 1715, 406, 1294, 465, 1164, 436, 571, 1070, 1149, 927, 745, 1071, 1837, 819, 854, 1620, 295, 1679, 847, 860, 74, 350, 1023, 113, 380, 409, 1652, 1848, 714, 1573, 92]. atlantica [1122]. atlanticus [126]. atlanticus-like [126]. Atmosphere [346]. Atmospheric [486]. ATP [1479]. ATPase [859, 327, 291, 307, 215]. atropurpurea [836]. Attachment [101, 637, 1131, 1269, 16, 689, 289]. Attention [1323]. Attenuate [1505]. Attenuated [282, 747, 1655, 998]. Aurantiochytrium [1190, 1425, 1711, 791, 1766]. aurantium [268]. aurata [430, 1192, 1919, 1118, 1075, 24, 203, 1302, 95, 1024, 1102, 1462, 1578, 1059]. auratus [1274, 1546, 1493, 1492, 1587, 1647, 1895]. Aurelia [505]. Aureobasidium [1487, 1339, 632, 1268, 706, 761]. AUREOCHROME [1167]. Aureococcus [143]. aureus [757, 1449, 590, 781]. Austenitic [537]. australasicus [1855]. Australian [524, 587, 236]. australiensis [767]. australis [1808]. Autolytic [994]. Autonomous [1257]. Autophagy [1757, 1559, 1065]. Autosomal [1837]. Autotetraploid [1493, 1492, 1647]. Autotomy [1914]. Autotriploid [1546]. Autotriploidy [1274]. Availability [1192, 1408]. avara [739]. Avoidance [68]. Axinella [377, 989]. Axis [1934]. AY1 [555]. Ayu [422]. Azido [1285]. Azooxanthellate [652, 416]. Azotobacter [811].

B [911, 281, 100, 557, 218, 1336, 319, 491, 760, 910, 345, 1082, 1510, 593]. B-1 [319]. B1 [859]. Babylonia [1445, 1946]. BAC [330, 1299, 972, 1019, 1499, 865, 887, 663, 710, 848]. Bacillus [259, 1189, 1248, 1691, 764, 590, 1751, 268, 1246, 792, 367, 1495, 1650, 331, 1217, 787].

Backcross [926]. Background [1279, 1460]. Bacteria [46, 511, 341, 722, 757, 1526, 1801, 876, 261, 793, 1224, 1171, 16, 713, 857, 872, 865, 725, 1153, 229, 1326, 1148, 1705, 1517, 1208, 990, 974, 484, 410, 1069, 782, 685, 289, 1137, 553, 1488, 1924, 788, 789]. Bacterial [1018, 1175, 652, 1089, 583, 1805, 286, 1320, 1084, 1314, 1802, 1777, 831, 858, 1410, 717, 1827, 142, 1615, 1121, 1499, 65, 1122, 239, 1439, 502, 915, 485, 1953, 742, 1899, 868, 564, 326, 362, 423, 1650, 283, 379, 1617, 113, 1464, 1177, 1394, 1455, 1082, 933, 1196, 710, 848]. Bacterially [1194, 698, 693]. Bacteriolysis [1189, 1248]. Bacteriolytic [1248]. Bacteriophage [217, 479, 179]. Bacterium [338, 845, 1468, 51, 534, 545, 1232, 1475, 945, 383, 457, 217, 579, 754, 801, 1673, 218, 187, 1442, 1936, 1311, 986, 453, 1037, 475, 1179, 365, 1296, 1021, 82, 154, 738, 1396]. Baculovirus [1669, 1325, 324]. Baculovirus-Mediated [324]. BAGM01 [1635]. Baikal [979, 1664]. baikalensis [979]. bairdi [1724]. Bait [1917]. Bal [1799]. Balaenoptera [540]. Balance [867]. Balanced [1337]. Balanus [258, 489, 1799, 1659]. Bali [1893]. Baltic [348, 782]. Banana [1674]. Bangia [836]. Bangiales [342]. Bangiophycean [1005]. Banks [760]. Bar [421]. barbatus [434, 208]. Barbour [1963]. barbouri [1963]. Barcoding [1226, 1022, 1323]. Barents [1449]. Barettin [1700]. Barfin [78]. Barnacle [1702, 1741, 258, 489, 457, 1484, 1659]. Barnacles [14, 977]. barretti [568]. Barrier [1264, 246]. Base [1433, 174, 1356]. Based [366, 463, 1018, 975, 37, 1870, 621, 1068, 1271, 866, 1445, 364, 1279, 1111, 661, 370, 1373, 451, 81, 1413, 171, 462, 1289, 490, 964, 390, 1784, 1952, 1744, 877, 1222, 1133, 585, 1593, 229, 1296, 374, 1253, 1008, 422, 431, 1833, 1944, 1567, 1507, 691, 1828, 1323, 496, 1322, 753, 728, 1226]. Basic [272, 465]. Basin [902]. Basis [242, 1745]. Bass [197, 392, 520, 539, 328, 716, 111, 1830, 1754, 124, 253, 1041, 230, 669, 1512, 906, 1077, 345, 310, 1388, 306, 1362, 1414, 1595]. Bastadin [1016]. bata [1531]. batch [497]. Bath [822]. Bathyal [385]. batrachus [1875, 30]. Bay [1351, 608, 679, 831, 1020, 548, 70]. BayesC [1456]. Bayesian [1282]. BBW831 [1811]. BC1 [331]. Be [866, 1129, 909]. Bearing [651, 838, 162, 231, 1937]. Beauty [971, 205]. Beginnings [182]. Behaviors [1617]. belcheri [487, 132, 321]. Beloniformes [236]. Benefit [1126, 8]. Benthic [618]. Benzyl [935]. Bering [619]. Beta [999, 1953]. Beta-Defensin [1953]. Between [1676, 866, 364, 609, 1855, 1212, 1341, 1889, 228, 1821, 683, 1945, 1242, 411, 381, 1393, 491, 743, 1824, 1172, 1179, 1624, 1124, 1521, 16, 113, 1412]. BG3 [692]. bgl [1329]. bgl-A [1329]. Bias [1401]. Biased [1918, 1733]. bicuspidata [1808]. Bicyclic [42, 109]. bicyclis [1055]. bidens [1812]. Big [1520]. Bighead [1357, 1798, 1634, 1677, 1556]. bimaculatus [1563]. Bind [1097]. binderi [987]. Bindin [937]. Binding [1099, 274, 1491, 1213, 1333, 1307, 1313, 816, 836, 937, 133, 1197, 1033]. Binds [1454, 218]. Bio [1125]. Bio-inspired [1125]. Bioaccumulation [1326]. Bioactive [1545, 1060, 882, 822, 1431, 403, 1134, 1224, 632, 398, 71, 761, 1623, 772, 1699, 1082]. Bioactivities [41]. Bioactivity [1698, 1321, 1711, 1008]. Bioadhesive [1799]. Bioassay [243]. Bioassays [1463]. Biocalcification [674, 1495]. Biocatalysis [1501]. Biocatalytic [1576]. Biochemical [1260, 1312, 1358, 1843, 207, 1009, 1445, 1416, 1543, 1322, 1745, 1755, 573, 932, 1100, 736]. Biochemistry [1806, 824]. Biocontainment [1403]. Bioconversion [1431, 1056]. Biodegradable [980]. Biodegradation [1340, 1477, 1173]. Biodiesel [1190, 1372, 1382]. Biodiversity [1127]. Bioelement [797]. Bioencapsulation [1437]. Biofilm [537, 1742, 1210, 1615, 1217]. Biofilms [792, 638]. Biofilter [1680]. Biofilters [1153]. Biofouling [1125]. Biofuel [1090]. Biogenesis [832]. Biogenic [763, 1029, 1196]. Biogeochemical [953]. Biohydrogen [445]. Biohydroxylation [1240]. Bioinformatic [861, 426]. Bioinspired [1772]. Biological [520, 1002, 683, 1437, 1923, 1430, 1483, 1580, 311, 855]. Biologically [361]. Biology [1804, 1073, 968, 1869]. Bioluminescence [265, 1618]. Biomarker [139, 222, 600, 690]. Biomarkers [1408, 1864, 1225, 1567, 1508, 505]. Biomass [1372, 1095, 1406, 50, 1195, 70, 1516]. Biomedical [1772]. Biomimetic [1125]. Biomineral [1158, 708]. Biomineralised [1448]. biomineralising [1113]. Biomineralization [613, 1376, 1000, 1540, 1108, 1151, 1324, 1572, 1714, 1086]. Biomineralization-Related [1151]. Biomineralizing [779]. Biominerals [1940]. Biopolymer [237]. Biopolymers [552]. Bioprocess [65]. Bioproducts [1322]. Bioprospecting [416]. Biopterin [28]. Biopterin- [28]. Bioreactor [65, 729]. Bioreactors [438]. Bioreduction [1053, 1285]. Bioremediation [828, 312, 691]. Biosciences [629]. Biosensor [366, 332]. Biosilica [1486, 1616, 1823]. Biosilicification [449]. Biosorption [50]. Biosurfactant [1635, 1198, 811]. Biosynthesis [611, 1742, 1760, 377, 1526, 1457, 1482, 1630, 1603, 1058, 1050, 1522, 1591, 1690]. Biosynthetic [1330, 1722, 1045, 1743, 1607]. Biotechnological [272, 72, 1529, 1745, 1517, 336]. Biotechnologically [1396]. Biotechnology [614, 1247, 49, 629, 588, 344, 1614, 1853, 1]. Biotransformation [545, 1579, 886, 1101, 1169, 1609]. Biotyper [1121]. Biowaste [1173]. Biphasic [536]. birdy [515]. Birnavirus [337, 97]. Bisexual [1588]. Bisphenol [1909]. Bivalve [1444, 454, 993, 1543, 1234]. Bivalves [1863, 1857]. Bivalvia [569, 923, 98]. Black [543, 1843, 1658, 270, 1868, 276, 461, 1364, 62, 1674, 117, 1391, 1612, 304]. Black-Footed [543]. Blacklip [524, 587]. Bladder [1929]. Blainville [299]. Blastema [1818]. Bleaching [734, 891]. Blenniidae [104]. Blennioidei [104]. Bloch [1953]. Block [342]. Blood [742]. Blooming [1049, 467]. Blooms [801, 1593]. Blotches [1514]. Blue [196, 1701, 1167, 7, 114, 1935, 1907]. Bluefin [315, 113]. Blunt [1567]. BMP [779]. BMP2 [779]. BMP2/4 [779]. Bodied [1278]. Body [1287, 1798, 1282, 317, 1404, 240, 1605, 562, 204, 1447, 131, 1538]. Body-Shape-Related [1538]. boliviensis [890]. Bombardment [191].

Bonamiosis [814]. Bone [1404, 1763, 1434, 1438, 1580, 1823, 1694, 1578]. Bony [57, 1542]. Boone [908]. borealis [540]. Boron [797]. Both [1175, 605, 1579, 258, 1094, 1188, 1488, 594]. Botryllus [330]. Bottarga [260]. Bottle [289]. Bottleneck [224]. Bovine [337]. Box [143]. bp [63, 94]. BPMS22 [1475]. brachiata [1213, 1091, 1184]. Brachionus [80, 1009, 1725, 585, 930, 758]. Brachymystax [1797, 1939]. Brain [1658, 1275, 140, 1114, 624, 1842, 1776]. Branched [804]. Branched-Chain [804]. Branches [1916]. Branchiostoma [487]. brasiliensis [1568]. Brazil [1653, 1845]. Brazilian [1240, 1148, 1169, 1609]. Bream [430, 1192, 270, 327, 598, 1118, 1720, 1111, 1834, 285, 95, 1665, 1407, 1447, 1462, 1112, 1567]. Breathing [1341, 1634]. Bred [821, 1130]. Breeders [1913]. Breeding [1596, 404, 335, 1119, 1810]. brevibarba [1525, 1552]. brevidactylus [332]. brevis [684, 508, 837]. Bright [1558]. Brine [1450, 1942]. British [144]. Brittany [394]. Brittle [87]. Broad [217, 1933, 1488]. Broad-Host-Range [217]. Broad-Spectrum [1488]. Bromoacetophenones [886]. Bromoperoxidase [1239]. Bromoperoxidases [1133]. Bromotyrosine [1215, 670]. Bromotyrosine-Derived [670]. Broodstock [921, 306]. Broodstocks [1149]. broughtonii [621]. Brown [1415, 37, 534, 224, 1143, 143, 1409, 853, 839, 293, 326, 1048, 689, 782, 1565, 1900, 1873]. Brown-Marbled [1565]. bruneus [1300]. bryoides [991]. BS [1198]. BS-15 [1198]. Bubble [802, 810]. Buccal [55]. Budget [1011]. Building [1480, 690]. Bulk [1721]. Bulked [1201, 1586, 1668]. Buried [631]. Bursatella [299]. Butcher [1917]. butcheri [605]. Butenolide [1488]. By-product [1730, 1420]. By-products [1529]. Bycatch [1698, 1653]. Byssal [119, 1168].

C [1337, 1337, 36, 103, 1143, 1289, 1159, 1154, 215, 496]. C-Terminal [1289, 1159]. C-type [103]. C. [583, 653, 1821]. C1 [425]. C16 [1132]. C16- [1132]. C18 [1132]. C18- [1132]. C20 [755]. C20- [1132]. C20-elongase [755]. C52 [986]. Ca [859, 838, 957]. CaCO [913, 1196, 594]. Cádiz [451]. Cadmium [699, 98, 336]. Calanus [386, 1040]. calcarifer [922, 538, 230, 669, 912, 145, 306]. Calcification [838, 1433, 786, 471]. Calcified [674]. Calcifying [1748, 1017, 997, 417]. Calcineurin [1064, 844]. Calcite [913]. Calcium [859, 720, 816, 659, 765, 1940]. Calcium/Calmodulin [765]. Calcium/Calmodulin-Dependent [765]. Caleosin [1205]. caliendrum [1277]. California [1260, 129, 54]. californica [129, 707]. californiensis [292]. Caligus [1820, 1851, 1918, 1290]. Callinectes [196, 7]. Callinectin [7]. Calmodulin [765, 1197]. Calmodulin-Binding [1197]. Calmodulin-Dependent [765]. Caloglossa [234]. CaLP [780]. calpain [253]. Calyptogena [1047]. Camelina [1331, 1331]. cAMP [1934]. cAMP-Mediated [1934]. campechanus [406, 212]. campechiensis [1729]. Camptothecin [1463]. camtschatica [106]. Can [1064, 909, 1129, 1572, 1578, 1575]. canadum [1831, 1727]. Canal [312]. canaliculatus [1482, 1522, 1591]. canaliculus [1441]. Cancer [1478, 1107, 319, 1365, 1515, 1897, 1080, 1255]. Candidate [1238, 1798, 719, 1212, 1553, 1170, 927, 664, 1634, 887, 1324, 1781, 1090, 898, 1710, 1712, 1565, 1874, 1113]. Cannery [312]. Capability [1416]. Capable [1199]. Capacity [596, 1192, 1797, 1245, 1947]. Capsosiphon [947, 1763, 1284]. Captive [1802, 1777, 831, 803]. Capture [1672, 113, 1196]. Capturing [939]. Carangoides [1953]. Carapace [332]. Carassius [1274, 1546, 1493, 1492, 1587, 1647, 1960, 1895]. Carbohydrate [1465, 1301, 952, 836]. Carbohydrate-Binding [836]. Carbon [867, 1225, 1817, 1322, 1261, 1828]. Carbonate [816, 659, 1940]. Carbonate-Binding [816]. Carbonic [1001, 258, 1328, 1392, 674, 1356]. Carboxylase [443, 1295, 1457]. Carboxypeptidases [106]. Carcass [317]. Carcasses [672]. Carcharhinidae [204]. Carcharhinus [204]. carchariae [279]. Carcinogen [895]. Carcinogenesis [185]. Cardiac [1192]. Cardiovascular [391]. Caribbean [1270, 1140, 990, 1127]. carnosus [1114]. Carotene [511]. Carotenoid [737, 1221, 1679]. Carotenoprotein [994]. carotovorum [1505]. Carp [6, 12, 1886, 1904, 1557, 1798, 1870, 1658, 1271, 317, 284, 858, 972, 1019, 852, 1531, 1076, 429, 1677, 1960, 1655, 1846, 1911, 942, 1003, 1117, 1496, 1612, 1776, 1736, 1458, 1556, 1895]. Carpet [126]. carpio [6, 12, 1271, 317, 284, 1211, 972, 1019, 942, 1117]. Carrageenan [991, 402, 313, 945, 1116, 1306, 1116]. Carrageenans [1469]. Carrageenase [945]. Carrageenin [1919]. Carrageenolytic [1468]. Carrageenophyta [1333, 1307, 1313]. Carries [582]. Carrying [199]. carterae [366, 523, 981]. Cartilage [1444, 325]. Cas-9 [1775]. Cas9 [1424, 1334, 1645, 1762, 1944, 1507, 1948]. Cas9-Mediated [1747, 1825]. Case [922, 104, 1588, 1323]. Caspase [961]. Caspase-Dependent [961]. Catalysis [120]. Catalysts [1384]. Catalytic [945]. Catalyze [935]. Catapulting [864]. catenella [510, 448, 746]. Catfish [1697, 1357, 1238, 1775, 68, 286, 1555, 1404, 40, 1721, 309, 190, 1435, 59, 245, 1119, 1275, 1634, 30, 1316, 426, 1471, 201, 870, 663, 1510, 1521, 1747, 1564, 1387]. Cathepsin [900, 487]. Cation [425]. Cationic [880, 967]. Caudal [1578, 912]. Caulerpa [936, 1633]. Caulerpenyne [936]. Causal [1752]. Causative [1458]. Cause [1620]. Caused [1055, 279, 1347, 918, 1689]. Causes [1287]. Causing [366, 1705]. Caution [420]. cava [702, 911]. Caves [1045]. CAZy [1465]. CAZyme [1468]. CAZyme-Rich [1468]. CAZys [1465]. CBMAI [1576, 1579, 1169, 1609]. CC [1602]. CCl [950]. Cd [541]. Cd-Stimulated [541]. Cdkn1 [1952]. cDNA [414, 63, 470, 494, 647, 651, 653, 1159, 1324, 428, 380, 580, 1036]. cDNAs [443, 527, 436, 954, 301, 1023]. Ceca [156]. cecilia [882]. Cecropin [281, 286, 910]. Cell [596, 169, 303, 539, 454, 610, 210, 1334, 41, 377, 325, 1183, 1740, 1613, 682, 1454, 878, 1413, 838, 1657, 448, 480, 1763, 218, 513, 1442, 1435, 1141, 1739, 319, 396, 398, 237, 915, 1597, 1640, 1425, 627, 669, 79, 503, 939, 429, 1669, 1236, 1436, 311, 166, 369, 1178, 989, 471, 257, 60, 1894, 1943, 836, 1422, 1578, 38, 1944, 1446, 1735, 781, 1948, 635, 561, 1912, 1574, 498]. Cell-Free [1442].

Cell-Labeling [60]. Cell-Less [1740]. Cell-Penetrating [1141]. Cell-Specific [1178]. Cell-to-Cell [915]. Cells [1104, 614, 90, 1804, 1481, 651, 971, 947, 1174, 684, 1732, 951, 205, 838, 1055, 424, 1478, 1344, 818, 1240, 237, 1123, 1665, 687, 742, 669, 1365, 122, 1674, 429, 1053, 1285, 1669, 282, 149, 369, 736, 591, 1245, 1578, 417, 121, 1136, 1065, 1298, 1500, 594, 1497, 1539, 336, 1609]. Cellular [1543, 1664, 989, 1277, 1422]. Cellularity [199]. Cellulase [929, 917, 55]. Cement [1702, 1741, 1799]. Centers [184]. Central [1582]. Centric [1663]. Centromere [680]. Centromeres [1092]. Centromeric [233]. Cephalopod [1058]. Cephalopoda [137]. Cephalothrix [1876]. Ceramiales [234]. cereus [787]. Cerianthus [645]. ceRNA [1852]. CF [1284]. cf. [1876]. CF4MRS [1437]. CFR26M [1173]. Chaetoceros [894, 1663, 355]. Chaetomium [359]. Chain [569, 1190, 141, 643, 934, 135, 296, 1760, 195, 1693, 667, 571, 153, 1645, 686, 115, 456, 564, 1785, 84, 671, 804, 200]. Challenge [49, 1338, 116, 1960]. Challenged [1165, 1675, 1608, 1186, 744]. Challenges [1672, 749]. Chambers [1208]. Change [1291, 1883, 1887, 276]. Changes [1477, 1938, 235, 1085, 1805, 207, 514, 1800, 146, 199, 1493, 1492, 896, 983, 956, 673]. changii [806, 1311]. Changing [1925]. Channa [1511]. Channel [1697, 1775, 1417, 68, 286, 40, 1721, 309, 190, 1435, 245, 1316, 426, 201, 870, 663, 1747, 1351]. Chanos [627, 252]. Chaperone [1459]. Char [199]. Character [1226, 1117, 1323]. Character-Based [1323, 1226]. Characterisation [1830, 952, 1150, 875, 795, 1223]. Characterised [1574]. Characteristics [1811, 1294, 540, 50, 378, 1437, 748, 1622, 573, 214, 1869, 1879]. Characterization [156, 991, 1692, 1227, 6, 12, 1487, 1243, 56, 845, 90, 1260, 11, 80, 1558, 543, 539, 805, 328, 1653, 606, 1723, 1358, 415, 1357, 401, 1241, 1321, 1731, 470, 643, 511, 1843, 1724, 1329, 1635, 1354, 323, 330, 1167, 138, 1255, 1286, 582, 1343, 1372, 1002, 1295, 258, 1486, 1288, 1232, 35, 720, 1678, 124, 767, 1868, 436, 945, 1314, 40, 832, 766, 645, 234, 320, 81, 590, 858, 461, 784, 1299, 1160, 1200, 1345, 268, 1827, 476, 1328, 339, 754, 1289, 351, 604, 187, 253, 18, 247, 435, 574, 554, 98, 730, 692, 972, 1019, 1416, 1352, 538, 1128]. Characterization [458, 986, 1043, 1834, 657, 837, 827, 664, 632, 203, 1923, 186, 1015, 125, 1470, 819, 706, 352, 1623, 295, 872, 1953, 1395, 627, 1483, 1037, 1046, 1258, 1755, 229, 759, 1512, 932, 333, 1100, 1659, 1236, 365, 423, 1592, 906, 1148, 1650, 1077, 221, 1447, 307, 1108, 1151, 1517, 192, 1147, 1006, 703, 310, 1327, 1462, 1663, 883, 1109, 487, 479, 663, 648, 787, 1124, 1062, 86, 39, 855, 1112, 1388, 685, 82, 1833, 851, 522, 1429, 673, 145, 1059, 150, 488, 546, 710, 679, 848, 1684, 635, 1310]. Characterize [1793, 1649]. Characterizes [1655]. Characters [826]. Charonia [1667]. Chelate [457]. Chemical [1061, 1321, 1371, 311, 1517, 1008]. Chemically [157]. Chemistry [734]. Chemokine [805, 1602]. Chemosystematics [603]. Chemotaxis [46]. Chemotherapy [575]. Chesapeake [831]. chevalieri [1545]. CHH [361]. Chilean [1489]. chilensis [1560]. Chimera [610]. Chimeras [1389]. China [421, 973, 661, 767, 1084, 902, 1229, 1903, 265, 788, 789, 621, 1022]. chinensis [1165, 378, 519, 1159, 580, 744, 1895]. Chinese [411, 378, 1138, 1914, 519, 1398, 1249, 1426, 1818, 1559, 1604, 1718, 1414, 1900, 1237]. Chinook [1600, 407, 116]. Chionoecetes [1724]. ChIP [1566]. ChIP-seq [1566]. Chitin [1925, 1871, 1431, 552, 984]. Chitinase [767, 89, 787]. Chiton [1737]. Chitooligosaccharides [1431, 591, 862]. Chitosan [1579, 1301, 1337, 332, 552, 1899, 466, 150]. Chitosan-Encapsulated [466]. Chitosan-Mediated [1301]. Chlamydomonas [1373, 541, 1051, 943, 1649]. Chlamys [1543, 1092, 1151, 1428, 1474, 131, 775, 655, 710, 688]. Chlorella [1227, 1297, 254, 995, 213]. Chloroperoxidase [1843]. Chlorophyceae [237]. Chlorophyta [737, 958, 1297, 1195, 16, 695]. Chlorophyte [1199]. Chloroplast [1457, 334, 1583, 1599, 1203]. Choice [622]. Cholesterol [331]. Cholesterol-Transforming [331]. Chondrocytes [1444]. Chondroitin [1265]. Chondrosia [1046, 1258]. Chordate [1135]. Christmas [1465]. chrLG18 [1503]. Chromatin [1135, 1769]. Chromatography [889]. Chromoprotein [1193]. Chromosomal [1525, 1552, 248, 663, 1627, 1626]. Chromosomal-scale [1627, 1626]. Chromosome [583, 858, 142, 1103, 1759, 1499, 1771, 710, 848, 1400]. Chromosome-Derived [1499]. Chromosome-level [1759]. Chromosome-Scale [1771]. Chromosomes [564, 479, 688]. Chronic [270, 1729, 1102]. Chronically [1813]. chroococcum [811]. chrysoplasta [617]. CHSE [1104]. CHSE-214 [1104]. chuatsi [920, 1237, 1879]. Chum [432, 619]. chymotrypsin [1260]. Cichlid [128]. ciliata [521]. Ciliate [1504]. Ciona [1135, 153, 1924]. circRNAs [1852]. Circulating [1864]. Cirripedia [1484]. Cis [1646]. Cis-acting [1646]. citreus [1189, 1248]. Citric [1295, 1094]. Citrinin [545]. citrinum [1579]. CK2 [954]. Clade [557]. Cladosiphon [379]. Cladosporium [1576]. Clam [1892, 861, 126, 1026, 427, 1750, 1093, 997, 1309, 1611, 1047, 1928, 817, 1932, 581, 1254, 1706]. Clamping [934]. Clams [889, 1865]. Clarias [1357, 1875, 1634, 30, 1959]. Clarified [1271]. Clarion [763]. clarkii [1880, 1947]. Class [386, 125, 128, 352, 592, 593]. Classical [911, 1745]. Classification [171, 1121]. Clathria [831]. Clathrina [973]. Clavularia [364]. Clawed [900, 1100]. Clearance [377]. Cleavage [120]. Climbing [1890]. Clipperton [763]. cloacae [1635]. Clock [1473, 1385]. Clonal [458, 725, 422]. Clone [1557, 766, 1110]. Cloned [645, 1091, 1184]. Clones [63, 887, 479, 710]. Cloning [156, 1487, 987, 382, 401, 395, 470, 452, 829, 1295, 436, 103, 832, 276, 387, 905, 1339, 579, 754, 439, 161, 604, 248, 247, 574, 1416, 143, 245, 458, 519, 1398, 45, 202, 125, 743, 706, 761, 349, 95, 301, 1037, 924, 78, 449, 1258, 1159, 759, 194, 910, 906, 307, 83, 1327, 883, 26, 625, 388, 488, 738, 494]. Closely [250, 573]. Clue [1733]. Cluster [1722, 1503, 398, 1591]. Clusters [1045, 1743, 1029]. Cnidaria [381, 442, 500, 13, 892]. Cnidarian [401, 645, 1422]. Cnidarians [709, 708]. CO [1615, 1704, 1904, 1870, 1802, 1777, 1750, 1839, 1342, 497, 1529, 1199, 1663, 1356, 1196, 1649]. Co-Culture [1802, 1777]. Co-culturing [1839]. Co-detection [1704]. Co-Expression [1750, 1342, 1870]. Co-Forms [1904]. CoA [1345, 1457, 1040, 1766]. coagulant [1284]. Coagulation [438]. Coalescent [1226]. Coast [406, 394, 665, 20, 129, 1903, 1022]. Coastal [809, 196, 101, 884, 97, 1431, 483, 331, 450, 691]. Coastline [889]. Coasts [1623]. Coatings [1125]. Cobetia [1232, 475]. Cobia [1831, 1727]. Cobitidae [349]. coccinea [1699, 1137]. Coccolith [651, 523, 838, 913, 1433]. Coccolith-Bearing [651, 838]. Coccolithophore [786, 1261]. Coccolithophores [44]. Coccolithophorid [320, 838, 913, 257, 607, 417]. Coccolithophorids [471]. Coccoliths [981]. Coccosphere [607]. Cochliobolus [1330]. Cochlodinium [448]. Cod [975, 928, 335, 1035, 465, 1070, 1149, 701, 74]. Coding [527, 768, 36, 1660, 442, 500, 1378, 1659, 988, 1338, 1610, 1767, 1769, 1652, 1412]. Codium [1245]. Codon [1177]. Codon-Optimized [1177]. Codons [565]. Coelomate [288]. Coenzyme [443]. Coexpression [1533, 1532]. Cognitive [1713]. Coho [1637, 1675, 1283]. COI [621, 490, 13]. coioides [533, 604, 627, 1651, 1527, 1833]. Cold [371, 1105, 826, 568, 1872, 717, 385, 52, 986, 628, 1100, 786, 1781, 1047, 996, 581, 738, 679, 1549, 561]. Cold-Adapted [826, 986, 628, 1100, 738, 679, 1047]. Cold-Pretreated [1549]. Cold-Seep [52]. Cold-Stress [1781]. Cold-Water [568]. coli [392, 511, 766, 1442, 1547, 1639, 755, 749, 282]. Collaboratively [1064]. Collagen [1625, 1002, 1483, 1580, 122, 1046]. Collected [876, 793, 52, 1008]. Collection [975]. Colonial [1277]. Colonisation [1217]. Colonization [537, 1437, 555, 638]. Colonized [1718]. Colony [1623]. Colony-Forming [1623]. Color [1536, 1279, 1855, 1554, 599, 240, 1660, 1752, 608, 1447, 1466, 1780, 1391, 1254, 1379, 264]. Color-Related [1254]. Colorectal [1436]. Colors [1193, 1690]. Columbia [144]. Columnaris [1564]. Combination [640, 1881]. Combined [37, 445]. Combining [814]. comes [1470]. Commercial [1451, 55]. Commercially [1746, 80, 321, 1370]. Common [1271, 1877, 317, 284, 972, 1019, 852, 1180, 265, 942, 1117, 1496]. Communication [915, 1574]. Communities [1175, 652, 1893, 1805, 1314, 831, 1615, 198, 1082]. Community [1018, 1089, 1680, 296, 1084, 568, 868, 933, 732]. Comparable [859]. Comparative [532, 768, 1925, 176, 1715, 1465, 1883, 1081, 370, 1209, 1142, 1902, 927, 1335, 1603, 1634, 745, 87, 182, 1597, 1640, 168, 704, 1484, 1382, 696, 1048, 1729, 1780, 4, 1717, 1791, 1426, 1521, 1891, 1412, 1908, 1790, 1689, 1733, 1575, 1078]. Compared [860]. Comparison [1190, 292, 1855, 523, 816, 831, 1165, 1242, 1520, 129, 1711, 1179, 1624, 1460, 744, 1226, 1556]. Compensatory [1887]. Competence [610]. Competitive [130, 410]. Complementary [1312, 19, 85, 317, 284, 36, 439, 5, 247, 98, 245, 491, 45, 202, 125, 78, 1235]. Complementation [511]. Complete [141, 1271, 1732, 538, 294, 745, 529, 493, 220, 800, 1879]. Complex [1016, 1469, 425, 13, 322, 1338, 128, 168, 592, 593, 1302]. Complexes [1613, 1410, 1507]. Complexity [618, 1905]. Component [898].

Components [1569, 781]. Composite [1806, 1580]. Composition [1287, 1018, 317, 207, 1009, 568, 1802, 1777, 791, 1937]. Compositions [1537]. Compound [338, 1463, 595, 1139]. Compounds [210, 1413, 1224, 391, 53, 398, 1623, 772, 1650, 729, 1859, 343, 1464, 1699, 796, 1276, 289]. Comprehensive [1238, 1754, 1685, 1928, 1783]. compressa [958]. Compromise [1341]. Concentrating [240, 1663]. Concentration [377, 337, 464]. Concentrations [1855]. Concomitant [1173]. Concrete [1495]. Condition [943]. Conditional [1485]. Conditions [1898, 1811, 1726, 743, 1050, 1840, 1666, 362, 786, 1489, 820]. Conference [1614, 1853]. Conferring [1547, 1639]. Confers [1184]. Confirmation [1862]. Confirms [111]. Confocal [400, 261]. Conformation [1404]. Conformational [352]. Congeners [1016]. Conjugated [1911]. Conjugates [146]. Connectivity [1892, 1761]. Connexin [92]. Conopeptide [1228]. Consensus [1117]. Conservation [292]. Conservative [118]. Conserved [18, 93, 1041, 1052, 771, 1178, 1769]. Consideration [485]. Consortia [1098, 138]. Constitutive [273, 36, 1457, 1766]. Constitutively [1104]. constricta [1892, 861, 1309, 1397]. Constructing [1504, 1511]. Construction [330, 1555, 332, 902, 858, 217, 1299, 1954, 954, 1366, 834, 964, 972, 1019, 825, 1499, 1338, 1397, 1097, 865, 3, 887, 1563, 776, 1203, 1027, 655, 710, 848]. Constructs [1161, 1435, 700]. Containing [1653, 1701, 1264, 68, 284, 364, 1337, 427, 30, 542, 887, 1259, 1388, 1139, 710, 1912]. Contains [211, 1082]. Contaminated [1585, 1175, 889, 1054]. Content [783, 1764, 227, 1411, 983, 956, 791, 1090, 1810, 1747]. Contents [1180, 1120, 1466, 1924]. continua [234]. Continuous [539, 1669, 497, 1009]. Contrasting [428]. Contributes [900, 898]. Contribution [111, 1860, 904]. Control [559, 1080, 421, 315, 406, 888, 1410, 801, 538, 1437, 869, 1432, 252, 4, 506]. Controlled [1469, 1131, 1281, 1638, 820]. Controlling [1116]. Conus [1228]. Convenient [14, 1293]. Convergent [772, 1360]. Conversion [787]. Converting [1929, 1778]. Cooperation [1892]. Coordinated [644, 536, 1574]. Cope [1192]. Copepod [993, 985, 1345, 386, 1040]. Copepoda [294]. Copepods [251]. Coping [1130]. Copper [1016, 1175, 958, 98, 1435]. Copper-Contaminated [1175]. Copper-Induced [958]. Copper-Sensitive [1435]. Copy [1355]. Coral [1001, 1723, 41, 139, 364, 1045, 618, 1802, 1777, 668, 1506, 1748, 666, 1657, 462, 803, 1673, 458, 1480, 765, 1453, 743, 968, 1916, 891, 1259, 690, 1114, 1617, 1277, 946, 459, 1542, 1494]. Coral-Associated [1673]. Coral-Derived [1259]. Coral-Reef [462]. coralliilyticus [1480, 1505, 1617]. Corallium [441]. Corals [532, 141, 91, 1673, 1137, 907, 779]. Corbicula [1865]. Core [1269]. Cores [641]. cornea [980]. corneum [1530]. coronopifolius [977]. Correction [1681, 1741, 1552, 1802, 1533, 1639, 1930, 1640, 1493, 1683, 1401, 1803, 1455, 1668, 1897, 1577, 1627]. Correlates [1620]. Correlation [1750, 1502, 1342]. Correlations [1391]. Corresponding [301]. Corrosion [537]. corrugata [377, 1632]. Cortez [129]. Cortisol [1251, 116, 1102, 1462]. Corynactis [707]. Cosmetics [1772]. Cosmid [267]. Cost [1704, 1817]. Cost-Effectiveness [1704]. Cost-Efficient [1817]. costatum [790, 27]. Count [1551]. Coupled [1725, 1319]. Coverage [907]. COVID [1835, 1638]. COVID-19 [1835]. COVID-19-Related [1638]. cp [1799]. cp-20 [1799]. cp23S [334]. CpG [1381]. Crab [675, 196, 1724, 1465, 332, 785, 7, 1914, 1608, 106, 1249, 1943, 1509, 1818, 1412, 1601]. Crabs [96]. Cranial [1823]. Craniella [767]. Crassostrea [90, 999, 583, 130, 1417, 1526, 1860, 1279, 1883, 1816, 228, 1057, 1852, 256, 653, 1305, 1821, 1926, 1687, 1030, 132, 193, 321, 490, 821, 390, 1571, 1645, 1660, 1844, 1533, 1532, 1566, 1017, 491, 529, 1728, 456, 1385, 1315, 1845, 1881, 1683, 1682, 347, 360, 885, 1466, 1467, 1534, 1767, 1780, 1440, 233, 479, 648, 1356, 1443, 1716, 937, 1391, 1908, 447, 1427, 1714, 518, 1157, 1342, 1310, 1690]. Crassostreas [756]. Crassotrea [216]. Crayfish [1757, 1712, 1947]. Cre [640, 1136]. Cre/loxP [1136]. CREM [1934]. CREM-MITF-TYR [1934]. cremoris [1292]. Crenarchaeote [246]. crispata [1469]. CRISPR [1424, 1775, 1334, 1645, 1747, 1762, 1944, 1507, 1825, 1948]. CRISPR/Cas [1775]. CRISPR/Cas-9 [1775]. CRISPR/Cas9 [1424, 1334, 1645, 1747, 1762, 1944, 1507, 1825, 1948]. CRISPR/Cas9-Mediated [1747, 1825]. Critical [1768, 892, 1927]. Critically [1718]. Croaker [1807, 1346, 1646, 1955, 1504, 1605, 680, 1956, 1636, 1758, 1781, 1110, 1850, 1762, 1219, 92, 1847, 1927, 1790, 1648, 1538]. crocea [1346, 427, 1955, 1504, 1605, 680, 1956, 1636, 1758, 1769, 1110, 1762, 1219, 136, 1847, 1549, 1648, 1538]. Cross [1772, 647, 1672, 1133]. Cross-linking [1772]. Cross-Species [647, 1133, 1672]. Crosses [491, 577]. Crosslinking [1702, 1741]. crossover [497]. Crowding [1251, 1144]. CrtW [511]. Crucial [1154]. Crucian [852, 1960, 1895]. Crude [138]. Cruise [1008]. Crustacea [219, 294, 413, 356, 521, 676]. Crustacean [1698, 64, 395, 478, 85, 1465, 361]. Crustaceans [552, 493, 1745, 162, 231, 431]. Crustins [280]. Cryopreservation [1123, 1674, 175]. Cryopreserved [1281, 468, 1718]. Cryptic [1323]. Cryptocaryon [1873, 1826, 1504, 1758, 1921, 1648]. Cryptocaryonosis [1956]. Cryptococcus [555, 781]. Crystal [659, 913, 594]. Crystalline [1234]. Crystallization [913, 322]. Crystallographic [981]. CSN423 [1349]. CspA [1329]. CT [104]. CTDA0820 [1824]. Ctenopharyngodon [1904, 1846, 1911, 1003, 1776, 1736]. Cu [339, 631]. Cu/Zn [339, 631]. Cu/Zn-Superoxide [339]. cucullata [321]. Cucumber [323, 1866, 1889, 476, 599, 102, 1338, 1426]. Cucumbers [462]. Culter [1875]. Cultivable [1148]. Cultivated [1804, 1845]. Cultivation [1018, 1089, 979, 568, 66, 1208, 933, 289, 654, 497]. Culturability [904]. Culturable [973, 876, 1229]. Culture [1089, 303, 882, 210, 534, 377, 325, 1406, 720, 383, 1802, 1777, 427, 1170, 227, 49, 398, 79, 1666, 369, 791, 1422, 946, 553]. Culture-Independent [1089]. Cultured [1374, 1448, 1898, 1174, 1011, 321, 424, 279, 356, 585, 1489].

Cultures [169, 1009, 682, 1195, 1322, 396, 241, 725, 122, 503, 939, 149, 355]. culturing [1839]. cumingii [807, 1536, 1502, 1954, 1901, 1867]. Cupped [228, 193, 490, 360]. cupressoides [1633]. Cupric [730]. Current [1126, 856, 575, 588, 976]. Currents [775]. Cusk [1560]. Cusk-Eel [1560]. Cutaneous [164]. Cuticle [1871]. Cuticular [1942]. Cuttlebone [1438]. Cuttlefish [1753]. Cuvier [1058]. cuvieri [1960]. Cuvierian [323, 1216]. CXC [805]. CXC-Like [805]. Cyan [813, 849]. Cyanidium [620]. Cyano [1576]. Cyanobactera [794]. Cyanobacteria [1243, 1623, 615, 1629, 1516]. Cyanobacterial [1419, 336]. Cyanobacterium [1060, 1454, 346, 127]. Cyclases [733]. Cycle [941, 88, 1764, 1738, 429, 89, 1803, 1786, 561]. Cycler [147]. Cyclic [1384, 268, 396]. Cyclin [143, 38, 919, 895]. Cyclin-Dependent [38]. Cycling [1353]. Cyclodepsipeptides [1336]. cydonium [71]. cylindrus [1744]. Cynoglossus [626, 777, 1031, 1252, 825, 1628, 1398, 887, 776, 1298, 1508, 1689, 1733]. cyp19a1a [1931]. cyp1a [1293, 202, 121]. CYP1A3 [22]. Cyprid [1702, 1741]. Cyprini [429]. Cypriniformes [349]. Cyprinodontiformes [93]. Cyprinus [6, 12, 1271, 317, 284, 1211, 972, 1019, 942, 1117]. Cyst [1902]. Cysteine [631]. Cystine [1844]. Cysts [237]. Cytochalasin [491]. Cytochrome [48, 94, 45, 202, 47, 43, 496]. Cytogenetic [1544]. Cytokine [1638]. Cytomegalovirus [700, 728]. Cytomegalovirus-U6 [728]. Cytometry [1793, 257]. cytoplasmic [780]. Cytosolic [829, 1301]. Cytotoxic [1568, 1701, 319, 484, 109]. Cytotoxicity [1486, 1912].

D [160, 978, 721, 900, 1842]. D-Galactose-Induced [1842]. DA0580 [1096]. DA11 [767]. dactylomela [1358]. DAHB1 [1217]. Damage [1187, 646, 180]. Damaged [760]. Damselfish [158]. Danio [275, 938, 36, 152, 247, 972, 1019, 1403, 1588, 1819, 110, 1957, 1078]. Dark [1738, 743]. Data [551, 922, 515, 1401, 1323]. Database [61, 964, 1962, 1114]. Date [1181]. dct [1770]. DD [1363]. DD-13 [1363]. ddRAD [1538]. Deacetylase [82]. Dead [903, 1740]. Deaminase [1015]. Death [1055, 1478, 1436, 1943]. Debaryomyces [555, 3, 31]. Debromohymenialdisine [989]. Decapod [493, 431, 1360]. Decapoda [1891, 908, 356, 676]. Decarboxycitrinin [545]. Decarboxylase [435]. Dechlorination [1351]. Decompression [486]. Decrease [1094]. Deep [652, 1843, 440, 582, 371, 1337, 945, 1174, 1314, 1713, 902, 766, 808, 770, 457, 1160, 641, 1198, 486, 1246, 187, 1353, 1383, 57, 485, 857, 904, 229, 1495, 1933, 1047, 883, 738, 1396, 723, 1113, 992]. Deep-Sea [1843, 440, 582, 1337, 1174, 1314, 902, 766, 457, 641, 486, 1246, 1353, 57, 485, 857, 904, 229, 1933, 883, 1396, 371, 945, 770, 1198, 187, 1047, 992]. Deepest [52]. Defect [1580, 1823]. Defective [162, 231, 1789]. Defects [1485]. Defense [1292, 1835, 1901, 1071, 1947]. Defensin [1953]. Deficient [786, 612]. Deformity [1302]. degradans [1296]. Degradation [1061, 197, 457, 1416, 970, 1023]. Degrade [1363]. Degrading [534, 138, 1410, 1936, 1505, 365, 423]. Dehalococcoides [1351].

Dehalogenase [1150]. Dehaloperoxidase [211]. Dehydration [1213]. Dehydration-Responsive [1213]. Dehydrogenase [382, 1751, 442, 500, 31, 109]. Dehydrogenases [439, 918]. Delayed [1813, 833]. Delivered [105, 1541]. Delivery [1141, 79, 466]. Delousing [1685]. Delta [617]. Delta-6 [617]. Deltamethrin [1290]. DeltaN [622]. Demethylation [1829]. Demographic [551]. Demonostrated [1591]. Demonstrated [1433]. Demonstration [221]. Demosponges [873, 772, 535]. Demospongiae [377, 100, 727, 449, 820]. Dendryphiella [1406]. Denitrification [1811]. Denitrifying [1811, 1137]. Density [1555, 1816, 1504, 1595, 1397, 1563, 1511, 133]. densum [1631]. Dental [662]. denticulata [1291, 1619]. Deoxy [1442]. Dependency [803, 1663]. Dependent [1098, 303, 1354, 1417, 1480, 765, 1133, 918, 1116, 38, 961, 957]. Depletion [1732, 1765, 1948]. Deployable [1796]. Deposition [720, 1813]. Depression [1535, 1577]. Deprivation [806]. Depsipeptides [268]. Depth [918]. Depths [723]. Derbesia [1195]. Derivative [1215, 978]. Derivatives [1576, 1049, 10, 1386]. Derivatization [1371]. Derived [1487, 1340, 1576, 1080, 1330, 828, 1465, 1083, 784, 595, 1134, 1478, 1756, 1094, 1499, 1240, 670, 1438, 879, 1229, 1493, 1492, 1587, 1053, 1285, 1259, 379, 1677, 1778, 1044, 1488, 1518, 1609, 1557, 492, 123]. Dermal [1002, 1187, 919, 895]. Desaturase [611, 436, 617, 905, 1482, 1383, 1058, 1665, 875, 906, 943]. Desaturases [1345]. Describing [535]. Deseasin [826]. Desiccation [1373, 1459]. Design [1191, 65, 877, 1430, 1125, 715]. Desmodesmus [1199]. Desorption [367, 1319]. Desorption/Ionization [367, 1319]. destructive [983, 956]. Detachment [1131]. Detect [130, 814]. Detecting [460, 179, 1915]. Detection [908, 366, 1060, 1032, 195, 547, 1083, 1164, 1367, 510, 711, 238, 884, 1300, 1673, 21, 752, 424, 73, 300, 686, 432, 1439, 115, 456, 564, 665, 1186, 246, 1935, 1857, 1795, 1907, 1066, 1704]. Detects [1796]. Deterioration [1347]. Determinant [1038]. Determination [777, 1238, 1931, 1004, 735, 1936, 1636, 889, 407, 1262, 1074, 1244, 887, 1427, 1556]. Determination/Differentiation [1238]. Determined [1089, 406, 1721, 96, 450]. Determining [834, 1267, 1176, 1377]. Detoxification [722, 1139]. Detrimental [1450]. Detritus [454]. Develop [1081, 998]. Developed [989]. Developing [69, 721, 30, 1669, 624, 1511]. Development [893, 1243, 975, 928, 303, 1851, 1904, 158, 1731, 1291, 1805, 716, 210, 327, 1334, 1812, 77, 1951, 1404, 1720, 1688, 1816, 1105, 1693, 1506, 550, 1672, 517, 1554, 1654, 238, 372, 480, 916, 1667, 132, 1366, 834, 513, 378, 1020, 1498, 1590, 1884, 508, 1194, 143, 59, 852, 1739, 644, 300, 1634, 1784, 1952, 1854, 1876, 239, 813, 53, 1616, 241, 1541, 627, 669, 515, 449, 67, 249, 1512, 1958, 1959, 1154, 1013, 530, 1683, 1682, 347, 636, 343, 548, 1114, 467, 1440, 624, 243, 1717, 1791, 648, 796, 898, 931, 1204, 1716, 1152, 1203, 1719, 926, 1293, 1235, 1935, 1915, 1446]. Development [895, 1795]. Development-Related [624]. Developmental [1804, 1812, 1118, 36, 387, 604, 1392, 1309, 1244, 1716]. Developmental-Stage-Specific [1118, 387]. Developments [344]. DGGE [732]. DHA [611, 1603, 1665, 1573, 851, 1496]. Diacarnus [1018]. Diagnosis [1705]. Diagnostic [96, 1785]. Diatom [790, 757, 894, 27, 1221, 1128, 1771, 1096, 1856, 1930, 1597, 1640, 1896, 1663, 1203]. Diatom-Specific [1597, 1640]. Diatoms [32, 1744, 53, 1824, 1263]. dicentrarchi [1072]. Dicentrarchus [197, 392, 520, 539, 328, 716, 1931, 111, 1830, 1864, 253, 1512, 906, 345, 1388, 1362]. Dicer [1671]. Dicistronic [69]. Dictyopteris [1409]. Dictyota [1322]. Dictyotales [1409, 1322]. Diel [127, 1602]. Diesel [1477]. Diet [197, 454, 1955, 854, 1531, 676, 1253, 1594, 1681]. Diet-Induced [1594, 1681]. Dietary [1904, 843, 1941, 1844, 1180, 600, 1202, 1859, 1819, 1695, 1573, 1947, 1496, 1895, 1078]. Diethylene [1330]. Diets [934, 1163, 1378]. Differ [821]. Difference [1676, 1520, 1937]. Differences [1697, 1290, 609, 1600, 1540]. Different [1585, 1904, 274, 1193, 1812, 291, 1764, 1026, 1750, 1779, 1103, 98, 285, 632, 1540, 794, 1180, 87, 1824, 1250, 918, 1610, 1130, 1489, 1460, 1490, 1110, 864, 1298, 522, 1612, 1946, 1690]. Differential [1287, 662, 1898, 557, 1402, 1376, 1283, 743, 1250, 1758, 759, 847, 650, 1709, 681, 597, 1939, 907, 1523, 1014]. Differentially [1068, 866, 1721, 558, 428, 89, 380, 1957, 1521, 899, 1412]. Differentiate [1578]. Differentiation [1804, 1238, 551, 770, 812, 1763, 863, 1571, 1628, 1834, 1074, 1428, 1277, 1079, 1509, 1443, 1500, 1427, 775, 1414]. Differentiation-Related [1277]. Differently [971]. Diffusion [1208]. DIGE [1078]. Digest [76, 213]. Digestibility [1900]. Digestion [1061, 1312, 900, 362, 379]. Digestive [1260, 415, 1941, 1745, 932, 1100]. Digests [326]. Digital [1413]. digitifera [1506, 765, 1617]. Dihydroquinolin [1259]. Dihydroquinolin-2 [1259]. Dihydrostilbene [1430]. dihydroxyphenoxy [863]. Diindol [1276]. Diindol-3-ylmethanes [1276]. Diketopiperazine [410]. Diketopiperazines [1189]. dimensional [730]. Dimethyl [24]. Dimorphic [1637, 1275]. Dimorphism [1628, 1644]. DING [1749]. Dinoflagellate [1210, 508, 837, 1188, 368, 725, 1436, 746, 525]. Dinoflagellates [1095, 510, 448, 318, 334]. Dinophyceae [480, 206, 467]. Dinophysis [884, 467]. Dioxide [1806]. Dioxin [177, 863]. Dipeptides [396]. Diploid [1357, 1305, 526, 680, 852, 1092]. Diploids [1031, 491]. Dipstick [1795]. Direct [148, 1367, 21, 1224, 1642, 1166]. Directed [987, 934, 631, 1289, 640, 560]. Directly [1000]. Disaccharides [1265]. Discharges [1548]. Discodermia [1089, 1140]. Discodermolide [1140]. Discolored [1451]. Discovery [928, 1290, 1739, 1547, 1639, 904, 974, 1127, 580, 1561]. Discrimination [569, 236, 257, 113, 516]. discus [816, 357, 526, 570, 577, 1446, 609, 1446]. Disease [1807, 1182, 286, 717, 1941, 1303, 181, 802, 810, 167, 1375, 1510, 714, 1564, 130]. Disease-affected [714]. Diseased [1879]. Diseases [869, 1179, 1650]. Disentangling [1686]. Dismutase [339, 631]. Dispersing [301]. Display [818, 761, 1097, 759, 747, 1792, 998]. Displayed [1308]. Displays [1485]. Disposal [445]. Disrupt [343]. Disruption [1476, 915, 1044, 1762, 1507, 1166]. Disrupts [1892]. Dissection [1642].

Disseminated [1107]. Dissimilarity [1520]. Dissimilatory [371, 641]. Dissociation [918]. dissoluta [1140]. Dissolved [284, 227]. Distal [1331]. Distance [1226]. Distant [1493, 1492, 1587, 1677]. Distinct [262, 1185, 772, 1040, 1578, 751]. Distinctions [381]. Distinguishing [250, 1912]. Distorted [1282]. Distributed [1622]. Distribution [598, 842, 1351, 641, 1621, 190, 794, 917, 455, 697, 1800, 1236, 1882, 1332, 1029]. Diterpene [58]. divaricata [1409]. Divergence [1212, 937, 1414, 1828]. Divergences [1013]. Diverse [1264, 757, 1421, 1224, 1137, 788, 789]. diversicolor [1678, 1299, 1694, 1027]. Diversification [1315]. Diversified [1332]. Diversity [533, 809, 1228, 1175, 1270, 652, 1291, 973, 1032, 34, 371, 641, 290, 1779, 17, 96, 193, 375, 411, 52, 633, 904, 206, 1229, 1054, 1659, 1845, 707, 1467, 988, 1256, 1218, 1440, 1082, 782, 1891, 1323]. Division [1454, 838, 1735]. Division-Associated [1735]. Dmc1 [1546]. Dmrt1 [1734]. dmrt1a [1931]. dmrt1b [1931]. DNA [496, 366, 675, 56, 463, 144, 80, 22, 768, 1330, 1557, 1701, 19, 719, 219, 504, 85, 327, 1491, 317, 284, 1866, 406, 514, 565, 1764, 1111, 661, 609, 460, 27, 451, 1305, 646, 372, 425, 462, 448, 251, 439, 17, 193, 226, 694, 834, 411, 381, 73, 247, 98, 357, 1498, 825, 508, 1814, 24, 245, 305, 1952, 294, 633, 208, 524, 587, 529, 493, 180, 45, 202, 125, 188, 432, 619, 260, 865, 1541, 1273, 600, 78, 1593, 229, 466, 1890, 1959, 1154, 334, 221, 151, 214, 347, 577, 191, 1709, 4, 108, 266, 374, 437, 960, 1079, 1829, 348]. DNA [422, 1921, 1112, 474, 1833, 1915, 751, 136, 302, 1226, 1861, 200, 1582, 1323]. DNA-Vaccinated [514]. DNAs [36, 5]. DNase [41, 957]. DNase-like [41]. DNMT [1498]. Do [985, 775]. Docosahexaenoic [414, 436, 370, 549, 227, 1603, 562, 542, 851]. Docosahexaenoic-Acid-Producing [227]. Dodecamer [742]. Does [605, 919]. dofleinii [1216]. Dollar [944]. Domain [527, 461, 1308, 1569, 1597, 1640, 334, 506]. Domains [1289, 1686, 1520, 1616, 1257]. domestic [516]. Domesticated [290]. domuncula [396, 311, 26]. Dop2 [1683, 1682]. Dop2/Invertebrate [1683, 1682]. Dop2/Invertebrate-Type [1683, 1682]. Dopamine [1683, 1682]. Doped [1308]. Dormant [718]. Dosidicus [1828]. Double [1207, 921, 560]. Double-Fluorescence [560]. Double-Stranded [921]. Doubled [870]. Doubling [1400]. Doubly [1867]. Dover [81]. Down [1911, 278, 1878]. Down-regulation [1911]. Draft [1691, 1875, 1396]. Drainage [312]. DREB [1213]. Dressing [1920]. dried [468]. Drive [775]. Driven [1724, 700]. Driver [893, 1568]. Droplet [1413]. Droplets [1205]. Drought [1184]. Drug [1691, 41, 1801, 1469, 1685]. Drug-Resistant [1691]. Drugs [588, 1464]. Drum [316, 1081, 1550]. dsRNA [1526, 1194, 1266, 712, 693, 930]. dsRNA-Expressing [1526]. dsRNA-Mediated [1194]. Dual [1730, 1888, 1765]. Dual-Peak [1888]. Dually [536]. Duct [1228, 656]. dumerili [1695]. Dunaliella [737, 835, 544, 1666, 695]. Dunker [931]. During [613, 197, 1189, 1248, 1757, 1680, 1477, 1822, 1291, 1658, 1805, 716, 1417, 139, 1688, 568, 1887, 550, 1954, 517, 238, 218, 486, 1498, 245, 1739, 644, 1392, 1858, 1338, 146, 1512, 1958, 1013, 1679, 847, 636, 89, 1218, 380, 1652, 1818, 446, 681, 855, 996, 1716, 546, 1684, 1905, 705, 275, 1628, 1815, 1462, 1429]. DUSP [1872]. Dust [976]. dUTPase [643]. Dwelling [502, 1542]. Dwells [656]. Dye [1907]. Dynamic [1820, 1938, 1491, 666]. Dynamics [537, 1680, 1528, 1498, 1273, 1462, 933]. Dysgenesis [888]. Dysidea [739]. Dystrophin [197].

Ear [1392]. Early [235, 1805, 1931, 1688, 735, 1149, 1498, 852, 1834, 854, 1154, 1013, 1683, 1682, 1087, 149, 1218, 1462, 820, 1500, 833]. Earning [588]. Earthen [1697]. East [265]. Eastern [1816, 216, 583, 130, 96, 529, 1881, 1440, 479, 648, 763, 447, 1714, 1157]. Easy [1551]. Echinodermata [323, 87, 102, 824]. Echinoderms [146]. Echinoidea [824]. Echinolittorina [1204]. echinulatum [1656]. Echiuran [914, 1042]. Ecklonia [702, 911, 1133]. Eco [1169]. Eco-friendly [1169]. Economic [1550]. Economics [968]. Ecosystem [1339]. Ecotoxicity [1368]. Ecotoxicology [400]. Ectocarpales [1390]. Ectoine [890]. Ectoines [1607]. Edge [1582]. Edible [1055, 1699]. Edited [1775]. Editing [1334, 1613, 1519, 1316, 1808, 1696, 1606, 1825]. Editor [272]. eDNA [1866, 1551]. edulis [114, 1158, 814, 98, 1107, 297]. Edwardsiella [1832, 1721, 1456, 1471, 898, 1783, 1950]. Edwardsiellosis [1437]. EDX [763]. Eel [48, 94, 169, 1922, 1560, 1661, 1703, 45, 202, 47, 43, 446, 1783]. Eels [1832, 1073]. Effect [1585, 1270, 281, 377, 317, 1866, 1009, 734, 1337, 950, 770, 1802, 1777, 1297, 646, 913, 721, 1671, 1434, 1797, 854, 1620, 925, 1359, 1390, 591, 1422, 895, 1865, 1400, 1895, 1350, 1417]. Effective [316, 1519, 700]. Effectiveness [1704, 1913, 1153]. Effects [1287, 702, 1831, 197, 328, 1723, 762, 976, 598, 101, 843, 1408, 489, 509, 808, 1187, 1838, 659, 205, 838, 812, 5, 469, 778, 1844, 1956, 1529, 1580, 1365, 1479, 1046, 1450, 1882, 957, 1102, 1937, 1662, 1559, 804, 1891, 1515, 1696, 1897, 1947, 389, 1927, 594, 1900, 561, 1350, 497]. Efficacy [1368, 1548]. Efficiency [647, 941, 205, 867, 1645, 875, 429, 728, 817, 1166]. Efficient [709, 559, 1701, 1334, 1613, 1817, 1268, 1856, 1930, 1214, 1765, 1203, 1944, 715]. Efficiently [1718]. Effluent [691]. EGFP [1425]. Egg [152, 1347, 1097, 771, 1362]. Eggs [1358, 468, 561]. Eggshell [33]. Eicosapentaenoic [414, 269, 436, 1231, 542]. Eisenia [1055]. Electrical [1281]. Electrochemiluminescence [1066]. Electrochemiluminescence-Molecular [1066]. Electrode [1131, 1281]. Electron [763]. Electrophoresis [730]. Electroporation [1896, 1316, 1606]. Electrospray [85]. elegans [804]. Element [1213, 248, 1466]. Elements [1103, 9, 846, 1769, 1891, 1389]. Elevated [611, 1356, 1649]. Elevates [1457]. Elicits [930, 1943]. Eliminates [565]. ELISA [21, 1439, 243]. ellipsoidea [254]. Ellisellidae [458]. Elongase [414, 436, 1132, 795, 755]. Elongases [819, 1132, 875]. Elongation [34, 95]. elongatus [1692]. ELOVL2 [819, 1603, 1747]. ELOVL2-like [819]. ELOVL5 [819, 1603]. ELOVL5- [819]. ELR [805]. Elsaied [601]. Elucidation [242, 1608]. EMB24 [1801].

Embryo [1317, 303, 1343, 1087, 149, 1139, 482, 718]. Embryogenesis [275, 1035, 1392, 546, 705]. Embryonic [539, 1851, 1424, 610, 550, 238, 53, 1958, 1942, 42, 314, 1152, 862, 926, 1500, 1497]. Embryos [278, 596, 216, 1220, 228, 438, 256, 1328, 721, 5, 1149, 30, 1244, 833, 1166, 1606]. Emerging [1247, 740]. Emiliania [320, 953, 913, 786, 1261, 417, 44]. Emission [1888]. Emissions [1432]. emphaeus [481]. Employing [709]. ENA [950]. Enantioselective [711]. Encapsulated [1734, 466, 1025]. Encephalitis [1063]. Encoded [1347]. Encoding [414, 141, 91, 19, 764, 1353, 706, 301, 194, 83, 970, 536, 487, 1878, 1044]. Encrusting [535]. Encysted [718]. End [776, 655, 1740]. Endangered [1866, 1718]. Endemic [1664]. Endless [939]. Endo [1116]. Endo-4S-iota-Carrageenan [1116]. Endocrine [774]. Endogenous [1321, 971, 1221, 1162]. Endolytic [1289]. Endonucleases [910]. Endopeptidase [900]. Endoplasmic [1425]. Endoskeletal [573]. Endoskeleton [674]. Endosymbioses [502]. Endosymbiosis [401]. Endosymbiotic [385, 351, 1188, 1205]. Endothelin [1270]. Ends [474]. Ene [1576]. Ene-Reduction [1576]. Energy [893, 1822, 1011]. Engineered [445]. Engineering [1444, 708, 1599, 1496]. Engraulidae [302]. Engraulis [302]. Enhance [1410, 1364, 1094]. Enhanced [1772, 1182, 286, 1295, 67, 1666, 729, 1859, 348, 1549]. Enhancement [120, 1866, 1372, 1373, 631, 1051, 199, 410, 1766, 961]. Enhances [1213, 1091, 1955, 1459, 1522, 1090, 1695, 1044, 1524]. Enhancing [1406, 1839]. Enigmatic [797]. Enlargement [471]. Enolase [1792]. Enope [1618]. Enriched [1012, 1072, 249]. Enriching [1562, 815]. Enrichment [1506, 1070]. ensis [681]. Enterobacter [1635, 469]. Enterococci [590]. Entire [613]. Envelope [1333, 1307, 1313]. Environment [101, 1570, 1351, 1125, 1840, 984, 811]. Environmental [596, 235, 155, 1866, 291, 739, 33, 1074, 1432, 201, 1786, 1515, 1897, 1915, 800, 505, 1803]. Environmentally [1724]. Environments [1802, 1777, 752, 485]. Enzymatic [702, 911, 1772, 1329, 1406, 1641, 646, 1730, 1520, 576, 978]. Enzyme [408, 1929, 1289, 1941, 1311, 1015, 1359, 1306, 362, 883, 26, 1778]. Enzyme-Assisted [1306]. Enzyme-Modified [1359]. Enzymes [1312, 1248, 1465, 1043, 628, 379, 773, 970, 787]. EPA [611, 1665, 1573, 1496]. Epicoccum [1006]. Epidermal [830]. Epigenetic [1083, 1587, 1959, 1780]. Epinephelus [533, 1734, 412, 1300, 604, 627, 333, 1651, 1527, 1542, 1833, 1565, 635, 1627, 1626]. Epiphytic [289]. Epipodial [1370]. Epithelial [951]. Epithelioma [429]. Epoxide [711]. EPRE [752]. EPS [261]. ERC [1827]. ERC-38 [1827]. Erie [774]. Erimacrus [675]. Eriocheir [785, 1249, 1818]. eriocheiris [1907]. Eriphia [1286]. ERK [1249, 1601]. EROD [420]. Errata [271, 601]. Erratum [287, 329, 1333, 1019, 983, 376, 397]. Erythrobacter [711]. ES-like [669]. Escherchia [392]. Escherichia [511, 766, 1442, 1547, 1639, 755, 749, 282]. EseD [898]. Esfenvalerate [1340]. Especially [1392]. Essential [1645, 1603, 1058]. EST [61, 1081, 964, 59, 1093, 515, 704, 1033, 648, 751, 724]. Established [1005, 1950]. Establishing [1657]. Establishment [61, 972, 1019, 1096, 1263, 635]. Ester [1330]. Esterase [872, 1933]. Esterases [1233, 784]. Esters [383]. Estimated [609]. Estimates [1391]. Estimating [647]. Estimation [1551]. Estradiol [1381, 673]. Estrogen [276, 1069]. Estrogens [33]. ESTs [975, 861, 964, 877, 1034]. Estuaries [605]. Estuarine [1351]. Estuary [1054]. ethanone [1053]. Ether [935]. Ethinylestradiol [1609]. Ethyl [1699]. Ethylenediaminetetraacetic [457]. Etychro [1442]. Eucheuma [402]. Eucyclogobius [186]. Eudistoma [1256]. Eukaryotes [917]. Eulachon [495]. Eunicea [759]. Euphausia [413, 600]. Euphyllia [1205]. European [328, 716, 1931, 1830, 1832, 1864, 814, 558, 1512, 1077, 960, 1783, 1388, 1458]. Eurotium [1545]. Euryarchaeote [246]. Euryhaline [138, 31, 1342]. Evaluate [1831, 1548, 208]. Evaluated [1464]. Evaluating [1866]. Evaluation [1415, 1704, 1317, 903, 1560, 1898, 843, 1032, 1486, 2, 1816, 1550, 239, 1670, 1430, 823, 1483, 1580, 79, 1382, 1656, 996, 1865]. evanescens [1143]. Event [1805]. Events [1956, 1758]. Evidence [395, 148, 551, 734, 850, 481, 960, 763]. Evolution [1271, 1004, 439, 1409, 1863, 1120, 701, 1332, 1534, 1106, 108, 623, 1891, 1360, 1117, 1684]. Evolutionary [768, 923, 1869]. Ex [1359, 933, 1294, 760]. Exaiptasia [1574]. Examination [13]. Examined [411, 732]. Example [57, 71, 431]. Excision [174]. Exclusive [111]. exertia [652]. Exhibit [240]. Exhibits [1459]. Exiguobacterium [1173]. Exist [1703]. Existing [1323]. EXO [1795, 924, 1795]. Exo- [924]. Exocellular [396, 512]. Exogenous [1243]. Exolytic [1289]. Exome [1672]. Exometabolites [1060]. Exon [1672, 1381]. Exopolysaccharide [1190, 1365, 389]. Exopolysaccharides [485, 198]. Exoproducts [1742]. Exosome [1536]. Exosomes [1508]. Expanding [904]. Expansin [1410]. Expansion [551, 481]. Experimental [1585, 975, 1625, 1470, 214, 170, 960, 1829]. Explained [909]. Explant [946]. Explants [882]. Explore [467, 1960]. Exploring [1961, 1743, 1071]. Exposed [537, 1290, 737, 222, 1288, 940, 1617, 1356, 1523]. Exposure [699, 98, 1361, 1042, 558, 1890, 930, 673]. Expressed [1104, 807, 1068, 61, 866, 256, 785, 1721, 916, 134, 584, 62, 508, 1194, 853, 704, 1076, 698, 693, 636, 426, 548, 428, 89, 624, 380, 1957, 746, 417, 931, 1521, 899, 405, 1412, 920]. Expresses [885]. Expressing [1886, 217, 1094, 1526]. Expression [443, 527, 1287, 611, 559, 699, 1243, 1807, 382, 1448, 1558, 1331, 543, 392, 328, 897, 606, 1918, 1357, 1481, 1676, 1938, 275, 135, 1843, 1085, 1724, 1329, 764, 716, 270, 1167, 1035, 1931, 750, 69, 291, 1118, 452, 119, 829, 651, 1288, 1252, 36, 1868, 1693, 1212, 1005, 639, 276, 1380, 1661, 1703, 27, 438, 909, 951, 1750, 461, 1165, 1646, 517, 557, 238, 254, 604, 309, 1242, 1366, 5, 105, 574, 1675, 98, 1402, 1376, 1457, 1416, 1498, 508, 1163, 1628, 245, 519, 630, 852, 1094, 1398, 644, 1275, 230, 203, 1540, 558, 1283, 1239, 745, 1122, 1188, 189, 813]. Expression [1012, 743, 1338, 1250, 706, 755, 1611, 1403, 95, 140, 1425, 33, 726, 1610, 1072, 544, 164, 919, 1479, 79, 1046, 1258, 1546, 1758, 466, 1909, 896, 940, 1592, 906, 847, 360, 891, 749, 1370, 1629, 149, 690, 225, 1467, 650, 695, 1880, 773, 1781, 1694, 60, 536, 310, 1102, 1327, 361, 624, 1462, 1602, 1109, 1253, 1272, 324, 1509, 487, 744, 1110, 1599, 1829, 1069, 681, 597, 996, 1716, 560, 1473, 1765, 1203, 1939, 1136, 1298, 1766, 625, 673, 1911, 1494, 907, 833, 895, 388, 546, 593, 738, 1606, 1684, 1342, 705, 1310, 1690, 779, 1014, 1870, 494, 1184]. Expressions [1904, 274, 1502, 1819]. Extends [1749]. Extinguishing [730]. Extracellular [1028, 1444, 1241, 1321, 383, 1748, 477, 261, 1173, 1675, 692, 1128, 1726, 1122, 917, 122, 900, 1006, 1196]. Extract [702, 911, 1000, 572, 1882, 1819, 1699, 1842]. Extracted [991, 1529, 687, 150]. Extraction [1701, 448, 464, 1858, 188, 1711]. Extracts [1444, 1063, 354, 489, 152, 646, 1434, 1670, 925, 1090, 174]. Extraordinary [1801]. Extreme [1346, 1570, 1091, 485, 904, 1659]. Extremely [955, 1584]. Extremes [596].

F [42, 314, 1351, 1144, 109]. F. [853]. F1 [317]. F2 [317, 284]. fabp2 [1280]. Facile [1593]. Facilitates [1787]. Facilitators [120]. Factor [401, 274, 275, 737, 999, 270, 147, 1288, 1213, 1057, 438, 1030, 194, 1024, 149, 121, 519, 34, 95]. Factor- [1057]. Factor-1 [34, 95]. Factor-Beta [999]. Factor-I [274, 275]. Factors [64, 308, 155, 849, 1473]. Facultatively [1050]. FADH [1098]. Families [1068, 821, 680, 1092, 1423, 214]. Family [478, 731, 85, 412, 661, 1352, 1863, 1051, 1223, 1106]. Far [1558]. Far-red [1558]. Farmed [1885, 1149, 555, 1862, 1531, 1130, 1218, 1272]. Farming [822]. Farnesyldiphosphate [759]. farreri [1543, 1092, 1151, 1474, 131, 775, 655, 710, 688]. fas [1198]. fasciata [773, 970, 576, 978]. fasciatus [1720, 1834, 1112]. fascicularis [41, 1480, 459]. Fast [1596, 1687, 188]. Fast-growing [1687]. Fasting [1658, 1473, 1604]. Fat [1804, 197]. fat1 [1207, 1496]. fat1-Transgenesis [1496]. fat2 [1207]. Fatty [414, 1227, 1760, 757, 1372, 941, 436, 370, 905, 1225, 1457, 1482, 1195, 1058, 819, 1132, 1207, 616, 983, 956, 791, 1040, 1537, 1155, 1747, 804, 943, 497, 795, 1808]. faveolata [139]. Favoring [1957]. favus [1082]. Faxonius [1757]. Features [1568, 731, 771, 1809, 1047, 1284]. Fed [1163, 1180, 1531, 860, 1253]. Feed [893, 1070]. Feeding [1364, 241, 1813, 1828, 1831]. Feeds [373, 466]. Feet [952, 1327]. Female [1697, 626, 1913, 1344, 825, 1274, 1683, 1682, 776, 1677, 1521, 1908, 1500, 1031, 1493, 1492, 1632]. Female-Specific [626, 825]. Females [1304]. Fenneropenaeus [1165, 378, 519, 657, 1159, 1674, 1894, 580, 744, 855]. Fensholt [839]. Feral [290]. Fermentation [1080, 949]. Fermentative [469]. Fermented [1607]. Ferric [457]. Fertile [1588]. Fertility [1105]. Fertilization [228, 668]. Fertilizing [736, 93]. fgf10a [1235]. fgf10b [1235]. Fiber [337, 1473, 942]. Fiber-Related [942]. fibrillar [1002]. Fibroblast [149, 42]. Fibroblasts [1002, 1187, 863, 314, 1655]. Fibrosarcoma [591]. ficiformis [449]. Field [1545, 769, 1463, 195, 1796]. Field-Deployable [1796]. Fight [1650]. Filamentous [359, 1060, 1623]. Filaments [269]. filiformis [87]. Fillet [1367]. Film [402]. Films [980]. Filter [337]. Fin [1931, 1578, 912]. Fine [419, 1553, 1595, 775, 1556]. Fine-Scale [775]. Finfish [922, 1294]. Finger [1220, 1316, 1023, 1176, 1166]. Fingerprinting [412]. finmarchicus [1040]. First [401, 783, 219, 434, 1247, 1502, 1595, 57, 1222, 1397, 1493, 1492, 824, 847, 1850, 1778, 1542, 1813]. Fish [527, 255, 614, 272, 699, 197, 1653, 64, 373, 22, 281, 647, 563, 1334, 61, 147, 173, 971, 514, 565, 88, 1688, 104, 177, 159, 1732, 1341, 438, 1280, 163, 339, 955, 248, 1121, 1142, 1206, 279, 1163, 630, 850, 49, 869, 57, 128, 181, 185, 713, 349, 1665, 1953, 1052, 164, 79, 165, 1074, 1088, 466, 399, 1650, 162, 231, 1013, 170, 282, 428, 1472, 312, 167, 174, 277, 898, 1677, 422, 179, 1813, 1065, 1389, 1551, 200, 1420, 920, 1787, 1825, 1948, 723, 480, 165, 688]. Fish-Gut [1650]. Fish-Pathogenic [1121]. Fish-Specific [955]. Fishes [184, 694, 180, 57, 182, 823, 175, 1937, 1542]. Fishing [472]. Fishmeal [1900]. Fistulifera [1771, 1096, 1824, 1856, 1930]. Fitness [1549]. Five [85, 461, 1315, 1048]. Five-Domain [461]. Fixed [468]. FJN [905]. FJN-10 [905]. flabellata [1633]. Flagellar [765]. Flanking [248, 92]. Flat [814, 1107]. Flatfish [4]. Flatworm [1800, 1949]. Flatworms [1621, 1937]. flavescens [774]. Flavobacteriaceae [293]. Flavobacterium [749, 1186]. Flavone [760]. Flesh [1620, 1679]. flesus [558]. flexibilis [803, 1011]. Flocculation [1090]. Florida [406, 508]. Flounder [63, 1282, 61, 1343, 103, 1010, 142, 254, 1945, 435, 1163, 1456, 1862, 558, 1665, 78, 244, 214, 1201, 522, 482, 546, 592, 561]. Flow [605, 1793, 1796, 739, 257, 1261, 1795]. Fluid [992, 299]. fluitans [50]. fluminea [1865]. Fluorescein [523, 1748, 257]. Fluorescein-4-Isothiocyanate-Labelled [1748]. Fluorescein-Isothiocyanate-Labeled [257]. Fluorescence [578, 1193, 568, 1888, 165, 257, 233, 479, 560]. Fluorescent [1558, 645, 666, 586, 813, 629, 1888, 726, 1214, 707, 690, 60, 849, 264, 1539]. Fluorescently [1950]. Fluorodeoxyuridine [1227]. Foam [553]. Focus [1223]. Focused [1597, 1640]. Follicle [245]. Follicle-Stimulating [245]. Followed [1688]. Following [1805, 750, 1715, 1528, 557, 1361, 189, 164, 466, 67, 1298, 116]. Food [1287, 893, 1312, 1772, 377, 445, 1445, 671]. Foot [824]. Footed [543]. Foreign [105, 24]. Forest [189, 79]. Form [751]. Formate [469]. Formation [1742, 613, 1536, 173, 258, 682, 1469, 1506, 1871, 951, 1643, 1502, 1954, 1901, 659, 838, 913, 844, 1352, 850, 189, 562, 1597, 1640, 1616, 1423, 607, 1217, 42, 1036, 1818, 1836, 1196, 594, 1749]. Formed [1460]. Forming [873, 794, 1623, 368]. Forms [1904, 43]. Formylglycine [1116]. Formylglycine-Dependent [1116]. forskali [323]. fortis [1215]. Forward [1691]. Fosmid [902, 776, 655, 330]. Fouling [1368, 101]. Found [1744]. Founders [726]. Four [184, 1115, 964, 129, 186, 108, 522, 305]. Fourier [85]. Fourth [1799]. FpVtg [1894]. Fraction [952, 1923]. Fractions [1415]. Fragilariopsis [1744]. fragile [1245]. Fragment [709, 196, 292, 608, 1785, 4]. Fragments [1557, 1159, 732]. Framework [1954]. France [394]. franciscana [1437, 1942]. Free [1579, 1442, 474]. Freeze [468]. Freeze-dried [468]. Freezing [24]. Frequency [356]. Fresh [200]. Freshwater [807, 1757, 395, 1676, 1938, 979, 1168, 1619, 1170, 30, 732, 1496, 1684]. Freund [1783]. friendly [1169]. fringe [566]. Frog [971]. Frond [871]. Fructofuranosidase [1339]. Fructosyl [452, 192]. Fructosyl-amine [452]. Fructosyl-Valine [192]. Fry [563]. FSH [1541]. FTO [1461]. ftsZ [1454]. Fucans [534]. fucata [1324, 1572, 1836, 1582, 1731, 780, 720, 909, 1643, 659, 844, 1376, 1352, 1423, 1460, 1234, 703, 536, 1036, 656, 388, 594, 1749]. Fucobacter [362]. Fucoglucuronomannan [326, 362]. Fucoidan [812, 293, 365, 423, 379]. Fucoidan-Degrading [365, 423]. Fucoidanase [423]. Fucose [937, 1912]. Fucose-Binding [937]. Fucose-Containing [1912]. Fucosterol [1187]. Fucoxanthin [1143, 1654]. Fucus [1143, 853]. Fugu [176, 194]. Full [1525, 1552, 1712]. Full-Length [1525, 1552, 1712]. fulvescens [947, 1763, 1284]. fulvidraco [1555, 1119, 1275]. fumigatus [1756]. Function [762, 842, 1280, 1763, 1630, 1634, 1459, 28, 697, 1859, 1694, 1573, 1235, 1257]. Functional [430, 1312, 1331, 805, 1822, 999, 1421, 436, 531, 639, 617, 905, 1619, 1751, 785, 254, 1289, 955, 552, 1686, 657, 1670, 1278, 1378, 47, 172, 1963, 771, 874, 906, 1013, 1325, 1277, 1040, 277, 1318, 841, 1420]. Functionality [671]. Functionally [1332]. Functioning [1261]. Functions [1404, 309, 1476, 1770]. Fundibacter [383]. Fundulus [307]. Fungal [1330, 1169]. Fungi [809, 1340, 1270, 359, 973, 1788, 1336, 935, 1240, 792, 1395, 1229, 1056, 1053, 1285, 1101]. Fungus [1028, 1576, 1545, 1080, 1241, 1330, 1321, 1449, 828, 905, 595, 886, 1259, 1006, 1656, 1230, 1518, 1609]. Furcellaran [1306]. Furcellaran-Like [1306]. Fusarium [1241, 1191]. fusca [1084, 1297, 759]. fuscoguttatus [1565]. fusifera [734]. fusiforme [1359]. Fusion [1918, 1380, 1490, 1789]. Future [856, 1691, 1247, 170].

G [1725]. G5 [1329]. G7a [781]. GA [209]. GA/GT [209]. gab3 [1794]. Gadidae [463]. Gadiformes [56, 463]. gaditana [815]. Gadus [975, 928, 335, 1035, 1070, 1149]. Gag [20]. Gal4 [303, 833, 1273, 1136]. Gal4-Gene-Dependent [303]. GAL4/UAS [833, 1273, 1136]. Galactose [1842]. Galactosidase [845]. Galaxea [41, 1480, 459]. galbana [1768, 355]. Galicia [126]. Galician [665]. galloprovincialis [1585, 1288, 1408, 1548, 954, 499, 665, 114, 988]. Galveston [1351]. Gametogenesis [1498]. gammarus [900, 1100]. Ganglia [140]. Ganglion [1055]. GAPDH [1013, 998]. Garfish [236]. garvieae [1292]. Gas [802, 810]. gasar [1845]. Gasterosteus [1041]. Gastric [1312]. Gastroenteritis [279]. Gastrointestinal [555]. Gastropod [1172, 1699]. Gastropoda [1113, 444, 1022]. Gayralia [1568]. GBLUP [1456]. GC [783]. GCC [1247]. GCRV [1655, 1846]. GDSL [872, 1933]. GDSL/SGNH [1933]. Gecarcoidea [1465]. Geitlerinema [1060]. Gel [313]. Gelatin [1772, 1929]. gelidicola [1200]. Gelidium [1530]. Gelling [1410]. gemmata [1033]. Gene [1287, 611, 969, 559, 48, 94, 430, 1487, 1243, 528, 1886, 1448, 278, 928, 1331, 303, 605, 1904, 606, 22, 1918, 196, 1357, 141, 91, 1481, 1676, 1938, 275, 19, 621, 148, 1775, 1724, 764, 716, 1167, 1035, 582, 315, 750, 69, 291, 126, 1295, 1740, 651, 1961, 565, 1288, 1220, 292, 1722, 1693, 103, 107, 1005, 1045, 1454, 639, 387, 905, 1380, 1661, 27, 550, 909, 951, 1750, 217, 1091, 1299, 1339, 1170, 1743, 557, 238, 1221, 254, 1945, 161, 604, 1242, 248, 1381, 1671, 93, 105, 435, 1675, 1402, 1476, 1482, 1645, 1956, 508, 1163, 1628, 640, 1094, 1398, 1533, 1532]. Gene [1603, 644, 837, 285, 99, 664, 230, 203, 1283, 1122, 524, 587, 102, 1012, 57, 398, 182, 743, 1547, 1639, 1459, 1597, 1640, 706, 761, 755, 349, 1378, 47, 157, 1072, 544, 1051, 164, 1385, 1479, 924, 199, 449, 1316, 1546, 1214, 466, 1909, 896, 194, 399, 940, 1592, 1024, 693, 847, 636, 891, 1370, 1629, 282, 1184, 690, 1467, 988, 1819, 650, 191, 84, 695, 1461, 773, 60, 89, 110, 1327, 1602, 1109, 1177, 1253, 982, 1272, 324, 744, 246, 1069, 681, 597, 996, 1716, 937, 714, 1747, 1044, 474, 1762, 1696, 1176, 625, 732, 92, 1507, 1494, 907, 833, 895]. Gene [546, 738, 1166, 1157, 1825, 1342, 1458, 1539, 1014, 1851, 829, 1338]. Gene-Edited [1775]. Gene-Gun-Mediated [110]. Gene-Targeting [1671]. Genein [1295]. Genera [193, 108, 1394, 1455]. Generate [1322]. Generated [807, 1858]. Generates [1520]. Generating [238]. Generation [1886, 1290, 1671, 1052, 1493, 1492, 1144, 264, 1293, 474, 1825]. Genes [255, 1227, 807, 543, 897, 1290, 1238, 1798, 1870, 719, 1264, 790, 958, 922, 1068, 1812, 61, 68, 286, 1931, 440, 32, 861, 1118, 1715, 371, 1404, 1754, 211, 1212, 370, 1553, 1373, 1750, 1926, 1502, 1721, 517, 571, 641, 579, 1725, 351, 309, 1198, 873, 1885, 483, 927, 1353, 1335, 1393, 1376, 1476, 1435, 1660, 386, 852, 1566, 1275, 774, 1634, 1863, 558, 1107, 1180, 1269, 1815, 743, 128, 1309, 140, 43, 301, 172, 1207, 318, 1369, 756, 940, 1244, 360, 887, 1108, 1151, 1324, 1428, 428, 83, 970, 350, 536, 310, 1581, 110, 624, 1462, 380, 409, 883, 1957, 1079, 264, 277]. Genes [580, 1036, 1204, 1426, 1586, 1668, 1647, 1710, 1473, 1521, 1712, 1573, 1412, 1565, 1618, 1226, 1446, 1513, 1735, 1874, 1911, 1524, 145, 1254, 1427, 1362, 488, 593, 710, 788, 789, 1086, 1342, 1782, 705, 1601, 1310, 1690, 1396, 496, 304, 1113]. Genetic [709, 533, 675, 605, 37, 1424, 384, 421, 626, 621, 1329, 404, 1282, 563, 34, 77, 1004, 1555, 434, 88, 1111, 1039, 263, 609, 460, 717, 540, 153, 163, 1300, 290, 353, 916, 96, 193, 226, 375, 321, 1366, 1504, 834, 411, 187, 959, 390, 1590, 1543, 825, 59, 570, 1119, 1268, 1784, 208, 1222, 619, 182, 260, 1096, 455, 407, 185, 1856, 1930, 1092, 1397, 1622, 741, 1074, 206, 881, 1493, 1492, 183, 333, 1263, 530, 1845, 1563, 1460, 1490, 1466, 481, 31, 1440, 266, 374, 437, 1204, 1511, 1599, 1391, 1219, 1642, 302, 447, 775, 518, 942, 1414, 1648, 1761, 441, 498]. Genetically [888, 199, 1959, 860]. Genetics [969, 1126, 165, 151, 244]. Genome [1468, 1691, 1557, 1798, 1334, 1404, 1613, 1722, 1525, 1552, 1348, 1519, 1363, 1209, 1299, 1646, 1687, 1146, 1725, 1605, 1103, 1267, 1142, 1206, 1671, 1393, 1936, 1962, 538, 1636, 1409, 1875, 1759, 1347, 413, 1771, 1863, 442, 500, 57, 823, 1052, 165, 521, 9, 874, 846, 1767, 1709, 1781, 1607, 1201, 1375, 1647, 1808, 560, 1542, 1561, 1513, 1874, 1847, 1117, 1497, 1582, 1606, 1684, 1787, 1825, 1379, 1648, 1761, 1387, 1458, 1538, 1627, 1626, 1879, 1396]. Genome-Encoded [1347]. Genome-Scale [1458]. Genome-Wide [1798, 1404, 1348, 1209, 1646, 1146, 1725, 1605, 1393, 1636, 1863, 1767, 1709, 1781, 1375, 1808, 1684, 1379, 1648, 1387, 1538, 1542, 1582]. Genomes [768, 1271, 1355, 1409, 1052, 696, 663, 431, 1034]. Genomic [430, 1098, 1331, 330, 1346, 111, 173, 1720, 1868, 1672, 1305, 142, 1955, 1636, 1456, 1550, 1499, 1634, 203, 1837, 349, 172, 1074, 1493, 1492, 1587, 194, 1659, 166, 1869, 1474, 1850, 1848, 1891, 1034, 1618, 1513, 1564, 1379, 724, 494]. Genomics [1126, 602, 176, 583, 1081, 927, 1073, 168, 696, 1067, 1414, 1575]. Genotypic [352]. Genotyping [458, 1093, 232, 1399, 1642]. Genotyping-by-Sequencing [1399]. Genus [556, 923, 412, 661, 118, 480, 96, 223, 374, 1194]. Genypterus [1560]. Geodia [568, 71]. Geographical [775]. Geographically [1022]. Geothermal [1803, 1786]. Gephyrocapsa [44]. Germ [1183, 1740, 1732, 1435, 1739, 1674, 1244, 166, 1178, 736, 1136, 1500, 1446, 1539]. Germ-Cell [166]. Germline [1500]. Germlings [1083]. Getting [922]. gfp [1250, 1293, 88, 1250, 416, 747, 833, 1122, 264]. GFP-like [416]. GFPxm [586]. GH [327, 145]. GH50 [1692]. GIA [1067]. Giant [395, 1938, 1806, 427, 1170, 1667, 17, 226, 262, 1383, 1542]. Gibel [1557]. Gigabase [1115]. Gigabase-Scale [1115]. Gigartina [1469]. gigas [90, 999, 583, 1417, 1526, 1860, 1279, 228, 1057, 653, 1305, 1821, 1926, 1687, 1030, 821, 390, 1571, 1645, 1660, 1844, 1566, 1017, 491, 493, 1728, 456, 1385, 1172, 1154, 756, 1683, 1682, 347, 360, 1466, 1467, 1534, 1767, 1780, 233, 1356, 1443, 1716, 1391, 1908, 1427, 1342, 1828, 1310, 1690]. Gill [1676, 134, 1402, 1283, 1130, 1185, 714, 1849]. Gills [291, 1815]. Gilthead [1192, 1919, 1118, 1075, 24, 203, 1302, 1462, 1059]. Give [1733]. glaberrima [102, 1338]. glabrescens [1205]. glacialis [701]. Gland [1702, 1741, 1344, 634]. Glands [85]. Glass [337, 982]. Glassin [1616]. glaucus [1080]. Global [876, 1852, 1311, 1102, 1008, 996]. Globally [1130]. Globular [535]. Glossomastix [617]. Glucan [133, 1025]. Glucanase [359, 879, 1044, 924]. glucans [721]. Glucose [90, 1337, 1838, 5, 1914]. Glucosidase [576, 978]. Glucosidases [1329]. glucoside [28]. Glucosylation [576, 978]. Glucosylmannosyl [403]. Glucotriose [1198]. Glucuronorhamnoxylan [1763]. Glucuronosyltransferase [1630]. Glutamate [1904, 730, 109]. Glutathione [11, 649]. Glutathione-S-transferases [11]. glutinosa [830]. Glycerol [31]. Glycerol-3-phosphate [31]. Glycerolipid [403]. Glycidyl [935]. Glycobiology [856]. Glycogen [1419, 90, 620]. Glycol [1330]. Glycoprotein [565, 1333, 1307, 1313, 220]. Glycoproteins [1245]. Glycosaminoglycan [15]. Glycoside [760]. Glycosylhydrolases [359]. Glycosylphosphatidylinositol [656]. Glycosylphosphatidylinositol-Anchored [656]. GM [860]. Gmelin [479, 447]. Gnotobiotic [1388]. GO [1806]. Gobioid [23]. Goby [186]. Golden [1752, 1466]. Gonad [999, 1304, 866, 1057, 378, 1628, 1610, 636, 1581, 1062, 1427]. Gonad-Inhibiting [1304]. Gonad-Specific [1057]. Gonadal [1291, 1754, 1105, 888, 1884, 1739, 1834, 1541, 67, 1512, 1509, 1446]. Gonadosomatic [1636]. Gonads [231, 1428, 1718]. Gonostoma [57]. Goose [1484]. Gorgonian [458, 1259]. gp120 [1333, 1307, 1313]. GPCR [1725]. GPR132 [1393]. GR [1102]. Gracilaria [1374, 806, 1311, 1197]. Gracilariopsis [1793, 1810]. gracile [57, 723]. gracilis [1372, 1663]. gracillimum [1137]. Graft [613]. Grafts [1438]. Gram [974, 1488]. Gram-Negative [1488]. Gram-Positive [974, 1488]. Graphene [1806, 1751, 1593]. GRAS [1642]. GRAS-Di [1642]. Grass [1904, 1870, 222, 1105, 1405, 858, 1655, 1846, 1911, 1003, 1776, 1736]. Grateloupia [1583]. Great [969, 1264, 246]. Green [1243, 1568, 1193, 541, 554, 586, 544, 881, 1214, 1629, 690, 970, 1284, 60, 849, 264, 1599, 901]. Green-Light-Regulated [1243]. Greenlip [1223, 1370]. Greenshell^{T M} [1441]. Grey [260]. Grilsing [1164]. Gross [962]. Group [667, 74, 1927]. Grouper [533, 1734, 1413, 1300, 604, 1541, 627, 333, 1527, 1542, 1833, 1565, 635, 1627, 1626]. Groupers [412]. Grouping [370]. Groups [353, 181, 685, 1828]. Growing [602, 1898, 1687]. Grown [1697, 1153, 1129, 1695]. Growth [1697, 1287, 1210, 303, 1904, 64, 1637, 1357, 1798, 1925, 274, 275, 790, 1596, 1775, 999, 270, 377, 68, 317, 147, 207, 947, 1830, 29, 2, 1039, 1887, 1553, 1057, 517, 1170, 838, 1955, 1300, 803, 254, 1945, 1941, 1242, 727, 867, 1595, 1093, 49, 664, 1480, 407, 1397, 368, 968, 687, 78, 199, 1088, 67, 1959, 1489, 1563, 1490, 149, 535, 1199, 1466, 1208, 1461, 1780, 201, 791, 1090, 1632, 1253, 1957, 1426, 1474, 1537, 1598, 1808, 1473, 1542, 1789, 1719, 1391, 1813, 820, 1565, 1219, 1513, 1714, 1549, 1612, 1900, 1538, 654, 1912, 1014]. Growth-Related [1357, 1039, 1300, 1242, 1595, 1397, 1563, 1466, 1426, 1565, 1014]. Grunt [411]. GSK- [1713]. GT [104, 209]. Guanylate [733]. Guided [1465, 1875]. guilliermondii [1501]. Gulf [537, 1635, 406, 212, 451, 881]. Gum [1630, 1817]. Gun [105, 191, 110]. Günther [1357, 325]. Guppy [1676, 290, 353]. Gut [1291, 1892, 1608, 1620, 1650, 1917, 1905]. Gutless [502]. GWAS [1503, 1510, 1710, 1565]. Gymnocorymbus [726]. Gymnodimine [889]. Gymnodimine-A [889]. Gymnodinium [195, 368]. Gynogenesis [384, 777, 1003]. Gynogenetic [1557, 1343, 40, 526, 680, 850, 852, 1092, 1587, 1735]. Gynogenetically [561]. Gyrodinium [389, 492].

H [314, 646, 1259]. H. [1632]. H2B [885]. Habitat [918]. Habitats [502, 428]. HaCeD [1551]. HaCeD-Seq [1551]. Haddock [596, 295]. haddoni [1229]. Haematococcus [237, 1599]. Haemorrhagic [1068, 1179]. Hagfish [662, 918, 830]. Hair [675]. Hairpin [700]. Hake [419, 551]. Half [626, 777, 1031, 825, 1628, 1092, 1244, 887, 776]. Half-Embryos [1244]. Half-Smooth [626, 777, 1031, 825, 1628, 887, 776]. Half-Tetrad [1092]. Halibut [1004, 745, 854, 1848]. Haliclona [298, 809, 793, 727, 867, 772]. Haliclonamides [298]. halioticoli [83, 84]. Haliotidae [496]. Haliotis [1061, 543, 1002, 35, 1678, 609, 816, 1299, 444, 375, 357, 526, 570, 555, 524, 587, 140, 122, 1236, 577, 1223, 1370, 1694, 450, 1699, 1632, 960, 1446, 150, 1027, 1014]. Haloacid [1150]. Haloarchaea [1575]. Halocarbon [1432]. Halocynthia [719, 268]. Halogenases [1098]. Halogenated [1518]. Halomonas [890]. Haloperoxidases [1432]. Halophilic [1584, 1607]. Halophyte [1213, 1091, 1184]. Halophytophthora [1231]. haloplanktis [801, 512]. Hamilton [1885, 1531, 1959, 1076]. Hammerhead [120]. Hampers [1506]. Handling [1085, 1560, 116]. hannai [816, 357, 526, 570, 577]. hansenii [555, 3, 31]. Haploid [418, 1222, 870, 1497]. Haplosclerida [727]. Haplosporidium [130, 21]. Haplotype [128, 455, 1551]. Haplotypes [1830, 432]. haptonemofera [651, 523, 838, 257, 607]. Haptophyceae [320, 786]. Haptophyta [584]. Haptophyte [651, 1654]. Hard [581]. hardwicki [421]. Hardwired [678]. Hare [129, 299, 55]. Harman [338]. Harmful [1877, 448, 801, 1049, 584, 1593]. harveyi [1922, 217, 218, 1689]. Hatched [23]. hatcheri [834]. Hatchery [1181, 454, 609, 518]. Hatching [486, 1902]. Having [94, 634]. Hawaiian [23, 333]. Hay [320]. HB226069 [1936]. HcKuSPI [1901]. HcN57 [1954]. Head [1012, 1072, 1298, 1790]. Headed [957]. Heads [994]. Healing [1625, 1920, 1823]. Health [139, 222, 1917]. Heat [1910, 1768, 235, 270, 139, 1860, 1961, 1165, 1030, 218, 821, 1571, 116, 1051, 67, 1223, 690, 1729, 1767, 1708, 970, 581, 1735, 1197, 1776]. Heat-inactivated [1165]. Heat-Responsive [1961]. Heat-Shock [218]. Heat-Shock-Induced [67]. Heavy [722, 1693, 1250, 1054, 597]. Hedgehog [343]. helgolandicus [386]. Hellisheidi [1803, 1786]. Hematopoietic [134, 115]. Heme [1728]. Heme-Peroxidase [1728]. Hemibarbus [1951]. Hemocyanin [1255, 1286, 742, 1159]. Hemocytes [807, 750, 256, 7, 1841, 756, 304]. Hemoglobin [938]. Hemolymph [133]. Hemorrhagic [969, 1671, 1088, 1785]. Hepatic [1831, 699, 270, 1301, 1163, 896, 1069, 656, 1813]. Hepatocyte [1024]. Hepatocytes [862]. Hepatopancreas [1291, 785, 1165, 106, 89, 681, 1712, 133, 1849]. Hepcidins [1919]. HepG2 [1174]. Heritability [1391]. Heritable [1176]. Hermaphroditic [1732, 93, 1861, 1874]. Herpesvirus [1417]. Herpesvirus-1 [1417]. heteroclitus [307]. Heterocyst [794]. Heterocyst-Forming [794]. Heterocystous [346]. Heterogametic [1262]. Heterogeneity [642, 316, 226, 266, 1736]. Heterokontophyta [853]. Heterologous [1843, 749]. Heteroplasmy [334]. Heterorhamnan [1568]. Heterosigma [366]. Heterosis [1821]. Heterotrophic [1691, 851]. HEW [596]. Hexactinellid [1616]. Hexactinellida [982]. Hexokinase [5]. Hexoseulose [1442]. HiCEP [907]. HIF [1860, 1288, 1712]. HIF-1 [1860, 1712]. Hif-prolyl [1288]. High [1727, 1228, 1680, 737, 1775, 764, 1555, 1816, 1333, 1307, 1313, 1553, 1297, 1339, 1955, 1300, 1504, 1645, 1533, 1532, 1595, 840, 889, 356, 1096, 1397, 865, 875, 1531, 1563, 1906, 1581, 576, 55, 663, 1511, 1810, 1939, 1561, 133, 612, 907, 355, 1761, 1310]. High-Affinity [1333, 1307, 1313]. High-Density [1555, 1816, 1504, 1595, 1397, 1563, 1511, 133].

High-Efficiency [875]. High-level [764]. High-Mannose [1333, 1307, 1313]. High-pH [1553]. High-Resolution [1300]. High-Salinity [1533, 1532]. High-Temperature-Induced [1939]. High-Throughput [1727, 1228, 1531, 1561]. High-Yielding [576, 355]. Higher [1957]. Higher-Growth [1957]. Highlights [1038]. Highly [1175, 341, 722, 1232, 1143, 916, 18, 819, 1824, 1856, 475, 54, 86, 781, 1574, 1930]. Hilsa [1815]. Hinge [1234]. Hippo [1818]. Hippocampus [1470, 1963]. Hippoglossus [1004, 745, 854, 1848]. Hirame [63]. Histidine [1308]. Histidine-Doped [1308]. Histocompatibility [330, 717, 128, 168, 592, 593]. Histological [90, 1951, 1540, 1849]. Histology [1884]. Histone [885]. Histopathologic [514, 171]. History [985, 87, 1840]. HIV [1333, 1307, 1313, 299]. Hiziki [1434]. Hizikia [1359]. HN2 [761]. HN2-3 [761]. Hokkaido [455]. Holobiont [1765]. Holoenzyme [954]. Holothuria [323, 1866, 1889, 102, 1338, 1216]. Holothuriidae [462]. Holothuroidea [323, 102]. Holoxea [973]. Homarid [372]. Homarus [900, 1100]. Homeobox [386]. Homeostasis [1892, 1573]. Homodiploid [1677]. Homogenate [427]. Homogeneity [316]. Homolog [1756, 690]. Homologs [1354]. Homologue [388]. Homologues [280, 622]. Homology [1133]. Homology-Based [1133]. Homozygous [1182, 870, 422]. Hong [1852, 381]. hongkongensis [1852]. Hormonal [387]. Hormone [1697, 1287, 1637, 395, 478, 274, 68, 317, 1304, 29, 254, 240, 262, 245, 407, 78, 199, 1316, 201, 361]. Hormone-Related [262]. Hormones [85, 1951, 301, 243]. Horse [1385]. Horseshoe [1943]. Hortaea [1843]. Host [1922, 1156, 217, 690, 1471, 1510, 714]. Host-Specific [690]. Hour [228]. Housekeeping [940]. Houston [1351]. Hox [102]. hoxH [488]. HSF1 [1860, 1566]. HSF1/HSP70 [1860]. HSP60 [235]. hsp70 [1250, 1860, 1223, 829, 1942, 1782]. Hsp72 [146]. Hsp90 [116]. HT1080 [591]. Hub [1815]. Human [1444, 1002, 1174, 438, 1187, 5, 1142, 1478, 1485, 319, 181, 687, 1365, 542, 591, 380, 167, 996, 862, 1912]. Humanized [1207]. huxleyi [320, 953, 913, 786, 1261, 417, 44]. Hyaluronate [1246]. Hybrid [163, 1411, 837, 185, 728, 1632, 960, 1389, 1861, 1874, 1387]. Hybridization [578, 647, 568, 251, 165, 1493, 1492, 1587, 428, 113, 409, 233, 479, 1036, 1677]. Hybrids [1821, 852, 1430, 1426, 1677, 1003]. Hydration [1527]. Hydrocarbon [138, 177, 225, 1293]. Hydrocarbon-Degrading [138]. Hydrocarbons [44]. Hydrofluoric [892]. Hydrogen [346, 469, 615, 773, 862, 612]. Hydrogenase [612]. Hydrogenase-Deficient [612]. Hydrogenases [794]. Hydrogenation [1518]. Hydroides [804]. Hydrolase [711]. Hydrolysate [799, 1929, 1730]. Hydrolysates [1698, 1653, 64, 1420]. Hydrolysis [818, 1730, 935, 1501]. hydrolyzing [1339]. hydrophila [998, 1608, 1429, 1790]. hydrophila-infected [1790]. Hydropuntia [980]. Hydrosoluble [978]. Hydrostatic [1400]. Hydrothermal [1843, 440, 1320, 1314, 483, 992, 485, 1395]. Hydroxy [1384, 1162, 1766, 1935]. Hydroxy-oxylipins [1162]. Hydroxyapatite [1483, 1580]. Hydroxybutyrate [1941]. Hydroxychalcones [1518].

Hydroxylase [1288, 1258]. Hydroxynaphthol [1907]. Hydroxyphenylacetic [1101]. Hydroxysteroid [109]. Hymeniacidon [820]. Hymenialdisine [989]. Hyper [945]. Hyper-production [945]. Hyperglycemic [395, 478, 85, 361]. Hyperoxygenated [802, 810]. Hypersalinity [773, 970]. Hypersalinity-Induced [970]. Hyperthermophilic [643, 574]. Hypertrophied [1344]. Hypervariable [1257, 506]. Hypnea [991]. Hypodermal [115]. Hypoosmotic [607]. Hypophthalmichthys [1798, 1556]. Hyporhamphus [236]. Hypothesized [535]. Hypotrophy [1473]. Hypoxia [1288, 938, 1030, 1402, 1393, 1180, 1712, 1847, 1790]. Hypoxia-Inducible [1288, 1030]. Hypoxia-Regulated [1180]. Hypoxia-tolerant [1790]. Hyriopsis [807, 1536, 1502, 1954, 1901, 1867].

I84 [1863]. IAG [1880]. Ianthelline [1215]. ianthina [427]. IB [717]. IbpA [218]. IbpA/B [218]. Ice [485, 747, 943, 943]. ICE-L [943]. Iceland [1803, 1786]. Ichthyophthirius [1870]. Ichthyosporean [672]. Ichthyotoxicity [1317]. Iciligorgia [652]. ICL1 [1094]. ICP11 [1439]. ictaluri [1721, 1471]. Ictalurus [1697, 1775, 286, 40, 190, 1435, 1316, 426, 201, 870, 663, 1747]. ide [1570]. idella [1846, 1911, 1776, 1736]. idellus [1904, 1003]. Identification [338, 443, 807, 1807, 1727, 144, 495, 528, 20, 626, 1031, 1731, 1870, 719, 556, 790, 958, 1449, 85, 1560, 1068, 61, 861, 1793, 866, 1004, 1465, 1753, 1279, 1720, 1754, 1864, 617, 816, 905, 1553, 1619, 27, 1373, 81, 1413, 1646, 1160, 1280, 251, 480, 1725, 321, 1103, 1267, 262, 949, 1121, 1442, 1376, 1590, 1352, 1128, 23, 852, 1043, 1251, 1566, 453, 300, 305, 827, 664, 1347, 1784, 1952, 1726, 558, 997, 1107, 1858, 1180, 102, 1837, 1470, 1058, 619, 748, 260, 622, 1133, 733, 1302, 1597, 1640, 1616, 755, 1309, 1262, 204, 585, 741, 1216, 1319, 608, 1531, 1179]. Identification [1382, 1592, 1033, 334, 1407, 1447, 1244, 887, 1151, 846, 369, 223, 1767, 213, 304, 1234, 672, 374, 380, 409, 1957, 758, 243, 1204, 1318, 1586, 1668, 841, 446, 1062, 1848, 416, 1565, 1377, 1446, 593, 710, 688, 920, 1508, 1086, 1396]. Identified [1168, 256, 1225, 1393, 233, 1458]. Identifies [927, 1335, 1636, 1375, 1427, 1379]. Identify [255, 1826, 1913, 260, 1881, 428]. Identifying [1061, 1418, 1503, 694, 1705]. IEC [947]. IEC-6 [947]. IgE [135]. IGF [274, 327, 145]. IGF-2 [145]. IGF-I [327]. IGFBP3 [387]. IgNAR [1308, 1097, 1257]. II [1016, 5, 352]. IIA [593]. IIB [592]. III [987]. IL-2 [911]. ilisha [1815]. Illuminating [1691]. Imaging [842, 261, 1916]. Imbalance [1078]. Imides [297]. Immature [231, 981]. Immobilization [1751, 1616, 1147]. Immobilized [1210, 1579]. Immune [966, 807, 1807, 861, 1038, 1715, 509, 965, 1380, 1838, 309, 1361, 391, 877, 1012, 1441, 1072, 1185, 1664, 1819, 1617, 1942, 409, 1573, 1947, 1846, 1524, 710, 1946, 961, 304]. Immune-Enriched [1012, 1072]. Immune-Related [1819, 304]. Immune-Relevant [409]. Immunisation [1294]. Immunity [843, 565, 1901, 1941, 1266, 1479, 1917, 1524, 1157, 1086]. Immunized [1832].

Immunoadjuvant [392]. Immunochromatographic [1796]. Immunoenhancing [1723]. Immunogens [29]. Immunolocalization [100, 659]. Immunologic [40]. Immunological [869]. Immunometabolic [1960]. Immunomodulatory [702, 762, 1530, 1171, 1365, 990]. Immunoprecipitation [1135]. Immunopurification [345]. immunostimulating [492]. Immunostimulation [843]. Immunostimulatory [1359, 1899]. Impact [1227, 1331, 1806, 1272]. Impacts [1803, 1786]. Impaired [1789]. Impairment [1822, 1713]. Impairs [1665, 1813]. Impeding [1908]. Implicated [1423]. Implication [832, 881, 1942, 1356]. Implications [316, 276, 1403, 638, 983, 956, 1869, 970, 1638, 623]. Importance [408, 953, 984, 1510]. Important [80, 844, 1863, 1370, 1644, 1356, 1496]. Imported [1272, 516]. Improve [985, 238, 1940, 728]. Improved [1775, 445, 1170, 1941, 403, 114]. Improvement [1481, 429]. Improves [1301, 1875, 1450, 1878, 1747]. Improving [1811, 826, 1202, 1349, 1917]. improvisus [1659]. impudicum [389, 492]. IMTA [1374]. Inactivated [985, 1165]. Inactivation [1176]. Inbred [153]. Inbreeding [1535, 1577]. Including [757]. Inclusion [1465]. Incorporated [466]. Incorporation [90, 1557, 1483]. Increase [1094]. Increased [1287, 550, 283, 1471]. Increases [859, 938, 1381, 1905]. Increasing [1615]. Independent [1089]. Index [1636]. India [1704, 259, 195]. Indian [1082]. Indicate [768, 1421]. Indicated [1510]. Indications [663]. Indicative [1864, 1352]. Indicator [1417]. Indicators [256, 1225, 1543]. indicum [359]. Indigenous [1796, 1664]. Individual [1312, 232]. Indole [410]. Indonesia [533]. Indoor [1095]. Induce [1734, 1478, 1392]. Induced [1227, 1594, 1681, 1330, 958, 1105, 950, 646, 163, 1726, 180, 1453, 185, 1250, 157, 1274, 67, 957, 983, 956, 970, 1581, 314, 597, 1939, 862, 1842, 289, 1003]. Inducer [1380]. Induces [911, 1872, 1909, 1436, 1589, 1894, 42]. Inducible [255, 1288, 878, 1030, 399]. Inducing [1951, 14, 161, 1452, 1749]. Induction [1031, 274, 565, 418, 1834, 239, 43, 1541, 116, 1087, 689, 31, 921, 410, 1065, 581, 482]. Industrial [1150, 1624, 1516]. Industrial-Scale [1624]. Infected [63, 1826, 1832, 134, 1441, 1069, 1655, 1790]. Infection [1873, 1161, 1886, 158, 1922, 1063, 1805, 130, 1417, 1068, 1528, 1413, 1721, 1841, 1012, 1439, 712, 1072, 164, 1758, 1705, 1651, 1471, 380, 580, 1201, 1783, 1921, 1950, 1298, 1429, 1689]. Infections [1153, 753]. Infectious [120, 134, 1796, 927, 189, 115]. infective [1691]. Inference [1111]. Inferred [675, 1594, 1681, 697, 1761]. Infertility [878, 888, 1709]. Infestation [1820]. Inflammation [1515, 1897]. Inflammatory [1415, 1919, 1939, 391]. Influence [537, 227, 1495, 201, 624, 1473, 1573]. Influenced [1188]. Influences [1074, 791, 1218]. Information [1126, 1863]. Ingested [1526]. Ingestion [1882, 1937]. ingrahamii [1150]. Ingredients [1670]. Inheritance [148, 1775, 1305, 357, 137, 1263, 214, 960, 1867, 1733]. Inhibit [1480, 977, 1088, 1572]. Inhibiting [1304, 1506, 262, 1224, 634, 243, 131, 581]. Inhibition [1063, 637, 1763, 778, 695, 297]. Inhibitor [6, 859, 606, 936, 1330, 258, 1475, 461, 1901, 778, 1863]. Inhibitors [642, 408, 1392, 363, 974, 1230]. Inhibitory [489, 770, 590, 1413, 1929, 812, 1671, 1730, 53, 1583, 591, 1778, 804]. Inhibits [1215, 863, 712, 879, 149, 1638, 1217, 1794]. iniae [1792]. Initial [330, 101, 1499, 776, 655]. Initiation [325, 71, 1433, 1400]. Injected [1951]. Injection [1344, 1266, 468, 191]. Injury [950]. Inland [107]. Innate [509, 965, 1380, 1901, 1524, 710, 1157]. Inner [1392]. Ino [570]. Inoculated [1783]. Insert [330, 205]. Insertion [1671, 1752, 1952]. Insertion-specific [1952]. Inserts [1499]. insig1 [1522]. Insight [1227, 1688, 945, 1779, 1533, 1532, 1603, 1755, 1185, 1315, 1846, 1761]. Insights [1904, 1773, 1951, 177, 1348, 1871, 1759, 1744, 1540, 939, 1959, 824, 1709, 1928, 1808, 1719, 1412, 1714, 1117, 1535, 1577]. Insoluble [658, 1061, 1168]. Inspired [1799, 1125]. Instability [166]. Insulin [274, 275, 270, 147]. Insulin-like [274, 275, 270, 147]. Intact [1289, 723]. Intake [1859, 1695]. Integrated [1585, 1374, 928, 1080, 1768, 1548, 448, 1660, 1737, 1611, 1052, 1655, 1027, 1782]. Integrating [1649]. Integration [1410, 254, 1543, 640, 1269, 560, 1706]. Integrin [668]. Intensely [348]. Intensification [65]. intensity [497]. Interact [1572]. Interaction [1338, 16, 109, 1124, 336]. Interactions [1378]. Interference [1851, 1304, 1526, 965, 1057, 1385, 930, 700, 728, 695]. Interferon [255, 1380, 194]. Interferon-Inducible [255]. Interferon-Stimulated [1380]. Intergenic [56, 1952, 442, 500, 701, 1318, 459]. Interleukin [392, 520]. Interleukin-1 [520]. Intermediate [1233]. intermedius [957]. Internal [250, 783, 219, 922, 136, 305]. Internally [93]. International [1007]. Interplay [1960]. Interpolyelectrolyte [1469]. Interrogation [1667]. interruptus [1260, 415]. Intersex [1428]. Interspecific [104]. Intertidal [1615, 853, 970]. Interval [814]. Intestinal [1294, 1941, 1797, 1218, 1217, 1947, 1924]. intestinalis [153]. Intestine [1331, 1528, 1341, 1171, 1378, 1326]. Intra [1234, 1277]. Intra-Colonial [1277]. Intra-Crystalline [1234]. Intracellular [859, 1477, 786, 471]. Intracytoplasmic [468]. Intraperitoneal [1183]. Intraspecific [34, 228, 1315]. Introduction [565, 854, 167, 962]. Introgression [960, 926]. Intron [34, 1467]. Introns [148, 288]. INU1 [1094]. Inulin [1339, 1094, 1268]. Inulin-hydrolyzing [1339]. Invasion [224]. Invasive [224, 1403, 839, 282, 1385]. Inverse [759]. Invertebrate [1135, 400, 547, 363, 1505, 503]. Invertebrate-Type [1683, 1682]. Invertebrates [439, 53, 939, 25]. Investigate [1422]. Investigating [1703, 506]. Investigation [934, 1931, 780, 1801, 1143, 1133, 864]. Investigations [892]. Involved [414, 807, 1536, 790, 436, 1901, 309, 1482, 852, 1566, 1634, 1728, 1058, 1616, 1309, 1679, 1460, 1249, 1522, 1591, 1462, 1932, 1573, 1446, 1362, 710, 943]. Involvement [401, 668, 1433, 957, 1152, 1861]. Iodine [684]. Iodoacetophenones [1056]. Ion [1716]. Ionian [266]. Ionization [85, 367, 1319]. Ions [1016]. iota [1116]. Ircinia [210]. iredalei [321]. IRF9 [1380]. IRF9-Stat2 [1380]. Iridovirus [1413, 1399]. iris [543]. Irish [809, 528, 884, 793]. Iron [1554]. Irradiance [1297, 544]. irradians [608, 548, 1020, 608, 548]. Irradiated [1187, 1003]. Irradiation [1576, 418]. irritans [1873, 1826, 1504, 1758, 1921, 1648]. ISA [927]. isenbeckii [675]. Island [1465]. Islands [1937]. Isochrysis [1768, 1839, 355]. Isoform [1402, 622, 1953, 1758]. Isoforms [156, 395, 98, 1423, 1013, 307, 225, 920]. Isogenic [860]. Isolate [549, 1206, 1395]. Isolated [809, 1193, 1843, 1491, 126, 1215, 1475, 370, 457, 1339, 1827, 913, 1631, 1138, 1268, 319, 1480, 792, 1037, 1495, 760, 1505, 756, 1326, 729, 1517, 1656, 811, 1607, 312, 1578, 1924, 679, 1350]. Isolates [1142, 367]. Isolation [443, 578, 48, 543, 626, 1167, 534, 138, 433, 258, 35, 36, 124, 1143, 427, 1200, 476, 599, 793, 954, 1308, 949, 187, 435, 554, 1590, 566, 630, 657, 1171, 186, 28, 295, 872, 172, 865, 1395, 627, 875, 333, 795, 365, 423, 1650, 331, 798, 298, 350, 703, 86, 39, 1112, 685, 851, 553, 154, 200]. Isomerase [1021]. Isothermal [1935, 1907]. Isothiocyanate [523, 1748, 257]. Isothiocyanate-Labeled [523]. Isotopes [1225, 1828]. Issue [962, 1853]. ITS-1 [219]. ITS-2 [460]. ITS1 [783, 381, 623]. ITS2 [381]. IV [109]. Ivory [1445]. Iwagaki [1883].

Jack [1278]. jadensis [383]. JAMB [945, 1490]. JAMB-A94 [945]. jamc [1490]. Japan [675, 1111, 52, 1622, 672, 1937]. Japanese [63, 1063, 176, 1282, 103, 1661, 1703, 1010, 909, 476, 599, 142, 411, 1163, 1456, 1862, 202, 455, 1665, 157, 741, 221, 530, 244, 214, 577, 676, 1602, 1937, 1201, 446, 1950, 1857, 522, 482, 673, 302, 592]. japonica [1167, 829, 1661, 1703, 929, 871, 45, 202, 209, 446]. japonicus [897, 610, 750, 1855, 1871, 476, 599, 424, 664, 294, 1171, 1266, 1278, 1338, 301, 636, 1244, 76, 676, 1719, 431, 625, 302]. Javanese [597]. javanicus [597]. jelly [505]. Jellyfish [265]. JJY [1006]. JJY-40 [1006]. JNK [1872, 1152]. Jones [131, 710]. JPCC [1096, 1824]. JSMB [1853]. JT01 [1396]. Juice [1312, 55]. juncea [458]. Junceella [458]. Juvenile [532, 1904, 1212, 1506, 1163, 802, 810, 1447, 225, 1426, 1559]. Juveniles [1806, 1889, 1662, 820, 654].

K-ATPase [307]. k-Carrageenan [313]. kaikoi [1047]. Kallman [160]. Kanagawa [1111, 1622]. Kappaphycus [1333, 1307, 1313]. Karenia [684, 508, 837]. Kaup [745, 1222, 1218]. Kazal [606, 778, 461]. Kazal-Type [606, 778, 461]. kDa [280, 235, 590, 1799]. Kelp [1764, 1300, 1627, 1626]. kenojei [1869]. Ketolases [511]. Ketones [1384, 1285]. Ketoreductases [1233]. Key [1812, 1335, 774, 1015, 1428, 1462, 1601]. KG03 [389, 492]. Kidney [514, 134, 1012, 1072, 1024, 1272, 1298, 1790]. Kidney-Specific [1272]. Killer [879, 1044]. Kinase [19, 36, 765, 1047, 1124, 38]. Kinetic [1384, 1233, 535]. Kinetics [1475]. kingsejongensis [1345]. kisutch [1637, 1283]. KL1 [1520]. KL2 [1520]. Klaus [160]. Klotho [1520]. KMM [845]. Knifejaw [1590, 1784]. Knock [278, 1878]. Knock-down [278, 1878].

Knockdown [1301, 1435, 1194, 1943, 1942]. Knockin [560]. Knockout [1740, 1671, 1645, 1603, 1214, 560]. Knockout/Knockin [560]. Knoevenagel [1576]. Known [159, 1458]. Kodamaea [692]. Koi [1677]. Kong [1852, 381]. König [760]. Korea [621]. Korean [19]. koreni [86]. Krill [413, 748, 600]. Kumamoto [1805]. Kunitz [1901]. Kunitz-Type [1901]. kurome [1133]. Kuroshio [976]. Kuruma [750, 1871, 664, 1171, 1266, 301, 636, 431, 1103].

L [1620, 234, 1442, 1755, 487, 943]. L-Asparaginase [1755]. L-Lactate [1442]. L. [520, 539, 37, 1192, 393, 903, 1271, 433, 111, 1118, 1528, 1294, 224, 1211, 1149, 972, 1019, 30, 208, 745, 854, 1302, 853, 199, 67, 906, 345, 847, 830, 1917, 942, 1117]. L1 [770, 812, 863]. L10a [855]. L37A [25]. Lab [1624]. Lab-Scale [1624]. Labeled [366, 523, 257]. Labeling [60, 1500, 498]. Labelled [1748]. Labeo [1929, 1885, 1531, 1076, 1951]. Laboratory [769, 1463]. labrax [197, 392, 520, 539, 328, 716, 1931, 111, 1830, 1864, 253, 1512, 906, 345, 1388, 1362]. Labyrinthulid [549]. Labyrinthulids [154]. Laccase [1487, 828]. Lacepède [1873]. Lachesis [1350]. Lack [1665, 266]. Lacking [1198]. Lactate [1442, 918]. Lactic [1171]. lactis [1292, 1437]. Lactococcus [1292, 1437]. Lactonase [1217]. Lactone [1354]. Lactone-Dependent [1354]. lactuca [1063]. laevigata [1236, 1223, 1370]. lagopoides [1371]. Lake [352, 9, 774, 352, 1664]. Lakes [969]. Lamarck [608]. lamB [217]. Lamellibrachia [385]. Lamina [1133]. Laminarales [871]. Laminaria [454, 1530, 829, 929, 871, 782]. Laminariaceae [829]. Laminariales [1167, 1133]. Lancelet [1520]. lanceolatus [1542]. Land [542]. Landlocked [1622]. Landscape [1687, 1769]. Large [1807, 1192, 196, 141, 330, 1346, 1111, 1646, 1955, 1504, 1605, 680, 1956, 1636, 1119, 1195, 356, 1338, 1758, 334, 847, 1781, 1803, 1786, 1110, 1318, 1850, 1762, 1219, 1847, 1927, 1790, 1648, 1538]. Large-Insert [330]. Large-Scale [1192, 1111, 1119, 1195, 1338, 847, 1803, 1786]. Largemouth [1754]. Largest [1542]. Larimichthys [1346, 1955, 1504, 1605, 1956, 1636, 1758, 1769, 1762, 1219, 1847, 1549, 1648, 1761, 1538]. Larvae [1018, 556, 1805, 1445, 547, 1883, 1380, 1070, 23, 230, 1453, 236, 95, 456, 1153, 8, 665, 232, 676, 1586, 1668, 38, 1388, 654]. Larval [1731, 934, 716, 327, 637, 1688, 489, 1026, 486, 1876, 854, 1309, 1611, 1172, 1202, 1087, 1716, 804, 1813, 1059, 499]. Larviculture [1437, 671]. Laser [400, 261, 367, 1319, 864]. Late [1164]. Lateolabrax [610, 1595, 1769, 625, 1684, 1414]. Lateral [1796, 1485, 1795]. Lates [922, 538, 230, 669, 912, 145, 306]. latipes [123, 1250, 267, 157, 741, 1950]. latissima [1764]. Latrunculiidae [100]. Latrunculin [100]. latus [1759, 1903]. Laurencia [1350]. Layer [1954, 322, 125, 1423]. Layers [1029]. LC [1445, 1482, 875, 1522, 1591]. LC3 [730]. ldhA [315]. Leach [524, 587]. leachii [299]. Leap [1691]. Learn [1575]. Learning [563]. Leaves [903]. Lectin [1099, 1358, 1333, 1307, 1313, 257, 937]. Lectins [523]. Left [1443]. lemaneiformis [1793, 1810, 1197]. Length [709, 196, 783, 1525, 1552, 871, 608, 252, 334, 1467, 912, 1712]. lenok [1797, 1939, 1797, 1939]. lens [1562]. lenticularis [725]. Lentivirus [1481]. Lentivirus-Mediated [1481]. Lepeophtheirus [751]. Leptogorgia [652]. Less [1740]. Lesser [159]. Lessons [1804]. Lethal [1403]. Leuciscus [1570]. leucospilota [1889]. Leukemia [41]. Leukocytes [63, 1909]. Level [1422, 764, 1759]. Levels [859, 737, 550, 1221, 774, 146, 1479, 201, 1327, 663, 1110, 1573, 673]. LexA [1454]. LF458 [1230]. LFD [1795]. lgi [1932]. lgi-miR-2d [1932]. LH [1477]. Libraries [255, 583, 330, 858, 1875, 704, 249, 887, 428, 710]. Library [902, 1299, 142, 1308, 972, 1019, 508, 1499, 1547, 1639, 1097, 267, 865, 746, 663, 776, 655, 848, 724]. Lice [1820, 1034]. licheniformis [590, 1217]. Lichens [762]. LICP [1234]. Life [797, 985, 88, 1764, 87, 1840, 1803, 1786]. Life-History [1840]. Lifestyle [1188]. Lifetime [1500]. Ligament [1234]. Light [1243, 1925, 737, 1167, 806, 803, 1738, 1645, 1120, 743, 940, 1629, 257, 1362, 497]. Light-Dependency [803]. Light-Regulated [1629]. Light/Dark [1738]. Lightened [240]. Like [611, 1374, 805, 1010, 1954, 1103, 622, 687, 1306, 9, 636, 1062, 1792, 274, 275, 470, 270, 41, 147, 126, 1868, 268, 189, 1058, 819, 368, 669, 121, 416, 1690]. lima [1935]. limacinum [562, 1425]. Limb [1914, 1818]. Limitation [737, 1445]. Limited [1726]. lincRNAs [1318]. Line [596, 539, 610, 1334, 41, 1170, 669, 1669, 1244, 860, 1944, 635, 498]. Lineages [1493, 1492, 1578, 1716]. Lines [866, 1613, 1657, 319, 627, 1436, 926, 1912]. Lingulodinium [525]. Link [1889, 683]. Linkage [1010, 1300, 353, 1504, 390, 1020, 825, 59, 570, 1041, 1595, 182, 128, 1397, 577, 1563, 422, 1377, 942, 1117]. Linked [1913, 433, 1279, 834, 1267, 959, 407, 608, 1586, 1668, 1389, 703]. Linking [639, 696, 1772]. Linnaeus [1113, 581]. Linoleic [1911]. Linolenic [436]. Lip [1452]. Lipase [932, 1147, 1110, 738]. Lipases [464]. Lipid [1227, 1831, 1372, 1095, 1174, 1297, 1198, 1476, 1195, 1180, 562, 1439, 1205, 791, 1090, 1110, 1813, 1911, 1900]. Lipid-Body [562]. Lipid-Metabolizing [1227]. Lipidic [839]. Lipids [1489, 995, 70]. Lipogenic [1911]. lipolytica [1295, 818, 1094, 761, 924]. Lipophrys [104]. Lipopolysaccharide [1075, 482]. Lipoprotein [133]. Liposome [1734, 425]. Liposome-Encapsulated [1734]. Lipovitellin [948]. Lipoxygenase [936]. Lipoxygenases [1744]. Lipped [1146]. Liquid [889, 44]. Liquid-Saturated [44]. Listeria [590]. Lithistid [1089]. Litopenaeus [356, 908, 280, 404, 1304, 1553, 1796, 742, 1479, 515, 1087, 1906, 634, 437, 1598, 1586, 1668, 1524, 848, 1849]. Litopeneaus [1377, 1513]. Live [75, 1070, 162, 231]. Live-Bearing [162]. Liver [1594, 1681, 1822, 1085, 1658, 1174, 950, 1482, 1814, 1600, 1283, 1180, 1012, 157, 1072, 1531, 860, 1769, 1102, 380, 1960, 1939, 1573, 1235, 1515, 1897, 1847, 895, 1612, 1776, 1523]. Liver-Specific [1531]. lividus [1328, 824, 1327]. Living [1131, 1281]. lncRNA [1338, 1378]. lncRNA-Coding [1378]. lncRNA-microRNA-Gene [1338]. lncRNAs [1338, 1783, 1807, 1852, 1338].

Loach [1341, 349]. Loaded [1899]. Loading [1617]. lobata [1558]. Lobophytum [1723]. Lobster [1260, 1312, 393, 415, 556, 676]. Lobsters [372, 900, 1100, 506]. Local [1622, 1903, 775]. Localization [444, 351, 1103, 837, 1837, 1916, 1178, 989, 849, 1949]. Localized [714]. Location [800]. Loci [419, 34, 1555, 434, 1164, 735, 124, 1553, 476, 1300, 1636, 458, 137, 129, 186, 1302, 1397, 1144, 333, 1179, 1407, 214, 1466, 209, 1201, 1375, 39, 1155, 1377]. Locus [330, 1279, 834, 1398, 1837, 352]. Lohman [320]. Loligo [137, 54]. Long [1190, 979, 551, 1760, 1560, 1832, 1503, 1266, 1338, 1123, 1767, 1652, 1783, 1412]. Long-Chain [1190]. Long-Tailed [551]. Long-Term [979]. Long-Time [1123]. longirostris [1529]. Look [797]. Loop [1935, 1907, 1257]. Loop-Mediated [1935, 1907]. Loss [1789]. Louse [1851, 1918, 1290, 1685, 751]. Low [605, 1680, 284, 1478, 1451, 1129, 471, 1663, 1276, 1808, 1387, 1310]. Low-Salt [1129]. Low-Toxicity [1276]. Lox [640]. loxP [1136]. LPS [1338]. LQ48 [840]. LR [1865]. LSK [1302]. Lubomirskia [979]. Luffariella [678]. Lugol [684]. lunatic [566]. lunatus [1330]. Lung [1436]. lusoria [993]. Luteinizing [245, 1316]. luteoculcitella [973]. Lutjanidae [661]. Lutjanus [406, 212, 661, 1909]. LuxI [1354]. LuxR [1354]. LvMMD2 [1598]. Lyase [929, 1265, 579, 754, 1289, 1246, 1936, 818, 1122, 1037, 326, 362, 83, 84, 883]. Lyases [1416]. Lymphocytes [911, 1365, 345]. Lymphoid [424, 1669]. Lysine [1904, 1352, 1078]. Lysine-Rich [1352]. Lysogens [239]. Lysozyme [103, 1097]. Lytic [801].

m [992, 723]. m-Passable [992]. M318 [1607]. Machinery [745]. Mackerel [1855, 1278, 1882, 516]. Macroalga [1195, 970]. Macroalgae [75, 1162, 1432]. Macroalgal [1406]. Macroarray [651]. Macrobiofouling [1215]. Macrobrachium [606, 395, 1773, 1806, 1838, 1170, 778, 1499, 493, 1589]. macrocephalus [1357, 1634]. macrodactyla [265]. Macrolactams [1491]. Macroorganisms [638]. Macrophages [762, 1950]. Macroscopic [1623]. Macruronus [551]. maculatus [161, 1595, 1769, 1684, 1414]. Made [1920]. magellanicus [551]. Magnetic [30, 113]. Magnetic-Capture [113]. magnifica [100]. magur [1959]. Main [737]. Major [135, 1418, 909, 717, 953, 964, 1858, 1071, 748, 128, 168, 536, 592, 593, 443, 1407, 1447, 1859]. Malabar [1953]. malabaricus [1953]. malaccana [1204]. Malacostraca [413]. Malaria [575]. Malaysia [1514, 206]. MALDI [842, 1121, 730, 1133]. MALDI-TOF [1133]. MALDI-TOF/TOF [1133]. Male [1697, 606, 878, 1242, 1364, 778, 1590, 850, 1119, 1784, 1862, 407, 1262, 1880, 1677, 1521, 1112, 1642, 1508, 1274, 1493, 1492, 1632]. Male-Linked [407]. Male-Specific [1590, 1784, 1112]. Males [1508]. maltophilia [1505]. Mammalian [388]. Man [855]. Management [316]. Mandarin [920]. Manganese [1029]. Mangrove [1241, 138, 1722, 1339, 1845]. Mangrove-Associated [1241]. mangrovi [1363]. Manila [1750, 997, 1706, 1932]. Manipulation [908, 843, 1119, 1735]. manjavacas [1725, 930]. Mannan [1099, 836]. Mannan-Binding [1099]. Mannan-Specific [836]. Mannitol [234, 501]. Mannitol-l-Phosphatase [234]. Mannogalactofucans [1478]. Mannose [1333, 1307, 1313]. Manoalide [678]. Mantis [748, 521]. Mantle [1536, 1158, 1156, 720, 951, 444, 1146, 1376, 1000, 1540, 122, 1236, 1108, 1151, 1714, 1582, 1113]. Mantles [1443]. Manufactured [1823]. Manzamine [1191]. Map [1555, 1504, 834, 825, 570, 1041, 1595, 1222, 182, 1397, 1052, 1563, 1544, 1027, 1117]. MAPK [947]. Mapping [578, 330, 433, 1555, 1039, 1553, 1010, 1504, 1211, 814, 959, 680, 1514, 825, 1411, 59, 871, 1303, 1398, 1595, 182, 165, 608, 1447, 1563, 1466, 1079, 912, 1399, 1537, 422, 1155, 1219, 688, 942, 1556]. Maps [1300, 353, 390, 1511]. Marbled [1565]. Marcos [178]. margaritifera [658, 1448]. Mariculture [1140, 615]. Mariculture-Raised [615]. Maricultured [1018]. marina [1232, 704, 362]. marincola [1839]. Marine [15, 1135, 1746, 1585, 602, 1487, 1368, 1243, 1419, 1317, 845, 1028, 1098, 1210, 1018, 1175, 259, 859, 1340, 1576, 649, 80, 1468, 1444, 46, 359, 1080, 1060, 856, 797, 1691, 400, 1330, 1321, 1384, 511, 783, 934, 790, 58, 958, 1449, 1625, 583, 828, 341, 722, 210, 1233, 598, 534, 545, 567, 973, 354, 1788, 1320, 32, 1811, 1372, 769, 1463, 452, 119, 34, 1486, 842, 1247, 547, 1232, 720, 876, 1835, 894, 211, 767, 1215, 1475, 394, 1045, 568, 682, 510, 617, 905, 383, 27, 711, 1145, 1373, 408, 217, 1209, 953, 1200, 1743, 1265, 337, 683]. Marine [1654, 1221, 261, 579, 793, 754, 1725, 439, 97, 218, 1738, 1281, 1336, 1173, 1162, 346, 552, 1206, 595, 1700, 575, 279, 1134, 692, 1138, 1482, 1630, 1128, 386, 1224, 986, 1043, 1094, 840, 453, 827, 49, 632, 1268, 1771, 1195, 935, 1240, 65, 628, 10, 319, 1122, 391, 239, 1529, 629, 1670, 1125, 1547, 1639, 1096, 917, 1856, 1930, 915, 706, 761, 485, 713, 755, 1150, 1050, 1896, 872, 368, 792, 66, 865, 367, 1438, 1231, 1483, 1580, 1365, 879, 924, 363, 475, 449, 1046, 1258, 1755, 638, 740, 760, 318, 1505, 1369, 3, 503, 939, 795, 886, 1056, 1053, 1285]. Marine [1140, 1382, 293, 365, 326, 423, 1920, 1436, 298, 1501, 1259, 379, 660, 473, 677, 25, 192, 989, 1401, 1823, 984, 344, 650, 1208, 1256, 1006, 1296, 1614, 974, 1656, 811, 1464, 1699, 576, 978, 410, 1261, 1177, 1021, 42, 314, 417, 787, 933, 38, 26, 44, 1230, 1765, 1044, 820, 685, 82, 1853, 289, 553, 389, 1488, 691, 775, 800, 781, 635, 1518, 1101, 1169, 1609, 1]. Marine-Based [1018]. Marine-Derived [1487, 1340, 1576, 1080, 1330, 828, 595, 1134, 1094, 1240, 1438, 879, 1053, 1285, 1044, 1488, 1518, 1609]. Marine-Isolated [1268]. marinum [1142]. Marker [928, 219, 88, 1720, 1590, 825, 1784, 33, 1177, 38, 506]. Markers [1181, 533, 807, 37, 393, 384, 626, 1031, 1238, 1913, 1282, 335, 433, 1279, 104, 609, 1619, 81, 717, 1300, 353, 916, 132, 96, 193, 375, 834, 411, 959, 357, 1020, 59, 871, 1251, 1739, 1093, 664, 1952, 87, 1222, 1262, 515, 608, 249, 530, 244, 347, 577, 149, 450, 1079, 648, 1204, 1511, 1710, 422, 1112, 447, 1027, 518, 504]. Market [54]. marmoratus [93]. Marsh [1054]. Marsupenaeus [1244, 897, 750, 664, 1171, 1266, 636, 431]. Marteilia [73]. martensii [1160, 1108, 1324, 1572, 931, 1836, 1582, 1086]. Masculinization [1732, 1581]. Masker [846]. masou [1622, 1747]. masoucida [1715]. Mass [1805, 85, 1788, 842, 733, 367, 1371, 1319, 1916, 864, 355]. Master [1308, 1176]. Masu [611, 1747]. MAT3885 [1265]. Material [1772, 683, 1438]. Materials [1173]. Maternal [111]. Matrices [1940, 573]. Matrix [642, 1061, 1448, 1444, 1168, 720, 118, 909, 1643, 1502, 1954, 444, 1352, 1000, 1017, 997, 125, 367, 122, 220, 591, 536, 589, 594, 1749, 1774]. Matrix-Assisted [367]. Mats [198, 691]. Matter [639]. Maturation [1697, 1804, 1734, 1164, 735, 1364, 1800, 1610, 1527, 921, 1509, 681, 1567]. Mature [981]. maxima [1146, 925, 1153, 1452, 150]. Maximum [901]. maximus [528, 1444, 384, 1068, 1528, 81, 877, 1012, 1072, 1179, 1079, 1917, 593, 1539, 1400]. Mayamaea [1824]. MB1 [545]. MBC2019 [1614]. MBR [1548]. mc4r [1775]. mccartyi [1351]. MCP [826]. MCP-01 [826]. Meal [1331, 1957]. Measured [257]. Measurement [420, 70]. Mechanically [1772]. Mechanism [1292, 1807, 945, 276, 730, 778, 1953, 1371, 1679, 1428, 1663, 1921, 1895]. Mechanisms [965, 1821, 1533, 1532, 1540, 391, 1178, 1527, 1717, 1791, 1139]. Medaka [123, 1220, 240, 248, 1250, 168, 47, 267, 741, 1214, 283, 166, 597, 1950, 1166, 1497, 157, 170]. Media [1129, 553]. Media-Supplemented [553]. Mediate [700]. Mediated [1481, 340, 941, 1301, 1519, 1934, 1645, 1194, 640, 352, 616, 110, 324, 1747, 1935, 1825, 1539, 1907, 888, 1589]. Mediates [1683, 1682, 1939, 1139]. Mediators [638]. Mediterranean [1415, 1408, 489, 988, 108, 1394, 1455]. Medium [1372, 1748, 427, 607]. Medway [1054]. Meeting [112, 1007]. Megalobrama [1274, 1493, 1492, 1567]. Meiogynogenetic [111]. Meiosis [1546]. Meiosis-Related [1546]. Melanin [766, 1934, 1926, 240, 1844, 1932]. Melanin-Concentrating [240]. Melanocortin [1775]. Melanocortin-4 [1775]. Melanogenesis [1770]. melanogenum [1487, 1268]. Melanogrammus [596, 295]. Melanoma [1568, 173, 1912, 273]. Melanomas [171]. Melatonin [1947]. melinii [1028]. Member [74]. Membrane [1832, 332, 1871, 1277, 899, 723]. Memory [962]. Mendelian [148, 129]. Menippidae [493]. menstrualis [1322]. Mercenaria [1809, 1729, 817, 581]. Mercinaria [11]. Mercury [341, 722, 752]. Meretrix [993, 1026, 1093, 1928, 1254, 1782]. merguiensis [657, 1674, 855]. Merluccius [419]. Meroterpenes [1631]. Mesocentrotus [1710]. Mesocosm [1548]. Mesocosms [1585]. Mesoderm [1244]. Mesohaline [1584]. Message [272]. Messenger [147, 190, 201]. Meta [1192]. Meta-analysis [1192]. Metabolic [328, 1192, 1477, 1496, 1649]. Metabolically [445]. Metabolism [1831, 1925, 1301, 1174, 1887, 501, 1838, 1191, 5, 1914, 1369, 1679, 1813, 1573, 1900]. Metabolite [1801, 618, 803, 1336, 1371, 904]. Metabolites [1248, 51, 882, 822, 489, 639, 1134, 1138, 30, 71, 670, 1395, 839, 1774]. Metabolizing [1227]. Metabolome [1445, 1889]. Metabolomic [1719, 1523]. Metabolomics [1788, 1476, 1441, 1567, 1649, 1865]. Metagenome [784, 872]. Metagenomic [902, 766, 1353, 1547, 1639, 992].

Metagenomics [1963]. Metal [828, 1054, 597]. Metal-Contaminated [1054]. Metal-Induced [597]. Metal-Tolerant [828]. Metalloproteinase [591]. Metalloproteinase-9 [591]. Metallothionein [483, 98, 1592, 597, 494]. Metals [722, 1250]. Metamorphosis [101, 1398, 1309, 1611]. Metamorphosis-Related [1398]. Metapenaeus [681]. Metastatic [1568]. Methane [440]. Methanogens [70]. Methanotrophic [582]. Methanotrophs [440]. Methicillin [590]. Methicillin-Resistant [590]. Methionyl [574]. Method [1135, 903, 1334, 944, 1793, 1519, 1654, 238, 448, 480, 21, 1000, 491, 1123, 564, 577, 369, 758, 946, 446, 864, 1765, 1935, 154, 1551, 516]. Methodologies [1711]. Methodology [709, 799, 251, 1173, 892, 890]. Methods [1704, 1089, 285, 1856, 1930, 1474, 1226]. Methoprene [343]. Methoxyphenyl [1053]. Methyl [903, 1579]. Methylation [1083, 1305, 1381, 1498, 1814, 1273, 1890, 1959, 1154, 1921, 1861, 1847, 1582]. Methylcholanthrene [43]. Methylcrotonyl [443]. Methylglutaryl [1766]. Methylmercury [1139]. Methylome [1780]. Methylomes [1709]. Methylphenylacetonitriles [1169]. Methyltransferase [1330]. Methyltransferases [1597, 1640]. methylurea [163]. Metolachlor [616]. Metolachlor-Mediated [616]. Metschnikowia [1808]. Mexico [537, 1635, 406, 212, 437]. Meyerozyma [1501]. Mg [838, 957]. mh0186 [791]. MHC [352, 168]. MHCII [350]. Mice [808, 1882]. Micro [1215]. Micro- [1215]. microadriaticum [557]. Microalga [366, 207, 1372, 617, 1200, 1654, 254, 795, 389, 492]. Microalgae [1562, 1032, 815, 646, 1457, 755, 1319, 875, 1382, 1916, 213, 1516, 851, 901, 497]. Microalgal [1009, 1365, 616, 1917, 336]. Microarray [966, 975, 1192, 647, 716, 1068, 1035, 866, 1038, 653, 508, 1163, 1266, 432, 619, 854, 172, 1072, 847, 1324, 580, 877, 1012]. Microarray-Based [1068, 866]. Microarrays [694, 380]. Microbacterium [403]. Microbes [904]. Microbial [1098, 46, 1680, 296, 985, 568, 408, 1641, 1743, 1431, 52, 198, 691]. Microbiology [1007]. Microbiome [1902, 1963, 1736]. Microbiota [1291, 1892, 1421, 1902, 1353, 1608, 1797, 1620, 1505, 1218, 1947, 1905]. Microbulbifer [1692, 945, 1363, 1936]. Microcapsules [1653]. Microcystin [1865]. Microcystis [1454]. Microencapsulated [854]. Microencapsulation [1294]. Microflora [385]. Microfossils [1029]. microfouling [839]. Microinjection [1519]. microlepis [20]. Microorganisms [1746, 1384, 1701, 1233, 711, 1131, 1281, 1043, 628, 992, 1501, 192]. Microparticles [1388]. Microplastics [1865]. Microplate [363]. Micropropagation [946]. Micropterus [1754]. microRNA [1338, 1886, 1918, 1671, 1685, 874, 1880, 1472, 1252, 1531, 1509, 1655, 1237]. MicroRNAs [1536, 1864, 1160, 1221, 1675, 1522]. Microsatellite [533, 807, 1727, 144, 393, 419, 384, 551, 335, 434, 212, 124, 609, 224, 242, 228, 81, 476, 599, 411, 357, 526, 680, 1020, 458, 1041, 137, 186, 1222, 249, 333, 530, 151, 244, 214, 347, 577, 232, 117, 450, 209, 437, 348, 422, 39, 302, 447, 1027, 518].

Microsatellites [37, 316, 404, 104, 684, 540, 356, 295, 54, 564, 426, 86, 145, 306]. Microscale [1317]. Microscopic [763]. Microscopy [400, 261]. Microspatial [261]. Microsynteny [972, 1019]. Microtiter [354]. Microtubule [561]. Microwave [1576]. micrura [1840]. Mictyris [332]. midae [555, 555, 1014]. Midkine [1756]. Midkine/Pleiotrophin [1756]. Migratory [1815]. mikimotoi [195, 368]. Mild [1061]. Milkfish [627, 252]. Milking [835]. Millennium [503]. Milli [253]. Milli-calpain [253]. Mimic [729]. Mimicking [669]. Mineral [1466]. Mineralization [1940, 1433]. Mineralogenic [594]. Mini [1835]. Mini-review [1835]. minimata [652]. minimum [801, 1066]. Mining [1098, 1722, 515, 426, 1607, 1036]. minor [1884]. minutum [195]. MIPS [1197]. miR [1932, 1770, 1522, 1591, 1651, 1298, 1836, 1601]. miR-10a-3p [1836]. miR-122 [1651]. miR-15 [1591]. miR-15/16 [1591]. miR-221 [1298]. miR-222 [1298]. miR-263 [1601]. miR-33 [1522]. MiR-382 [1770]. miR-9 [1601]. mirabilis [944]. MiRNA [1528, 1706, 1939]. MiRNA-mRNA [1706]. MiRNAs [1604, 1706, 1768, 1731, 1961, 1852]. MiRNAs-Modulation [1604]. miRNome [1768]. miRomics [1689]. Mirror [1458]. Miscellaneous [391]. Misgurnus [1341, 349]. MITF [1934, 1932]. Mitigating [1548]. Mitochondria [1425]. Mitochondrial [675, 56, 463, 569, 91, 421, 768, 1271, 406, 1111, 661, 292, 1525, 1552, 550, 451, 372, 462, 17, 226, 538, 1347, 294, 413, 633, 208, 524, 587, 442, 500, 529, 493, 57, 432, 619, 260, 455, 521, 252, 1024, 221, 4, 108, 266, 374, 817, 960, 459, 431, 1034, 751, 496, 506]. Mitogenomes [1315]. Mitogenomic [731, 1570]. Mitogynogenetic [1031]. Mitophagy [1867]. Mitosis [581]. Mitten [785, 1914, 1249, 1818]. mizolepis [349]. Mizuhopecten [530]. MJ [1289, 1037]. MJ-3 [1289, 1037]. moara [1627, 1626]. Model [1585, 1317, 1291, 563, 1280, 1194, 1670, 1120, 1125, 1052, 1207, 1580, 1359, 122, 449, 1920, 166, 796, 1388, 1515, 1897, 1574, 1672]. Modeling [1692, 987, 986]. Models [272, 177, 159, 181, 535, 167, 1850, 179]. modes [497]. modestus [1868]. Modification [1268, 1439, 429, 1666]. Modified [1287, 888, 1359, 199, 1316, 860, 289]. Modulate [1378]. Modulated [1728]. Modulates [270, 1069]. Modulation [1820, 941, 1174, 1145, 1322, 896, 1819, 1277, 1717, 1791, 1604]. Module [836]. Mohler [320]. Moina [1840]. Molasses [1817]. Molecular [156, 1181, 1742, 1704, 1831, 987, 463, 845, 569, 606, 626, 470, 556, 219, 1913, 1760, 1491, 139, 222, 1032, 440, 582, 963, 829, 364, 547, 1720, 1883, 1868, 436, 103, 107, 40, 242, 387, 1852, 462, 480, 351, 439, 955, 193, 375, 321, 490, 604, 1478, 1590, 143, 23, 519, 1398, 644, 1784, 1744, 1540, 558, 1015, 236, 301, 865, 1953, 78, 449, 1046, 1258, 1159, 573, 771, 194, 1382, 1592, 910, 665, 1916, 221, 1679, 311, 505, 1809, 307, 369, 223, 1781, 676, 1106, 467, 672, 1040, 4, 1443, 26, 1833, 1544, 625, 673, 388, 593, 738, 1066]. molecule [1787]. Molecules [1545, 1559, 961]. molitrix [1556]. Mollusc [1061, 1002]. Mollusca [1749, 1323, 569, 444, 98, 137, 1113]. Molluscan [622, 122, 220]. Molluscs [589]. Mollusk [993, 135, 1543, 1000, 122, 1694, 1356]. Mollusks [454, 1106]. Molt [262, 89]. Molt-Inhibiting [262]. Molting [1925]. Monitoring [1866, 568]. Monocells [75]. Monoclonal [513, 300]. Monocrotophos [1890]. monocytogenes [590]. monodon [1161, 470, 85, 461, 17, 226, 262, 1364, 62, 1674, 1669, 698, 117, 304, 89, 361, 921, 243]. Monogalactosyldiacylglycerol [1143]. Monomethylmonothioarsonic [1858]. Monooxygenase [440]. monophosphate [1015]. Monopterus [1732]. Montastraea [139, 13]. Months [1502]. moon [505, 633]. Moraxella [545]. Moray [169]. mordax [896, 724]. morhua [975, 928, 335, 1035, 1070, 1149]. Moritella [1396]. Morone [124, 263, 18, 1041, 771]. Morphogenetic [1694]. Morpholino [278]. Morphologic [16]. Morphological [1282, 381, 1622, 892, 981, 736]. Morphologically [204]. Morphology [1941, 1884, 1828]. Mortality [1805, 866, 1038, 1370, 1928, 1217]. moseri [78]. mossambicus [274, 291]. Most [1698, 1033]. Motif [805]. Motifs [22]. Motile [1510]. Motility [765]. Mouse [22, 205, 1359, 779]. Mouth [1813]. mpeg1.1 [1950]. mpeg1.1-Transgenic [1950]. MrIAG [1589]. mRNA [667, 774, 1611, 116, 1076, 1880, 1706, 1655]. mRNA-microRNA [1655]. mRNA-seq [1076]. mRNAs [1731, 1852, 1783]. MRPINK [778]. MRSA [757]. MS [1445, 730]. MS-Based [1445]. MS/MS [1445]. MS/MS-Based [1445]. MSAP [1083]. MSP [220]. MSP-1 [220]. MSTN [1830, 1945, 1878, 1762]. MSX [130]. MT [483]. MT-10 [483]. mtDNA [315, 701, 960]. Mucosal [843, 1736]. Mucous [1689]. Mucus [830]. Mud [96, 1608, 349, 1509, 1412, 1601]. muelleri [355]. Mugilidae [260]. Mullet [434, 208, 260]. Mullus [434, 208]. Multi [1374, 757, 1896, 1928]. Multi-omics [1928]. Multi-Pulse [1896]. Multi-resistant [757]. Multi-trophic [1374]. multifiliis [1870]. Multifunctional [414]. Multigene [1121, 1022]. Multimodular [1686]. Multiomics [1960]. Multiple [393, 41, 1271, 34, 29, 1613, 1363, 1936, 1319, 771, 896, 191, 1226, 1387]. Multiple-Drug [41]. Multiplex [528, 1594, 1681, 335, 204, 758]. Multipoint [353]. multistriata [1263]. multitentaculata [1800, 1949]. Multivalent [998]. Multivariate [296]. MuRF [1023]. Muricids [1022]. Murine [702]. muscarum [210]. Muscle [1804, 197, 1424, 1904, 1192, 514, 1855, 517, 1075, 1945, 1885, 253, 105, 1485, 1600, 1958, 1490, 76, 1023, 310, 1878, 1473, 1559, 1604, 1789, 546, 942, 920, 1612, 1237, 1895, 1078]. Muscle-Regulatory [310, 546]. Muscle-Related [920]. Muscle-Specific [1023, 310]. Muscles [662, 1695, 723]. Muscular [191]. Mussel [807, 1016, 567, 582, 1168, 1408, 224, 954, 881, 114, 660, 689, 988, 672, 499]. Mussels [118, 483, 1441]. MUT [1545]. muta [1350]. Mutagenesis [1227, 987, 340, 631, 1289, 166, 1349, 1810, 1497]. Mutagenesis-Based [1289]. Mutant [1945, 554, 640, 612, 781]. Mutants [343, 215]. Mutated [1193]. Mutation [1775]. Mutations [1393, 157, 1490, 179]. muticum [839]. Mx [1104, 1380]. Mycalazal [882]. Mycalazal-Type [882]. Mycale [882, 1491]. Mycelia [1579, 1518]. Mycobacterium [1121, 1142, 1206]. Mycosporine [1374, 687]. Mycosporine-Like [1374, 687]. Mycteroperca [20]. Myf [277]. Myf-5 [277]. Myf5 [310]. MYH [1693]. myh14 [1693]. mykiss [1181, 1292, 144, 1596, 735, 550, 517, 717, 1381, 134, 105, 1378, 1052, 1958, 874, 1708, 1318, 1770, 1176]. Mylopharyngodon [1658, 1612]. Myoblast [1490, 1789]. MyoD [1381, 546]. Myogenic [1473]. Myogenin [310]. myomaker [1490]. myomiRNAs [1332]. myomiRs [1332]. Myomixer [1789]. Myonecrosis [1796]. Myopathy [1789]. Myosin [1693, 1645]. Myostatin [604, 203, 1878, 625]. Myotubes [1337]. Mystus [1521]. Mytilid [672]. Mytillus [665]. Mytilus [1585, 528, 119, 1158, 1288, 1408, 1548, 118, 954, 98, 499, 114, 988, 297]. Mytimycin [988]. MytM [988]. Myxine [830]. Myxinidin [830].

n [1747, 1354, 826, 163, 1289, 703, 82, 297]. N- [1289]. n-3 [1747]. N-Acetylglucosamine-6-phosphate [82]. N-Acylhomoserine [1354]. N-linked [703]. N-Nitroso-N-methylurea [163]. N-oxide [826]. N-Substituted [297]. N1 [452]. N1-1 [452]. N16 [1423]. N19 [1423]. N38 [1749]. Na/K [307]. Na/K-ATPase [307]. NA1 [643, 574]. NaCl [31]. Nacre [642, 658, 1536, 1954, 1554, 1842, 1836, 150, 594, 1774]. Nacrein [697, 703]. Nacrein-Related [697]. Nacreous [1954, 322, 125, 1423, 122]. NADH [320, 442, 500]. Naked [651, 191]. namaycush [352, 9]. nanhaiensi [1751]. Nannochloropsis [815, 941, 554, 1624, 983, 956, 1090]. Nanomaterials [1368]. Nanomatrices [1319]. Nanoparticle [1806]. Nanoparticles [1141, 1453, 1899]. Nanos1 [1739]. Nanos2 [1739]. Nanos2/3 [1739]. nanos3 [1178, 1539]. Nanosheets [1593]. Nanostructurally [944]. Naphthol [1935]. Narcotic [58]. Nardo [1046]. Nass [144]. natalis [1465]. Natural [1746, 1317, 934, 58, 567, 222, 609, 224, 1700, 575, 172, 904, 1140, 1650, 891, 689, 960, 1127]. Naupliar [251]. NDP [19]. Near [860]. Near-Isogenic [860]. Necessary [172]. Necessity [892]. Necrosis [120, 134, 1303, 189, 115, 1375, 1792, 1794]. Nectin [1327]. Negative [1488]. Negombata [100]. nelsoni [130, 21]. Nematocyst [798]. Nematocysts [1120, 864]. Nemertea [1876]. Neocaridina [1291, 1619]. Neogastropoda [1022]. neogracile [894]. Neomale [1913, 1242]. Neoparamoeba [1194]. Neoplasia [163, 1107]. Neopyropia [1738]. Neotropical [823]. Nephrops [393]. Nereididae [250]. nerka [504]. Nervous [1303, 391, 1880, 1375, 1792, 1794]. Nested [456]. Network [1870, 1892, 1750, 1533, 1532, 1338, 1342]. Networks [1852, 1815, 1604, 1655]. Neural [149, 624]. Neuropeptide [190]. Neurotoxin [311]. Neutral [1090]. newberryi [186]. Newer [1691, 1743]. Newly [370, 1827, 23, 1037, 989]. Next [1290]. Next-Generation [1290]. NF [911, 1510]. NF- [911, 1510]. NgAgo [1424]. niacini [1246]. Nibea [1550, 1769]. Niche [729, 1828]. Niche-Mimic [729]. NIES [1896]. NIES-4635 [1896]. niger [1756]. nigrum [1006]. Nile [433, 1183, 1418, 959, 1402, 1393, 468, 1581, 1878, 1473, 1561, 1523]. niloticus [433, 1183, 36, 959, 128, 67, 846, 1878, 1561, 1523]. Nine [333]. NIOTVJ5 [1495]. niphobles [1105, 1405]. nippona [1883]. nipponense [1773]. Nisin [1292]. Nitrate [320]. Nitratireductor [1854]. Nitric [1245]. Nitrification [1680]. Nitriles [1501]. Nitrogen [1811, 346, 1726, 1322, 1824, 1369, 1390, 1828]. Nitrogen-Limited [1726]. Nitrogenase [612]. Nitrogenases [794]. Nitroimidazole [258]. Nitroreductase [888]. Nitroreductase-mediated [888]. Nitroso [163]. Nitzschia [1896, 1744, 1263]. NKBG [1419]. nm23 [19]. NNV [1696, 1792]. No [94, 228, 240, 474]. nobilis [1798, 1428, 1556]. Nocardiopsis [679]. Nocodazole [581]. Noctiluca [686]. Nodules [1807, 763, 1029]. Non [1368, 1568, 1002, 177, 1672, 1280, 1941, 1381, 1660, 837, 1266, 1338, 1610, 1385, 983, 956, 1767, 1769, 1652, 1412, 948, 860]. Non-coding [1660, 1338, 1610, 1767, 1769, 1652, 1412]. Non-Cytotoxic [1568]. Non-destructive [983, 956]. Non-fibrillar-Related [1002]. non-GM [860]. Non-invasive [1385]. Non-Model [1280, 1672]. Non-ribosomal [837]. Non-self [948]. Non-specific [1941, 1266]. Non-target [1368]. Non-traditional [177]. Nonbioluminescent [416]. Noncoding [768, 1560, 1832, 442, 500, 1318, 1783]. Nonfibrillar [1046]. Nonhydrolyzable [237]. Nonhydrothermal [502]. Nonmodel [696]. Nonprostanoid [364]. Nonprostanoid-Containing [364]. Nonribosomal [788, 789]. Nonselectable [32]. Nonspicule [873]. Nonspicule-forming [873]. Nori [1015]. Normal [838, 1242, 607, 1790, 1736]. Normalization [922]. Normalizing [922]. North [540, 18, 619]. Northern [1809, 144, 11, 421, 315, 1729, 113, 1511, 1915]. Northwest [495]. Northwestern [556, 229]. norvegicus [393]. Norway [393]. Norzoanthamine [842]. Nostoc [612]. Note [1707]. nov [1491]. Noval [998]. Novel [1368, 1098, 1468, 1358, 1241, 790, 335, 138, 41, 637, 61, 1931, 1773, 1463, 1156, 1168, 826, 103, 1045, 1748, 645, 1643, 1339, 1954, 1901, 801, 1289, 1198, 949, 595, 1398, 1383, 840, 453, 827, 1275, 1017, 997, 65, 202, 1470, 1125, 1547, 1639, 755, 1050, 1685, 1423, 267, 1953, 1037, 1205, 79, 1116, 1755, 674, 616, 1933, 1669, 1033, 1451, 1447, 1683, 1682, 347, 577, 1785, 1199, 74, 830, 946, 1710, 246, 1778, 656, 1203, 685, 1139, 492, 1551, 306, 691, 594, 788, 789, 943, 1496, 1387]. Novo [1357, 1840, 1223, 1204]. NP20 [1756]. NPY [1836]. Nrf2 [1604]. NRPS [1045, 788, 789]. Nuclear [141, 768, 315, 1764, 1111, 30, 1896, 1024, 225, 849, 1177]. Nuclease [1028, 1316]. Nucleases [1220, 1176, 1166]. Nucleation [1954, 747]. Nuclei [152]. Nucleic [934, 1141, 753]. Nucleic-acid-based [753]. Nucleo [780]. Nucleo-cytoplasmic [780]. Nucleotide [250, 1830, 1070, 538, 1251, 1659, 1467, 1957, 459, 1561]. Nudibranchs [798]. nudus [1099, 1710, 1062]. Null [242]. Number [601, 1355, 356]. Numbers [1525, 1552]. Nutraceutical [1641]. Nutrient [1192, 737, 1341, 1297, 1353, 1237]. Nutrition [1889, 1225, 542]. Nutritional [1009, 1887, 1859]. Nutritive [781]. NV [969]. NZ2114 [1917].

O [646, 359]. O-Glycosylhydrolases [359]. O2 [579]. ob [808]. ob/ob [808]. obesity [808]. obscurus [204]. Observed [99, 836]. Obtain [882]. Obtained [1698, 1193, 107, 1923]. Occluded [1616]. Occurrence [97, 982]. Ocean [876, 1809, 1356, 1662, 1716, 1494, 619, 485, 229, 1082, 763]. Oceanic [1314]. oceanica [903, 1624, 44]. Oceanimonas [1475]. ocellatus [316, 1081]. ochroleuca [1530]. Octocoral [1011, 759]. Octocorallia [458, 442, 500]. Octocorallians [573]. Octocorals [652, 58]. Octopus [1884, 1058]. Octylphenol [673]. oculata [727, 867, 554, 1090]. Odontesthes [834]. odorabile [1264, 933]. Oestradiol [1145]. Oestradiol-17 [1145]. off [1828]. officinalis [1753]. Offspring [1405, 1955, 1242]. ohmeri [692]. Oil [373, 1477, 138, 464, 1322, 1824]. Oil-Accumulative [1824]. Oil-based [1322]. Oils [1190, 1163, 542]. Oily [1548]. Okamejei [1869]. okamuranus [379]. Okinawan [364, 1623]. Old [228]. Oleaginous [1457, 1771, 1856, 1930]. oleracea [1819]. Olfactory [1283]. Oligo [877, 1012, 1072]. Oligo-Microarray [1072, 877, 1012]. Oligoalginate [1289, 1416, 1037]. Oligomers [1330]. Oligonucleotide [975, 120, 288]. Oligonucleotides [278]. Oligosaccharide [945]. Oligosaccharides [1337, 1641, 379, 703]. olivaceus [63, 1282, 1343, 103, 1010, 142, 1945, 435, 1163, 1456, 1862, 244, 214, 1201, 546, 592, 705, 561]. Olive [1343, 1945, 561]. Oman [991]. Omega [1190, 1411, 1207, 1461, 1155, 943]. Omega-3 [1190, 1411, 1207, 1461, 1155, 943]. Omega-3/-6 [1461]. omics [1928]. OmpA [1832]. On-site [1367, 723]. Onboard [446]. Oncorhynchus [1181, 1292, 144, 1637, 1596, 504, 735, 550, 517, 717, 1381, 134, 105, 1600, 1283, 407, 1378, 1622, 1052, 1958, 874, 1708, 1318, 1770, 1747, 1176]. One [1374, 1751, 1259]. One-Pot [1751]. Only [549]. onnurineus [643]. Ontogenetic [624]. Ontogeny [327, 598, 230, 1462]. Onychostoma [1822, 1525, 1552]. Oocyte [1527]. Oogenesis [1647, 1908]. Oogenesis-Related [1647]. Oomycete [1231]. opalescens [54]. Open [71]. Opening [1813]. Opens [416]. Operation [65]. Ophiuroidea [87]. Opine [439]. Oplegnathus [1720, 1590, 1834, 1784, 1952, 1112]. OPMS [1654]. Opportunities [1247]. Opsariichthys [1812]. Opsonic [948]. Optical [983, 956]. Optimal [429]. Optimise [799]. Optimization [1135, 497, 296, 1877, 1372, 1406, 1173, 986, 890]. Optimized [1737, 1177]. Oral [1294, 466, 698, 1077, 1388, 1917]. Orally [1541]. Orange [533, 645, 604, 1527, 1833]. Orange-Spotted [533, 1527, 1833]. Oreochromis [274, 433, 1183, 291, 36, 959, 1514, 128, 67, 846, 1878, 1561, 1523]. orfA [919]. Organ [1071, 1338]. Organelle [1425]. Organelle-Specific [1425]. Organic [642, 1061, 1384, 545, 1940, 573, 1920, 331, 1774]. Organism [563, 796]. Organisms [259, 354, 876, 683, 1138, 915, 677]. Organization [1424, 413, 524, 587, 57, 521, 92, 561]. Organized [944]. Organogenesis [1152]. Organs [443, 527, 98]. Oriental [1773]. Origin [1292, 408, 260, 467, 763, 800]. Originally [1622]. Ornamental [1619, 726, 1825]. ornata [211]. Ornithine [435]. Ortholog [779]. Oryzias [123, 1250, 267, 157, 741, 597, 1950]. OsHV [1417]. OsHV-1 [1417]. Osmerus [896, 724]. Osmolyte [826]. Osmoregulation [1815, 31, 1360]. Osmotic [1815, 1810]. Osteoblastic [1940]. Osteoclastic [1763]. Osteogenic [1929]. Osteoinduction [1678]. Osteoinductive [1694]. Osteoporotic [1763]. Ostrea [814, 1107]. Ostreid [1417]. Ostreidae [1962]. Ostreopsis [725, 206]. Other [1252, 104, 177, 391, 740, 1244, 1859, 771]. Oulactis [1569]. our [1707]. Outdoor [1095, 355]. Outer [1832, 951, 899, 589]. Output [1851]. Ovarian [1661, 644, 657, 1890, 1894, 1943, 921, 1717, 1791, 681, 1567]. Ovaries [687]. Ovary [1951, 1335, 852, 260, 1428, 1412, 1601]. ovata [206]. Over-expression [1184]. Overexpressing [240]. Overexpression [1487, 780, 1295, 938, 1485, 924, 121]. Overlooked [1323]. Oversulfated [10]. Overview [1641, 1420]. Oviparous [669]. Ovulation [245]. oxalicum [1609]. Oxidase [452, 496]. Oxidation [970]. Oxidative [1702, 1741, 1408, 339, 1055, 1453, 773, 1559, 1604, 862]. Oxide [1806, 1751, 1593, 1245, 826]. Oxidizing [1084, 1137]. Oxidoreductase [914, 1042]. Oxime [1430]. Oxygen [284, 1408, 227, 1909, 1387]. oxylipins [1162]. oxysporum [1241]. Oyster [642, 658, 90, 216, 1731, 999, 1805, 130, 866, 1860, 1156, 1279, 720, 1883, 1039, 1816, 460, 1348, 1057, 909, 256, 1643, 1305, 1160, 1687, 659, 1146, 1554, 1030, 132, 21, 814, 821, 390, 418, 844, 1376, 1571, 1645, 1660, 1844, 1566, 1017, 1107, 529, 1728, 125, 1423, 456, 1385, 1315, 1154, 756, 1845, 1881, 1683, 1682, 347, 360, 885, 1108, 1324, 1460, 1466, 1467, 1534, 536, 1780, 1452, 1440, 233, 479, 648, 931, 1036, 1356, 1443, 1716, 656, 1391, 1842, 1836, 447, 1507, 1427, 1714, 150, 388, 518, 594, 1157, 1582, 1086, 1310, 1690]. OysterDB [1962]. Oysters [923, 583, 1417, 1038, 1115, 228, 1852, 1934, 1926, 193, 321, 490, 1709, 1829].

P [1886, 1671, 30, 202, 701, 1615]. P-450 [202]. P. [780]. P1 [479]. p10 [594]. P16 [1268]. P2 [25]. p21 [780]. p32 [247]. P450 [48, 94, 45, 47, 43]. p53 [275]. P6 [1339]. p63 [622]. p63/p73 [622]. p63/p73-Like [622]. p73-Like [622]. pachyptila [1383]. Pacific [619, 763, 578, 496, 495, 90, 1270, 556, 999, 404, 583, 1417, 866, 1038, 1860, 1004, 1279, 1039, 609, 1855, 1348, 1519, 1553, 1057, 1305, 1934, 1926, 1687, 540, 1030, 1941, 873, 821, 357, 390, 418, 526, 1571, 1645, 1660, 1844, 570, 1566, 1017, 1728, 456, 1479, 229, 1909, 756, 1683, 1682, 347, 1087, 360, 1325, 1466, 1467, 1534, 113, 1780, 437, 233, 1356, 1586, 1668, 1829, 1443, 1716, 1391, 1915, 751, 1377, 1446, 1513, 1507, 1524, 1427, 848, 1310, 1690]. pacifica [672]. pacificus [495]. Packaging [1772]. Pagrus [443, 1407, 1447, 1859]. pagurus [1255]. Paints [1463]. Pair [577]. Paired [1443]. Palaemonidae [493]. Palaeomonetes [222]. Palaeonemertea [1876]. Palisada [1406]. pallida [1574]. Palythoa [1229]. Panama [1270, 437]. Pancreas [1235]. Pancreatic [120, 189]. Panel [513, 1848]. Panels [1440]. Pantropic [164, 162, 231]. Panulirus [1260, 415, 556, 676]. Papulosum [429]. Paracentrotus [1328, 824, 1327]. parahaemolyticus [1941, 1782]. Paralichthys [63, 1282, 1343, 103, 1010, 142, 1945, 435, 1163, 1456, 1862, 244, 214, 1201, 546, 592, 705, 561]. Paralithodes [106]. Paralogs [1280]. Paralytic [313]. paramamosain [1925, 1608, 1509, 1717, 1791, 1412, 1601]. Parameters [1087, 1166]. Parapeneus [1529]. Parapristipoma [411]. Parasite [1504, 1179, 1648]. Paratope [1257]. Parentage [232, 1440]. Parents [1405]. Park [1803, 1786]. Part [1920]. Partial [381, 986, 453]. Partially [373]. Participate [1064]. Participates [1643, 1954, 1901, 1836]. Participation [1919]. Particle [241]. Particles [169, 189, 368, 113, 1792, 1025]. Particulate [440]. Parts [204]. parvum [584]. Passable [992]. Past [170]. Paste [1607]. pastoris [361]. Patagonian [834]. Patella [1113]. Patellidae [1113]. Paternal [960]. Paternity [393]. Pathogen [1264, 21, 1480, 1088, 1617, 996, 1298, 1782]. Pathogenesis [279]. Pathogenic [1121, 1206, 1441, 1450, 1069, 1488, 1879]. Pathogenicity [969, 1922]. Pathogens [1704, 1691, 963, 1868, 1953, 283]. Pathway [911, 1860, 947, 1713, 1872, 1827, 1958, 343, 1510, 1601]. Pathways [1445, 1715, 1926, 1353, 1335, 1607, 1521, 1712, 1874]. Patinopecten [1335, 1498, 1540, 687, 220, 221, 1544, 1379, 1535, 1577]. Pattern [27, 203, 137, 1110, 1761, 1458]. Patterns [699, 606, 1724, 923, 716, 639, 909, 951, 1750, 1628, 1250, 1963, 1610, 127, 891, 1534, 546, 1684, 1342]. Paul [962]. Pavillard [1066]. Pavlova [1654, 755, 795]. Pax [140]. Pay [1323]. PcarnBase [1114]. PCC [612]. PCC9438 [425]. PCDD [1351]. PCDD/F [1351]. PCP [305]. PCR [516, 366, 528, 1594, 1681, 130, 147, 1279, 684, 510, 288, 1413, 884, 448, 21, 1784, 407, 204, 741, 759, 758, 446, 431, 518]. PCR-Based [366, 1279, 1784, 431]. Pdp11 [1898]. Pds [737, 695]. Peak [1888]. pealeii [137]. Pearl [613, 807, 1448, 1731, 923, 1156, 720, 460, 1115, 909, 951, 1643, 1160, 1502, 659, 1146, 1554, 844, 1376, 1423, 1108, 1324, 1460, 536, 1452, 931, 1036, 656, 1842, 1836, 388, 594, 1582, 1086]. Pearlin [125]. Pearls [909, 125]. Pecten [528, 1444]. Pectinaria [86]. Pectinidae [633, 623]. Pectobacterium [1505]. Pedunculata [1484]. Pejerrey [834]. Pelodiscus [1900]. Pelteobagrus [1555, 1119, 1275]. Pen [949]. Penaeid [1126, 280, 1481, 1841, 964, 644, 115, 880, 967]. Penaeidae [908]. Penaeidin [470]. Penaeidin-like [470]. Penaeus [1161, 578, 470, 85, 34, 292, 1871, 461, 1941, 17, 226, 262, 424, 1364, 62, 356, 301, 1674, 1669, 698, 1244, 76, 117, 304, 89, 361, 921, 243, 133]. Penelope [1103]. Penelope-Like [1103]. Penetrating [1141]. penicillatus [1894]. Penicillium [1028, 1579, 1295, 1395, 1656, 1518, 1609]. Pennate [1096, 1896]. Pentanucleotide [578]. Peptic [213]. Peptidase [606, 778]. Peptidases [1312, 1745]. Peptide [280, 470, 934, 29, 531, 7, 837, 632, 1470, 761, 988, 830, 1234, 42, 314, 1599, 1917, 788, 789]. Peptides [262, 1141, 512, 298, 76, 213, 1583, 880, 967, 634, 109, 1778, 1420]. Peptidoglycan [750]. Perca [774]. Perch [610, 774, 1890, 771, 1559, 1604, 625, 1237]. Perciform [1013]. Percula [1873]. perforata [1406]. Performance [1904, 1811, 1170, 1955, 889, 877, 67, 1202, 1632, 1813, 1927, 1900]. Peridinin [1210]. Peridinin-Producing [1210]. Perinereis [250]. Periodontitis [1434]. Peritrophic [1871]. Peritrophin [657]. Perkinsus [126]. perlevis [820]. Permeable [781]. Perna [569, 224, 1441, 881, 689]. Peroxidase [1728, 1184]. Peroxide [773, 862]. Peroxisomal [1184]. Persistence [613]. Perspective [603, 575, 1745]. Perspectives [856, 1570, 1247, 416]. pertusa [16]. Peru [1828, 1909]. Pesticide [1340, 558]. petechialis [1928]. Petrosia [449]. PfMSX [1572]. PfSMAD1 [1572]. PfSMAD1/5 [1572]. PfSMAD4 [1572]. Ph [1433, 1553, 1308, 840, 215]. pH-Sensitive [1308]. Ph1 [1546]. Phaeobacter [1153]. Phaeodactylum [757, 1221, 1128, 1203]. Phaeophyceae [1167, 1409, 1322, 853, 1390]. Phaeophyta [987, 829, 871]. Phage [1097]. Phagocytic [1941]. Phagocytosis [1950]. Phakellia [1084]. Pharmaceutical [408, 1923]. Pharmacology [391]. Pharmacy [473]. PHB2 [1867]. Phenocopies [343]. Phenoloxidase [1016, 297]. Phenotype [1642, 1825]. Phenotypes [1346]. Phenotypic [1144]. Phenylacetonitrile [1101]. Phenylacrylaldehyde [1579]. Phenylalanine [1751]. Phenylpropanamide [1576]. phenylpyruvicus [154]. Pheromone [1452]. Pheromones [1753]. Philasterides [1072]. philippinarum [1961, 997, 1706, 1932]. Philippine [402, 363]. Phlorotannin [1415]. Phlorotannin-Rich [1415]. Phlorotannins [572]. pholis [104]. Phosphatase [1232, 234, 659, 1128, 475, 656]. Phosphate [1506, 1748, 786, 31, 82]. Phosphate-Deficient [786]. Phosphoinositides [813]. Phospholipase [713]. Phosphoproteomic [1826]. Phosphorodiamidate [278]. Phosphorus [30]. Phosphorus-Containing [30]. Phosphorylation [780, 1713, 817]. Photoautotrophic [995, 1516]. Photobiological [615, 612]. Photobiomodulation [1920]. Photobioreactor [1210, 715]. Photobioreactors [1095, 901]. Photodimer [342]. Photolytic [343]. Photon [261]. Photoperiod [1188]. Photoproduction [445, 346]. Photoprotection [544]. Photoprotective [1656]. Photosynthesis [893, 155, 1824]. Photosynthetic [941, 1261, 901]. Photosynthetically [1638]. Phthalate [1330]. Phyllosoma [556, 676]. Phylogenesis [4]. Phylogenetic [496, 809, 1270, 219, 973, 1271, 364, 371, 661, 451, 641, 1336, 1121, 1043, 202, 1229, 334, 108, 623, 246, 26, 488]. Phylogenetically [1421, 1137]. Phylogenetics [463]. Phylogeny [1271, 1280, 462, 223, 374, 1008, 1022]. Phylogeographic [1761, 506]. Physical [330, 433, 959]. Physicochemical [980]. Physio [1806]. Physiologic [28]. Physiological [1594, 1681, 1926, 1881, 1516, 1356, 1849, 1895]. Physiology [1331]. Phytase [692]. Phytochelatin [541]. Phytochemical [1415]. Phytoene [1221]. Phytopathogenic [1198, 792]. Phytoplankton [667, 953]. Phytoplankton-Group [667]. PI3K [1713]. PI3K/Akt/GSK [1713]. PI3K/Akt/GSK- [1713]. piceus [1658, 1612]. Pichia [452, 361]. Pictet [1191]. Piezotolerant [1495]. Pif [1106]. Pigment [301, 983, 956]. Pigment-Dispersing [301]. Pigmentation [1526, 1889, 1540, 1728, 1620, 1679, 1460, 1706, 1690]. Pigmented [1558].

Pigments [1517]. pilchardus [266]. Pilot [1061]. Pinctada [658, 1448, 1731, 923, 720, 909, 1643, 1160, 659, 1146, 844, 1376, 1352, 1423, 1108, 1324, 1572, 1460, 1234, 703, 536, 1452, 931, 1036, 656, 1836, 150, 388, 594, 1582, 1749, 1086]. Pink [1558]. Pink-Pigmented [1558]. pinnatifida [1478]. Pioneering [1744]. piRNA [1062]. piRNA-Like [1062]. piRNAs [1733]. Pisces [605, 104]. piscicida [1950]. Piscine [805, 183, 498]. Piscirickettsia [1675]. pistillata [1001]. Pituitary [201]. PKS [1045, 1476]. PL10 [636]. Plankton [456]. Planktonic [1224]. Planocera [1800, 1949]. Planocerid [1937]. Plant [1463, 1955, 760, 1253]. Plants [768, 542]. Plaque [118]. Plasma [1675, 116, 1508]. Plasmid [217, 187, 466, 3, 728, 179, 800]. Plasmids [1734]. Plasticity [173, 1829, 1574]. Plastid [1366, 1409, 467]. Plastids [525]. Plate [354, 1433, 114, 355]. Plated [756]. Platelet [363]. platensis [425, 28, 213, 1803, 1786, 507]. Platforms [1228, 1594, 1681]. Platichthys [558]. Platygyra [1114]. Platyhelminthes [1621, 1949]. Plays [844, 1644]. Playtygyra [381]. Plecoglossus [422]. Pleiotrophin [1756]. Pleurochrysis [651, 523, 838, 981, 257, 607]. pleuronectes [633]. Pleuronectiformes [4, 451, 108]. Plexaurella [734]. plicatilis [80, 1009]. Ploidy [908, 1793]. plumbeus [204]. Pluripotency [610]. Pluripotent [669]. pluvialis [237, 1599]. Pmel17 [1752]. pMP1 [800]. PmRab7 [712]. Pocilloporid [1277]. Pocilloporidae [731]. Poecilia [382, 1676, 290, 353, 170]. Pogonophoran [351]. Poison [313]. Polar [762, 210, 131, 943]. Pollachius [799]. Pollicipes [1484]. Pollution [699, 1431]. Pollution-Affected [699]. Poly [1941]. Poly- [1941]. polyactis [1761]. Polyadenylation [622]. Polyamines [1196]. Polychaeta [250]. Polychaete [1364, 86, 804]. Polychaets [250]. Polycladida [1621, 1949]. Polycyclic [1491]. polydactyla [618]. Polyextremophilic [1449]. Polyglucans [620]. Polyguluronate [83]. Polyhydroxyalkanaotes [1050]. Polyhydroxyalkanoate [1607]. Polyketide [987, 1080, 837, 318]. polykrikoides [448]. Polymerase [569, 141, 643, 934, 135, 296, 195, 27, 667, 571, 153, 686, 115, 456, 564, 1629, 84, 200, 1795]. Polymeric [261]. Polymetallic [763, 1029]. Polymorphic [34, 916, 193, 59, 356, 86]. Polymorphism [709, 196, 1083, 490, 1251, 352, 608, 221, 360, 117, 1598, 348, 459, 1561, 592, 593]. Polymorphisms [1830, 538, 1076, 151, 209, 1957]. Polynesian [198]. Polyp [1617]. Polypeptidic [1753]. Polyploid [908, 1677]. Polysaccharide [1723, 1241, 1321, 1063, 477, 1936, 1923, 319, 1433, 1390, 1284, 1912, 492]. Polysaccharide-Degrading [1936]. Polysaccharide-Rich [1923]. Polysaccharides [1530, 947, 523, 1363, 838, 913, 10, 1451, 1048, 1006, 1633]. Polyunsaturated [1760, 370, 1207, 616]. Polyurethane [553]. Polyvinyl [980]. Pomc [1644]. Pomfret [1887, 1872]. Pond [1624]. Ponds [1697, 802, 810]. Population [1181, 709, 675, 144, 495, 605, 316, 421, 621, 219, 406, 1571, 1837, 455, 1622, 333, 221, 530, 151, 244, 481, 117, 450, 1440, 437, 1204, 1903, 1551, 775, 1414]. Populations [37, 976, 1271, 315, 1793, 1084, 609, 224, 153, 290, 17, 1533, 1532, 208, 87, 129, 172, 1845, 360, 348, 302]. porasa [325]. Porgy [276]. Porifera [449, 982]. Porites [1558]. Porphyra [1530, 1005, 813, 1015, 342, 1051, 849, 1109]. Porphyridium [620]. PORT2 [1692]. Portulaca [1819]. Portunus [1361]. Posidonia [903]. Positioning [1092]. Positive [741, 1046, 974, 1488]. Possesses [768]. Possessing [286, 1289, 1132]. Possible [392, 842, 1678, 682]. Post [1831, 1675, 1133, 1679]. Post-feeding [1831]. Post-Smolt [1675, 1679]. Post-translational [1133]. Postbiotic [1898]. Posterior [1341]. Postglacial [481]. Postmortem [197]. Postovulatory [1347]. Postprandial [1831]. Pot [1751]. Potency [1801]. Potent [1333, 1307, 1313, 885, 1259, 1276, 1386]. Potential [1415, 845, 259, 1545, 1358, 373, 269, 308, 235, 828, 1491, 1898, 1715, 1004, 1111, 1475, 1351, 256, 784, 1821, 1673, 1131, 1281, 73, 1902, 72, 1335, 1923, 1670, 1150, 1623, 1365, 1325, 1428, 974, 1656, 484, 1699, 1638, 1127, 1509, 1717, 1791, 1921, 1861, 1427, 705, 1257, 336, 1633, 506]. Potential-Controlled [1131, 1281]. Potentially [1683, 1682, 1932, 781, 1086, 1912]. Potentials [723]. POU [140]. POU2F1 [1728]. Ppar [1482, 1591]. Practical [465, 431]. Prawn [606, 395, 1938, 1773, 1806, 1170, 262, 664, 1499, 301, 76, 1894, 431]. Pre [1346, 1224]. Pre-screen [1224]. Pre-Selection [1346]. Precoated [337]. Precocious [1642]. Precursor [771]. Predator [68]. Predetermined [1370]. Predict [1812, 467]. Predicted [1731]. Predicting [1474]. Prediction [1744, 1338, 1318, 1850, 1564]. Prefecture [1111, 1622]. Preferential [228, 1239, 1109]. Preferentially [1718]. Preliminary [783, 764, 902, 570, 422]. Premise [1772]. Premium [1562, 815]. prenanti [1429]. Preparation [454, 1653, 1241, 944, 929, 1306, 746, 55, 1386, 1388]. Preparations [985]. Preparing [1438]. Prerequisite [322]. Prerigor [197]. Presence [1703, 322, 237, 1915]. Present [170]. Preserved [684, 200]. Pressure [486, 918, 1810, 1400]. Preston [131, 710]. Pretreated [1549]. Prevention [575]. Prevents [1450, 1842]. Previously [1394, 1455, 1458]. Primary [893, 303, 210, 720, 682, 424, 122, 1669, 220, 1617, 1422]. Primed [153]. Primer [577]. Primers [569, 91, 288, 372, 93, 1915]. Primitive [1703]. Primmorphs [979, 1123, 449]. Primordial [1732, 1435, 1136, 1500, 1539]. Prince [971]. Principles [1125]. Printed [1823]. Prints [1216]. Prior [1661]. Pro [1529]. Pro-apoptotic [1529]. Probe [366, 77, 1748, 1066, 499]. Probes [73, 467]. Probing [1700]. Probionts [555]. Probiotic [1673, 1171, 1153, 1218]. Probiotics [1450, 671]. Problem [178]. Procambarus [1880, 1947]. Procedure [1735]. Process [613, 101, 1212, 1173, 1630, 1543, 1012, 1072, 994, 1249, 1296, 1652, 1427, 715]. Processed [260]. Processes [1858, 896, 943]. Processing [1730, 1133, 1882]. Prochloron [155]. Produce [399]. Produced [338, 1189, 1248, 1060, 1321, 258, 1339, 827, 1278, 879, 1144, 231, 1006, 131, 679]. Producer [1096]. Producers [359]. Produces [766, 1198, 549, 1344, 1584]. Producing [1210, 1449, 1635, 1877, 1156, 370, 1336, 227, 1224, 616, 343, 685, 713]. Product [58, 567, 904, 1127, 1730, 1420]. Production [1287, 414, 911, 893, 1292, 1487, 1419, 1374, 1576, 46, 539, 1080, 1190, 1182, 1913, 75, 678, 1372, 941, 734, 1295, 1095, 1740, 1232, 1406, 1405, 436, 383, 1641, 477, 1339, 1265, 1654, 1929, 803, 1336, 1173, 469, 403, 1442, 986, 1094, 1730, 1817, 285, 632, 1268, 1195, 65, 491, 342, 398, 562, 71, 761, 1953, 1231, 875, 1755, 1450, 740, 1909, 542, 1666, 1140, 1382, 615, 729, 162, 283, 1451, 1129, 1489, 995, 1090, 811, 1516, 1803, 1786, 1021, 890, 870, 1349, 289, 851, 1766, 612, 781, 497, 945]. Productivity [1009, 1624]. Products [1746, 1317, 1330, 575, 260, 689, 1261, 1529]. Profile [1227, 750, 1801, 309, 519, 1600, 1620, 1298]. Profiles [370, 1661, 1165, 1242, 1402, 1012, 1133, 1072, 1244, 1880, 1509, 744, 1494, 1733]. Profiling [1727, 1448, 1893, 1357, 1290, 1938, 1035, 651, 1252, 806, 639, 618, 1779, 1738, 821, 1628, 1338, 1371, 1890, 1130, 1185, 1959, 860, 1581, 1652, 1008, 1765, 926, 1573, 907, 1362, 1535, 1577, 705]. Progenies [1913]. Program [404]. Progression [1568, 158, 1434, 1905]. Prokaryotes [1054]. Prokaryotic [1893, 759]. Prolactin [29, 201]. Prolidase [26]. prolifera [831]. Proliferated [1718]. Proliferation [919, 38]. proliferative [1365]. Prolyl [1258, 1288]. Promise [1691]. Promising [903, 769, 1835, 1070, 1508]. Promoter [430, 35, 566, 349, 47, 429, 1325, 700, 728, 310, 1109, 277, 522, 482]. Promoter-based [728]. Promoter-Driven [700]. Promoters [101, 350, 1833, 522]. Promotes [1207, 1090, 1500]. Promoting [615]. Promotion [1616]. Propagation [281]. Properties [1410, 980, 1899, 420, 106, 1679, 1325, 983, 956, 1516, 1420]. Prophage [239]. Proposed [171]. Prorocentrum [801, 1935, 1066]. Prospecting [1743]. Prostaglandin [1233]. Prostanoid [364]. Prostanoid-Containing [364]. Prostate [1478]. Protandrous [276]. Protease [826, 1901, 1173, 632, 1863, 706, 761, 1100, 76, 787, 1349, 507, 990]. Proteases [415, 637, 1726]. Protect [1292, 698, 862]. Protection [1161, 339, 1453, 1125, 687, 796, 1025]. Protective [565, 950, 1187, 646, 1434, 1350]. Protects [1055]. Protein [527, 1104, 1448, 1558, 1698, 1653, 64, 274, 235, 148, 1354, 270, 119, 720, 1832, 894, 832, 816, 1380, 118, 645, 590, 1643, 1799, 1165, 1502, 1954, 1955, 1725, 955, 218, 247, 1686, 1352, 453, 99, 586, 1863, 813, 322, 125, 1120, 765, 1133, 1597, 1640, 1616, 33, 1888, 726, 1051, 1205, 1214, 299, 1433, 1033, 220, 885, 1223, 690, 970, 1694, 1023, 60, 1253, 817, 264, 855, 747, 1388, 133, 1420, 594, 781, 920, 1197, 1749, 1539, 1690]. Protein-3 [274]. Protein-Based [1253]. Protein-Coupled [1725]. Proteinaceous [864]. Proteinase [6, 642, 461, 487]. Proteins [1227, 658, 1061, 22, 1702, 1741, 1158, 1168, 909, 444, 1200, 666, 218, 1000, 1141, 1017, 997, 745, 629, 733, 1616, 697, 1216, 771, 1484, 707, 749, 25, 1583, 1106, 536, 849, 416, 1716, 899, 589, 499]. Proteoglycans [122]. Proteolysis [197]. Proteolytic [464]. proteolyticus [1607]. Proteome [662, 1085, 1202, 1489, 1428, 718, 1078]. Proteomic [1061, 1312, 1145, 953, 499, 997, 1269, 1133, 860, 1296]. Proteomics [532, 658, 1826, 629, 802, 810]. Protist [672]. Protists [477, 740]. Protocol [1070, 188, 989, 1927]. Protogynous [1732, 1541]. Protoplast [1083, 929, 607]. Protoplast-Derived [1083]. Protoplasts [836, 55]. Provide [1045, 1540, 1709, 1928, 1719]. Provides [1238, 1773, 1348, 1533, 1532, 1759, 1269, 1315, 1959, 1808, 1765, 1412, 1846, 1714, 1117, 1535, 1577, 1761]. Proximal [277]. pRSVrtGH1 [284]. Prydz [679]. Prymnesiophyceae [584]. Prymnesiophyte [127]. Prymnesium [584]. PS [1433]. PS-2 [1433]. Psetta [1153]. Pseudo [1508, 1744, 1263]. Pseudo-Males [1508]. Pseudo-nitzschia [1744, 1263]. Pseudoalteromonas [1742, 845, 1264, 1354, 1722, 801, 453, 1122, 512, 1008, 1349]. Pseudobagrus [1262]. Pseudocarcinus [493]. Pseudogene [407]. Pseudomonas [15, 259, 51, 1811, 10, 319, 1755, 1933, 1664, 747, 738]. Pseudoperkinsus [672]. pseudorotundata [1923]. Pseudosciaena [680, 1110]. Pseudosquilla [521]. Pseudosuberites [241]. PsHSP70 [1051]. PSII [544]. PSII-Repair [544]. Psy [737]. Psychrobacter [154]. Psychromonas [1150]. Psychrophile [512, 1150]. Psychrophilic [826, 1345]. psychrophilum [749, 1186]. Psychrotrophic [187, 229, 1808, 82, 738]. PtDRG1 [1459]. PtDRG2 [1373]. Pteraeolidia [427]. Pteriidae [923]. PTM [1491]. PU14 [1643]. PUFA [1482, 464, 875, 1129, 1522, 1591]. PUFAs [740]. Puffer [1105, 1405]. Pufferfish [176, 194, 223, 1602, 1642]. Puget [481]. pugio [222]. pulchella [832]. Pullulan [1268]. pullulans [632, 706, 761]. Pulse [1896]. pumilus [1189, 1248, 268]. punctatus [1697, 1775, 286, 40, 190, 1435, 1590, 1784, 1952, 1316, 426, 201, 870, 663, 1747]. pUR288 [179]. Pure [1199]. Purification [6, 12, 1260, 11, 392, 1358, 1321, 234, 1751, 1339, 253, 692, 453, 632, 1616, 299, 932, 362, 106, 989, 634, 525, 787, 82, 679]. Purified [1475]. Purple [1193, 299]. Purslane [1819]. Putative [1827, 1191, 519, 1296, 898, 388]. putrefaciens [1898]. Pyloric [156]. Pylorus [1600]. Pyramimonas [1200]. Pyrethroid [1340]. Pyrethroids [769]. Pyridoxine [576]. Pyrolysis [44]. Pyropia [1373, 1366, 1239, 1459, 1451, 1177]. Pyrosequencing [1059]. Pyrroloiminoquinones [1082]. Pyruvate [443, 1295]. Pyruvoyl [382].

Qiongdongnan [902]. QM [388]. qPCR [922]. QTL [1418, 1039, 1164, 735, 1503, 814, 1393, 1514, 1411, 871, 1303, 1595, 1093, 1563, 1186, 912, 1399, 1537, 1219, 1387, 1556]. QTL-Seq [1418, 1503]. QTLs [1553, 1646, 1211, 942]. Quahog [11, 1809, 1729]. Quality [1448, 909, 1202, 1878, 1362, 1895]. Quantification [1329, 1032, 510, 1162, 98, 686, 150]. Quantify [130]. Quantitative [147, 1164, 735, 1553, 517, 667, 1300, 889, 1302, 1397, 1144, 1179, 1407, 1705, 1466, 1296, 1261, 1201, 38, 1155]. Quenching [1217]. quernus [333]. Quick [446]. Quinolinic [311, 1656]. Quinone [1042, 914]. quinqueradiata [1278]. Quorum [1354, 1224, 677, 974, 1217]. Quorum-Quenching [1217].

R [1110]. R&D [615]. Rab [832]. Rabbitfish [1591]. Raceway [1624]. Rachycentron [1831, 1727]. RACK [36]. Radial [1338]. Radical [572]. RADseq [1881]. Rainbow [1181, 1292, 278, 1182, 1596, 1913, 843, 317, 1740, 735, 1212, 242, 550, 517, 717, 1242, 1381, 134, 5, 105, 1251, 1347, 1378, 1052, 1130, 1185, 249, 1144, 896, 1958, 874, 225, 1708, 60, 1186, 1272, 1318, 841, 1069, 1770, 926, 1176, 724]. Raised [615]. raistrickii [1518]. Raja [325]. Raman [1367]. Ran [1064]. Random [1227, 1349, 1642]. Randomized [1097]. Randomly [193]. Range [217]. RAPD [37, 412, 81, 353, 96, 226, 87]. RAPD-Based [81]. Raphidophytes [1095]. Rapid [516, 1424, 1291, 790, 251, 686, 432, 267, 701, 204, 1492, 1587, 564, 758, 1515, 1897, 1935, 200, 1795, 1907, 1574, 1493]. Rapid-Growth [790]. Rapidly [1796]. Rare [1588]. Rat [5, 42, 314]. Rate [377, 207, 727, 926]. Rates [1612]. Rather [1912]. Ratio [327, 378, 1119, 1461]. Rational [1191]. Rats [950, 1779, 1580, 1823]. RAW264.7 [1245]. Razor [1892, 1309, 1611]. rbcL [127]. rDNA [528, 490]. Re [1647, 1874]. Re-sequencing [1647, 1874]. Reaction [569, 141, 643, 934, 135, 296, 195, 667, 571, 153, 1191, 686, 115, 456, 564, 84, 1249, 200]. Reactions [1169]. Reactive [1909]. Reactivity [135]. Reactors [355]. Real [147, 510, 571, 884, 686, 407, 446, 1787]. Real-Time [147, 510, 571, 686, 407, 446, 884, 1787]. Reared [1126]. Rearing [1070, 1153]. Rearrangements [57]. Rec8 [1734]. RecA [340]. RecA-Mediated [340]. Receptor [275, 1775, 1167, 273, 177, 36, 1725, 225, 1867, 1293, 948, 779]. Receptors [1270, 1868, 276, 1010]. Recipient [1460, 1718]. Recipients [1740]. Reciprocal [1821]. Recognition [22, 948]. Recombinant [1487, 1886, 520, 1702, 1741, 1232, 929, 461, 1671, 189, 1616, 1953, 199, 1755, 1669, 1325]. Recombinants [1333, 1307, 1313]. Recombinase [640, 1795]. Recombination [161]. Recommendations [181]. Reconstitution [152]. Reconstruction [623]. Recorded [723]. Recovery [973, 590, 994, 1394, 1927, 1455]. Recretohalophyte [1371]. Recycling [832]. Red [443, 366, 316, 269, 1560, 1734, 1793, 1465, 406, 434, 100, 212, 35, 1111, 107, 1081, 1373, 234, 1410, 1738, 1514, 508, 852, 1600, 980, 208, 1239, 813, 342, 1665, 368, 1888, 742, 726, 977, 1493, 1492, 1587, 868, 1909, 1451, 1407, 1447, 1859, 620, 1583, 849, 1109, 1177, 264, 1778, 1521, 1712, 1947, 1782, 635, 1350, 1558, 1274, 492]. Red-Shell [1782]. Red-Spotted [1734]. Red-Tail [1521]. red-tide [492]. Red-Tide-Forming [368]. Redox [1453, 1573]. Redtail [1894]. Reduced [1192, 1751]. Reduces [1568, 1892, 1217, 1829, 1911]. Reducing [1054]. Reductase [196, 371, 320, 641, 1766]. Reductases [1442, 1040]. Reduction [1576, 1221, 1431, 1547, 1639, 1908, 1924]. Reductive [1351]. Reef [1045, 462, 1480, 690, 1542, 1264, 246]. Reef-Building [1480, 690]. Reef-Dwelling [1542]. Refeeding [1838, 1237]. Reference [922, 571, 1875, 352, 705, 1627, 1626]. Reference-Guided [1875]. Refined [238]. refringens [73]. Regenerating [1578]. Regeneration [1145, 1914, 1338, 146, 836, 1818]. Regenerative [1574]. Regimes [739]. Region [94, 430, 421, 315, 406, 1720, 905, 538, 566, 1952, 1122, 349, 823, 47, 252, 221, 4, 459, 92, 1379, 1539, 1556, 506, 305]. Regional [1272]. Regionally [1130]. Regions [768, 1314, 248, 1267, 1540, 442, 500, 1097, 1659, 277, 1513]. Regrowth [607]. Regulate [1572, 1601]. Regulated [1243, 501, 558, 1180, 1629, 1473, 833]. Regulates [1934, 1818]. Regulating [1427]. Regulation [1444, 1757, 1918, 1822, 790, 1760, 1715, 1860, 1961, 1693, 1064, 1887, 387, 666, 1482, 644, 765, 544, 127, 1959, 1522, 1591, 1110, 1604, 1770, 1846, 1714, 1649, 1895, 1685, 970, 1911]. Regulations [1402]. Regulator [1356]. Regulators [1023]. Regulatory [22, 737, 1454, 1885, 566, 1122, 1815, 1611, 194, 310, 1527, 1706, 1655, 546]. reinhardtii [541, 1649]. Related [1227, 463, 1904, 897, 606, 1357, 1290, 250, 1798, 1870, 1264, 1002, 1555, 1754, 1039, 1164, 1553, 1926, 1300, 1242, 1605, 262, 1211, 778, 1335, 1398, 1595, 1347, 1107, 1569, 1815, 733, 1397, 697, 1205, 1546, 573, 1909, 1151, 1563, 1466, 1819, 1277, 1472, 624, 1638, 1079, 1426, 1647, 1710, 1559, 1565, 1567, 1254, 942, 920, 1157, 1086, 1689, 1538, 1014, 304]. Related-Genes [1107]. Relatedness [460, 153]. Relation [650]. Relationship [412, 1336, 1700, 183, 1467, 26, 1783]. Relationships [1742, 364, 661, 451, 128, 1245, 108]. Relative [1084, 174]. Release [1294, 1469]. Releases [1111]. Releasing [1245]. Relevance [915]. Relevant [1126, 1290, 1107, 409, 1516, 1396]. Reliable [188, 1551]. Relief [541]. Remarkable [240]. Removal [1811]. reniformis [1046, 1258]. Reoxygenation [1712]. Repair [544, 760, 174]. Repeat [1678, 1300, 249, 846]. Repeated [1085, 103]. Repeats [578, 81, 1209, 937, 1389]. Repellents [46]. Repertoire [1724, 874]. Repetitive [1799, 18, 846, 1948]. Replace [373]. Replacement [1220, 631]. Replication [1671, 247, 162, 231, 1651, 1794, 800]. Replication-Defective [162, 231]. Report [1722, 1007, 1778]. Reporter [32, 752, 1122, 164, 60, 110, 264]. Reports [181]. Represent [1333, 1307, 1313]. Represses [121]. Repressible [1435]. Reproducing [1263]. Reproduction [897, 606, 778, 1683, 1682, 1942]. Reproduction-Related [897, 606, 778]. Reproductive [1851, 599, 1048, 1880]. Required [969, 742]. Requirement [838]. Requisite [16]. rerio [275, 938, 36, 152, 247, 972, 1019, 1403, 1588, 1819, 110, 1957, 1078]. Rescued [215]. Research [272, 465, 874, 1008, 1787]. Resequencing [1561, 1761]. Reshuffling [663]. Resident [1212]. Residue [1882]. Resilience [1809, 1370]. Resilient [1716]. Resistance [537, 1807, 1870, 1182, 286, 1038, 717, 1191, 1941, 1504, 814, 927, 1956, 1303, 1456, 918, 1179, 283, 1407, 1201, 1375, 1399, 1510, 1696, 1846, 1494, 1524, 592, 1564, 1648]. Resistance/Susceptibility [592]. Resistant [1691, 341, 722, 41, 866, 1801, 590, 1821, 1326, 1501, 757]. Resolution [1384, 1233, 1300, 180]. Resonance [30]. Resorption [1763]. Resource [184, 950, 1672, 1667, 1451, 974]. Resources [583, 1431, 1034, 724].

Respiration [817]. Response [1873, 1910, 807, 1919, 1768, 799, 1596, 1085, 1560, 1773, 1038, 1860, 1252, 1806, 1868, 965, 806, 1373, 256, 1075, 1030, 1173, 1402, 1361, 1571, 1311, 1251, 1042, 1266, 1459, 116, 853, 1890, 1144, 1664, 1906, 1729, 1781, 1471, 1942, 1632, 1462, 890, 580, 1783, 1939, 899, 1712, 1946, 961, 1237, 1574]. Responses [966, 1831, 1594, 1681, 963, 1883, 1852, 1838, 1615, 1814, 1824, 1441, 1685, 1840, 1185, 1881, 1356, 1523]. Responsible [780, 734, 1554, 1015]. Responsive [1961, 1213, 1885, 910, 1204]. Resting [723]. Restored [1585]. Restricted [833]. Restriction [196, 1237, 305]. Result [919]. Resulting [1805, 44]. Results [1732, 1789]. Retaining [826, 1425]. reticulata [382, 1676, 1083, 290, 353]. Reticulum [1425]. Retinal [1055]. Retinoblastoma [157]. RETRACTED [1255]. Retrieval [1281]. Retrieved [1320]. Retro [1191]. Retroviral [169, 164, 162, 231]. Retroviruses [183]. Reveal [1807, 1768, 1883, 1926, 1275, 1685, 896, 1647, 1710, 714, 1874, 1849, 1895]. Revealed [1228, 419, 421, 1676, 1922, 34, 412, 1832, 540, 153, 193, 226, 1885, 1093, 1634, 1815, 1076, 1490, 1781, 1527, 1696, 1500, 1690]. Revealing [1371, 1324]. Reveals [1918, 1798, 1822, 1925, 1938, 1788, 1715, 514, 1404, 1889, 228, 1852, 1363, 1750, 1821, 1571, 1628, 1224, 1041, 1863, 1611, 1963, 1262, 1130, 1185, 1659, 1428, 1906, 1708, 1780, 1607, 1717, 1791, 1426, 1521, 1921, 1706, 1712, 1891, 1573, 1908, 1157, 1342, 1414, 1648, 1387, 1113]. Revelation [15]. Reversal [1834, 1428, 1589]. Reversal/Differentiation [1428]. Reverse [969, 135]. Reversed [1862]. Review [1625, 822, 575, 358, 964, 398, 485, 399, 671, 708, 1835]. Reviewers [1707]. Revision [319]. RFLP [516, 226, 208, 585, 518]. rfp [264]. Rhabdovirus [63, 1886]. Rhamnolipid [1635]. Rhizochalin [957]. Rhodobacter [445]. Rhodococcus [1477, 1198]. Rhodomonas [1562]. Rhodophyta [1374, 1406, 806, 1373, 234, 1311, 1015, 342, 1459, 836]. Rhodophyte [1468]. Rhodophyte-Associated [1468]. Rhodopsin [277]. Rhogocyte [1236]. Rhopaloeides [1264, 933]. Rhythmic [1473]. Ribbon [1876]. Ribonucleoprotein [1613, 1507]. Ribonucleotide [196]. Ribosomal [366, 141, 91, 250, 148, 783, 219, 126, 364, 460, 107, 462, 448, 351, 381, 745, 229, 334, 25, 246, 855, 1765, 136, 837, 305]. Ribosome [1908]. Ribosylation [401]. Ribozyme [120]. Ribozymes [120]. Rich [1415, 1468, 955, 1352, 1141, 1923, 237, 994, 215, 405, 1157, 1033]. Richardson [1119]. Riftia [1383]. Right [922, 1443]. rigida [1923]. RING [1023]. ringgoldianum [572]. Risks [1111]. River [1773, 1806, 1458]. Rivers [144]. rivoliana [1688]. Rivulidae [93]. Rivulus [93]. RNA [123, 1831, 1727, 1851, 120, 141, 91, 1658, 327, 1560, 147, 1773, 1304, 1526, 126, 107, 965, 1057, 1721, 351, 190, 1361, 57, 1611, 1610, 1088, 1077, 1629, 930, 1401, 700, 728, 695, 1708, 201, 921, 1253, 246, 1765, 1696, 1782]. RNA-Depletion [1765]. RNA-Encoding [91]. RNA-Seq [1658, 1721, 1361, 1765, 1560, 1773, 1401, 1708, 1696, 1782]. RNA-Sequencing [1831]. RNAi [1833, 1782]. RNAs [1560, 1832, 1660, 1347, 1338, 1378, 1767, 1652, 1318, 1062, 1783, 1412]. RNAseq [1881]. robusta [1924]. Rock [1720, 1834, 1112]. Rockfish [151, 481, 39]. rogercresseyi [1820, 1851, 1918, 1290]. rohita [1929, 1885, 1076]. Rohu [1885, 1076]. Role [1312, 1820, 1768, 1832, 1678, 1871, 1030, 1902, 844, 1000, 1171, 1023, 1509, 1644, 1867, 1808, 1908]. Roles [993, 1686, 1352, 765, 1490, 1651, 1706, 1782]. Roller [289]. roretzi [719]. rosenbergii [606, 395, 1806, 1838, 1170, 778, 1499, 493, 1589]. Roseobacter [8]. Rosettes [1283]. rostrata [1922]. Rotifer [1725]. Rotifera [930]. Rotifers [80, 1562, 815, 1141, 585, 758]. rotundiformis [80]. royi [993]. RPA [1795, 1795]. RPA-EXO [1795]. RPA-LFD [1795]. rps6kb2 [1398]. rrm1 [1696]. rRNA [496, 440, 77, 292, 370, 1008, 732]. rRNA-Based [1008]. RSIVD [1407]. RT [922]. RT-qPCR [922]. rtn4al [1485]. rubens [1295, 952]. Rubisco [1188, 582, 667, 351]. rubra [524, 587]. rubripes [1826, 176, 1405, 194, 1602, 1921, 1642]. rubrum [441]. ruckeri [910, 1069]. Ruditapes [1961, 1706, 1932]. Ruegeria [1673, 396, 1480]. rufescens [35, 1632]. Russia [621]. Russian [1426]. rv [919, 445, 895]. rv-cyclin [919, 895]. Ryukyu [1937].

S [962, 11, 649]. S-Transferases [649]. S. [1855]. S1 [1028]. S1-Type [1028]. S27E [25]. S31603 [537]. Sac [1448, 909, 951, 1502, 1086]. Saccharification [1406]. Saccharina [1167, 1764, 454, 782]. Saccharophagus [1296]. Saccostrea [193, 321, 490]. Sacs [1460]. SAF [498]. SAF-1 [498]. Safety [2]. Saithe [799]. salar [1331, 1772, 1594, 1681, 1715, 1294, 1164, 436, 927, 99, 1071, 1837, 1620, 847, 860, 1023, 714, 1573]. Salicornia [1213, 1091, 1184]. salina [737, 835, 544, 795, 695]. Saline [1938, 1411]. Saline-Alkaline [1938]. Salinities [291]. Salinity [1892, 270, 598, 1773, 1883, 1418, 1852, 1533, 1532, 1759, 1283, 201]. Salinivibrio [1607]. Salmo [1331, 1772, 1594, 1681, 37, 1715, 1294, 1164, 436, 927, 99, 1071, 1837, 1620, 847, 860, 1023, 348, 714, 1573]. salmoides [1754]. Salmon [1287, 611, 1820, 1331, 1594, 1681, 373, 1637, 1290, 1760, 1715, 1294, 1164, 436, 1367, 571, 1675, 927, 464, 1600, 1283, 1071, 1837, 819, 432, 619, 407, 1620, 1622, 33, 116, 1679, 847, 860, 350, 1023, 380, 409, 1652, 714, 1747, 1573, 751, 504]. salmonicida [1715, 1012, 1185, 380, 1795]. Salmonid [68, 1613]. salmonis [1675, 751]. Salt [737, 501, 1091, 1503, 1371, 1054, 729, 1501, 1129, 1184]. Salt-Regulated [501]. Salt-Resistant [1501]. Salt-Tolerant [729]. Salvaged [723]. Salvelinus [352, 199, 9]. Same [1460]. samea [1326, 215, 405]. Samples [195, 157, 1483, 456, 363]. Sampling [876, 446]. San [178]. Sand [944]. sapidus [196, 7]. sarba [327, 598, 285]. Sarcoma [919, 895]. Sarcomere [1424]. sarcophytoides [1723]. Sarcothalia [1469]. Sardina [266]. Sargassum [987, 50, 572, 839, 1524]. Sasaki [1884]. Satellite [80, 233]. sativa [1331]. Saturated [44]. saturnus [879, 1044, 924]. saxatilis [124, 1041]. SbASR [1091]. SbASR-1 [1091]. SbpAPX [1184]. scabra [1866]. Scaffolds [1788, 1823]. Scale [1192, 1111, 1115, 1119, 1771, 1195, 1338, 1624, 847, 1803, 1786, 775, 1458, 1627, 1626]. Scales [200]. Scallop [1064, 1020, 1335, 1498, 1739, 633, 1540, 1269, 455, 1092, 687, 608, 8, 220, 221, 530, 1151, 1428, 548, 1474, 131, 1544, 775, 655, 710, 688, 1379, 1535, 1577, 305]. Scallops [1857, 1861, 1874]. Scan [1201, 1513]. Scanning [400, 261]. Scapharca [621]. Scaphechinus [944]. Scattering [257]. Scavengers [572]. Scavenging [1421]. Scheme [171]. Schiller [1066]. Schisandra [1895]. Schizochytrium [1190, 1476, 1726, 562]. Schizophrenia [1464]. Schizothorax [1429]. schlegeli [276]. schlosseri [330]. Schneider [1953]. schrammi [652]. Schrenck [621]. schröederi [1912]. Sciaenops [316, 1081]. Science [683]. scintillans [686, 1618]. Sclareol [1240]. Sclareolide [1240]. Scleractina [768]. Scleractinia [731, 381, 13]. Scleractinian [91, 731, 891, 459]. Sclerites [573]. Sclerosis [1485]. Sclerotioramine [1386]. Scomber [1855]. Scophthalmus [384, 1068, 1528, 81, 877, 1012, 1072, 1179, 1079, 1917, 593, 1539, 1400]. Scopularide [1336]. Scopulariopsis [1259]. Scorable [88]. Scraper [1868]. Screen [1877, 796, 1224]. Screened [363]. Screening [46, 269, 354, 711, 1413, 1299, 1097, 267, 1395, 585, 564, 192, 1778, 1810, 851, 1567, 788, 789, 1497]. scRNA [1841]. scRNA-seq [1841]. Scyliorhinus [19, 494]. Scylla [1925, 96, 1608, 1509, 1717, 1791, 1412, 1601]. Scyphozoa [864]. sdY [1837]. Sea [430, 197, 1820, 303, 1099, 392, 520, 539, 1851, 328, 1192, 1918, 610, 1193, 719, 235, 1843, 716, 323, 270, 327, 598, 1866, 440, 582, 111, 1118, 100, 1830, 1111, 1337, 1174, 1314, 1713, 952, 531, 902, 766, 1889, 808, 457, 427, 641, 476, 599, 1328, 462, 721, 97, 486, 1246, 253, 1686, 1353, 1383, 1595, 285, 230, 129, 102, 57, 1120, 1338, 733, 485, 857, 95, 1685, 1665, 669, 904, 588, 229, 868, 299, 1495, 1512, 1933, 906, 824, 1077, 1407, 1447, 345, 957, 1327, 1462, 883, 55, 1426, 1710, 348, 1062, 1388, 1915, 1034, 625, 306, 1362]. Sea [738, 1414, 1396, 1574, 723, 371, 945, 770, 1198, 187, 1047, 421, 1449, 973, 661, 767, 1084, 107, 902, 129, 619, 992, 1229, 782, 265, 788, 789]. Seabass [1931, 1864, 538, 1303, 1461, 1375, 1399, 1537, 1662, 1155, 1794, 145]. Seabream [443, 1919, 1075, 1814, 24, 203, 1302, 1859, 1903, 1059]. Seafood [1804, 358]. Seagrass [704]. Seahorse [1963]. Seascapes [1903]. Seasonal [1793, 394, 727, 1225, 774, 896]. Seawater [537, 1676, 667, 337, 97, 1673, 116, 732]. Seaweed [1568, 1055, 1366, 293, 1451, 689, 1350, 1633]. Seaweeds [1415, 1835, 853, 1048, 336]. Sebastes [151, 481, 39]. Second [1052]. Secondary [1801, 618, 803, 1336, 1134, 1138, 1395, 904, 623, 1774]. Secretagogues [64]. Secreted [952]. Secretion [720, 1000, 1638]. Secretions [261]. securis [665]. Sediment [371, 1722, 902, 766, 784, 641, 1547, 1639, 1054, 1495, 331]. Sediments [1585, 1175, 138, 1351, 1045, 52, 1369, 70]. Seed [75]. Seep [371, 385, 52]. Segment [189, 221]. Segregant [1721, 1201, 1586, 1668]. Segregation [526, 577]. Sei [540]. Selected [866, 550, 1162, 1171, 1959, 447]. Selection [1346, 1687, 1955, 1456, 1550, 1144, 616, 1177, 1474, 560]. Selective [1596, 335, 801, 1093, 1501]. Selectively [821]. Selectivity [835]. Selenium [1855, 1859, 1139]. Selenium-Containing [1139]. Selenka [102]. Selenoneine [1882, 1859, 1695, 1139]. Self [1799, 1671, 93, 948]. Self-assembling [1799]. Self-fertilizing [93]. Self-Inhibitory [1671]. Semi [1009, 620]. Semi-Amylopectin [620]. Semi-continuous [1009]. semilaevis [626, 777, 1031, 1252, 825, 1628, 1398, 887, 776, 1298, 1508, 1689, 1733]. Seminal [1508]. Semiquantitative [400]. Semliki [189, 79]. Senegal [802, 810]. senegalensis [1104, 1085, 745, 1222, 802, 810, 1218]. Senegalese [1085, 745, 1222, 1202, 1218]. Sensing [1354, 1224, 1593, 677, 974]. Sensitive [1083, 1308, 1435, 1593, 38, 1293]. Sensitivities [1661]. Sensitivity [1704, 215, 817]. Sensory [1737]. Separation [864]. Sepia [1753]. Septicaemia [1068, 1179]. Septicemia [969, 1671, 1088, 1785, 1510]. Septifer [1592]. SEQ [1077, 1658, 1418, 1721, 1361, 1765, 1551, 1560, 1773, 1503, 1841, 1566, 1076, 1401, 1708, 1696, 1782]. Sequence [48, 675, 807, 56, 1358, 1691, 141, 719, 85, 61, 126, 364, 406, 292, 1363, 256, 1799, 785, 1209, 462, 1300, 916, 134, 584, 381, 18, 93, 62, 190, 508, 538, 203, 294, 745, 524, 587, 529, 748, 260, 853, 252, 704, 221, 426, 548, 83, 266, 624, 746, 960, 417, 233, 931, 459, 405, 1655, 136, 800, 1396, 494]. Sequences [63, 94, 463, 80, 250, 421, 621, 315, 661, 103, 451, 13, 102, 493, 1076, 334, 988, 776, 246, 751, 655, 1948, 496]. Sequencing [1831, 1727, 1228, 196, 861, 34, 1619, 1672, 1160, 579, 448, 1885, 1335, 45, 202, 1611, 1531, 883, 1253, 1204, 1399, 1586, 1668, 431, 1655, 1642, 1847, 1787, 1538, 1647, 1874]. Sequester [313]. SERCA [859]. seriata [1051]. Seriatopora [1277]. Serine [461, 1901]. Seriola [1688, 1278, 1695]. Serpin [12]. Serranidae [412, 333]. Serratia [949]. serratus [853]. Serum [6, 12, 337]. Serves [1356]. Services [184]. Sesquiterpene [936, 734]. Set [1753, 745, 1269, 1157, 1400, 1597, 1640]. setiferus [280]. Seto [107]. Settlement [101, 489, 14, 977, 804, 654]. Seven [412, 1550]. Several [1898]. Severe [1789]. Sex [897, 1918, 1290, 626, 777, 1031, 1238, 1913, 1931, 433, 1004, 1555, 1753, 1720, 1754, 1864, 276, 834, 1267, 959, 378, 1335, 1590, 1636, 1119, 1834, 774, 1784, 1952, 1862, 407, 1262, 741, 1074, 1244, 887, 1428, 1589, 1581, 1079, 1509, 1586, 1668, 1710, 1848, 1176, 1377, 1427, 1733, 1556]. Sex-Associated [1862]. Sex-Biased [1918, 1733]. Sex-Determination [887]. Sex-Determining [834, 1267, 1176, 1377]. Sex-Linked [1913, 433, 1267, 959, 1586, 1668]. Sex-Related [1290, 1555, 1754, 1335, 1710]. Sex-Reversed [1862]. Sex-Specific [1031, 1720, 1262, 1079, 1586, 1668, 774]. Sex-Undifferentiated [1581]. Sexes [605]. Sexing [1511]. Sexual [1697, 1164, 1628, 1512, 1644]. Sexually [1637, 1275, 1263]. SFGM [326]. Sfp1 [1686]. SGIV [1651]. SGNH [1933]. shad [1815]. Shallow [1320, 1353]. Shallow-Water [1353]. Shanny [104]. Shape [1798, 317, 1605, 1538]. Shape-Related [1605]. Shaped [1903]. Shared [434]. Sharing [1126]. Shark [19, 494, 1257]. Sharks [308, 204]. Shedding [607]. Sheds [1120, 1362]. Shell [1061, 1448, 1444, 621, 1526, 126, 1445, 258, 1279, 720, 909, 1643, 1502, 1901, 444, 844, 1660, 1352, 1000, 1737, 1017, 997, 608, 220, 1572, 1466, 536, 1780, 1391, 1706, 1932, 1842, 1254, 1749, 1379, 1782, 1690]. Shelled [1900]. Shellfish [313, 75]. Shells [1554, 589]. Shewanella [1329, 1898]. Ship [1351]. Shock [235, 270, 1860, 1105, 1030, 218, 821, 116, 1051, 67, 1223, 970, 996, 581, 1735, 1197, 561, 1400]. Short [1288, 700, 1473]. Short-term [1473]. Short-Time [1288]. Shortage [1287]. Shotgun [953, 1963]. Showing [221, 1549]. Shrimp [1161, 578, 908, 1126, 966, 280, 1481, 1291, 470, 404, 222, 750, 963, 34, 965, 1553, 1871, 461, 1941, 17, 226, 1841, 486, 1103, 1173, 1796, 1364, 62, 964, 378, 519, 644, 657, 1171, 1266, 356, 748, 115, 712, 1479, 515, 521, 1450, 1674, 932, 693, 636, 1087, 753, 1705, 1325, 994, 117, 880, 967, 1607, 1942, 634, 437, 243, 1217, 1586, 1668, 681, 855, 133, 1377, 1513, 1524, 848, 1905, 961, 1025, 304]. Shrimps [1619, 1360]. shRNA [1301]. SHRV [1886]. Shuttling [780]. Siganus [1482, 1522, 1591]. Sigmops [723]. sign [1274, 1493, 1492, 1632]. Signaling [1860, 947, 273, 733, 1958, 1683, 1682, 343, 1510, 1818, 1521, 1559, 1604]. Signals [1425]. Signature [1675, 1237]. Signatures [1687, 1780]. Significance [1178]. Significant [1393]. Silencing [1851, 1304, 693, 1880, 1589, 1894, 1794]. Silica [1597, 1640, 1616]. Silicatein [873, 449, 982]. Silicates [1046]. Silico [1875, 1470, 1583, 1464, 478]. siliquosum [1524]. Silk [1954]. Silk-Like [1954]. Sillago [673, 1787]. Silver [327, 598, 1887, 1872, 1146, 285, 1452, 1556]. Silver-Lipped [1146]. sima [1822]. Similar [768, 1485, 204]. Simple [1210, 944, 1209, 1300, 188]. Simplex [741]. Simplified [1811]. simula [1876]. simulans [809, 793]. Simultaneous [335, 204]. Simultaneously [826, 1524]. sinensis [785, 1249, 1818, 1900]. SINEs [123]. Singapore [1413, 290]. Single [454, 85, 1279, 1830, 684, 666, 448, 1344, 538, 1251, 352, 1785, 1957, 1850, 1561, 781, 1787]. Single-Cell [448, 781]. Single-Chain [1785]. Single-Locus [1279]. Single-molecule [1787]. Single-Nucleotide [538, 1251, 1957]. Single-Step [1850]. Single-Stranded [352]. sinica [1787, 718]. Siniperca [920, 1237, 1879]. Sinonovacula [1892, 861, 1309, 1397]. Sinularia [618, 803, 1011]. Sinus [85, 634]. Siphonodictyal [859]. siRNA [1589, 1894, 1943, 1124, 1025]. siRNA-Mediated [1589]. Site [987, 780, 1454, 631, 1289, 640, 622, 1395, 560, 1367, 723]. Site-Directed [987, 631, 1289, 640, 560]. Sites [640, 504]. Situ [578, 197, 261, 165, 84, 479, 933, 553, 568, 351, 233]. Six [421, 335, 1634, 899, 496]. Size [205, 823, 1644, 1927]. Sizes [316]. SJ4 [1879]. Skate [325]. Skeena [144]. Skeletal [197, 1192, 662, 1075, 1302, 1958, 1559, 1604, 1237, 723]. Skeleton [1506, 1748]. Skeletonema [790, 27]. Skin [1625, 1187, 1145, 1752, 1071, 760, 1920, 1842]. Skirt [726]. skottsbergii [1469]. slc45a2 [1825]. SLDP [1205]. Sleeping [971, 205]. Slow [1424]. Slug [427]. Small [91, 563, 126, 1299, 218, 483, 1347, 1278, 1611, 1062, 961]. Small-Bodied [1278]. Smelt [495, 896, 724]. Smolt [1675, 1679]. Smoltification [1212, 847, 1652]. Smooth [626, 777, 1031, 825, 1628, 887, 776]. SMRT [1787]. Snail [1667, 1204]. Snake [1350]. Snakehead [1886, 1511]. Snapper [406, 212, 1909]. Snappers [661]. Snout [1567]. SNP [807, 1238, 1816, 1146, 1862, 1447, 1440, 1253, 1848, 1561]. SNPs [1798, 1346, 1267, 1565]. snRNP [99]. snRNP-Specific [99]. sockeye [504]. Sodium [1663]. Soft [532, 1723, 364, 618, 803, 1900]. Soft-Shelled [1900]. Softness [719]. Soil [1722]. Solar [1638]. solaris [1771, 1824, 1856, 1930]. Soldier [332]. Sole [626, 777, 1031, 1085, 451, 81, 825, 1628, 1398, 745, 1180, 1222, 1202, 802, 810, 887, 1218, 776]. Solea [1104, 1085, 81, 745, 1180, 1222, 802, 810, 1218]. Soleidae [451, 108]. Soles [108]. Solexa [1160]. solidissima [1910]. Solitary [153]. Soluble [642, 658, 766, 1120, 1046, 1390]. Solution [684, 607]. Solvent [1384, 545, 331]. Solvent-Tolerant [1384, 545, 331]. Somatic [662, 1754, 110]. Somatolactin [430, 328, 201]. Some [259, 439, 106, 995]. Sonic [343]. Sonication [24]. Sorbicillinoid [1449]. Sorbicillinoid-Producing [1449]. Sound [481]. Source [1746, 1691, 269, 308, 1233, 1134, 1670, 1650, 1699, 1496]. Sources [1625, 1817, 632, 1623, 1110]. South [902, 421, 973, 661, 767, 1084, 555, 1229, 1903, 788, 789]. Southeast [20]. Southern [1729, 485, 236, 1845, 348]. Southwest [195]. Sox [140]. SOX5 [1728]. Sox9b [630]. Soy [860]. Soybean [1957]. sp [1419, 845, 1576, 1060, 1691, 1190, 1321, 1477, 1701, 1354, 545, 973, 842, 1801, 767, 905, 645, 1339, 1265, 1654, 1827, 579, 425, 754, 1289, 385, 1198, 486, 1173, 1206, 595, 574, 730, 1476, 1630, 986, 1817, 840, 827, 555, 1854, 935, 1188, 1569, 1480, 1096, 1824, 1441, 1050, 1132, 1896, 1037, 1495, 1711, 1933, 1259, 1664, 1390, 983, 956, 1199, 989, 791, 1021, 1778, 1230, 1349, 1766, 612, 738, 679, 943, 1396, 1321, 1782, 1475]. SP-32 [1321]. sp. [1329]. Spacer [56, 783, 219, 442, 500, 701, 136, 305]. Spacers [250]. Spain [665, 126, 451]. Sparidae [605, 1759]. Sparus [430, 1192, 1919, 327, 598, 1118, 1075, 24, 285, 203, 1302, 95, 1024, 1102, 1462, 1578, 1059]. Spatial [419, 316, 881]. Spatiotemporal [559]. Spatoglossum [1912]. Spawn [1951, 1452]. Spawn-Inducing [1951, 1452]. Spawned [736]. Spawning [1181, 921, 1567, 504]. Spdsx [1791]. Special [485, 962, 1853]. Specialization [1958]. Speciation [1315]. Species [1368, 280, 259, 569, 402, 51, 556, 647, 563, 207, 452, 364, 258, 412, 263, 510, 1115, 1802, 1777, 451, 953, 1280, 666, 884, 916, 1049, 132, 96, 346, 1121, 381, 403, 18, 1134, 23, 453, 300, 319, 1858, 13, 260, 1133, 1403, 515, 1450, 573, 1315, 1909, 8, 1013, 298, 1178, 223, 1401, 1106, 4, 374, 446, 937, 265, 82, 1915, 1494, 1496, 1574, 496, 1672, 305, 331]. Species-Diagnostic [96]. Species-Specific [569, 510, 132, 300, 1915]. Specific [1287, 569, 626, 1031, 1118, 119, 77, 273, 1720, 1333, 1307, 1313, 510, 387, 1057, 667, 955, 1673, 132, 604, 513, 62, 1590, 825, 300, 99, 1784, 1597, 1640, 1425, 116, 1262, 456, 1758, 1531, 1785, 1178, 690, 1023, 310, 836, 1272, 1079, 1586, 1668, 1112, 507, 1915, 779, 1941, 774, 1952, 1266, 1736]. Specificities [1040]. Specificity [1104, 1933]. Specifies [184]. Specimens [1495]. Speckled [1569]. Spectra [1925]. Spectral [420, 707]. Spectrometry [85, 1788, 842, 733, 367, 1319, 1916, 864]. Spectrophotometric [46]. Spectroscopy [1367]. Spectrum [1488]. Speed [1927]. Spengler [1191]. Sperm [1557, 152, 24, 1278, 733, 468, 736, 1452, 1003]. Sperm-derived [1557]. Spermatogenesis [1812, 1064]. Spermatogonia [1718]. Spermatozoa [957]. Spfoxl [1717]. Spfoxl-2 [1717]. Sphaerococcus [977]. sphaeroides [445]. Sphingan [1630, 1817]. Sphingolipid [957]. Sphingomonas [1289, 1630, 1817, 1037]. Spicule [873]. Spicule-Forming [873]. spine [1943]. Spined [1041]. Spines [944]. Spiny [1260, 415, 556, 676, 506]. Spiroplasma [1907]. Spirulina [1638, 1803, 1786, 425, 28, 925, 213, 507, 488]. Spisula [1910]. Spleen [1826, 1715, 309, 1012, 1072, 1429, 1846]. splendidus [1442]. Splenocytes [702]. Splice [841]. Splicing [1820, 395, 1832, 1348, 1956, 1758, 1906, 1708, 1471]. Sponge [1746, 1098, 859, 1270, 1545, 1089, 51, 1893, 882, 1321, 1701, 1264, 210, 1491, 377, 603, 678, 1486, 100, 767, 1084, 1215, 489, 568, 682, 831, 793, 727, 867, 403, 1123, 1616, 872, 670, 241, 865, 367, 449, 1046, 1258, 1140, 298, 369, 989, 650, 1208, 1007, 484, 1082, 246, 933, 26, 1765, 820, 654]. Sponge-Associated [1746, 1098, 1270, 1545, 51, 1491, 403, 1208, 484, 872]. Sponge-Cell [369]. Sponges [1018, 979, 973, 822, 1421, 1225, 1281, 1353, 739, 398, 71, 1125, 857, 66, 1483, 868, 1920, 1148, 473, 1823, 990, 982, 1394, 1455, 1127, 788, 789]. Spongin [1483]. Spontaneous [185]. Sporophyte [1109]. Sporophyte-Preferential [1109]. Sporophytes [1239]. Spot [196, 993, 1165, 698, 693, 580, 1524, 1025]. Spotted [533, 1734, 1855, 604, 1590, 1595, 1784, 1527, 1833, 635]. spp [1742, 528, 1089, 936, 711, 1191, 1902, 1514, 792, 1650, 505]. Spring [1886]. SPS [1284]. SpyCatcher [1792]. SpyTag [1792]. SpyTag/SpyCatcher [1792]. Squalene [227, 1766]. Squid [949, 1730, 137, 54, 1618]. Squirt [719]. SR21 [562]. srebp1 [1522]. srsf3 [1952]. SS1 [555]. SSCP [585]. SSR [1031, 1238, 1619, 871, 1093, 515, 648]. SSR/SNP [1238]. SSRs [1081]. Stability [631, 840, 1905, 498]. Stable [1557, 254, 1225, 105, 1933, 1177, 1828]. Stage [197, 1118, 387, 604, 814, 1813, 1059]. Stage-Specific [604]. Stages [1812, 1764, 1026, 251, 1610, 1218]. Staining [257]. Stainless [537]. Stalked [457, 1484]. Standard [528]. Standards [420]. Staphylococcus [1449, 757, 590]. Star [952, 1686, 87, 1915]. Starch [1297, 1531]. Stars [733]. Starvation [1822, 1688, 1297, 1838, 1814, 1824, 1559, 1065]. Stat2 [1380]. State [737, 650]. States [20]. Statistical [1372, 1406]. Status [1126, 856, 1866, 603, 1009, 1361, 588, 1857]. Steady [737]. Stearoyl [1345]. Steel [537]. Steelhead [144, 1212]. Stem [610, 669, 1497]. Stemness [939]. stenolepis [1004]. Stenotrophomonas [1505]. Step [1850]. Stereoselective [1233, 1053, 1285]. Sterile [1048]. Sterility [1472, 1861, 1874]. Sterilization [1435]. Sterilize [1316]. Sterol [227].

Stevensine [377]. Stichopus [476, 599]. Stickleback [1041]. Stimulate [947]. Stimulated [1380, 541]. Stimulates [786, 471, 1524]. Stimulating [245]. Stimulation [1637, 750, 687]. Stimulatory [1337]. Stimuli [1252]. STKMTI.2 [1722]. Stochastic [1905]. Stock [144, 495, 20, 1866, 434, 619]. Stocks [419, 609, 184, 266, 518]. Stomatopoda [521]. Stomiiformes [57]. Stop [565]. Storage [90, 1123, 620, 1261]. Stories [170]. Storm [1638]. Storms [976]. Strain [969, 1487, 1419, 1637, 1264, 1449, 1801, 550, 1827, 425, 453, 1171, 1122, 396, 1096, 1896, 616, 293, 362, 423, 1501, 1489, 791, 1607, 1021, 1230, 747, 1391, 389, 1169, 492]. Strains [427, 1687, 290, 1480, 1130, 1185, 1505, 729, 995, 883, 1008, 1810, 851, 447, 445]. Strand [352]. Strand-Mediated [352]. Stranded [352, 921]. Strategies [141, 814, 87, 1130, 399, 1828]. Strategy [559, 1061, 1080, 1811, 1403, 1385]. Streblonema [1390]. Strength [402, 1495]. Strengthening [602]. Streptococcus [1792]. Streptomyces [1491, 258, 1801, 767, 754, 827]. Stress [1910, 1768, 737, 1085, 1892, 1560, 139, 1773, 1961, 1408, 1213, 1864, 1872, 1348, 1852, 1373, 1750, 256, 557, 339, 1055, 1885, 1571, 1814, 1566, 1759, 1815, 1453, 1459, 1371, 1051, 1840, 1130, 704, 1144, 940, 786, 1683, 1682, 1906, 1184, 690, 956, 1729, 1767, 1708, 773, 1781, 1462, 1422, 1272, 1204, 1559, 1604, 862, 1847, 1776, 1849, 1310, 505, 983]. Stress-Induced [956, 983]. Stress-Responsive [1885]. Stresses [1938, 1310]. Stressful [1811]. Stressors [222, 853]. striatum [1489]. Strikes [1257]. Striostrea [193, 490]. Striped [124, 1041, 208, 310]. Stromatolites [261]. Strombus [1172]. Strongylocentrotus [1099, 957, 1062]. Structural [527, 1227, 1723, 531, 1469, 1687, 319, 697, 1006, 1284, 623, 92]. Structurally [772]. Structure [430, 675, 144, 495, 419, 421, 621, 1255, 1286, 406, 434, 1084, 1868, 118, 1700, 778, 99, 203, 1744, 1569, 455, 918, 1116, 1755, 194, 220, 1845, 360, 117, 1245, 450, 1440, 1942, 437, 623, 1903, 1386, 775]. Structure-Activity [1245]. Structures [1103, 379, 988, 732]. Structuring [605, 506]. Stryphnus [1215]. Studies [1625, 335, 1491, 769, 1465, 571, 964, 30, 1470, 172, 449, 573, 940, 151, 244, 1464, 1204]. Study [533, 1061, 1798, 999, 764, 922, 1526, 1404, 1548, 104, 1170, 1605, 1636, 1603, 1744, 743, 1441, 1074, 122, 362, 530, 847, 1048, 1869, 1253, 1375, 1867, 422, 488, 1379, 1387, 1458, 1538, 1078]. Studying [1194, 166]. Sturgeon [1718]. stutzeri [1811]. stylirostris [292, 515]. Stylophora [1001]. Stypopodium [689]. Subcalicoblastic [1748]. subcordiformis [1917]. Suberites [396, 311, 26]. suberitoides [1893]. Subfamily [1043]. Subjected [284, 704]. Submerged [672]. Submersed [383]. Submitted [558]. subsalsum [355]. Subsequent [864]. subsp [1292, 1491, 1715, 1437]. Subspecies [113]. Substance [827, 342]. Substances [1701, 14, 114, 298]. Substituted [297]. Substrate [949, 1933, 1040]. Substrates [1946]. subtilis [764]. Subtle [551, 1520]. Subtraction [255]. Subtractive [428, 409, 1036]. Subtropical [976, 1673]. Subunit [196, 141, 91, 126, 1832, 954, 334, 215, 496]. Subunits [291, 442, 500]. Success [682]. Successful [1765]. Succulent [1213]. Sucrose [1817]. suecica [207]. Sugar [1764]. Suggestive [717]. Suggests [266]. Suitable [59]. Suites [1573]. Sulfatase [1116]. Sulfate [1265, 1054]. Sulfate-Reducing [1054]. Sulfated [1568, 1063, 534, 1530, 1763, 319, 326, 362, 1245, 389, 1350, 1633, 1912, 492]. Sulfation [15]. Sulfide [914, 1042]. Sulfite [371, 641]. Sulfoxide [24]. Sulphated [760, 1284]. Suminoe [518]. Summer [866, 1038, 1370, 1928]. Sun [796]. Sundarbans [729]. Super [1031, 1119]. Super-female [1031]. Super-Male [1119]. superba [413, 600]. Supercritical [1529]. Superior [352]. Superoxide [339, 631]. Supplementation [1844, 1202, 1859, 1218]. Supplemented [1163, 553]. Supply [822]. Support [751]. Supports [599]. Suppress [368]. Suppression [712, 428, 409, 1036, 714]. Suppressive [255, 895]. Suppressor [275, 388]. Surface [1448, 799, 876, 1173, 513, 818, 761, 229, 257, 890, 747, 289, 1792, 998]. Surfaces [119]. Surfactin [268]. Surfactin-like [268]. Surfclam [1910]. surmuletus [208]. Surrogate [1278]. Survey [794]. Surveys [446]. Survival [284, 486, 8, 1087, 1186, 1069, 820, 654]. Susceptibility [185, 592, 635]. Susceptible [866, 1716]. Sustainable [893, 1657, 1529, 398, 71, 355]. SVCV [1886]. Swabs [876]. Swamp [1712, 1947]. Swiftia [652]. Swim [1929]. Swimming [1211, 1927]. SY9 [555]. sydneiensis [1326, 215, 405]. sydowii [935, 886, 1169]. Symbiodinium [1210, 427, 557, 1188, 1205, 1172, 940]. Symbiodinum [401]. Symbiont [582, 1916]. Symbionts [1320, 1084, 1172]. Symbiosome [832]. Symbiotic [709, 532, 16, 1172, 334, 1199, 650, 907]. Sympatric [1421, 772]. Symposium [1247, 1007, 1067]. Syndrome [196, 993, 719, 1165, 698, 693, 580, 1524, 1025]. Syndrome-Associated [719]. Synechococcus [1419, 976]. Synergistic [1416]. Synteny [1041, 1052]. Synthase [987, 382, 1871, 1221, 837, 759, 318]. Synthesis [1576, 1444, 618, 1934, 1926, 541, 1476, 1844, 819, 1665, 1207, 122, 311, 1932, 133]. Synthesis-Related [1926]. Synthesized [268, 772]. Synthetase [837, 788, 789]. Synthetic [281, 670]. syringae [747]. System [969, 1243, 1374, 1424, 415, 1481, 235, 1354, 335, 139, 222, 651, 1548, 509, 1005, 1366, 1121, 187, 1442, 1194, 1854, 1122, 813, 1439, 1096, 1273, 79, 1683, 1682, 1785, 1325, 1629, 1617, 1203, 1944, 1360, 833, 710, 1948]. Systematic [1625]. Systems [49, 391, 71, 1136].

T [701]. T3SS [898]. T7 [1629]. TAACC [578]. Tachyplesin [764]. Tachyplesin-Encoding [764]. Tag [256, 134, 508, 1616, 853, 704, 746, 405]. Tags [807, 61, 785, 916, 584, 62, 426, 548, 624, 417, 931]. Tail [1931, 733, 1521, 474]. Tailed [551, 577]. Taiwanese [851]. Takifugu [1826, 1105, 1405, 1563, 223, 1602, 1921, 1642]. Talaromyces [1230]. TALEN [1519]. TALEN-Mediated [1519]. TALENs [1214]. tamarense [1877, 510]. Tandem [764, 1389]. Tandem-Arranged [764]. Tandemly [123]. Tanner [1724]. tapetis [672]. TaqMan [147]. TAR [161]. tarda [1456, 898]. Target [1368, 1016, 1961, 1566, 1783, 1948]. Targeted [340, 77, 878, 1221, 1457, 399, 1136, 1176, 1507]. Targeting [1886, 440, 1671, 1482, 1425, 1522, 1591, 1943, 1770, 1932, 1836]. Targeting/Retaining [1425]. Targets [1912]. Taste [1015]. Tau [1713]. Taxonomic [412, 682, 1621, 1963]. Taxonomically [1468]. Taxonomy [490]. Tc [9]. Tear [402]. Technique [366, 1156, 81, 1590, 1784]. Techniques [1795]. Technologies [740]. Technology [1304, 1316, 1833]. Teleost [598, 291, 1703, 1145, 1280, 93, 1482, 1610, 1332, 1178, 663, 38]. Teleostei [57]. Teleosts [771, 696, 1332, 1522]. Tellinoidea [1226]. Tellurite [1547, 1639]. Telomerase [424]. Telomeric [578]. Temperate [1369]. Temperature [1750, 227, 1480, 704, 471, 791, 1581, 624, 1808, 1939, 1310]. Temperature-Dependent [1480]. Temperature-Induced [1581]. Tempo [768]. Temporal [316, 1938, 274, 1675, 386, 881]. Temporary [952]. Ten [451]. tenera [1373, 1459]. Tensile [402]. Tentacle [1569, 1370]. Tenualosa [1815]. tenuis [1657]. tenuissima [1195]. Term [979, 1473]. Terminal [1289, 1159]. ternetzi [726]. Terpene [734]. Terpenes [977]. Terrestrial [792]. tert [673]. tertiolecta [1666]. Test [1927]. Testes [1812, 1472, 1718]. Testicular [1364, 630, 736]. Testing [814, 1915]. Testis [1812, 1734, 878, 1335, 1428, 1412]. testudienus [1890]. testudinum [760]. Tetra [726]. Tetrad [1092]. Tetrahydropterin [382]. Tetramate [1491]. Tetrameric [918]. Tetranucleotide [295]. Tetraodontidae [223]. Tetraodontiform [823]. Tetraodontiformes [223]. Tetraploid [491, 131, 581, 1389]. Tetraploidization [1400]. Tetraploids [491]. Tetraselmis [207, 1917]. Tetrodotoxin [1621, 1800, 1937, 1857, 1949]. Tetrodotoxins [1876]. Texas [1351, 178]. Tg [1250, 1293]. Thailand [533, 132, 96, 193, 226, 375, 321, 490, 633, 881, 117, 450]. Thalassia [760]. Thalassiolin [760]. Thaleichthys [495]. Thallasoma [421]. Thamnaconus [1868]. Thank [1707]. Thapsigargin [859]. Their [443, 527, 1018, 1270, 250, 1731, 1805, 434, 104, 1868, 370, 1852, 1802, 1777, 1143, 1923, 10, 1569, 629, 748, 1745, 530, 1767, 1006, 1023, 1127, 1509, 1426, 1391, 1891, 306, 1740, 1961, 1442, 745, 736]. Theonellamide [42, 314]. Theonellamides [109]. Therapeutic [845, 1886]. Therapeutics [753]. Therapy [1638]. Thermal [147, 1885, 1566, 1130, 940, 1617, 817, 1422, 1272, 1204, 1829]. Thermally [550, 1453]. Thermo [1821, 1933]. Thermo-Alkaline [1933]. Thermo-Resistant [1821]. Thermococcus [643, 574]. Thermosensitive [781]. Thermostability [826]. Thermostable [1692, 1635, 828, 1147]. Thermotolerance [1839]. thermotolerans [945, 1936]. These [1000, 787]. Thiaminase [1403]. Thiazolidinedione [1049]. Thioarsenical [1858]. Thioautotrophic [857]. thompsoni [151]. Thraustochytrid [477, 227, 188, 740, 1584]. Thraustochytrids [373, 1190, 370, 72, 1726, 917, 1129]. Thraustochytrium [1190, 905, 1132, 1489]. Three [1415, 543, 790, 1314, 523, 17, 321, 1041, 1093, 1863, 515, 729, 170, 83, 1040, 4, 374, 1394, 1455, 937, 1828]. Three-Spined [1041]. Threeline [411]. Thrombospondin [1678]. Throughout [88]. Throughput [1727, 1228, 1531, 1561]. Thunberg [390, 233]. Thunnus [148, 263, 113]. Thymine [342]. Thymus [514]. thynnus [113]. Tibial [1580]. Tide [366, 107, 508, 143, 368, 492]. Tidewater [186]. Tiger [470, 19, 1405, 461, 17, 226, 262, 1364, 62, 1674, 117, 243, 1642, 304]. Tigriopus [1345, 294]. Tilapia [274, 433, 1183, 1418, 2, 36, 1503, 959, 1355, 1402, 1393, 1514, 1411, 1752, 468, 466, 67, 846, 1581, 624, 1644, 1878, 1473, 1561, 1523]. Time [147, 1288, 510, 571, 686, 1123, 407, 446, 1927, 1400, 884, 1787]. Timing [1716]. TiO [1806]. Tissue [1287, 382, 1448, 598, 1158, 720, 36, 387, 427, 1146, 351, 604, 62, 95, 1800, 116, 919, 199, 1758, 1890, 1130, 1236, 906, 311, 360, 191, 708, 174, 1949, 1736, 1574, 1113]. Tissue-Specific [1287, 387, 604, 62, 116, 1758, 1736]. Tissues [1558, 1192, 1754, 134, 424, 877, 1048, 225, 1880, 110, 714, 1298, 1612, 779, 1931]. Titanium [1806]. TKU006 [787]. TLRs [1010]. Tn5 [425]. TNF [1638, 482]. TNF- [1638]. Tobacco [1213, 1091, 1184]. toealensis [1256]. TOF [730, 1133]. TOF/TOF [1133]. Tokyo [70]. Tol2 [248]. Tolerance [341, 1418, 938, 1213, 1553, 1373, 1091, 1503, 821, 1393, 1547, 1639, 1459, 1371, 1051, 1184, 1957, 1829, 1782, 1387]. Tolerant [1384, 1635, 828, 545, 331, 729, 1790]. Tolerate [596]. Toll [1868, 1010]. Toll-Like [1010, 1868]. Tongue [626, 777, 1031, 825, 1628, 1398, 887, 776]. tonsa [985]. Tool [709, 1831, 467, 60, 1440, 1422]. Tools [321, 73]. Topical [1656]. torazame [19, 494]. Total [1855, 983, 956]. Toxic [722, 1454, 510, 837, 1800, 725, 746, 1350]. Toxicants [341, 1515, 1897]. Toxicity [1317, 541, 1392, 1276, 1139]. Toxicology [177, 336]. Toxification [1857]. Toxin [1877, 100, 879, 1937, 42, 314, 1044]. Toxin-Bearing [1937]. Toxin-Producing [1877]. Toxins [531, 864]. Trachurus [1278, 374]. Trachypenaeus [515]. Tract [1260, 555]. traditional [177]. Trait [1164, 735, 1553, 1300, 1504, 1302, 1397, 1144, 1179, 1407, 912, 1201, 1155, 1513]. Trait-Related [1553]. Traits [1282, 985, 1830, 1039, 1242, 1605, 1211, 1636, 1550, 1595, 1397, 1563, 1466, 1474, 1537, 1598, 1542, 1391, 1219, 1538]. Transactivates [1024]. Transcribed [250, 783, 219, 136, 305]. Transcript [737]. Transcriptase [135]. Transcription [1918, 1213, 1454, 1945, 386, 127, 849, 121]. Transcriptional [1192, 1860, 1806, 1482, 1273, 1878]. Transcriptome [699, 1228, 1768, 1357, 1290, 1238, 1676, 1822, 1925, 1922, 1658, 1343, 1158, 1715, 1465, 1528, 1883, 1754, 1688, 806, 639, 618, 1525, 1552, 1115, 1619, 1889, 1852, 1373, 1821, 1146, 1779, 1738, 1885, 1149, 821, 964, 1335, 1884, 1311, 1628, 1814, 1532, 1600, 1071, 1815, 1309, 1685, 1130, 1185, 1512, 1202, 1244, 1108, 1151, 1428, 1223, 1617, 1102, 1581, 1652, 1717, 1791, 1204, 1426, 1586, 1668, 1647, 1521, 714, 926, 1712, 1573, 1412, 1908, 1655, 1429, 1446, 1846, 1874, 1254, 1427, 1714, 1157, 1612, 1790, 1787, 1946, 1086, 1895, 1113, 1533]. Transcriptome-Based [1373]. Transcriptome-Derived [1465]. Transcriptome-Guided [1465]. Transcriptomes [1812, 1569, 1662]. Transcriptomic [1873, 1910, 1727, 1938, 1596, 1951, 1026, 1341, 1703, 1926, 1075, 1667, 1364, 1275, 1634, 1269, 1840, 1114, 1527, 1632, 1719, 1618, 1059, 1362, 1549, 1776, 1849, 1535, 1523, 1577]. Transcriptomics [1729, 1567, 1649]. Transcripts [1304, 1421, 866, 550, 667, 99, 1077, 1178, 474]. Transducing [1669]. Transduction [1325]. Transfection [559, 216, 429]. Transfer [1481, 248, 285, 57, 282, 110, 123]. Transfer-RNA-Derived [123]. Transferases [649, 11]. Transform [85]. Transformant [1547, 1639]. Transformation [32, 425, 161, 1366, 187, 1854, 1096, 1856, 1930, 1896, 3, 1263, 981, 191, 1177, 1203]. Transformation-Associated [161]. Transformed [254]. Transforming [999, 1057, 331, 231]. Transgene [438, 1263]. Transgenerational [1733]. Transgenesis [349, 1207, 277, 1747, 1496]. Transgenic [1697, 1637, 1182, 68, 286, 317, 284, 88, 2, 1213, 1091, 240, 1435, 640, 285, 1250, 47, 726, 919, 875, 67, 542, 399, 162, 231, 283, 1184, 60, 264, 179, 1293, 1950, 1515, 1897, 522, 833, 895, 1606]. Transient [514, 1005, 813, 1109]. Translation [565, 1362]. Translational [745, 1133]. Translocator [225]. Translocon [898]. Transmission [993]. Transparency [1447]. Transparent [238]. Transplantation [1183, 1740, 1156]. Transplanted [1214]. Transport [1716]. Transporter [5]. Transposable [248, 9, 1891, 1389]. Transposase [425]. Transposases [971]. Transposition [205]. Transposon [205, 425]. Transposons [971, 472]. Treatment [1105, 1661, 1090, 1604, 895, 336]. Treatments [308, 24, 491]. Trench [52]. Trends [682, 503]. Trevally [1953]. Tri [1943]. Tri-spine [1943]. Trials [1463]. tribblei [1228]. Trichoderma [935]. Trichomonas [1788]. tricornutum [757, 1221, 1128, 1203]. Tridacna [427, 136]. Tridacnid [427]. Triggered [1919, 780]. Triglyceride [1096, 1110]. trihydroxydibenzo [863]. trihydroxyphenoxy [863]. trilineatum [411]. Trimethylamine [826]. Triploid [1357, 1105, 1405, 1305, 378, 1588, 1709, 1472, 1719, 131, 1908, 1895]. Triploidy [67, 1087]. Trisaccharides [326]. Triton [1667]. tritonis [1667]. trituberculatus [1361]. Trivalent [1554]. tRNAVal [292]. Trojan [1385]. Trophic [1828, 1374]. Tropical [377, 1866, 132, 375, 1162, 140, 241, 1382, 450]. Tropomyosin [135]. Tropomyosins [748]. Trout [1181, 1292, 144, 278, 37, 22, 1182, 1596, 1913, 843, 317, 1740, 735, 1212, 242, 550, 517, 717, 1242, 1381, 134, 5, 105, 1251, 1347, 1797, 352, 1378, 1052, 1130, 1185, 9, 249, 1144, 1958, 874, 225, 1708, 60, 1186, 1272, 1318, 841, 348, 1069, 1770, 1939, 926, 1176]. Truncated [747]. trutta [37, 348]. Trypsin [1061, 1475, 74]. Trypsinogen [156]. Trypsins [465]. tshawytscha [1600, 407]. tsingtauense [487]. Tsinling [1797, 1939]. tsinlingensis [1797, 1939]. Tsitsikamma [1082]. Tsitsikammamine [1082]. TtVgR [1943]. TU37 [346]. TU37-1 [346]. Tubastraea [1137]. Tube [952, 645, 502, 824, 1327]. Tube-Dwelling [502]. tuberculata [1699, 960, 1061, 1002, 444, 122, 960, 150]. Tubeworm [385, 1383]. Tubules [323, 1216]. Tudor [527]. Tulamben [1893]. Tumor [169, 158, 308, 159, 161, 157, 183, 1436, 895, 388, 1912]. Tumor-Associated [183]. Tumor-Inducing [161]. Tumorigenesis [1515, 1897]. Tuna [148, 315, 113]. Tunicate [338]. Tunicate-Associated [338]. Tuning [1666]. Tunisia [889]. Turbot [384, 1068, 1528, 81, 877, 1012, 1072, 1153, 1179, 1079, 1917, 593, 1539, 1400]. Turn [1522]. Turtle [1900]. Twelve [39]. Twenty [476, 1634]. Twenty-Six [1634]. Twice [1257]. Two [156, 123, 1228, 280, 395, 211, 1868, 1802, 1777, 858, 784, 1280, 261, 1103, 262, 814, 381, 730, 1416, 386, 1533, 1532, 87, 43, 1132, 301, 1610, 627, 1159, 868, 573, 1185, 772, 106, 906, 1013, 307, 887, 1460, 957, 213, 243, 264, 1677, 685, 1494, 710, 1795]. Two-color [264]. Two-dimensional [730]. Two-Headed [957]. Two-Photon [261]. Two-Stage [814]. Type [1028, 606, 882, 1358, 1678, 668, 461, 1901, 1328, 916, 5, 778, 630, 1683, 1682, 426, 620, 215, 103, 987, 5]. Types [1868, 523, 794, 1236, 620, 83, 1677, 864, 899]. Typing [441]. TYR [1934]. Tyrosinase [1690]. Tyrosinase-like [1690]. Tyrosine [1844].

U.S. [272]. U2 [99]. U6 [728, 1833]. UAS [1273, 1136, 833]. UAS-Regulated [833]. Ubiquitin [146]. Ubiquitination [1362]. UBOCC [1843]. UBOCC-A-208029 [1843]. Ueda [1015]. UK [1054]. Ulkenia [1190]. Ultrasonificated [925]. Ultraviolet [1227]. Ulva [1063, 958, 1083, 1923, 16, 773, 970]. Ulvan [1923, 1899]. Ulvan-Loaded [1899]. Umami [1015]. umbilicalis [1530]. Uncontaminated [1175]. Unconventional [1691]. Uncovering [965, 1711]. Uncovers [1312]. Uncultured [1394, 1455]. Undaria [1478]. Undergoing [891]. Underlying [1821, 1895]. Underpinnings [1781]. Understanding [1543, 1409, 1269, 1745, 660]. Underwater [683]. Undifferentiated [1581]. Unicellular [544, 1599]. unicinctus [914, 1042]. Uniparental [1867]. Unique [56, 731, 1045, 170]. United [20]. Units [1737]. Universal [91]. Unraveled [1801]. Unraveling [1691]. Unravels [1108]. UNS [537]. Unsaturated [1143, 819, 1808]. Untaps [1722]. Untranslated [1539]. Unusual [945, 1954]. Unveiling [1820]. Up-Regulation [544, 970, 1911]. Upaccumulated [1417]. upon [1938]. Upregulates [1522, 1524]. Upregulation [773]. Upstream [94, 905, 566]. Uptake [1294, 1337, 1341, 612]. Urchin [1099, 1328, 721, 824, 957, 1327, 1710, 1062]. Urchins [303]. Urechis [914, 1042]. ureilytica [949]. Urochordate [513]. Uronic [1442]. Use [1886, 1653, 176, 130, 335, 1740, 104, 263, 826, 1081, 332, 217, 372, 1529, 629, 743, 1403, 115, 1899, 429, 1656, 380, 446, 1304]. Used [420, 1177]. Useful [945, 781]. Usefulness [349]. Uses [1442]. Using [1161, 578, 1585, 533, 366, 1727, 1228, 1061, 569, 278, 46, 1594, 1681, 384, 1031, 934, 1282, 545, 1346, 1035, 1931, 1579, 842, 29, 1613, 1418, 1083, 609, 1367, 523, 1619, 1748, 427, 1503, 1654, 261, 1300, 251, 480, 248, 1131, 1281, 949, 752, 680, 640, 871, 1456, 1817, 1122, 491, 87, 13, 1012, 1670, 619, 1133, 407, 204, 1072, 456, 468, 1116, 1316, 1214, 1531, 892, 577, 114, 1325, 989, 84, 70, 113, 450, 978, 409, 1109, 836, 890, 55, 264, 277, 1399, 1586, 1668, 422, 446, 289, 1515, 1565, 1792, 1897, 1176, 154, 1551, 302, 447, 1377, 907, 833]. Using [1027, 1003, 1538, 1539, 504, 81, 305, 1642]. ussuriensis [1262]. Utility [250, 104, 460, 1672, 1191, 172, 1226]. Utilization [1835, 1889, 217, 1875]. Utilizes [293]. Utilizing [1406, 615, 192, 864]. UTR [1752]. UV [941, 418, 342, 180, 185, 687]. UV-A [941]. UV-Absorbing [342]. UV-Induced [180, 185]. UVB [1187, 760]. UVB-Damaged [760]. UVB-Irradiated [1187]. Uwa [97].

V [334]. V. [1922, 1077]. Vaccinated [514]. Vaccine [1715, 565, 1832, 1077, 898, 998]. Vacuolar [215]. Vacuolar-Type [215]. Vacuoles [42, 314]. vaginalis [1788]. Validating [1537]. Validation [975, 1146, 571, 940, 1561, 705]. Valine [192]. Value [1451]. Vanadium [1133, 1432, 1326, 1033, 215, 405, 1924]. Vanadium-Accumulating [1924]. Vanadium-binding [1033]. Vanadium-Dependent [1133]. Vanadium-Resistant [1326]. Vanadium-Rich [215, 405, 1033]. Vanadocytes [405]. Vancomycin [590]. Vancomycin-Resistant [590]. vannamei [578, 908, 280, 404, 1304, 34, 292, 1553, 1941, 1796, 356, 742, 1479, 515, 1087, 1906, 634, 437, 1598, 1586, 1668, 133, 1377, 1513, 1524, 848, 1849]. Var [1493, 1492, 1587, 1274, 1808, 355, 1633]. variabilis [678]. Variability [37, 421, 404, 460, 153, 252]. Variable [1308, 1097, 54, 1785, 1257]. Variables [1406]. Variant [1752, 1888, 771, 841]. Variants [119, 599, 586, 622, 906, 1781, 1655]. Variation [1873, 1181, 675, 419, 719, 783, 1793, 406, 1764, 212, 263, 394, 224, 1703, 1305, 1687, 1225, 208, 442, 500, 128, 881, 1315, 1489, 620, 108, 266, 1602, 437, 136, 518]. Variations [148, 1083, 1355, 1253]. Varied [1494]. Various [341, 1365, 1106]. Vasa [1851, 636, 1539]. vasa-Like [636]. Vaterite [1749]. VDAC [1417]. Vector [1366, 1203, 1833, 1944, 998]. Vector-Based [1833, 1944]. Vectors [217, 164, 3, 162, 231, 728, 560]. Vegetable [1163]. Vegetal [1244]. Vegetative [237]. Venerupis [997]. Venom [1228, 1569, 1120, 1350]. Venom-Related [1569]. Vent [1843, 440, 502, 1395, 904]. Vents [1320, 1314, 485]. Verasper [78]. vermiculophylla [1374]. veronii [1960, 1879]. verrucosa [1286]. Versatility [1363]. Vertebrate [1779, 1591]. Vertical [355]. Vesicles [1675]. vesiculosus [853]. Vestimentiferan [385]. VHb [938]. VHS [1179]. VHSV [1068, 1671, 1088]. Via [798, 1560, 1713, 1872, 1191, 631, 1763, 813, 334, 695, 1943, 1770, 1911]. Viability [1413, 1087]. Viability-Based [1413]. Viable [1746, 870]. Vibrio [1922, 1805, 1528, 1252, 217, 1165, 579, 1941, 218, 279, 1442, 555, 1480, 1441, 1050, 1153, 1450, 1505, 1705, 83, 84, 1617, 883, 1021, 1217, 744, 1201, 996, 899, 747, 592, 800, 998, 1689, 1782]. VII [438]. Violacein [685]. Violet [903]. Viral [969, 1068, 963, 514, 565, 1333, 1307, 1313, 1671, 1303, 1088, 1179, 1785, 753, 1375]. Viremia [1886]. virens [799]. virgatus [1592]. virginica [216, 583, 130, 1816, 256, 653, 529, 1881, 885, 1440, 479, 648, 447, 1714, 1157]. viridis [235, 364, 755, 881]. virilis [1757]. Virulence [1264, 1505, 898]. Virulence-Related [1264]. Virulent [1655]. Virus [969, 1886, 196, 993, 120, 1063, 1068, 1165, 1841, 134, 1671, 189, 115, 368, 627, 919, 79, 1088, 1179, 698, 693, 1785, 580, 1792, 1794, 1524, 895, 635, 1025].

Virus-Like [1792, 189, 368]. Viruses [281]. Visceral [1807, 1813]. Visible [1193]. Visual [1795, 1907]. Visualisation [1748]. Visualization [813, 1425, 84, 849, 1539]. Visualized [1950]. Vital [683]. Vitamin [1202, 978]. Vitellogenesis [634, 243, 855]. Vitellogenesis-Inhibiting [634, 243]. Vitellogenin [673, 243]. Vitellogenins [1918, 771]. Vitreoscilla [938]. Vitro [120, 1702, 1741, 377, 122, 1656, 1911, 1633, 1270, 281, 762, 1143, 659, 913, 1763, 253, 1478, 1194, 1529, 1483, 1359, 1365, 1433, 1249, 1694, 1217]. Viviparous [669]. Vivo [1359, 492, 1886, 328, 1294, 1057, 1763, 1478, 241, 1580, 700, 243, 179, 1293, 1539]. VLPs [1792]. vma5 [215]. Voltage [1417]. Voltage-Dependent [1417]. Volume [601]. Vote [1465]. VP28 [693, 1025]. vp28-siRNA [1025]. VP28dsRNA [698]. vs [768]. vulgaris [1227, 1058, 213]. vulgata [1113]. vulnificus [996, 800].

W7 [1808]. W7-5 [1808]. WA2 [1292]. WA2-67 [1292]. Walbaum [1292]. waleckii [1570]. Wall [237, 1597, 1640, 836]. Walleye [919, 895]. Wangia [986]. Warburg [1417]. Warm [1632]. Warm-Acclimated [1632]. Warning [235]. Warren [601]. Wastes [445, 787]. Wastewater [1548, 312, 691, 336]. Watasenia [1618]. Water [642, 652, 1320, 1337, 1174, 1713, 568, 766, 808, 770, 717, 1353, 1390, 1549]. Water-Soluble [642, 766, 1390]. Waterborne [1704]. Waters [809, 528, 196, 876, 884, 793, 236, 229, 450]. Wax [383]. Wb800 [764]. WC91 [879, 1044]. WC91-2 [879, 1044]. Weight [1502, 1478, 865]. Weights [1282]. WelK [1630]. Were [1193, 1278]. werneckii [1843]. Western [540]. WG [1630, 1817]. WGCNA [1883]. Whale [672]. Whales [540]. which [1522]. White [908, 1807, 197, 196, 993, 404, 1553, 1165, 1941, 253, 1600, 1097, 726, 1479, 771, 698, 693, 1087, 1325, 437, 580, 1586, 1668, 1960, 855, 1377, 1513, 1524, 848, 920, 1025]. White-Nodules [1807]. Whiteleg [578, 932, 634]. Whole [1852, 1672, 480, 1206, 1240, 1053, 1285, 796, 1647, 1561, 1874, 1847, 1761, 1609]. Whole-Cell [480]. Whole-Genome [1647, 1561, 1874, 1847, 1761]. Whole-organism [796]. Wide [1798, 1404, 1348, 1209, 1646, 1146, 1725, 1605, 1393, 1636, 1863, 1767, 1709, 1781, 1375, 1808, 1497, 1684, 1379, 1648, 1387, 1538, 1542, 1582]. Widely [1005]. Widespread [1351, 1022]. Width [871]. Wild [1018, 1793, 831, 153, 1149, 1533, 1532, 1845, 518]. Wille [1917]. Williopsis [879, 924, 1044]. Winter [61]. Within [261, 290, 772, 690, 1029, 100]. Without [147, 224, 1129]. WL [1630, 1817]. Working [181]. World [1749]. Worm [211, 914, 1042, 1876]. Worms [502]. Wound [1625, 1920]. Wrasse [421]. WSSV [1161, 1608, 1266, 1439, 712, 580, 744, 1524, 1025]. WW [1031]. wyckioides [1521].

X153 [453]. Xanthine [382]. Xanthophyll [941, 554]. Xenograft [1156]. Xenopus [152]. Xenostrobus [665]. Xestospongia [1701]. xiamenensis [1722]. Xiphophorus [273, 173, 178, 163, 171, 161, 180, 182, 185, 170, 174].

Xmrk [273]. Xmrk2 [161]. XRE [94]. XX [1242]. XY [1242, 1119, 1021]. XY-214 [1021]. Xylan [1021]. Xylose [1021]. Xylulose [1021].

Yangtze [1718]. Yarrowia [1295, 818, 1094, 761, 924]. Year [1374, 97]. Yeast [1487, 1843, 452, 1295, 1339, 1308, 692, 1094, 632, 1268, 706, 761, 879, 924, 3, 31, 215, 1808, 781]. Yeast-Displayed [1308]. Yeasts [312]. Yellow [1807, 1238, 1346, 1555, 1646, 1554, 1955, 1504, 1605, 680, 1956, 1636, 1119, 1550, 1275, 774, 1758, 1460, 1781, 1110, 1850, 1762, 1219, 1847, 1927, 1790, 1648, 1538, 1458]. Yellowfin [148, 1814, 1903]. Yellowtail [527, 1267, 1278]. Yendo [839]. Yersinia [910, 1069]. Yesso [1540, 1379, 1535, 1577]. yessoensis [1335, 1498, 1540, 687, 220, 221, 530, 1544, 1379, 1535, 1577]. yezoensis [1005, 1738, 1366, 1239, 813, 1015, 342, 1451, 849, 1109, 1177]. YHV [712]. Yield [317, 901]. Yielding [576, 355]. Yields [55]. ylmethanes [1276]. Yolk [771]. YY [1119].

Z [1292]. Zealand [543, 1441]. Zebra [567, 1168, 88]. Zebrafish [414, 611, 559, 1317, 1424, 275, 340, 563, 69, 938, 36, 1693, 509, 878, 888, 387, 1380, 152, 238, 631, 1779, 752, 247, 972, 1019, 1485, 566, 640, 1603, 1392, 1670, 1403, 1207, 1273, 919, 1588, 149, 343, 1819, 700, 728, 191, 110, 1253, 1957, 324, 264, 121, 560, 1789, 1152, 1136, 1293, 474, 1139, 1235, 1515, 1897, 522, 482, 1500, 1944, 833, 895, 1606, 1078]. zhangjiangensis [1839]. zhaozhouensis [1491]. Zhikong [1092, 1151, 1474, 131, 775, 655, 710, 688]. Zinc [1220, 98, 1316, 1176, 1166]. Zinc-Finger [1176]. ZLE [121]. Zn [631]. Zn-Superoxide [339]. Zoanthid [842, 1229]. Zoanthids [892]. Zoanthus [842]. Zoea [1586, 1668]. Zona [444]. zonale [689]. Zone [763]. Zoospore [686]. Zooxanthellae [734]. Zooxanthellate [803, 1011]. Zostera [704]. ZW [1499].

References

Chen:1999:MB

[1] Thomas T. Chen, Yves LeGal, and Shigetoh Miyachi. Marine Biotechnology. Marine Biotechnology, 1(1):1, January 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

      springer.com/article/10.1007/BF03391712.

Guillen:1999:SET

[2] Isabel Guillén, Jorge Berlanga, Carmen M. Valenzuela, Antonio Morales, José Toledo, Mario P. Estrada, Pedro Puentes, Orlando Hayes, and José de la Fuente. Safety evaluation of transgenic tilapia with accelerated growth. Marine Biotechnology, 1(1):2–14, January 1999. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

      link.springer.com/article/10.1007/PL00011746.

Ricaurte:1999:CPV

[3] Martha L. Ricaurte and Nadathur S. Govind. Construction of plasmid vectors and transformation of the marine yeast Debaryomyces hansenii. Marine Biotechnology, 1(1):15–19, January 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/PL00011745.

Tinti:1999:CAM

[4] Fausto Tinti, Andrea Colombari, Maria Vallisneri, Corrado Piccinetti, and Anna Maria Stagni. Comparative analysis of a mitochondrial DNA control region fragment amplified from three Adriatic flatfish species and molecular phylogenesis of pleuronectiformes. Marine Biotechnology, 1 (1):20–24, January 1999. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/PL00011747.

Krasnov:1999:EHG

[5] Aleksei Krasnov, Tiina I. Pitkänen, Mika Reinisalo, and Hannu Mölsä. Expression of human glucose transporter type 1 and rat hexokinase Type II complementary DNAs in rainbow trout embryos: Effects on glucose metabolism. Marine Biotechnology, 1(1):25–32, January 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/PL00011748.

Aranishi:1999:PCP

[6] Futoshi Aranishi. Purification and characterization of α_{1}-proteinase inhibitor from carp (Cyprinus carpio) serum. Marine Biotechnology, 1 (1):33–43, January 1999. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/PL00011749.

Khoo:1999:CAP

[7] Lester Khoo, David W. Robinette, and Edward J. Noga. Callinectin, an antibacterial peptide from blue crab, Callinectes sapidus, hemocytes. Marine Biotechnology, 1(1):44–51, January 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/PL00011750.

Ruiz-Ponte:1999:BRS

[8] C. Ruiz-Ponte, J. F. Samain, J. L. Sánchez, and J. L. Nicolas. The benefit of a Roseobacter species on the survival of scallop larvae. Marine Biotechnology, 1(1):52–59, January 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/PL00011751.

Reed:1999:GTL

[9] Kent M. Reed. Tc 1-like transposable elements in the genome of lake trout (Salvelinus namaycush). Marine Biotechnology, 1(1):60–67, January 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
PL00011752.

Matsuda:1999:AAM

[10] Masahiro Matsuda, Shiro Shigeta, and Koichi Okutani. Antiviral activities of marine Pseudomonas polysaccharides and their oversulfated derivatives. Marine Biotechnology, 1(1):68–73, January 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/PL00011753.

Blanchette:1999:PCG

[11] Brian N. Blanchette and Bal Ram Singh. Purification and characterization of the glutathione-S-transferases from the Northern quahog Mercinaria mercinaria. Marine Biotechnology, 1(1):74–80, January 1999. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/PL00011754.

Aranishi:1999:PCS

[12] Futoshi Aranishi. Purification and characterization of serum serpin from carp (Cyprinus carpio). Marine Biotechnology, 1(1):81–88, January 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic).

URL https://link.springer.com/article/10.1007/PL00011755.

Medina:1999:EMA

[13] Monica Medina, Ernesto Weil, and Alina M. Szmant. Examination of the Montastraea annularis species complex (Cnidaria: Scleractinia) using ITS and COI sequences. Marine Biotechnology, 1(1):89–97, January 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
PL00011756.

Kawahara:1999:CAS

[14] Hiroyuki Kawahara, Ryo Tamura, Seiko Ajioka, and Yoshikazu Shizuri. Convenient assay for settlement inducing substances of barnacles. Marine Biotechnology, 1(1):98–101, January 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/PL00011757.

Ahmad:1999:RAA

[15] Ahmad Shamsuddin Ahmad, Masahiro Matsuda, Shiro Shigeta, and Koichi Okutani. Revelation of antiviral activities by artificial sulfation of a glycosaminoglycan from a marine Pseudomonas. Marine Biotechnology, 1(1):102–106, January 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/PL00011743.

Nakanishi:1999:RMI

[16] Koichi Nakanishi, Miyuki Nishijima, Ann M. Nomoto, Ayano Yamazaki, and Naotsune Saga. Requisite morphologic interaction for attachment between Ulva pertusa (Chlorophyta) and symbiotic bacteria. Marine Biotechnology, 1(1):107–111, January 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/PL00011744.

Klinbunga:1999:MDD

[17] S. Klinbunga, D. J. Penman, B. J. McAndrew, and A. Tassanakajon. Mitochondrial DNA diversity in three populations of the giant tiger shrimp Penaeus monodon. Marine Biotechnology, 1(2):113–121, March 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic).

URL https://link.springer.com/article/10.1007/PL00011758.

Leclerc:1999:CHR

[18] Gilles M. Leclerc, Kaiping Han, Guy J. Leclerc, and Bert Ely. Characterization of a highly repetitive sequence conserved among the North American Morone species. Marine Biotechnology, 1(2):122–130, March 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic).

URL https://link.springer.com/article/10.1007/PL00011759.

Cho:1999:CDE

[19] Jung Jong Cho, Jae Hyung Lee, Se-Kwon Kim, Tae-Jin Choi, and Young Tae Kim. Complementary DNA encoding nm23 /NDP kinase gene from the Korean tiger shark Scyliorhinus torazame. Marine Biotechnology, 1(2):131–136, March 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/PL00011760.

Chapman:1999:SIG

[20] Robert W. Chapman, George R. Sedberry, Christopher C. Koenig, and Brandon M. Eleby. Stock identification of gag, Mycteroperca microlepis, along the Southeast Coast of the United States. Marine Biotechnology, 1(2):137–146, March 1999. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/PL00011761.

Ko:1999:PEM

[21] Yuan-Tih Ko, Marion Man-Ying Chan, Susan E. Ford, and Dunne Fong. A PCR–ELISA method for direct detection of the oyster pathogen Haplosporidium nelsoni. Marine Biotechnology, 1(2):147–154, March 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic).

URL https://link.springer.com/article/10.1007/PL00011762.

Carvan:1999:TGC

[22] Michael J. Carvan III, Larissa V. Ponomareva, and Daniel W. Nebert. Trout CYP1A3 gene: Recognition of fish DNA motifs by mouse regulatory proteins. Marine Biotechnology, 1(2):155–166, March 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/PL00011763.

Lindstrom:1999:MSI

[23] Daniel P. Lindstrom. Molecular species identification of newly hatched Hawaiian amphidromous gobioid larvae. Marine Biotechnology, 1(2): 167–174, March 1999. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/PL00011764.

Liu:1999:AFD

[24] X. Y. Liu, Y. Zohar, and W. Knibb. Association of foreign DNA with sperm of gilthead seabream, Sparus aurata, after sonication, freezing, and dimethyl sulfoxide treatments. Marine Biotechnology, 1(2): 175–183, March 1999. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/PL00011765.

Snyder:1999:RPS

[25] Mark J. Snyder. Ribosomal proteins S27E, P2, and L37A from marine invertebrates. Marine Biotechnology, 1(2):184–190, March 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/PL00011766.

Wiens:1999:PMS

[26] Matthias Wiens, Claudia Koziol, Renato Batel, and Werner E. G. Müller. Prolidase in the marine sponge Suberites domuncula: Enzyme activity, molecular cloning, and phylogenetic relationship. Marine Biotechnology, 1(2):191–199, March 1999. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/PL00011767.

Hwang:1999:IEP

[27] Sheng-Ping L. Hwang, Shing Kou Wang, Su Fen Wei, Li-Chuan Cheng, and Jeng Chang. Identification and expression pattern of DNA polymerase α gene in a marine diatom, Skeletonema costatum. Marine Biotechnology, 1(2):200–206, March 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/PL00011768.

Noguchi:1999:IBG

[28] Yukinori Noguchi, Asako Ishii, Ayako Matsushima, Daisuke Haishi, Ken ichi Yasumuro, Tomohisa Moriguchi, Takeshi Wada, Yoh Kodera, Misao Hiroto, Hiroyuki Nishimura, Mitsuo Sekine, and Yuji Inada. Isolation of Biopterin-α-glucoside from Spirulina (Arthrospira) platensis and its physiologic function. Marine Biotechnology, 1(2):207–210, March 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic).

URL https://link.springer.com/article/10.1007/PL00011769.

Gonzalez-Villasenor:1999:AGH

[29] Lucía Irene González-Villaseñor and Thomas T. Chen. Antibodies for growth hormone and prolactin using multiple antigen peptide immunogens. Marine Biotechnology, 1(3):211–220, May 1999. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/PL00011770.

Majumdar:1999:NMR

[30] K. C. Majumdar, K. Nasaruddin, K. Ravinder, C. S. Sundaram, P. Manickam, and S. Shivaji. ^{31}P nuclear magnetic resonance studies on the phosphorus-containing metabolites of the developing embryos of a freshwater catfish, Clarias batrachus (L.). Marine Biotechnology, 1 (3):221–229, May 1999. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/PL00011771.

Thome:1999:OGI

[31] Patricia E. Thomé and Robert K. Trench. Osmoregulation and the genetic induction of glycerol-3-phosphate dehydrogenase by NaCl in the euryhaline yeast Debaryomyces hansenii. Marine Biotechnology, 1(3): 230–238, May 1999. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/PL00011772.

Falciatore:1999:TNR

[32] Angela Falciatore, Raffaella Casotti, Catherine Leblanc, Chiara Abrescia, and Chris Bowler. Transformation of nonselectable reporter genes in marine diatoms. Marine Biotechnology, 1(3):239–251, May 1999. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/PL00011773.

Oppen-Berntsen:1999:SEP

[33] Dag O. Oppen-Berntsen, Augustine Arukwe, Fekadu Yadetie, James B. Lorens, and Rune Male. Salmon eggshell protein expression: a marker for environmental estrogens. Marine Biotechnology, 1(3):252–260, May 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic).

URL https://link.springer.com/article/10.1007/PL00011774.

France:1999:IGD

[34] Scott C. France, Nicholas Tachino, Thomas F. Duda, Jr., Robert A. Shleser, and Stephen R. Palumbi. Intraspecific genetic diversity in the marine shrimp Penaeus vannamei: Multiple polymorphic elongation Factor-1α loci revealed by intron sequencing. Marine Biotechnology, 1 (3):261–268, May 1999. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/PL00011775.

Gomez-Chiarri:1999:ICA

[35] Marta Gómez-Chiarri, Vicky L. Kirby, and Dennis A. Powers. Isolation and characterization of an actin promoter from the red abalone (Haliotis rufescens). Marine Biotechnology, 1(3):269–278, May 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/PL00011776.

Hamilton:1999:ICD

[36] Lorraine C. Hamilton and Jonathan M. Wright. Isolation of complementary DNAs coding for a receptor for activated C kinase (RACK) from zebrafish (Danio rerio) and tilapia (Oreochromis niloticus): Constitutive developmental and tissue expression. Marine Biotechnology, 1 (3):279–285, May 1999. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/PL00011777.

Cagigas:1999:CAG

[37] M. E. Cagigas, E. Vazquez, G. Blanco, and J. A. Sánchez. Combined assessment of genetic variability in populations of brown trout (Salmo trutta L.) based on allozymes, microsatellites, and RAPD markers. Marine Biotechnology, 1(3):286–296, May 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/PL00011778.

Westerman:1999:QAC

[38] Mark E. Westerman, G. Joan Holt, and Leonard DiMichele. Quantitative assay of cyclin-dependent kinase activity as a sensitive marker of cell proliferation in marine teleost larvae. Marine Biotechnology, 1 (3):297–310, May 1999. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/PL00011779.

Wimberger:1999:ICT

[39] Peter Wimberger, Jenni Burr, Andy Gray, Andres Lopez, and Paul Bentzen. Isolation and characterization of twelve microsatellite loci for rockfish (Sebastes). Marine Biotechnology, 1(3):311–315, May 1999. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/PL00011780.

Hogan:1999:MIC

[40] Robert J. Hogan, Geoffrey C. Waldbieser, Cheryl A. Goudie, Aurita Antao, Ulla B. Godwin, Melanie R. Wilson, Norman W. Miller, L. William Clem, Thomas J. McConnell, William R. Wolters, and V. Gregory Chinchar. Molecular and immunologic characterization of gynogenetic channel catfish (Ictalurus punctatus). Marine Biotechnology, 1 (4):317–327, July 1999. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/PL00011781.

Ding:1999:NBC

[41] J. L. Ding, F. M. Y. Fung, G. W. S. Ng, and L. M. Chou. Novel bioactivities from a coral, Galaxea fascicularis: DNase-like activity and apoptotic activity against a multiple-drug–resistant leukemia cell line. Marine Biotechnology, 1(4):328–336, July 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/PL00011782.

Wada:1999:TFB

[42] Shun ichi Wada, Shigeki Matsunaga, Nobuhiro Fusetani, and Shugo Watabe. Theonellamide f, a bicyclic peptide marine toxin, induces formation of vacuoles in 3Y1 rat embryonic fibroblast. Marine Biotechnology, 1(4):337–341, July 1999. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/PL00011783.

Ogino:1999:ITF

[43] Yukiko Ogino, Takao Itakura, Ryoichi Mitsuo, and Mamoru Sato. Induction of two forms of eel cytochrome P450 1A genes by 3methylcholanthrene. Marine Biotechnology, 1(4):342–345, July 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic).

URL https://link.springer.com/article/10.1007/PL00011784.

Wu:1999:LSH

[44] Qingyu Wu, Yoshihiro Shiraiwa, Hiroshi Takeda, Guoying Sheng, and Jiamo Fu. Liquid-saturated hydrocarbons resulting from pyrolysis of the marine coccolithophores Emiliania huxleyi and Gephyrocapsa oceanica. Marine Biotechnology, 1(4):346–352, July 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/PL00011785.

Mitsuo:1999:CSC

[45] Ryoichi Mitsuo, Takao Itakura, and Mamoru Sato. Cloning and sequencing of cytochrome P450 1A complementary DNA in eel (Anguilla japonica). Marine Biotechnology, 1(4):353–358, July 1999. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/PL00011786.

Boyd:1999:SMB

[46] Kenneth G. Boyd, Andrew Mearns-Spragg, and J. Grant Burgess. Screening of marine bacteria for the production of microbial repellents using a spectrophotometric chemotaxis assay. Marine Biotechnology, 1 (4):359–363, July 1999. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/PL00011787.

Ogino:1999:FAP

[47] Yukiko Ogino, Takao Itakura, Hironori Kato, Jun ya Aoki, and Mamoru Sato. Functional analysis of promoter region from eel cytochrome P450 1A1 gene in transgenic medaka. Marine Biotechnology, 1(4):

364–370, July 1999. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/PL00011788.

Aoki:1999:ISA

[48] Jun ya Aoki, Takao Itakura, Hironori Kato, and Mamoru Sato. Isolation and sequence analysis of the eel cytochrome P450 1A1 gene. Marine Biotechnology, 1(4):371–375, July 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/PL00011789.

Lyndon:1999:FGM

[49] A. R. Lyndon. Fish growth in marine culture systems: a challenge for biotechnology. Marine Biotechnology, 1(4):376–379, July 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/PL00011790.

Lee:1999:CAB

[50] Hak Sung Lee and Bohumil Volesky. Characteristics of aluminum biosorption by Sargassum fluitans biomass. Marine Biotechnology, 1 (4):380–383, July 1999. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/PL00011791.

Bultel-Ponce:1999:MSA

[51] Valérie Bultel-Poncé, Jean-Pascal Berge, Cécile Debitus, Jean-Louis Nicolas, and Michèle Guyot. Metabolites from the sponge-associated bacterium Pseudomonas species. Marine Biotechnology, 1(4):384–390, July 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
PL00011792.

Li:1999:MDS

[52] Lina Li, Chiaki Kato, and Koki Horikoshi. Microbial diversity in sediments collected from the deepest cold-seep area, the Japan Trench. Marine Biotechnology, 1(4):391–400, July 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/PL00011793.

Miralto:1999:EDI

[53] A. Miralto, A. Ianora, S. A. Poulet, G. Romano, I. Buttino, and S. Scala. Embryonic development in invertebrates is arrested by inhibitory compounds in diatoms. Marine Biotechnology, 1(4):401–402, July 1999. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/PL00011794.

Reichow:1999:HVM

[54] Denise Reichow and Michael J. Smith. Highly variable microsatellites in the California market squid Loligo opalescens. Marine Biotechnology, 1(4):403–406, July 1999. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/PL00011795.

Wakabayashi:1999:PHY

[55] Tomoaki Wakabayashi, Toru Kuboi, Takuya Tuboi, Masayuki Kaji, and Masakazu Hara. Preparation of high yields of algal protoplasts using buccal juice of sea hare and commercial cellulase. Marine Biotechnology, 1(4):407–410, July 1999. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/PL00011796.

Bakke:1999:SCU

[56] Ingrid Bakke, Gerald F. Shields, and Steinar Johansen. Sequence characterization of a unique intergenic spacer in Gadiformes mitochondrial DNA. Marine Biotechnology, 1(5):411–415, September 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/PL00011797.

Miya:1999:OMG

[57] Masaki Miya and Mutsumi Nishida. Organization of the mitochondrial genome of a deep-sea fish, Gonostoma gracile (Teleostei: Stomiiformes): First example of transfer RNA gene rearrangements in bony fishes. Marine Biotechnology, 1(5):416–426, September 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/PL00011798.

Clare:1999:ANA

[58] Anthony S. Clare, Daniel Rittschof, Donald J. Gerhart, Irving R. Hooper, and Joseph Bonaventura. Antisettlement and narcotic action of analogues of diterpene marine natural product antifoulants from octocorals. Marine Biotechnology, 1(5):427–436, September 1999. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/PL00011799.

Liu:1999:DPE

[59] Zhanjiang Liu, Attila Karsi, and Rex A. Dunham. Development of polymorphic EST markers suitable for genetic linkage mapping of catfish. Marine Biotechnology, 1(5):437–447, September 1999. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/PL00011800.

Takeuchi:1999:GFP

[60] Yutaka Takeuchi, Goro Yoshizaki, and Toshio Takeuchi. Green fluorescent protein as a cell-labeling tool and a reporter of gene expression in transgenic rainbow trout. Marine Biotechnology, 1(5):448– 457, September 1999. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/PL00011801.

Douglas:1999:WFE

[61] Susan E. Douglas, Jeffrey W. Gallant, Charles E. Bullerwell, Camille Wolff, Janet Munholland, and Michael E. Reith. Winter flounder expressed sequence tags: Establishment of an EST database and identification of novel fish genes. Marine Biotechnology, 1(5):458–464, September 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
PL00011802.

Lehnert:1999:TSE

[62] Sigrid A. Lehnert, Kate J. Wilson, Keren Byrne, and Stephen S. Moore. Tissue-specific expressed sequence tags from the black tiger shrimp Penaeus monodon. Marine Biotechnology, 1(5):465–476, September 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic).

URL https://link.springer.com/article/10.1007/PL00011803.

Aoki:1999:SCC

[63] Takashi Aoki, Bo-Hye Nam, Ikuo Hirono, and Eiichi Yamamoto. Sequences of 596 cDNA clones (565,977 bp) of Japanese flounder (Paralichthys olivaceus) leukocytes infected with hirame rhabdovirus. Marine Biotechnology, 1(5):477–488, September 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/PL00011804.

Cancre:1999:SGF

[64] Isabelle Cancre, Rozenn Ravallec, Alain Van Wormhoudt, Even Stenberg, Asbjoern Gildberg, and Yves Le Gal. Secretagogues and growth factors in fish and crustacean protein hydrolysates. Marine Biotechnology, 1(5):489–494, September 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/PL00011805.

Marwick:1999:BIP

[65] Jamie D. Marwick, Phillip C. Wright, and J. Grant Burgess. Bioprocess intensification for production of novel marine bacterial antibiotics through bioreactor operation and design. Marine Biotechnology, 1(5): 495–507, September 1999. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/PL00011806.

Osinga:1999:CMS

[66] Ronald Osinga, Johannes Tramper, and René H. Wijffels. Cultivation of marine sponges. Marine Biotechnology, 1(6):509–532, November 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic).

URL https://link.springer.com/article/10.1007/PL00011807.

Razak:1999:GPG

[67] Shaharudin Abdul Razak, Gyu-Lin Hwang, M. Azizur Rahman, and Norman Maclean. Growth performance and gonadal development of growth enhanced transgenic tilapia Oreochromis niloticus (L.) following heat-shock-induced triploidy. Marine Biotechnology, 1(6):533–544, November 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-

2236 (electronic). URL https://link.springer.com/article/10.
1007/PL00011808.

Dunham:1999:PAT

[68] Rex A. Dunham, Chanagun Chitmanat, Amy Nichols, Brad Argue, Dennis A. Powers, and Thomas T. Chen. Predator avoidance of transgenic channel catfish containing salmonid growth hormone genes. Marine Biotechnology, 1(6):545–551, November 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/PL00011809.

Fahrenkrug:1999:DGE

[69] Scott C. Fahrenkrug, Karl J. Clark, Mark O. Dahlquist, and Perry B. Hackett, Jr. Dicistronic gene expression in developing zebrafish. Marine Biotechnology, 1(6):552–561, November 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/PL00011810.

Sunamura:1999:BMM

[70] Michinari Sunamura, Yosuke Koga, and Kouichi Ohwada. Biomass measurement of methanogens in the sediments of Tokyo Bay using archaeol lipids. Marine Biotechnology, 1(6):562–568, November 1999. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/PL00011811.

Muller:1999:IAS

[71] Werner E. G. Müller, Wolfram Wimmer, Wolfgang Schatton, Markus Böhm, Renato Batel, and Zelimir Filic. Initiation of an aquaculture of sponges for the sustainable production of bioactive metabolites in open systems: Example, Geodia cydonium. Marine Biotechnology, 1(6): 569–579, November 1999. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/PL00011812.

Lewis:1999:BPT

[72] Tom E. Lewis, Peter D. Nichols, and Thomas A. McMeekin. The biotechnological potential of thraustochytrids. Marine Biotechnology, 1(6):580– 587, November 1999. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/PL00011813.

LeRoux:1999:DPP

[73] Frédérique Le Roux, Corinne Audemard, Antoine Barnaud, and Franck Berthe. DNA probes as potential tools for the detection of Marteilia refringens. Marine Biotechnology, 1(6):588–597, November 1999. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/PL00011814.

Spilliaert:1999:GAC

[74] Rémi Spilliaert and Ágústa Gudmundsdóttir. Atlantic cod trypsin Y — member of a novel trypsin group. Marine Biotechnology, 1(6):598– 607, November 1999. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/PL00011815.

Dai:2000:ALM

[75] Jixun Dai, Hai Wang, Baoqin Han, Jingjie Hu, and Zhenmin Bao. Application of live monocells from macroalgae to shellfish seed production. Marine Biotechnology, 2(1):1–4, January 2000. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s101269900001.

Suetsuna:2000:APP

[76] Kunio Suetsuna. Antioxidant peptides from the protease digest of prawn (Penaeus japonicus) muscle. Marine Biotechnology, 2(1):5–10, January 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s101269900002.

Frischer:2000:DAS

[77] Marc E. Frischer, Jean M. Danforth, Libby C. Tyner, Jay R. Leverone, Dan C. Marelli, William S. Arnold, and Norman J. Blake. Development of an Argopecten- specific 18S rRNA targeted genetic probe. Marine Biotechnology, 2(1):11–20, January 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s101269900003.

Peyush:2000:MCG

[78] Punia Peyush, Shunsuke Moriyama, Akiyoshi Takahashi, and Hiroshi Kawauchi. Molecular cloning of growth hormone complementary DNA in barfin flounder (Verasper moseri). Marine Biotechnology, 2(1): 21–26, January 2000. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101269900004.

Phenix:2000:CCE

[79] K. V. Phenix, B. McKenna, R. Fitzpatrick, L. Vaughan, G. Atkins, P. Liljestrom, and D. Todd. Cell culture evaluation of the semliki forest virus expression system as a novel approach for antigen delivery and expression in fish. Marine Biotechnology, 2(1):27–37, January 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s101269900005.

Boehm:2000:CSD

[80] E. W. A. Boehm, O. Gibson, and E. Lubzens. Characterization of satellite DNA sequences from the commercially important marine rotifers Brachionus rotundiformis and Brachionus plicatilis. Marine Biotechnology, 2(1):38–48, January 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s101269900006.

Iyengar:2000:IMR

[81] Arati Iyengar, Sanit Piyapattanakorn, David M. Stone, Diana A. Heipel, Bari R. Howell, Stephen M. Baynes, and Norman Maclean. Identification of microsatellite repeats in turbot (Scophthalmus maximus) and Dover sole (Solea solea) using a RAPD-based technique: Characterization of microsatellite markers in Dover sole. Marine Biotechnology, 2 (1):49–56, January 2000. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101269900007.

Yamano:2000:PCA

[82] Naoko Yamano, Noriko Higashida, Chieko Endo, Nami Sakata, Shizu Fujishima, Akihiko Maruyama, and Takanori Higashihara. Purification and characterization of N-acetylglucosamine-6-phosphate deacetylase from a psychrotrophic marine bacterium, Alteromonas species. Marine Biotechnology, 2(1):57–64, January 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s101269900008.

Sugimura:2000:CSA

[83] Itsuro Sugimura, Tomoo Sawabe, and Yoshio Ezura. Cloning and sequence analysis of Vibrio halioticoli genes encoding three types of polyguluronate lyase. Marine Biotechnology, 2(1):65–73, January 2000. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s101269900010.

Sugimura:2000:SPC

[84] Itsuro Sugimura, Tomoo Sawabe, and Yoshio Ezura. In situ polymerase chain reaction visualization of Vibrio halioticoli using alginate lyase gene AlyVG2. Marine Biotechnology, 2(1):74–79, January 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s101269900009.

Davey:2000:FCH

[85] M. L. Davey, M. R. Hall, R. H. Willis, R. W. A. Oliver, M. J. Thurn, and K. J. Wilson. Five crustacean hyperglycemic family hormones of Penaeus monodon: Complementary DNA sequence and identification in single sinus glands by electrospray ionization–Fourier transform mass spectrometry. Marine Biotechnology, 2(1):80–91, January 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s101269900011.

Weinmayr:2000:ICH

[86] Gudrun Weinmayr, Dominique Vautrin, and Michel Solignac. Isolation and characterization of highly polymorphic microsatellites from the polychaete Pectinaria koreni. Marine Biotechnology, 2(1):92–99, January 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s101269900012.

McCormack:2000:CAT

[87] Grace P. McCormack, Richard Powell, and Brendan F. Keegan. Comparative analysis of two populations of the brittle star Amphiura filiformis (Echinodermata: Ophiuroidea) with different life history strategies using RAPD markers. Marine Biotechnology, 2(1):100–106, January 2000. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s101269900013.

Gibbs:2000:GGM

[88] P. D. L. Gibbs and M. C. Schmale. GFP as a genetic marker scorable throughout the life cycle of transgenic zebra fish. Marine Biotechnology, 2(2):107–125, March 2000. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101269900014.

Tan:2000:PMC

[89] Siok Hwee Tan, Bernard M. Degnan, and Sigrid A. Lehnert. The Penaeus monodon chitinase 1 gene is differentially expressed in the hepatopancreas during the molt cycle. Marine Biotechnology, 2(2): 126–135, March 2000. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101269900016.

Berthelin:2000:HCG

[90] Clothilde Berthelin, Kristell Kellner, and Michel Mathieu. Histological characterization and glucose incorporation into glycogen of the Pacific oyster Crassostrea gigas storage cells. Marine Biotechnology, 2(2): 136–145, March 2000. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101269900017.

Chen:2000:UPA

[91] Chaolun Allen Chen and Jr-Kai Yu. Universal primers for amplification of mitochondrial small subunit ribosomal RNA-encoding gene in scleractinian corals. Marine Biotechnology, 2(2):146–153, March 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s101269900018.

Yoshizaki:2000:SOG

[92] G. Yoshizaki, W. Jin, R. Patiño, P. Thomas, and L. Janecek. Structural organization of the Atlantic croaker connexin 32.2 gene and its 5′ flanking region. Marine Biotechnology, 2(2):154–160, March 2000. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s101269900019.

Lee:2000:ISF

[93] Jae-Seong Lee. The internally self-fertilizing hermaphroditic teleost Rivulus marmoratus (Cyprinodontiformes, Rivulidae) β-actin gene: Amplification and sequence analysis with conserved primers. Marine Biotechnology, 2(2):161–166, March 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s101269900020.

Aoki:2000:ECP

[94] Jun ya Aoki, Mamoru Sato, and Takao Itakura. An eel cytochrome P450 1A gene having no XRE sequences in its 5′ upstream region of 1600 bp. Marine Biotechnology, 2(2):167–172, March 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s101269900021.

Nowell:2000:CEE

[95] Mari A. Nowell, Deborah M. Power, Pedro M. Guerreiro, Lynda Llewellyn, Vimi Ramsurn, Trevor Wigham, and Glen E. Sweeney. Cloning and expression of an elongation Factor-1α in sea bream (Sparus aurata) larvae and adult tissue. Marine Biotechnology, 2(2):173–179, March 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s101269900022.

Klinbunga:2000:GDS

[96] S. Klinbunga, A. Boonyapakdee, and B. Pratoomchat. Genetic diversity and species-diagnostic markers of mud crabs (genus Scylla) in Eastern Thailand determined by RAPD analysis. Marine Biotechnology, 2 (2):180–187, March 2000. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101269900023.

Kitamura:2000:OMB

[97] Shin-Ichi Kitamura and Satoru Suzuki. Occurrence of marine birnavirus through the year in coastal seawater in the uwa sea. Marine Biotechnology, 2(2):188–194, March 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s101269900025.

Lemoine:2000:MIM

[98] S. Lemoine, Y. Bigot, D. Sellos, R. P. Cosson, and M. Laulier. Metallothionein isoforms in Mytilus edulis (Mollusca, Bivalvia): Complementary DNA characterization and quantification of expression in different organs after exposure to cadmium, zinc, and copper. Marine Biotechnology, 2 (2):195–203, March 2000. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101269900026.

Lundin:2000:GSU

[99] Maria Lundin, Birthe Mikkelsen, Monica Gudim, and Mohasina Syed. Gene structure of the U2 snRNP-specific A′ protein gene from Salmo salar: Alternative transcripts observed. Marine Biotechnology, 2(2): 204–211, March 2000. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101269900015.

Gillor:2000:ITL

[100] O. Gillor, S. Carmeli, Y. Rahamim, Z. Fishelson, and M. Ilan. Immunolocalization of the toxin latrunculin B within the red sea sponge Negombata magnifica (Demospongiae, Latrunculiidae). Marine Biotechnology, 2(3):213–223, May 2000. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101260000026.

Delort:2000:AIF

[101] Eric Delort, Naoharu Watanabe, Hideo Etoh, Kanzo Sakata, and Hubert Jean Ceccaldi. Analysis of initial fouling process in coastal environment: Effects of settlement, attachment, and metamorphosis promoters. Marine Biotechnology, 2(3):224–230, May 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s101269900024.

Mendez:2000:IGH

[102] Ana T. Méndez, José L. Roig-López, Pedro Santiago, Carlos Santiago, and José E. García-Arrarás. Identification of Hox gene sequences in the sea cucumber Holothuria glaberrima Selenka (Holothuroidea: Echinodermata). Marine Biotechnology, 2(3):231–240, May 2000. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s101269900027.

Hikima:2000:MCN

[103] Jun ichi Hikima, Ikuo Hirono, and Takashi Aoki. Molecular cloning and novel repeated sequences of a C-type lysozyme gene in Japanese flounder (Paralichthys olivaceus). Marine Biotechnology, 2(3):241–247, May 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s101269900028.

Guillemaud:2000:IUM

[104] T. Guillemaud, F. Almada, R. Serrão Santos, and M. L. Cancela. Interspecific utility of microsatellites in fish: a case study of (CT)_{n} and (GT)_{n} markers in the shanny Lipophrys pholis (Pisces: Blenniidae) and their use in other blennioidei. Marine Biotechnology, 2(3):248–253, May 2000. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s101269900029.

Lee:2000:SEF

[105] Jong-Young Lee, Ikuo Hirono, and Takashi Aoki. Stable expression of a foreign gene, delivered by gene gun, in the muscle of rainbow trout Oncorhynchus mykiss. Marine Biotechnology, 2(3):254–258, May 2000. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s101269900030.

Sakharov:2000:PSP

[106] Ivan Sakharov and Gloria Astrid Prieto. Purification and some properties of two carboxypeptidases from the hepatopancreas of the crab Paralithodes camtschatica. Marine Biotechnology, 2(3):259–266, May 2000. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s101269900031.

Hirashita:2000:MAR

[107] Takashi Hirashita, Kazuhiko Ichimi, Shigeru Montani, Mika Nomura, and Shigeyuki Tajima. Molecular analysis of ribosomal RNA gene of red tide algae obtained from the Seto Inland Sea. Marine Biotechnology, 2 (3):267–273, May 2000. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101269900034.

Tinti:2000:MDV

[108] Fausto Tinti, Corrado Piccinetti, Stefano Tommasini, and Maria Vallisneri. Mitochondrial DNA variation, phylogenetic relationships, and evolution of four Mediterranean genera of soles (Soleidae, Pleuronectiformes). Marine Biotechnology, 2(3):274–284, May 2000. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s101269900035.

Wada:2000:ICB

[109] Shun ichi Wada, Shigeki Matsunaga, Nobuhiro Fusetani, and Shugo Watabe. Interaction of cytotoxic bicyclic peptides, theonellamides A and F, with glutamate dehydrogenase and 17β-hydroxysteroid dehydrogenase IV. Marine Biotechnology, 2(3):285–292, May 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s101260000006.

Torgersen:2000:GGM

[110] Jacob Torgersen, Philippe Collas, and Peter Aleström. Gene-gun-mediated transfer of reporter genes to somatic zebrafish (Danio rerio) tissues. Marine Biotechnology, 2(3):293–300, May 2000. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s101260000002.

Felip:2000:AAC

[111] Alicia Felip, Gonzalo Martínez-Rodríguez, Francesc Piferrer, Manuel Carrillo, and Silvia Zanuy. AFLP analysis confirms exclusive maternal genomic contribution of meiogynogenetic sea bass (Dicentrarchus labrax L.). Marine Biotechnology, 2(3):301–306, May 2000. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s101260000000.

Anonymous:2000:MA

[112] Anonymous. Meeting announcements. Marine Biotechnology, 2(3):307, May 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s101260000027.

Takeyama:2000:DBG

[113] Haruko Takeyama, Hisahito Tsuzuki, Seinen Chow, Hideki Nakayama, and Tadashi Matsunaga. Discrimination between Atlantic and Pacific subspecies of Northern bluefin tuna (Thunnus thynnus) by magneticcapture hybridization using bacterial magnetic particles. Marine Biotechnology, 2(4):309–313, July 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.com/

article/10.1007/s10126000010.

Sera:2000:IPA

[114] Yutaka Sera, Shigeo Iida, Kyoko Adachi, and Yoshikazu Shizuri. Improved plate assay for antifouling substances using blue mussel Mytilus edulis galloprovincialis. Marine Biotechnology, 2(4):314–318, July 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s101260000011.

Nunan:2000:UPC

[115] L. M. Nunan, B. T. Poulos, and D. V. Lightner. Use of polymerase chain reaction for the detection of infectious hypodermal and hematopoietic necrosis virus in penaeid shrimp. Marine Biotechnology, 2(4): 319–328, July 2000. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101260000003.

Palmisano:2000:TSI

[116] Aldo N. Palmisano, James R. Winton, and Walton W. Dickhoff. Tissuespecific induction of Hsp90 mRNA and plasma cortisol response in Chinook salmon following heat shock, seawater challenge, and handling challenge. Marine Biotechnology, 2(4):329–338, July 2000. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s101260000005.

Supungul:2000:MPP

[117] Premruethai Supungul, Pitak Sootanan, Sirawut Klinbunga, Wongpathom Kamonrat, Padermsak Jarayabhand, and Anchalee Tassanakajon. Microsatellite polymorphism and the population structure of the black tiger shrimp (Penaeus monodon) in Thailand. Marine Biotechnology, 2(4):339–347, July 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.com/

article/10.1007/s101260000012.

Inoue:2000:CSP

[118] Koji Inoue, Kei Kamino, Futaba Sasaki, Satoshi Odo, and Shigeaki Harayama. Conservative structure of the plaque matrix protein of mussels in the genus Mytilus. Marine Biotechnology, 2(4):348–351, July 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s101260000013.

Floriolli:2000:MSE

[119] Renee Y. Floriolli, Johannes von Langen, and J. Herbert Waite. Marine surfaces and the expression of specific byssal adhesive protein variants in Mytilus. Marine Biotechnology, 2(4):352–363, July 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s101269900032.

Chen:2000:IPN

[120] Jyh-Yih Chen, Jyh-Yeu Chen, Wan-Ken Chu, Huey-Nan Wu, Ya-Li Hsu, and Jen-Leih Wu. Infectious pancreatic necrosis virus RNA cleavage in vitro by hammerhead ribozymes and enhancement of ribozyme catalysis by oligonucleotide facilitators. Marine Biotechnology, 2(4): 364–375, July 2000. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101269900033.

Wang:2000:OZA

[121] Wen-Der Wang, Jun-Chyi Wu, Hwei-Jan Hsu, Zwe-Ling Kong, and Chin-Hwa Hu. Overexpression of a zebrafish ARNT2-like factor represses CYP1A transcription in ZLE cells. Marine Biotechnology, 2 (4):376–386, July 2000. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101260000001.

Poncet:2000:VSP

[122] J.-M. Poncet, A. Serpentini, B. Thiébot, C. Villers, J. Bocquet, E. Boucaud-Camou, and J.-M. Lebel. In vitro synthesis of proteoglycans and collagen in primary cultures of mantle cells from the nacreous mollusk, Haliotis tuberculata: a new model for study of molluscan extracellular matrix. Marine Biotechnology, 2(4):387–398, July 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/PL00021685.

Amano:2000:TTA

[123] Naoko Amano, Haruko Takeyama, Takehiko Kusama, Mitsuru Sakaizumi, Takashi Oshiro, and Tadashi Matsunaga. Two tandemly arrayed Transfer-RNA-Derived SINEs of the medaka (Oryzias latipes). Marine Biotechnology, 2(4):399–403, July 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/PL00021686.

Han:2000:ICM

[124] Kaiping Han, Li Li, Gilles M. Leclerc, Alison M. Hays, and Bert Ely. Isolation and characterization of microsatellite loci for striped bass (Morone saxatilis). Marine Biotechnology, 2(5):405–408, September 2000. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s101260000014.

Miyashita:2000:CDC

[125] T. Miyashita, R. Takagi, M. Okushima, S. Nakano, H. Miyamoto, E. Nishikawa, and A. Matsushiro. Complementary DNA cloning and characterization of pearlin, a new class of matrix protein in the nacreous layer of oyster pearls. Marine Biotechnology, 2(5):409–418, September 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
PL00021687.

Figueras:2000:SSS

[126] Antonio Figueras, Gema Lorenzo, M. Camino Ordás, Manolo Gouy, and Beatriz Novoa. Sequence of the small subunit ribosomal RNA gene of Perkinsus atlanticus-like isolated from carpet shell clam in Galicia, Spain. Marine Biotechnology, 2(5):419–428, September 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s101260000015.

Paul:2000:DPR

[127] John H. Paul, Jordan B. Kang, and F. Robert Tabita. Diel patterns of regulation of rbcL transcription in a cyanobacterium and a prymnesiophyte. Marine Biotechnology, 2(5):429–436, September 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s101260000016.

Murray:2000:LRH

[128] Brent W. Murray, Philomeen Nilsson, Zofia Zaleska-Rutczynska, Holger Sültmann, and Jan Klein. Linkage relationships and haplotype variation of the major histocompatibility complex class I A genes in the cichlid fish Oreochromis niloticus. Marine Biotechnology, 2(5):437– 448, September 2000. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/PL00021683.

Medina:2000:CFM

[129] Mónica Medina and Patrick J. Walsh. Comparison of four Mendelian loci of the California sea hare (Aplysia californica) from populations of the Coast of California and the Sea of Cortez. Marine Biotechnology, 2(5): 449–455, September 2000. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101260000020.

Day:2000:UCP

[130] J. Michael Day, Dean E. Franklin, and Bonnie L. Brown. Use of competitive PCR to detect and quantify Haplosporidium nelsoni infection (MSX disease) in the Eastern oyster (Crassostrea virginica). Marine Biotechnology, 2(5):456–465, September 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s101260000021.

Yang:2000:TTZ

[131] Huiping Yang, Fusui Zhang, and Ximing Guo. Triploid and tetraploid Zhikong scallop, Chlamys farreri Jones et Preston, produced by inhibiting polar body I. Marine Biotechnology, 2(5):466–475, September 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic).

URL https://link.springer.com/article/10.1007/PL00021688.

Klinbunga:2000:DSS

[132] S. Klinbunga, P. Ampayup, A. Tassanakajon, P. Jarayabhand, and W. Yoosukh. Development of species-specific markers of the tropical oyster (Crassostrea belcheri) in Thailand. Marine Biotechnology, 2(5): 476–484, September 2000. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101260000028.

Yepiz-Plascencia:2000:SHH

[133] Gloria Yepiz-Plascencia, Teresa Gollas Galván, Francisco Vargas-Albores, and Mónica García-Bañuelos. Synthesis of hemolymph highdensity lipoprotein β-glucan binding protein by Penaeus vannamei shrimp hepatopancreas. Marine Biotechnology, 2(5):485–492, September 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s101260000030.

Kono:2000:EST

[134] Tomoya Kono, Masahiro Sakai, and Scott E. LaPatra. Expressed sequence tag analysis of kidney and gill tissues from rainbow trout (Oncorhynchus mykiss) infected with infectious hematopoietic necrosis virus. Marine Biotechnology, 2(5):493–498, September 2000. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s101260000032.

Chu:2000:TMM

[135] Ka Hou Chu, Shun Hang Wong, and Patrick S. C. Leung. Tropomyosin is the major mollusk allergen: Reverse transcriptase polymerase chain reaction, expression and IgE reactivity. Marine Biotechnology, 2(5): 499–509, September 2000. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101260000035.

Yu:2000:SVR

[136] Eizadora T. Yu, Ma. Antonette Juinio-Meñez, and Virginia D. Monje. Sequence variation in the ribosomal DNA internal transcribed spacer of Tridacna crocea. Marine Biotechnology, 2(6):511–516, November 2000. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s101260000033.

Maxwell:2000:PIM

[137] Michael R. Maxwell, Kendra M. Buresch, and Roger T. Hanlon. Pattern of inheritance of microsatellite loci in the squid Loligo pealeii (Mollusca: Cephalopoda). Marine Biotechnology, 2(6):517–521, November 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s101260000036.

Diaz:2000:ICN

[138] María Piedad Díaz, Steve J. W. Grigson, Chris J. Peppiatt, and J. Grant Burgess. Isolation and characterization of novel hydrocarbon-degrading euryhaline consortia from crude oil and mangrove sediments. Marine Biotechnology, 2(6):522–532, November 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s101260000037.

Downs:2000:MBS

[139] Craig A. Downs, Erich Mueller, Susan Phillips, John E. Fauth, and Cheryl M. Woodley. A molecular biomarker system for assessing the health of coral (Montastraea faveolata) during heat stress. Marine Biotechnology, 2(6):533–544, November 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s101260000038.

OBrien:2000:EGP

[140] Elizabeth K. O’Brien and Bernard M. Degnan. Expression of POU, Sox, and Pax genes in the brain ganglia of the tropical abalone Haliotis asinina. Marine Biotechnology, 2(6):545–557, November 2000. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s101260000039.

Chen:2000:SAP

[141] Chaolun Allen Chen, Carden C. Wallace, Jr-Kai Yu, and Nuwei Vivian Wei. Strategies for amplification by polymerase chain reaction of the complete sequence of the gene encoding nuclear large subunit ribosomal RNA in corals. Marine Biotechnology, 2(6):558–570, November 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s101260000040.

Katagiri:2000:GBA

[142] Takayuki Katagiri, Shuichi Asakawa, Ikuo Hirono, Takashi Aoki, and Nobuyoshi Shimizu. Genomic bacterial artificial chromosome library of the Japanese flounder Paralichthys olivaceus. Marine Biotechnology, 2(6):571–576, November 2000. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s101260000041.

Lin:2000:MCA

[143] Senjie Lin, Erika Magaletti, and Edward J. Carpenter. Molecular cloning and antiserum development of cyclin box in the brown tide alga Aureococcus anophagefferens. Marine Biotechnology, 2(6):577– 586, November 2000. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101260000044.

Beacham:2000:MDP

[144] Terry D. Beacham, Susan Pollard, and Khai D. Le. Microsatellite DNA population structure and stock identification of steelhead trout (Oncorhynchus mykiss) in the Nass and Skeena Rivers in Northern British Columbia. Marine Biotechnology, 2(6):587–600, November 2000. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s101260000045.

Yue:2001:CMG

[145] Genhua Yue, Yang Li, and Laszlo Orban. Characterization of microsatellites in the IGF-2 and GH genes of Asian seabass (Lates calcarifer). Marine Biotechnology, 3(1):1–3, January 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s101260000043.

Patruno:2001:CUC

[146] M. Patruno, M. C. Thorndyke, M. D. Candia Carnevali, F. Bonasoro, and P. Beesley. Changes in Ubiquitin conjugates and Hsp72 levels during arm regeneration in echinoderms. Marine Biotechnology, 3 (1):4–15, January 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101260000018.

Dyer:2001:QTP

[147] Anthony Dyer, Kathleen Soole, and George Matsumoto. Quantitative TaqMan PCR without a real-time thermal cycler: an assay for fish insulin-like growth factor I messenger RNA. Marine Biotechnology, 3 (1):16–21, January 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101260000029.

Chow:2001:DEM

[148] S. Chow, V. P. Scholey, A. Nakazawa, D. Margulies, J. B. Wexler, R. J. Olson, and K. Hazama. Direct evidence for Mendelian inheritance of the variations in the ribosomal protein gene introns in yellowfin tuna (Thunnus albacares). Marine Biotechnology, 3(1):22–26, January 2001. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s101260000031.

Singh:2001:FGF

[149] Natalia N. Singh, Kay Fischer, Olaf Hedstrom, and David W. Barnes. Fibroblast growth factor inhibits expression of neural markers in cultures of zebrafish early embryo cells. Marine Biotechnology, 3(1): 27–35, January 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101260000034.

Zentz:2001:CQC

[150] Frédéric Zentz, Laurent Bédouet, Maria José Almeida, Christian Milet, Evelyne Lopez, and Michel Giraud. Characterization and quantification of chitosan extracted from nacre of the abalone Haliotis tuberculata and the oyster Pinctada maxima. Marine Biotechnology, 3(1): 36–44, January 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101260000042.

Sekino:2001:AMD

[151] Masashi Sekino, Norimasa Takagi, Motoyuki Hara, and Hideaki Takahashi. Analysis of microsatellite DNA polymorphisms in rockfish Sebastes thompsoni and application to population genetics studies. Marine Biotechnology, 3(1):45–52, January 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s101260000046.

Imahara:2001:RSN

[152] Jungo Imahara, Mika Tokumoto, Yoshitaka Nagahama, Katsutoshi Ishikawa, and Toshinobu Tokumoto. Reconstitution of sperm nuclei of zebrafish (Danio rerio) in Xenopus egg extracts. Marine Biotechnology, 3(1):53–57, January 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101260000047.

Kano:2001:GRV

[153] Shungo Kano, Shota Chiba, and Noriyuki Satoh. Genetic relatedness and variability in inbred and wild populations of the solitary ascidian Ciona intestinalis revealed by arbitrarily primed polymerase chain reaction. Marine Biotechnology, 3(1):58–67, January 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s101260000048.

Yokochi:2001:NIM

[154] Toshihiro Yokochi, Toro Nakahara, Takanori Higashihara, Masakazu Yamaoka, and Ryuuichirou Kurane. A new isolation method for labyrinthulids using a bacterium, Psychrobacter phenylpyruvicus. Marine Biotechnology, 3(1):68–73, January 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s101260000054.

Dionisio-Sese:2001:PPA

[155] Maribel L. Dionisio-Sese, Tadashi Maruyama, and Shigetoh Miyachi. Photosynthesis of Prochloron as affected by environmental factors. Marine Biotechnology, 3(1):74–79, January 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s101260000062.

Ahsan:2001:MCC

[156] Md. Nazmul Ahsan, Daisuke Funabara, and Shugo Watabe. Molecular cloning and characterization of two isoforms of trypsinogen from anchovy pyloric ceca. Marine Biotechnology, 3(1):80–90, January 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s101260000055.

Ostrander:2001:RGM

[157] Gary K. Ostrander, Jeanette M. Rotchell, Elcin Ulnal, and Rebecca J. Van Beneden. Retinoblastoma gene mutations in chemically induced liver tumor samples of Japanese medaka (Oryzias latipes). Marine Biotechnology, 3(1):S44–S49, January 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-001-0026-1.

Campbell:2001:PIT

[158] C. E. Campbell, P. D. L. Gibbs, and M. C. Schmale. Progression of infection and tumor development in damselfish. Marine Biotechnology, 3(1): S107–S114, January 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-001-00323.

Harshbarger:2001:LKA

[159] John C. Harshbarger and Marilyn S. Slatick. Lesser known aquarium fish tumor models. Marine Biotechnology, 3(1):S115–S129, January 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-001-0033-2.

Atz:2001:AKD

[160] James W. Atz and Steven Kazianis. In appreciation of Klaus D. Kallman. Marine Biotechnology, 3(1):S3–S5, January 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-001-0021-6.

Kirchner:2001:TAR

[161] Jakob M. Kirchner, Vesselina Ivanova, Vladimir N. Noskov, Autum Samson, Jean-Nicolas Volff, Michael A. Resnick, and Ronald B. Walter. Transformation-associated recombination (TAR) cloning of tumorinducing Xmrk2 gene from Xiphophorus maculatus. Marine Biotechnology, 3(1):S168–S176, January 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s1012601-0039-9.

Sarmasik:2001:PTL

[162] Aliye Sarmasik, Thomas T. Chen, C. Z. Chun, In-Kwon Jang, and J. K. Lu. Production of transgenic live-bearing fish and crustaceans with replication-defective pantropic retroviral vectors. Marine Biotechnology, 3(1):S177–S184, January 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-001-0040-3.

Kazianis:2001:GAN

[163] Steven Kazianis, Irma Gimenez-Conti, David Trono, Anthony Pedroza, Lori B. Chovanec, Donald C. Morizot, Rodney S. Nairn, and Ronald B. Walter. Genetic analysis of neoplasia induced by N-nitroso-N-methylurea in Xiphophorus hybrid fish. Marine Biotechnology, 3(1):S37–S43, January 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126001-0025-2.

Paul:2001:RGE

[164] T. A. Paul, J. C. Burns, H. Shike, R. Getchell, P. R. Bowser, K. E. Whitlock, and J. W. Casey. Reporter gene expression in fish following cutaneous infection with pantropic retroviral vectors. Marine Biotechnology, 3(1):S81–S87, January 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-001-0029-y.

Phillips:2001:AFS

[165] Ruth B. Phillips. Application of fluorescence in situ hybridization (FISH) to fish genetics and genome mapping. Marine Biotechnology, 3(1):S145– S152, January 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-001-0036-z.

Shima:2001:MMS

[166] Akihiro Shima and Atsuko Shimada. The medaka as a model for studying germ-cell mutagenesis and genomic instability. Marine Biotechnology, 3(1):S162–S167, January 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-001-0038-X.

Walter:2001:IAF

[167] Ronald B. Walter. Introduction: Aquaria fish models of human disease. Marine Biotechnology, 3(1):S1–S2, January 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-001-0020-7.

Nonaka:2001:CGM

[168] Masaru Nonaka, Kiyoshi Naruse, Megumi Matsuo, and Akihiro Shima. Comparative genomics of medaka: The major histocompatibility complex (MHC). Marine Biotechnology, 3(1):S141–S144, January 2001. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-001-0035-0.

Buck:2001:CCR

[169] Charles Buck, Charles Walsh, Raymond Davis, Araz Toumadje, Kenichi Kusamoto, Angela Helmrich, Christine Chapline, Patricia Mericko, and David Barnes. Cell cultures and retroviral particles from a tumor of a moray eel. Marine Biotechnology, 3(1):S196–S202, January 2001. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-001-0042-1.

Setlow:2001:TUE

[170] Richard B. Setlow and Avril D. Woodhead. Three unique experimental fish stories: Poecilia (the past), Xiphophorus (the present), and Medaka (the future). Marine Biotechnology, 3(1):S17–S23, January 2001. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s1012601-0023-4.

Kazianis:2001:PCS

[171] Steven Kazianis, Richard B. Setlow, John C. Harshbarger, Arvil D. Woodhead, Ronald B. Walter, Rodney S. Nairn, and Irma Gimenez-Conti. A proposed classification scheme for Xiphophorus melanomas based on histopathologic analyses. Marine Biotechnology, 3(1):S100– S106, January 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126001-0031-4.

Oleksiak:2001:UNP

[172] Marjorie F. Oleksiak, Kevin J. Kolell, and Douglas L. Crawford. Utility of natural populations for microarray analyses: Isolation of genes necessary for functional genomic studies. Marine Biotechnology, 3(1): S203–S211, January 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-001-0043-0.

Froschauer:2001:GPM

[173] Alexander Froschauer, Cornelia Körting, Wolfgang Bernhardt, Indrajit Nanda, Michael Schmid, Manfred Schartl, and Jean-Nicolas Volff. Genomic plasticity and melanoma formation in the fish Xiphophorus. Marine Biotechnology, 3(1):S72–S80, January 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-001-0049-7.

Walter:2001:RBE

[174] Ronald B. Walter, Huang-Mo Sung, Rebecca D. Obermoeller, David L. Mitchell, Gabriel W. Intano, and Christi A. Walter. Relative base excision repair in Xiphophorus fish tissue extracts. Marine Biotechnology, 3 (1):S50–S60, January 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-001-0027-0.

Tiersch:2001:CAF

[175] Terrence R. Tiersch. Cryopreservation in aquarium fishes. Marine Biotechnology, 3(1):S212–S223, January 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126001-0044-z.

Clark:2001:UGJ

[176] Melody S. Clark, Sarah F. Smith, and Greg Elgar. Use of the Japanese pufferfish (Fugu rubripes) in comparative genomics. Marine Biotechnology, 3(1):S130–S140, January 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-001-0034-1.

Hahn:2001:DTA

[177] Mark E. Hahn. Dioxin toxicology and the aryl hydrocarbon receptor: Insights from fish and other non-traditional models. Marine Biotechnology, 3(1):S224–S238, January 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-001-0045-Y.

Kallman:2001:HXP

[178] Klaus D. Kallman. How the Xiphophorus problem arrived in San Marcos, Texas. Marine Biotechnology, 3(1):S6–S16, January 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-001-0022-5.

Winn:2001:BPP

[179] Richard N. Winn, Michelle Norris, Stacy Muller, Cecilia Torres, and Kathryn Brayer. Bacteriophage λ and plasmid pUR288 transgenic fish models for detecting In Vivo mutations. Marine Biotechnology, 3(1):

S185–S195, January 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-001-0041-2.

Mitchell:2001:RUI

[180] David L. Mitchell, Jarah A. Meador, Michelle Byrom, and Ronald B. Walter. Resolution of UV-induced DNA damage in Xiphophorus fishes. Marine Biotechnology, 3(1):S61–S71, January 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/PL00021684.

Nairn:2001:AFM

[181] Rodney S. Nairn, Michael C. Schmale, John Stegman, Richard N. Winn, and Ronald B. Walter. Aquaria fish models of human disease: Reports and recommendations from the working groups. Marine Biotechnology, 3(1):S249–S258, January 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-001-0047-9.

Morizot:2001:XGL

[182] Donald C. Morizot, Rodney S. Nairn, Padmaja Simhambhatla, Luis Della Coletta, David Trono, Lori Chovanec, Ronald B. Walter, and Steven Kazianis. Xiphophorus genetic linkage map: Beginnings of comparative gene mapping in fishes. Marine Biotechnology, 3(1):S153– S161, January 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126001-0037-Y.

Quackenbush:2001:GRT

[183] Sandra L. Quackenbush, Joel Rovnak, Rufina N. Casey, Thomas A. Paul, Paul R. Bowser, Claudia Sutton, and James W. Casey. Genetic relationship of tumor-associated piscine retroviruses. Marine Biotechnology, 3 (1):S88–S99, January 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-01-0030-5.

Hawkins:2001:FRC

[184] William E. Hawkins, Melody S. Clark, Akihiro Shima, Ronald B. Walter, Richard N. Winn, and Monte Westerfield. Four resource centers for fishes: Specifies, stocks, and services. Marine Biotechnology, 3(1): S239–S248, January 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-001-0046-X.

Nairn:2001:GAS

[185] Rodney S. Nairn, Steven Kazianis, Luis Della Coletta, David Trono, Andrew P. Butler, Ronald B. Walter, and Donald C. Morizot. Genetic analysis of susceptibility to spontaneous and UV-induced carcinogenesis in Xiphophorus hybrid fish. Marine Biotechnology, 3(1): S24–S36, January 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s1012601-0004-7.

Mendonca:2001:ICF

[186] Holly Mendonca, Jerry Smith, and Chris Brinegar. Isolation and characterization of four microsatellite loci in the tidewater goby (Eucyclogobius newberryi). Marine Biotechnology, 3(2):91–95, March 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s101260000051.

Kurusu:2001:GTS

[187] Yasurou Kurusu, Satoshi Yoshimura, Mitsuko Tanaka, Takamichi Nakamura, Akihiko Maruyama, and Takanori Higashihara. Genetic transformation system for a psychrotrophic deep-sea bacterium: Isolation and characterization of a psychrotrophic plasmid. Marine Biotechnology, 3 (2):96–99, March 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101260000067.

Mo:2001:SRF

[188] Caiqing Mo and Baruch Rinkevich. A simple, reliable, and fast protocol for thraustochytrid DNA extraction. Marine Biotechnology, 3(2): 100–102, March 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101260000069.

McKenna:2001:FIP

[189] Bronagh M. McKenna, Richard M. Fitzpatrick, Kerry V. Phenix, Daniel Todd, Laurence M. Vaughan, and Gregory J. Atkins. Formation of infectious pancreatic necrosis virus-like particles following expression of segment A by recombinant semliki forest virus. Marine Biotechnology, 3 (2):103–110, March 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101260000049.

Leonard:2001:NSM

[190] Jill B. K. Leonard, Geoffrey C. Waldbieser, and Jeffrey T. Silverstein. Neuropeptide Y sequence and messenger RNA distribution in channel catfish (Ictalurus punctatus). Marine Biotechnology, 3(2):111–118, March 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s101260000050.

Sudha:2001:MTT

[191] Puttur Mudumana Sudha, Sharon Low, Jimmy Kwang, and Zhiyuan Gong. Multiple tissue transformation in adult zebrafish by gene gun bombardment and muscular injection of naked DNA. Marine Biotechnology, 3(2):119–125, March 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101260000056.

Sode:2001:SCF

[192] Koji Sode, Fumimasa Ishimura, and Wakako Tsugawa. Screening and characterization of fructosyl-valine–utilizing marine microorganisms. Marine Biotechnology, 3(2):126–132, March 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s101260000065.

Klinbunga:2001:GDM

[193] S. Klinbunga, P. Ampayup, A. Tassanakajon, P. Jarayabhand, and W. Yoosukh. Genetic diversity and molecular markers of cupped oysters (genera Crassostrea, Saccostrea, and Striostrea) in Thailand revealed by randomly amplified polymorphic DNA analysis. Marine Biotechnology, 3(2):133–144, March 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101260000057.

Richardson:2001:MCG

[194] Michael P. Richardson, Boon Hui Tay, Boon Yeong Goh, Byrappa Venkatesh, and Sydney Brenner. Molecular cloning and genomic structure of a gene encoding Interferon regulatory factor in the pufferfish (Fugu rubripes). Marine Biotechnology, 3(2):145–151, March 2001. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s101260000066.

Godhe:2001:PCR

[195] Anna Godhe, S. K. Otta, Ann-Sofi Rehnstam-Holm, Indrani Karunasagar, and Iddya Karunasagar. Polymerase chain reaction in detection of Gymnodinium mikimotoi and Alexandrium minutum in field samples from Southwest India. Marine Biotechnology, 3(2):152–162, March 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s101260000052.

Chang:2001:SAR

[196] Yun-Shiang Chang, Shao-En Peng, Han-Ching Wang, Hui-Chen Hsu, Ching-Hui Ho, Chung-Hsiung Wang, Sho-Ya Wang, Chu-Fang Lo, and Guang-Hsiung Kou. Sequencing and amplified restriction fragment length polymorphism analysis of ribonucleotide reductase large subunit gene of the white spot syndrome virus in blue crab (Callinectes sapidus) from American coastal waters. Marine Biotechnology, 3(2): 163–171, March 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101260000058.

Bonnal:2001:PDW

[197] C. Bonnal, F. Raynaud, C. Astier, M. C. Lebart, A. Marcilhac, D. Coves, G. Corraze, A. Gélineau, J. Fleurence, C. Roustan, and Y. Benyamin. Postmortem degradation of white fish skeletal muscle (sea bass, Dicentrarchus labrax): Fat diet effects on in situ dystrophin proteolysis during the prerigor stage. Marine Biotechnology, 3(2):172–180, March 2001. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s101260000059.

Rougeaux:2001:MCE

[198] Hélène Rougeaux, Muriel Guezennec, Lydie Mao Che, Claude Payri, Eric Deslandes, and Jean Guezennec. Microbial communities and exopolysaccharides from Polynesian mats. Marine Biotechnology, 3(2): 181–187, March 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101260000063.

Pitkanen:2001:CTC

[199] Tiina I. Pitkänen, Sheila Q. Xie, Aleksei Krasnov, Paul S. Mason, Hannu Mölsä, and Neil C. Stickland. Changes in tissue cellularity are associated with growth enhancement in genetically modified Arctic char (Salvelinus alpinus L.) carrying recombinant growth hormone gene. Marine Biotechnology, 3(2):188–197, March 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s101260000060.

Yue:2001:RID

[200] Gen Hua Yue and Laszlo Orban. Rapid isolation of DNA from fresh and preserved fish scales for polymerase chain reaction. Marine Biotechnology, 3(3):199–204, May 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.com/

article/10.1007/s10126-001-0010-9.

Tang:2001:IES

[201] Yanlin Tang, Brian S. Shepherd, Amy J. Nichols, Rex Dunham, and Thomas T. Chen. Influence of environmental salinity on messenger RNA levels of growth hormone, prolactin, and somatolactin in pituitary of the channel catfish (Ictalurus punctatus). Marine Biotechnology, 3 (3):205–217, May 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101260000061.

Mitsuo:2001:CSP

[202] Ryoichi Mitsuo, Mamoru Sato, and Takao Itakura. Cloning, sequencing, and phylogenetic analysis of complementary DNA of novel cytochrome P-450 CYP1A in Japanese eel (Anguilla japonica). Marine Biotechnology, 3(3):218–223, May 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101260000070.

Maccatrozzo:2001:CMG

[203] Lisa Maccatrozzo, Luca Bargelloni, Giuseppe Radaelli, Francesco Mascarello, and Tomaso Patarnello. Characterization of the myostatin gene in the gilthead seabream (Sparus aurata): Sequence, genomic structure, and expression pattern. Marine Biotechnology, 3(3):224–230, May 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s101260000064.

Pank:2001:RSI

[204] Melissa Pank, Michael Stanhope, Lisa Natanson, Nancy Kohler, and Mahmood Shivji. Rapid and simultaneous identification of body parts from the morphologically similar sharks Carcharhinus obscurus and Carcharhinus plumbeus (Carcharhinidae) using multiplex PCR. Marine Biotechnology, 3(3):231–240, May 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s101260000071.

Karsi:2001:EIS

[205] Attila Karsi, Boaz Moav, Perry Hackett, and Zhanjiang Liu. Effects of insert size on transposition efficiency of the Sleeping Beauty transposon in mouse cells. Marine Biotechnology, 3(3):241–245, May 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s101260000072.

Pin:2001:GDO

[206] Leaw Chui Pin, Lim Po Teen, Asmat Ahmad, and Gires Usup. Genetic diversity of Ostreopsis ovata (Dinophyceae) from Malaysia. Marine Biotechnology, 3(3):246–255, May 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s101260000073.

Fabregas:2001:GRM

[207] Jaime Fábregas, Ana Otero, Adolfo Domínguez, and Manuel Patiño. Growth rate of the microalga Tetraselmis suecica changes the biochemical composition of Artemia species. Marine Biotechnology, 3(3): 256–263, May 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s101260000074.

Mamuris:2001:RAM

[208] Zissis Mamuris, Costas Stamatis, Katerina A. Moutou, Apostolos P. Apostolidis, and Costas Triantaphyllidis. RFLP analysis of mitochondrial DNA to evaluate genetic variation in striped red mullet (Mullus surmuletus L.) and red mullet (Mullus barbatus L.) populations. Marine Biotechnology, 3(3):264–274, May 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s101260000075.

Tseng:2001:PGG

[209] Mei-Chen Tseng, Chaolun Allen Chen, Hsiao-Wei Kao, Wann-Nian Tzeng, and Sin-Che Lee. Polymorphisms of GA/GT microsatellite loci from Anguilla japonica. Marine Biotechnology, 3(3):275–280, May 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s101260000076.

DeRosa:2001:DPC

[210] Salvatore De Rosa, Salvatore De Caro, Giuseppina Tommonaro, Krasimir Slantchev, Kamen Stefanov, and Simeon Popov. Development in a primary cell culture of the marine sponge Ircinia muscarum and analysis of the polar compounds. Marine Biotechnology, 3(3):281–286, May 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-001-0001-x.

Han:2001:AOM

[211] Kaiping Han, Sarah A. Woodin, David E. Lincoln, Kevin T. Fielman, and Bert Ely. Amphitrite ornata, a marine worm, contains two dehaloperoxidase genes. Marine Biotechnology, 3(3):287–292, May 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-001-0003-8.

Gold:2001:MVA

[212] John R. Gold, Elena Pak, and Linda R. Richardson. Microsatellite variation among red snapper (Lutjanus campechanus) from the Gulf of Mexico. Marine Biotechnology, 3(3):293–304, May 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-001-0004-7.

Suetsuna:2001:IAP

[213] Kunio Suetsuna and Jiun-Rong Chen. Identification of antihypertensive peptides from peptic digest of two microalgae, Chlorella vulgaris and Spirulina platensis. Marine Biotechnology, 3(4):305–309, July 2001. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-001-0012-7.

Sekino:2001:ICM

[214] Masashi Sekino and Motoyuki Hara. Inheritance characteristics of microsatellite DNA loci in experimental families of Japanese flounder Paralichthys olivaceus. Marine Biotechnology, 3(4):310–315, July 2001. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-001-0052-z.

Ueki:2001:SEC

[215] Tatsuya Ueki, Taro Uyama, Kan Kanamori, and Hitoshi Michibata. Subunit C of the vacuolar-type ATPase from the vanadium-rich ascidian Ascidia sydneiensis samea rescued the pH sensitivity of yeast vma5 mutants. Marine Biotechnology, 3(4):316–321, July 2001. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s1012601-0054-x.

Buchanan:2001:TGE

[216] John T. Buchanan, Amy D. Nickens, Richard K. Cooper, and Terrence R. Tiersch. Transfection of Eastern oyster (Crassotrea virginica) embryos. Marine Biotechnology, 3(4):322–335, July 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-001-0002-9.

Jasiecki:2001:CUB

[217] Jacek Jasiecki, Agata Czyż, Magdalena Gabig, and Grzegorz Wȩgrzyn. Construction and use of a broad-host-range plasmid expressing the lamB gene for utilization of bacteriophage λ vectors in the marine bacterium Vibrio harveyi. Marine Biotechnology, 3(4):336–345, July 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-001-0005-6.

Klein:2001:IBS

[218] Gracjana Klein, Ewa Laskowska, Alina Taylor, and Barbara Lipińska. IbpA/B small heat-shock protein of marine bacterium Vibrio harveyi binds to proteins aggregated in a cell during heat shock. Marine Biotechnology, 3(4):346–354, July 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.com/

article/10.1007/s10126001-0009-2.

Chu:2001:FIT

[219] K. H. Chu, C. P. Li, and H. Y. Ho. The first internal transcribed spacer (ITS-1) of ribosomal DNA as a molecular marker for phylogenetic and population analyses in crustacea. Marine Biotechnology, 3 (4):355–361, July 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126001-0014-5.

Sarashina:2001:CPS

[220] Isao Sarashina and Kazuyoshi Endo. The complete primary structure of molluscan shell protein 1 (MSP-1), an acidic glycoprotein in the shell matrix of the scallop Patinopecten yessoensis. Marine Biotechnology, 3 (4):362–369, July 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-001-0013-6.

Sato:2001:MCM

[221] Maremi Sato and Koji Nagashima. Molecular characterization of a mitochondrial DNA segment from the Japanese scallop (Patinopecten yessoensis): Demonstration of a region showing sequence polymorphism in the population. Marine Biotechnology, 3(4):370–379, July 2001. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126001-0015-4.

Downs:2001:AHG

[222] Craig A. Downs, John E. Fauth, and Cheryl M. Woodley. Assessing the health of grass shrimp (Palaeomonetes pugio) exposed to natural and anthropogenic stressors: a molecular biomarker system. Marine Biotechnology, 3(4):380–397, July 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-001-0008-3.

Song:2001:MPS

[223] Linsheng Song, Baozhong Liu, Jianhai Xiang, and Pei-Yuan Qian. Molecular phylogeny and species identification of pufferfish of the genus Takifugu (Tetraodontiformes, Tetraodontidae). Marine Biotechnology, 3 (4):398–406, July 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-001-0006-5.

Holland:2001:IBM

[224] Brenden S. Holland. Invasion without a bottleneck: Microsatellite variation in natural and invasive populations of the brown mussel Perna perna (L.). Marine Biotechnology, 3(5):407–415, September 2001. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s1012601-0060-Z.

Sojka:2001:EAH

[225] Katherine Marks Sojka and Richard S. Pollenz. Expression of aryl hydrocarbon receptor nuclear translocator (ARNT) isoforms in juvenile and adult rainbow trout tissues. Marine Biotechnology, 3(5):416– 427, September 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s1012601-0011-8.

Klinbunga:2001:GHG

[226] S. Klinbunga, D. Siludjai, W. Wudthijinda, A. Tassanakajon, P. Jarayabhand, and P. Menasveta. Genetic heterogeneity of the giant tiger shrimp (Penaeus monodon) in Thailand revealed by RAPD and mitochondrial DNA RFLP analyses. Marine Biotechnology, 3(5):428– 438, September 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-001-0055-9.

Lewis:2001:SSC

[227] Tom E. Lewis, Peter D. Nichols, and Tom A. McMeekin. Sterol and squalene content of a docosahexaenoic-acid-producing thraustochytrid: Influence of culture age, temperature, and dissolved oxygen. Marine Biotechnology, 3(5):439–447, September 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-001-0016-3.

Huvet:2001:MAH

[228] A. Huvet, K. Balabaud, N. Bierne, and P. Boudry. Microsatellite analysis of 6-hour-old embryos reveals no preferential intraspecific fertilization between cupped oysters Crassostrea gigas and Crassostrea angulata. Marine Biotechnology, 3(5):448–453, September 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-001-0017-2.

Radjasa:2001:CPB

[229] Ocky Karna Radjasa, Hidetoshi Urakawa, Kumiko Kita-Tsukamoto, and Kouichi Ohwada. Characterization of psychrotrophic bacteria in the surface and deep-sea waters from the northwestern Pacific Ocean based on 16S ribosomal DNA analysis. Marine Biotechnology, 3(5): 454–462, September 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-001-0050-1.

Ma:2001:OAG

[230] P. Ma, B. Sivaloganathan, P. K. Reddy, W. K. Chan, and T. J. Lam. Ontogeny of α-amylase gene expression in sea bass larvae (Lates calcarifer). Marine Biotechnology, 3(5):463–469, September 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-001-0018-1.

Sarmasik:2001:TLB

[231] Aliye Sarmasik, In-Kwon Jang, C. Z. Chun, J. K. Lu, and Thomas T. Chen. Transgenic live–bearing fish and crustaceans produced by transforming immature gonads with replication-defective pantropic retroviral vectors. Marine Biotechnology, 3(5):470–477, September 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126001-0019-0.

Selvamani:2001:MGI

[232] Maria John P. Selvamani, Sandie M. Degnan, and Bernard M. Degnan. Microsatellite genotyping of individual abalone larvae: Parentage assignment in aquaculture. Marine Biotechnology, 3(5):478–485, September 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-

2236 (electronic). URL https://link.springer.com/article/10.
1007/s1012601-0062-X.

Wang:2001:CSS

[233] Yongping Wang, Zhe Xu, and Ximing Guo. A centromeric satellite sequence in the Pacific oyster (Crassostrea gigas Thunberg) identified by fluorescence In Situ hybridization. Marine Biotechnology, 3(5):486– 492, September 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-001-0063-3.

Iwamoto:2001:PCM

[234] Koji Iwamoto, Hideaki Kawanobe, Yoshihiro Shiraiwa, and Tomoyoshi Ikawa. Purification and characterization of mannitol-l-phosphatase in the red alga Caloglossa continua (Ceramiales, Rhodophyta). Marine Biotechnology, 3(5):493–500, September 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-001-0068-4.

Choresh:2001:KHS

[235] Omer Choresh, Eliora Ron, and Yossi Loya. The 60-kDa heat shock protein (HSP60) of the sea anemone Anemonia viridis: a potential early warning system for environmental changes. Marine Biotechnology, 3(5): 501–508, September 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-001-0007-4.

Noell:2001:MDG

[236] Craig J. Noell, S. Donnellan, R. Foster, and L. Haigh. Molecular discrimination of garfish Hyporhamphus (Beloniformes) larvae in Southern Australian waters. Marine Biotechnology, 3(6):509–514, November 2001. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-001-0059-5.

Montsant:2001:PNB

[237] Anton Montsant, Aliza Zarka, and Sammy Boussiba. Presence of a nonhydrolyzable biopolymer in the cell wall of vegetative cells and astaxanthin-rich cysts of Haematococcus pluvialis (Chlorophyceae). Marine Biotechnology, 3(6):515–521, November 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s1012601-0051-0.

Karlsson:2001:GTZ

[238] Johnny Karlsson, Jonas von Hofsten, and Per-Erik Olsson. Generating transparent zebrafish: a refined method to improve detection of gene expression during embryonic development. Marine Biotechnology, 3(6): 522–527, November 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s1012601-0053-4.

McDaniel:2001:EMB

[239] L. McDaniel, D. W. Griffin, J. Crespo-Gomez, M. R. McLaughlin, and J. H. Paul. Evaluation of marine bacterial lysogens for development of a marine prophage induction assay. Marine Biotechnology, 3(6):528– 535, November 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-001-0056-8.

Kinoshita:2001:TMO

[240] Masato Kinoshita, Takami Morita, Haruhiko Toyohara, Takashi Hirata, Morihiko Sakaguchi, Masao Ono, Koji Inoue, Yuko Wakamatsu, and Kenjiro Ozato. Transgenic medaka overexpressing a melaninconcentrating hormone exhibit lightened body color but no remarkable abnormality. Marine Biotechnology, 3(6):536–543, November 2001. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-001-0061-Y.

Osinga:2001:DVS

[241] Ronald Osinga, Roelco Kleijn, Esther Groenendijk, Patrick Niesink, Johannes Tramper, and Rene H. Wijffels. Development of In Vivo sponge cultures: Particle feeding by the tropical sponge Pseudosuberites aff. andrewsi. Marine Biotechnology, 3(6):544–554, November 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s1012601-0078-2.

Holm:2001:EMB

[242] Lars-Erik Holm, Volker Loeschcke, and Christian Bendixen. Elucidation of the molecular basis of a null allele in a rainbow trout microsatellite. Marine Biotechnology, 3(6):555–560, November 2001. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-001-0057-7.

Vincent:2001:DVE

[243] Samuel G. Prakash Vincent, Rainer Keller, and T. Subramoniam. Development of Vitellogenin–ELISA, an In Vivo bioassay, and identification of two vitellogenesis-inhibiting hormones of the tiger shrimp Penaeus monodon. Marine Biotechnology, 3(6):561–571, November 2001. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s1012601-0066-6.

Sekino:2001:AMM

[244] Masashi Sekino and Motoyuki Hara. Application of microsatellite markers to population genetics studies of Japanese flounder Paralichthys olivaceus. Marine Biotechnology, 3(6):572–589, November 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-001-0064-8.

Liu:2001:CCF

[245] Zhanjiang Liu, Soonhag Kim, and Attila Karsi. Channel catfish folliclestimulating hormone and luteinizing hormone: Complementary DNA cloning and expression during ovulation. Marine Biotechnology, 3(6): 590–599, November 2001. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s1012601-0067-5.

Webster:2001:DPA

[246] Nicole S. Webster, Joy E. M. Watts, and Russell T. Hill. Detection and phylogenetic analysis of novel crenarchaeote and euryarchaeote 16S ribosomal RNA gene sequences from a Great Barrier Reef sponge. Marine Biotechnology, 3(6):600–608, November 2001. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-001-0065-7.

Lee:2002:CCR

[247] Jae-Seong Lee and Yong-Sung Lee. Cloning and characterization of replication protein A p32 complementary DNA in zebrafish (Danio rerio). Marine Biotechnology, 4(1):1–5, January 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-001-0077-3.

Koga:2002:GTC

[248] Akihiko Koga, Hiroshi Hori, and Mitsuru Sakaizumi. Gene transfer and cloning of flanking chromosomal regions using the medaka fish Tol2 transposable element. Marine Biotechnology, 4(1):6–11, January 2002. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-001-0086-2.

Rexroad:2002:DRT

[249] C. E. Rexroad III, R. L. Coleman, and J. Killefer. Development of rainbow trout microsatellite markers from repeat enriched libraries. Marine Biotechnology, 4(1):12–16, January 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-001-0058-6.

Chen:2002:NSR

[250] Chaolun Allen Chen, Chang-Po Chen, Tung-Yung Fan, Jr-Kai Yu, and Hwey-Lian Hsieh. Nucleotide sequences of ribosomal internal transcribed spacers and their utility in distinguishing closely related Perinereis polychaets (Annelida; Polychaeta; Nereididae). Marine Biotechnology, 4(1): 17–29, January 2002. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-001-0069-3.

Kiesling:2002:RIA

[251] Traci L. Kiesling, Eric Wilkinson, Jean Rabalais, Peter B. Ortner, Mead M. McCabe, and Jack W. Fell. Rapid identification of adult and naupliar stages of copepods using DNA hybridization methodology. Marine Biotechnology, 4(1):30–39, January 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/PL00021689.

Ravago:2002:LSV

[252] Rachel G. Ravago, Virginia D. Monje, and Marie Antonette Juinio-Meñez. Length and sequence variability in mitochondrial control region of the milkfish, Chanos chanos. Marine Biotechnology, 4(1): 40–50, January 2002. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-001-0076-4.

Ladrat:2002:MCS

[253] Christine Ladrat, Véronique Verrez-Bagnis, Joël Noël, and Joël Fleurence. Milli-calpain from sea bass (Dicentrarchus labrax) white muscle: Purification, characterization of its activity and activation In Vitro. Marine Biotechnology, 4(1):51–62, January 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s1012601-0075-5.

Kim:2002:SIF

[254] Dae-Hyun Kim, Young Tae Kim, Jung Jong Cho, Jin-Hee Bae, Sung-Bum Hur, Inhwan Hwang, and Tae-Jin Choi. Stable integration and functional expression of flounder growth hormone gene in transformed microalga, Chlorella ellipsoidea. Marine Biotechnology, 4(1):63–73, January 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s1012601-0070-x.

Alonso:2002:SSL

[255] Marta Alonso and Jo-Ann Leong. Suppressive subtraction libraries to identify Interferon-inducible genes in fish. Marine Biotechnology, 4(1): 74–80, January 2002. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-001-0001-9.

Jenny:2002:PIS

[256] Matthew J. Jenny, Amy H. Ringwood, Eric R. Lacy, Alan J. Lewitus, Jason W. Kempton, Paul S. Gross, Gregory W. Warr, and Robert W. Chapman. Potential indicators of stress response identified by expressed sequence tag analysis of hemocytes and embryos from the American oyster, Crassostrea virginica. Marine Biotechnology, 4 (1):81–93, January 2002. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-001-0072-8.

Takahashi:2002:DCS

[257] Jun ichiro Takahashi, Shoko Fujiwara, Masami Kikyo, Yasutaka Hirokawa, and Mikio Tsuzuki. Discrimination of the cell surface of the coccolithophorid Pleurochrysis haptonemofera from light scattering and fluorescence after fluorescein-isothiocyanate-labeled lectin staining measured by flow cytometry. Marine Biotechnology, 4(1):94–101, January 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-001-0083-5.

Fukushima:2002:ICN

[258] Mari Fukushima, Noriaki Ozaki, Hiroyuki Ikeda, Keiko Furihata, Yoichi Hayakawa, Shohei Sakuda, and Hiromichi Nagasawa. Isolation and characterization of 2-nitroimidazole produced by Streptomyces species as an inhibitor of both carbonic anhydrase and shell formation in the barnacle Balanus amphitrite. Marine Biotechnology, 4(2):103–110, March 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-001-0091-5.

Bhosale:2002:APS

[259] S. H. Bhosale, V. L. Nagle, and T. G. Jagtap. Antifouling potential of some marine organisms from India against species of Bacillus and Pseudomonas. Marine Biotechnology, 4(2):111–118, March 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-001-0087-1.

Murgia:2002:GIG

[260] Rosalba Murgia, Gabriella Tola, Simon N. Archer, Silvana Vallerga, and Joe Hirano. Genetic identification of grey mullet species (Mugilidae) by analysis of mitochondrial DNA sequence: Application to identify the origin of processed ovary products (Bottarga). Marine Biotechnology, 4 (2):119–126, March 2002. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-001-0074-6.

Kawaguchi:2002:SMI

[261] Tomohiro Kawaguchi and Alan W. Decho. In situ microspatial imaging using two-photon and confocal laser scanning microscopy of bacteria and extracellular polymeric secretions (EPS) within marine stromatolites. Marine Biotechnology, 4(2):127–131, March 2002. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-001-0073-7.

Krungkasem:2002:ITD

[262] Chatchadaporn Krungkasem, Tsuyoshi Ohira, Wei-Jun Yang, Rosman Abdullah, Hiromichi Nagasawa, and Katsumi Aida. Identification of two distinct molt-inhibiting hormone-related peptides from the giant tiger prawn Penaeus monodon. Marine Biotechnology, 4(2):132–140, March 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-001-0079-1.

Han:2002:UAA

[263] Kaiping Han and Bert Ely. Use of AFLP analyses to assess genetic variation in Morone and Thunnus species. Marine Biotechnology, 4(2): 141–145, March 2002. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-001-0080-8.

Wan:2002:GTC

[264] Haiyan Wan, Jiangyan He, Bensheng Ju, Tie Yan, Toong Jin Lam, and Zhiyuan Gong. Generation of two-color transgenic zebrafish using the green and red fluorescent protein reporter genes gfp and rfp. Marine Biotechnology, 4(2):146–154, March 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-001-0085-3.

Xia:2002:BAM

[265] Ning-Shao Xia, Wen-Xin Luo, Jun Zhang, Xiao-Yan Xie, Hai-Jie Yang, Shao-Wei Li, Ming Chen, and Mun-Hon Ng. Bioluminescence of Aequorea macrodactyla, a common jellyfish species in the East China Sea. Marine Biotechnology, 4(2):155–162, March 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-001-0081-7.

Tinti:2002:MDS

[266] Fausto Tinti, Cesare Di Nunno, Ilaria Guarniero, Monia Talenti, Stefano Tommasini, Elena Fabbri, and Corrado Piccinetti. Mitochondrial DNA sequence variation suggests the lack of genetic heterogeneity in the Adriatic and Ionian stocks of Sardina pilchardus. Marine Biotechnology, 4 (2):163–172, March 2002. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-002-0003-3.

Ohtsuka:2002:RSN

[267] Masato Ohtsuka, Natsuko Kikuchi, Masahiro Nogami, Hidetoshi Inoko, Kenjiro Ozato, and Minoru Kimura. Rapid screening of a novel arrayed medaka (Oryzias latipes) cosmid library. Marine Biotechnology, 4(2): 173–178, March 2002. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-001-0088-0.

Kalinovskaya:2002:CSL

[268] Natalie I. Kalinovskaya, Tatyana A. Kuznetsova, Elena P. Ivanova, Ludmila A. Romanenko, Valery G. Voinov, Felix Huth, and Hartmut Laatsch. Characterization of surfactin-like cyclic depsipeptides synthesized by Bacillus pumilus from ascidian Halocynthia aurantium. Marine Biotechnology, 4(2):179–188, March 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-001-0084-4.

Chen:2002:SRA

[269] Chih-Yu Chen and Hong-Nong Chou. Screening of red algae filaments as a potential alternative source of eicosapentaenoic acid. Marine Biotechnology, 4(2):189–192, March 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-002-0002-4.

Deane:2002:CSA

[270] Eddie E. Deane, Scott P. Kelly, James C. Y. Luk, and Norman Y. S. Woo. Chronic salinity adaptation modulates hepatic heat shock protein and insulin-like growth factor I expression in black sea bream. Marine Biotechnology, 4(2):193–205, March 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/PL00021690.

Anonymous:2002:Ea

[271] Anonymous. Errata. Marine Biotechnology, 4(2):206, March 2002. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-002-0026-9.

Anonymous:2002:MGU

[272] Anonymous. A message from the U.S. Editor: Fish as models for basic research and biotechnological application. Marine Biotechnology, 4(3):207, June 2002. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/PL00021693.

Gomez:2002:CAS

[273] Ana Gómez, Claudia Wellbrock, and Manfred Schartl. Constitutive activation and specific signaling of the xmrk receptor in xiphophorus melanoma. Marine Biotechnology, 4(3):208–217, June 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/PL00021691.

Cheng:2002:DTE

[274] Ronshan Cheng, Kay-Min Chang, and Jen-Leih Wu. Different temporal expressions of tilapia (Oreochromis mossambicus) insulin-like growth factor-i and IGF binding protein-3 after growth hormone induction. Marine Biotechnology, 4(3):218–225, June 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-002-0014-0.

Cheng:2002:GEI

[275] Ronshan Cheng and Jen-Leih Wu. Gene expression of insulin-like growth factor-i receptor and p53 suppressor during zebrafish (Danio rerio) embryogenesis. Marine Biotechnology, 4(3):226–235, June 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-002-0015-z.

Huang:2002:CEE

[276] Yu-Shan Huang, Wen-Shiun Yueh, Jing-Duan Huang, Jin-Lien Du, Lian-Tien Sun, Yoshitaka Nagahama, and Ching-Fong Chang. Cloning and expression of estrogen receptors in the protandrous black porgy (Acanthopagrus schlegeli): Implications of sex change mechanism. Marine Biotechnology, 4(3):236–246, June 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-002-0019-8.

Wang:2002:FAP

[277] Tzu-Ming Wang, Yau-Hung Chen, Chia-Feng Liu, and Huai-Jen Tsai. Functional analysis of the proximal promoter regions of fish rhodopsin and myf-5 genes using transgenesis. Marine Biotechnology, 4(3):247– 255, June 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-

2236 (electronic). URL https://link.springer.com/article/10.
1007/s10126-002-0016-y.

Boonanuntanasarn:2002:GKR

[278] Surintorn Boonanuntanasarn, Goro Yoshizaki, Yutaka Takeuchi, Tetsuro Morita, and Toshio Takeuchi. Gene knock-down in rainbow trout embryos using antisense morpholino phosphorodiamidate oligonucleotides. Marine Biotechnology, 4(3):256–266, June 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-002-0017-x.

Lee:2002:PGC

[279] K. K. Lee, P. C. Liu, and W. H. Chuang. Pathogenesis of gastroenteritis caused by Vibrio carchariae in cultured marine fish. Marine Biotechnology, 4(3):267–277, June 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-002-0018-9.

Bartlett:2002:CHK

[280] Thomas C. Bartlett, Brandon J. Cuthbertson, Eleanor F. Shepard, Robert W. Chapman, Paul S. Gross, and Gregory W. Warr. Crustins, homologues of an 11.5-kDa antibacterial peptide, from two species of penaeid shrimp, Litopenaeus vannamei and Litopenaeus setiferus. Marine Biotechnology, 4(3):278–293, June 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-002-0020-2.

Chiou:2002:ECB

[281] Peter P. Chiou, Chun-Mean Lin, Luis Perez, and Thomas T. Chen. Effect of cecropin B and a synthetic analogue on propagation of fish viruses In Vitro. Marine Biotechnology, 4(3):294–302, June 2002. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-002-0021-1.

Simon:2002:GTF

[282] Benjamin E. Simon and Jo-Ann C. Leong. Gene transfer to fish cells by attenuated invasive Escherichia coli. Marine Biotechnology, 4(3): 303–309, June 2002. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-002-0022-0.

Sarmasik:2002:PTM

[283] Aliye Sarmasik, Gregory Warr, and Thomas T. Chen. Production of transgenic medaka with increased resistance to bacterial pathogens. Marine Biotechnology, 4(3):310–322, June 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-002-0023-z.

Dunham:2002:SFT

[284] Rex A. Dunham, Nagaraj Chatakondi, Amy Nichols, Thomas T. Chen, D. A. Powers, and Huseyin Kucuktas. Survival of F2 transgenic common carp (Cyprinus carpio) containing pRSVrtGH1 complementary DNA when subjected to low dissolved oxygen. Marine Biotechnology, 4(3): 323–327, June 2002. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-002-0025-x.

Lu:2002:PTS

[285] Jenn-Kan Lu, Bo-Hua Fu, Jen-Leh Wu, and Thomas T. Chen. Production of transgenic silver sea bream (Sparus sarba) by different gene transfer methods. Marine Biotechnology, 4(3):328–337, June 2002. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-002-0027-8.

Dunham:2002:EBD

[286] Rex A. Dunham, Gregory W. Warr, Amy Nichols, Patricia L. Duncan, Brad Argue, Darlene Middleton, and Huseyin Kucuktas. Enhanced bacterial disease resistance of transgenic channel catfish Ictalurus punctatus possessing cecropin genes. Marine Biotechnology, 4(3):

338–344, June 2002. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-002-0024-y.

Anonymous:2002:Eb

[287] Anonymous. Erratum. Marine Biotechnology, 4(3):345, June 2002. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/PL00021692.

Jarman:2002:OPP

[288] Simon N. Jarman, Robert D. Ward, and Nicholas G. Elliott. Oligonucleotide primers for PCR amplification of coelomate introns. Marine Biotechnology, 4(4):347–355, September 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-002-0029-6.

Yan:2002:SAI

[289] Liming Yan, Kenneth G. Boyd, and J. Grant Burgess. Surface attachment induced production of antimicrobial compounds by marine epiphytic bacteria using modified roller bottle cultivation. Marine Biotechnology, 4(4):356–366, September 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-002-0041-x.

Khoo:2002:GDW

[290] Gideon Khoo, Kok Fang Lim, Damien K. Y. Gan, Fan Chen, Woon-Khiong Chan, Tit Meng Lim, and Violet P. E. Phang. Genetic diversity within and among feral populations and domesticated strains of the guppy (Poecilia reticulata) in Singapore. Marine Biotechnology, 4(4): 367–378, September 2002. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-002-0007-z.

Feng:2002:GEK

[291] Shin-Huey Feng, Jiann-Horng Leu, Chi-Hwa Yang, Mei-Jane Fang, Chang-Jen Huang, and Pung-Pung Hwang. Gene expression of Na+-K+- ATPase a1 and a3 subunits in gills of the teleost Oreochromis mossambicus, adapted to different environmental salinities. Marine Biotechnology, 4(4):379–391, September 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-002-0006-0.

Gutierrez-Millan:2002:SCR

[292] Luis Enrique Gutiérrez-Millán, Alma Beatriz Peregrino-Uriarte, Rogerio Sotelo-Mundo, Francisco Vargas-Albores, and Gloria Yepiz-Plascencia. Sequence and conservation of a rRNA and tRNAVal mitochondrial gene fragment from Penaeus californiensis and comparison with Penaeus vannamei and Penaeus stylirostris. Marine Biotechnology, 4(4):392– 398, September 2002. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-002-0008-y.

Sakai:2002:MSF

[293] Takeshi Sakai, Hitomi Kimura, and Ikunoshin Kato. A marine strain of flavobacteriaceae utilizes brown seaweed fucoidan. Marine Biotechnology, 4(4):399–405, September 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-002-0032-y.

Machida:2002:CMD

[294] Ryuji J. Machida, Masaki U. Miya, Mutsumi Nishida, and Shuhei Nishida. Complete mitochondrial DNA sequence of Tigriopus japonicus (Crustacea: Copepoda). Marine Biotechnology, 4(4):406–417, September 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-002-0033-x.

OReilly:2002:ICT

[295] Patrick T. O’Reilly, Arran A. McPherson, Ellen Kenchington, Chris Taggart, Mathew W. Jones, and Paul Bentzen. Isolation and characterization of tetranucleotide microsatellites from Atlantic haddock (Melanogrammus aeglefinus). Marine Biotechnology, 4(4):418–422, September 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-

2236 (electronic). URL https://link.springer.com/article/10.
1007/s10126-002-0010-4.

Dahllof:2002:MOP

[296] Ingela Dahllöf and Staffan Kjelleberg. Multivariate optimization of polymerase chain reaction for microbial community analysis. Marine Biotechnology, 4(4):423–430, September 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-002-0011-3.

Zentz:2002:AAS

[297] Frédéric Zentz, Claire Hellio, Alain Valla, Denis De La Broise, Graham Bremer, and Roger Labia. Antifouling activities of N-substituted imides: Antimicrobial activities and inhibition of Mytilus edulis phenoloxidase. Marine Biotechnology, 4(4):431–440, September 2002. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-002-0004-2.

Sera:2002:IHN

[298] Yutaka Sera, Kyoko Adachi, Kiyonaga Fujii, and Yoshikazu Shizuri. Isolation of haliclonamides: New peptides as antifouling substances from a marine sponge species, haliclona. Marine Biotechnology, 4(5): 441–446, October 2002. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-001-0082-6.

Rajaganapathi:2002:PAH

[299] J. Rajaganapathi, K. Kathiresan, and T. P. Singh. Purification of anti-HIV protein from purple fluid of the sea hare Bursatella leachii de Blainville. Marine Biotechnology, 4(5):447–453, October 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-002-0012-2.

Lopata:2002:DMA

[300] Andreas L. Lopata, Thomas Luijx, Bartha Fenemore, Neville A. Sweijd, and Peter A. Cook. Development of a monoclonal antibody detection assay for species-specific identification of abalone. Marine Biotechnology, 4(5):454–462, October 2002. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-002-0009-x.

Ohira:2002:MCC

[301] Tsuyoshi Ohira, Hiromichi Nagasawa, and Katsumi Aida. Molecular cloning of cDNAs encoding two pigment-dispersing hormones and two corresponding genes from the kuruma prawn (Penaeus japonicus). Marine Biotechnology, 4(5):463–470, October 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-002-0042-9.

Yu:2002:GAP

[302] Hon-Tsen Yu, Yann-Jium Lee, Shiao-Wei Huang, and Tai-Sheng Chiu. Genetic analysis of the populations of Japanese anchovy (Engraulidae:

Engraulis japonicus) using microsatellite DNA. Marine Biotechnology, 4 (5):471–479, October 2002. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-002-0035-8.

Bulgakov:2002:GGD

[303] V. P. Bulgakov, N. A. Odintsova, S. V. Plotnikov, K. V. Kiselev, E. V. Zacharov, and Y. N. Zhuravlev. Gal4-gene-dependent alterations of embryo development and cell growth in primary culture of sea urchins. Marine Biotechnology, 4(5):480–486, October 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-002-0036-7.

Supungul:2002:IIR

[304] Premruethai Supungul, Sirawut Klinbunga, Rath Pichyangkura, Sarawut Jitrapakdee, Ikuo Hirono, Takashi Aoki, and Anchalee Tassanakajon. Identification of immune-related genes in hemocytes of black tiger shrimp (Penaeus monodon). Marine Biotechnology, 4(5):487–494, October 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-002-0043-8.

Lopez-Pinon:2002:IFS

[305] M. J. López-Piñón, A. Insua, and J. Méndez. Identification of four scallop species using PCP and restriction analysis of the ribosomal DNA internal transcribed spacer region. Marine Biotechnology, 4(5): 495–502, October 2002. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-002-0030-0.

Yue:2002:NMG

[306] Gen Yue, Yang Li, Tien Chao, Renee Chou, and Laszlo Orban. Novel microsatellites from Asian sea bass (Lates calcarifer) and their application to broodstock analysis. Marine Biotechnology, 4(5):503–511, October 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-002-0037-6.

Semple:2002:MCC

[307] Jeffrey W. Semple, Howard J. Green, and Patricia M. Schulte. Molecular cloning and characterization of two Na/K-ATPase isoforms in Fundulus heteroclitus. Marine Biotechnology, 4(5):512–519, October 2002. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-002-0031-z.

Cho:2002:SPS

[308] Jung Cho and Young Kim. Sharks: a potential source of antiangiogenic factors and tumor treatments. Marine Biotechnology, 4(6):521– 525, December 2002. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-002-0064-3.

Kocabas:2002:EPC

[309] Arif M. Kocabas, Ping Li, Dongfeng Cao, Attila Karsi, Chongbo He, Andrea Patterson, Zhenlin Ju, Rex A. Dunham, and Zhanjiang Liu. Expression profile of the channel catfish spleen: Analysis of genes involved in immune functions. Marine Biotechnology, 4(6):526–536, December 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-002-0067-0.

Tan:2002:CMR

[310] Xungang Tan, Lily Hoang, and Shao Du. Characterization of muscleregulatory genes, Myf5 and Myogenin, from striped bass and promoter analysis of muscle-specific expression. Marine Biotechnology, 4(6):537– 545, December 2002. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-002-0013-1.

Schroder:2002:SNQ

[311] Heinz-Christoph Schröder, Sebastian Sudek, Salvatore Caro, Salvatore Rosa, Sanja Perovic, Renate Steffen, Isabel M. Möller, and Werner E. G. Möller. Synthesis of the neurotoxin quinolinic acid in apoptotic tissue from Suberites domuncula: Cell biological, molecular biological, and chemical analyses. Marine Biotechnology, 4(6):546–558, December 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-

2236 (electronic). URL https://link.springer.com/article/10.
1007/s10126-002-0028-7.

Urano:2002:BFC

[312] Naoto Urano, Eigo Sasaki, Ryohei Ueno, Hidehiro Namba, and Yusuke Shida. Bioremediation of fish cannery wastewater with yeasts isolated from a drainage canal. Marine Biotechnology, 4(6):559–564, December 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-002-0045-6.

Canete:2002:KCG

[313] Socrates Jose P. Cañete and Marco Nemesio E. Montaño. k-carrageenan gel as agent to sequester paralytic shellfish poison. Marine Biotechnology, 4(6):565–570, December 2002. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-002-0054-5.

Wada:2002:AHV

[314] Shun ichi Wada, Sachi Kantha, Takahiro Yamashita, Shigeki Matsunaga, Nobuhiro Fusetani, and Shugo Watabe. Accumulation of h+ in vacuoles induced by a marine peptide toxin, theonellamide f, in rat embryonic 3Y1 fibroblasts. Marine Biotechnology, 4(6):571–582, December 2002. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-002-0044-7.

Ely:2002:ANL

[315] B. Ely, D. S. Stoner, Alvarado J. R. Bremer, J. M. Dean, P. Addis, A. Cau, E. J. Thelen, W. J. Jones, D. E. Black, L. Smith, K. Scott, I. Naseri, and J. M. Quattro. Analyses of nuclear ldhA gene and mtDNA control region sequences of Atlantic Northern bluefin tuna populations. Marine Biotechnology, 4(6):583–588, December 2002. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-002-0040-y.

Chapman:2002:SHT

[316] Robert W. Chapman, A. O. Ball, and Lisa R. Mash. Spatial homogeneity and temporal heterogeneity of red drum (Sciaenops ocellatus) microsatellites: Effective population sizes and management implications. Marine Biotechnology, 4(6):589–603, December 2002. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-002-0038-5.

Dunham:2002:ERT

[317] R. A. Dunham, N. Chatakondi, A. J. Nichols, H. Kucuktas, T. T. Chen, D. A. Powers, J. D. Weete, K. Cummins, and R. T. Lovell. Effect of rainbow trout growth hormone complementary DNA on body shape, carcass yield, and carcass composition of F1 and F2 transgenic common carp (Cyprinus carpio). Marine Biotechnology, 4(6): 604–611, December 2002. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-002-0034-9.

Rein:2003:PSG

[318] K. S. Rein, P. D. L. Gibbs, A. Palacios, L. Abiy, R. Dickey, R. V. Snyder, and J. V. Lopez. Polyketide synthase genes from marine dinoflagellates. Marine Biotechnology, 5(1):1–12, February 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-002-0077-y.

Matsuda:2003:SRS

[319] Masahiro Matsuda, Takao Yamori, Mikihiko Naitoh, and Koichi Okutani. Structural revision of sulfated polysaccharide B-1 isolated from a marine Pseudomonas species and its cytotoxic activity against human cancer cell lines. Marine Biotechnology, 5(1):13–19, February 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-002-0046-5.

Iwamoto:2003:CNN

[320] Koji Iwamoto and Yoshihiro Shiraiwa. Characterization of NADH:Nitrate reductase from the coccolithophorid Emiliania huxleyi (Lohman) Hay & Mohler (Haptophyceae). Marine Biotechnology, 5(1):20–26, February 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-002-0051-8.

Klinbunga:2003:MGI

[321] S. Klinbunga, N. Khamnamtong, A. Tassanakajon, N. Puanglarp, P. Jarayabhand, and W. Yoosukh. Molecular genetic identification tools for three commercially cultured oysters (Crassostrea belcheri, Crassostrea iredalei, and Saccostrea cucullata) in Thailand. Marine Biotechnology, 5 (1):27–36, February 2003. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-002-0047-4.

Miyamoto:2003:PPC

[322] Hiroshi Miyamoto, Kohichi Morimoto, Akio Tanaka, Koki Sato, Aizo Matsushiro, Tomoyuki Miyashita, and Ben’ichiro Tonomura. Presence of protein complex is prerequisite for aragonite crystallization in the nacreous layer. Marine Biotechnology, 5(1):37–44, February 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-002-0048-3.

DeMoor:2003:CAC

[323] Sonia DeMoor, Herbert J. Waite, Michel J. Jangoux, and Patrick J. Flammang. Characterization of the adhesive from cuvierian tubules of the sea cucumber Holothuria forskali (Echinodermata, Holothuroidea). Marine Biotechnology, 5(1):45–57, February 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-002-0049-2.

Wagle:2003:BMG

[324] Mahendra Wagle and Suresh Jesuthasan. Baculovirus-mediated gene expression in zebrafish. Marine Biotechnology, 5(1):58–63, February 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-002-0050-9.

Fan:2003:ICC

[325] Ting-Jun Fan, Ling-Yun Jin, and Xiao-Feng Wang. Initiation of cartilage cell culture from skate (Raja porasa Günther). Marine Biotechnology, 5 (1):64–69, February 2003. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-002-0055-4.

Sakai:2003:MBS

[326] Takeshi Sakai, Hitomi Kimura, Kaoru Kojima, Kazuo Shimanaka, Katsushige Ikai, and Ikunoshin Kato. Marine bacterial sulfated fucoglucuronomannan (SFGM) lyase digests brown algal SFGM into trisaccharides. Marine Biotechnology, 5(1):70–78, February 2003. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-002-0056-3.

Deane:2003:LDS

[327] E. E. Deane, S. P. Kelly, P. M. Collins, and N. Y. S. Woo. Larval development of silver sea bream (Sparus sarba): Ontogeny of RNA– DNA ratio, GH, IGF-I, and Na^{+}-K^{+}-ATPase. Marine Biotechnology, 5 (1):79–91, February 2003. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-002-0052-7.

Calduch-Giner:2003:ECG

[328] Vega-Rubön J. A. Calduch-Giner, Vega-Rubön J. Pörez-Sönchez, Vega-Rubön S. de Celis, Vega-Rubön P. Gömez, and Vega-Rubön F. Mödale. Expression and characterization of European sea bass (Dicentrarchus labrax) somatolactin: Assessment of In Vivo metabolic effects. Marine Biotechnology, 5(1):92–101, February 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-002-0053-6.

Anonymous:2003:E

[329] Anonymous. Erratum. Marine Biotechnology, 5(1):102, February 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-002-0095-9.

DeTomaso:2003:CCL

[330] Anthony W. De Tomaso and Irving L. Weissman. Construction and characterization of large-insert genomic libraries (BAC and Fosmid) from the ascidian Botryllus schlosseri and initial physical mapping of a histocompatibility locus. Marine Biotechnology, 5(2):103–115, April 2003. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-002-0071-4.

Sardessai:2003:IOS

[331] Yogita Sardessai and Saroj Bhosle. Isolation of an organic–solventtolerant cholesterol-transforming Bacillus species, BC1, from coastal sediment. Marine Biotechnology, 5(2):116–118, April 2003. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-002-0069-y.

Hsieh:2003:UCM

[332] Bo-Chuan Hsieh, Tzong-Jih Cheng, Tzu-Yu Wang, and Richie L. C. Chen. Use of chitosan membrane from the carapace of the soldier crab Mictyris brevidactylus for biosensor construction. Marine Biotechnology, 5(2):119–125, April 2003. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-002-0094-x.

Rivera:2003:ICN

[333] Malia Ana J. Rivera, Glenn C. Graham, and George K. Roderick. Isolation and characterization of nine microsatellite loci from the Hawaiian grouper Epinephelus quernus (Serranidae) for population genetic analyses. Marine Biotechnology, 5(2):126–129, April 2003. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-002-0093-y.

Santos:2003:PIS

[334] Scott R. Santos, Carla Gutierrez-Rodriguez, and Mary Alice Coffroth. Phylogenetic identification of symbiotic dinoflagellates via length heteroplasmy in domain V of chloroplast large subunit (cp23S) — ribosomal DNA sequences. Marine Biotechnology, 5(2):130–140, April 2003. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-002-0076-z.

Delghandi:2003:SAS

[335] Madjid Delghandi, Atle Mortensen, and Jon-Ivar Westgaard. Simultaneous analysis of six microsatellite markers in Atlantic cod (Gadus morhua): a novel multiplex assay system for use in selective breeding studies. Marine Biotechnology, 5(2):141–148, April 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-002-0086-x.

daCosta:2003:CIM

[336] Antonio Carlos Augusto da Costa and Francisca Pessôa de França. Cadmium interaction with microalgal cells, cyanobacterial cells, and seaweeds; toxicology and biotechnological potential for wastewater treatment. Marine Biotechnology, 5(2):149–156, April 2003. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-002-0109-7.

Kamata:2003:CMB

[337] Shin-Ichiro Kamata and Satoru Suzuki. Concentration of marine birnavirus from seawater with a glass fiber filter precoated with bovine serum albumin. Marine Biotechnology, 5(2):157–162, April 2003. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-002-0057-2.

Aassila:2003:IHA

[338] H. Aassila, M. L. Bourguet-Kondracki, S. Rifai, A. Fassouane, and M. Guyot. Identification of harman as the antibiotic compound produced by a tunicate-associated bacterium. Marine Biotechnology, 5 (2):163–166, April 2003. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-002-0060-7.

Ken:2003:CFC

[339] Chuian-Fu Ken, Chi-Tsai Lin, Jei-Fu Shaw, and Jen-Leih Wu. Characterization of fish Cu/Zn-superoxide dismutase and its protection from oxidative stress. Marine Biotechnology, 5(2):167–173, April 2003. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-002-0058-1.

Cui:2003:RMT

[340] Zongbin Cui, Ying Yang, Christopher D. Kaufman, Dritan Agalliu, and Perry B. Hackett. RecA-mediated, targeted mutagenesis in zebrafish. Marine Biotechnology, 5(2):174–184, April 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-002-0059-0.

De:2003:TVT

[341] Jaysankar De, N. Ramaiah, A. Mesquita, and X. N. Verlekar. Tolerance to various toxicants by marine bacteria highly resistant to mercury. Marine Biotechnology, 5(2):185–193, April 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-002-0061-6.

Misonou:2003:UAS

[342] Taku Misonou, Junko Saitoh, Saori Oshiba, Yukiko Tokitomo, Miyuki Maegawa, Yukio Inoue, Hirokazu Hori, and Takeki Sakurai. UVabsorbing substance in the red alga Porphyra yezoensis (Bangiales, Rhodophyta) block thymine photodimer production. Marine Biotechnology, 5(2):194–200, April 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-002-0065-2.

Smith:2003:MPC

[343] Denice G. Smith, Claudia Wilburn, and Robert A. McCarthy. Methoprene photolytic compounds disrupt zebrafish development, producing phenocopies of mutants in the sonic hedgehog signaling pathway. Marine Biotechnology, 5(2):201–212, April 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-002-0062-5.

Springer-Verlag:2003:NDM

[344] Springer-Verlag. New developments in marine biotechnology. Marine Biotechnology, 5(3):213, June 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-003-0003-y.

Scapigliati:2003:IBL

[345] G. Scapigliati, S. Meloni, F. Buonocore, P. Bossö, D. Prugnoli, and C. J. Secombes. Immunopurification of B lymphocytes from sea bass Dicentrarchus labrax (L.). Marine Biotechnology, 5(3):214–221, June 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-002-0100-3.

Kumazawa:2003:PHM

[346] Shuzo Kumazawa. Photoproduction of hydrogen by the marine heterocystous cyanobacterium Anabaena species TU37-1 under a nitrogen atmosphere. Marine Biotechnology, 5(3):222–226, June 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-002-0106-x.

Sekino:2003:DNM

[347] Masashi Sekino, Masami Hamaguchi, Futoshi Aranishi, and Kenji Okoshi. Development of novel microsatellite DNA markers from the Pacific oyster Crassostrea gigas. Marine Biotechnology, 5(3):227–233, June 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-002-0104-z.

Was:2003:MDP

[348] Anna Was and Roman Wenne. Microsatellite DNA polymorphism in intensely enhanced populations of sea trout (Salmo trutta) in the Southern Baltic. Marine Biotechnology, 5(3):234–243, June 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-002-0068-z.

Noh:2003:GCM

[349] Jae Koo Noh, Kyu-Nam Cho, Eun Hwa Han, AeRi Kim, Jae-Seong Lee, Dong Soo Kim, and Chul Geun Kim. Genomic cloning of mud loach Misgurnus mizolepis (Cypriniformes, Cobitidae) β-actin gene and usefulness of its promoter region for fish transgenesis. Marine Biotechnology, 5 (3):244–252, June 2003. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-002-0066-1.

Syed:2003:IPG

[350] Mohasina Syed, Olav Vestrheim, Birthe Mikkelsen, and Maria Lundin. Isolation of the promoters of Atlantic salmon MHCII genes. Marine Biotechnology, 5(3):253–260, June 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-002-0063-4.

Kimura:2003:MCS

[351] Hiroyuki Kimura, Makoto Sato, Yuichi Sasayama, and Takeshi Naganuma. Molecular characterization and In Situ localization of endosymbiotic 16S ribosomal RNA and RuBisCO genes in the pogonophoran tissue. Marine Biotechnology, 5(3):261–269, June 2003. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-002-0073-2.

Noakes:2003:GCM

[352] Marc A. Noakes, Tara Reimer, and Ruth B. Phillips. Genotypic characterization of an MHC class II locus in lake trout (Salvelinus namaycush) from Lake Superior by single-stranded conformational polymorphism analysis and reference strand-mediated conformational analysis. Marine Biotechnology, 5(3):270–278, June 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-002-0079-9.

Khoo:2003:GLM

[353] Gideon Khoo, Meng Huat Lim, Haridas Suresh, Damien K. Y. Gan, Kok Fang Lim, Fan Chen, Woon-Khiong Chan, Tit Meng Lim, and Violet P. E. Phang. Genetic linkage maps of the guppy (Poecilia reticulata): Assignment of RAPD markers to multipoint linkage groups. Marine Biotechnology, 5(3):279–293, June 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-002-0072-3.

Dunlap:2003:MPA

[354] Walter Dunlap, Lyndon Llewellyn, Jason Doyle, and Yorihiro Yamamoto. A microtiter plate assay for screening antioxidant activity in extracts of marine organisms. Marine Biotechnology, 5(3):294–301, June 2003. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-002-0074-1.

Zhang:2003:SHY

[355] C. W. Zhang and A. Richmond. Sustainable, high-yielding outdoor mass cultures of Chaetoceros muelleri var. subsalsum and Isochrysis galbana in vertical plate reactors. Marine Biotechnology, 5(3):302–310, June 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-002-0075-0.

Meehan:2003:HFL

[356] Dawn Meehan, Zhenkang Xu, Gladys Zuniga, and Acacia Alcivar-Warren. High frequency and large number of polymorphic microsatellites in cultured shrimp, Penaeus (Litopenaeus) vannamei [Crustacea: Decapoda]. Marine Biotechnology, 5(4):311–330, August 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-002-0092-z.

Li:2003:IMD

[357] Qi Li, Choulji Park, Toshimasa Kobayashi, and Akihiro Kijima. Inheritance of microsatellite DNA markers in the Pacific abalone Haliotis discus hannai. Marine Biotechnology, 5(4):331–338, August 2003. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-002-0116-8.

Lehrer:2003:SAA

[358] S. B. Lehrer, R. Ayuso, and G. Reese. Seafood allergy and allergens: a review. Marine Biotechnology, 5(4):339–348, August 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-002-0082-1.

Burtseva:2003:FMF

[359] Yu. V. Burtseva, N. S. Verigina, V. V. Sova, M. V. Pivkin, and T. N. Zvyagintseva. Filamentous marine fungi as producers of Oglycosylhydrolases: β-1,3-glucanase from Chaetomium indicum. Marine Biotechnology, 5(4):349–359, August 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-002-0070-2.

Sellos:2003:SAG

[360] D. Sellos, J. Moal, L. Degremont, A. Huvet, J.-Y. Daniel, S. Nicoulaud, P. Boudry, J.-F. Samain, and A. Van Wormhoudt. Structure of amylase genes in populations of Pacific cupped oyster (Crassostrea gigas): Tissue expression and allelic polymorphism. Marine Biotechnology, 5 (4):360–372, August 2003. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-002-0089-7.

Treerattrakool:2003:EBA

[361] Supattra Treerattrakool, Apinunt Udomkit, Lily Eurwilaichitr, Burachai Sonthayanon, and Sakol Panyim. Expression of biologically active crustacean hyperglycemic hormone (CHH) of Penaeus monodon in Pichia pastoris. Marine Biotechnology, 5(4):373–379, August 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-002-0078-x.

Sakai:2003:PSF

[362] Takeshi Sakai, Hitomi Kimura, and Ikunoshin Kato. Purification of sulfated fucoglucuronomannan lyase from bacterial strain of Fucobacter marina and study of appropriate conditions for its enzyme digestion. Marine Biotechnology, 5(4):380–387, August 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-002-0083-0.

Pimentel:2003:PAI

[363] Sheila Marie V. Pimentel, Zenaida P. Bojo, Amy V. D. Roberto, Jose Enrico H. Lazaro, Gina C. Mangalindan, Leila M. Florentino, Pilar Lim-Navarro, Deniz Tasdemir, Chris M. Ireland, and Gisela P. Concepcion. Platelet aggregation inhibitors from Philippine marine invertebrate samples screened in a new microplate assay. Marine Biotechnology, 5(4): 395–400, August 2003. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-002-0080-3.

Fujiwara:2003:MPR

[364] Shoko Fujiwara, Kazuyuki Yasui, Kinzo Watanabe, Takako Wakabayashi, Mikio Tsuzuki, and Kazuo Iguchi. Molecular phylogenetic relationships between prostanoid-containing Okinawan soft coral (Clavularia viridis) and nonprostanoid-containing Clavularia species based on ribosomal ITS sequence. Marine Biotechnology, 5(4):401–407, August 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-002-0084-z.

Sakai:2003:ICF

[365] Takeshi Sakai, Kumiko Ishizuka, and Ikunoshin Kato. Isolation and characterization of a fucoidan-degrading marine bacterium. Marine Biotechnology, 5(5):409–416, October 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-002-0118-6.

Asai:2003:PBR

[366] Ryoichi Asai, Kozue Ootani, Yoko Nomura, Chikashi Nakamura, Kazunori Ikebukuro, Yoshiko Arikawa, Jun Miyake, and Isao Karube. PCR-based ribosomal DNA detection technique for microalga (Heterosigma carterae) causing red tide and its application to a biosensor using labeled probe. Marine Biotechnology, 5(5):417–423, October 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-002-0081-2.

Pabel:2003:AAM

[367] Christian T. Pabel, Joachim Vater, Christopher Wilde, Peter Franke, Jörgen Hofemeister, Barbara Adler, Gerhard Bringmann, Jörg Hacker, and Ute Hentschel. Antimicrobial activities and matrix-assisted laser desorption/ionization mass spectrometry of Bacillus isolates from the marine sponge Aplysina aerophoba. Marine Biotechnology, 5(5):424– 434, October 2003. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-002-0088-8.

Onji:2003:VLP

[368] Masashi Onji, Shin ichi Nakano, and Satoru Suzuki. Virus-like particles suppress growth of the red-tide-forming marine dinoflagellate Gymnodinium mikimotoi. Marine Biotechnology, 5(5):435–442, October 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-002-0085-y.

Sipkema:2003:SCC

[369] Detmer Sipkema, Hans G. H. J. Heilig, Antoon D. L. Akkermans, Ronald Osinga, Johannes Onji, and Renö H. Wijffels. Sponge-cell culture? A molecular identification method for sponge cells. Marine Biotechnology, 5 (5):443–449, October 2003. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-002-0090-1.

Huang:2003:GNI

[370] Jianzhong Huang, Tsunehiro Aki, Toshihiro Yokochi, Toro Nakahara, Daiske Honda, Seiji Kawamoto, Seiko Shigeta, Kazuhisa Ono, and Osamu Suzuki. Grouping newly isolated docosahexaenoic acid-producing thraustochytrids based on their polyunsaturated fatty acid profiles and comparative analysis of 18S rRNA genes. Marine Biotechnology, 5(5): 450–457, October 2003. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-002-0110-1.

Fukuba:2003:PDD

[371] Tatsuhiro Fukuba, Mari Ogawa, Teruo Fujii, and Takeshi Naganuma. Phylogenetic diversity of dissimilatory sulfite reductase genes from deepsea cold seep sediment. Marine Biotechnology, 5(5):458–468, October 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-002-0091-0.

Katsares:2003:DMD

[372] Vassilios Katsares, Apostolos Apostolidis, Alexandros Triantafyllidis, Anastasia Kouvatsi, and Costas Triantaphyllidis. Development of mitochondrial DNA primers for use with homarid lobsters. Marine Biotechnology, 5(5):469–479, October 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-002-0097-7.

Carter:2003:PTP

[373] C. G. Carter, M. P. Bransden, T. E. Lewis, and P. D. Nichols. Potential of thraustochytrids to partially replace fish oil in Atlantic salmon feeds. Marine Biotechnology, 5(5):480–492, October 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-002-0096-8.

Triantafyllidis:2003:GIP

[374] Alexandros Triantafyllidis, Anastasia Kouvatsi, Costas Triantaphyllidis, Nikoletta Karaiskou, and Apostolos P. Apostolidis. Genetic identification and phylogeny of three species of the genus Trachurus based on mitochondrial DNA analysis. Marine Biotechnology, 5(5):493–504, October 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-002-0099-5.

Klinbunga:2003:GDM

[375] S. Klinbunga, P. Pripue, N. Khamnamtong, N. Puanglarp, A. Tassanakajon, P. Jarayabhand, I. Hirono, T. Aoki, and P. Menasveta. Genetic diversity and molecular markers of the tropical abalone (Haliotis asinina) in Thailand. Marine Biotechnology, 5(5):505–517, October 2003. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-002-0108-8.

Springer-Verlag:2003:E

[376] Springer-Verlag. Erratum. Marine Biotechnology, 5(5):518, October 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-003-0030-x.

Duckworth:2003:VCT

[377] Alan R. Duckworth, Gail A. Samples, Amy E. Wright, and Shirley A. Pomponi. In vitro culture of the tropical sponge Axinella corrugata (Demospongiae): Effect of food cell concentration on growth, clearance rate, and biosynthesis of stevensine. Marine Biotechnology, 5(6): 519–527, December 2003. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-002-0111-0.

Li:2003:GDC

[378] Fuhua Li, Jianhai Xiang, Xiaojun Zhang, Linghua Zhou, Chengsong Zhang, and Changgong Wu. Gonad development characteristics and sex ratio in triploid Chinese shrimp (Fenneropenaeus chinensis). Marine Biotechnology, 5(6):528–535, December 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-002-0103-0.

Shimanaka:2003:SOD

[379] Kazuo Shimanaka, Katsushige Ikai, Ikunoshin Kato, Takeshi Sakai, and Kumiko Ishizuka. Structures of oligosaccharides derived from Cladosiphon okamuranus fucoidan by digestion with marine bacterial enzymes. Marine Biotechnology, 5(6):536–544, December 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-002-0107-9.

Tsoi:2003:UHC

[380] Stephen C. M. Tsoi, Jacqueline M. Cale, Ian M. Bird, Vanya Ewart, Laura L. Brown, and Susan Douglas. Use of human cDNA microarrays for identification of differentially expressed genes in Atlantic salmon liver during Aeromonas salmonicida infection. Marine Biotechnology, 5(6): 545–554, December 2003. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-002-0112-z.

Lam:2003:MIP

[381] Katherine Lam and Brian Morton. Morphological and ITS1, 5.8S, and partial ITS2 ribosomal DNA sequence distinctions between two species Playtygyra (Cnidaria: Scleractinia) from Hong Kong. Marine Biotechnology, 5(6):555–567, December 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-002-0114-x.

Ben:2003:CTE

[382] Jin Ben, Tit Meng Lim, Violet P. E. Phang, and Woon-Khiong Chan. Cloning and tissue expression of 6-pyruvoyl tetrahydropterin synthase and xanthine dehydrogenase from Poecilia reticulata. Marine Biotechnology, 5(6):568–578, December 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-002-0121-y.

Hulsch:2003:SCP

[383] R. Hulsch, J. O. Metzger, L. Berthe-Corti, and R. Bredemeier. Submersed culture production of extracellular wax esters by the marine bacterium Fundibacter jadensis. Marine Biotechnology, 5(6):579–583, December 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-

2236 (electronic). URL https://link.springer.com/article/10.
1007/s10126-003-0012-x.

Castro:2003:GAU

[384] J. Castro, C. Bouza, L. Sönchez, R. M. Cal, F. Piferrer, and P. Martönez. Gynogenesis assessment using microsatellite genetic markers in turbot (Scophthalmus maximus). Marine Biotechnology, 5(6):584–592, December 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-003-0004-x.

Kimura:2003:EMV

[385] Hiroyuki Kimura, Yukimasa Higashide, and Takeshi Naganuma. Endosymbiotic microflora of the vestimentiferan tubeworm (Lamellibrachia sp.) from a bathyal cold seep. Marine Biotechnology, 5(6):593–603, December 2003. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-002-0117-7.

Lindeque:2003:TTT

[386] Penelope K. Lindeque and Gary R. Smerdon. Temporal transcription of two Antennapedia class homeobox genes in the marine copepod Calanus helgolandicus. Marine Biotechnology, 5(6):604–615, December 2003. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-002-0119-5.

Huang:2004:MCT

[387] Wei-Tung Huang, Cheng-Wen Liu, Cho-Fat Hui, Thomas T. Chen, Jen-Leih Wu, Jyh-Yih Chen, and Jian-Chyi Chen. Molecular cloning and tissue-specific, developmental-stage-specific, and hormonal regulation of IGFBP3 gene in zebrafish. Marine Biotechnology, 6(1):1–7, February 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-002-0115-9.

Zhang:2004:MCE

[388] Yong Zhang, Jie Huang, Qingxiong Meng, Tiemin Jiang, Liping Xie, Zhao Wang, and Rongqing Zhang. Molecular cloning and expression of a pearl oyster (Pinctada fucata) homologue of mammalian putative tumor suppressor QM. Marine Biotechnology, 6(1):8–16, February 2004. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-002-0120-z.

Yim:2004:AES

[389] Joung Han Yim, Sung Jin Kim, Se Hun Ahn, Chong Kyo Lee, Ki Tae Rhie, and Hong Kum Lee. Antiviral effects of sulfated exopolysaccharide from the marine microalga Gyrodinium impudicum strain KG03. Marine Biotechnology, 6(1):17–25, February 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-003-0002-z.

Li:2004:ABG

[390] Li Li and Ximing Guo. AFLP-based genetic linkage maps of the Pacific oyster Crassostrea gigas Thunberg. Marine Biotechnology, 6(1): 26–36, February 2004. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-003-0001-0.

Mayer:2004:MPM

[391] Alejandro M. S. Mayer and Mark T. Hamann. Marine pharmacology in 2000: Marine compounds with antibacterial, anticoagulant, antifungal, anti-inflammatory, antimalarial, antiplatelet, antituberculosis, and antiviral activities; affecting the cardiovascular, immune, and nervous systems and other miscellaneous mechanisms of action. Marine Biotechnology, 6(1):37–52, February 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-003-0007-7.

Buonocore:2004:EEC

[392] Francesco Buonocore, Massimo Mazzini, Maria Forlenza, Elisa Randelli, Christopher J. Secombes, Jun Zou, and Giuseppe Scapigliati. Expression in Escherchia coli and purification of sea bass (Dicentrarchus labrax) interleukin 1β, a possible immunoadjuvant in aquaculture. Marine Biotechnology, 6(1):53–59, February 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-003-0011-y.

Cancela:2004:MPG

[393] M. L. Cancela, R. Streiff, S. Mira, and M. Castro. Multiple paternity in Norway lobster (Nephrops norvegicus L.) assessed with microsatellite markers. Marine Biotechnology, 6(1):60–66, February 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-003-0015-7.

Hellio:2004:SVA

[394] Claire Hellio, Jean-Philippe Marechal, Benoöt Vöron, Graham Bremer, Anthony S. Clare, and Yves Le Gal. Seasonal variation of antifouling activities of marine algae from the Brittany Coast (France). Marine Biotechnology, 6(1):67–82, February 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-003-0020-x.

Chen:2004:CTC

[395] Shu-Hwa Chen, Chung-Yen Lin, and Ching Ming Kuo. Cloning of two crustacean hyperglycemic hormone isoforms in freshwater giant prawn (Macrobrachium rosenbergii): Evidence of alternative splicing. Marine Biotechnology, 6(1):83–94, February 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-003-0014-8.

Mitova:2004:ECD

[396] Maya Mitova, Giuseppina Tommonaro, Ute Hentschel, Werner E. G. Möller, and Salvatore De Rosa. Exocellular cyclic dipeptides from a Ruegeria strain associated with cell cultures of Suberites domuncula. Marine Biotechnology, 6(1):95–103, February 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-003-0018-4.

Springer-Verlag:2004:E

[397] Springer-Verlag. Erratum. Marine Biotechnology, 6(1):104, February 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-004-1500-3.

Moller:2004:SPB

[398] Werner E. G. Möller, Vladislav A. Grebenjuk, Gaöl Le Pennec, Heinz-C. Schröder, Franz Brömmer, Ute Hentschel, Isabel M. Möller, and Hans-J. Breter. Sustainable production of bioactive compounds by sponges — cell culture and gene cluster approach: a review. Marine Biotechnology, 6(2):105–117, April 2004. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-002-0098-6.

Rocha:2004:AIT

[399] A. Rocha, S. Ruiz, A. Estepa, and J. M. Coll. Application of inducible and targeted gene strategies to produce transgenic fish: a review. Marine Biotechnology, 6(2):118–127, April 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-003-0013-9.

Chandler:2004:SCL

[400] G. Thomas Chandler and David C. Volz. Semiquantitative confocal laser scanning microscopy applied to marine invertebrate ecotoxicology. Marine Biotechnology, 6(2):128–137, April 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-002-0105-y.

Chen:2004:CCF

[401] Ming-Chyuan Chen, Ying-Min Cheng, Li-Hsueh Wang, Chorng-Horng Lin, Xing-Yan Huang, Ming-Chin Liu, Ping-Jyun Sung, and Lee-Shing Fang. Cloning and characterization of the first cnidarian ADP– Ribosylation factor, and its involvement in the Aiptasia-Symbiodinum endosymbiosis. Marine Biotechnology, 6(2):138–147, April 2004. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-002-0113-y.

Briones:2004:TTS

[402] Annabelle V. Briones, Wilhelmina O. Ambal, Romulo R. Estrella, Rolando Pangilinan, Carlos J. De Vera, Raymund L. Pacis, Ner Rodriguez, and Merle A. Villanueva. Tensile and tear strength of carrageenan film from Philippine Eucheuma species. Marine Biotechnology, 6(2):148–151, April 2004. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-003-0005-9.

Lang:2004:IPB

[403] Siegmund Lang, Winfried Beil, Harukumi Tokuda, Corinna Wicke, and Verena Lurtz. Improved production of bioactive glucosylmannosyl– glycerolipid by sponge-associated Microbacterium species. Marine Biotechnology, 6(2):152–156, April 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-003-0009-5.

Cruz:2004:GVA

[404] Pedro Cruz, Ana M. Ibarra, Humberto Mejia-Ruiz, Patrick M. Gaffney, and Ricardo Pörez-Enröquez. Genetic variability assessed by microsatellites in a breeding program of Pacific white shrimp (Litopenaeus vannamei). Marine Biotechnology, 6(2):157–164, April 2004. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-003-0017-5.

Yamaguchi:2004:EST

[405] Nobuo Yamaguchi, Kei Kamino, Tatsuya Ueki, and Hitoshi Michibata. Expressed sequence tag analysis of vanadocytes in a vanadiumrich ascidian, Ascidia sydneiensis samea. Marine Biotechnology, 6(2): 165–174, April 2004. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-003-0024-6.

Garber:2004:PSV

[406] Amber F. Garber, Michael D. Tringali, and Kenneth C. Stuck. Population structure and variation in red snapper (Lutjanus campechanus) from the Gulf of Mexico and Atlantic Coast of Florida as determined from mitochondrial DNA control region sequence. Marine Biotechnology, 6 (2):175–185, April 2004. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-003-0023-7.

Nagler:2004:DGS

[407] James J. Nagler, Tim Cavileer, Kirk Steinhorst, and Robert H. Devlin. Determination of genetic sex in Chinook salmon (Oncorhynchus tshawytscha) using the male-linked growth hormone pseudogene by realtime PCR. Marine Biotechnology, 6(2):186–191, April 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-003-0022-8.

Imada:2004:EIM

[408] Chiaki Imada. Enzyme inhibitors of marine microbial origin with pharmaceutical importance. Marine Biotechnology, 6(3):193–198, June 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-003-0027-3.

Tsoi:2004:IIR

[409] S. C. M. Tsoi, K. V. Ewart, S. Penny, K. Melville, R. S. Liebscher, L. L. Brown, and S. E. Douglas. Identification of immune-relevant genes from Atlantic salmon using suppression subtractive hybridization. Marine Biotechnology, 6(3):199–214, June 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-002-0101-2.

Trischman:2004:CIE

[410] Jacqueline A. Trischman, Richard E. Oeffner, Michelle G. de Luna, and Martin Kazaoka. Competitive induction and enhancement of indole and a diketopiperazine in marine bacteria. Marine Biotechnology, 6(3):215–220, June 2004. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-003-0010-z.

Kumagai:2004:GDB

[411] Keita Kumagai, Anna A. Barinova, Masamichi Nakajima, and Nobuhiko Taniguchi. Genetic diversity between Japanese and Chinese threeline grunt (Parapristipoma trilineatum) examined by microsatellite DNA markers. Marine Biotechnology, 6(3):221–228, June 2004. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-003-0006-8.

Govindaraju:2004:TRA

[412] G. S. Govindaraju and P. Jayasankar. Taxonomic relationship among seven species of groupers (genus Epinephelus; family Serranidae) as revealed by RAPD fingerprinting. Marine Biotechnology, 6(3):229–237, June 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-003-0021-9.

Machida:2004:OMG

[413] Ryuji J. Machida, Masaki U. Miya, Mitsugu M. Yamauchi, Mutsumi Nishida, and Shuhei Nishida. Organization of the mitochondrial genome of Antarctic krill Euphausia superba (Crustacea: Malacostraca). Marine Biotechnology, 6(3):238–250, June 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-003-0016-6.

Agaba:2004:ZCE

[414] Morris Agaba, Douglas R. Tocher, Cathryn A. Dickson, James R. Dick, and Alan J. Teale. Zebrafish cDNA encoding multifunctional fatty acid elongase involved in production of eicosapentaenoic (20:5n-3) and docosahexaenoic (22:6n-3) acids. Marine Biotechnology, 6(3): 251–261, June 2004. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-003-0029-1.

Celis-Guerrero:2004:CPD

[415] Laura E. Celis-Guerrero, Fernando L. García-Carreño, and M. Angeles Navarrete del Toro. Characterization of proteases in the digestive system of spiny lobster (Panulirus interruptus). Marine Biotechnology, 6(3):262–269, June 2004. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-003-0032-6.

Wiedenmann:2004:IGL

[416] Jörg Wiedenmann, Sergey Ivanchenko, Franz Oswald, and G. Ulrich. Nienhaus. Identification of GFP-like proteins in nonbioluminescent, azooxanthellate Anthozoa opens new perspectives for bioprospecting. Marine Biotechnology, 6(3):270–277, June 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-004-3006-4.

Wahlund:2004:AES

[417] Thomas M. Wahlund, Ahmad R. Hadaegh, Robin Clark, Binh Nguyen, Michael Fanelli, and Betsy A. Read. Analysis of expressed sequence tags from calcifying cells of marine coccolithophorid (Emiliania huxleyi). Marine Biotechnology, 6(3):278–290, June 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-003-0035-3.

Li:2004:IHA

[418] Qi Li, Tomoko Hisatsune, and Akihiro Kijima. Induction of haploid androgenesis in Pacific oyster by UV irradiation. Marine Biotechnology, 6(3):291–297, June 2004. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-003-0034-4.

Castillo:2004:FSS

[419] Ana G. F. Castillo, Jose L. Martinez, and Eva Garcia-Vazquez. Fine spatial structure of Atlantic hake (Merluccius merluccius) stocks revealed by variation at microsatellite loci. Marine Biotechnology, 6(4): 299–306, August 2004. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-3027-z.

Radenac:2004:MEA

[420] G. Radenac, G. Coteur, B. Danis, Ph. Dubois, and M. Warnau. Measurement of EROD activity: Caution on spectral properties of standards used. Marine Biotechnology, 6(4):307–311, August 2004. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-004-3014-4.

Chen:2004:PSG

[421] Chaolun Allen Chen, Maria Carmen Anonuevo Ablan, John Williams McManus, Johann Diepernk Bell, Vo Si Tuan, Annadel Sarmiento Cabanban, and Kwang-Tsao Shao. Population structure and genetic variability of six bar wrasse (Thallasoma hardwicki) in Northern South China Sea revealed by mitochondrial control region sequences. Marine Biotechnology, 6(4):312–326, August 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-003-0028-2.

Watanabe:2004:PSL

[422] Tomohisa Watanabe, Hironobu Fujita, Kimio Yamasaki, Shingo Seki, and Nobuhiko Taniguchi. Preliminary study on linkage mapping based on microsatellite DNA and AFLP markers using homozygous clonal fish in ayu (Plecoglossus altivelis). Marine Biotechnology, 6(4):327–334, August 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-003-0026-4.

Sakai:2004:ICF

[423] Takeshi Sakai, Takashi Kawai, and Ikunoshin Kato. Isolation and characterization of a fucoidan-degrading marine bacterial strain and its fucoidanase. Marine Biotechnology, 6(4):335–346, August 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-003-0033-5.

Lang:2004:DTA

[424] Gang hua Lang, Yong Wang, Nakao Nomura, and Masatoshi Matsumura. Detection of telomerase activity in tissues and primary cultured lymphoid cells of Penaeus japonicus. Marine Biotechnology, 6(4): 347–354, August 2004. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-003-0038-0.

Kawata:2004:TSP

[425] Yoshikazu Kawata, Shin’ichi Yano, Hiroyuki Kojima, and Masaaki Toyomizu. Transformation of Spirulina platensis strain C1 (Arthrospira sp. PCC9438) with Tn5 transposase–transposon DNA–cation liposome complex. Marine Biotechnology, 6(4):355–363, August 2004. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-003-0037-1.

Serapion:2004:BMT

[426] Jerry Serapion, Huseyin Kucuktas, Jinian Feng, and Zhanjiang Liu. Bioinformatic mining of type I microsatellites from expressed sequence tags of channel catfish (Ictalurus punctatus). Marine Biotechnology, 6 (4):364–377, August 2004. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-003-0039-z.

Ishikura:2004:INS

[427] Masaharu Ishikura, Kiyoshi Hagiwara, Kiyotaka Takishita, Miyuki Haga, Kenji Iwai, and Tadashi Maruyama. Isolation of new Symbiodinium strains from tridacnid giant clam (Tridacna crocea) and sea slug (Pteraeolidia ianthina) using culture medium containing giant clam tissue homogenate. Marine Biotechnology, 6(4):378–385, August 2004. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-004-1800-7.

Straub:2004:SSH

[428] Peter F. Straub, Mary L. Higham, Arnaud Tanguy, Brenda J. Landau, William C. Phoel, L. Stanton Hales, and Theodore K. M. Thwing. Suppression subtractive hybridization cDNA libraries to identify differentially expressed genes from contrasting fish habitats. Marine Biotechnology, 6(4):386–399, August 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-004-3146-6.

Rocha:2004:ITE

[429] A. Rocha, S. Ruiz, and J. M. Coll. Improvement of transfection efficiency of epithelioma papulosum cyprini carp cells by modification of cell cycle and use of an optimal promoter. Marine Biotechnology, 6(5): 401–410, October 2004. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-003-0008-6.

Astola:2004:GSF

[430] Antonio Astola, Josep A. Calduch-Giner, Manuela Ortiz, Jaume Pörez-Sönchez, and Manuel M. Valdivia. Genomic structure and functional analysis of promoter region of somatolactin gene of sea bream (Sparus aurata). Marine Biotechnology, 6(5):411–418, October 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-004-3210-2.

Yamauchi:2004:PBA

[431] Mitsugu M. Yamauchi, Masaki U. Miya, Ryuji J. Machida, and Mutsumi Nishida. PCR-based approach for sequencing mitochondrial genomes of decapod crustaceans, with a practical example from kuruma prawn (Marsupenaeus japonicus). Marine Biotechnology, 6(5):419–429, October 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-003-0036-2.

Moriya:2004:DMR

[432] Shogo Moriya, Shigehiko Urawa, Osamu Suzuki, Akihisa Urano, and Syuiti Abe. DNA microarray for rapid detection of mitochondrial DNA haplotypes of chum salmon. Marine Biotechnology, 6(5):430– 434, October 2004. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-1100-2.

Ezaz:2004:IPM

[433] M. Tariq Ezaz, Simon C. Harvey, Chuta Boonphakdee, Alan J. Teale, Brendan J. McAndrew, and David J. Penman. Isolation and physical mapping of sex-linked AFLP markers in Nile tilapia (Oreochromis niloticus L.). Marine Biotechnology, 6(5):435–445, October 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-004-3004-6.

Garoia:2004:FML

[434] Flavio Garoia, Ilaria Guarniero, Corrado Piccinetti, and Fausto Tinti. First microsatellite loci of red mullet (Mullus barbatus) and their application to genetic structure analysis of Adriatic shared stock. Marine Biotechnology, 6(5):446–452, October 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-004-3045-x.

Lee:2004:ICO

[435] Jae Hyung Lee, Mi Young Son, Moon-Young Yoon, Jung-Do Choi, and Young Tae Kim. Isolation and characterization of ornithine decarboxylase gene from flounder (Paralichthys olivaceus). Marine Biotechnology, 6 (5):453–462, October 2004. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-4100-3.

Hastings:2004:MCF

[436] Nicola Hastings, Morris K. Agaba, Douglas R. Tocher, Xiaozhong Zheng, Cathryn A. Dickson, James R. Dick, and Alan J. Teale. Molecular cloning and functional characterization of fatty acyl desaturase and elongase cDNAs involved in the production of eicosapentaenoic and docosahexaenoic acids from α-linolenic acid in Atlantic salmon (Salmo salar). Marine Biotechnology, 6(5):463–474, October 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-004-3002-8.

Valles-Jimenez:2004:PGS

[437] R. Valles-Jimenez, P. Cruz, and R. Perez-Enriquez. Population genetic structure of Pacific white shrimp (Litopenaeus vannamei) from Mexico to Panama: Microsatellite DNA variation. Marine Biotechnology, 6(5): 475–484, October 2004. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-3138-6.

Hwang:2004:FBT

[438] Gyulin Hwang, Ferenc Möller, M. Aziz. Rahman, Darren W. Williams, Paul J. Murdock, K. John Pasi, Geoffrey Goldspink, Hamid Farahmand, and Norman Maclean. Fish as bioreactors: Transgene expression of human coagulation factor VII in fish embryos. Marine Biotechnology, 6(5): 485–492, October 2004. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-3121-2.

Kimura:2004:CDC

[439] Tomohiro Kimura, Toshiki Nakano, Toshiyasu Yamaguchi, Minoru Sato, Tomohisa Ogawa, Koji Muramoto, Takehiko Yokoyama, Nobuhiro Kanno, Eizou Nagahisa, Frank Janssen, and Manfred K. Grieshaber. Complementary DNA cloning and molecular evolution of opine dehydrogenases in some marine invertebrates. Marine Biotechnology, 6(5): 493–502, October 2004. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-2700-6.

Elsaied:2004:MAD

[440] Hosam Easa Elsaied, Toru Hayashi, and Takeshi Naganuma. Molecular analysis of deep-sea hydrothermal vent aerobic methanotrophs by targeting genes of 16S rRNA and particulate methane monooxygenase. Marine Biotechnology, 6(5):503–509, October 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-004-3042-0.

delGaudio:2004:GTC

[441] D. del Gaudio, G. Fortunato, M. Borriello, J. M. Gili, P. Buono, G. Calcagno, F. Salvatore, and L. Sacchetti. Genetic typing of Corallium rubrum. Marine Biotechnology, 6(6):511–515, December 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-004-3001-9.

McFadden:2004:VCN

[442] Catherine S. McFadden, Ian D. Tullis, M. Breton Hutchinson, Katherine Winner, and Jill A. Sohm. Variation in coding (NADH dehydrogenase subunits 2, 3, and 6) and noncoding intergenic spacer regions of the mitochondrial genome in Octocorallia (Cnidaria: Anthozoa). Marine Biotechnology, 6(6):516–526, December 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-002-0102-1.

Abe:2004:IIM

[443] Shunnosuke Abe, Chhoun Chamnan, Kenichi Miyamoto, Yasutaka Minamino, and Makoto Nouda. Isolation and identification of 3methylcrotonyl coenzyme A carboxylase cDNAs and pyruvate carboxylase, and their expression in red seabream (Pagrus major) organs. Marine Biotechnology, 6(6):527–540, December 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-004-4203-x.

Jolly:2004:ZLS

[444] Cécile Jolly, Sophie Berland, Christian Milet, Sandrine Borzeix, Evelyne Lopez, and Dominique Doumenc. Zona localization of shell matrix proteins in mantle of Haliotis tuberculata (Mollusca, Gastropoda). Marine Biotechnology, 6(6):541–551, December 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-004-3129-7.

Franchi:2004:MER

[445] Elisabetta Franchi, Claudio Tosi, Giuseppe Scolla, Gino Della Penna, Francesco Rodriguez, and Paola Maria Pedroni. Metabolically engineered Rhodobacter sphaeroides RV strains for improved biohydrogen photoproduction combined with disposal of food wastes. Marine Biotechnology, 6(6):552–565, December 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-004-1007-y.

Watanabe:2004:QMS

[446] Shun Watanabe, Yuki Minegishi, Tatsuki Yoshinaga, Jun Aoyama, and Katsumi Tsukamoto. A quick method for species identification of Japanese eel (Anguilla japonica) using real-time PCR: an onboard application for use during sampling surveys. Marine Biotechnology, 6(6): 566–574, December 2004. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-1000-5.

Yu:2004:GAS

[447] Ziniu Yu and Ximing Guo. Genetic analysis of selected strains of Eastern oyster (Crassostrea virginica Gmelin) using AFLP and microsatellite markers. Marine Biotechnology, 6(6):575–586, December 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-004-3600-5.

Ki:2004:IMS

[448] Jang-Seu Ki, Gi Young Jang, and Myung-Soo Han. Integrated method for single-cell DNA extraction, PCR amplification, and sequencing of ribosomal DNA from harmful dinoflagellates Cochlodinium polykrikoides and Alexandrium catenella. Marine Biotechnology, 6(6):587–593, December 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-004-1700-x.

Pozzolini:2004:MCS

[449] Marina Pozzolini, Laura Sturla, Carlo Cerrano, Giorgio Bavestrello, Laura Camardella, Anna Maria Parodi, Federica Raheli, Umberto Benatti, Werner E. G. Müller, and Marco Giovine. Molecular cloning of silicatein gene from marine sponge Petrosia ficiformis (Porifera, Demospongiae) and development of primmorphs as a model for biosilicification studies. Marine Biotechnology, 6(6):594–603, December 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-004-3036-y.

Tang:2004:PST

[450] S. Tang, A. Tassanakajon, S. Klinbunga, P. Jarayabhand, and P. Menasveta. Population structure of tropical abalone (Haliotis asinina) in coastal waters of Thailand determined using microsatellite markers. Marine Biotechnology, 6(6):604–611, December 2004. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-004-2300-5.

Infante:2004:PRA

[451] Carlos Infante, Gaetano Catanese, and Manuel Manchado. Phylogenetic relationships among ten sole species (Soleidae, Pleuronectiformes) from the Gulf of Cádiz (Spain) based on mitochondrial DNA sequences. Marine Biotechnology, 6(6):612–624, December 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-004-3081-6.

Ferri:2004:CEF

[452] Stefano Ferri, Seiji Miura, Akane Sakaguchi, Fumimasa Ishimura, Wakako Tsugawa, and Koji Sode. Cloning and expression of fructosylamine oxidase from marine yeast Pichia species N1-1. Marine Biotechnology, 6(6):625–632, December 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-004-0001-8.

Longeon:2004:PPI

[453] Arlette Longeon, Jean Peduzzi, Michel Barthelemy, Sophie Corre, Jean-Louis Nicolas, and Michele Guyot. Purification and partial identification of novel antimicrobial protein from marine bacterium Pseudoalteromonas species strain X153. Marine Biotechnology, 6(6):633– 641, December 2004. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-3009-1.

Camacho:2004:PSC

[454] Pérez Camacho, J. M. Salinas, C. Fuertes, and M. Delgado. Preparation of single cell detritus from Laminaria saccharina as a hatchery diet for bivalve mollusks. Marine Biotechnology, 6(6):642–649, December 2004. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-004-2901-z.

Nagashima:2005:GSG

[455] Koji Nagashima, Maremi Sato, Kenji Kawamata, Aoi Nakamura, and Tomoki Ohta. Genetic structure of Japanese scallop population in Hokkaido, analyzed by mitochondrial haplotype distribution. Marine Biotechnology, 7(1):1–10, February 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-004-3046-9.

Patil:2005:SDG

[456] Jawahar G. Patil, Rasanthi M. Gunasekera, Bruce E. Deagle, and Nicholas J. Bax. Specific detection of Pacific oyster (Crassostrea gigas) larvae in plankton samples using nested polymerase chain reaction. Marine Biotechnology, 7(1):11–20, February 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-004-0034-z.

Imada:2005:DFC

[457] Chiaki Imada, Yohei Harada, Takeshi Kobayashi, Naoko Hamada-Sato, and Etsuo Watanabe. Degradation of ferric chelate of ethylenediaminetetraacetic acid by bacterium isolated from deep-sea stalked barnacle. Marine Biotechnology, 7(1):21–25, February 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-004-3700-2.

Liu:2005:CCM

[458] Shang-Yin Vanson Liu, Chang-Feng Dai, Tung-Yung Fan, and Hon-Tsen Yu. Cloning and characterization of microsatellite loci in a gorgonian coral, Junceella juncea (Anthozoa; Octocorallia; Ellisellidae) and its application in clonal genotyping. Marine Biotechnology, 7(1): 26–32, February 2005. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-0015-2.

Watanabe:2005:PNS

[459] Toshiki Watanabe, Mutsumi Nishida, Katsutoshi Watanabe, Defny S. Wewengkang, and Michio Hidaka. Polymorphism in nucleotide sequence of mitochondrial intergenic region in scleractinian coral (Galaxea fascicularis). Marine Biotechnology, 7(1):33–39, February 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-004-3200-4.

He:2005:VRD

[460] Maoxian He, Liangmin Huang, Jianhua Shi, and Yinping Jiang. Variability of ribosomal DNA ITS-2 and its utility in detecting genetic relatedness of pearl oyster. Marine Biotechnology, 7(1):40–45, February 2005. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-004-0003-6.

Jarasrassamee:2005:REC

[461] Boonyarin Jarasrassamee, Premruethai Supungul, Sakol Panyim, Sirawut Klinbunga, Vichien Rimphanichayakit, and Anchalee Tassanakajon. Recombinant expression and characterization of five-domain Kazaltype serine proteinase inhibitor of black tiger shrimp (Penaeus monodon). Marine Biotechnology, 7(1):46–52, February 2005. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-004-0100-6.

Kerr:2005:MPC

[462] Alexander M. Kerr, Daniel A. Janies, Ronald M. Clouse, Yves Samyn, Jeni Kuszak, and Junhyong Kim. Molecular phylogeny of coral-reef sea cucumbers (Holothuriidae: Aspidochirotida) based on 16S mitochondrial ribosomal DNA sequence. Marine Biotechnology, 7(1):53– 60, February 2005. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-0019-y.

Bakke:2005:MPG

[463] Ingrid Bakke and Steinar D. Johansen. Molecular phylogenetics of Gadidae and related Gadiformes based on mitochondrial DNA sequences. Marine Biotechnology, 7(1):61–69, February 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-004-3131-0.

Linder:2005:PES

[464] Michel Linder, Jacques Fanni, and Michel Parmentier. Proteolytic extraction of salmon oil and PUFA concentration by lipases. Marine Biotechnology, 7(1):70–76, February 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-004-0149-2.

Gudmundsdottir:2005:GAC

[465] Ágústa Gudmundsdóttir and Helga Margrét Pálsdóttir. Atlantic cod trypsins: From basic research to practical applications. Marine Biotechnology, 7(2):77–88, April 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.com/

article/10.1007/s10126-004-0061-9.

Ramos:2005:GET

[466] Erwin A. Ramos, Jenne Liza V. Relucio, and Celia Aurora T. Torres-Villanueva. Gene expression in tilapia following oral delivery of chitosanencapsulated plasmid DNA incorporated into fish feeds. Marine Biotechnology, 7(2):89–94, April 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.com/

article/10.1007/s10126-004-3018-0.

Takahashi:2005:DMP

[467] Yoshiaki Takahashi, Kiyotaka Takishita, Kazuhiko Koike, Tadashi Maruyama, Takeshi Nakayama, Atsushi Kobiyama, and Takehiko Ogata. Development of molecular probes for Dinophysis (Dinophyceae) plastid: a tool to predict blooming and explore plastid origin. Marine Biotechnology, 7(2):95–103, April 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.com/

article/10.1007/s10126-004-0482-5.

Poleo:2005:ISI

[468] Germán A. Poleo, Robert R. Godke, and Terrence R. Tiersch. Intracytoplasmic sperm injection using cryopreserved, fixed, and freezedried sperm in eggs of Nile tilapia. Marine Biotechnology, 7(2):104– 111, April 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-

2236 (electronic). URL https://link.springer.com/article/10.
1007/s10126-004-0162-5.

Kurokawa:2005:EFF

[469] Tatsuo Kurokawa and Shigeharu Tanisho. Effects of formate on fermentative hydrogen production by Enterobacter aerogenes. Marine Biotechnology, 7(2):112–118, April 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.com/

article/10.1007/s10126-004-3088-z.

Chiou:2005:MCC

[470] Tze-Ting Chiou, Jen-Leih Wu, Thomas T. Chen, and Jenn-Kan Lu. Molecular cloning and characterization of cDNA of penaeidin-like antimicrobial peptide from tiger shrimp (Penaeus monodon). Marine Biotechnology, 7(2):119–127, April 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.com/

article/10.1007/s10126-004-3164-4.

Sorrosa:2005:LTS

[471] Joy M. Sorrosa, Manami Satoh, and Yoshihiro Shiraiwa. Low temperature stimulates cell enlargement and intracellular calcification of coccolithophorids. Marine Biotechnology, 7(2):128–133, April 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-004-0478-1.

Wadman:2005:FAT

[472] Shannon A. Wadman, Karl J. Clark, and Perry B. Hackett. Fishing for answers with transposons. Marine Biotechnology, 7(3):135–141, June 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-004-0068-2.

Sipkema:2005:MSP

[473] Detmer Sipkema, Maurice C. R. Franssen, Ronald Osinga, Johannes Tramper, and René H. Wijffels. Marine sponges as pharmacy. Marine Biotechnology, 7(3):142–162, June 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-004-0405-5.

Yamakoshi:2005:GAT

[474] Kimi Yamakoshi, Yuji Shishido, and Nobuyoshi Shimoda. Generation of aberrant transcripts of and free DNA ends in zebrafish no tail gene. Marine Biotechnology, 7(3):163–172, June 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-004-3500-8.

Plisova:2005:HAA

[475] E. Yu Plisova, L. A. Balabanova, E. P. Ivanova, V. B. Kozhemyako, V. V. Mikhailov, E. V. Agafonova, and V. A. Rasskazov. A highly active alkaline phosphatase from the marine bacterium Cobetia. Marine Biotechnology, 7(3):173–178, June 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-004-3022-4.

Kanno:2005:ICT

[476] Manami Kanno, Qi Li, and Akihiro Kijima. Isolation and characterization of twenty microsatellite loci in Japanese sea cucumber (Stichopus japonicus). Marine Biotechnology, 7(3):179–183, June 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-004-0006-3.

Jain:2005:EPP

[477] Ruchi Jain, Seshagiri Raghukumar, R. Tharanathan, and N. B. Bhosle. Extracellular polysaccharide production by thraustochytrid protists. Marine Biotechnology, 7(3):184–192, June 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-004-4025-x.

Chen:2005:SAC

[478] S. H. Chen, C. Y. Lin, and C. M. Kuo. In Silico analysis of crustacean hyperglycemic hormone family. Marine Biotechnology, 7(3): 193–206, June 2005. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-0020-5.

Wang:2005:CGE

[479] Yongping Wang, Zhe Xu, James C. Pierce, and Ximing Guo. Characterization of Eastern oyster (Crassostrea virginica Gmelin) chromosomes by fluorescence in situ hybridization with bacteriophage P1 clones. Marine Biotechnology, 7(3):207–214, June 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-004-0051-y.

Kim:2005:DMI

[480] Choong-Jae Kim, Chang-Hoon Kim, and Yoshihiko Sako. Development of molecular identification method for genus Alexandrium (Dinophyceae) using whole-cell FISH. Marine Biotechnology, 7(3):215–222, June 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-004-0424-2.

Sotka:2005:GEP

[481] Erik E. Sotka, Jennifer A. Hempelmann, and Christiane H. Biermann. Genetic evidence of postglacial population expansion in Puget Sound rockfish (Sebastes emphaeus). Marine Biotechnology, 7(3):223–230, June 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-004-0437-x.

Yazawa:2005:IGJ

[482] Ryosuke Yazawa, Ikuo Hirono, Tsuyoshi Ohira, and Takashi Aoki. Induction of Japanese flounder TNF promoter activity by lipopolysaccharide in zebrafish embryo. Marine Biotechnology, 7(3):231–235, June 2005. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-004-0403-7.

Leignel:2005:SMM

[483] V. Leignel, Y. Hardivillier, and M. Laulier. Small metallothionein MT-10 genes in coastal and hydrothermal mussels. Marine Biotechnology, 7 (3):236–244, June 2005. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-0135-8.

Thakur:2005:AAC

[484] Archana N. Thakur, Narsinh L. Thakur, Madhavi M. Indap, Reena A. Pandit, Vrushali V. Datar, and Werner E. G. Müller. Antiangiogenic, antimicrobial, and cytotoxic potential of sponge-associated bacteria. Marine Biotechnology, 7(3):245–252, June 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-004-4085-y.

Nichols:2005:BEE

[485] C. A. Mancuso Nichols, J. Guezennec, and J. P. Bowman. Bacterial exopolysaccharides from extreme marine environments with special consideration of the Southern Ocean, sea ice, and deep-sea hydrothermal vents: a review. Marine Biotechnology, 7(4):253–271, August 2005. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-004-5118-2.

Koyama:2005:SDS

[486] Sumihiro Koyama, Takahiko Nagahama, Noriyuki Ootsu, Tomoji Takayama, Masae Horii, Satoshi Konishi, Tetsuya Miwa, Yoichi Ishikawa, and Masuo Aizawa. Survival of deep-sea shrimp (Alvinocaris sp.) during decompression and larval hatching at atmospheric pressure. Marine Biotechnology, 7(4):272–278, August 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-004-3050-0.

Wang:2005:CEG

[487] Yongjun Wang, Shicui Zhang, Zhenhui Liu, Hongyan Li, and Lei Wang. Characterization and expression of AmphiCL encoding cathepsin L proteinase from amphioxus Branchiostoma belcheri tsingtauense. Marine Biotechnology, 7(4):279–286, August 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-004-4084-9.

Zhang:2005:CCG

[488] Xiaohui Zhang, Xuecheng Zhang, Yoshihiro Shiraiwa, Yunxiang Mao, Zhenghong Sui, and Jinjie Liu. Cloning and characterization of hoxH genes from Arthrospira and Spirulina and application in phylogenetic study. Marine Biotechnology, 7(4):287–296, August 2005. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-004-3127-9.

Hellio:2005:IEG

[489] Claire Hellio, Maria Tsoukatou, Jean-Philippe Maréchal, Nick Aldred, Claude Beaupoil, Anthony S. Clare, Constantinos Vagias, and Vassilios Roussis. Inhibitory effects of Mediterranean sponge extracts and metabolites on larval settlement of the barnacle Balanus amphitrite. Marine Biotechnology, 7(4):297–305, August 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-004-3150-x.

Klinbunga:2005:MTC

[490] S. Klinbunga, B. Khamnamtong, N. Puanglarp, P. Jarayabhand, W. Yoosukh, and P. Menasveta. Molecular taxonomy of cupped oysters (Crassostrea, Saccostrea, and Striostrea) in Thailand based on COI, 16S, and 18S rDNA polymorphism. Marine Biotechnology, 7(4): 306–317, August 2005. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-0036-x.

McCombie:2005:CMP

[491] Helen McCombie, Christophe Ledu, Pascal Phelipot, Sylvie Lapègue, Pierre Boudry, and André Gérard. A complementary method for production of tetraploid Crassostrea gigas using crosses between diploids and tetraploids with cytochalasin B treatments. Marine Biotechnology, 7(4): 318–330, August 2005. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-0440-2.

Yim:2005:NSP

[492] Joung Han Yim, Eunwha Son, Suhkneung Pyo, and Hong Kum Lee. Novel sulfated polysaccharide derived from red-tide microalga Gyrodinium impudicum strain KG03 with immunostimulating activity in vivo. Marine Biotechnology, 7(4):331–338, August 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-004-0404-6.

Miller:2005:CMD

[493] Adam D. Miller, Nicholas P. Murphy, Christopher P. Burridge, and Christopher M. Austin. Complete mitochondrial DNA sequences of the decapod crustaceans Pseudocarcinus gigas (Menippidae) and Macrobrachium rosenbergii (Palaemonidae). Marine Biotechnology, 7(4): 339–349, August 2005. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-4077-8.

Cho:2005:SST

[494] Young Sun Cho, Buyl Nim Choi, En-Mi Ha, Ki Hong Kim, Sung Koo Kim, Dong Soo Kim, and Yoon Kwon Nam. Shark (Scyliorhinus torazame) metallothionein: cDNA cloning, genomic sequence, and expression analysis. Marine Biotechnology, 7(4):350–362, August 2005. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-004-0043-y.

Beacham:2005:PSS

[495] Terry D. Beacham, Douglas E. Hay, and Khai D. Le. Population structure and stock identification of eulachon (Thaleichthys pacificus), an anadromous smelt, in the Pacific northwest. Marine Biotechnology, 7 (4):363–372, August 2005. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-4075-0.

An:2005:PAS

[496] Hye-Suck An, Young-Ju Jee, Kwang-Sik Min, Bong-Lae Kim, and Seok jung Han. Phylogenetic analysis of six species of Pacific abalone (Haliotidae) based on DNA sequences of 16s rRNA and cytochrome c oxidase subunit I mitochondrial genes. Marine Biotechnology, 7(4): 373–380, August 2005. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-4405-2.

Carvalho:2005:OFA

[497] Ana P. Carvalho and F. Xavier Malcata. Optimization of ω-3 fatty acid production by microalgae: crossover effects of CO_{2} and light intensity under batch and continuous cultivation modes. Marine Biotechnology, 7 (4):381–388, August 2005. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-4047-4.

Bejar:2005:PSC

[498] Julia Béjar, Javier Porta, J. José Borrego, and M. Carmen Alvarez. The piscine SAF-1 cell line: genetic stability and labeling. Marine Biotechnology, 7(4):389–395, August 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-004-4083-0.

Lopez:2005:PAP

[499] José Luis López, Silvia Lorenzo Abalde, and José Fuentes. Proteomic approach to probe for larval proteins of the mussel Mytilus galloprovincialis. Marine Biotechnology, 7(4):396–404, August 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-004-4029-6.

McFadden:2005:VCN

[500] Catherine S. McFadden, Ian D. Tullis, M. Breton Hutchinson, Katherine Winner, and Jill A. Sohm. Variation in coding (NADH dehydrogenase subunits 2, 3, and 6) and noncoding intergenic spacer regions of the mitochondrial genome in Octocorallia (Cnidaria: Anthozoa). Marine Biotechnology, 7(4):405–406, August 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-005-1000-0.

Iwamoto:2005:SRM

[501] Koji Iwamoto and Yoshihiro Shiraiwa. Salt-regulated mannitol metabolism in algae. Marine Biotechnology, 7(5):407–415, October 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-005-0029-4.

Naganuma:2005:BEG

[502] Takeshi Naganuma, Hosam E. Elsaied, Daiki Hoshii, and Hiroyuki Kimura. Bacterial endosymbioses of gutless tube-dwelling worms in nonhydrothermal vent habitats. Marine Biotechnology, 7(5):416–428, October 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-004-5089-3.

Rinkevich:2005:MIC

[503] Baruch Rinkevich. Marine invertebrate cell cultures: New millennium trends. Marine Biotechnology, 7(5):429–439, October 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-004-0108-y.

Corley-Smith:2005:ASS

[504] Graham E. Corley-Smith, Liv Wennerberg, Joy A. Schembri, Chinten J. Lim, Karen L. Cooper, and Bruce P. Brandhorst. Assignment of sockeye salmon (Oncorhynchus nerka) to spawning sites using DNA markers. Marine Biotechnology, 7(5):440–448, October 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-004-4065-z.

Schroth:2005:MBA

[505] Werner Schroth, Andrea Ender, and Bernd Schierwater. Molecular biomarkers and adaptation to environmental stress in moon jelly (Aurelia spp.). Marine Biotechnology, 7(5):449–461, October 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-004-4095-9.

Diniz:2005:HDM

[506] Fabio M. Diniz, Norman Maclean, Masayoshi Ogawa, Israel H. A. Cintra, and Paul Bentzen. The hypervariable domain of the mitochondrial control region in Atlantic spiny lobsters and its potential as a marker for investigating phylogeographic structuring. Marine Biotechnology, 7 (5):462–473, October 2005. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-4062-5.

Yada:2005:ASP

[507] Etsuko Yada, Hiroyuki Nagata, Yukinori Noguchi, Yoh Kodera, Hiroyuki Nishimura, Yuji Inada, and Ayako Matsushima. An arginine specific protease from Spirulina platensis. Marine Biotechnology, 7(5): 474–480, October 2005. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-4115-9.

Lidie:2005:GEG

[508] Kristy B. Lidie, James C. Ryan, Michele Barbier, and Frances M. Van Dolah. Gene expression in Florida red tide dinoflagellate Karenia brevis: Analysis of an expressed sequence tag library and development of DNA microarray. Marine Biotechnology, 7(5):481–493, October 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-004-4110-6.

Hermann:2005:EAZ

[509] Andrea C. Hermann and Carol H. Kim. Effects of arsenic on zebrafish innate immune system. Marine Biotechnology, 7(5):494–505, October 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-004-4109-7.

Hosoi-Tanabe:2005:SSD

[510] Shoko Hosoi-Tanabe and Yoshihiko Sako. Species-specific detection and quantification of toxic marine dinoflagellates Alexandrium tamarense and A. catenella by real-time PCR assay. Marine Biotechnology, 7(5): 506–514, October 2005. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-4128-4.

Choi:2005:CCK

[511] Seon kang Choi, Yasuhiro Nishida, Satoru Matsuda, Kyoko Adachi, Hiroaki Kasai, Xue Peng, Sadao Komemushi, Wataru Miki, and Norihiko Misawa. Characterization of β-carotene ketolases, CrtW, from marine bacteria by complementation analysis in Escherichia coli. Marine Biotechnology, 7(5):515–522, October 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-004-5100-z.

Mitova:2005:EPG

[512] Maya Mitova, Maria Luisa Tutino, Giuseppe Infusini, Gennaro Marino, and Salvatore De Rosa. Exocellular peptides from Antarctic psychrophile Pseudoalteromonas haloplanktis. Marine Biotechnology, 7(5): 523–531, October 2005. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-5098-2.

Lapidot:2005:DPM

[513] Ziva Lapidot and Baruch Rinkevich. Development of panel of monoclonal antibodies specific to urochordate cell surface antigens. Marine Biotechnology, 7(5):532–539, October 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-004-5067-9.

Garver:2005:ADV

[514] Kyle A. Garver, Carla M. Conway, Diane G. Elliott, and Gael Kurath. Analysis of DNA-vaccinated fish reveals viral antigen in muscle, kidney and thymus, and transient histopathologic changes. Marine Biotechnology, 7(5):540–553, October 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-004-5129-z.

Perez:2005:DES

[515] Franklin Pérez, Juan Ortiz, Mariuxi Zhinaula, Cesar Gonzabay, Jorge Calderón, and Filip A. M. J. Volckaert. Development of EST–SSR markers by data mining in three species of shrimp: Litopenaeus vannamei, Litopenaeus stylirostris, and Trachypenaeus birdy. Marine Biotechnology, 7(5):554–569, October 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.com/

article/10.1007/s10126-004-5099-1.

Aranishi:2005:RPR

[516] Futoshi Aranishi. Rapid PCR–RFLP method for discrimination of imported and domestic mackerel. Marine Biotechnology, 7(6):571– 575, December 2005. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-4102-1.

Johansen:2005:QEA

[517] Katherine A. Johansen and Ken Overturf. Quantitative expression analysis of genes affecting muscle growth during development of rainbow trout (Oncorhynchus mykiss). Marine Biotechnology, 7(6):576– 587, December 2005. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-5133-3.

Zhang:2005:GVW

[518] Qian Zhang, Standish K. Allen, Jr., and Kimberly S. Reece. Genetic variation in wild and hatchery stocks of suminoe oyster (Crassostrea ariakensis) assessed by PCR–RFLP and microsatellite markers. Marine Biotechnology, 7(6):588–599, December 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-004-5105-7.

Liu:2005:MCE

[519] Fengsong Liu, Yichen Liu, Fuhua Li, Bo Dong, and Jianhai Xiang. Molecular cloning and expression profile of putative antilipopolysaccharide factor in Chinese shrimp (Fenneropenaeus chinensis). Marine Biotechnology, 7(6):600–608, December 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-005-5006-4.

Buonocore:2005:BAS

[520] Francesco Buonocore, Maria Forlenza, Elisa Randelli, Stefania Benedetti, Paola Bossù, Sabrina Meloni, Christopher J. Secombes, Massimo Mazzini, and Giuseppe Scapigliati. Biological activity of sea bass (Dicentrarchus labrax L.) recombinant interleukin-1β. Marine Biotechnology, 7(6): 609–617, December 2005. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-5131-5.

Podsiadlowski:2005:OMG

[521] Lars Podsiadlowski and Thomas Bartolomaeus. Organization of the mitochondrial genome of mantis shrimp Pseudosquilla ciliata (Crustacea: Stomatopoda). Marine Biotechnology, 7(6):618–624, December 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-005-0017-8.

Yazawa:2005:CPA

[522] Ryosuke Yazawa, Ikuo Hirono, and Takashi Aoki. Characterization of promoter activities of four different Japanese flounder promoters in transgenic zebrafish. Marine Biotechnology, 7(6):625–633, December 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-005-0011-1.

Hirokawa:2005:TTA

[523] Yasutaka Hirokawa, Shoko Fujiwara, and Mikio Tsuzuki. Three types of acidic polysaccharides associated with coccolith of Pleurochrysis haptonemofera: Comparison with Pleurochrysis carterae and analysis using fluorescein–isothiocyanate-labeled lectins. Marine Biotechnology, 7(6): 634–644, December 2005. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-5148-9.

Maynard:2005:MDS

[524] Ben T. Maynard, Lyndal J. Kerr, Joanne M. McKiernan, Eliza S. Jansen, and Peter J. Hanna. Mitochondrial DNA sequence and gene organization in the Australian blacklip abalone Haliotis rubra (Leach). Marine Biotechnology, 7(6):645–658, December 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-005-0013-z.

Wang:2005:PPD

[525] Yunling Wang, Tyler MacKenzie, and David Morse. Purification of plastids from the dinoflagellate Lingulodinium. Marine Biotechnology, 7(6): 659–668, December 2005. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-5126-2.

Li:2005:SMA

[526] Qi Li and Akihiro Kijima. Segregation of microsatellite alleles in gynogenetic diploid Pacific abalone (Haliotis discus hannai). Marine Biotechnology, 7(6):669–676, December 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-004-5119-1.

Abe:2005:SAC

[527] Shunnosuke Abe, Pi-Lin Wang, Fuminori Takahashi, and Eiji Sasaki. Structural analysis of cDNAs coding for 4SNc-Tudor domain protein from fish and their expression in yellowtail organs. Marine Biotechnology, 7(6): 677–686, December 2005. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-5137-z.

Bendezu:2005:IMS

[528] Ivan F. Bendezu, John W. Slater, and Brian F. Carney. Identification of Mytilus spp. and Pecten maximus in Irish waters by standard PCR of the 18S rDNA gene and multiplex PCR of the 16S rDNA gene. Marine Biotechnology, 7(6):687–696, December 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-004-0124-y.

Milbury:2005:CMD

[529] Coren A. Milbury and Patrick M. Gaffney. Complete mitochondrial DNA sequence of the Eastern oyster Crassostrea virginica. Marine Biotechnology, 7(6):697–712, December 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-005-0004-0.

Sato:2005:DMM

[530] Maremi Sato, Kenji Kawamata, Nadezhda Zaslavskaya, Aoi Nakamura, Tomoki Ohta, Takafumi Nishikiori, Vladimir Brykov, and Koji Nagashima. Development of microsatellite markers for Japanese scallop (Mizuhopecten yessoensis) and their application to a population genetic study. Marine Biotechnology, 7(6):713–728, December 2005. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-004-0127-8.

Honma:2006:PTS

[531] Tomohiro Honma and Kazuo Shiomi. Peptide toxins in sea anemones: Structural and functional aspects. Marine Biotechnology, 8(1):1–10, January 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-005-5093-2.

Barneah:2006:CPS

[532] O. Barneah, Y. Benayahu, and V. M. Weis. Comparative proteomics of symbiotic and aposymbiotic juvenile soft corals. Marine Biotechnology, 8(1):11–16, January 2006. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-004-5120-8.

Antoro:2006:SGD

[533] Suci Antoro, Uthairat Na-Nakorn, and Worawut Koedprang. Study of genetic diversity of orange-spotted grouper, Epinephelus coioides, from Thailand and Indonesia using microsatellite markers. Marine Biotechnology, 8(1):17–26, January 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-005-5026-0.

Descamps:2006:ICM

[534] Valérie Descamps, Sébastien Colin, Marc Lahaye, Murielle Jam, Christophe Richard, Philippe Potin, Tristan Barbeyron, Jean-Claude Yvin, and Bernard Kloareg. Isolation and culture of a marine bacterium degrading the sulfated fucans from marine brown algae. Marine Biotechnology, 8(1):27–39, January 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-005-5107-0.

Sipkema:2006:HKM

[535] Detmer Sipkema, Nejla A. M. Yosef, Marcin Adamczewski, Ronald Osinga, Dominick Mendola, Johannes Tramper, and René H. Wijffels. Hypothesized kinetic models for describing the growth of globular and encrusting demosponges. Marine Biotechnology, 8(1):40–51, January 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-005-5002-8.

Takeuchi:2006:BDC

[536] Takeshi Takeuchi and Kazuyoshi Endo. Biphasic and dually coordinated expression of the genes encoding major shell matrix proteins in the pearl oyster Pinctada fucata. Marine Biotechnology, 8(1):52–61, January 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-005-5037-x.

Acuna:2006:BCD

[537] Narciso Acuña, Benjamin Otto Ortega-Morales, and Alex Valadez-González. Biofilm colonization dynamics and its influence on the corrosion resistance of austenitic UNS S31603 stainless steel exposed to Gulf of Mexico seawater. Marine Biotechnology, 8(1):62–70, January 2006. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-005-5145-7.

Lin:2006:CMG

[538] G. Lin, L. C. Lo, Z. Y. Zhu, F. Feng, R. Chou, and G. H. Yue. The complete mitochondrial genome sequence and characterization of singlenucleotide polymorphisms in the control region of the Asian seabass (Lates calcarifer). Marine Biotechnology, 8(1):71–79, January 2006. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-005-5051-z.

Buonocore:2006:PCC

[539] Francesco Buonocore, Angelo Libertini, Daniela Prugnoli, Massimo Mazzini, and Giuseppe Scapigliati. Production and characterization of a continuous embryonic cell line from sea bass (Dicentrarchus labrax L.). Marine Biotechnology, 8(1):80–85, January 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-005-5032-2.

Kanda:2006:GCG

[540] N. Kanda, M. Goto, and L. A. Pastene. Genetic characteristics of Western North Pacific sei whales, Balaenoptera borealis, as revealed by microsatellites. Marine Biotechnology, 8(1):86–93, January 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-005-5130-1.

Kobayashi:2006:RAT

[541] Isao Kobayashi, Shoko Fujiwara, Hirotaka Saegusa, Masahiro Inouhe, Hiroko Matsumoto, and Mikio Tsuzuki. Relief of arsenate toxicity by Cd-stimulated phytochelatin synthesis in the green alga Chlamydomonas reinhardtii. Marine Biotechnology, 8(1):94–101, January 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-005-5092-3.

Robert:2006:PED

[542] Stanley S. Robert. Production of eicosapentaenoic and docosahexaenoic acid-containing oils in transgenic land plants for human and aquaculture nutrition. Marine Biotechnology, 8(2):103–109, April 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-005-5142-x.

Bryant:2006:ICE

[543] Maxine J. Bryant, Heather J. Flint, and Frank Y. T. Sin. Isolation, characterization, and expression analysis of three actin genes in the New Zealand black-footed abalone, Haliotis iris. Marine Biotechnology, 8 (2):110–119, April 2006. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-005-5139-5.

Park:2006:RPP

[544] Seunghye Park, Juergen E. W. Polle, Anastasios Melis, Taek Kyun Lee, and EonSeon Jin. Up-regulation of photoprotection and PSII-repair gene expression by irradiance in the unicellular green alga Dunaliella salina. Marine Biotechnology, 8(2):120–128, April 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-005-5030-4.

Devi:2006:BCD

[545] Prabha Devi, Chandrakant Govind Naik, and Celina Rodrigues. Biotransformation of citrinin to decarboxycitrinin using an organic solventtolerant marine bacterium, Moraxella sp. MB1. Marine Biotechnology, 8(2):129–138, April 2006. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-005-5021-5.

Zhang:2006:CMR

[546] Yuqing Zhang, Xungang Tan, Pei-Jun Zhang, and Yongli Xu. Characterization of muscle-regulatory gene, MyoD, from flounder (Paralichthys olivaceus) and analysis of its expression patterns during embryogenesis. Marine Biotechnology, 8(2):139–148, April 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-005-5042-0.

Goffredi:2006:MDM

[547] Shana K. Goffredi, William J. Jones, Christopher A. Scholin, Roman Marin, and Robert C. Vrijenhoek. Molecular detection of marine invertebrate larvae. Marine Biotechnology, 8(2):149–160, April 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-005-5016-2.

Song:2006:DES

[548] Linsheng Song, Wei Xu, Chenhua Li, Honglei Li, Longtao Wu, Jianhai Xiang, and Ximing Guo. Development of expressed sequence tags from the Bay scallop, Argopecten irradians irradians. Marine Biotechnology, 8(2):161–169, April 2006. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-005-0126-4.

Kumon:2006:NLI

[549] Yasuyuki Kumon, Rinka Yokoyama, Zinia Haque, Toshihiro Yokochi, Daisuke Honda, and Toro Nakahara. A new labyrinthulid isolate that produces only docosahexaenoic acid. Marine Biotechnology, 8(2): 170–177, April 2006. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-005-5098-x.

Ikeguchi:2006:ILM

[550] Koki Ikeguchi, Toshinao Ineno, Shiro Itoi, Hidehiro Kondo, Shigeharu Kinoshita, and Shugo Watabe. Increased levels of mitochondrial gene transcripts in the thermally selected rainbow trout (Oncorhynchus mykiss) strain during embryonic development. Marine Biotechnology, 8 (2):178–188, April 2006. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-005-5110-5.

DAmato:2006:DES

[551] María Eugenia D’Amato. Demographic expansion and subtle differentiation in the long-tailed hake Macruronus magellanicus: Evidence from microsatellite data. Marine Biotechnology, 8(2):189–201, April 2006. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-005-5075-4.

Kurita:2006:CCF

[552] Keisuke Kurita. Chitin and chitosan: Functional biopolymers from marine crustaceans. Marine Biotechnology, 8(3):203–226, June 2006. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-005-0097-5.

Yasumoto-Hirose:2006:IMB

[553] Mina Yasumoto-Hirose, Miyuki Nishijima, Metiek Kimie Ngirchechol, Kaneo Kanoh, Yoshikazu Shizuri, and Wataru Miki. Isolation of marine bacteria by in situ culture on media-supplemented polyurethane foam. Marine Biotechnology, 8(3):227–237, June 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-005-5015-3.

Lee:2006:ICX

[554] Mi-Young Lee, Byung-Sul Min, Chung-Soon Chang, and EonSeon Jin. Isolation and characterization of a xanthophyll aberrant mutant of the green alga Nannochloropsis oculata. Marine Biotechnology, 8(3): 238–245, June 2006. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-006-5078-9.

Macey:2006:CGT

[555] Brett M. Macey and Vernon E. Coyne. Colonization of the gastrointestinal tract of the farmed South African abalone Haliotis midae by the probionts Vibrio midae SY9 , Cryptococcus sp. SS1, and Debaryomyces hansenii AY1 . Marine Biotechnology, 8(3):246–259, June 2006. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-005-0113-9.

Chow:2006:MSI

[556] Seinen Chow, Nobuaki Suzuki, Hideyuki Imai, and Taku Yoshimura. Molecular species identification of spiny lobster phyllosoma larvae of the genus Panulirus from the northwestern Pacific. Marine Biotechnology, 8(3):260–267, June 2006. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-005-5151-9.

Karako-Lampert:2006:DGE

[557] S. Karako-Lampert, G. Hershkovits, N. Stambler, N. Simon-Blecher, Y. Achituv, Z. Dubinsky, and D. J. Katcoff. Differential gene expression in Symbiodinium microadriaticum clade B following stress. Marine Biotechnology, 8(3):268–274, June 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-005-5008-2.

Marchand:2006:MIE

[558] J. Marchand, A. Tanguy, G. Charrier, L. Quiniou, E. Plee-Gauthier, and J. Laroche. Molecular identification and expression of differentially regulated genes of the European flounder, Platichthys flesus, submitted to pesticide exposure. Marine Biotechnology, 8(3):275–294, June 2006. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-005-0099-3.

Ando:2006:ETS

[559] Hideki Ando and Hitoshi Okamoto. Efficient transfection strategy for the spatiotemporal control of gene expression in zebrafish. Marine Biotechnology, 8(3):295–303, June 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.com/

article/10.1007/s10126-005-5138-6.

Wu:2006:IDF

[560] Yuping Wu, Guangxian Zhang, Qian Xiong, Fang Luo, Caimei Cui, Wei Hu, Yanhong Yu, Jin Su, Anlong Xu, and Zuoyan Zhu. Integration of double-fluorescence expression vectors into zebrafish genome for the selection of site-directed knockout/knockin. Marine Biotechnology, 8 (3):304–311, June 2006. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-006-5116-7.

Zhu:2006:ECS

[561] Xiang-Ping Zhu, Feng You, Pei-Jun Zhang, Yong-Li Xu, and Jian-He Xu. Effects of cold shock on microtubule organization and cell cycle in gynogenetically activated eggs of olive flounder (Paralichthys olivaceus). Marine Biotechnology, 8(3):312–318, June 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-006-5128-3.

Morita:2006:DAP

[562] Eiko Morita, Yasuyuki Kumon, Toro Nakahara, Satoshi Kagiwada, and Tetsuko Noguchi. Docosahexaenoic acid production and lipid-body formation in Schizochytrium limacinum SR21 . Marine Biotechnology, 8 (3):319–327, June 2006. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-005-5060-y.

Dahm:2006:LSF

[563] Ralf Dahm and Robert Geisler. Learning from small fry: The zebrafish as a genetic model organism for aquaculture fish species. Marine Biotechnology, 8(4):329–345, August 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-006-5139-0.

Rodriguez:2006:PCR

[564] M. F. Rodriguez, S. A. Gahr, C. E. Rexroad, and Y. Palti. A polymerase chain reaction screening method for rapid detection of microsatellites in bacterial artificial chromosomes. Marine Biotechnology, 8(4):

346–350, August 2006. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-005-5064-7.

Garver:2006:ITS

[565] Kyle A. Garver, Carla M. Conway, and Gael Kurath. Introduction of translation stop codons into the viral glycoprotein gene in a fish DNA vaccine eliminates induction of protective immunity. Marine Biotechnology, 8(4):351–356, August 2006. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-005-5154-6.

Liu:2006:URR

[566] Jing Liu, Yong-Hua Sun, Na Wang, Ya-Ping Wang, and Zuo-Yan Zhu. Upstream regulatory region of zebrafish lunatic fringe: Isolation and promoter analysis. Marine Biotechnology, 8(4):357–365, August 2006. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-005-5125-y.

Diers:2006:ZMA

[567] Jeffrey A. Diers, John J. Bowling, Stephen O. Duke, Subagus Wahyuono, Michelle Kelly, and Mark T. Hamann. Zebra mussel antifouling activity of the marine natural product aaptamine and analogs. Marine Biotechnology, 8(4):366–372, August 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-005-6055-4.

Hoffmann:2006:MMC

[568] Friederike Hoffmann, Hans Tore Rapp, and Joachim Reitner. Monitoring microbial community composition by fluorescence In Situ hybridization during cultivation of the marine cold-water sponge Geodia barretti. Marine Biotechnology, 8(4):373–379, August 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-006-5152-3.

Blair:2006:MDP

[569] D. Blair, M. Waycott, L. Byrne, G. Dunshea, C. Smith-Keune, and K. M. Neil. Molecular discrimination of Perna (Mollusca: Bivalvia) species using the polymerase chain reaction and species-specific mitochondrial primers. Marine Biotechnology, 8(4):380–385, August 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-005-6121-y.

Liu:2006:PGL

[570] Xiande Liu, Xiao Liu, Ximing Guo, Qikang Gao, Hongen Zhao, and Guofan Zhang. A preliminary genetic linkage map of the Pacific abalone Haliotis discus hannai Ino. Marine Biotechnology, 8(4):386–397, August 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-005-6133-7.

Jorgensen:2006:VRG

[571] Sven Martin Jorgensen, Ellen Johanne Kleveland, Unni Grimholt, and Tor Gjoen. Validation of reference genes for real-time polymerase chain reaction studies in Atlantic salmon. Marine Biotechnology, 8(4): 398–408, August 2006. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-005-5164-4.

Nakai:2006:PRS

[572] Masaaki Nakai, Norihiko Kageyama, Koichi Nakahara, and Wataru Miki. Phlorotannins as radical scavengers from the extract of Sargassum ringgoldianum. Marine Biotechnology, 8(4):409–414, August 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-005-6168-9.

Rahman:2006:STC

[573] M. Azizur Rahman, Y. Isa, and T. Uehara. Studies on two closely related species of octocorallians: Biochemical and molecular characteristics of the organic matrices of endoskeletal sclerites. Marine Biotechnology, 8 (4):415–424, August 2006. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-005-6150-6.

Lee:2006:CEC

[574] H. S. Lee, Y. J. Kim, S. S. Bae, J. H. Jeon, J. K. Lim, B. C. Jeong, S. G. Kang, and J.-H. Lee. Cloning, expression, and characterization of a methionyl aminopeptidase from a hyperthermophilic archaeon Thermococcus sp. NA1. Marine Biotechnology, 8(4):425–432, August 2006. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-005-6124-8.

Laurent:2006:AMN

[575] Dominique Laurent and Francesco Pietra. Antiplasmodial marine natural products in the perspective of current chemotherapy and prevention of malaria. A review. Marine Biotechnology, 8(5):433–447, October 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-006-6100-y.

Tramice:2006:HYE

[576] Annabella Tramice, Assunta Giordano, Giuseppina Andreotti, Ernesto Mollo, and Antonio Trincone. High-yielding enzymatic α-glucosylation of pyridoxine by marine α-glucosidase from Aplysia fasciata. Marine Biotechnology, 8(5):448–452, October 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-005-6144-4.

Sekino:2006:SLA

[577] Masashi Sekino, Toshimasa Kobayashi, and Motoyuki Hara. Segregation and linkage analysis of 75 novel microsatellite DNA markers in pair crosses of Japanese abalone (Haliotis discus hannai) using the 5′-tailed primer method. Marine Biotechnology, 8(5):453–466, October 2006. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-005-6179-6.

Alcivar-Warren:2006:IMT

[578] Acacia Alcivar-Warren, Dawn Meehan-Meola, Yongping Wang, Ximing Guo, Linghua Zhou, Jianhai Xiang, Shaun Moss, Steve Arce, William Warren, Zhenkang Xu, and Kireina Bell. Isolation and mapping of telomeric pentanucleotide (TAACC)_{n} repeats of the Pacific whiteleg shrimp, Penaeus vannamei, using fluorescence in situ hybridization. Marine Biotechnology, 8(5):467–480, October 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-005-6031-z.

Kawamoto:2006:CSA

[579] Hitoshi Kawamoto, Akio Horibe, Yasunari Miki, Takayuki Kimura, Katsunori Tanaka, Tsuyoshi Nakagawa, Makoto Kawamukai, and Hideyuki Matsuda. Cloning and sequencing analysis of alginate lyase genes from the marine bacterium Vibrio sp. O2. Marine Biotechnology, 8(5): 481–490, October 2006. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-005-6157-z.

Wang:2006:DGR

[580] Bing Wang, Fuhua Li, Bo Dong, Xiaojun Zhang, Chengsong Zhang, and Jianhai Xiang. Discovery of the genes in response to white spot syndrome virus (WSSV) infection in Fenneropenaeus chinensis through cDNA microarray. Marine Biotechnology, 8(5):491–500, October 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-005-6136-4.

Yang:2006:TII

[581] Huiping Yang and Ximing Guo. Tetraploid induction by inhibiting mitosis I with heat shock, cold shock, and nocodazole in the hard clam Mercenaria mercenaria (Linnaeus, 1758). Marine Biotechnology, 8(5): 501–510, October 2006. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-005-6183-x.

Elsaied:2006:MCD

[582] Hosam Easa Elsaied, Ryo Kaneko, and Takeshi Naganuma. Molecular characterization of a deep-sea methanotrophic mussel symbiont that carries a RuBisCO gene. Marine Biotechnology, 8(5):511–520, October 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-005-6135-5.

Cunningham:2006:NRM

[583] Charles Cunningham, Jun ichi Hikima, Matthew J. Jenny, Robert W. Chapman, Guang-Chen Fang, Chris Saski, Mats L. Lundqvist, Rod A. Wing, Pauline M. Cupit, Paul S. Gross, Greg W. Warr, and Jeff P. Tomkins. New resources for marine genomics: Bacterial artificial chromosome libraries for the Eastern and Pacific oysters (Crassostrea virginica and C. gigas). Marine Biotechnology, 8(5):521–533, October 2006. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-006-6013-9.

LaClaire:2006:AES

[584] John W. La Claire II. Analysis of expressed sequence tags from the harmful alga, Prymnesium parvum (Prymnesiophyceae, Haptophyta). Marine Biotechnology, 8(5):534–546, October 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-005-5182-2.

Papakostas:2006:ICB

[585] Spiros Papakostas, Stefania Dooms, Marianna Christodoulou, Alexander Triantafyllidis, Ilias Kappas, Kristof Dierckens, Peter Bossier, Patrick Sorgeloos, and Theodore J. Abatzopoulos. Identification of cultured Brachionus rotifers based on RFLP and SSCP screening. Marine Biotechnology, 8(5):547–559, October 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-005-6181-z.

Luo:2006:VGF

[586] Wen-Xin Luo, Tong Cheng, Bao-Quan Guan, Shao-Wei Li, Ji Miao, Jun Zhang, and Ning-Shao Xia. Variants of green fluorescent protein GFPxm. Marine Biotechnology, 8(5):560–566, October 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-006-6006-8.

Maynard:2006:MDS

[587] Ben T. Maynard, Lyndal J. Kerr, Joanne M. McKiernan, Eliza S. Jansen, and Peter J. Hanna. Mitochondrial DNA sequence and gene organization in the Australian blacklip abalone Haliotis rubra (Leach). Marine Biotechnology, 8(5):567, October 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-005-5704-y.

Pietra:2006:ESC

[588] Francesco Pietra. Earning from the sea: The current status of the biotechnology of drugs from the sea. Marine Biotechnology, 8(6):569– 571, December 2006. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-006-6065-x.

Zhang:2006:MPO

[589] Cen Zhang and Rongqing Zhang. Matrix proteins in the outer shells of molluscs. Marine Biotechnology, 8(6):572–586, December 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-005-6029-6.

Jamal:2006:RCK

[590] Mamdoh T. Jamal, Peter C. Morris, Rasmus Hansen, Derek J. Jamieson, J. Grant Burgess, and Brian Austin. Recovery and characterization of a 30.7-kDa protein from Bacillus licheniformis associated with inhibitory activity against methicillin-resistant Staphylococcus aureus, vancomycinresistant enterococci, and Listeria monocytogenes. Marine Biotechnology, 8(6):587–592, December 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-005-6160-4.

Ta:2006:IEC

[591] Quang Van Ta, Moon-Moo Kim, and Se-Kwon Kim. Inhibitory effect of chitooligosaccharides on matrix metalloproteinase-9 in human fibrosarcoma cells (HT1080). Marine Biotechnology, 8(6):593–599, December 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-006-6031-7.

Zhang:2006:MHC

[592] Y. X. Zhang, S. L. Chen, Y. G. Liu, Z. X. Sha, and Z. J. Liu. Major histocompatibility complex class IIB allele polymorphism and its association with resistance/susceptibility to Vibrio anguillarum in Japanese flounder (Paralichthys olivaceus). Marine Biotechnology, 8(6): 600–610, December 2006. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-005-6185-8.

Zhang:2006:MIP

[593] Yu-Xi Zhang and Song-Lin Chen. Molecular identification, polymorphism, and expression analysis of major histocompatibility complex class IIA and B genes of turbot (Scophthalmus maximus). Marine Biotechnology, 8(6):611–623, December 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-005-6174-y.

Zhang:2006:NMP

[594] Cen Zhang, Shuo Li, Zhuojun Ma, Liping Xie, and Rongqing Zhang. A novel matrix protein p10 from the nacre of pearl oyster (Pinctada fucata) and its effects on both CaCO_{3} crystal formation and mineralogenic cells. Marine Biotechnology, 8(6):624–633, December 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-006-6037-1.

Kwong:2006:NAA

[595] Theresa Fuk Ning Kwong, Li Miao, Xiancui Li, and Pei Yuan Qian. Novel antifouling and antimicrobial compound from a marine-derived fungus Ampelomyces sp. Marine Biotechnology, 8(6):634–640, December 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-005-6146-2.

Bryson:2006:CLH

[596] Steve P. Bryson, Elizabeth M. Joyce, D. John Martell, Lucy E. J. Lee, Shawn E. Holt, Steve C. Kales, Kazuhiro Fujiki, Brian Dixon, and Niels C. Bols. A cell line (HEW) from embryos of haddock (Melanogrammus aeglefinius) and its capacity to tolerate environmental extremes. Marine Biotechnology, 8(6):641–653, December 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-005-6163-1.

Woo:2006:HMI

[597] Seonock Woo, Seungshic Yum, Jee Hyun Jung, Won Joon Shim, Chang-Hoon Lee, and Taek-Kyun Lee. Heavy metal-induced differential gene expression of metallothionein in Javanese medaka, Oryzias javanicus. Marine Biotechnology, 8(6):654–662, December 2006. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-006-6046-0.

Deane:2006:TDE

[598] Eddie E. Deane and Norman Y. S. Woo. Tissue distribution, effects of salinity acclimation, and ontogeny of aquaporin 3 in the marine teleost, silver sea bream (Sparus sarba). Marine Biotechnology, 8(6): 663–671, December 2006. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-006-6001-0.

Kanno:2006:MAG

[599] Manami Kanno, Yoshihisa Suyama, Qi Li, and Akihiro Kijima. Microsatellite analysis of Japanese sea cucumber, Stichopus (Apostichopus) japonicus, supports reproductive isolation in color variants. Marine Biotechnology, 8(6):672–685, December 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-006-6014-8.

Passmore:2006:DDB

[600] A. J. Passmore, S. N. Jarman, K. M. Swadling, S. Kawaguchi, A. McMinn, and S. Nicol. DNA as a dietary biomarker in Antarctic krill, Euphausia superba. Marine Biotechnology, 8(6):686–696, December 2006. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-005-6088-8.

Anonymous:2006:VNE

[601] Anonymous. Volume 8, number 5 errata for Alcivar–Warren et al. and Elsaied et al. Marine Biotechnology, 8(6):697–698, December 2006. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-006-6802-1.

Anonymous:2007:GSM

[602] Anonymous. Growing and strengthening marine genomics. Marine Biotechnology, 9(1):1, January 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-007-0900-6.

Erpenbeck:2007:SPS

[603] Dirk Erpenbeck and Rob W. M. van Soest. Status and perspective of sponge chemosystematics. Marine Biotechnology, 9(1):2–19, January 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-005-6109-7.

Ko:2007:MCM

[604] Chi-Fong Ko, Tzu-Ting Chiou, Thomas T. Chen, Jen-Leih Wu, Jiann-Chu Chen, and Jenn-Kan Lu. Molecular cloning of myostatin gene and characterization of tissue-specific and developmental stage-specific expression of the gene in orange spotted grouper, Epinephelus coioides. Marine Biotechnology, 9(1):20–32, January 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-006-6059-8.

Burridge:2007:PGS

[605] Christopher P. Burridge and Vincent L. Versace. Population genetic structuring in Acanthopagrus butcheri (Pisces: Sparidae): Does low gene flow among estuaries apply to both sexes? Marine Biotechnology, 9 (1):33–44, January 2007. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-006-6023-7.

Cao:2007:MRR

[606] Jun-Xia Cao, Jie-Qiong Dai, Zhong-Min Dai, Guo-Li Yin, and Wei-Jun Yang. A male reproduction-related kazal-type peptidase inhibitor gene in the prawn, Macrobrachium rosenbergii: Molecular characterization and expression patterns. Marine Biotechnology, 9(1):45–55, January 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-006-6026-4.

Takayanagi:2007:PFC

[607] Takatoshi Takayanagi, Yasutaka Hirokawa, Maki Yamamoto, Toshichika Ohki, Shoko Fujiwara, and Mikio Tsuzuki. Protoplast formation of the coccolithophorid Pleurochrysis haptonemofera in hypoosmotic K^{+} solution: Shedding of the coccosphere and regrowth of the protoplast in normal medium. Marine Biotechnology, 9(1):56–65, January 2007. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-006-6058-9.

Qin:2007:IMA

[608] Yanjie Qin, Xiao Liu, Haibin Zhang, Guofan Zhang, and Ximing Guo. Identification and mapping of amplified fragment length polymorphism markers linked to shell color in bay scallop, Argopecten irradians irradians (Lamarck, 1819). Marine Biotechnology, 9(1):66–73, January 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-006-6076-7.

Hara:2007:GDB

[609] Motoyuki Hara and Masashi Sekino. Genetic differences between hatchery stocks and natural populations in Pacific abalone (Haliotis discus) estimated using microsatellite DNA markers. Marine Biotechnology, 9 (1):74–81, January 2007. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-006-6060-2.

Chen:2007:PCC

[610] Song-Lin Chen, Zhen-Xia Sha, Han-Qing Ye, Yang Liu, Yong-Sheng Tian, Yunhan Hong, and Qi-Sheng Tang. Pluripotency and chimera competence of an embryonic stem cell line from the sea perch (Lateolabrax japonicus). Marine Biotechnology, 9(1):82–91, January 2007. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-006-6050-1.

Alimuddin:2007:EMS

[611] Alimuddin, Goro Yoshizaki, Viswanath Kiron, Shuichi Satoh, and Toshio Takeuchi. Expression of masu salmon Δ 5-desaturase-like gene elevated EPA and DHA biosynthesis in zebrafish. Marine Biotechnology, 9(1): 92–100, January 2007. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-006-6003-y.

Yoshino:2007:HPH

[612] Fuminori Yoshino, Hiroshi Ikeda, Hajime Masukawa, and Hidehiro Sakurai. High photobiological hydrogen production activity of a Nostoc sp. PCC 7422 uptake hydrogenase-deficient mutant with high nitrogenase activity. Marine Biotechnology, 9(1):101–112, January 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-006-6035-3.

Arnaud-Haond:2007:PFP

[613] S. Arnaud-Haond, E. Goyard, V. Vonau, C. Herbaut, J. Prou, and D. Saulnier. Pearl formation: Persistence of the graft during the entire process of biomineralization. Marine Biotechnology, 9(1):113– 116, January 2007. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-006-6033-5.

Alvarez:2007:FEC

[614] M. Carmen Alvarez, Julia Béjar, Songlin Chen, and Yunhan Hong. Fish ES cells and applications to biotechnology. Marine Biotechnology, 9 (2):117–127, March 2007. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-006-6034-4.

Sakurai:2007:PRP

[615] Hidehiro Sakurai and Hajime Masukawa. Promoting R&D in photobiological hydrogen production utilizing mariculture-raised cyanobacteria. Marine Biotechnology, 9(2):128–145, March 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-006-6073-x.

Robert:2007:MMS

[616] Stanley Robert, Maged P. Mansour, and Susan I. Blackburn. Metolachlormediated selection of a microalgal strain producing novel polyunsaturated fatty acids. Marine Biotechnology, 9(2):146–153, March 2007. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-006-6102-9.

Hsiao:2007:IFA

[617] Tracy Y. Hsiao, Bradley Holmes, and Harvey W. Blanch. Identification and functional analysis of a Delta-6 desaturase from the marine microalga Glossomastix chrysoplasta. Marine Biotechnology, 9(2): 154–165, March 2007. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-006-6075-8.

Hoover:2007:PTC

[618] Cindi A. Hoover, Marc Slattery, and Adam G. Marsh. Profiling transcriptome complexity and secondary metabolite synthesis in a benthic soft coral, Sinularia polydactyla. Marine Biotechnology, 9(2): 166–178, March 2007. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-006-6048-y.

Moriya:2007:GSI

[619] Shogo Moriya, Shunpei Sato, Tomonori Azumaya, Osamu Suzuki, Shigehiko Urawa, Akihisa Urano, and Syuiti Abe. Genetic stock identification of chum salmon in the Bering Sea and North Pacific Ocean using mitochondrial DNA microarray. Marine Biotechnology, 9(2): 179–191, March 2007. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-006-6079-4.

Shimonaga:2007:VSP

[620] Takahiro Shimonaga, Shoko Fujiwara, Miki Kaneko, Asako Izumo, Satoko Nihei, Perigio B. Francisco, Aya Satoh, Naoko Fujita, Yasunori Nakamura, and Mikio Tsuzuki. Variation in storage α-polyglucans of red algae: Amylose and semi-amylopectin types in Porphyridium and glycogen type in Cyanidium. Marine Biotechnology, 9(2):192–202, March 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-006-6104-7.

Cho:2007:PGS

[621] Eun-Seob Cho, Choon-Goon Jung, Sang-Gyu Sohn, Chul-Won Kim, and Seock-Jung Han. Population genetic structure of the ark shell Scapharca broughtonii schrenck from Korea, China, and Russia based on COI gene sequences. Marine Biotechnology, 9(2):203–216, March 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-006-6057-x.

Muttray:2007:IDI

[622] Annette F. Muttray, Rachel L. Cox, Carol L. Reinisch, and Susan A. Baldwin. Identification of DeltaN isoform and polyadenylation site choice variants in molluscan p63/p73 -like homologues. Marine Biotechnology, 9(2):217–230, March 2007. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-006-6045-1.

Wang:2007:ASS

[623] Shi Wang, Zhenmin Bao, Ning Li, Lingling Zhang, and Jingjie Hu. Analysis of the secondary structure of ITS1 in Pectinidae: Implications for phylogenetic reconstruction and structural evolution. Marine Biotechnology, 9(2):231–242, March 2007. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-006-6113-6.

Tsai:2007:ITO

[624] C. L. Tsai, L. H. Wang, Y. L. Shiue, and T. Y. Chao. Influence of temperature on the ontogenetic expression of neural development-related genes from developing tilapia brain expressed sequence tags. Marine Biotechnology, 9(2):243–261, March 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-006-6089-2.

Ye:2007:MCE

[625] Han-Qing Ye, Song-Lin Chen, Zhen-Xia Sha, and Yang Liu. Molecular cloning and expression analysis of the myostatin gene in sea perch (Lateolabrax japonicus). Marine Biotechnology, 9(2):262–272, March 2007. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-006-6093-6.

Chen:2007:IFS

[626] Song-Lin Chen, Jing Li, Si-Ping Deng, Yong-Sheng Tian, Qing-Yin Wang, Zhi-Meng Zhuang, Zhen-Xia Sha, and Jian-Yong Xu. Isolation of female-specific AFLP markers and molecular identification of genetic sex in half-smooth tongue sole (Cynoglossus semilaevis). Marine Biotechnology, 9(2):273–280, March 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-006-6081-x.

Parameswaran:2007:DCT

[627] V. Parameswaran, V. P. Ishaq Ahmed, Ravi Shukla, R. R. Bhonde, and A. S. Sahul Hameed. Development and characterization of two new cell lines from milkfish (Chanos chanos) and grouper (Epinephelus coioides) for virus isolation. Marine Biotechnology, 9(2):281–291, March 2007. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-006-6110-9.

Marx:2007:CAE

[628] J-C. Marx, T. Collins, S. D’Amico, G. Feller, and C. Gerday. Coldadapted enzymes from marine Antarctic microorganisms. Marine Biotechnology, 9(3):293–304, May 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-006-6103-8.

Mocz:2007:FPT

[629] Gabor Mocz. Fluorescent proteins and their use in marine biosciences, biotechnology, and proteomics. Marine Biotechnology, 9(3):305–328, May 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-006-7145-7.

Liu:2007:IEA

[630] Jifang Liu, Shaojun Liu, Min Tao, Wei Li, and Yun Liu. Isolation and expression analysis of testicular type sox9b in allotetraploid fish. Marine Biotechnology, 9(3):329–334, May 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-006-6123-4.

Ken:2007:RBC

[631] Chuian-Fu Ken, Chi-Tsai Lin, Yu-Der Wen, and Jen-Leih Wu. Replacement of buried cysteine from zebrafish Cu/Zn superoxide dismutase and enhancement of its stability via site-directed mutagenesis. Marine Biotechnology, 9(3):335–342, May 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-006-0143-y.

Ma:2007:PCA

[632] Chunling Ma, Xiumei Ni, Zhenming Chi, Liyan Ma, and Lingmei Gao. Purification and characterization of an alkaline protease from the marine yeast Aureobasidium pullulans for bioactive peptide production from different sources. Marine Biotechnology, 9(3):343–351, May 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-006-6105-6.

Mahidol:2007:MDD

[633] Chulabhorn Mahidol, Uthairat Na-Nakorn, Srijanya Sukmanomon, Nobuhiko Taniguchi, and Thuy T. T. Nguyen. Mitochondrial DNA diversity of the Asian Moon scallop, Amusium pleuronectes (Pectinidae), in Thailand. Marine Biotechnology, 9(3):352–359, May 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-006-6137-y.

Tsutsui:2007:PSG

[634] Naoaki Tsutsui, Tsuyoshi Ohira, Ichiro Kawazoe, Akiyoshi Takahashi, and Marcy N. Wilder. Purification of sinus gland peptides having vitellogenesis-inhibiting activity from the whiteleg shrimp Litopenaeus vannamei. Marine Biotechnology, 9(3):360–369, May 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-006-6151-0.

Zhou:2007:ECV

[635] Guang-Zhou Zhou, Zheng-Qiu Li, Xiu-Ping Yuan, and Qi-Ya Zhang. Establishment, characterization, and virus susceptibility of a new marine cell line from red spotted grouper (Epinephelus akaara). Marine Biotechnology, 9(3):370–376, May 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-006-7165-3.

Sellars:2007:PVL

[636] Melony J. Sellars, Russell E. Lyons, Peter M. Grewe, Tony Vuocolo, Lisa Leeton, Greg J. Coman, Bernard M. Degnan, and Nigel P. Preston. A PL10 vasa-like gene in the kuruma shrimp, Marsupenaeus japonicus, expressed during development and in adult gonad. Marine Biotechnology, 9(3):377–387, May 2007. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-007-6118-9.

Dobretsov:2007:NAI

[637] Sergey Dobretsov, Hairong Xiong, Ying Xu, Lisa A. Levin, and Pei-Yuan Qian. Novel antifoulants: Inhibition of larval attachment by proteases. Marine Biotechnology, 9(3):388–397, May 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-007-7091-z.

Qian:2007:MBM

[638] P.-Y. Qian, S. C. K. Lau, H.-U. Dahms, S. Dobretsov, and T. Harder. Marine biofilms as mediators of colonization by marine macroorganisms: Implications for antifouling and aquaculture. Marine Biotechnology, 9 (4):399–410, August 2007. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-007-9001-9.

Hoover:2007:FAT

[639] Cindi A. Hoover, Marc Slattery, and Adam G. Marsh. A functional approach to transcriptome profiling: Linking gene expression patterns to metabolites that matter. Marine Biotechnology, 9(4):411–419, August 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-007-9008-2.

Liu:2007:SDG

[640] Wei yi Liu, Yun Wang, Yao Qin, Ya ping Wang, and Zuo yan Zhu. Sitedirected gene integration in transgenic zebrafish mediated by Cre recombinase using a combination of mutant Lox sites. Marine Biotechnology, 9 (4):420–428, August 2007. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-007-9000-x.

Kaneko:2007:PDD

[641] Ryo Kaneko, Toru Hayashi, Manabu Tanahashi, and Takeshi Naganuma. Phylogenetic diversity and distribution of dissimilatory sulfite reductase genes from deep-sea sediment cores. Marine Biotechnology, 9(4): 429–436, August 2007. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-007-9003-7.

Bedouet:2007:HPI

[642] Laurent Bédouet, Denis Duplat, Arul Marie, Lionel Dubost, Sophie Berland, Marthe Rousseau, Christian Milet, and Evelyne Lopez. Heterogeneity of proteinase inhibitors in the water-soluble organic matrix from the oyster nacre. Marine Biotechnology, 9(4):437–449, August 2007. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-007-7120-y.

Cho:2007:CDH

[643] Y. Cho, H. S. Lee, Y. J. Kim, S. G. Kang, S.-J. Kim, and J.-H. Lee. Characterization of a dUTPase from the hyperthermophilic archaeon Thermococcus onnurineus NA1 and its application in polymerase chain reaction amplification. Marine Biotechnology, 9(4):450–458, August 2007. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-007-9002-8.

Lo:2007:MCR

[644] Ting Sze Lo, Zhaoxia Cui, Janice L. Y. Mong, Queenie W. L. Wong, Siu-Ming Chan, Hoi Shan Kwan, and Ka Hou Chu. Molecular coordinated regulation of gene expression during ovarian development in the penaeid shrimp. Marine Biotechnology, 9(4):459–468, August 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-007-9006-4.

Ip:2007:CNO

[645] Denis Tsz-Ming Ip, Kam-Bo Wong, and David Chi-Cheong Wan. Characterization of novel orange fluorescent protein cloned from cnidarian tube anemone Cerianthus sp. Marine Biotechnology, 9(4):469–478, August 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-007-9005-5.

Karawita:2007:PEE

[646] Rohan Karawita, Mahinda Senevirathne, Yasantha Athukorala, Abu Affan, Young-Jae Lee, Se-Kwon Kim, Joon-Baek Lee, and You-Jin Jeon. Protective effect of enzymatic extracts from microalgae against DNA damage induced by H_{2}O_{2}. Marine Biotechnology, 9(4):479–490, August 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-007-9007-3.

Cohen:2007:EEF

[647] Raphael Cohen, Vered Chalifa-Caspi, Timothy D. Williams, Meirav Auslander, Stephen G. George, James K. Chipman, and Moshe Tom. Estimating the efficiency of fish cross-species cDNA microarray hybridization. Marine Biotechnology, 9(4):491–499, August 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-007-9010-8.

Wang:2007:DCE

[648] Yongping Wang and Ximing Guo. Development and characterization of EST–SSR markers in the Eastern oyster Crassostrea virginica. Marine Biotechnology, 9(4):500–511, August 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-007-9011-7.

Blanchette:2007:MGT

[649] Brian Blanchette, Xia Feng, and Bal Ram Singh. Marine glutathione S-transferases. Marine Biotechnology, 9(5):513–542, October 2007. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-007-9034-0.

Steindler:2007:DGE

[650] Laura Steindler, Silvia Schuster, Micha Ilan, Adi Avni, Carlo Cerrano, and Sven Beer. Differential gene expression in a marine sponge in relation to its symbiotic state. Marine Biotechnology, 9(5):543– 549, October 2007. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-007-9024-2.

Fujiwara:2007:GEP

[651] Shoko Fujiwara, Yasutaka Hirokawa, Yukiko Takatsuka, Kunihiro Suda, Erika Asamizu, Takatoshi Takayanagi, Daisuke Shibata, Satoshi Tabata, and Mikio Tsuzuki. Gene expression profiling of coccolith-bearing cells and naked cells in haptophyte Pleurochrysis haptonemofera with a cDNA macroarray system. Marine Biotechnology, 9(5):550–560, October 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-007-9039-8.

Bruck:2007:DBC

[652] Thomas B. Brück, Wolfram M. Brück, Lory Z. Santiago-Vázquez, Peter J. McCarthy, and Russell G. Kerr. Diversity of the bacterial communities associated with the azooxanthellate deep water octocorals Leptogorgia minimata, Iciligorgia schrammi, and Swiftia exertia. Marine Biotechnology, 9(5):561–576, October 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-007-9009-1.

Jenny:2007:CMC

[653] Matthew J. Jenny, Robert W. Chapman, Annalaura Mancia, Yian A. Chen, David J. McKillen, Hal Trent, Paul Lang, Jean-Michel Escoubas, Evelyne Bachere, Viviane Boulo, Z. John Liu, Paul S. Gross, Charles Cunningham, Pauline M. Cupit, Arnaud Tanguy, Ximing Guo, Dario Moraga, Isabelle Boutet, Arnaud Huvet, Sylvain De Guise, Jonas S. Almeida, and Gregory W. Warr. A cDNA microarray for Crassostrea virginica and C. gigas. Marine Biotechnology, 9(5):577–591, October 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-007-9041-1.

deCaralt:2007:CSL

[654] Sònia de Caralt, Henri Otjens, María J. Uriz, and René H. Wijffels. Cultivation of sponge larvae: Settlement, survival, and growth of juveniles. Marine Biotechnology, 9(5):592–605, October 2007. CODEN MABIFW.

ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-007-9013-5.

Zhang:2007:FLC

[655] Lingling Zhang, Zhenmin Bao, Jie Cheng, Hui Li, Xiaoting Huang, Shi Wang, Can Zhang, and Jingjie Hu. Fosmid library construction and initial analysis of end sequences in Zhikong scallop (Chlamys farreri). Marine Biotechnology, 9(5):606–612, October 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-007-9014-4.

Xie:2007:NGA

[656] Li-Ping Xie, Yuan-Tai Wu, Yi-Ping Dai, Qing Li, and Rong-Qing Zhang. A novel glycosylphosphatidylinositol-anchored alkaline phosphatase dwells in the hepatic duct of the pearl oyster, Pinctada fucata. Marine Biotechnology, 9(5):613–623, October 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-007-9015-3.

Loongyai:2007:IFC

[657] Wiriya Loongyai, Jean-Christophe Avarre, Martine Cerutti, Esther Lubzens, and Wilaiwan Chotigeat. Isolation and functional characterization of a new shrimp ovarian peritrophin with antimicrobial activity from Fenneropenaeus merguiensis. Marine Biotechnology, 9(5): 624–637, October 2007. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-007-9019-z.

Bedouet:2007:PAN

[658] Laurent Bédouet, Arul Marie, Lionel Dubost, Jean Péduzzi, Denis Duplat, Sophie Berland, Marion Puisségur, Hélène Boulzaguet, Marthe Rousseau, Christian Milet, and Evelyne Lopez. Proteomics analysis of the nacre soluble and insoluble proteins from the oyster Pinctada margaritifera. Marine Biotechnology, 9(5):638–649, October 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-007-9017-1.

Jing:2007:IAP

[659] Gu Jing, Zhenguang Yan, Yu Li, Liping Xie, and Rongqing Zhang. Immunolocalization of an acid phosphatase from pearl oyster (Pinctada fucata) and its In Vitro effects on calcium carbonate crystal formation. Marine Biotechnology, 9(5):650–659, October 2007. CODEN MABIFW.

ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-007-9018-0.

Silverman:2007:UMM

[660] Heather G. Silverman and Francisco F. Roberto. Understanding marine mussel adhesion. Marine Biotechnology, 9(6):661–681, December 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-007-9053-x.

Guo:2007:PRG

[661] Yusong Guo, Zhongduo Wang, Chuwu Liu, Li Liu, and Yun Liu. Phylogenetic relationships of South China Sea snappers (genus Lutjanus; family Lutjanidae) based on mitochondrial DNA sequences. Marine Biotechnology, 9(6):682–688, December 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-007-9012-6.

Chiu:2007:DPA

[662] Kuo-Hsun Chiu, Hurng-Wern Huang, and Hin-Kiu Mok. Differential proteome analysis of hagfish dental and somatic skeletal muscles. Marine Biotechnology, 9(6):689–700, December 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-007-9020-6.

Wang:2007:CBL

[663] Shaolin Wang, Peng Xu, Jim Thorsen, Baoli Zhu, Pieter J. de Jong, Geoff Waldbieser, Huseyin Kucuktas, and Zhanjiang Liu. Characterization of a BAC library from channel catfish Ictalurus punctatus: Indications of high levels of chromosomal reshuffling among teleost genomes. Marine Biotechnology, 9(6):701–711, December 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-007-9021-5.

Lyons:2007:CAM

[664] R. E. Lyons, L. M. Dierens, S. H. Tan, N. P. Preston, and Y. Li. Characterization of AFLP markers associated with growth in the kuruma prawn, Marsupenaeus japonicus, and identification of a candidate gene. Marine Biotechnology, 9(6):712–721, December 2007. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-007-9022-4.

Santaclara:2007:MDX

[665] Francisco J. Santaclara, Montserrat Espiñeira, and Juan M. Vieites. Molecular detection of Xenostrobus securis and Mytillus galloprovincialis larvae in Galician Coast (Spain). Marine Biotechnology, 9(6): 722–732, December 2007. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-007-9023-3.

Kao:2007:DRF

[666] Hung-Teh Kao, Shelby Sturgis, Rob DeSalle, Julia Tsai, Douglas Davis, David F. Gruber, and Vincent A. Pieribone. Dynamic regulation of fluorescent proteins from a single species of coral. Marine Biotechnology, 9(6):733–746, December 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-007-9025-1.

John:2007:PGS

[667] David E. John, Stacey S. Patterson, and John H. Paul. Phytoplanktongroup specific quantitative polymerase chain reaction assays for Ru-BisCO mRNA transcripts in seawater. Marine Biotechnology, 9(6): 747–759, December 2007. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-007-9027-z.

Iguchi:2007:AIT

[668] Akira Iguchi, Luis M. Márquez, Brent Knack, Chuya Shinzato, Madeleine J. H. van Oppen, Bette L. Willis, Kate Hardie, Julian Catmull, and David J. Miller. Apparent involvement of a β1 type integrin in coral fertilization. Marine Biotechnology, 9(6):760–765, December 2007. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-007-9026-0.

Parameswaran:2007:DPL

[669] V. Parameswaran, Ravi Shukla, Ramesh Bhonde, and A. S. Sahul Hameed. Development of a pluripotent ES-like cell line from Asian sea bass (Lates calcarifer) — an oviparous stem cell line mimicking viviparous ES cells. Marine Biotechnology, 9(6):766–775, December 2007. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-007-9028-y.

Ortlepp:2007:AAB

[670] Sofia Ortlepp, Martin Sjögren, Mia Dahlström, Horst Weber, Rainer Ebel, RuAngelie Edrada, Carsten Thoms, Peter Schupp, Lars Bohlin, and Peter Proksch. Antifouling activity of bromotyrosine-derived sponge metabolites and synthetic analogues. Marine Biotechnology, 9(6):776– 785, December 2007. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-007-9029-x.

Tinh:2008:RFP

[671] Nguyen Thi Ngoc Tinh, Kristof Dierckens, Patrick Sorgeloos, and Peter Bossier. A review of the functionality of probiotics in the larviculture food chain. Marine Biotechnology, 10(1):1–12, January 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-007-9054-9.

Takishita:2008:MII

[672] Kiyotaka Takishita, Yoshihiro Fujiwara, Masaru Kawato, Natsuki Kakizoe, Masayuki Miyazaki, and Tadashi Maruyama. Molecular identification of the ichthyosporean protist “ Pseudoperkinsus tapetis “ from the mytilid mussel Adipicola pacifica associated with submerged whale carcasses in Japan. Marine Biotechnology, 10(1):13–18, January 2008. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-007-9032-2.

Yoon:2008:MCG

[673] Sung Ho Yoon, Yasuo Itoh, Gen Kaneko, Makiko Nakaniwa, Masataka Ohta, and Shugo Watabe. Molecular characterization of Japanese sillago vitellogenin and changes in its expression levels on exposure to 17β-estradiol and 4-tert-octylphenol. Marine Biotechnology, 10(1): 19–30, January 2008. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-007-9055-8.

Rahman:2008:CAC

[674] M. Azizur Rahman, Tamotsu Oomori, and Tsuyoshi Uehara. Carbonic anhydrase in calcified endoskeleton: Novel activity in biocalcification in alcyonarian. Marine Biotechnology, 10(1):31–38, January 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-007-9030-4.

Azuma:2008:GVP

[675] Noriko Azuma, Yasushi Kunihiro, Jun Sasaki, Eiji Mihara, Yukio Mihara, Tomoaki Yasunaga, Deuk-Hee Jin, and Syuiti Abe. Genetic variation and population structure of hair crab (Erimacrus isenbeckii) in Japan inferred from mitochondrial DNA sequence analysis. Marine Biotechnology, 10(1):39–48, January 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-007-9033-1.

Suzuki:2008:MDA

[676] Nobuaki Suzuki, Kouichi Hoshino, Keisuke Murakami, Haruko Takeyama, and Seinen Chow. Molecular diet analysis of phyllosoma larvae of the Japanese spiny lobster Panulirus japonicus (Decapoda: Crustacea). Marine Biotechnology, 10(1):49–55, January 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-007-9038-9.

Skindersoe:2008:QSA

[677] Mette Elena Skindersoe, Piers Ettinger-Epstein, Thomas Bovbjerg Rasmussen, Thomas Bjarnsholt, Rocky de Nys, and Michael Givskov. Quorum sensing antagonism from marine organisms. Marine Biotechnology, 10(1):56–63, January 2008. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-007-9036-y.

Ettinger-Epstein:2008:PMA

[678] Piers Ettinger-Epstein, Dianne M. Tapiolas, Cherie A. Motti, Anthony D. Wright, Christopher N. Battershill, and Rocky de Nys. Production of manoalide and its analogues by the sponge Luffariella variabilis is hardwired. Marine Biotechnology, 10(1):64–74, January 2008. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-007-9037-x.

Zhang:2008:PCC

[679] Jin-Wei Zhang and Run-Ying Zeng. Purification and characterization of a cold-adapted α-amylase produced by Nocardiopsis sp. 7326 isolated from Prydz Bay, Antarctic. Marine Biotechnology, 10(1): 75–82, January 2008. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-007-9035-z.

Li:2008:MCM

[680] Yiyun Li, Mingyi Cai, ZhiYong Wang, Wei Guo, Xiande Liu, Xiaoqing Wang, and Yue Ning. Microsatellite–centromere mapping in large yellow croaker (Pseudosciaena crocea) using gynogenetic diploid families. Marine Biotechnology, 10(1):83–90, January 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-007-9040-2.

Wong:2008:DGE

[681] Queenie W. L. Wong, Wai Yan Mak, and Ka Hou Chu. Differential gene expression in hepatopancreas of the shrimp Metapenaeus ensis during ovarian maturation. Marine Biotechnology, 10(1):91–98, January 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-007-9042-0.

Holmes:2008:PTT

[682] Bradley Holmes and Harvey Blanch. Possible taxonomic trends in the success of primary aggregate formation in marine sponge cell cultures. Marine Biotechnology, 10(1):99–109, January 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-007-9047-8.

Kamino:2008:UAM

[683] Kei Kamino. Underwater adhesive of marine organisms as the vital link between biological science and material science. Marine Biotechnology, 10 (2):111–121, March 2008. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-007-9076-3.

Henrichs:2008:PAM

[684] D. W. Henrichs, M. A. Renshaw, C. A. Santamaria, B. Richardson, J. R. Gold, and L. Campbell. PCR amplification of microsatellites from single cells of Karenia brevis preserved in Lugol’s iodine solution. Marine Biotechnology, 10(2):122–127, March 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-007-9044-y.

Yada:2008:ICT

[685] Shuichi Yada, Yi Wang, Yanshuang Zou, Keiko Nagasaki, Kakushi Hosokawa, Issey Osaka, Ryuichi Arakawa, and Keiich Enomoto. Isolation and characterization of two groups of novel marine bacteria producing violacein. Marine Biotechnology, 10(2):128–132, March 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-007-9046-9.

Miyaguchi:2008:RTP

[686] Hideo Miyaguchi, Norio Kurosawa, and Tatsuki Toda. Real-time polymerase chain reaction assays for rapid detection and quantification of Noctiluca scintillans zoospore. Marine Biotechnology, 10(2):133–140, March 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-007-9031-3.

Oyamada:2008:MLA

[687] Chiaki Oyamada, Masaki Kaneniwa, Koji Ebitani, Masakazu Murata, and Kenji Ishihara. Mycosporine-like amino acids extracted from scallop (Patinopecten yessoensis) ovaries: UV protection and growth stimulation activities on human cells. Marine Biotechnology, 10(2):141–150, March 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-007-9043-z.

Zhang:2008:FMI

[688] Lingling Zhang, Zhenmin Bao, Shi Wang, Xiaoli Hu, and Jingjie Hu. FISH mapping and identification of Zhikong scallop (Chlamys farreri) chromosomes. Marine Biotechnology, 10(2):151–157, March 2008. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-007-9045-x.

Soares:2008:IAM

[689] Angélica R. Soares, Bernardo A. P. da Gama, Andrea P. da Cunha, Valéria L. Teixeira, and Renato C. Pereira. Induction of attachment of the mussel Perna perna by natural products from the brown seaweed Stypopodium zonale. Marine Biotechnology, 10(2):158–165, March 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-007-9048-7.

Smith-Keune:2008:GEG

[690] C. Smith-Keune and S. Dove. Gene expression of a green fluorescent protein homolog as a host-specific biomarker of heat stress within a reefbuilding coral. Marine Biotechnology, 10(2):166–180, March 2008. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-007-9049-6.

Zamora-Castro:2008:NAB

[691] J. Zamora-Castro, J. Paniagua-Michel, and C. Lezama-Cervantes. A novel approach for bioremediation of a coastal marine wastewater effluent based on artificial microbial mats. Marine Biotechnology, 10(2): 181–189, March 2008. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-007-9050-0.

Li:2008:PCE

[692] Xiaoyu Li, Zhenming Chi, Zhiqiang Liu, Jing Li, Xianghong Wang, and Nalini Yasoda Hirimuthugoda. Purification and characterization of extracellular phytase from a marine yeast Kodamaea ohmeri BG3 . Marine Biotechnology, 10(2):190–197, March 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-007-9051-z.

Sarathi:2008:SVG

[693] M. Sarathi, Martin C. Simon, V. P. Ishaq Ahmed, S. Rajesh Kumar, and A. S. Sahul Hameed. Silencing VP28 gene of white spot syndrome virus of shrimp by bacterially expressed dsRNA. Marine Biotechnology, 10 (2):198–206, March 2008. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-007-9052-y.

Kochzius:2008:DMI

[694] M. Kochzius, M. Nölte, H. Weber, N. Silkenbeumer, S. Hjörleifsdottir, G. O. Hreggvidsson, V. Marteinsson, K. Kappel, S. Planes, F. Tinti, A. Magoulas, E. Garcia Vazquez, C. Turan, C. Hervet, D. Campo Falgueras, A. Antoniou, M. Landi, and D. Blohm. DNA microarrays for identifying fishes. Marine Biotechnology, 10(2):207–217, March 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-007-9068-3.

Sun:2008:IGP

[695] Guohua Sun, Xuecheng Zhang, Zhenghong Sui, and Yunxiang Mao. Inhibition of pds gene expression via the RNA interference approach in Dunaliella salina (Chlorophyta). Marine Biotechnology, 10(3):219– 226, May 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-

2236 (electronic). URL https://link.springer.com/article/10.
1007/s10126-007-9056-7.

Sarropoulou:2008:LGN

[696] E. Sarropoulou, D. Nousdili, A. Magoulas, and G. Kotoulas. Linking the genomes of nonmodel teleosts through comparative genomics. Marine Biotechnology, 10(3):227–233, May 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-007-9066-5.

Norizuki:2008:DFN

[697] Michiko Norizuki and Testuro Samata. Distribution and function of the nacrein-related proteins inferred from structural analysis. Marine Biotechnology, 10(3):234–241, May 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-007-9061-x.

Sarathi:2008:OAB

[698] M. Sarathi, Martin C. Simon, C. Venkatesan, and A. S. Sahul Hameed. Oral administration of bacterially expressed VP28dsRNA to protect Penaeus monodon from white spot syndrome virus. Marine Biotechnology, 10(3):242–249, May 2008. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-007-9057-6.

Auslander:2008:PAF

[699] M. Auslander, Y. Yudkovski, V. Chalifa-Caspi, B. Herut, R. Ophir, R. Reinhardt, P. M. Neumann, and M. Tom. Pollution-affected fish hepatic transcriptome and its expression patterns on exposure to cadmium. Marine Biotechnology, 10(3):250–261, May 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-007-9060-y.

Su:2008:CPD

[700] Jianguo Su, Zuoyan Zhu, Yaping Wang, Feng Xiong, and Jun Zou. The cytomegalovirus promoter-driven short hairpin RNA constructs mediate effective RNA interference in zebrafish In Vivo. Marine Biotechnology, 10(3):262–269, May 2008. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-007-9059-4.

Palsson:2008:REI

[701] Snæbjörn Pálsson, Jonas Paulsen, and Einar Árnason. Rapid evolution of the intergenic T–P spacer in the mtDNA of Arctic cod Arctogadus glacialis. Marine Biotechnology, 10(3):270–277, May 2008. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-007-9058-5.

Ahn:2008:IEE

[702] Ginnae Ahn, Insun Hwang, Eunjin Park, Jinhe Kim, You-Jin Jeon, Jehee Lee, Jae Woo Park, and Youngheun Jee. Immunomodulatory effects of an enzymatic extract from Ecklonia cava on murine splenocytes. Marine Biotechnology, 10(3):278–289, May 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-007-9062-9.

Takakura:2008:ICL

[703] Daisuke Takakura, Michiko Norizuki, Fumio Ishikawa, and Tetsuro Samata. Isolation and characterization of the N-linked oligosaccharides in nacrein from Pinctada fucata. Marine Biotechnology, 10(3): 290–296, May 2008. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-007-9063-8.

Reusch:2008:CAE

[704] Thorsten B. H. Reusch, Amelie S. Veron, Christoph Preuss, January Weiner, Lothar Wissler, Alfred Beck, Sven Klages, Michael Kube, Richard Reinhardt, and Erich Bornberg-Bauer. Comparative analysis of expressed sequence tag (EST) libraries in the seagrass Zostera marina subjected to temperature stress. Marine Biotechnology, 10 (3):297–309, May 2008. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-007-9065-6.

Zhong:2008:EPV

[705] Qiwang Zhong, Quanqi Zhang, Zhigang Wang, Jie Qi, Yanjie Chen, Shuo Li, Yeying Sun, Chunmei Li, and Xun Lan. Expression profiling and validation of potential reference genes during Paralichthys olivaceus embryogenesis. Marine Biotechnology, 10(3):310–318, May 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-007-9064-7.

Ni:2008:CCE

[706] Xiumei Ni, Zhenming Chi, Chunling Ma, and Catherine Madzak. Cloning, characterization, and expression of the gene encoding alkaline protease in the marine yeast Aureobasidium pullulans 10. Marine Biotechnology, 10(3):319–327, May 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-007-9067-4.

Schnitzler:2008:SDF

[707] Christine E. Schnitzler, Robert J. Keenan, Robert McCord, Artur Matysik, Lynne M. Christianson, and Steven H. D. Haddock. Spectral diversity of fluorescent proteins from the Anthozoan Corynactis californica. Marine Biotechnology, 10(3):328–342, May 2008. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-007-9072-7.

Vago:2008:CBT

[708] Razi Vago. Cnidarians biomineral in tissue engineering: a review. Marine Biotechnology, 10(4):343–349, July 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-008-9103-z.

Amar:2008:EAF

[709] Keren-Or Amar, Jacob Douek, Claudette Rabinowitz, and Baruch Rinkevich. Employing of the amplified fragment length polymorphism (AFLP) methodology as an efficient population genetic tool for symbiotic cnidarians. Marine Biotechnology, 10(4):350–357, July 2008. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-007-9069-2.

Zhang:2008:CCT

[710] Yang Zhang, Xiaojun Zhang, Chantel F. Scheuring, Hong-Bin Zhang, Pin Huan, Fuhua Li, and Jianhai Xiang. Construction and characterization of two bacterial artificial chromosome libraries of Zhikong scallop, Chlamys farreri Jones et Preston, and identification of BAC clones containing the genes involved in its innate immune system. Marine Biotechnology, 10 (4):358–365, July 2008. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-007-9071-8.

Hwang:2008:SEE

[711] Young-Ok Hwang, Sung Gyun Kang, Jung-Hee Woo, Kye Kyung Kwon, Takako Sato, Eun Yeol Lee, Myong Soo Han, and Sang-Jin Kim. Screening enantioselective epoxide hydrolase activities from marine microorganisms: Detection of activities in Erythrobacter spp. Marine Biotechnology, 10(4):366–373, July 2008. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-007-9070-9.

Ongvarrasopone:2008:SPD

[712] Chalermporn Ongvarrasopone, Mayuree Chanasakulniyom, Kallaya Sritunyalucksana, and Sakol Panyim. Suppression of PmRab7 by dsRNA inhibits WSSV or YHV infection in shrimp. Marine Biotechnology, 10 (4):374–381, July 2008. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-007-9073-6.

Nishihara:2008:NPP

[713] Masaaki Nishihara, Masazumi Kamata, Tomoyuki Koyama, and Kazunaga Yazawa. New phospholipase A_{1}-producing bacteria from a marine fish. Marine Biotechnology, 10(4):382–387, July 2008. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-007-9074-5.

Wynne:2008:TAA

[714] James W. Wynne, Maree G. O’Sullivan, Mathew T. Cook, Glenn Stone, Barbara F. Nowak, David R. Lovell, and Nicholas G. Elliott. Transcriptome analyses of amoebic gill disease-affected Atlantic salmon (Salmo salar) tissues reveal localized host gene suppression. Marine Biotechnology, 10(4):388–403, July 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-007-9075-4.

Zijffers:2008:DPA

[715] Jan-Willem F. Zijffers, Marcel Janssen, Johannes Tramper, and René H. Wijffels. Design process of an area-efficient photobioreactor. Marine Biotechnology, 10(4):404–415, July 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-007-9077-2.

Darias:2008:GEP

[716] M. J. Darias, J. L. Zambonino-Infante, K. Hugot, C. L. Cahu, and D. Mazurais. Gene expression patterns during the larval development of European sea bass (Dicentrarchus labrax) by microarray analysis. Marine Biotechnology, 10(4):416–428, July 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-007-9078-1.

Johnson:2008:SAM

[717] Nathan A. Johnson, Roger L. Vallejo, Jeffrey T. Silverstein, Timothy J. Welch, Gregory D. Wiens, Eric M. Hallerman, and Yniv Palti. Suggestive association of major histocompatibility IB genetic markers with resistance to bacterial cold water disease in rainbow trout (Oncorhynchus mykiss). Marine Biotechnology, 10(4):429–437, July 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-007-9080-7.

Zhou:2008:EDE

[718] Qian Zhou, Changgong Wu, Bo Dong, Fengqi Liu, and Jianhai Xiang. The encysted dormant embryo proteome of Artemia sinica. Marine Biotechnology, 10(4):438–446, July 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-007-9079-0.

Cho:2008:ISS

[719] Hyun Kook Cho, Bo-Hye Nam, Hee Jeong Kong, Hyon Sob Han, Young Baek Hur, Tae Jin Choi, Yung Hyun Choi, Woo Jin Kim, and JaeHun Cheong. Identification of softness syndrome-associated candidate genes and DNA sequence variation in the sea squirt, Halocynthia roretzi. Marine Biotechnology, 10(4):447–456, July 2008. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-008-9084-y.

Gong:2008:CCD

[720] Ningping Gong, Zhuojun Ma, Qing Li, Qi Li, Zhenguang Yan, Liping Xie, and Rongqing Zhang. Characterization of calcium deposition and shell matrix protein secretion in primary mantle tissue culture from the marine pearl oyster Pinctada fucata. Marine Biotechnology, 10 (4):457–465, July 2008. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9081-1.

Kiseleva:2008:EDG

[721] Marina I. Kiseleva, Larissa A. Balabanova, Valery A. Rasskazov, and Tatiana N. Zvyagintseva. Effect of 1,3;1,6-β-D-glucans on developing sea urchin embryos. Marine Biotechnology, 10(4):466–470, July 2008. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-008-9082-0.

De:2008:DTH

[722] Jaysankar De, N. Ramaiah, and L. Vardanyan. Detoxification of toxic heavy metals by marine bacteria highly resistant to mercury. Marine Biotechnology, 10(4):471–477, July 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-008-9083-z.

vonWegner:2008:RMP

[723] Frederic von Wegner, Sumihiro Koyama, Tetsuya Miwa, and Oliver Friedrich. Resting membrane potentials recorded on-site in intact skeletal muscles from deep sea fish (Sigmops gracile) salvaged from depths up to 1.000 m. Marine Biotechnology, 10(4):478–486, July 2008. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-008-9085-x.

vonSchalburg:2008:RSO

[724] K. R. von Schalburg, J. Leong, G. A. Cooper, A. Robb, M. R. Beetz-Sargent, R. Lieph, R. A. Holt, R. Moore, K. V. Ewart, W. R. Driedzic, B. F. H. ten Hallers, B. Zhu, P. J. de Jong, W. S. Davidson, and B. F. Koop. Rainbow smelt (Osmerus mordax) genomic library and EST resources. Marine Biotechnology, 10(5):487–491, September 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-008-9089-6.

Perez-Guzman:2008:BAT

[725] Lumarie Pérez-Guzmán, Ana E. Pérez-Matos, William Rosado, Thomas R. Tosteson, and Nadathur S. Govind. Bacteria associated with toxic clonal cultures of the dinoflagellate Ostreopsis lenticularis. Marine Biotechnology, 10(5):492–496, September 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-008-9088-7.

Pan:2008:OER

[726] Xiufang Pan, Huiqing Zhan, and Zhiyuan Gong. Ornamental expression of red fluorescent protein in transgenic founders of white skirt tetra (Gymnocorymbus ternetzi). Marine Biotechnology, 10(5):497– 501, September 2008. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9094-9.

Koopmans:2008:SGR

[727] Marieke Koopmans and René H. Wijffels. Seasonal growth rate of the sponge Haliclona oculata (Demospongiae: Haplosclerida). Marine Biotechnology, 10(5):502–510, September 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-008-9086-9.

Su:2008:HCU

[728] Jianguo Su, Zuoyan Zhu, Feng Xiong, and Yaping Wang. Hybrid cytomegalovirus-u6 promoter-based plasmid vectors improve efficiency of RNA interference in zebrafish. Marine Biotechnology, 10(5):511– 517, September 2008. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9087-8.

Sarkar:2008:EPA

[729] Sreyashi Sarkar, Malay Saha, Debashis Roy, Parasuraman Jaisankar, Satadal Das, Lalita Gauri Roy, Ratan Gachhui, Tuhinadri Sen, and Joydeep Mukherjee. Enhanced production of antimicrobial compounds by three salt-tolerant actinobacterial strains isolated from the sundarbans in a niche-mimic bioreactor. Marine Biotechnology, 10(5):518– 526, September 2008. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9090-0.

Li:2008:CCG

[730] Hao Li, Jinlai Miao, Fengxia Cui, and Guangyou Li. Characterization of cupric glutamate extinguishing mechanism of Alexandrium sp. LC3 with two-dimensional electrophoresis and MALDI–TOF MS. Marine Biotechnology, 10(5):527–537, September 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-008-9091-z.

Chen:2008:UMF

[731] Chienhsun Chen, Chih-Yung Chiou, Chang-Feng Dai, and Chaolun Allen Chen. Unique mitogenomic features in the scleractinian family pocilloporidae (Scleractinia: Astrocoeniina). Marine Biotechnology, 10(5): 538–553, September 2008. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9093-x.

Yoshida:2008:ACS

[732] Akihiro Yoshida, Yuna Seo, Shuhei Suzuki, Tomohiko Nishino, Takeshi Kobayashi, Naoko Hamada-Sato, Kazuhiro Kogure, and Chiaki Imada. Actinomycetal community structures in seawater and freshwater examined by DGGE analysis of 16S rRNA gene fragments. Marine Biotechnology, 10(5):554–563, September 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-008-9092-y.

Nakachi:2008:IGC

[733] Mia Nakachi, Midori Matsumoto, Philip M. Terry, Ronald L. Cerny, and Hideaki Moriyama. Identification of guanylate cyclases and related signaling proteins in sperm tail from sea stars by mass spectrometry. Marine Biotechnology, 10(5):564–571, September 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-008-9096-7.

Frenz-Ross:2008:EBT

[734] Jamie L. Frenz-Ross, Julie J. Enticknap, and Russell G. Kerr. The effect of bleaching on the terpene chemistry of Plexaurella fusifera: Evidence that zooxanthellae are not responsible for sesquiterpene production. Marine Biotechnology, 10(5):572–578, September 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-008-9095-8.

Haidle:2008:DQT

[735] Lisa Haidle, Jennifer E. Janssen, Karim Gharbi, Hooman K. Moghadam, Moira M. Ferguson, and Roy G. Danzmann. Determination of quantitative trait loci (QTL) for early maturation in rainbow trout (Oncorhynchus mykiss). Marine Biotechnology, 10(5):579–592, September 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-008-9098-5.

Suphamungmee:2008:MBA

[736] Worawit Suphamungmee, Wattana Weerachatyanukul, Tanes Poomtong, Peter Hanna, and Prasert Sobhon. Morphological and biochemical alterations of abalone testicular germ cells and spawned sperm and their fertilizing ability. Marine Biotechnology, 10(5):593–601, September 2008. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-008-9097-6.

Coesel:2008:NLM

[737] Sacha Nicole Coesel, Alexandra Cordeiro Baumgartner, Licia Marlene Teles, Ana Alexandra Ramos, Nuno Miguel Henriques, Leonor Cancela, and João Carlos Serafim Varela. Nutrient limitation is the main regulatory factor for carotenoid accumulation and for Psy and Pds steady state transcript levels in Dunaliella salina (Chlorophyta) exposed to high light and salt stress. Marine Biotechnology, 10(5):602– 611, September 2008. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9100-2.

Zhang:2008:MCE

[738] Jin wei Zhang and Run ying Zeng. Molecular cloning and expression of a cold-adapted lipase gene from an Antarctic deep sea psychrotrophic bacterium Pseudomonas sp. 7323. Marine Biotechnology, 10(5):612– 621, September 2008. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9099-4.

Mendola:2008:EFR

[739] Dominick Mendola, Sonia de Caralt, Maria J. Uriz, Fred van den End, Johan L. Van Leeuwen, and René H. Wijffels. Environmental flow regimes for Dysidea avara sponges. Marine Biotechnology, 10(5):622– 630, September 2008. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9102-0.

Raghukumar:2008:TMP

[740] Seshagiri Raghukumar. Thraustochytrid marine protists: Production of PUFAs and other emerging technologies. Marine Biotechnology, 10(6): 631–640, December 2008. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9135-4.

Patil:2008:SPA

[741] Jawahar G. Patil and Susan J. Hinze. Simplex PCR assay for positive identification of genetic sex in the Japanese medaka, Oryzias latipes. Marine Biotechnology, 10(6):641–644, December 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-008-9106-9.

Pan:2008:DRA

[742] Jian yi Pan, Yue ling Zhang, San ying Wang, and Xuan xian Peng. Dodecamer is required for agglutination of Litopenaeus vannamei hemocyanin with bacterial cells and red blood cells. Marine Biotechnology, 10(6): 645–652, December 2008. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9115-8.

Moya:2008:CUC

[743] Aurélie Moya, Sylvie Tambutté, Guillaume Béranger, Béatrice Gaume, Jean-Claude Scimeca, Denis Allemand, and Didier Zoccola. Cloning and use of a coral 36B4 gene to study the differential expression of coral genes between light and dark conditions. Marine Biotechnology, 10(6): 653–663, December 2008. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9101-1.

Wang:2008:CGE

[744] Bing Wang, Fuhua Li, Wei Luan, Yusu Xie, Chengsong Zhang, Zhan Luo, Lang Gui, Hui Yan, and Jianhai Xiang. Comparison of gene expression profiles of Fenneropenaeus chinensis challenged with WSSV and Vibrio. Marine Biotechnology, 10(6):664–675, December 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-008-9105-x.

Matsuoka:2008:TMG

[745] Makoto P. Matsuoka, Carlos Infante, Michael Reith, Jose Pedro Cañavate, Susan E. Douglas, and Manuel Manchado. Translational machinery of Senegalese sole (Solea senegalensis Kaup) and Atlantic halibut (Hippoglossus hippoglossus L.): Comparative sequence analysis of the complete set of 60S ribosomal proteins and their expression. Marine Biotechnology, 10(6):676–691, December 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-008-9104-y.

Uribe:2008:PAE

[746] Paulina Uribe, Daniela Fuentes, Jorge Valdés, Amir Shmaryahu, Alicia Zúñiga, David Holmes, and Pablo D. T. Valenzuela. Preparation and analysis of an expressed sequence tag library from the toxic dinoflagellate Alexandrium catenella. Marine Biotechnology, 10(6):692– 700, December 2008. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9107-8.

Xu:2008:SDG

[747] Yuzhou Xu, Qin Liu, Lingyun Zhou, Zhao Yang, and Yuanxing Zhang. Surface display of GFP by Pseudomonas syringae truncated ice nucleation protein in attenuated Vibrio anguillarum strain. Marine Biotechnology, 10(6):701–708, December 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-008-9108-7.

Motoyama:2008:ITM

[748] Kanna Motoyama, Yota Suma, Shoichiro Ishizaki, Yuji Nagashima, Ying Lu, Hideki Ushio, and Kazuo Shiomi. Identification of tropomyosins as major allergens in Antarctic krill and mantis shrimp and their amino acid sequence characteristics. Marine Biotechnology, 10(6):709– 718, December 2008. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9109-6.

Shah:2008:CAH

[749] Devendra H. Shah, Kenneth D. Cain, Gregory D. Wiens, and Douglas R. Call. Challenges associated with heterologous expression of Flavobacterium psychrophilum proteins in Escherichia coli. Marine Biotechnology, 10(6):719–730, December 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-008-9111-z.

Fagutao:2008:GEP

[750] Fernand F. Fagutao, Motoshige Yasuike, Christopher Marlowe Caipang, Hidehiro Kondo, Ikuo Hirono, Yukinori Takahashi, and Takashi Aoki. Gene expression profile of hemocytes of kuruma shrimp, Marsupenaeus japonicus following peptidoglycan stimulation. Marine Biotechnology, 10(6):731–740, December 2008. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.com/

article/10.1007/s10126-008-9110-0.

Yazawa:2008:EMD

[751] Ryosuke Yazawa, Motoshige Yasuike, Jong Leong, Kristian R. von Schalburg, Glenn A. Cooper, Marianne Beetz-Sargent, Adrienne Robb, William S. Davidson, Simon R. M. Jones, and Ben F. Koop. EST and mitochondrial DNA sequences support a distinct Pacific form of salmon louse, Lepeophtheirus salmonis. Marine Biotechnology, 10(6): 741–749, December 2008. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9112-y.

Kusik:2008:DMA

[752] Brandon W. Kusik, Michael J. Carvan III, and Ava J. Udvadia. Detection of mercury in aquatic environments using EPRE reporter zebrafish. Marine Biotechnology, 10(6):750–757, December 2008. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-008-9113-x.

Shekhar:2009:ANA

[753] Mudagandur S. Shekhar and Yuanan Lu. Application of nucleic-acid-based therapeutics for viral infections in shrimp aquaculture. Marine Biotechnology, 11(1):1–9, February 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-008-9155-0.

Kim:2009:CCA

[754] Dong Eun Kim, Eun Yeol Lee, and Hee Sook Kim. Cloning and characterization of alginate lyase from a marine bacterium Streptomyces sp. ALG-5. Marine Biotechnology, 11(1):10–16, February 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-008-9114-9.

Niu:2009:INC

[755] Yan Niu, Jian Kong, Longyun Fu, Jing Yang, and Yi Xu. Identification of a novel C20-elongase gene from the marine microalgae Pavlova viridis and its expression in Escherichia coli. Marine Biotechnology, 11 (1):17–23, February 2009. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9116-7.

Roberts:2009:AGI

[756] Steven Roberts, Giles Goetz, Samuel White, and Frederick Goetz. Analysis of genes isolated from plated hemocytes of the Pacific oyster, Crassostreas gigas. Marine Biotechnology, 11(1):24–44, February 2009. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-008-9117-6.

Desbois:2009:FAD

[757] Andrew P. Desbois, Andrew Mearns-Spragg, and Valerie J. Smith. A fatty acid from the diatom Phaeodactylum tricornutum is antibacterial against diverse bacteria including multi-resistant Staphylococcus aureus (MRSA). Marine Biotechnology, 11(1):45–52, February 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-008-9118-5.

Vasileiadou:2009:MPM

[758] Kalliopi Vasileiadou, Spiros Papakostas, Alexander Triantafyllidis, Ilias Kappas, and Theodore J. Abatzopoulos. A multiplex PCR method for rapid identification of Brachionus rotifers. Marine Biotechnology, 11(1): 53–61, February 2009. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9119-4.

Ranzer:2009:NPF

[759] Llanie K. Ranzer, Thomas B. Brück, Wolfram M. Brück, Jose V. Lopez, and Russell G. Kerr. A new prokaryotic farnesyldiphosphate synthase from the octocoral Eunicea fusca: Differential display, inverse PCR, cloning, and characterization. Marine Biotechnology, 11(1):

62–73, February 2009. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9120-y.

Regalado:2009:RUD

[760] Erik L. Regalado, María Rodríguez, Roberto Menéndez, Ángel A. Concepción, Clara Nogueiras, Abilio Laguna, Armando A. Rodríguez, David E. Williams, Patricia Lorenzo-Luaces, Olga Valdés, and Yasnay Hernandez. Repair of UVB-damaged skin by the antioxidant sulphated flavone glycoside thalassiolin B isolated from the marine plant Thalassia testudinum banks ex König. Marine Biotechnology, 11(1): 74–80, February 2009. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9123-8.

Ni:2009:APG

[761] Xiumei Ni, Lixi Yue, Zhenming Chi, Jing Li, Xianghong Wang, and Catherine Madzak. Alkaline protease gene cloning from the marine yeast Aureobasidium pullulans HN2-3 and the protease surface display on Yarrowia lipolytica for bioactive peptide production. Marine Biotechnology, 11(1):81–89, February 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-008-9122-9.

Choi:2009:IEP

[762] Hye-Sook Choi, Joung Han Yim, Hong Kum Lee, and Suhkneung Pyo. Immunomodulatory effects of polar lichens on the function of macrophages In Vitro. Marine Biotechnology, 11(1):90–98, February 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-008-9121-x.

Wang:2009:BOP

[763] Xiaohong Wang, Ute Schloßmacher, Matthias Wiens, Heinz C. Schröder, and Werner E. G. Müller. Biogenic origin of polymetallic nodules from the Clarion–Clipperton Zone in the Eastern Pacific Ocean: Electron microscopic and EDX evidence. Marine Biotechnology, 11(1):99– 108, February 2009. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9124-7.

Dai:2009:PSH

[764] Jian guo Dai, Hai wei Xie, Gang Jin, Wei guang Wang, Yan Zhang, and Yong Guo. Preliminary study on high-level expression of tandemarranged tachyplesin-encoding gene in Bacillus subtilis wb800 and its antibacterial activity. Marine Biotechnology, 11(1):109–117, February 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-008-9125-6.

Morita:2009:RCC

[765] Masaya Morita, Akira Iguchi, and Akihiro Takemura. Roles of calmodulin and calcium/calmodulin-dependent protein kinase in flagellar motility regulation in the coral Acropora digitifera. Marine Biotechnology, 11 (1):118–123, February 2009. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9127-4.

Huang:2009:CDS

[766] Yali Huang, Xintian Lai, Xiaocui He, Lixiang Cao, Zhirui Zeng, Jiong Zhang, and Shining Zhou. Characterization of a deep-sea sediment metagenomic clone that produces water-soluble melanin in Escherichia coli. Marine Biotechnology, 11(1):124–131, February 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-008-9128-3.

Han:2009:CAC

[767] Yue Han, Bingjie Yang, Fengli Zhang, Xiaoling Miao, and Zhiyong Li. Characterization of antifungal chitinase from marine Streptomyces sp. DA11 associated with South China Sea sponge Craniella australiensis. Marine Biotechnology, 11(1):132–140, February 2009. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-008-9126-5.

Chen:2009:CAC

[768] I.-Ping Chen, Chung-Yu Tang, Chih-Yung Chiou, Jia-Ho Hsu, Nuwei Vivian Wei, Carden C. Wallace, Paul Muir, Henry Wu, and Chaolun Allen Chen. Comparative analyses of coding and noncoding DNA regions indicate that Acropora (Anthozoa: Scleractina) possesses a similar evolutionary tempo of nuclear vs. mitochondrial genomes as in plants. Marine Biotechnology, 11(1):141–152, February 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-008-9129-2.

Feng:2009:PPM

[769] Danqing Feng, Caihuan Ke, Shaojing Li, Changyi Lu, and Feng Guo. Pyrethroids as promising marine antifoulants: Laboratory and field studies. Marine Biotechnology, 11(2):153–160, April 2009. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-008-9130-9.

Hwang:2009:IED

[770] Hee Sun Hwang, Seon Hwa Kim, Yung Geun Yoo, Yong Shik Chu, Yun Hee Shon, Kyung Soo Nam, and Jong Won Yun. Inhibitory effect of deep-sea water on differentiation of 3T3-L1 adipocytes. Marine Biotechnology, 11(2):161–168, April 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-008-9131-8.

Reading:2009:CVM

[771] Benjamin J. Reading, Naoshi Hiramatsu, Sayumi Sawaguchi, Takahiro Matsubara, Akihiko Hara, Mark O. Lively, and Craig V. Sullivan. Conserved and variant molecular and functional features of multiple egg yolk precursor proteins (vitellogenins) in white perch (Morone americana) and other teleosts. Marine Biotechnology, 11(2):169–187, April 2009. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-008-9133-6.

Roper:2009:CAA

[772] K. E. Roper, H. Beamish, M. J. Garson, G. A. Skilleter, and B. M. Degnan. Convergent antifouling activities of structurally distinct bioactive compounds synthesized within two sympatric Haliclona demosponges. Marine Biotechnology, 11(2):188–198, April 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-008-9132-7.

Sung:2009:HHP

[773] Ming-Shiuan Sung, Yi-Ting Hsu, Yuan-Ting Hsu, Tzure-Meng Wu, and Tse-Min Lee. Hypersalinity and hydrogen peroxide upregulation of gene expression of antioxidant enzymes in Ulva fasciata against oxidative stress. Marine Biotechnology, 11(2):199–209, April 2009. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-008-9134-5.

Lynn:2009:SSS

[774] S. G. Lynn, K. A. Powell, D. F. Westneat, and B. S. Shepherd. Seasonal and sex-specific mRNA levels of key endocrine genes in adult yellow perch (Perca flavescens) from Lake Erie. Marine Biotechnology, 11 (2):210–222, April 2009. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9136-3.

Zhan:2009:FSP

[775] Aibin Zhan, Jingjie Hu, Xiaoli Hu, Zunchun Zhou, Min Hui, Shi Wang, Wei Peng, Mingling Wang, and Zhenmin Bao. Fine-scale population genetic structure of Zhikong scallop (Chlamys farreri): Do local marine currents drive geographical differentiation? Marine Biotechnology, 11 (2):223–235, April 2009. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9138-1.

Wang:2009:FLC

[776] Xubo Wang, Quanqi Zhang, Xiaohua Sun, Yanjie Chen, Teng Zhai, Wei Zhuang, Jie Qi, and Zhigang Wang. Fosmid library construction and initial analysis of end sequences in female half-smooth tongue sole (Cynoglossus semilaevis). Marine Biotechnology, 11(2):236–242, April 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-008-9137-2.

Chen:2009:AGS

[777] Song-Lin Chen, Yong-Sheng Tian, Jing-Feng Yang, Chang-Wei Shao, Xiang-Shan Ji, Jie-Ming Zhai, Xiao-Lin Liao, Zhi-Meng Zhuang, Peng-Zhi Su, Jian-Yong Xu, Zhen-Xia Sha, Peng-Fei Wu, and Na Wang. Artificial gynogenesis and sex determination in half-smooth tongue sole (Cynoglossus semilaevis). Marine Biotechnology, 11(2):243–251, April 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-008-9139-0.

Li:2009:IME

[778] Ye Li, Ye-Qing Qian, Wen-Ming Ma, and Wei-Jun Yang. Inhibition mechanism and the effects of structure on activity of male reproduction-related peptidase inhibitor kazal-type (MRPINK) of Macrobrachium rosenbergii. Marine Biotechnology, 11(2):252–259, April 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-008-9140-7.

Zoccola:2009:SEB

[779] Didier Zoccola, Aurélie Moya, Guillaume E. Béranger, Eric Tambutté, Denis Allemand, Georges F. Carle, and Sylvie Tambutté. Specific expression of BMP2/4 ortholog in biomineralizing tissues of corals and action on mouse BMP receptor. Marine Biotechnology, 11(2):260–269, April 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-008-9141-6.

Fang:2009:IPS

[780] Z. Fang, Q. Wang, W. Cao, Q. Feng, C. Li, L. Xie, and R. Zhang. Investigation of phosphorylation site responsible for CaLP (P. fucata) nucleo-cytoplasmic shuttling triggered by overexpression of p21^{Cip1}. Marine Biotechnology, 11(2):270–279, April 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-008-9142-5.

Zhang:2009:HTP

[781] Tong Zhang, Zhenming Chi, and Jun Sheng. A highly thermosensitive and permeable mutant of the marine yeast Cryptococcus aureus G7a potentially useful for single-cell protein production and its nutritive components. Marine Biotechnology, 11(2):280–286, April 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-008-9144-3.

Wiese:2009:DAA

[782] Jutta Wiese, Vera Thiel, Kerstin Nagel, Tim Staufenberger, and Johannes F. Imhoff. Diversity of antibiotic-active bacteria associated with the brown alga Laminaria saccharina from the Baltic Sea. Marine Biotechnology, 11(2):287–300, April 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-008-9143-4.

Chow:2009:PAL

[783] S. Chow, Y. Ueno, M. Toyokawa, I. Oohara, and H. Takeyama. Preliminary analysis of length and GC content variation in the ribosomal first internal transcribed spacer (ITS1) of marine animals. Marine Biotechnology, 11(3):301–306, June 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-008-9153-2.

Jeon:2009:CPA

[784] Jeong Ho Jeon, Jun-Tae Kim, Sung Gyun Kang, Jung-Hyun Lee, and Sang-Jin Kim. Characterization and its potential application of two esterases derived from the Arctic sediment metagenome. Marine Biotechnology, 11(3):307–316, June 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-008-9145-2.

Jiang:2009:FAA

[785] Hui Jiang, Yi-Mei Cai, Li-Qiao Chen, Xiao-Wei Zhang, Song-Nian Hu, and Qun Wang. Functional annotation and analysis of expressed sequence tags from the hepatopancreas of mitten crab (Eriocheir sinensis). Marine Biotechnology, 11(3):317–326, June 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-008-9146-1.

Satoh:2009:CSS

[786] Manami Satoh, Koji Iwamoto, Iwane Suzuki, and Yoshihiro Shiraiwa. Cold stress stimulates intracellular calcification by the coccolithophore, Emiliania huxleyi (Haptophyceae) under phosphate-deficient conditions. Marine Biotechnology, 11(3):327–333, June 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-008-9147-0.

Wang:2009:PCP

[787] San-Lang Wang, Chia-Hsing Chao, Tzu-Wen Liang, and Chung-Chih Chen. Purification and characterization of protease and chitinase from Bacillus cereus TKU006 and conversion of marine wastes by these enzymes. Marine Biotechnology, 11(3):334–344, June 2009. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-008-9149-y.

Zhang:2009:SABa

[788] Wei Zhang, Zhiyong Li, Xiaoling Miao, and Fengli Zhang. The screening of antimicrobial bacteria with diverse novel nonribosomal peptide synthetase (NRPS) genes from South China Sea sponges. Marine Biotechnology, 11(3):345, June 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.com/

article/10.1007/s10126-009-9183-4.

Zhang:2009:SABb

[789] Wei Zhang, Zhiyong Li, Xiaoling Miao, and Fengli Zhang. The screening of antimicrobial bacteria with diverse novel nonribosomal peptide synthetase (NRPS) genes from South China Sea sponges. Marine Biotechnology, 11(3):346–355, June 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-008-9148-z.

Chung:2009:ITN

[790] Chih-Ching Chung, Sheng-Ping L. Hwang, and Jeng Chang. The identification of three novel genes involved in the rapid-growth regulation in a marine diatom, Skeletonema costatum. Marine Biotechnology, 11 (3):356–367, June 2009. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9150-5.

Taoka:2009:ICT

[791] Yousuke Taoka, Naoki Nagano, Yuji Okita, Hitoshi Izumida, Shinichi Sugimoto, and Masahiro Hayashi. Influences of culture temperature on the growth, lipid content and fatty acid composition of Aurantiochytrium sp. strain mh0186. Marine Biotechnology, 11(3):368–374, June 2009. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-008-9151-4.

Ortega-Morales:2009:ABS

[792] B. O. Ortega-Morales, F. N. Ortega-Morales, J. Lara-Reyna, S. C. oDe la Rosa-García, A. Martínez-Hernández, and Jorge Montero-M. Antagonism of Bacillus spp. isolated from marine biofilms against terrestrial phytopathogenic fungi. Marine Biotechnology, 11(3):375–383, June 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-008-9152-3.

Kennedy:2009:IAB

[793] Jonathan Kennedy, Paul Baker, Clare Piper, Paul D. Cotter, Marcella Walsh, Marlies J. Mooij, Marie B. Bourke, Mary C. Rea, Paula M. O’Connor, R. Paul Ross, Colin Hill, Fergal O’Gara, Julian R. Marchesi, and Alan D. W. Dobson. Isolation and analysis of bacteria with antimicrobial activities from the marine sponge Haliclona simulans collected from Irish waters. Marine Biotechnology, 11(3):384–396, June 2009. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-008-9154-1.

Masukawa:2009:SDD

[794] Hajime Masukawa, Xiaohui Zhang, Emi Yamazaki, Syunsuke Iwata, Kensuke Nakamura, Mari Mochimaru, Kazuhito Inoue, and Hidehiro Sakurai. Survey of the distribution of different types of nitrogenases and hydrogenases in heterocyst-forming cyanobactera. Marine Biotechnology, 11(3):397–409, June 2009. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9156-z.

Robert:2009:ICF

[795] Stanley S. Robert, James R. Petrie, Xue-Rong Zhou, Maged P. Mansour, Susan I. Blackburn, Allan G. Green, Surinder P. Singh, and Peter D. Nichols. Isolation and characterisation of a Δ5-fatty acid elongase from the marine microalga Pavlova salina. Marine Biotechnology, 11 (3):410–418, June 2009. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9157-y.

Wang:2009:DWO

[796] Yun-Hsin Wang, Chi-Chung Wen, Zhi-Shiang Yang, Chien-Chung Cheng, Jen-Ning Tsai, Chia-Chen Ku, Hsin-Ju Wu, and Yau-Hung Chen. Development of a whole-organism model to screen new compounds for Sun protection. Marine Biotechnology, 11(3):419–429, June 2009. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-008-9159-9.

Carrano:2009:BML

[797] Carl J. Carrano, Stephen Schellenberg, Shady A. Amin, David H. Green, and Frithjof C. Küpper. Boron and marine life: a new look at an enigmatic bioelement. Marine Biotechnology, 11(4):431–440, July 2009. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-009-9191-4.

Schlesinger:2009:ANI

[798] Ami Schlesinger, Esti Kramarsky-Winter, and Yossi Loya. Active nematocyst isolation via nudibranchs. Marine Biotechnology, 11(4): 441–444, July 2009. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9175-9.

Chabeaud:2009:ARS

[799] A. Chabeaud, P. Dutournié, F. Guérard, L. Vandanjon, and P. Bourseau. Application of response surface methodology to optimise the antioxidant activity of a saithe (Pollachius virens) hydrolysate. Marine Biotechnology, 11(4):445–455, July 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-008-9158-x.

Zhang:2009:CSP

[800] Ruifu Zhang and Ji-Dong Gu. Complete sequence of plasmid pMP1 from the marine environmental Vibrio vulnificus and location of its replication origin. Marine Biotechnology, 11(4):456–462, July 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-008-9160-3.

Kim:2009:SCP

[801] Jeong-Dong Kim, Ji-Young Kim, Jae-Kweon Park, and Choul-Gyun Lee. Selective control of the Prorocentrum minimum harmful algal blooms by a novel algal-lytic bacterium Pseudoalteromonas haloplanktis AFMB-008041 . Marine Biotechnology, 11(4):463–472, July 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-008-9167-9.

Salas-Leiton:2009:PJGa

[802] E. Salas-Leiton, B. Cánovas-Conesa, R. Zerolo, J. López-Barea, J. P. Cañavate, and J. Alhama. Proteomics of juvenile Senegal sole (Solea senegalensis) affected by gas bubble disease in hyperoxygenated ponds. Marine Biotechnology, 11(4):473–487, July 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-008-9168-8.

Khalesi:2009:LDG

[803] Mohammad K. Khalesi, H. H. Beeftink, and R. H. Wijffels. Lightdependency of growth and secondary metabolite production in the captive zooxanthellate soft coral Sinularia flexibilis. Marine Biotechnology, 11(4):488–494, July 2009. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9164-z.

Xu:2009:IEB

[804] Ying Xu, Honglei Li, Xiancui Li, Xiang Xiao, and Pei-Yuan Qian. Inhibitory effects of a branched-chain fatty acid on larval settlement of the polychaete Hydroides elegans. Marine Biotechnology, 11(4): 495–504, July 2009. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9161-2.

Cai:2009:FCE

[805] Zhonghua Cai, Chunping Gao, Yong Zhang, and Kezhi Xing. Functional characterization of the ELR motif in piscine ELR^{+} CXC-like chemokine. Marine Biotechnology, 11(4):505–512, July 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-008-9165-y.

Ho:2009:PTG

[806] Chai-Ling Ho, Seddon Teoh, Swee-Sen Teo, Raha Abdul Rahim, and Siew-Moi Phang. Profiling the transcriptome of Gracilaria changii (Rhodophyta) in response to light deprivation. Marine Biotechnology, 11(4):513–519, July 2009. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9166-x.

Bai:2009:IGI

[807] Zhiyi Bai, Yuxin Yin, Songnian Hu, Guiling Wang, Xiaowei Zhang, and Jiale Li. Identification of genes involved in immune response, microsatellite, and SNP markers from expressed sequence tags generated from hemocytes of freshwater pearl mussel (Hyriopsis cumingii). Marine Biotechnology, 11(4):520–530, July 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-008-9163-0.

Hwang:2009:AOA

[808] Hee Sun Hwang, Hyun Ah Kim, Sung Hak Lee, and Jong Won Yun. Anti-obesity and antidiabetic effects of deep sea water on ob/ob mice. Marine Biotechnology, 11(4):531–539, July 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-008-9171-0.

Baker:2009:PDA

[809] Paul W. Baker, Jonathan Kennedy, Alan D. W. Dobson, and Julian R. Marchesi. Phylogenetic diversity and antimicrobial activities of fungi associated with Haliclona simulans isolated from Irish coastal waters. Marine Biotechnology, 11(4):540–547, July 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-008-9169-7.

Salas-Leiton:2009:PJGb

[810] E. Salas-Leiton, B. Cánovas-Conesa, R. Zerolo, J. López-Barea, J. P. Cañavate, and J. Alhama. Proteomics of juvenile Senegal sole (Solea senegalensis) affected by gas bubble disease in hyperoxygenated ponds. Marine Biotechnology, 11(4):548–549, July 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-009-9190-5.

Thavasi:2009:BPA

[811] R. Thavasi, V. R. M. Subramanyam Nambaru, S. Jayalakshmi, T. Balasubramanian, and Ibrahim M. Banat. Biosurfactant production by Azotobacter chroococcum isolated from the marine environment. Marine Biotechnology, 11(5):551–556, October 2009. CODEN MABIFW.

ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-008-9162-1.

Kim:2009:IEF

[812] Mi-Ja Kim, Un-Jae Chang, and Jin-Sil Lee. Inhibitory effects of fucoidan in 3T3-L1 adipocyte differentiation. Marine Biotechnology, 11 (5):557–562, October 2009. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9170-1.

Mikami:2009:VPD

[813] Koji Mikami, Toshiki Uji, Lin Li, Megumu Takahashi, Hajime Yasui, and Naotsune Saga. Visualization of phosphoinositides via the development of the transient expression system of a cyan fluorescent protein in the red alga Porphyra yezoensis. Marine Biotechnology, 11(5): 563–569, October 2009. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9172-z.

Lallias:2009:CTS

[814] D. Lallias, L. Gomez-Raya, C. S. Haley, I. Arzul, S. Heurtebise, A. R. Beaumont, P. Boudry, and S. Lapègue. Combining two-stage testing and interval mapping strategies to detect QTL for resistance to bonamiosis in the European flat oyster Ostrea edulis. Marine Biotechnology, 11(5): 570–584, October 2009. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-008-9173-y.

Ferreira:2009:ERP

[815] Martiña Ferreira, Paula Coutinho, Pedro Seixas, Jaime Fábregas, and Ana Otero. Enriching rotifers with “Premium” microalgae. Nannochloropsis gaditana. Marine Biotechnology, 11(5):585–595, October 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-008-9174-x.

Huang:2009:ICA

[816] Jing Huang, Hongzhong Wang, Yu Cui, Guiyou Zhang, Guilan Zheng, Shiting Liu, Liping Xie, and Rongqing Zhang. Identification and comparison of amorphous calcium carbonate-binding protein and acetylcholinebinding protein in the abalone, Haliotis discus hannai. Marine Biotechnology, 11(5):596–607, October 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-008-9176-8.

Ulrich:2009:TSM

[817] P. N. Ulrich and A. G. Marsh. Thermal sensitivity of mitochondrial respiration efficiency and protein phosphorylation in the clam Mercenaria mercenaria. Marine Biotechnology, 11(5):608–618, October 2009. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-009-9177-2.

Liu:2009:SDA

[818] Guanglei Liu, Lixi Yue, Zhe Chi, Wengong Yu, Zhenming Chi, and Catherine Madzak. The surface display of the alginate lyase on the cells of Yarrowia lipolytica for hydrolysis of alginate. Marine Biotechnology, 11 (5):619–626, October 2009. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-009-9178-1.

Morais:2009:HUF

[819] Sofia Morais, Oscar Monroig, Xiaozhong Zheng, Michael J. Leaver, and Douglas R. Tocher. Highly unsaturated fatty acid synthesis in Atlantic salmon: Characterization of ELOVL5- and ELOVL2-like elongases. Marine Biotechnology, 11(5):627–639, October 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-009-9179-0.

Xue:2009:GSE

[820] Lingyun Xue and Wei Zhang. Growth and survival of early juveniles of the marine sponge Hymeniacidon perlevis (Demospongiae) under controlled conditions. Marine Biotechnology, 11(5):640–649, October 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-009-9180-7.

Lang:2009:TPS

[821] R. Paul Lang, Christopher J. Bayne, Mark D. Camara, Charles Cunningham, Matthew J. Jenny, and Christopher J. Langdon. Transcriptome profiling of selectively bred Pacific oyster Crassostrea gigas families that differ in tolerance of heat shock. Marine Biotechnology, 11(5): 650–668, October 2009. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-009-9181-6.

Duckworth:2009:FSS

[822] Alan Duckworth. Farming sponges to supply bioactive metabolites and bath sponges: a review. Marine Biotechnology, 11(6):669–679, December 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-009-9213-2.

Noleto:2009:GSE

[823] Rafael Bueno Noleto, Fernando de Souza Fonseca Guimarães, Katia Sabrina Paludo, Marcelo Ricardo Vicari, Roberto Ferreira Artoni, and Marta Margarete Cestari. Genome size evaluation in Tetraodontiform fishes from the neotropical region. Marine Biotechnology, 11(6):680– 685, December 2009. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-009-9215-0.

Santos:2009:FIB

[824] R. Santos, G. da Costa, C. Franco, P. Gomes-Alves, P. Flammang, and A. V. Coelho. First insights into the biochemistry of tube foot adhesive from the sea urchin Paracentrotus lividus (Echinoidea, Echinodermata). Marine Biotechnology, 11(6):686–698, December 2009. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-009-9182-5.

Liao:2009:CGL

[825] Xiaolin Liao, Hong-Yu Ma, Gen-Bo Xu, Chang-Wei Shao, Yong-Sheng Tian, Xiang-Shan Ji, Jing-Feng Yang, and Song-Lin Chen. Construction of a genetic linkage map and mapping of a female-specific DNA marker in half-smooth tongue sole (Cynoglossus semilaevis). Marine Biotechnology, 11(6):699–709, December 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-009-9184-3.

He:2009:NUO

[826] Hai-Lun He, Xiu-Lan Chen, Xi-Ying Zhang, Cai-Yun Sun, Bai-Chen Zou, and Yu-Zhong Zhang. Novel use for the osmolyte trimethylamine N-oxide: Retaining the psychrophilic characters of cold-adapted protease deseasin MCP-01 and simultaneously improving its thermostability. Marine Biotechnology, 11(6):710–716, December 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-009-9185-2.

Lu:2009:CIN

[827] Yingjian Lu, Xin Dong, Shu Liu, and Xiaomei Bie. Characterization and identification of a novel marine Streptomyces sp. produced antibacterial substance. Marine Biotechnology, 11(6):717–724, December 2009. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-009-9186-1.

DSouza-Ticlo:2009:TMT

[828] Donna D’Souza-Ticlo, Deepak Sharma, and Chandralata Raghukumar. A thermostable metal-tolerant laccase with bioremediation potential from a marine-derived fungus. Marine Biotechnology, 11(6):725– 737, December 2009. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-009-9187-0.

Fu:2009:MCE

[829] Wandong Fu, Jianting Yao, Xiuliang Wang, Fuli Liu, Gang Fu, and Delin Duan. Molecular cloning and expression analysis of a cytosolic Hsp70 gene from Laminaria japonica (Laminariaceae, Phaeophyta). Marine Biotechnology, 11(6):738–747, December 2009. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-009-9188-z.

Subramanian:2009:MNA

[830] Sangeetha Subramanian, Neil W. Ross, and Shawna L. MacKinnon. Myxinidin, a novel antimicrobial peptide from the epidermal mucus of hagfish, Myxine glutinosa L. Marine Biotechnology, 11(6):748– 757, December 2009. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-009-9189-y.

Isaacs:2009:CBC

[831] LeLeng To Isaacs, Jinjun Kan, Linh Nguyen, Patrick Videau, Matthew A. Anderson, Toby L. Wright, and Russell T. Hill. Comparison of the bacterial communities of wild and captive sponge Clathria prolifera from the Chesapeake Bay. Marine Biotechnology, 11(6):758– 770, December 2009. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-009-9192-3.

Hong:2009:CCA

[832] Ming-Cheng Hong, Yung-Sen Huang, Pa-Ching Song, Wen-Wen Lin, Lee-Shing Fang, and Ming-Chyuan Chen. Cloning and characterization of ApRab4, a recycling Rab protein of Aiptasia pulchella, and its implication in the symbiosome biogenesis. Marine Biotechnology, 11(6): 771–785, December 2009. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-009-9193-2.

Zhan:2010:DRE

[833] Huiqing Zhan and Zhiyuan Gong. Delayed and restricted expression of UAS-Regulated GFP gene in early transgenic zebrafish embryos by using the GAL4/UAS system. Marine Biotechnology, 12(1): 1–7, February 2010. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-009-9217-y.

Koshimizu:2010:CGM

[834] Eriko Koshimizu, Carlos Augusto Strüssmann, Nobuaki Okamoto, Hideo Fukuda, and Takashi Sakamoto. Construction of a genetic map and development of DNA markers linked to the sex-determining locus in the Patagonian pejerrey (Odontesthes hatcheri). Marine Biotechnology, 12 (1):8–13, February 2010. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-009-9194-1.

Kleinegris:2010:SMD

[835] Dorinde M. M. Kleinegris, Marcel Janssen, Willem A. Brandenburg, and René H. Wijffels. The selectivity of milking of Dunaliella salina. Marine Biotechnology, 12(1):14–23, February 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-009-9195-0.

Umemoto:2010:CWR

[836] Yoshiaki Umemoto and Toshiyoshi Araki. Cell wall regeneration in Bangia atropurpurea (Rhodophyta) protoplasts observed using a mannanspecific carbohydrate-binding module. Marine Biotechnology, 12(1): 24–31, February 2010. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-009-9196-z.

Lopez-Legentil:2010:CLH

[837] Susanna López-Legentil, Bongkeun Song, Michael DeTure, and Daniel G. Baden. Characterization and localization of a hybrid non-ribosomal peptide synthetase and polyketide synthase gene from the toxic dinoflagellate Karenia brevis. Marine Biotechnology, 12(1):32–41, February 2010. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-009-9197-y.

Katagiri:2010:ECM

[838] Fumio Katagiri, Yukiko Takatsuka, Shoko Fujiwara, and Mikio Tsuzuki. Effects of Ca and Mg on growth and calcification of the coccolithophorid Pleurochrysis haptonemofera: Ca requirement for cell division in coccolith-bearing cells and for normal coccolith formation with acidic polysaccharides. Marine Biotechnology, 12(1):42–51, February 2010. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-009-9198-x.

Plouguerne:2010:AMA

[839] Erwan Plouguerné, Efstathia Ioannou, Panagiota Georgantea, Constantinos Vagias, Vassilios Roussis, Claire Hellio, Edouard Kraffe, and Valérie Stiger-Pouvreau. Anti-microfouling activity of lipidic metabolites from the invasive brown alga Sargassum muticum (Yendo) Fensholt. Marine Biotechnology, 12(1):52–61, February 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-009-9199-9.

Long:2010:NAH

[840] Mengxian Long, Ziniu Yu, and Xun Xu. A novel β-agarase with high pH stability from marine Agarivorans sp. LQ48. Marine Biotechnology, 12 (1):62–69, February 2010. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-009-9200-7.

Wanna:2010:IFS

[841] Warapond Wanna, Caird E. Rexroad III, and Jianbo Yao. Identification of a functional splice variant of 14-3-3E1 in rainbow trout. Marine Biotechnology, 12(1):70–80, February 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-009-9201-6.

Genji:2010:DPF

[842] Takahisa Genji, Seketsu Fukuzawa, and Kazuo Tachibana. Distribution and possible function of the marine alkaloid, norzoanthamine, in the zoanthid Zoanthus sp. using MALDI imaging mass spectrometry. Marine Biotechnology, 12(1):81–87, February 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-009-9202-5.

Donate:2010:EIT

[843] Carmen Doñate, Joan Carles Balasch, Agnes Callol, Julien Bobe, Lluis Tort, and Simon MacKenzie. The effects of immunostimulation through dietary manipulation in the rainbow trout; evaluation of mucosal immunity. Marine Biotechnology, 12(1):88–99, February 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-009-9203-4.

Li:2010:CPI

[844] Changzhong Li, Yilin Hu, Jian Liang, Yawei Kong, Jing Huang, Qiaoli Feng, Shuo Li, Guiyou Zhang, Liping Xie, and Rongqing Zhang. Calcineurin plays an important role in the shell formation of pearl oyster (Pinctada fucata). Marine Biotechnology, 12(1):100–110, February 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-009-9204-3.

Balabanova:2010:MCT

[845] Larissa A. Balabanova, Irina Yu. Bakunina, Olga I. Nedashkovskaya, Ilona D. Makarenkova, Tatiana S. Zaporozhets, Natalia N. Besednova, Tatiana N. Zvyagintseva, and Valery A. Rasskazov. Molecular characterization and therapeutic potential of a marine bacterium Pseudoalteromonas sp. KMM 701 α-galactosidase. Marine Biotechnology, 12(1): 111–120, February 2010. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-009-9205-2.

Shirak:2010:IRE

[846] Andrey Shirak, Manfred Grabherr, Federica Di Palma, Kerstin Lindblad-Toh, Gideon Hulata, Micha Ron, Tom D. Kocher, and Eyal Seroussi. Identification of repetitive elements in the genome of Oreochromis niloticus: Tilapia repeat masker. Marine Biotechnology, 12(2):121–125, April 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-009-9236-8.

Seear:2010:DGE

[847] Paul J. Seear, Stephen N. Carmichael, Richard Talbot, John B. Taggart, James E. Bron, and Glen E. Sweeney. Differential gene expression during smoltification of Atlantic salmon (Salmo salar L.): a first large-scale microarray study. Marine Biotechnology, 12(2):126–140, April 2010. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-009-9218-x.

Zhang:2010:CCB

[848] Xiaojun Zhang, Yang Zhang, Chantel Scheuring, Hong-Bin Zhang, Pin Huan, Bing Wang, Chengzhang Liu, Fuhua Li, Bin Liu, and Jianhai Xiang. Construction and characterization of a bacterial artificial chromosome (BAC) library of Pacific white shrimp, Litopenaeus vannamei. Marine Biotechnology, 12(2):141–149, April 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-009-9209-y.

Uji:2010:VNL

[849] Toshiki Uji, Megumu Takahashi, Naotsune Saga, and Koji Mikami. Visualization of nuclear localization of transcription factors with cyan and green fluorescent proteins in the red alga Porphyra yezoensis. Marine Biotechnology, 12(2):150–159, April 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-009-9210-5.

Liu:2010:EFM

[850] Shaojun Liu, Qinbo Qin, Yuequn Wang, Hong Zhang, Rurong Zhao, Chun Zhang, Jing Wang, Wei Li, Lin Chen, Jun Xiao, Kaikun Luo, Min Tao, Wei Duan, and Yun Liu. Evidence for the formation of the male gynogenetic fish. Marine Biotechnology, 12(2):160–172, April 2010. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-009-9219-9.

Yang:2010:ICG

[851] Huey-Lang Yang, Chung-Kuang Lu, Shu-Fen Chen, Young-Mao Chen, and Yi-Min Chen. Isolation and characterization of Taiwanese heterotrophic microalgae: Screening of strains for docosahexaenoic acid (DHA) production. Marine Biotechnology, 12(2):173–185, April 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic).

URL https://link.springer.com/article/10.1007/s10126-009-
9207-0.

Liu:2010:IEA

[852] Dong Liu, Shaojun Liu, Cuiping You, Lin Chen, Zhen Liu, Liangguo Liu, Jing Wang, and Yun Liu. Identification and expression analysis of genes involved in early ovary development in diploid gynogenetic hybrids of red crucian carp × common carp. Marine Biotechnology, 12 (2):186–194, April 2010. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-009-9212-3.

Pearson:2010:EST

[853] Gareth A. Pearson, Galice Hoarau, Asuncion Lago-Leston, James A. Coyer, Michael Kube, Richard Reinhardt, Kolja Henckel, Ester T. A.

Serrão, Erwan Corre, and Jeanine L. Olsen. An expressed sequence tag analysis of the intertidal brown seaweeds Fucus serratus (L.) and F. vesiculosus (L.) (Heterokontophyta, Phaeophyceae) in response to abiotic stressors. Marine Biotechnology, 12(2):195–213, April 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-009-9208-z.

Murray:2010:EEI

[854] H. M. Murray, S. P. Lall, R. Rajaselvam, L. A. Boutilier, R. M. Flight, B. Blanchard, S. Colombo, V. Mohindra, M. Yúfera, and S. E. Douglas. Effect of early introduction of microencapsulated diet to larval Atlantic halibut, Hippoglossus hippoglossus L. assessed by microarray analysis. Marine Biotechnology, 12(2):214–229, April 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-009-9211-4.

Wonglapsuwan:2010:CBA

[855] Monwadee Wonglapsuwan, Teruo Miyazaki, Wiriya Loongyai, and Wilaiwan Chotigeat. Characterization and biological activity of the ribosomal protein L10a of the white shrimp: Fenneropenaeus merguiensis De Man during vitellogenesis. Marine Biotechnology, 12(2):230–240, April 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-009-9220-3.

Caldwell:2010:MGC

[856] Gary S. Caldwell and Helen E. Pagett. Marine glycobiology: Current status and future perspectives. Marine Biotechnology, 12(3): 241–252, June 2010. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9263-5.

Nishijima:2010:ATB

[857] Miyuki Nishijima, Dhugal J. Lindsay, Junko Hata, Aoi Nakamura, Hiroaki Kasai, Yuji Ise, Charles R. Fisher, Yoshihiro Fujiwara, Masaru Kawato, and Tadashi Maruyama. Association of thioautotrophic bacteria with deep-sea sponges. Marine Biotechnology, 12(3):253–260, June 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-009-9253-7.

Jang:2010:CCT

[858] Songhun Jang, Hang Liu, Jianguo Su, Feng Dong, Feng Xiong, Lanjie Liao, Yaping Wang, and Zuoyan Zhu. Construction and characterization of two bacterial artificial chromosome libraries of grass carp. Marine Biotechnology, 12(3):261–266, June 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-010-9268-0.

Bickmeyer:2010:SBM

[859] Ulf Bickmeyer, Achim Grube, Karl-Walter Klings, Joseph R. Pawlik, and Matthias Köck. Siphonodictyal B1 from a marine sponge increases intracellular calcium levels comparable to the Ca^{2+}-ATPase (SERCA) inhibitor thapsigargin. Marine Biotechnology, 12(3):267–272, June 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-009-9206-1.

Sissener:2010:PPL

[860] Nini H. Sissener, Samuel A. M. Martin, Phillip Cash, Ernst M. Hevrøy, Monica Sanden, and Gro-Ingunn Hemre. Proteomic profiling of liver from Atlantic salmon (Salmo salar) fed genetically modified soy compared to the near-isogenic non-GM line. Marine Biotechnology, 12(3): 273–281, June 2010. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-009-9214-1.

Feng:2010:IIG

[861] Bingbing Feng, Lingli Dong, Donghong Niu, Shanshan Meng, Bing Zhang, Dabo Liu, Songnian Hu, and Jiale Li. Identification of immune genes of the agamaki clam (Sinonovacula constricta) by sequencing and bioinformatic analysis of ESTs. Marine Biotechnology, 12(3): 282–291, June 2010. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-009-9216-z.

Xu:2010:CPH

[862] Qingsong Xu, Pan Ma, Weiting Yu, Chengyu Tan, Hongtao Liu, Chuannan Xiong, Ying Qiao, and Yuguang Du. Chitooligosaccharides protect human embryonic hepatocytes against oxidative stress induced by hydrogen peroxide. Marine Biotechnology, 12(3):292–298, June 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-009-9222-1.

Kong:2010:DTT

[863] Chang-Suk Kong, Jung-Ae Kim, Byul-Nim Ahn, Thanh Sang Vo, Na-Young Yoon, and Se-Kwon Kim. 1-(3′, 5′-dihydroxyphenoxy)-7- (2”,4”,6-trihydroxyphenoxy)-2,4,9-trihydroxydibenzo-1,4-dioxin inhibits adipocyte differentiation of 3T3-L1 fibroblasts. Marine Biotechnology, 12(3):299–307, June 2010. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-009-9224-z.

Wiebring:2010:NMS

[864] Annika Wiebring, Heike Helmholz, Ilka Sötje, Stephan Lassen, Andreas Prange, and Henry Tiemann. A new method for the separation of different types of nematocysts from Scyphozoa and investigation of proteinaceous toxins utilizing laser catapulting and subsequent mass spectrometry. Marine Biotechnology, 12(3):308–317, June 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-010-9261-7.

Ouyang:2010:IHM

[865] Yongchang Ouyang, Shikun Dai, Lianwu Xie, M. S. Ravi Kumar, Wei Sun, Huimin Sun, Danling Tang, and Xiang Li. Isolation of high molecular weight DNA from marine sponge bacteria for BAC library construction. Marine Biotechnology, 12(3):318–325, June 2010. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-009-9223-0.

Fleury:2010:MBI

[866] Elodie Fleury, Jeanne Moal, Viviane Boulo, Jean-Yves Daniel, David Mazurais, Alain Hénaut, Charlotte Corporeau, Pierre Boudry, Pascal Favrel, and Arnaud Huvet. Microarray-based identification of gonad transcripts differentially expressed between lines of Pacific oyster selected to be resistant or susceptible to summer mortality. Marine Biotechnology, 12(3):326–339, June 2010. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-009-9227-9.

Koopmans:2010:GEC

[867] Marieke Koopmans, Dirk Martens, and Rene H. Wijffels. Growth efficiency and carbon balance for the sponge Haliclona oculata. Marine Biotechnology, 12(3):340–349, June 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-009-9228-8.

Radwan:2010:BCA

[868] Mona Radwan, Amro Hanora, Jindong Zan, Naglaa M. Mohamed, Dina M. Abo-Elmatty, Soad H. Abou-El-Ela, and Russell T. Hill. Bacterial community analyses of two red sea sponges. Marine Biotechnology, 12(3):350–360, June 2010. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-009-9239-5.

Magnadottir:2010:ICF

[869] Bergljot Magnadottir. Immunological control of fish diseases. Marine Biotechnology, 12(4):361–379, August 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-010-9279-x.

Waldbieser:2010:PVH

[870] Geoffrey C. Waldbieser, Brian G. Bosworth, and Sylvie M. A. Quiniou. Production of viable homozygous, doubled haploid channel catfish (Ictalurus punctatus). Marine Biotechnology, 12(4):380–385, August 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-009-9221-2.

Liu:2010:QMF

[871] Fuli Liu, Zhanru Shao, Haining Zhang, Jidong Liu, Xiuliang Wang, and Delin Duan. QTL mapping for frond length and width in Laminaria japonica Aresch (Laminarales, Phaeophyta) using AFLP and SSR markers. Marine Biotechnology, 12(4):386–394, August 2010. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-009-9229-7.

Okamura:2010:ICG

[872] Yoshiko Okamura, Tomonori Kimura, Hiroko Yokouchi, Macarena Meneses-Osorio, Masaya Katoh, Tadashi Matsunaga, and Haruko Takeyama. Isolation and characterization of a GDSL esterase from the metagenome of a marine sponge-associated bacteria. Marine Biotechnology, 12(4):395–402, August 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-009-9226-x.

Kozhemyako:2010:SGS

[873] Valeri B. Kozhemyako, Galina N. Veremeichik, Yuri N. Shkryl, Svetlana N. Kovalchuk, Vladimir B. Krasokhin, Valeri A. Rasskazov, Yuri N. Zhuravlev, Victor P. Bulgakov, and Yuri N. Kulchin. Silicatein genes in spicule-forming and nonspicule-forming Pacific demosponges. Marine Biotechnology, 12(4):403–409, August 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-009-9225-y.

Salem:2010:MRF

[874] Mohamed Salem, Caide Xiao, Jonah Womack, Caird E. Rexroad, and Jianbo Yao. A MicroRNA repertoire for functional genome research in rainbow trout (Oncorhynchus mykiss). Marine Biotechnology, 12(4): 410–429, August 2010. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-009-9232-z.

Petrie:2010:ICH

[875] James R. Petrie, Qing Liu, Anne M. Mackenzie, Pushkar Shrestha, Maged P. Mansour, Stan S. Robert, Dion F. Frampton, Susan I. Blackburn, Peter D. Nichols, and Surinder P. Singh. Isolation and characterisation of a high-efficiency desaturase and elongases from microalgae for transgenic LC–PUFA production. Marine Biotechnology, 12(4): 430–438, August 2010. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-009-9230-1.

Gram:2010:AAM

[876] Lone Gram, Jette Melchiorsen, and Jesper Bartholin Bruhn. Antibacterial activity of marine culturable bacteria collected from a global sampling of ocean surface waters and surface swabs of marine organisms. Marine Biotechnology, 12(4):439–451, August 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-009-9233-y.

Millan:2010:DPT

[877] Adrián Millán, Antonio Gómez-Tato, Carlos Fernández, Belén G. Pardo, José A. Álvarez-Dios, Manuel Calaza, Carmen Bouza, María Vázquez, Santiago Cabaleiro, and Paulino Martínez. Design and performance of a turbot (Scophthalmus maximus) oligo-microarray based on ESTs from immune tissues. Marine Biotechnology, 12(4):452–465, August 2010. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-009-9231-0.

Hsu:2010:IMI

[878] Chia-Chun Hsu, Min-Fon Hou, Jiann-Ruey Hong, Jen-Leih Wu, and Guor Mour Her. Inducible male infertility by targeted cell ablation in zebrafish testis. Marine Biotechnology, 12(4):466–478, August 2010. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-009-9248-4.

Peng:2010:GIA

[879] Ying Peng, Zhenming Chi, Xianghong Wang, and Jing Li. β-1,3-glucanase inhibits activity of the killer toxin produced by the marinederived yeast Williopsis saturnus WC91-2 . Marine Biotechnology, 12 (4):479–485, August 2010. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-009-9243-9.

Tassanakajon:2010:CAP

[880] Anchalee Tassanakajon, Piti Amparyup, Kunlaya Somboonwiwat, and Premruethai Supungul. Cationic antimicrobial peptides in penaeid shrimp. Marine Biotechnology, 12(5):487–505, October 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-010-9288-9.

Prakoon:2010:STG

[881] Warangkana Prakoon, Suriyan Tunkijjanukij, Thuy T. T. Nguyen, and Uthairat Na-Nakorn. Spatial and temporal genetic variation of green mussel, Perna viridis in the Gulf of Thailand and implication for aquaculture. Marine Biotechnology, 12(5):506–515, October 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-009-9234-x.

Carballo:2010:CES

[882] Jose L. Carballo, Benjamin Yañez, Eva Zubía, Maria J. Ortega, and Cristina Vega. Culture of explants from the sponge Mycale cecilia to obtain bioactive mycalazal-type metabolites. Marine Biotechnology, 12 (5):516–525, October 2010. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-009-9235-9.

Uchimura:2010:CSA

[883] Kohsuke Uchimura, Masayuki Miyazaki, Yuichi Nogi, Tohru Kobayashi, and Koki Horikoshi. Cloning and sequencing of alginate lyase genes from deep-sea strains of Vibrio and Agarivorans and characterization of a new Vibrio enzyme. Marine Biotechnology, 12(5):526–533, October 2010. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-009-9237-7.

Kavanagh:2010:RTP

[884] Siobhán Kavanagh, Claire Brennan, Louise O’Connor, Siobhán Moran, Rafael Salas, Josephine Lyons, Joe Silke, and Majella Maher. Realtime PCR detection of Dinophysis species in Irish coastal waters. Marine Biotechnology, 12(5):534–542, October 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-009-9238-6.

Seo:2010:GAO

[885] Jung-Kil Seo, Jeana Stephenson, J. Myron Crawford, Kathryn L. Stone, and Edward J. Noga. American oyster, Crassostrea virginica, expresses a potent antibacterial histone H2B protein. Marine Biotechnology, 12 (5):543–551, October 2010. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-009-9240-z.

Rocha:2010:BBM

[886] Lenilson Coutinho Rocha, Hercules Vicente Ferreira, Eli Fernando Pimenta, Roberto Gomes Souza Berlinck, Maria Olímpia Oliveira Rezende, Maria Diva Landgraf, Mirna Helena Regali Seleghim, Lara Durães Sette, and André Luiz Meleiro Porto. Biotransformation of αbromoacetophenones by the marine fungus Aspergillus sydowii. Marine Biotechnology, 12(5):552–557, October 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-009-9241-y.

Shao:2010:CTB

[887] Chang-Wei Shao, Song-Lin Chen, Chantel F. Scheuring, Jian-Yong Xu, Zhen-Xia Sha, Xiao-Li Dong, and Hong-Bin Zhang. Construction of two BAC libraries from half-smooth tongue sole Cynoglossus semilaevis and identification of clones containing candidate sex-determination genes. Marine Biotechnology, 12(5):558–568, October 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-009-9242-x.

Hu:2010:NMG

[888] Shao-Yang Hu, Pei-Yu Lin, Chia-Hsuan Liao, Hong-Yi Gong, Gen-Hwa Lin, Koichi Kawakami, and Jen-Leih Wu. Nitroreductase-mediated gonadal dysgenesis for infertility control of genetically modified zebrafish. Marine Biotechnology, 12(5):569–578, October 2010. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-009-9244-8.

Marrouchi:2010:QDG

[889] Riadh Marrouchi, Faten Dziri, Nawel Belayouni, Asma Hamza, Evelyne Benoit, Jordi Molgó, and Riadh Kharrat. Quantitative determination of gymnodimine-A by high performance liquid chromatography in contaminated clams from Tunisia coastline. Marine Biotechnology, 12(5): 579–585, October 2010. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-009-9245-7.

Van-Thuoc:2010:EPH

[890] Doan Van-Thuoc, Héctor Guzmán, Mai Thi-Hang, and Rajni Hatti-Kaul. Ectoine production by Halomonas boliviensis: Optimization using response surface methodology. Marine Biotechnology, 12(5):586– 593, October 2010. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-009-9246-6.

Seneca:2010:PGE

[891] Francois O. Seneca, Sylvain Forêt, Eldon E. Ball, Carolyn Smith-Keune, David J. Miller, and Madeleine J. H. van Oppen. Patterns of gene expression in a scleractinian coral undergoing natural bleaching. Marine Biotechnology, 12(5):594–604, October 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-009-9247-5.

Reimer:2010:UHA

[892] James Davis Reimer, Shu Nakachi, Mamiko Hirose, Euichi Hirose, and Shinji Hashiguchi. Using hydrofluoric acid for morphological investigations of zoanthids (Cnidaria: Anthozoa): a critical assessment of methodology and necessity. Marine Biotechnology, 12(5):605–617, October 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-009-9249-3.

Anemaet:2010:APP

[893] Ida G. Anemaet, Martijn Bekker, and Klaas J. Hellingwerf. Algal photosynthesis as the primary driver for a sustainable development in energy, feed, and food production. Marine Biotechnology, 12(6):619– 629, November 2010. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9311-1.

Gwak:2010:APG

[894] In Gyu Gwak, Woong sic Jung, Hak Jun Kim, Sung-Ho Kang, and Eon-Seon Jin. Antifreeze protein in Antarctic marine diatom, Chaetoceros neogracile. Marine Biotechnology, 12(6):630–639, November 2010. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-009-9250-x.

Zhan:2010:TEW

[895] Huiqing Zhan, Jan M. Spitsbergen, Wei Qing, Yi Lian Wu, Thomas A. Paul, James W. Casey, Guor Muor Her, and Zhiyuan Gong. Transgenic expression of walleye dermal sarcoma virus rv-cyclin gene in zebrafish and its suppressive effect on liver tumor development after carcinogen treatment. Marine Biotechnology, 12(6):640–649, November 2010. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-009-9251-9.

Richards:2010:SCH

[896] Robert C. Richards, Connie E. Short, William R. Driedzic, and K. Vanya Ewart. Seasonal changes in hepatic gene expression reveal modulation of multiple processes in rainbow smelt (Osmerus mordax). Marine Biotechnology, 12(6):650–663, November 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-009-9252-8.

Callaghan:2010:ESR

[897] Tamera R. Callaghan, Bernard M. Degnan, and Melony J. Sellars. Expression of sex and reproduction-related genes in Marsupenaeus japonicus. Marine Biotechnology, 12(6):664–677, November 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-009-9254-6.

Wang:2010:EPT

[898] Bo Wang, Zhao Lan Mo, Peng Xiao, Jie Li, Yu Xia Zou, Bin Hao, and Gui Yang Li. EseD, a putative T3SS translocon component of Edwardsiella tarda, contributes to virulence in fish and is a candidate for vaccine development. Marine Biotechnology, 12(6):678–685, November 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-

2236 (electronic). URL https://link.springer.com/article/10.
1007/s10126-009-9255-5.

Xiong:2010:DEO

[899] Xiao-Peng Xiong, Chao Wang, Ming-Zhi Ye, Tian-Ci Yang, Xuan-Xian Peng, and Hui Li. Differentially expressed outer membrane proteins of Vibrio alginolyticus in response to six types of antibiotics. Marine Biotechnology, 12(6):686–695, November 2010. CODEN MABIFW.

ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-009-9256-4.

Rojo:2010:ACD

[900] Liliana Rojo, Adriana Muhlia-Almazan, Reinhard Saborowski, and Fernando García-Carreño. Aspartic cathepsin D endopeptidase contributes to extracellular digestion in clawed lobsters Homarus americanus and Homarus gammarus. Marine Biotechnology, 12(6):696–707, November 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-010-9257-3.

Zijffers:2010:MPY

[901] Jan-Willem F. Zijffers, Klaske J. Schippers, Ke Zheng, Marcel Janssen, Johannes Tramper, and René H. Wijffels. Maximum photosynthetic yield of green microalgae in photobioreactors. Marine Biotechnology, 12(6): 708–718, November 2010. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9258-2.

Hu:2010:CPA

[902] Yongfei Hu, Chengzhang Fu, Yeshi Yin, Gong Cheng, Fang Lei, Xi Yang, Jing Li, Elizabeth Jane Ashforth, Lixin Zhang, and Baoli Zhu. Construction and preliminary analysis of a deep-sea sediment metagenomic fosmid library from Qiongdongnan Basin, South China Sea. Marine Biotechnology, 12(6):719–727, November 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-010-9259-1.

Cengiz:2010:PEM

[903] Sevilay Cengiz and Levent Cavas. A promising evaluation method for dead leaves of Posidonia oceanica (L.) in the adsorption of methyl violet. Marine Biotechnology, 12(6):728–736, November 2010. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-010-9260-8.

Pettit:2011:CSM

[904] Robin K. Pettit. Culturability and secondary metabolite diversity of extreme microbes: Expanding contribution of deep sea and deep-sea vent microbes to natural product discovery. Marine Biotechnology, 13 (1):1–11, February 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9294-y.

Huang:2011:CFI

[905] Jian-Zhong Huang, Xian-Zhang Jiang, Xiao-Feng Xia, Ai-Qun Yu, Ruo-Yu Mao, Xiao-Feng Chen, and Bao-Yu Tian. Cloning and functional identification of Δ5 fatty acid desaturase gene and its 5′-upstream region from marine fungus Thraustochytrium sp. FJN-10. Marine Biotechnology, 13(1):12–21, February 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-010-9262-6.

Santigosa:2011:CTE

[906] Ester Santigosa, Florian Geay, Thierry Tonon, Herve Le Delliou, Heiner Kuhl, Richard Reinhardt, Laurent Corcos, Chantal Cahu, José Luis Zambonino-Infante, and David Mazurais. Cloning, tissue expression analysis, and functional characterization of two Δ6-desaturase variants of sea bass (Dicentrarchus labrax L.). Marine Biotechnology, 13(1): 22–31, February 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9264-4.

Yuyama:2011:PDG

[907] Ikuko Yuyama, Toshiki Watanabe, and Yoshio Takei. Profiling differential gene expression of symbiotic and aposymbiotic corals using a high coverage gene expression profiling (HiCEP) analysis. Marine Biotechnology, 13(1):32–40, February 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-010-9265-3.

Aloise:2011:PMP

[908] Débora de Almeida Aloise, Francisco de Assis Maia-Lima, Ruth Medeiros de Oliveira, Thiago de Melo Cabral, and Wagner Franco Molina. Ploidy manipulation and polyploid detection in the white shrimp Litopenaeus vannamei (Boone 1931) (Decapoda, Penaeidae). Marine Biotechnology, 13(1):41–47, February 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9266-2.

Inoue:2011:CQP

[909] Nariaki Inoue, Ryo Ishibashi, Takashi Ishikawa, Takashi Atsumi, Hideo Aoki, and Akira Komaru. Can the quality of pearls from the Japanese pearl oyster (Pinctada fucata) be explained by the gene expression patterns of the major shell matrix proteins in the pearl sac? Marine Biotechnology, 13(1):48–55, February 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-010-9267-1.

Sallum:2011:MCC

[910] Ulysses W. Sallum and Thomas T. Chen. Molecular cloning of cecropin B responsive endonucleases in Yersinia ruckeri. Marine Biotechnology, 13 (1):56–65, February 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9269-z.

Ahn:2011:EEE

[911] Ginnae Ahn, Eunjin Park, Won-Woo Lee, Jin-Won Hyun, Ki-Wan Lee, Taekyun Shin, You-Jin Jeon, and Youngheun Jee. Enzymatic extract from Ecklonia cava induces the activation of lymphocytes by IL-2 production through the classical NF-κ B pathway. Marine Biotechnology, 13 (1):66–73, February 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9270-6.

Wang:2011:MQA

[912] C. M. Wang, L. C. Lo, Z. Y. Zhu, H. Y. Pang, H. M. Liu, J. Tan, H. S. Lim, R. Chou, L. Orban, and G. H. Yue. Mapping QTL for an adaptive trait: The length of caudal fin in Lates calcarifer. Marine Biotechnology, 13(1):74–82, February 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9271-5.

Kayano:2011:ECP

[913] Keisuke Kayano, Kazuko Saruwatari, Toshihiro Kogure, and Yoshihiro Shiraiwa. Effect of coccolith polysaccharides isolated from the coccolithophorid, Emiliania huxleyi, on calcite crystal formation in In Vitro CaCO_{3} crystallization. Marine Biotechnology, 13(1):83–92, February 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-010-9272-4.

Ma:2011:SQO

[914] Yu-Bin Ma, Zhi-Feng Zhang, Ming-Yu Shao, Kyoung-Ho Kang, Zhi Tan, and Jin-Long Li. Sulfide:quinone oxidoreductase from echiuran worm Urechis unicinctus. Marine Biotechnology, 13(1):93–107, February 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-010-9273-3.

Natrah:2011:DBC

[915] F. M. I. Natrah, Tom Defoirdt, Patrick Sorgeloos, and Peter Bossier. Disruption of bacterial cell-to-cell communication by marine organisms and its relevance to aquaculture. Marine Biotechnology, 13(2): 109–126, April 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9346-3.

Kim:2011:DTG

[916] Woo-Jin Kim, Hyungtaek Jung, and Patrick M. Gaffney. Development of type I genetic markers from expressed sequence tags in highly polymorphic species. Marine Biotechnology, 13(2):127–132, April 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-010-9280-4.

Nagano:2011:DEC

[917] Naoki Nagano, Shou Matsui, Tomoyo Kuramura, Yousuke Taoka, Daiske Honda, and Masahiro Hayashi. The distribution of extracellular cellulase activity in marine eukaryotes, thraustochytrids. Marine Biotechnology, 13(2):133–136, April 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9297-8.

Nishiguchi:2011:DPR

[918] Yoshikazu Nishiguchi, Fumiyoshi Abe, and Mitsumasa Okada. Different pressure resistance of lactate dehydrogenases from hagfish is dependent on habitat depth and caused by tetrameric structure dissociation. Marine Biotechnology, 13(2):137–141, April 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-010-9299-6.

Paul:2011:TEW

[919] Thomas A. Paul, Joel Rovnak, Sandra L. Quackenbush, Kathleen Whitlock, Huiqing Zhan, Zhiyuan Gong, Jan Spitsbergen, Paul R. Bowser, and James W. Casey. Transgenic expression of walleye dermal sarcoma virus rv-cyclin (orfA) in zebrafish does not result in tissue proliferation. Marine Biotechnology, 13(2):142–150, April 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-010-9274-2.

Zhang:2011:IAM

[920] Guoqiang Zhang, Wuying Chu, Songnian Hu, Tao Meng, Linlin Pan, Renxue Zhou, Zhen Liu, and Jianshe Zhang. Identification and analysis of muscle-related protein isoforms expressed in the white muscle of the mandarin fish (Siniperca chuatsi). Marine Biotechnology, 13 (2):151–162, April 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9275-1.

Treerattrakool:2011:IOM

[921] Supattra Treerattrakool, Sakol Panyim, and Apinunt Udomkit. Induction of ovarian maturation and spawning in Penaeus monodon broodstock by double-stranded RNA. Marine Biotechnology, 13(2):163–169, April 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-010-9276-0.

DeSantis:2011:NRQ

[922] Christian De Santis, Carolyn Smith-Keune, and Dean R. Jerry. Normalizing RT-qPCR data: Are we getting the right answers? An appraisal of normalization approaches and internal reference genes from a case study in the finfish Lates calcarifer. Marine Biotechnology, 13 (2):170–180, April 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9277-z.

Cunha:2011:EPP

[923] Regina L. Cunha, Françoise Blanc, François Bonhomme, and Sophie Arnaud-Haond. Evolutionary patterns in pearl oysters of the genus Pinctada (Bivalvia: Pteriidae). Marine Biotechnology, 13(2):181–192, April 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-010-9278-y.

Peng:2011:CEG

[924] Ying Peng, Guang-Lei Liu, Xin-Jun Yu, Xiang-Hong Wang, Li Jing, and Zhen-Ming Chi. Cloning of exo-β-1,3-glucanase gene from a marine yeast Williopsis saturnus and its overexpression in Yarrowia lipolytica. Marine Biotechnology, 13(2):193–204, April 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-010-9281-3.

Oh:2011:EUE

[925] Sung-Ho Oh, Juhee Ahn, Do-Hyung Kang, and Hyeon-Yong Lee. The effect of ultrasonificated extracts of Spirulina maxima on the anticancer activity. Marine Biotechnology, 13(2):205–214, April 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-010-9282-2.

Xu:2011:TPE

[926] Peng Xu, Lauren M. McIntyre, Julie Scardina, Paul A. Wheeler, Gary H. Thorgaard, and Krista M. Nichols. Transcriptome profiling of embryonic development rate in rainbow trout advanced backcross introgression lines. Marine Biotechnology, 13(2):215–231, April 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-010-9283-1.

Li:2011:CGI

[927] Jieying Li, Keith A. Boroevich, Ben F. Koop, and William S. Davidson. Comparative genomics identifies candidate genes for infectious salmon anemia (ISA) resistance in Atlantic salmon (Salmo salar). Marine Biotechnology, 13(2):232–241, April 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-010-9284-0.

Bowman:2011:IAG

[928] Sharen Bowman, Sophie Hubert, Brent Higgins, Cynthia Stone, Jennifer Kimball, Tudor Borza, Jillian Tarrant Bussey, Gary Simpson, Catherine Kozera, Bruce A. Curtis, Jennifer R. Hall, Tiago S. Hori, Charles Y. Feng, Marlies Rise, Marije Booman, A. Kurt Gamperl, Edward Trippel, Jane Symonds, Stewart C. Johnson, and Matthew L. Rise. An integrated approach to gene discovery and marker development in Atlantic cod (Gadus morhua). Marine Biotechnology, 13(2):242–255, April 2011. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-010-9285-z.

Inoue:2011:PPL

[929] Akira Inoue, Chieco Mashino, Teina Kodama, and Takao Ojima. Protoplast preparation from Laminaria japonica with recombinant alginate lyase and cellulase. Marine Biotechnology, 13(2):256–263, April 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-010-9290-2.

Snell:2011:EDE

[930] Terry W. Snell, Tonya L. Shearer, and Hilary A. Smith. Exposure to dsRNA elicits RNA interference in Brachionus manjavacas (Rotifera). Marine Biotechnology, 13(2):264–274, April 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-010-9295-x.

Wang:2011:DES

[931] Aimin Wang, Yan Wang, Zhifeng Gu, Sifa Li, Yaohua Shi, and Ximing Guo. Development of expressed sequence tags from the pearl oyster, Pinctada martensii Dunker. Marine Biotechnology, 13(2):275–283, April 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-010-9296-9.

Rivera-Perez:2011:PBC

[932] Crisalejandra Rivera-Pérez, Fernando L. García-Carreño, and Reinhard Saborowski. Purification and biochemical characterization of digestive lipase in whiteleg shrimp. Marine Biotechnology, 13(2):284–295, April 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-010-9298-7.

Webster:2011:BCD

[933] Nicole S. Webster, Rose E. Cobb, Rochelle Soo, Shelley L. Anthony, Christopher N. Battershill, Steve Whalan, and Elizabeth Evans-Illidge. Bacterial community dynamics in the marine sponge Rhopaloeides odorabile under in situ and ex situ cultivation. Marine Biotechnology, 13 (2):296–304, April 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9300-4.

Chow:2011:IND

[934] Seinen Chow, Sayaka Suzuki, Tadashi Matsunaga, Shane Lavery, Andrew Jeffs, and Haruko Takeyama. Investigation on natural diets of larval marine animals using peptide nucleic acid-directed polymerase chain reaction clamping. Marine Biotechnology, 13(2):305–313, April 2011. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-010-9301-3.

Martins:2011:MFA

[935] Mariana Provedel Martins, Ana Maria Mouad, Letícia Boschini, Mirna Helena Regali Seleghim, Lara Durães Sette, and André Luiz Meleiro Porto. Marine fungi Aspergillus sydowii and Trichoderma sp. catalyze the hydrolysis of benzyl glycidyl ether. Marine Biotechnology, 13(2): 314–320, April 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9302-2.

Cengiz:2011:SCC

[936] Sevilay Cengiz, Levent Cavas, Kadir Yurdakoc, and Georg Pohnert. The sesquiterpene caulerpenyne from Caulerpa spp. is a lipoxygenase inhibitor. Marine Biotechnology, 13(2):321–326, April 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-010-9303-1.

Wu:2011:CAB

[937] Qi Wu, Li Li, and Guofan Zhang. Crassostrea angulata bindin gene and the divergence of fucose-binding lectin repeats among three species of Crassostrea. Marine Biotechnology, 13(2):327–335, April 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-010-9304-0.

Guan:2011:VHV

[938] Bo Guan, Hong Ma, Yaping Wang, Yuanlei Hu, Zhongping Lin, Zuoyan Zhu, and Wei Hu. Vitreoscilla hemoglobin (VHb) overexpression increases hypoxia tolerance in zebrafish (Danio rerio). Marine Biotechnology, 13(2):336–344, April 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-010-9305-z.

Rinkevich:2011:CCM

[939] Baruch Rinkevich. Cell cultures from marine invertebrates: New insights for capturing endless stemness. Marine Biotechnology, 13(3): 345–354, June 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9354-3.

Rosic:2011:VHG

[940] Nedeljka N. Rosic, Mathieu Pernice, Mauricio Rodriguez-Lanetty, and Ove Hoegh-Guldberg. Validation of housekeeping genes for gene expression studies in Symbiodinium exposed to thermal and light stress. Marine Biotechnology, 13(3):355–365, June 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-010-9308-9.

Forjan:2011:UMM

[941] Eduardo Forján, Inés Garbayo, Marta Henriques, Jorge Rocha, José M. Vega, and Carlos Vílchez. UV-A mediated modulation of photosynthetic efficiency, xanthophyll cycle and fatty acid production of Nannochloropsis. Marine Biotechnology, 13(3):366–375, June 2011. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-010-9306-y.

Zhang:2011:GLM

[942] Yan Zhang, Peng Xu, Cuiyun Lu, Youyi Kuang, Xiaofeng Zhang, Dingchen Cao, Chao Li, Yumei Chang, Ning Hou, Hengde Li, Shu Wang, and Xiaowen Sun. Genetic linkage mapping and analysis of muscle fiberrelated QTLs in common carp (Cyprinus carpio L.). Marine Biotechnology, 13(3):376–392, June 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-010-9307-x.

Zhang:2011:NOF

[943] Pengying Zhang, Shenghao Liu, Bailin Cong, Guangting Wu, Chenlin Liu, Xuezheng Lin, Jihong Shen, and Xiaohang Huang. A novel omega-3 fatty acid desaturase involved in acclimation processes of polar condition from Antarctic ice algae Chlamydomonas sp. ICE-L. Marine Biotechnology, 13(3):393–401, June 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-010-9309-8.

Ehrlich:2011:SMP

[944] Herman Ehrlich, Yury N. Elkin, Alexandr A. Artoukov, Valentin A. Stonik, Peter P. Safronov, Vasily V. Bazhenov, Denis V. Kurek, Valery P. Varlamov, René Born, Heike Meissner, and Gert Richter. Simple method for preparation of nanostructurally organized spines of sand dollar Scaphechinus mirabilis (Agassiz, 1863). Marine Biotechnology, 13 (3):402–410, June 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9310-2.

Hatada:2011:HPC

[945] Yuji Hatada, Masahiro Mizuno, Zhijun Li, and Yukari Ohta. Hyperproduction and characterization of the ι-carrageenase useful for ιcarrageenan oligosaccharide production from a deep-sea bacterium, Microbulbifer thermotolerans JAMB-A94^{T}, and insight into the unusual catalytic mechanism. Marine Biotechnology, 13(3):411–422, June 2011. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-010-9312-0.

Vizel:2011:NMC

[946] Maya Vizel, Yossi Loya, Craig A. Downs, and Esti Kramarsky-Winter. A novel method for coral explant culture and micropropagation. Marine Biotechnology, 13(3):423–432, June 2011. CODEN MABIFW.

ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-010-9313-z.

Go:2011:PCF

[947] Hiroe Go, Hye-Jung Hwang, and Taek-Jeong Nam. Polysaccharides from Capsosiphon fulvescens stimulate the growth of IEC-6 cells by activating the MAPK signaling pathway. Marine Biotechnology, 13 (3):433–440, June 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9314-y.

Zhang:2011:LNS

[948] Jun Zhang and Shicui Zhang. Lipovitellin is a non-self recognition receptor with opsonic activity. Marine Biotechnology, 13(3):441–450, June 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-010-9315-x.

Kuo:2011:IIN

[949] Yao-Haur Kuo, Tzu-Wen Liang, Kao-Cheng Liu, Ya-Wen Hsu, Hsiu-Ching Hsu, and San-Lang Wang. Isolation and identification of a novel antioxidant with antitumour activity from Serratia ureilytica using squid pen as fermentation substrate. Marine Biotechnology, 13 (3):451–461, June 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9316-9.

Hong:2011:PEE

[950] Il-Hwa Hong, Hoon Ji, Sung-Yong Hwa, Won-Il Jeong, Da-Hee Jeong, Sun-Hee Do, Ji-Min Kim, Mi-Ran Ki, Jin-Kyu Park, Moon-Jung Goo, Ok-Kyung Hwang, Kyung-Sook Hong, Jung-Youn Han, Hae-Young Chung, and Kyu-Shik Jeong. The protective effect of ENA actimineral resource A on CCl_{4}-induced liver injury in rats. Marine Biotechnology, 13(3):462–473, June 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9317-8.

Inoue:2011:GEP

[951] Nariaki Inoue, Ryo Ishibashi, Takashi Ishikawa, Takashi Atsumi, Hideo Aoki, and Akira Komaru. Gene expression patterns in the outer mantle epithelial cells associated with pearl sac formation. Marine Biotechnology, 13(3):474–483, June 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-010-9318-7.

Hennebert:2011:CCF

[952] Elise Hennebert, Ruddy Wattiez, and Patrick Flammang. Characterisation of the carbohydrate fraction of the temporary adhesive secreted by the tube feet of the sea star Asterias rubens. Marine Biotechnology, 13(3):484–495, June 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9319-6.

Jones:2011:SPA

[953] Bethan M. Jones, Richard J. Edwards, Paul J. Skipp, C. David O’Connor, and M. Debora Iglesias-Rodriguez. Shotgun proteomic analysis of Emiliania huxleyi, a marine phytoplankton species of major biogeochemical importance. Marine Biotechnology, 13(3):496–504, June 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-010-9320-0.

Kolaiti:2011:ICS

[954] Regina-Maria Kolaiti, Andrea Baier, Ryszard Szyszka, and Sophia Kouyanou-Koutsoukou. Isolation of a CK2α subunit and the holoenzyme from the mussel Mytilus galloprovincialis and construction of the CK2α and CK2β cDNAs. Marine Biotechnology, 13(3):505–516, June 2011. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-010-9321-z.

Kinoshita:2011:MFA

[955] Shigeharu Kinoshita, Eriko Katsumi, Hiroshi Yamamoto, Kazuharu Takeuchi, and Shugo Watabe. Molecular and functional analyses of aspolin, a fish-specific protein extremely rich in aspartic acid. Marine Biotechnology, 13(3):517–526, June 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-010-9322-y.

Solovchenko:2011:SIC

[956] Alexei Solovchenko, Inna Khozin-Goldberg, Lee Recht, and Sammy Boussiba. Stress-induced changes in optical properties, pigment and fatty acid content of Nannochloropsis sp.: Implications for nondestructive assay of total fatty acids. Marine Biotechnology, 13(3): 527–535, June 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9323-x. See erratum [983].

Sibirtsev:2011:CMD

[957] Juriy T. Sibirtsev, Valeria V. Shastina, Natalia I. Menzorova, Tatyana N. Makarieva, and Valeriy A. Rasskazov. Ca^{2+}, Mg^{2+}-dependent DNase involvement in apoptotic effects in spermatozoa of sea urchin Strongylocentrotus intermedius induced by two-headed sphingolipid rhizochalin. Marine Biotechnology, 13(3):536–543, June 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-010-9324-9.

Contreras-Porcia:2011:ICI

[958] Loretto Contreras-Porcia, Geraldine Dennett, Alberto González, Eva Vergara, Cristóbal Medina, Juan A. Correa, and Alejandra Moenne. Identification of copper-induced genes in the marine alga Ulva compressa (Chlorophyta). Marine Biotechnology, 13(3):544–556, June 2011. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-010-9325-8.

Lee:2011:GPM

[959] Bo-Young Lee, Jean-Pierre Coutanceau, Catherine Ozouf-Costaz, Helena D’Cotta, Jean-Francois Baroiller, and Thomas D. Kocher. Genetic and physical mapping of sex-linked AFLP markers in Nile tilapia (Oreochromis niloticus). Marine Biotechnology, 13(3):557–562, June 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-010-9326-7.

VanWormhoudt:2011:MDI

[960] Alain Van Wormhoudt, Valérie Roussel, Gercende Courtois, and Sylvain Huchette. Mitochondrial DNA introgression in the European abalone Haliotis tuberculata tuberculata: Evidence for experimental mtDNA paternal inheritance and a natural hybrid sequence. Marine Biotechnology, 13(3):563–574, June 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9327-6.

Zhi:2011:ESA

[961] Bin Zhi, Wen Tang, and Xiaobo Zhang. Enhancement of shrimp antiviral immune response through caspase-dependent apoptosis by small molecules. Marine Biotechnology, 13(3):575–583, June 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-010-9328-5.

Warr:2011:ISI

[962] Gregory W. Warr. Introduction to a special issue in memory of Paul S. Gross. Marine Biotechnology, 13(4):585–586, August 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-010-9293-z.

Flegel:2011:SMR

[963] T. W. Flegel and Kallaya Sritunyalucksana. Shrimp molecular responses to viral pathogens. Marine Biotechnology, 13(4):587–607, August 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-010-9287-x.

Leu:2011:RMP

[964] Jiann-Horng Leu, Shu-Hwa Chen, Yu-Bin Wang, Yen-Chen Chen, Sheng-Yao Su, Chung-Yen Lin, Jan-Ming Ho, and Chu-Fang Lo. A review of the major penaeid shrimp EST studies and the construction of a shrimp transcriptome database based on the ESTs from four penaeid shrimp. Marine Biotechnology, 13(4):608–621, August 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-010-9286-y.

Hirono:2011:UMS

[965] Ikuo Hirono, Fernand F. Fagutao, Hidehiro Kondo, and Takashi Aoki. Uncovering the mechanisms of shrimp innate immune response by RNA interference. Marine Biotechnology, 13(4):622–628, August 2011. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-010-9292-0.

Aoki:2011:MAS

[966] Takashi Aoki, Han-Ching Wang, Sasimanas Unajak, Mudjekeewis D. Santos, Hidehiro Kondo, and Ikuo Hirono. Microarray analyses of shrimp immune responses. Marine Biotechnology, 13(4):629–638, August 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-010-9291-1.

Tassanakajon:2011:CAP

[967] Anchalee Tassanakajon, Piti Amparyup, Kunlaya Somboonwiwat, and Premruethai Supungul. Cationic antimicrobial peptides in penaeid shrimp. Marine Biotechnology, 13(4):639–657, August 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-011-9381-8.

Osinga:2011:BEC

[968] Ronald Osinga, Miriam Schutter, Ben Griffioen, René H. Wijffels, Johan A. J. Verreth, Shai Shafir, Stéphane Henard, Maura Taruffi, Claudia Gili, and Silvia Lavorano. The biology and economics of coral growth. Marine Biotechnology, 13(4):658–671, August 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-011-9382-7.

Ammayappan:2011:RGS

[969] Arun Ammayappan, Gael Kurath, Tarin M. Thompson, and Vikram N. Vakharia. A reverse genetics system for the Great Lakes strain of viral hemorrhagic septicemia virus: the NV gene is required for pathogenicity. Marine Biotechnology, 13(4):672–683, August 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-010-9329-4.

Sung:2011:IRG

[970] Ming-Shiuan Sung, Yuan-Ting Hsu, Kuan-Lin Ho, and Tse-Min Lee. Implications of the up-regulation of genes encoding protein degradation enzymes and heat shock protein 90 for intertidal green macroalga Ulva fasciata against hypersalinity-induced protein oxidation. Marine Biotechnology, 13(4):684–694, August 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-010-9330-y.

Gallardo-Galvez:2011:ETA

[971] Jose Braulio Gallardo-Gálvez, Teresa Méndez, Julia Béjar, and M. Carmen Alvarez. Endogenous transposases affect differently Sleeping Beauty and Frog Prince transposons in fish cells. Marine Biotechnology, 13(4): 695–705, August 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9331-x.

Li:2011:CCB

[972] Yan Li, Peng Xu, Zixia Zhao, Jian Wang, Yan Zhang, and Xiao-Wen Sun. Construction and characterization of the BAC library for common carp Cyprinus carpio L. and establishment of microsynteny with zebrafish Danio rerio. Marine Biotechnology, 13(4):706–712, August 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/

s10126-010-9332-9. See erratum [1019].

Ding:2011:RPD

[973] Bo Ding, Ying Yin, Fengli Zhang, and Zhiyong Li. Recovery and phylogenetic diversity of culturable fungi associated with marine sponges Clathrina luteoculcitella and Holoxea sp. in the South China Sea. Marine Biotechnology, 13(4):713–721, August 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-010-9333-8.

Teasdale:2011:GPM

[974] Margaret E. Teasdale, Kellye A. Donovan, Stephanie R. Forschner-Dancause, and David C. Rowley. Gram-positive marine bacteria as a potential resource for the discovery of quorum sensing inhibitors. Marine Biotechnology, 13(4):722–732, August 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-010-9334-7.

Booman:2011:DEV

[975] Marije Booman, Tudor Borza, Charles Y. Feng, Tiago S. Hori, Brent Higgins, Adrian Culf, Daniel Léger, Ian C. Chute, Anissa Belkaid, Marlies Rise, A. Kurt Gamperl, Sophie Hubert, Jennifer Kimball, Rodney J. Ouellette, Stewart C. Johnson, Sharen Bowman, and Matthew L. Rise. Development and experimental validation of a 20K Atlantic cod (Gadus morhua) oligonucleotide microarray based on a collection of over 150,000 ESTs. Marine Biotechnology, 13(4):733–750, August 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-010-9335-6.

Chung:2011:EGA

[976] Chih-Ching Chung, Jeng Chang, Gwo-Ching Gong, Shih-Chieh Hsu, Kuo-Ping Chiang, and Chia-Wen Liao. Effects of Asian dust storms on Synechococcus populations in the subtropical Kuroshio Current. Marine Biotechnology, 13(4):751–763, August 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-010-9336-5.

Piazza:2011:TRA

[977] Veronica Piazza, Vassilios Roussis, Francesca Garaventa, Giuliano Greco, Vangelis Smyrniotopoulos, Constantinos Vagias, and Marco Faimali. Terpenes from the red alga Sphaerococcus coronopifolius inhibit the settlement of barnacles. Marine Biotechnology, 13(4):764–772, August 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-010-9337-4.

Tramice:2011:HAE

[978] Annabella Tramice, Giuseppina Andreotti, and Antonio Trincone. Hydrosoluble antioxidants by enzymatic glucosylation of a vitamin E derivative using marine α-d-glucosidase from Aplysia fasciata. Marine Biotechnology, 13(4):773–781, August 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-010-9339-2.

Chernogor:2011:LTC

[979] Lubov I. Chernogor, Natalia N. Denikina, Sergey I. Belikov, and Alexander V. Ereskovsky. Long-term cultivation of primmorphs from freshwater Baikal sponges Lubomirskia baikalensis. Marine Biotechnology, 13 (4):782–792, August 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9340-9.

Madera-Santana:2011:PPB

[980] Tomás J. Madera-Santana, Daniel Robledo, and Yolanda Freile-Pelegrín. Physicochemical properties of biodegradable polyvinyl alcohol–agar films from the red algae Hydropuntia cornea. Marine Biotechnology, 13(4): 793–800, August 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9341-8.

Saruwatari:2011:MCT

[981] Kazuko Saruwatari, Seiji Nagasaka, Noriaki Ozaki, and Hiromichi Nagasawa. Morphological and crystallographic transformation from immature to mature coccoliths, Pleurochrysis carterae. Marine Biotechnology, 13 (4):801–809, August 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9342-7.

Veremeichik:2011:OSG

[982] Galina N. Veremeichik, Yuri N. Shkryl, Victor P. Bulgakov, Sergey V. Shedko, Valery B. Kozhemyako, Svetlana N. Kovalchuk, Vladimir B. Krasokhin, Yuri N. Zhuravlev, and Yuri N. Kulchin. Occurrence of a silicatein gene in glass sponges (Hexactinellida: Porifera). Marine Biotechnology, 13(4):810–819, August 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-010-9343-6.

Solovchenko:2011:ESI

[983] Alexei Solovchenko, Inna Khozin-Goldberg, Lee Recht, and Sammy Boussiba. Erratum: Stress-induced changes in optical properties, pigment and fatty acid content of Nannochloropsis sp.: Implications for non-destructive assay of total fatty acids. Marine Biotechnology, 13(4): 820–821, August 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9338-3. See [956].

Souza:2011:ICM

[984] Claudiana P. Souza, Bianca C. Almeida, Rita R. Colwell, and Irma N. G. Rivera. The importance of chitin in the marine environment. Marine Biotechnology, 13(5):823–830, October 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-011-9388-1.

Drillet:2011:DIM

[985] Guillaume Drillet, Tahina Rabarimanantsoa, Stéphane Frouël, Jacob S. Lamson, Anette M. Christensen, Sandra Kim-Tiam, and Benni W. Hansen. Do inactivated microbial preparations improve life history traits of the copepod Acartia tonsa? Marine Biotechnology, 13(5): 831–836, October 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9348-1.

Liu:2011:PCA

[986] Jianguo Liu, Zhiqiang Zhang, Zhiqiang Liu, Hu Zhu, Hongyue Dang, Jianren Lu, and Zhanfeng Cui. Production of cold-adapted amylase by marine bacterium Wangia sp. C52: Optimization, modeling, and partial characterization. Marine Biotechnology, 13(5):837–844, October 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-010-9360-5.

Baharum:2011:MCM

[987] Hariyanti Baharum, Hiroyuki Morita, Akifumi Tomitsuka, Fong-Chin Lee, Kim-Yong Ng, Raha Abdul Rahim, Ikuro Abe, and Chai-Ling Ho. Molecular cloning, modeling, and site-directed mutagenesis of Type III polyketide synthase from Sargassum binderi (Phaeophyta). Marine Biotechnology, 13(5):845–856, October 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-010-9344-5.

Sonthi:2011:DCS

[988] Molruedee Sonthi, Mylène Toubiana, Alberto Pallavicini, Paola Venier, and Philippe Roch. Diversity of coding sequences and gene structures of the antifungal peptide mytimycin (MytM) from the Mediterranean mussel, Mytilus galloprovincialis. Marine Biotechnology, 13(5): 857–867, October 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9345-4.

Song:2011:CLD

[989] Yue-Fan Song, Yi Qu, Xu-Peng Cao, and Wei Zhang. Cellular localization of debromohymenialdisine and hymenialdisine in the marine sponge Axinella sp. using a newly developed cell purification protocol. Marine Biotechnology, 13(5):868–882, October 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-010-9347-2.

Tabares:2011:API

[990] Paula Tabares, Sheila M. Pimentel-Elardo, Tanja Schirmeister, Thomas Hünig, and Ute Hentschel. Anti-protease and immunomodulatory activities of bacteria associated with Caribbean sponges. Marine Biotechnology, 13(5):883–892, October 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-010-9349-0.

Al-Alawi:2011:CCE

[991] Ahmed Ali Al-Alawi, Insaaf Mohammed Al-Marhubi, Mohammed Said Moosa Al-Belushi, and Bassam Soussi. Characterization of carrageenan extracted from Hypnea bryoides in Oman. Marine Biotechnology, 13(5):893–899, October 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-010-9350-7.

Nakai:2011:MAM

[992] Ryosuke Nakai, Takashi Abe, Haruko Takeyama, and Takeshi Naganuma. Metagenomic analysis of 0.2-μ m-passable microorganisms in Deep-Sea hydrothermal fluid. Marine Biotechnology, 13(5):900–908, October 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-010-9351-6.

Chang:2011:ARC

[993] Yun-Shiang Chang, Tsan-Chi Chen, Wang-Jing Liu, Jiang-Shiou Hwang, Guang-Hsiung Kou, and Chu-Fang Lo. Assessment of the roles of copepod Apocyclops royi and bivalve mollusk Meretrix lusoria in white spot syndrome virus transmission. Marine Biotechnology, 13(5):909– 917, October 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9352-5.

Sowmya:2011:APR

[994] R. Sowmya, K. Rathinaraj, and N. M. Sachindra. An autolytic process for recovery of antioxidant activity rich carotenoprotein from shrimp heads. Marine Biotechnology, 13(5):918–927, October 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-010-9353-4.

Sirisansaneeyakul:2011:PPL

[995] Sarote Sirisansaneeyakul, Somruethai Singhasuwan, Wanna Choorit, Natapas Phoopat, Jose Luis Garcia, and Yusuf Chisti. Photoautotrophic production of lipids by some Chlorella strains. Marine Biotechnology, 13 (5):928–941, October 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9355-2.

Wood:2011:EGG

[996] Raphael R. Wood and Cova R. Arias. Evaluation of global gene expression during cold shock in the human pathogen Vibrio vulnificus. Marine Biotechnology, 13(5):942–954, October 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-010-9356-1.

Marie:2011:PIN

[997] Benjamin Marie, Nolwenn Trinkler, Isabelle Zanella-Cleon, Nathalie Guichard, Michel Becchi, Christine Paillard, and Frédéric Marin. Proteomic identification of novel proteins from the calcifying shell matrix of the Manila clam Venerupis philippinarum. Marine Biotechnology, 13 (5):955–962, October 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9357-0.

Zhao:2011:SDA

[998] Yan Zhao, Qin Liu, Xiuhua Wang, Lingyun Zhou, Qiyao Wang, and Yuanxing Zhang. Surface display of Aeromonas hydrophila GAPDH in attenuated Vibrio anguillarum to develop a noval multivalent vector vaccine. Marine Biotechnology, 13(5):963–970, October 2011. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-010-9359-y.

Corporeau:2011:FST

[999] Charlotte Corporeau, Agnès Groisillier, Alexandra Jeudy, Tristan Barbeyron, Elodie Fleury, Caroline Fabioux, Mirjam Czjzek, and Arnaud Huvet. A functional study of transforming growth factor-beta from the gonad of Pacific oyster Crassostrea gigas. Marine Biotechnology, 13(5): 971–980, October 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-010-9361-4.

Liu:2011:NME

[1000] Xiaojun Liu, Chang Liu, Lei Chen, Juan Sun, Yujuan Zhou, Qi Li, Guilan Zheng, Guiyou Zhang, Hongzhong Wang, Liping Xie, and Rongqing Zhang. A new method to extract matrix proteins directly from the secretion of the mollusk mantle and the role of these proteins in shell biomineralization. Marine Biotechnology, 13(5):981–991, October 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-011-9362-y.

Bertucci:2011:NCC

[1001] Anthony Bertucci, Sylvie Tambutté, Claudiu T. Supuran, Denis Allemand, and Didier Zoccola. A new coral carbonic anhydrase in Stylophora pistillata. Marine Biotechnology, 13(5):992–1002, October 2011. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-011-9363-x.

Fleury:2011:CNF

[1002] Christophe Fleury, Antoine Serpentini, Magdalini Kypriotou, Emmanuelle Renard, Philippe Galéra, and Jean-Marc Lebel. Characterization of a non-fibrillar-related collagen in the mollusc Haliotis tuberculata and its biological activity on human dermal fibroblasts. Marine Biotechnology, 13(5):1003–1016, October 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-011-9364-9.

Zhang:2011:IGG

[1003] Hong Zhang, ShaoJun Liu, Chun Zhang, Min Tao, LiangYue Peng, CuiPing You, Jun Xiao, Yi Zhou, GongJian Zhou, KaiKun Luo, and Yun Liu.

Induced gynogenesis in grass carp (Ctenopharyngodon idellus) using irradiated sperm of allotetraploid hybrids. Marine Biotechnology, 13(5): 1017–1026, October 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-011-9365-8.

Galindo:2011:GSI

[1004] Heather M. Galindo, Timothy Loher, and Lorenz Hauser. Genetic sex identification and the potential evolution of sex determination in Pacific halibut (Hippoglossus stenolepis). Marine Biotechnology, 13(5): 1027–1037, October 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-011-9366-7.

Hirata:2011:TGE

[1005] Ryo Hirata, Megumu Takahashi, Naotsune Saga, and Koji Mikami. Transient gene expression system established in Porphyra yezoensis is widely applicable in bangiophycean algae. Marine Biotechnology, 13(5): 1038–1047, October 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-011-9367-6.

Sun:2011:SCE

[1006] Hai-Hong Sun, Wen-Jun Mao, Jie-Ying Jiao, Jia-Chao Xu, Hong-Yan Li, Yin Chen, Xiao-Hui Qi, Yan-Li Chen, Jian Xu, Chun-Qi Zhao, Yu-Jiao Hou, and Yu-Pin Yang. Structural characterization of extracellular polysaccharides produced by the marine fungus Epicoccum nigrum JJY-40 and their antioxidant activities. Marine Biotechnology, 13(5): 1048–1055, October 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-011-9368-5.

Taylor:2011:MRI

[1007] Michael W. Taylor, Russell T. Hill, and Ute Hentschel. Meeting report: 1st international symposium on sponge microbiology. Marine Biotechnology, 13(6):1057–1061, December 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-011-9397-0.

Vynne:2011:BCP

[1008] Nikolaj G. Vynne, Maria Månsson, Kristian F. Nielsen, and Lone Gram. Bioactivity, chemical profiling, and 16S rRNA-based phylogeny of Pseudoalteromonas strains collected on a global research cruise. Marine Biotechnology, 13(6):1062–1073, December 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-011-9369-4.

Ferreira:2011:ENS

[1009] Martiña Ferreira, Pedro Seixas, Paula Coutinho, Jaime Fábregas, and Ana Otero. Effect of the nutritional status of semi-continuous microalgal cultures on the productivity and biochemical composition of Brachionus plicatilis. Marine Biotechnology, 13(6):1074–1085, December 2011. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-011-9370-y.

Hwang:2011:LMT

[1010] Seong Don Hwang, Kanako Fuji, Tomokazu Takano, Takashi Sakamoto, Hidehiro Kondo, Ikuo Hirono, and Takashi Aoki. Linkage mapping of toll-like receptors (TLRs) in Japanese flounder, Paralichthys olivaceus. Marine Biotechnology, 13(6):1086–1091, December 2011. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-011-9371-x.

Khalesi:2011:EBC

[1011] Mohammad K. Khalesi, H. H. Beeftink, and R. H. Wijffels. Energy budget for the cultured, zooxanthellate octocoral Sinularia flexibilis. Marine Biotechnology, 13(6):1092–1098, December 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-011-9373-8.

Millan:2011:GEP

[1012] Adrián Millán, Antonio Gómez-Tato, Belén G. Pardo, Carlos Fernández, Carmen Bouza, Manuel Vera, José Antonio Alvarez-Dios, Santiago Cabaleiro, Jesús Lamas, Manuel L. Lemos, and Paulino Martínez. Gene expression profiles of the spleen, liver, and head kidney in turbot (Scophthalmus maximus) along the infection process with Aeromonas salmonicida using an immune-enriched oligo-microarray. Marine Biotechnology, 13(6):1099–1114, December 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-011-9374-7.

Sarropoulou:2011:FDG

[1013] Elena Sarropoulou, Dimitra Nousdili, Georgios Kotoulas, and Antonios Magoulas. Functional divergences of GAPDH isoforms during early development in two Perciform fish species. Marine Biotechnology, 13(6): 1115–1124, December 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-011-9375-6.

vanderMerwe:2011:DGR

[1014] Mathilde van der Merwe, Paolo Franchini, and Rouvay Roodt-Wilding. Differential growth-related gene expression in abalone (Haliotis midae). Marine Biotechnology, 13(6):1125–1139, December 2011. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-011-9376-5.

Minami:2011:MCA

[1015] Seiko Minami, Minoru Sato, Yoshihiro Shiraiwa, and Koji Iwamoto. Molecular characterization of adenosine 5′-monophosphate deaminase — the key enzyme responsible for the umami taste of nori (Porphyra yezoensis Ueda, Rhodophyta). Marine Biotechnology, 13(6):1140– 1147, December 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-011-9377-4.

Bayer:2011:ABC

[1016] Mirko Bayer, Claire Hellio, Jean-Philippe Maréchal, Walter Frank, Wenhan Lin, Horst Weber, and Peter Proksch. Antifouling bastadin congeners target mussel phenoloxidase and complex copper(II) ions. Marine Biotechnology, 13(6):1148–1158, December 2011. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-011-9378-3.

Marie:2011:NPC

[1017] Benjamin Marie, Isabelle Zanella-Cléon, Nathalie Guichard, Michel Becchi, and Frédéric Marin. Novel proteins from the calcifying shell matrix of the Pacific oyster Crassostrea gigas. Marine Biotechnology, 13(6): 1159–1168, December 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-011-9379-2.

Bergman:2011:MBC

[1018] Oded Bergman, Markus Haber, Boaz Mayzel, Matthew A. Anderson, Muki Shpigel, Russell T. Hill, and Micha Ilan. Marine-based cultivation of Diacarnus sponges and the bacterial community composition of wild and maricultured sponges and their larvae. Marine Biotechnology, 13(6): 1169–1182, December 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-011-9391-6.

Li:2011:ECC

[1019] Yan Li, Peng Xu, Zixia Zhao, Jian Wang, Yan Zhang, and Xiao-Wen Sun. Erratum to: Construction and characterization of the BAC library for common carp Cyprinus carpio L. and establishment of microsynteny with zebrafish Danio rerio. Marine Biotechnology, 13(6): 1183, December 2011. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-011-9404-5. See [972].

Li:2012:DLA

[1020] Hongjun Li, Xiao Liu, and Guofan Zhang. Development and linkage analysis of 104 new microsatellite markers for Bay scallop (Argopecten irradians). Marine Biotechnology, 14(1):1–9, February 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-011-9383-6.

Umemoto:2012:XIM

[1021] Yoshiaki Umemoto, Toshiyuki Shibata, and Toshiyoshi Araki. d-xylose isomerase from a marine bacterium, Vibrio sp. strain XY-214, and dxylulose production from β-1,3-xylan. Marine Biotechnology, 14(1): 10–20, February 2012. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-011-9380-9.

Zou:2012:MBP

[1022] Shanmei Zou, Qi Li, and Lingfeng Kong. Multigene barcoding and phylogeny of geographically widespread muricids (Gastropoda: Neogastropoda) along the Coast of China. Marine Biotechnology, 14(1): 21–34, February 2012. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-011-9384-5.

Tacchi:2012:MSR

[1023] Luca Tacchi, Ralph Bickerdike, Christopher J. Secombes, and Samuel A. M. Martin. Muscle-specific RING finger (MuRF) cDNAs in Atlantic salmon (Salmo salar) and their role as regulators of muscle protein degradation. Marine Biotechnology, 14(1):35–45, February 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-011-9385-4.

Salgado:2012:HNF

[1024] María C. Salgado, Isidoro Metón, Ida G. Anemaet, J. Diego González, Felipe Fernández, and Isabel V. Baanante. Hepatocyte nuclear factor 4α transactivates the mitochondrial alanine aminotransferase gene in the kidney of Sparus aurata. Marine Biotechnology, 14(1):46–62, February 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-

2236 (electronic). URL https://link.springer.com/article/10.
1007/s10126-011-9386-3.

Zhu:2012:PSA

[1025] Fei Zhu and Xiaobo Zhang. Protection of shrimp against white spot syndrome virus (WSSV) with β-1,3-d-glucan-encapsulated vp28-siRNA particles. Marine Biotechnology, 14(1):63–68, February 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-011-9387-2.

Huan:2012:TAC

[1026] Pin Huan, Hongxia Wang, and Baozhong Liu. Transcriptomic analysis of the clam Meretrix meretrix on different larval stages. Marine Biotechnology, 14(1):69–78, February 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-011-9389-0.

Zhan:2012:CIM

[1027] Xin Zhan, Feilong Fan, Weiwei You, Jinjin Yu, and Caihuan Ke. Construction of an integrated map of Haliotis diversicolor using microsatellite markers. Marine Biotechnology, 14(1):79–86, February 2012. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-011-9390-7.

Balabanova:2012:EST

[1028] Larissa A. Balabanova, Yury M. Gafurov, Mikhael V. Pivkin, Natalya A. Terentyeva, Galina N. Likhatskaya, and Valery A. Rasskazov. An extracellular S1-type nuclease of marine fungus Penicillium melinii. Marine Biotechnology, 14(1):87–95, February 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-011-9392-5.

Wang:2012:DMW

[1029] Xiaohong Wang, Lu Gan, Matthias Wiens, Ute Schloßmacher, Heinz C. Schröder, and Werner E. G. Müller. Distribution of microfossils within polymetallic nodules: Biogenic clusters within manganese layers. Marine Biotechnology, 14(1):96–105, February 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-011-9393-4.

Kawabe:2012:RHI

[1030] Shinya Kawabe and Yoshihiro Yokoyama. Role of hypoxia-inducible factor α in response to hypoxia and heat shock in the Pacific oyster Crassostrea gigas. Marine Biotechnology, 14(1):106–119, February 2012. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-011-9394-3.

Chen:2012:IMD

[1031] Song-Lin Chen, Xiang-Shan Ji, Chang-Wei Shao, Wen-Long Li, Jing-Feng Yang, Zuo Liang, Xiao-Lin Liao, Gen-Bo Xu, Ying Xu, and Wen-Tao Song. Induction of mitogynogenetic diploids and identification of WW super-female using sex-specific SSR markers in half-smooth tongue sole (Cynoglossus semilaevis). Marine Biotechnology, 14(1):120– 128, February 2012. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-011-9395-2.

Ebenezer:2012:MDQ

[1032] Vinitha Ebenezer, Linda K. Medlin, and Jang-Seu Ki. Molecular detection, quantification, and diversity evaluation of microalgae. Marine Biotechnology, 14(2):129–142, April 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-011-9427-y.

Samino:2012:INV

[1033] Setijono Samino, Hitoshi Michibata, and Tatsuya Ueki. Identification of a novel vanadium-binding protein by EST analysis on the most vanadium-rich ascidian, Ascidia gemmata. Marine Biotechnology, 14 (2):143–154, April 2012. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-011-9396-1.

Yasuike:2012:GRS

[1034] Motoshige Yasuike, Jong Leong, Stuart G. Jantzen, Kristian R. von Schalburg, Frank Nilsen, Simon R. M. Jones, and Ben F. Koop. Genomic resources for sea lice: Analysis of ESTs and mitochondrial genomes. Marine Biotechnology, 14(2):155–166, April 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-011-9398-z.

Drivenes:2012:GEP

[1035] Øyvind Drivenes, Geir Lasse Taranger, and Rolf B. Edvardsen. Gene expression profiling of Atlantic cod (Gadus morhua) embryogenesis using microarray. Marine Biotechnology, 14(2):167–176, April 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-011-9399-y.

Wang:2012:MPF

[1036] Ning Wang, Shigeharu Kinoshita, Naoko Nomura, Chihiro Riho, Kaoru Maeyama, Kiyohito Nagai, and Shugo Watabe. The mining of pearl formation genes in pearl oyster Pinctada fucata by cDNA suppression subtractive hybridization. Marine Biotechnology, 14(2):177–188, April 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-011-9400-9.

Park:2012:CCN

[1037] Hwan Hee Park, Natania Kam, Eun Yeol Lee, and Hee Sook Kim. Cloning and characterization of a novel oligoalginate lyase from a newly isolated bacterium Sphingomonas sp. MJ-3. Marine Biotechnology, 14 (2):189–202, April 2012. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-011-9402-7.

Fleury:2012:MAH

[1038] Elodie Fleury and Arnaud Huvet. Microarray analysis highlights immune response of Pacific oysters as a determinant of resistance to summer mortality. Marine Biotechnology, 14(2):203–217, April 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-011-9403-6.

Guo:2012:GMQ

[1039] Xiang Guo, Qi Li, Qing Z. Wang, and Ling F. Kong. Genetic mapping and QTL analysis of growth-related traits in the Pacific oyster. Marine Biotechnology, 14(2):218–226, April 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-011-9405-4.

Teerawanichpan:2012:MFA

[1040] Prapapan Teerawanichpan and Xiao Qiu. Molecular and functional analysis of three fatty Acyl–CoA reductases with distinct substrate specificities in copepod Calanus finmarchicus. Marine Biotechnology, 14 (2):227–236, April 2012. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-011-9406-3.

Liu:2012:MLM

[1041] Sixin Liu, Caird E. Rexroad III, and Craig V. Sullivan. A microsatellite linkage map of striped bass (Morone saxatilis) reveals conserved synteny with the three-spined stickleback (Gasterosteus aculeatus). Marine Biotechnology, 14(2):237–244, April 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-011-9407-2.

Ma:2012:RSQ

[1042] Yu-Bin Ma, Zhi-Feng Zhang, Ming-Yu Shao, Kyoung-Ho Kang, Xiao-Li Shi, Ying-Ping Dong, and Jin-Long Li. Response of sulfide:quinone oxidoreductase to sulfide exposure in the echiuran worm Urechis unicinctus. Marine Biotechnology, 14(2):245–251, April 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-011-9408-1.

Liu:2012:IPC

[1043] Yang Liu, Yin Lei, Xuecheng Zhang, Yi Gao, Yazhong Xiao, and Hui Peng. Identification and phylogenetic characterization of a new subfamily of α-amylase enzymes from marine microorganisms. Marine Biotechnology, 14(3):253–260, June 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-011-9414-3.

Xu:2012:DGE

[1044] Jin-Li Xu, Xue Zhang, Huai-Yong Sun, and Zhen-Ming Chi. Disruption of the gene encoding β-1, 3-glucanase in marine-derived Williopsis saturnus WC91-2 enhances its killer toxin activity. Marine Biotechnology, 14(3):261–269, June 2012. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-011-9409-0.

Hodges:2012:UAM

[1045] Tyler W. Hodges, Marc Slattery, and Julie B. Olson. Unique actinomycetes from marine caves and coral reef sediments provide novel PKS and NRPS biosynthetic gene clusters. Marine Biotechnology, 14 (3):270–280, June 2012. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-011-9410-7.

Pozzolini:2012:MCN

[1046] Marina Pozzolini, Federica Bruzzone, Valentina Berilli, Francesca Mussino, Carlo Cerrano, Umberto Benatti, and Marco Giovine. Molecular characterization of a nonfibrillar collagen from the marine sponge Chondrosia reniformis Nardo 1847 and positive effects of soluble silicates on its expression. Marine Biotechnology, 14(3):281–293, June 2012. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-011-9415-2.

Suzuki:2012:CAF

[1047] Tomohiko Suzuki, Kentaro Yamamoto, Hiroshi Tada, and Kouji Uda. Cold-adapted features of arginine kinase from the deep-sea clam Calyptogena kaikoi. Marine Biotechnology, 14(3):294–303, June 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-011-9411-6.

Skriptsova:2012:CSP

[1048] Anna V. Skriptsova, Nataliya M. Shevchenko, Daria V. Tarbeeva, and Tatiana N. Zvyagintseva. Comparative study of polysaccharides from reproductive and sterile tissues of five brown seaweeds. Marine Biotechnology, 14(3):304–311, June 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-011-9413-4.

Kim:2012:AAT

[1049] Yeon-Mi Kim, Ying Wu, Thi Uyen Duong, Seul-Gi Jung, Si Wouk Kim, Hoon Cho, and EonSeon Jin. Algicidal activity of thiazolidinedione derivatives against harmful algal blooming species. Marine Biotechnology, 14(3):312–322, June 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-011-9412-5.

Numata:2012:BPN

[1050] Keiji Numata and Yoshiharu Doi. Biosynthesis of polyhydroxyalkanaotes by a novel facultatively anaerobic Vibrio sp. under marine conditions. Marine Biotechnology, 14(3):323–331, June 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-011-9416-1.

Park:2012:HSP

[1051] Hong-Sil Park, Won-Joong Jeong, EuiCheol Kim, Youngja Jung, Jong Min Lim, Mi Sook Hwang, Eun-Jeong Park, Dong-Soo Ha, and Dong-Woog Choi. Heat shock protein gene family of the Porphyra seriata and enhancement of heat stress tolerance by PsHSP70 in Chlamydomonas. Marine Biotechnology, 14(3):332–342, June 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-011-9417-0.

Palti:2012:SGI

[1052] Yniv Palti, Carine Genet, Guangtu Gao, Yuqin Hu, Frank M. You, Mekki Boussaha, Caird E. Rexroad, and Ming-Cheng Luo. A second generation integrated map of the rainbow trout (Oncorhynchus mykiss) genome: Analysis of conserved synteny with model fish genomes. Marine Biotechnology, 14(3):343–357, June 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-011-9418-z.

Rocha:2012:SBM

[1053] Lenilson C. Rocha, Hercules V. Ferreira, Rodrigo F. Luiz, Lara D. Sette, and André L. M. Porto. Stereoselective bioreduction of 1-(4-methoxyphenyl)ethanone by whole cells of marine-derived fungi. Marine Biotechnology, 14(3):358–362, June 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-011-9419-y.

Quillet:2012:ADA

[1054] Laurent Quillet, Ludovic Besaury, Milka Popova, Sandrine Paissé, Julien Deloffre, and Baghdad Ouddane. Abundance, diversity and activity of sulfate-reducing prokaryotes in heavy metal-contaminated sediment from a salt marsh in the Medway Estuary (UK). Marine Biotechnology, 14 (3):363–381, June 2012. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-011-9420-5.

Kim:2012:ESE

[1055] Kyung-A Kim, Sang Min Kim, Suk Woo Kang, Sang Il Jeon, Byung Hun Um, and Sang Hoon Jung. Edible seaweed, Eisenia bicyclis, protects retinal ganglion cells death caused by oxidative stress. Marine Biotechnology, 14(4):383–395, August 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-012-9459-y.

Rocha:2012:BIM

[1056] Lenilson C. Rocha, Rodrigo F. Luiz, Isac G. Rosset, Cristiano Raminelli, Mirna H. R. Seleghim, Lara Durães Sette, and André L. M. Porto. Bioconversion of iodoacetophenones by marine fungi. Marine Biotechnology, 14(4):396–401, August 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-012-9463-2.

Huvet:2012:VRI

[1057] Arnaud Huvet, Elodie Fleury, Charlotte Corporeau, Virgile Quillien, Jean Yves Daniel, Guillaume Riviere, Pierre Boudry, and Caroline Fabioux. In Vivo RNA interference of a gonad-specific transforming growth factor-β in the Pacific oyster Crassostrea gigas. Marine Biotechnology, 14 (4):402–410, August 2012. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-011-9421-4.

Monroig:2012:ILF

[1058] Óscar Monroig, Juan C. Navarro, James R. Dick, Frederic Alemany, and Douglas R. Tocher. Identification of a Δ5-like fatty acyl desaturase from the cephalopod Octopus vulgaris (Cuvier 1797) involved in the biosynthesis of essential fatty acids. Marine Biotechnology, 14(4): 411–422, August 2012. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-011-9423-2.

Yufera:2012:TCL

[1059] Manuel Yúfera, Silke Halm, Sergi Beltran, Berta Fusté, Josep V. Planas, and Gonzalo Martínez-Rodríguez. Transcriptomic characterization of the larval stage in gilthead seabream (Sparus aurata) by 454 pyrosequencing. Marine Biotechnology, 14(4):423–435, August 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-011-9422-3.

Caicedo:2012:DBE

[1060] Nelson H. Caicedo, Jolanta Kumirska, Jennifer Neumann, Stefan Stolte, and Jorg Thöming. Detection of bioactive exometabolites produced by the filamentous marine cyanobacterium Geitlerinema sp. Marine Biotechnology, 14(4):436–445, August 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-011-9424-1.

Bedouet:2012:PSI

[1061] Laurent Bédouet, Arul Marie, Sophie Berland, Benjamin Marie, Stéphanie Auzoux-Bordenave, Frédéric Marin, and Christian Milet. Proteomic strategy for identifying mollusc shell proteins using mild chemical degradation and trypsin digestion of insoluble organic shell matrix: a pilot study on Haliotis tuberculata. Marine Biotechnology, 14 (4):446–458, August 2012. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-011-9425-0.

Wei:2012:ICP

[1062] Zhenlin Wei, Xiaolin Liu, and Huilin Zhang. Identification and characterization of piRNA-like small RNAs in the gonad of sea urchin (Strongylocentrotus nudus). Marine Biotechnology, 14(4):459–467, August 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-011-9426-z.

Chiu:2012:IGJ

[1063] Ya-Huang Chiu, Yi-Lin Chan, Tsung-Lin Li, and Chang-Jer Wu. Inhibition of Japanese encephalitis virus infection by the sulfated polysaccharide extracts from Ulva lactuca. Marine Biotechnology, 14(4):468– 478, August 2012. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-011-9428-x.

Hino:2012:RCC

[1064] Hirotsugu Hino, Kana Arimoto, Michio Yazawa, Yota Murakami, and Akiko Nakatomi. Ran and calcineurin can participate collaboratively in the regulation of spermatogenesis in scallop. Marine Biotechnology, 14 (4):479–490, August 2012. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-011-9429-9.

Yabu:2012:IAA

[1065] Takeshi Yabu, Shintaro Imamura, Nanami Mizusawa, Ken Touhata, and Michiaki Yamashita. Induction of autophagy by amino acid starvation in fish cells. Marine Biotechnology, 14(4):491–501, August 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-012-9432-9.

Zhu:2012:DPM

[1066] Xia Zhu, Yu Zhen, Tiezhu Mi, and Zhigang Yu. Detection of Prorocentrum minimum (Pavillard) Schiller with an electrochemiluminescencemolecular probe assay. Marine Biotechnology, 14(4):502–511, August 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-012-9431-x.

Sarropoulou:2012:GGA

[1067] Elena Sarropoulou, Jorge Fernandes, and Zhanjiang John Liu. GIA 2011: Genomics in Aquaculture 2011 Symposium. Marine Biotechnology, 14 (5):513–514, October 2012. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-012-9468-x.

Diaz-Rosales:2012:MBI

[1068] P. Díaz-Rosales, A. Romero, P. Balseiro, S. Dios, B. Novoa, and A. Figueras. Microarray-based identification of differentially expressed genes in families of turbot (Scophthalmus maximus) after infection with viral haemorrhagic septicaemia virus (VHSV). Marine Biotechnology, 14 (5):515–529, October 2012. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-012-9465-0.

Wenger:2012:EMH

[1069] Michael Wenger, Aleksei Krasnov, Stanko Skugor, Elinor Goldschmidt-Clermont, Ursula Sattler, Sergey Afanasyev, and Helmut Segner. Estrogen modulates hepatic gene expression and survival of rainbow trout infected with pathogenic bacteria Yersinia ruckeri. Marine Biotechnology, 14(5):530–543, October 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-012-9473-0.

Lanes:2012:NEL

[1070] Carlos Frederico Ceccon Lanes, Sylvie Bolla, Jorge M. O. Fernandes, Ove Nicolaisen, Viswanath Kiron, and Igor Babiak. Nucleotide enrichment of live feed: a promising protocol for rearing of Atlantic cod Gadus morhua larvae. Marine Biotechnology, 14(5):544–558, October 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-012-9458-z.

Micallef:2012:ETG

[1071] Giulia Micallef, Ralph Bickerdike, Caroline Reiff, Jorge M. O. Fernandes, Alan S. Bowman, and Samuel A. M. Martin. Exploring the transcriptome of Atlantic salmon (Salmo salar) skin, a major defense organ. Marine Biotechnology, 14(5):559–569, October 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-012-9447-2.

Pardo:2012:GEP

[1072] Belén G. Pardo, Adrián Millán, Antonio Gómez-Tato, Carlos Fernández, Carmen Bouza, José Antonio Alvarez-Dios, Santiago Cabaleiro, Jesús Lamas, José M. Leiro, and Paulino Martínez. Gene expression profiles of spleen, liver, and head kidney in turbot (Scophthalmus maximus) along the infection process with Philasterides dicentrarchi using an immune-enriched oligo-microarray. Marine Biotechnology, 14(5):

570–582, October 2012. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-012-9440-9.

Minegishi:2012:GET

[1073] Yuki Minegishi, Christiaan V. Henkel, Ron P. Dirks, and Guido E. E. J. M. van den Thillart. Genomics in eels — towards aquaculture and biology. Marine Biotechnology, 14(5):583–590, October 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-012-9444-5.

Piferrer:2012:GAS

[1074] Francesc Piferrer, Laia Ribas, and Noelia Díaz. Genomic approaches to study genetic and environmental influences on fish sex determination and differentiation. Marine Biotechnology, 14(5):591–604, October 2012. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-012-9445-4.

Kaitetzidou:2012:TRS

[1075] Elisavet Kaitetzidou, Diego Crespo, Yoryia Vraskou, Efthimia Antonopoulou, and Josep V. Planas. Transcriptomic response of skeletal muscle to lipopolysaccharide in the gilthead seabream (Sparus aurata). Marine Biotechnology, 14(5):605–619, October 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-012-9469-9.

Robinson:2012:ESP

[1076] Nicholas Robinson, Pramoda K. Sahoo, Matthew Baranski, Kanta Das Mahapatra, Jatindra N. Saha, Sweta Das, Yashowant Mishra, Paramananda Das, Hirak K. Barman, and Ambekar E. Eknath. Expressed sequences and polymorphisms in rohu carp (Labeo rohita, Hamilton) revealed by mRNA-seq. Marine Biotechnology, 14(5):620–633, October 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-012-9433-8.

Sarropoulou:2012:CGE

[1077] E. Sarropoulou, J. Galindo-Villegas, A. García-Alcázar, P. Kasapidis, and V. Mulero. Characterization of European sea bass transcripts by RNA SEQ after oral vaccine against V. anguillarum. Marine Biotechnology, 14(5):634–642, October 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-012-9466-z.

deVareilles:2012:DLI

[1078] Mahaut de Vareilles, Luis E. C. Conceição, Pedro Gómez-Requeni, Katerina Kousoulaki, Nadège Richard, Pedro M. Rodrigues, Kari E. Fladmark, and Ivar Rønnestad. Dietary Lysine imbalance affects muscle proteome in zebrafish (Danio rerio): a comparative 2D-DIGE study. Marine Biotechnology, 14(5):643–654, October 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-012-9462-3.

Vinas:2012:MDS

[1079] Ana Viñas, Xoana Taboada, Luis Vale, Diego Robledo, Miguel Hermida, Manel Vera, and Paulino Martínez. Mapping of DNA sex-specific markers and genes related to sex differentiation in turbot (Scophthalmus maximus). Marine Biotechnology, 14(5):655–663, October 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-012-9451-6.

Cai:2012:ICS

[1080] Menghao Cai, Xiangshan Zhou, Jian Lu, Weimin Fan, Jiushun Zhou, Chuanpeng Niu, Li Kang, Xueqian Sun, and Yuanxing Zhang. An integrated control strategy for the fermentation of the marine-derived fungus Aspergillus glaucus for the production of anti-cancer polyketide. Marine Biotechnology, 14(6):665–671, December 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-012-9435-6.

Hollenbeck:2012:UCG

[1081] Christopher M. Hollenbeck, David S. Portnoy, and John R. Gold. Use of comparative genomics to develop EST–SSRs for red drum (Sciaenops ocellatus). Marine Biotechnology, 14(6):672–680, December 2012. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-012-9449-0.

Walmsley:2012:DBC

[1082] Tara A. Walmsley, Gwynneth F. Matcher, Fan Zhang, Russell T. Hill, Michael T. Davies-Coleman, and Rosemary A. Dorrington. Diversity of bacterial communities associated with the Indian Ocean sponge Tsitsikamma favus that contains the bioactive pyrroloiminoquinones, tsitsikammamine A and B. Marine Biotechnology, 14(6):681–691, December 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-012-9430-y.

Gupta:2012:DEV

[1083] Vishal Gupta, A. J. Bijo, Manoj Kumar, C. R. K. Reddy, and Bhavanath Jha. Detection of epigenetic variations in the protoplast-derived germlings of Ulva reticulata using methylation sensitive amplification polymorphism (MSAP). Marine Biotechnology, 14(6):692–700, December 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-012-9434-7.

Han:2012:BAS

[1084] Minqi Han, Fang Liu, Fengli Zhang, Zhiyong Li, and Houwen Lin. Bacterial and archaeal symbionts in the South China Sea sponge Phakellia fusca: Community structure, relative abundance, and ammoniaoxidizing populations. Marine Biotechnology, 14(6):701–713, December 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-012-9436-5.

Cordeiro:2012:CLP

[1085] Odete D. Cordeiro, Tomé S. Silva, Ricardo N. Alves, Benjamin Costas, Tune Wulff, Nadège Richard, Mahaut de Vareilles, Luís E. C. Conceição, and Pedro M. Rodrigues. Changes in liver proteome expression of Senegalese sole (Solea senegalensis) in response to repeated handling stress. Marine Biotechnology, 14(6):714–729, December 2012. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-012-9437-4.

Zhao:2012:IGP

[1086] Xiaoxia Zhao, Qingheng Wang, Yu Jiao, Ronglian Huang, Yuewen Deng, Huan Wang, and Xiaodong Du. Identification of genes potentially related to biomineralization and immunity by transcriptome analysis of pearl sac in pearl oyster Pinctada martensii. Marine Biotechnology, 14(6): 730–739, December 2012. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-012-9438-3.

Sellars:2012:TIG

[1087] M. J. Sellars, S. M. Arce, and P. L. Hertzler. Triploidy induction in the Pacific white shrimp Litopenaeus vannamei: an assessment of induction agents and parameters, embryo viability, and early larval survival. Marine Biotechnology, 14(6):740–751, December 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-012-9439-2.

Punnarak:2012:RAI

[1088] Porntep Punnarak, Mudjekeewis D. Santos, Seong Don Hwang, Hidehiro Kondo, Ikuo Hirono, Yo Kikuchi, and Takashi Aoki. RNA aptamers inhibit the growth of the fish pathogen viral hemorrhagic septicemia virus (VHSV). Marine Biotechnology, 14(6):752–761, December 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-012-9448-1.

Bruck:2012:BCL

[1089] Wolfram M. Brück, John K. Reed, and Peter J. McCarthy. The bacterial community of the lithistid sponge Discodermia spp. as determined by cultivation and culture-independent methods. Marine Biotechnology, 14(6): 762–773, December 2012. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-012-9443-6.

Taylor:2012:TAE

[1090] Rebecca L. Taylor, Jonathan D. Rand, and Gary S. Caldwell. Treatment with algae extracts promotes flocculation, and enhances growth and neutral lipid content in Nannochloropsis oculata — a candidate for biofuel production. Marine Biotechnology, 14(6):774–781, December 2012. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-012-9441-8.

Jha:2012:GSG

[1091] Bhavanath Jha, Sanjay Lal, Vivekanand Tiwari, Sweta Kumari Yadav, and Pradeep K. Agarwal. The SbASR-1 gene cloned from an extreme halophyte Salicornia brachiata enhances salt tolerance in transgenic tobacco. Marine Biotechnology, 14(6):782–792, December 2012. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-012-9442-7.

Nie:2013:GPC

[1092] Hongtao Nie, Qi Li, Xuelin Zhao, and Lingfeng Kong. Genetic positioning of centromeres through half-tetrad analysis in gynogenetic diploid families of the Zhikong scallop (Chlamys farreri). Marine Biotechnology, 15(1):1–15, February 2013. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-012-9454-3.

Lu:2013:TES

[1093] Xia Lu, Hongxia Wang, Baozhong Liu, and Jianhai Xiang. Three EST– SSR markers associated with QTL for the growth of the clam Meretrix meretrix revealed by selective genotyping. Marine Biotechnology, 15(1): 16–25, February 2013. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-012-9453-4.

Liu:2013:BDG

[1094] Xiao-Yan Liu, Zhe Chi, Guang-Lei Liu, Catherine Madzak, and Zhen-Ming Chi. Both decrease in ACL1 gene expression and increase in ICL1 gene expression in marine-derived yeast Yarrowia lipolytica expressing INU1 gene enhance citric acid production from inulin. Marine Biotechnology, 15(1):26–36, February 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-012-9452-5.

Fuentes-Grunewald:2013:BLP

[1095] C. Fuentes-Grünewald, E. Garcés, E. Alacid, S. Rossi, and J. Camp. Biomass and lipid production of dinoflagellates and raphidophytes in indoor and outdoor photobioreactors. Marine Biotechnology, 15(1): 37–47, February 2013. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-012-9450-7.

Muto:2013:EGT

[1096] Masaki Muto, Yorikane Fukuda, Michiko Nemoto, Tomoko Yoshino, Tadashi Matsunaga, and Tsuyoshi Tanaka. Establishment of a genetic transformation system for the marine pennate diatom Fistulifera sp. strain JPCC DA0580 — a high triglyceride producer. Marine Biotechnology, 15(1):48–55, February 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-012-9457-0.

Ohtani:2013:CAR

[1097] Maki Ohtani, Jun ichi Hikima, Tae Sung Jung, Hidehiro Kondo, Ikuo Hirono, and Takashi Aoki. Construction of an artificially randomized IgNAR phage display library: Screening of variable regions that bind to hen egg white lysozyme. Marine Biotechnology, 15(1):56– 62, February 2013. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-012-9456-1.

Bayer:2013:GMN

[1098] Kristina Bayer, Matthias Scheuermayer, Lars Fieseler, and Ute Hentschel. Genomic mining for novel FADH_{2}-dependent halogenases in marine sponge-associated microbial consortia. Marine Biotechnology, 15 (1):63–72, February 2013. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-012-9455-2.

Bulgakov:2013:MBL

[1099] Aleksandr A. Bulgakov, Marina G. Eliseikina, Svetlana N. Kovalchuk, Irina Yu Petrova, Galina N. Likhatskaya, Ekaterina V. Shamshurina, and Valery A. Rasskazov. Mannan-binding lectin of the sea urchin Strongylocentrotus nudus. Marine Biotechnology, 15(1):73–86, February 2013. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-012-9460-5.

Rojo:2013:CAD

[1100] Liliana Rojo, Fernando García-Carreño, and Maria de los Angeles Navarrete del Toro. Cold-adapted digestive aspartic protease of the clawed lobsters Homarus americanus and Homarus gammarus: Biochemical characterization. Marine Biotechnology, 15(1):87–96, February 2013. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-012-9461-4.

deOliveira:2013:BPH

[1101] Julieta Rangel de Oliveira, Carolina Megumi Mizuno, Mirna Helena Regali Seleghim, Darci Consolação Diniz Javaroti, Maria Olímpia Oliveira Rezende, Maria Diva Landgraf, Lara Durães Sette, and André Luiz Meleiro Porto. Biotransformation of phenylacetonitrile to 2hydroxyphenylacetic acid by marine fungi. Marine Biotechnology, 15 (1):97–103, February 2013. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-012-9464-1.

Teles:2013:ECC

[1102] Mariana Teles, Sebastian Boltaña, Felipe Reyes-López, Maria Ana Santos, Simon Mackenzie, and Lluis Tort. Effects of chronic cortisol administration on global expression of GR and the liver transcriptome in Sparus aurata. Marine Biotechnology, 15(1):104–114, February 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-012-9467-y.

Koyama:2013:ITP

[1103] Takashi Koyama, Hidehiro Kondo, Takashi Aoki, and Ikuo Hirono. Identification of two Penelope-like elements with different structures and chromosome localization in Kuruma shrimp genome. Marine Biotechnology, 15(1):115–123, February 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-012-9474-z.

Alvarez-Torres:2013:ASS

[1104] Daniel Alvarez-Torres, Esther Garcia-Rosado, M. Alejandra Fernandez-Trujillo, Julia Bejar, M. Carmen Alvarez, Juan J. Borrego, and M. Carmen Alonso. Antiviral specificity of the Solea senegalensis Mx protein constitutively expressed in CHSE-214 cells. Marine Biotechnology, 15 (2):125–132, April 2013. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-012-9478-8.

Hamasaki:2013:GDF

[1105] Masaomi Hamasaki, Yutaka Takeuchi, Kadoo Miyaki, and Goro Yoshizaki. Gonadal development and fertility of triploid grass puffer Takifugu niphobles induced by cold shock treatment. Marine Biotechnology, 15(2):133–144, April 2013. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-012-9470-3.

Suzuki:2013:MEP

[1106] Michio Suzuki, Ai Iwashima, Mariko Kimura, Toshihiro Kogure, and Hiromichi Nagasawa. The molecular evolution of the Pif family proteins in various species of mollusks. Marine Biotechnology, 15(2): 145–158, April 2013. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-012-9471-2.

Martin-Gomez:2013:IRC

[1107] Laura Martín-Gómez, Antonio Villalba, Maria Jesús Carballal, and Elvira Abollo. Identification of relevant cancer related-genes in the flat oyster Ostrea edulis affected by disseminated neoplasia. Marine Biotechnology, 15(2):159–174, April 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-012-9472-1.

Shi:2013:CPO

[1108] Yaohua Shi, Chengcheng Yu, Zhifeng Gu, Xin Zhan, Yan Wang, and Aimin Wang. Characterization of the pearl oyster (Pinctada martensii) mantle transcriptome unravels biomineralization genes. Marine Biotechnology, 15(2):175–187, April 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-012-9476-x.

Uji:2013:CSP

[1109] Toshiki Uji, Hiroyuki Mizuta, and Naotsune Saga. Characterization of the sporophyte-preferential gene promoter from the red alga Porphyra yezoensis using transient gene expression. Marine Biotechnology, 15 (2):188–196, April 2013. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-012-9475-y.

Wang:2013:ATL

[1110] Xinxia Wang, Yizhen Wang, and Yongjin Li. Adipose triglyceride lipase (ATGL) clone, expression pattern, and regulation by different lipid sources and lipid levels in large yellow croaker (Pseudosciaena crocea r.). Marine Biotechnology, 15(2):197–205, April 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-012-9477-9.

Gonzalez:2013:IPG

[1111] Enrique Blanco Gonzalez, Masato Aritaki, Shigeru Sakurai, and Nobuhiko Taniguchi. Inference of potential genetic risks associated with large-scale releases of red sea bream in Kanagawa Prefecture, Japan based on nuclear and mitochondrial DNA analysis. Marine Biotechnology, 15(2):206–220, April 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-012-9479-7.

Xu:2013:ICM

[1112] Dongdong Xu, Bao Lou, Hanxiang Xu, Sanlei Li, and Zhi Geng. Isolation and characterization of male-specific DNA markers in the rock bream Oplegnathus fasciatus. Marine Biotechnology, 15(2):221–229, April 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-012-9480-1.

Werner:2013:ADT

[1113] Gijsbert D. A. Werner, Patrick Gemmell, Stefanie Grosser, Rebecca Hamer, and Sebastian M. Shimeld. Analysis of a deep transcriptome from the mantle tissue of Patella vulgata Linnaeus (Mollusca: Gastropoda: Patellidae) reveals candidate biomineralising genes. Marine Biotechnology, 15(2):230–243, April 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-012-9481-0.

Sun:2013:PDT

[1114] Jin Sun, Qian Chen, Janice C. Y. Lun, Jianliang Xu, and Jian-Wen Qiu. PcarnBase: Development of a transcriptomic database for the brain coral Platygyra carnosus. Marine Biotechnology, 15(2):244–251, April 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-012-9482-z.

Huang:2013:GST

[1115] Xian-De Huang, Mi Zhao, Wen-Guang Liu, Yun-Yan Guan, Yu Shi, Qi Wang, Shan-Zeng Wu, and Mao-Xian He. Gigabase-scale transcriptome analysis on four species of pearl oysters. Marine Biotechnology, 15 (3):253–264, June 2013. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-012-9484-x.

Prechoux:2013:CCS

[1116] Aurélie Préchoux, Sabine Genicot, Hélène Rogniaux, and William Helbert. Controlling carrageenan structure using a novel formylglycinedependent sulfatase, an Endo-4S-iota-Carrageenan sulfatase. Marine Biotechnology, 15(3):265–274, June 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-012-9483-y.

Zhang:2013:CLM

[1117] Xiaofeng Zhang, Yan Zhang, Xianhu Zheng, Youyi Kuang, Zixia Zhao, Lan Zhao, Chao Li, Li Jiang, Dingchen Cao, Cuiyun Lu, Peng Xu, and Xiaowen Sun. A consensus linkage map provides insights on genome character and evolution in common carp (Cyprinus carpio L.). Marine Biotechnology, 15(3):275–312, June 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-012-9485-9.

Fernandez:2013:EDS

[1118] Carmen García Fernández, Chrysoula Roufidou, Efthimia Antonopoulou, and Elena Sarropoulou. Expression of developmental-stage-specific genes in the gilthead sea bream Sparus aurata L. Marine Biotechnology, 15 (3):313–320, June 2013. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-012-9486-8.

Liu:2013:GMS

[1119] Hanqin Liu, Bo Guan, Jiang Xu, Changchun Hou, Hua Tian, and Hongxi Chen. Genetic manipulation of sex ratio for the large-scale breeding of YY super-male and XY all-male yellow catfish (Pelteobagrus fulvidraco (Richardson)). Marine Biotechnology, 15(3):321–328, June 2013. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-012-9487-7.

Moran:2013:ASP

[1120] Yehu Moran, Daniela Praher, Ami Schlesinger, Ari Ayalon, Yossi Tal, and Ulrich Technau. Analysis of soluble protein contents from the nematocysts of a model sea anemone sheds light on venom evolution. Marine Biotechnology, 15(3):329–339, June 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-012-9491-y.

Kurokawa:2013:BCF

[1121] Satoru Kurokawa, Jun Kabayama, Tsuguaki Fukuyasu, Seong Don Hwang, Chan-Il Park, Seong-Bin Park, Carmelo S. del Castillo, Jun ichi Hikima, Tae-Sung Jung, Hidehiro Kondo, Ikuo Hirono, Haruko Takeyama, and Takashi Aoki. Bacterial classification of fish-pathogenic Mycobacterium species by multigene phylogenetic analyses and MALDI biotyper identification system. Marine Biotechnology, 15(3):340–348, June 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-012-9492-x.

Matsushima:2013:AEA

[1122] Ryoji Matsushima, Ryuichi Watanabe, Masataka Tsuda, and Toshiyuki Suzuki. Analysis of extracellular alginate lyase (alyA) expression and its regulatory region in a marine bacterial strain, Pseudoalteromonas atlantica AR06 , using a gfp gene reporter system. Marine Biotechnology, 15(3):349–356, June 2013. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-012-9488-6.

Mussino:2013:PCN

[1123] Francesca Mussino, Marina Pozzolini, Laura Valisano, Carlo Cerrano, Umberto Benatti, and Marco Giovine. Primmorphs cryopreservation: a new method for long-time storage of sponge cells. Marine Biotechnology, 15(3):357–367, June 2013. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-012-9490-z.

Wang:2013:CIB

[1124] Jiajia Wang, Kai Yang, and Xiaobo Zhang. Characterization of the interaction between arginine kinase and siRNA. Marine Biotechnology, 15(3):368–374, June 2013. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-012-9489-5.

Muller:2013:PBP

[1125] Werner E. G. Müller, Xiaohong Wang, Peter Proksch, Carole C. Perry, Ronald Osinga, Johan Gardères, and Heinz C. Schröder. Principles of biofouling protection in marine sponges: a model for the design of novel biomimetic and bio-inspired coatings in the marine environment? Marine Biotechnology, 15(4):375–398, August 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-013-9497-0.

Andriantahina:2013:CSG

[1126] Farafidy Andriantahina, Xiaolin Liu, Tingting Feng, and Jianhai Xiang. Current status of genetics and genomics of reared penaeid shrimp: Information relevant to access and benefit sharing. Marine Biotechnology, 15 (4):399–412, August 2013. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-013-9500-9.

Vicente:2013:BAA

[1127] Jan Vicente, Allison Stewart, Bongkeun Song, Russell T. Hill, and Jeffrey L. Wright. Biodiversity of actinomycetes associated with Caribbean sponges and their potential for natural product discovery. Marine Biotechnology, 15(4):413–424, August 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-013-9493-4.

Lin:2013:ICE

[1128] Hung-Yun Lin, Chi-Yu Shih, Hung-Chun Liu, Jeng Chang, Ying-Lan Chen, Yet-Ran Chen, Han-Tso Lin, Yu-Yung Chang, Chun-Hua Hsu, and Han-Jia Lin. Identification and characterization of an extracellular alkaline phosphatase in the marine diatom Phaeodactylum tricornutum. Marine Biotechnology, 15(4):425–436, August 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-013-9494-3.

Shabala:2013:TCG

[1129] Lana Shabala, Tom McMeekin, and Sergey Shabala. Thraustochytrids can be grown in low-salt media without affecting PUFA production. Marine Biotechnology, 15(4):437–444, August 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-013-9499-y.

Rebl:2013:TPG

[1130] Alexander Rebl, Marieke Verleih, Judith M. Köbis, Carsten Kühn, Klaus Wimmers, Bernd Köllner, and Tom Goldammer. Transcriptome profiling of gill tissue in regionally bred and globally farmed rainbow trout strains reveals different strategies for coping with thermal stress. Marine Biotechnology, 15(4):445–460, August 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-013-9501-8.

Koyama:2013:ADL

[1131] Sumihiro Koyama, Masa aki Konishi, Yukari Ohta, Tetsuya Miwa, Yuji Hatada, Takashi Toyofuku, Tadashi Maruyama, Yuichi Nogi, Chiaki Kato, and Taishi Tsubouchi. Attachment and detachment of living microorganisms using a potential-controlled electrode. Marine Biotechnology, 15(4):461–475, August 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-013-9495-2.

Ohara:2013:TFA

[1132] Junichiro Ohara, Keishi Sakaguchi, Yuji Okita, Nozomu Okino, and Makoto Ito. Two fatty acid elongases possessing C18-Δ6/C18-Δ9/C20-Δ5 or C16-Δ9 elongase activity in Thraustochytrium sp. ATCC 26185. Marine Biotechnology, 15(4):476–486, August 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-013-9496-1.

Nagai:2013:IPP

[1133] Kouhei Nagai, Koichi Morimoto, Haruka Ikegami, Hajime Kimura, and Norishige Yotsukura. Investigation of proteomic profiles of lamina of Ecklonia kurome (Laminariales): Homology-based cross-species protein identification and analysis of the post-translational processing of vanadiumdependent bromoperoxidases using MALDI-TOF/TOF. Marine Biotechnology, 15(4):487–498, August 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-013-9498-z.

Lee:2013:MDA

[1134] Yoon Mi Lee, Min Jeong Kim, Huayue Li, Ping Zhang, Baoquan Bao, Ka Jeong Lee, and Jee H. Jung. Marine-derived Aspergillus species as a source of bioactive secondary metabolites. Marine Biotechnology, 15 (5):499–519, October 2013. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-013-9506-3.

Aihara:2013:OMC

[1135] Hitoshi Aihara, Lavanya Katikala, Robert W. Zeller, Anna Di Gregorio, and Yutaka Nibu. Optimization of a method for chromatin immunoprecipitation assays in the marine invertebrate chordate Ciona. Marine Biotechnology, 15(5):520–525, October 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-013-9504-5.

Xiong:2013:TEZ

[1136] Feng Xiong, Zhi-Qiang Wei, Zuo-Yan Zhu, and Yong-Hua Sun. Targeted expression in zebrafish primordial germ cells by Cre/loxP and Gal4/UAS systems. Marine Biotechnology, 15(5):526–539, October 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-013-9505-4.

Yang:2013:PDD

[1137] Shan Yang, Wei Sun, Fengli Zhang, and Zhiyong Li. Phylogenetically diverse denitrifying and ammonia-oxidizing bacteria in corals Alcyonium gracillimum and Tubastraea coccinea. Marine Biotechnology, 15 (5):540–551, October 2013. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-013-9503-6.

Li:2013:AAS

[1138] Yong-Xin Li, Hui-Xian Wu, Ying Xu, Chang-Lun Shao, Chang-Yun Wang, and Pei-Yuan Qian. Antifouling activity of secondary metabolites isolated from Chinese marine organisms. Marine Biotechnology, 15 (5):552–558, October 2013. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-013-9502-7.

Yamashita:2013:SNS

[1139] Michiaki Yamashita, Yumiko Yamashita, Tamami Suzuki, Yoko Kani, Nanami Mizusawa, Shintaro Imamura, Kenji Takemoto, Tatsuro Hara, Md. Anwar Hossain, Takeshi Yabu, and Ken Touhata. Selenoneine, a novel selenium-containing compound, mediates detoxification mechanisms against methylmercury accumulation and toxicity in zebrafish embryo. Marine Biotechnology, 15(5):559–570, October 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-013-9508-1.

Ruiz:2013:MNP

[1140] Cesar Ruiz, Katherine Valderrama, Sven Zea, and Leonardo Castellanos. Mariculture and natural production of the antitumoural (+)- discodermolide by the Caribbean marine sponge Discodermia dissoluta. Marine Biotechnology, 15(5):571–583, October 2013. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-013-9510-7.

Liu:2013:DNA

[1141] Betty Revon Liu, Ji-Sing Liou, Yung-Jen Chen, Yue-Wern Huang, and Han-Jung Lee. Delivery of nucleic acids, proteins, and nanoparticles by arginine-rich cell-penetrating peptides in rotifers. Marine Biotechnology, 15(5):584–595, October 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-013-9509-0.

Kurokawa:2013:CGA

[1142] Satoru Kurokawa, Jun Kabayama, Seong Don Hwang, Seong-Won Nho, Jun ichi Hikima, Tae-Sung Jung, Masahiro Sakai, Hidehiro Kondo, Ikuo Hirono, and Takashi Aoki. Comparative genome analysis of fish and human isolates of Mycobacterium marinum. Marine Biotechnology, 15 (5):596–605, October 2013. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-013-9511-6.

Imbs:2013:IFH

[1143] Tatiana I. Imbs, Svetlana P. Ermakova, Sergey A. Fedoreyev, Stanislav D. Anastyuk, and Tatiana N. Zvyagintseva. Isolation of fucoxanthin and highly unsaturated monogalactosyldiacylglycerol from brown alga Fucus evanescens C agardh and In Vitro investigation of their antitumor activity. Marine Biotechnology, 15(5):606–612, October 2013. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-013-9507-2.

Rexroad:2013:QTL

[1144] Caird E. Rexroad, Roger L. Vallejo, Sixin Liu, Yniv Palti, and Gregory M. Weber. Quantitative trait loci affecting response to crowding stress in an F_{2} generation of rainbow trout produced through phenotypic selection. Marine Biotechnology, 15(5):613–627, October 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-013-9512-5.

Ibarz:2013:PAS

[1145] Antoni Ibarz, Patricia I. S. Pinto, and Deborah M. Power. Proteomic approach to skin regeneration in a marine teleost: Modulation by Oestradiol-17β. Marine Biotechnology, 15(6):629–646, December 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-013-9513-4.

Jones:2013:GWS

[1146] David B. Jones, Dean R. Jerry, Sylvain Forêt, Dmitry A. Konovalov, and Kyall R. Zenger. Genome-wide SNP validation and mantle tissue transcriptome analysis in the silver-lipped pearl oyster, Pinctada maxima. Marine Biotechnology, 15(6):647–658, December 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-013-9514-3.

Song:2013:ICT

[1147] Chongfu Song, Liangquan Sheng, and Xiaobo Zhang. Immobilization and characterization of a thermostable lipase. Marine Biotechnology, 15(6): 659–667, December 2013. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-013-9515-2.

Santos-Gandelman:2013:CCB

[1148] Juliana F. Santos-Gandelman, Olinda C. S. Santos, Paula V. M. Pontes, Cleyton Lage Andrade, Elisa Korenblum, Guilherme Muricy, Marcia Giambiagi-deMarval, and Marinella S. Laport. Characterization of cultivable bacteria from Brazilian sponges. Marine Biotechnology, 15(6): 668–676, December 2013. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-013-9518-z.

Lanes:2013:TGA

[1149] Carlos Frederico Ceccon Lanes, Teshome Tilahun Bizuayehu, Jorge Manuel de Oliveira Fernandes, Viswanath Kiron, and Igor Babiak. Transcriptome of Atlantic cod (Gadus morhua L.) early embryos from farmed and wild broodstocks. Marine Biotechnology, 15(6):677–694, December 2013. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-013-9527-y.

Novak:2013:CHD

[1150] Halina R. Novak, Christopher Sayer, Jana Panning, and Jennifer A. Littlechild. Characterisation of an l-haloacid dehalogenase from the marine psychrophile Psychromonas ingrahamii with potential industrial application. Marine Biotechnology, 15(6):695–705, December 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-013-9522-3.

Shi:2013:CZS

[1151] Mingjun Shi, Ya Lin, Guangrui Xu, Liping Xie, Xiaoli Hu, Zhenmin Bao, and Rongqing Zhang. Characterization of the Zhikong scallop (Chlamys farreri) mantle transcriptome and identification of biomineralization-related genes. Marine Biotechnology, 15(6):706–715, December 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-

2236 (electronic). URL https://link.springer.com/article/10.
1007/s10126-013-9517-0.

Xiao:2013:IJE

[1152] Yamei Xiao, Yonghua Zhou, Zhen Xiong, Lijun Zou, Minggui Jiang, Zhongwen Luo, Sheng Wen, Wenbin Liu, Shaojun Liu, and Wancheng Li. Involvement of JNK in the embryonic development and organogenesis in zebrafish. Marine Biotechnology, 15(6):716–725, December 2013. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-013-9520-5.

Prol-Garcia:2013:EPP

[1153] María J. Prol-García and José Pintado. Effectiveness of probiotic Phaeobacter bacteria grown in biofilters against Vibrio anguillarum infections in the rearing of turbot (Psetta maxima) larvae. Marine Biotechnology, 15(6):726–738, December 2013. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-013-9521-4.

Riviere:2013:DMC

[1154] Guillaume Riviere, Guan-Chung Wu, Alexandre Fellous, Didier Goux, Pascal Sourdaine, and Pascal Favrel. DNA methylation is crucial for the early development in the oyster C . gigas. Marine Biotechnology, 15(6): 739–753, December 2013. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-013-9523-2.

Xia:2014:MQT

[1155] Jun Hong Xia, Grace Lin, Xiaoping He, Bu Yunping, Peng Liu, Feng Liu, Fei Sun, Rongjian Tu, and Gen Hua Yue. Mapping quantitative trait loci for omega-3 fatty acids in Asian seabass. Marine Biotechnology, 16(1):1–9, February 2014. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-013-9524-1.

Fukushima:2014:XMT

[1156] Ei Fukushima, Toshiharu Iwai, Chiemi Miura, Fritzie T. Celino, Shintarou Urasaki, and Takeshi Miura. A xenograft mantle transplantation technique for producing a novel pearl in an akoya oyster host. Marine Biotechnology, 16(1):10–16, February 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-013-9525-0.

Zhang:2014:TAR

[1157] Linlin Zhang, Li Li, Yabing Zhu, Guofan Zhang, and Ximing Guo. Transcriptome analysis reveals a rich gene set related to innate immunity in the Eastern oyster (Crassostrea virginica). Marine Biotechnology, 16 (1):17–33, February 2014. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-013-9526-z.

Freer:2014:BPM

[1158] Andy Freer, Stephen Bridgett, Jiahong Jiang, and Maggie Cusack. Biomineral proteins from Mytilus edulis mantle tissue transcriptome. Marine Biotechnology, 16(1):34–45, February 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-013-9516-1.

Qiu:2014:MCH

[1159] Chuwen Qiu, Jie Sun, Mei Liu, Baojie Wang, Keyong Jiang, Shujuan Sun, Xiaolin Meng, Zuoyong Luo, and Lei Wang. Molecular cloning of hemocyanin cDNA from Fenneropenaeus chinensis and antimicrobial analysis of two C-terminal fragments. Marine Biotechnology, 16 (1):46–53, February 2014. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-013-9519-y.

Jiao:2014:ICM

[1160] Yu Jiao, Zhe Zheng, Xiaodong Du, Qingheng Wang, Ronglian Huang, Yuewen Deng, Shangli Shi, and Xiaoxia Zhao. Identification and characterization of MicroRNAs in pearl oyster Pinctada martensii by solexa deep sequencing. Marine Biotechnology, 16(1):54–62, February 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-013-9528-x.

Ahanger:2014:PSP

[1161] Sajad Ahanger, Supriyanka Sandaka, Deepika Ananad, Madhu K. Mani, Ravinder Kondadhasula, Chandra Sekhar Reddy, Makesh Marappan, Rajendran K. Valappil, Kshitish C. Majumdar, and Rakesh K. Mishra. Protection of shrimp Penaeus monodon from WSSV infection using antisense constructs. Marine Biotechnology, 16(1):63–73, February 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-

2236 (electronic). URL https://link.springer.com/article/10.
1007/s10126-013-9529-9.

Kumari:2014:QSE

[1162] Puja Kumari, Radhakrishnan Reddy, and Bhavanath Jha. Quantification of selected endogenous hydroxy-oxylipins from tropical marine macroalgae. Marine Biotechnology, 16(1):74–87, February 2014. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-013-9533-0.

Limtipsuntorn:2014:MAH

[1163] Ubonrat Limtipsuntorn, Yutaka Haga, Hidehiro Kondo, Ikuo Hirono, and Shuichi Satoh. Microarray analysis of hepatic gene expression in juvenile Japanese flounder Paralichthys olivaceus fed diets supplemented with fish or vegetable oils. Marine Biotechnology, 16(1):88– 102, February 2014. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-013-9535-y.

Gutierrez:2014:DQT

[1164] Alejandro P. Gutierrez, Krzysztof P. Lubieniecki, Steve Fukui, Ruth E. Withler, Bruce Swift, and William S. Davidson. Detection of quantitative trait loci (QTL) related to grilsing and late sexual maturation in Atlantic salmon (Salmo salar). Marine Biotechnology, 16(1): 103–110, February 2014. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-013-9530-3.

Jiang:2014:CPE

[1165] Hao Jiang, Fuhua Li, Jiquan Zhang, Jinkang Zhang, Bingxin Huang, Yang Yu, and Jianhai Xiang. Comparison of protein expression profiles of the hepatopancreas in Fenneropenaeus chinensis challenged with heat-inactivated Vibrio anguillarum and white spot syndrome virus. Marine Biotechnology, 16(1):111–123, February 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-013-9538-8.

Zhang:2014:PED

[1166] Xi Zhang, Guijun Guan, Jianbin Chen, Kiyoshi Naruse, and Yunhan Hong. Parameters and efficiency of direct gene disruption by zinc finger nucleases in medaka embryos. Marine Biotechnology, 16 (2):125–134, April 2014. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-013-9556-6.

Deng:2014:IEC

[1167] Yunyan Deng, Jianting Yao, Gang Fu, Hui Guo, and Delin Duan. Isolation, expression, and characterization of blue light receptor AU-REOCHROME gene from Saccharina japonica (Laminariales, Phaeophyceae). Marine Biotechnology, 16(2):135–143, April 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-013-9539-7.

Gantayet:2014:NPI

[1168] Arpita Gantayet, David J. Rees, and Eli D. Sone. Novel proteins identified in the insoluble byssal matrix of the freshwater zebra mussel. Marine Biotechnology, 16(2):144–155, April 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-013-9537-9.

deOliveira:2014:BMG

[1169] Julieta Rangel de Oliveira, Mirna Helena Regali Seleghim, and André Luiz Meleiro Porto. Biotransformation of methylphenylacetonitriles by Brazilian marine fungal strain Aspergillus sydowii CBMAI 934: Ecofriendly reactions. Marine Biotechnology, 16(2):156–160, April 2014. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-013-9534-z.

Jung:2014:CGA

[1170] Hyungtaek Jung, Russell E. Lyons, Yutao Li, Nguyen Minh Thanh, Hung Dinh, David A. Hurwood, Krishna R. Salin, and Peter B. Mather. A candidate gene association study for growth performance in an improved giant freshwater prawn (Macrobrachium rosenbergii) culture line. Marine Biotechnology, 16(2):161–180, April 2014. CODEN MABIFW.

ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-013-9555-7.

Maeda:2014:ILA

[1171] M. Maeda, A. Shibata, G. Biswas, H. Korenaga, T. Kono, T. Itami, and M. Sakai. Isolation of lactic acid bacteria from kuruma shrimp (Marsupenaeus japonicus) intestine and assessment of immunomodulatory role of a selected strain as probiotic. Marine Biotechnology, 16 (2):181–192, April 2014. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-013-9532-1.

Ramos:2014:SAB

[1172] Maribel García Ramos and Anastazia T. Banaszak. Symbiotic association between Symbiodinium and the gastropod Strombus gigas: Larval acquisition of symbionts. Marine Biotechnology, 16(2):193–201, April 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-013-9536-x.

Kumar:2014:BSB

[1173] P. K. Anil Kumar and P. V. Suresh. Biodegradation of shrimp biowaste by marine Exiguobacterium sp. CFR26M and concomitant production of extracellular protease and antioxidant materials: Production and process optimization by response surface methodology. Marine Biotechnology, 16 (2):202–218, April 2014. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-013-9531-2.

He:2014:MLM

[1174] Shan He, Jiejie Hao, Weibing Peng, Peiju Qiu, Chunxia Li, and Huashi Guan. Modulation of lipid metabolism by deep-sea water in cultured human liver (HepG2) cells. Marine Biotechnology, 16(2): 219–229, April 2014. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-013-9540-1.

Besaury:2014:AAD

[1175] Ludovic Besaury, Jean-François Ghiglione, and Laurent Quillet. Abundance, activity, and diversity of archaeal and bacterial communities in both uncontaminated and highly copper-contaminated marine sediments. Marine Biotechnology, 16(2):230–242, April 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-013-9542-z.

Yano:2014:HTI

[1176] Ayaka Yano, Barbara Nicol, Elodie Jouanno, and Yann Guiguen. Heritable targeted inactivation of the rainbow trout (Oncorhynchus mykiss) master sex-determining gene using zinc-finger nucleases. Marine Biotechnology, 16(2):243–250, April 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-013-9546-8.

Uji:2014:COB

[1177] Toshiki Uji, Ryo Hirata, Satoru Fukuda, Hiroyuki Mizuta, and Naotsune Saga. A codon-optimized bacterial antibiotic gene used as selection marker for stable nuclear transformation in the marine red alga Pyropia yezoensis. Marine Biotechnology, 16(3):251–255, June 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-013-9549-5.

Skugor:2014:CMG

[1178] Adrijana Škugor, Krasimir Slanchev, Jacob Seilø Torgersen, Helge Tveiten, and Øivind Andersen. Conserved mechanisms for germ cellspecific localization of nanos3 transcripts in teleost species with aquaculture significance. Marine Biotechnology, 16(3):256–264, June 2014. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-013-9543-y.

Rodriguez-Ramilo:2014:IQT

[1179] Silvia T. Rodríguez-Ramilo, Roberto De La Herrán, Carmelo Ruiz-Rejón, Miguel Hermida, Carlos Fernández, Patricia Pereiro, Antonio Figueras, Carmen Bouza, Miguel A. Toro, Paulino Martínez, and Jesús Fernández. Identification of quantitative trait loci associated with resistance to viral haemorrhagic septicaemia (VHS) in turbot (Scophthalmus maximus): a comparison between bacterium, parasite and virus diseases. Marine Biotechnology, 16(3):265–276, June 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-013-9544-x.

Mazurais:2014:IHR

[1180] David Mazurais, Serena Ferraresso, Pier Paolo Gatta, Elisabeth Desbruyères, Armelle Severe, Charlotte Corporeau, Guy Claireaux, Luca Bargelloni, and Jose-Luis Zambonino-Infante. Identification of hypoxiaregulated genes in the liver of common sole (Solea solea) fed different dietary lipid contents. Marine Biotechnology, 16(3):277–288, June 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-013-9545-9.

Allen:2014:MMV

[1181] M. S. Allen, M. M. Ferguson, and R. G. Danzmann. Molecular markers for variation in spawning date in a hatchery population of rainbow trout (Oncorhynchus mykiss). Marine Biotechnology, 16(3):289– 298, June 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-

2236 (electronic). URL https://link.springer.com/article/10.
1007/s10126-013-9547-7.

Chiou:2014:PHT

[1182] Pinwen Peter Chiou, Maria J. Chen, Chun-Mean Lin, Jenny Khoo, Jon Larson, Rich Holt, Jo-Ann Leong, Gary Thorgarrd, and Thomas T. Chen. Production of homozygous transgenic rainbow trout with enhanced disease resistance. Marine Biotechnology, 16(3):299–308, June 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-013-9550-z.

Farlora:2014:IGC

[1183] Rodolfo Farlora, Shoko Hattori-Ihara, Yukata Takeuchi, Makoto Hayashi, Anna Octavera, Alimuddin, and Goro Yoshizaki. Intraperitoneal germ cell transplantation in the Nile tilapia Oreochromis niloticus. Marine Biotechnology, 16(3):309–320, June 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-013-9551-y.

Singh:2014:EPA

[1184] Natwar Singh, Avinash Mishra, and Bhavanath Jha. Over-expression of the peroxisomal ascorbate peroxidase (SbpAPX ) gene cloned from halophyte Salicornia brachiata confers salt and drought stress tolerance in transgenic tobacco. Marine Biotechnology, 16(3):321–332, June 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-013-9548-6.

Rebl:2014:TPR

[1185] Alexander Rebl, Tomáš Korytář, Judith M. Köbis, Marieke Verleih, Aleksei Krasnov, Joanna Jaros, Carsten Kühn, Bernd Köllner, and Tom Goldammer. Transcriptome profiling reveals insight into distinct immune responses to Aeromonas salmonicida in gill of two rainbow trout strains. Marine Biotechnology, 16(3):333–348, June 2014. CODEN MABIFW.

ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-013-9552-x.

Vallejo:2014:DQR

[1186] Roger L. Vallejo, Yniv Palti, Sixin Liu, Jason P. Evenhuis, Guangtu Gao, Caird E. Rexroad, and Gregory D. Wiens. Detection of QTL in rainbow trout affecting survival when challenged with Flavobacterium psychrophilum. Marine Biotechnology, 16(3):349–360, June 2014. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-013-9553-9.

Hwang:2014:PEF

[1187] Eunson Hwang, Sang-Yong Park, Zheng wang Sun, Heon-Sub Shin, Don-Gil Lee, and Tae Hoo Yi. The protective effects of fucosterol against skin damage in UVB-irradiated human dermal fibroblasts. Marine Biotechnology, 16(3):361–370, June 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-013-9554-8.

Mayfield:2014:RED

[1188] Anderson B. Mayfield, Yi-Yuong Hsiao, Hung-Kai Chen, and Chii-Shiarng Chen. Rubisco expression in the dinoflagellate Symbiodinium sp. is influenced by both photoperiod and endosymbiotic lifestyle. Marine Biotechnology, 16(4):371–384, August 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-014-9558-z.

Brack:2014:DPB

[1189] Christiane Brack, Annett Mikolasch, and Frieder Schauer. 2,5diketopiperazines produced by Bacillus pumilus during bacteriolysis of Arthrobacter citreus. Marine Biotechnology, 16(4):385–395, August 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-014-9559-y.

Chang:2014:CTA

[1190] Kim Jye Lee Chang, Carol Mancuso Nichols, Susan I. Blackburn, Graeme A. Dunstan, Anthony Koutoulis, and Peter D. Nichols. Comparison of thraustochytrids Aurantiochytrium sp., Schizochytrium sp., Thraustochytrium sp., and Ulkenia sp. for production of biodiesel, longchain omega-3 oils, and exopolysaccharide. Marine Biotechnology, 16 (4):396–411, August 2014. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-014-9560-5.

Kasanah:2014:MRF

[1191] Noer Kasanah, Lorelei Lucas Farr, Abbas Gholipour, David E. Wedge, and Mark T. Hamann. Metabolism and resistance of Fusarium spp. to the manzamine alkaloids via a putative retro Pictet–Spengler reaction and utility of the rational design of antimalarial and antifungal agents. Marine Biotechnology, 16(4):412–422, August 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-014-9557-0.

Calduch-Giner:2014:TAM

[1192] Josep A. Calduch-Giner, Yann Echasseriau, Diego Crespo, Daniel Baron, Josep V. Planas, Patrick Prunet, and Jaume Pérez-Sánchez. Transcriptional assessment by microarray analysis and large-scale meta-analysis of the metabolic capacity of cardiac and skeletal muscle tissues to cope with reduced nutrient availability in gilthead sea bream (Sparus aurata L.). Marine Biotechnology, 16(4):423–435, August 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-014-9562-3.

Chiang:2014:DVC

[1193] Cheng-Yi Chiang, Yi-Lin Chen, and Huai-Jen Tsai. Different visible colors and green fluorescence were obtained from the mutated purple chromoprotein isolated from sea anemone. Marine Biotechnology, 16(4): 436–446, August 2014. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-014-9563-2.

Lima:2014:DVM

[1194] Paula C. Lima, Natasha A. Botwright, James O. Harris, and Mathew Cook. Development of an In Vitro model system for studying bacterially expressed dsRNA-mediated knockdown in Neoparamoeba genus. Marine Biotechnology, 16(4):447–455, August 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-014-9561-4.

Magnusson:2014:BLF

[1195] Marie Magnusson, Leonardo Mata, Rocky de Nys, and Nicholas A. Paul. Biomass, lipid and fatty acid production in large-scale cultures of the marine macroalga Derbesia tenuissima (Chlorophyta). Marine Biotechnology, 16(4):456–464, August 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-014-9564-1.

Yasumoto:2014:BPC

[1196] Ko Yasumoto, Mina Yasumoto-Hirose, Jun Yasumoto, Ryo Murata, Shun ichi Sato, Megumi Baba, Kanami Mori-Yasumoto, Mitsuru Jimbo, Yasukatsu Oshima, Takenori Kusumi, and Shugo Watabe. Biogenic polyamines capture CO_{2} and accelerate extracellular bacterial CaCO_{3} formation. Marine Biotechnology, 16(4):465–474, August 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-014-9566-z.

Zhang:2014:MCB

[1197] Xuan Zhang, Huiyue Zhou, Xiaonan Zang, Le Gong, Hengyi Sun, and Xuecheng Zhang. MIPS: a calmodulin-binding protein of Gracilaria lemaneiformis under heat shock. Marine Biotechnology, 16(4):475–483, August 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-014-9565-0.

Konishi:2014:DSR

[1198] Masaaki Konishi, Shinro Nishi, Tokuma Fukuoka, Dai Kitamoto, Tomo o Watsuji, Yuriko Nagano, Akinori Yabuki, Satoshi Nakagawa, Yuji Hatada, and Jun ichi Horiuchi. Deep-sea Rhodococcus sp. BS-15, lacking the phytopathogenic fas genes, produces a novel glucotriose lipid biosurfactant. Marine Biotechnology, 16(4):484–493, August 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-014-9568-x.

Solovchenko:2014:DSN

[1199] Alexei Solovchenko, Olga Gorelova, Irina Selyakh, Larisa Semenova, Olga Chivkunova, Olga Baulina, and Elena Lobakova. Desmodesmus sp. 3Dp86E-1 — a novel symbiotic chlorophyte capable of growth on pure CO_{2}. Marine Biotechnology, 16(5):495–501, October 2014. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-014-9572-1.

Jung:2014:ICA

[1200] Woongsic Jung, Yunho Gwak, Peter L. Davies, Hak Jun Kim, and Eon-Seon Jin. Isolation and characterization of antifreeze proteins from the Antarctic marine microalga Pyramimonas gelidicola. Marine Biotechnology, 16(5):502–512, October 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-014-9567-y.

Wang:2014:GSQ

[1201] Lei Wang, Caixia Fan, Yang Liu, Yingping Zhang, Shoutang Liu, Deqiang Sun, Han Deng, Ying Xu, Yongsheng Tian, Xiaolin Liao, Mingshu Xie, Wenlong Li, and Songlin Chen. A genome scan for quantitative trait loci associated with Vibrio anguillarum infection resistance in Japanese flounder (Paralichthys olivaceus) by bulked segregant analysis. Marine Biotechnology, 16(5):513–521, October 2014. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-014-9569-9.

Richard:2014:DSV

[1202] Nadège Richard, Ignacio Fernández, Tune Wulff, Kristin Hamre, Leonor Cancela, Luis E. C. Conceição, and Paulo J. Gavaia. Dietary supplementation with vitamin K affects transcriptome and proteome of Senegalese sole, improving larval performance and quality. Marine Biotechnology, 16 (5):522–537, October 2014. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-014-9571-2.

Xie:2014:CNC

[1203] Wei-Hong Xie, Cong-Cong Zhu, Nai-Sheng Zhang, Da-Wei Li, Wei-Dong Yang, Jie-Sheng Liu, Ramalingam Sathishkumar, and Hong-Ye Li. Construction of novel chloroplast expression vector and development of an efficient transformation system for the diatom Phaeodactylum tricornutum. Marine Biotechnology, 16(5):538–546, October 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-014-9570-3.

Wang:2014:NTS

[1204] Wei Wang, Jerome H. L. Hui, Ting Fung Chan, and Ka Hou Chu. De Novo transcriptome sequencing of the snail Echinolittorina malaccana: Identification of genes responsive to thermal stress and development of genetic markers for population studies. Marine Biotechnology, 16(5): 547–559, October 2014. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-014-9573-0.

Pasaribu:2014:SNP

[1205] Buntora Pasaribu, I-Ping Lin, Jason T. C. Tzen, Guang-Yuh Jauh, Tung-Yung Fan, Yu-Min Ju, Jing-O Cheng, Chii-Shiarng Chen, and Pei-Luen Jiang. SLDP: a novel protein related to caleosin is associated with the endosymbiotic Symbiodinium lipid droplets from Euphyllia glabrescens.

Marine Biotechnology, 16(5):560–571, October 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-014-9574-z.

Kurokawa:2014:WGA

[1206] Satoru Kurokawa, Jun Kabayama, Seong Don Hwang, Seong Won Nho, Jun ichi Hikima, Tae Sung Jung, Hidehiro Kondo, Ikuo Hirono, Haruko Takeyama, Tetsushi Mori, and Takashi Aoki. Whole genome analyses of marine fish pathogenic isolate, Mycobacterium sp. 012931. Marine Biotechnology, 16(5):572–579, October 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-014-9576-x.

Pang:2014:DTH

[1207] Shao-Chen Pang, Hou-Peng Wang, Kuo-Yu Li, Zuo-Yan Zhu, Jing X. Kang, and Yong-Hua Sun. Double transgenesis of humanized fat1 and fat2 genes promotes omega-3 polyunsaturated fatty acids synthesis in a zebrafish model. Marine Biotechnology, 16(5):580–593, October 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-014-9577-9.

Steinert:2014:ADG

[1208] Georg Steinert, Susanna Whitfield, Michael W. Taylor, Carsten Thoms, and Peter J. Schupp. Application of diffusion growth chambers for the cultivation of marine sponge-associated bacteria. Marine Biotechnology, 16(5):594–603, October 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-014-9575-y.

Jiang:2014:GWA

[1209] Qun Jiang, Qi Li, Hong Yu, and Lingfeng Kong. Genome-wide analysis of simple sequence repeats in marine animals — a comparative approach. Marine Biotechnology, 16(5):604–619, October 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-014-9580-1.

Benstein:2014:IGP

[1210] Ruben Maximilian Benstein, Zehra Çebi, Björn Podola, and Michael Melkonian. Immobilized growth of the peridinin-producing marine dinoflagellate Symbiodinium in a simple biofilm photobioreactor. Marine Biotechnology, 16(6):621–628, December 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-014-9581-0.

Laghari:2014:MQS

[1211] Muhammad Younis Laghari, Punhal Lashari, Xiaofeng Zhang, Peng Xu, Naeem Tariq Narejo, Yongxin Liu, Shahid Mehboob, K. Al-Ghanim, Yan Zhang, and Xiaowen Sun. Mapping QTLs for swimming ability related traits in Cyprinus carpio L. Marine Biotechnology, 16(6):629– 637, December 2014. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-014-9578-8.

Hecht:2014:DEC

[1212] Benjamin C. Hecht, Madeline E. Valle, Frank P. Thrower, and Krista M. Nichols. Divergence in expression of candidate genes for the smoltification process between juvenile resident rainbow and anadromous steelhead trout. Marine Biotechnology, 16(6):638–656, December 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-014-9579-7.

Gupta:2014:DRE

[1213] Kapil Gupta, Bhavanath Jha, and Pradeep K. Agarwal. A dehydrationresponsive element binding (DREB) transcription factor from the succulent halophyte Salicornia brachiata enhances abiotic stress tolerance in transgenic tobacco. Marine Biotechnology, 16(6):657–673, December 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-014-9582-z.

Qiu:2014:EKT

[1214] Chao Qiu, Bin Cheng, Yunsheng Zhang, Rong Huang, Lanjie Liao, Yongming Li, Daji Luo, Wei Hu, and Yaping Wang. Efficient knockout of transplanted green fluorescent protein gene in medaka using TALENs. Marine Biotechnology, 16(6):674–683, December 2014. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-014-9584-x.

Hanssen:2014:BDI

[1215] Kine Ø. Hanssen, Gunnar Cervin, Rozenn Trepos, Julie Petitbois, Tor Haug, Espen Hansen, Jeanette H. Andersen, Henrik Pavia, Claire Hellio, and Johan Svenson. The bromotyrosine derivative ianthelline isolated from the Arctic marine sponge Stryphnus fortis inhibits marine micro- and macrobiofouling. Marine Biotechnology, 16(6):684– 694, December 2014. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-014-9583-y.

Peng:2014:IPA

[1216] Yong Y. Peng, Veronica Glattauer, Timothy D. Skewes, Andrew McDevitt, Christopher M. Elvin, Jerome A. Werkmeister, Lloyd D. Graham, and John A. M. Ramshaw. Identification of proteins associated with adhesive prints from Holothuria dofleinii cuvierian tubules. Marine Biotechnology, 16(6):695–706, December 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-014-9586-8.

Vinoj:2014:QQA

[1217] G. Vinoj, B. Vaseeharan, S. Thomas, A. J. Spiers, and S. Shanthi. Quorum-quenching activity of the AHL–Lactonase from Bacillus licheniformis DAHB1 inhibits Vibrio biofilm formation In Vitro and reduces shrimp intestinal colonisation and mortality. Marine Biotechnology, 16 (6):707–715, December 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.com/

article/10.1007/s10126-014-9585-9.

Tapia-Paniagua:2014:PSI

[1218] Silvana Tapia-Paniagua, Carmen Lobo, Xabier Moreno-Ventas, Inés García de la Banda, Miguel A. Moriñigo, and M. Carmen Balebona. Probiotic supplementation influences the diversity of the intestinal microbiota during early stages of farmed Senegalese sole (Solea senegalensis, Kaup 1858). Marine Biotechnology, 16(6):716–728, December 2014. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-014-9588-6.

Ye:2014:GMQ

[1219] Hua Ye, Yang Liu, Xiande Liu, Xiaoqing Wang, and Zhiyong Wang. Genetic mapping and QTL analysis of growth traits in the large yellow croaker Larimichthys crocea. Marine Biotechnology, 16(6):729– 738, December 2014. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-014-9590-z.

Guan:2014:GRZ

[1220] Guijun Guan, Xi Zhang, Kiyoshi Naruse, Yoshitaka Nagahama, and Yunhan Hong. Gene replacement by zinc finger nucleases in medaka embryos. Marine Biotechnology, 16(6):739–747, December 2014. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-014-9587-7.

Kaur:2015:RCL

[1221] Simrat Kaur and Charles Spillane. Reduction in carotenoid levels in the marine diatom Phaeodactylum tricornutum by artificial MicroRNAs targeted against the endogenous phytoene synthase gene. Marine Biotechnology, 17(1):1–7, February 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-014-9593-9.

Molina-Luzon:2015:FHG

[1222] Ma. Jesús Molina-Luzón, Miguel Hermida, Rafael Navajas-Pérez, Francisca Robles, José Ignacio Navas, Carmelo Ruiz-Rejón, Carmen Bouza, Paulino Martínez, and Roberto de la Herrán. First haploid genetic map based on microsatellite markers in Senegalese sole (Solea senegalensis, Kaup 1858). Marine Biotechnology, 17(1):8–22, February 2015. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-014-9589-5.

Shiel:2015:NCG

[1223] Brett P. Shiel, Nathan E. Hall, Ira R. Cooke, Nicholas A. Robinson, and Jan M. Strugnell. De Novo characterisation of the greenlip abalone transcriptome (Haliotis laevigata) with a focus on the heat shock protein 70 (HSP70) family. Marine Biotechnology, 17(1):23–32, February 2015. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-014-9591-y.

Linthorne:2015:DPS

[1224] Jamie S. Linthorne, Barbara J. Chang, Gavin R. Flematti, Emilio L. Ghisalberti, and David C. Sutton. A direct pre-screen for marine bacteria producing compounds inhibiting quorum sensing reveals diverse planktonic bacteria that are bioactive. Marine Biotechnology, 17(1): 33–42, February 2015. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-014-9592-x.

Koopmans:2015:SVF

[1225] Marieke Koopmans, Pieter van Rijswijk, Henricus T. S. Boschker, Houtekamer Marco, Dirk Martens, and Rene H. Wijffels. Seasonal variation of fatty acids and stable carbon isotopes in sponges as indicators for nutrition: Biomarkers in sponges identified. Marine Biotechnology, 17 (1):43–54, February 2015. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-014-9594-8.

Yu:2015:UDB

[1226] Zhenzhen Yu, Qi Li, Lingfeng Kong, and Hong Yu. Utility of DNA barcoding for tellinoidea: a comparison of distance, coalescent and character-based methods on multiple genes. Marine Biotechnology, 17 (1):55–65, February 2015. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-014-9596-6.

Anthony:2015:UFI

[1227] Josephine Anthony, Vijaya Raghavan Rangamaran, Dharani Gopal, Kumar T. Shivasankarasubbiah, Mary Leema J. Thilagam, Magesh Peter Dhassiah, Divya Shridhar M. Padinjattayil, VinithKumar N. Valsalan, Vijayakumaran Manambrakat, Sivakumar Dakshinamurthy, Sivaraman Thirunavukkarasu, and Kirubagaran Ramalingam. Ultraviolet and 5′ fluorodeoxyuridine induced random mutagenesis in Chlorella vulgaris and its impact on fatty acid profile: a new insight on lipid-metabolizing genes and structural characterization of related proteins. Marine Biotechnology, 17(1):66–80, February 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-014-9597-5.

Barghi:2015:HCD

[1228] Neda Barghi, Gisela P. Concepcion, Baldomero M. Olivera, and Arturo O. Lluisma. High conopeptide diversity in Conus tribblei revealed through analysis of venom duct transcriptome using two highthroughput sequencing platforms. Marine Biotechnology, 17(1):81–98, February 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-

2236 (electronic). URL https://link.springer.com/article/10.
1007/s10126-014-9595-7.

Qin:2015:PDA

[1229] Xiao-Yan Qin, Kai-Lin Yang, Jing Li, Chang-Yun Wang, and Chang-Lun Shao. Phylogenetic diversity and antibacterial activity of culturable fungi derived from the zoanthid Palythoa haddoni in the South China Sea. Marine Biotechnology, 17(1):99–109, February 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-014-9598-4.

Wu:2015:AIM

[1230] Bin Wu, Birgit Ohlendorf, Vanessa Oesker, Jutta Wiese, Susann Malien, Rolf Schmaljohann, and Johannes F. Imhoff. Acetylcholinesterase inhibitors from a marine fungus Talaromyces sp. strain LF458. Marine Biotechnology, 17(1):110–119, February 2015. CODEN MABIFW.

ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-014-9599-3.

Pang:2015:PAE

[1231] Ka-Lai Pang, Han-Jia Lin, Hung-Yun Lin, Yu-Fen Huang, and Yi-Min Chen. Production of arachidonic and eicosapentaenoic acids by the marine oomycete Halophytophthora. Marine Biotechnology, 17(2): 121–129, April 2015. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-014-9600-1.

Golotin:2015:RPC

[1232] Vasily Golotin, Larissa Balabanova, Galina Likhatskaya, and Valery Rasskazov. Recombinant production and characterization of a highly active alkaline phosphatase from marine bacterium Cobetia marina. Marine Biotechnology, 17(2):130–143, April 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-014-9601-0.

DeVitis:2015:MMS

[1233] Valerio De Vitis, Benedetta Guidi, Martina Letizia Contente, Tiziana Granato, Paola Conti, Francesco Molinari, Elena Crotti, Francesca Mapelli, Sara Borin, Daniele Daffonchio, and Diego Romano. Marine microorganisms as source of stereoselective esterases and ketoreductases: Kinetic resolution of a prostaglandin intermediate. Marine Biotechnology, 17(2):144–152, April 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.com/

article/10.1007/s10126-014-9602-z.

Suzuki:2015:ILI

[1234] Michio Suzuki, Toshihiro Kogure, Shohei Sakuda, and Hiromichi Nagasawa. Identification of ligament intra-crystalline peptide (LICP) from the hinge ligament of the bivalve, Pinctada fucata. Marine Biotechnology, 17(2):153–161, April 2015. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-014-9603-y.

Yan:2015:ZFC

[1235] Chuan Yan, Weiling Zheng, and Zhiyuan Gong. Zebrafish fgf10b has a complementary function to fgf10a in liver and pancreas development. Marine Biotechnology, 17(2):162–167, April 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-014-9604-x.

Sairi:2015:DCR

[1236] Fareed Sairi, Peter Valtchev, Vincent G. Gomes, and Fariba Dehghani. Distribution and characterization of rhogocyte cell types in the mantle tissue of Haliotis laevigata. Marine Biotechnology, 17(2):168–179, April 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-014-9605-9.

Zhu:2015:MSR

[1237] Xin Zhu, Dunxue Chen, Yi Hu, Ping Wu, Kaizuo Wang, Junzhi Zhang, Wuying Chu, and Jianshe Zhang. The microRNA signature in response to nutrient restriction and refeeding in skeletal muscle of Chinese perch (Siniperca chuatsi). Marine Biotechnology, 17(2):180–189, April 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-014-9606-8.

Chen:2015:CTP

[1238] Xin Chen, Jie Mei, Junjie Wu, Jing Jing, Wenge Ma, Jin Zhang, Cheng Dan, Weimin Wang, and Jian-Fang Gui. A comprehensive transcriptome provides candidate genes for sex determination/differentiation and SSR/SNP markers in yellow catfish. Marine Biotechnology, 17(2): 190–198, April 2015. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-014-9607-7.

Matsuda:2015:PEB

[1239] Ryuya Matsuda, Rengin Ozgur, Yuya Higashi, Katsuaki Takechi, Hiroyoshi Takano, and Susumu Takio. Preferential expression of a bromoperoxidase in sporophytes of a red alga, Pyropia yezoensis. Marine Biotechnology, 17(2):199–210, April 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-014-9608-6.

Martins:2015:BAS

[1240] Mariana P. Martins, Jamal Ouazzani, Guillaume Arcile, Alex H. Jeller, João P. F. de Lima, Mirna H. R. Seleghim, Ana Lígia L. Oliveira, Hosana M. Debonsi, Tiago Venâncio, Nair S. Yokoya, Mutue T. Fujii, and André L. M. Porto. Biohydroxylation of (-)-Ambrox(R), (-)- Sclareol, and (+)-Sclareolide by whole cells of Brazilian marine-derived fungi. Marine Biotechnology, 17(2):211–218, April 2015. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-015-9610-7.

Chen:2015:PCN

[1241] Yan-Li Chen, Wen-Jun Mao, Hong-Wen Tao, Wei-Ming Zhu, Meng-Xia Yan, Xue Liu, Tian-Tian Guo, and Tao Guo. Preparation and characterization of a novel extracellular polysaccharide with antioxidant activity, from the mangrove-associated fungus Fusarium oxysporum. Marine Biotechnology, 17(2):219–228, April 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-015-9611-6.

Kocmarek:2015:CGR

[1242] Andrea L. Kocmarek, Moira M. Ferguson, and Roy G. Danzmann. Comparison of growth-related traits and gene expression profiles between the offspring of neomale (XX) and normal male (XY) rainbow trout. Marine Biotechnology, 17(2):229–243, April 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-015-9612-5.

Badary:2015:DCE

[1243] Amr Badary, Koichi Abe, Stefano Ferri, Katsuhiro Kojima, and Koji Sode. The development and characterization of an exogenous greenlight-regulated gene expression system in marine cyanobacteria. Marine Biotechnology, 17(3):245–251, June 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-015-9616-1.

Sellars:2015:TPP

[1244] Melony J. Sellars, Carolyn Trewin, Sean M. McWilliam, R. S. E. Glaves, and Philip L. Hertzler. Transcriptome profiles of Penaeus (Marsupenaeus) japonicus animal and vegetal half-embryos: Identification of sex determination, germ line, mesoderm, and other developmental genes. Marine Biotechnology, 17(3):252–265, June 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-015-9613-4.

Tabarsa:2015:SAR

[1245] Mehdi Tabarsa, Geun-Mook Park, Il-Shik Shin, EunJung Lee, Jin-Kyung Kim, and SangGuan You. Structure-activity relationships of sulfated glycoproteins from Codium fragile on nitric oxide releasing capacity from RAW264.7 cells. Marine Biotechnology, 17(3):266–276, June 2015. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-015-9615-2.

Kurata:2015:HLD

[1246] Atsushi Kurata, Mioko Matsumoto, Tohru Kobayashi, Shigeru Deguchi, and Noriaki Kishimoto. Hyaluronate lyase of a deep-sea Bacillus niacini. Marine Biotechnology, 17(3):277–284, June 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-015-9618-z.

Goddard:2015:FGM

[1247] Stephen Goddard, Madjid Delghandi, Sergey Dobretsov, Hamed Al-Oufi, Saoud Al-Habsi, and J. Grant Burgess. The first GCC marine biotechnology symposium: Emerging opportunities and future perspectives. Marine Biotechnology, 17(3):285–289, June 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-015-9617-0.

Brack:2015:AMB

[1248] Christiane Brack, Annett Mikolasch, Rabea Schlueter, Andreas Otto, Dörte Becher, Uwe Wegner, Dirk Albrecht, Katharina Riedel, and Frieder Schauer. Antibacterial metabolites and bacteriolytic enzymes produced by Bacillus pumilus during bacteriolysis of Arthrobacter citreus. Marine Biotechnology, 17(3):290–304, June 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-015-9614-3.

Sun:2015:EIP

[1249] Wen-Juan Sun, Ming Zhu, Yuan-Li Wang, Qing Li, Hong-Dan Yang, Ze-Lin Duan, Lin He, and Qun Wang. ERK is involved in the process of acrosome reaction In Vitro of the Chinese mitten crab, Eriocheir sinensis. Marine Biotechnology, 17(3):305–316, June 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-015-9619-y.

Ng:2015:DGE

[1250] Grace Hwee Boon Ng, Hongyan Xu, Na Pi, Barry C. Kelly, and Zhiyuan Gong. Differential GFP expression patterns induced by different heavy metals in Tg(hsp70:gfp) transgenic medaka (Oryzias latipes). Marine Biotechnology, 17(3):317–327, June 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-015-9620-5.

Liu:2015:ISN

[1251] Sixin Liu, Roger L. Vallejo, Guangtu Gao, Yniv Palti, Gregory M. Weber, Alvaro Hernandez, and Caird E. Rexroad. Identification of single-nucleotide polymorphism markers associated with cortisol response to crowding in rainbow trout. Marine Biotechnology, 17(3): 328–337, June 2015. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-015-9621-4.

Gong:2015:EPA

[1252] Guangye Gong, Zhenxia Sha, Songlin Chen, Chao Li, Hui Yan, Yadong Chen, and Tianzi Wang. Expression profiling analysis of the microRNA response of Cynoglossus semilaevis to Vibrio anguillarum and other stimuli. Marine Biotechnology, 17(3):338–352, June 2015. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-015-9623-2.

Ulloa:2015:RSS

[1253] Pilar E. Ulloa, Gonzalo Rincón, Alma Islas-Trejo, Cristian Araneda, Patricia Iturra, Roberto Neira, and Juan F. Medrano. RNA sequencing to study gene expression and SNP variations associated with growth in zebrafish fed a plant protein-based diet. Marine Biotechnology, 17 (3):353–363, June 2015. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-015-9624-1.

Yue:2015:TAS

[1254] Xin Yue, Qing Nie, Guoqiang Xiao, and Baozhong Liu. Transcriptome analysis of shell color-related genes in the clam Meretrix meretrix. Marine Biotechnology, 17(3):364–374, June 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-015-9625-0.

Dolashki:2015:RAS

[1255] Aleksandar Dolashki, Mariana Radkova, Elena Todorovska, Martin Ivanov, Stefan Stevanovic, Laura Molin, Piero Traldi, Wolfgang Voelter, and Pavlina Dolashka. RETRACTED ARTICLE: Structure and characterization of Cancer pagurus hemocyanin. Marine Biotechnology, 17(3):375, June 2015. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-014-9609-5.

Steinert:2015:DAA

[1256] Georg Steinert, Michael W. Taylor, and Peter J. Schupp. Diversity of Actinobacteria associated with the marine ascidian Eudistoma toealensis. Marine Biotechnology, 17(4):377–385, August 2015. CODEN MABIFW.

ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-015-9622-3.

Zielonka:2015:SST

[1257] Stefan Zielonka, Martin Empting, Doreen Könning, Julius Grzeschik, Simon Krah, Stefan Becker, Stephan Dickgießer, and Harald Kolmar. The shark strikes twice: Hypervariable loop 2 of shark IgNAR antibody variable domains and its potential to function as an autonomous paratope. Marine Biotechnology, 17(4):386–392, August 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-015-9642-z.

Pozzolini:2015:MCC

[1258] Marina Pozzolini, Sonia Scarfì, Francesca Mussino, Sara Ferrando, Lorenzo Gallus, and Marco Giovine. Molecular cloning, characterization, and expression analysis of a prolyl 4-hydroxylase from the marine sponge Chondrosia reniformis. Marine Biotechnology, 17(4): 393–407, August 2015. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-015-9630-3.

Shao:2015:PAM

[1259] Chang-Lun Shao, Ru-Fang Xu, Chang-Yun Wang, Pei-Yuan Qian, Kai-Ling Wang, and Mei-Yan Wei. Potent antifouling marine dihydroquinolin-2(1 H)-one-containing alkaloids from the gorgonian coral-derived fungus Scopulariopsis sp. Marine Biotechnology, 17(4): 408–415, August 2015. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-015-9628-x.

Bibo-Verdugo:2015:CDT

[1260] Betsaida Bibo-Verdugo, Liliana Rojo-Arreola, Maria A. Navarrete del Toro, and Fernando García-Carreño. A chymotrypsin from the digestive tract of California spiny lobster, Panulirus interruptus: Purification and biochemical characterization. Marine Biotechnology, 17(4): 416–427, August 2015. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-015-9626-z.

Tsuji:2015:QAC

[1261] Yoshinori Tsuji, Masatoshi Yamazaki, Iwane Suzuki, and Yoshihiro Shiraiwa. Quantitative analysis of carbon flow into photosynthetic products functioning as carbon storage in the marine coccolithophore, Emiliania huxleyi. Marine Biotechnology, 17(4):428–440, August 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-015-9632-1.

Pan:2015:ISS

[1262] Zheng-Jun Pan, Xi-Yin Li, Feng-Jian Zhou, Xiao-Gang Qiang, and Jian-Fang Gui. Identification of sex-specific markers reveals male heterogametic sex determination in Pseudobagrus ussuriensis. Marine Biotechnology, 17(4):441–451, August 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-015-9631-2.

Sabatino:2015:EGT

[1263] Valeria Sabatino, Monia Teresa Russo, Shrikant Patil, Giuliana d’Ippolito, Angelo Fontana, and Maria Immacolata Ferrante. Establishment of genetic transformation in the sexually reproducing diatoms Pseudo-nitzschia multistriata and Pseudo-nitzschia arenysensis and inheritance of the transgene. Marine Biotechnology, 17(4):452–462, August 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-015-9633-0.

Choudhury:2015:PGG

[1264] Jayanta D. Choudhury, Arnab Pramanik, Nicole S. Webster, Lyndon E. Llewellyn, Ratan Gachhui, and Joydeep Mukherjee. The pathogen of the Great Barrier Reef sponge Rhopaloeides odorabile is a new strain of Pseudoalteromonas agarivorans containing abundant and diverse virulencerelated genes. Marine Biotechnology, 17(4):463–478, August 2015. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-015-9627-y.

Kale:2015:CLM

[1265] Varsha Kale, Ólafur Fridjónsson, Jón Óskar Jónsson, Hördur G. Kristinsson, Sesselja Ómarsdóttir, and Gudmundur Ó. Hreggvidsson. Chondroitin lyase from a marine Arthrobacter sp. MAT3885 for the production of chondroitin sulfate disaccharides. Marine Biotechnology, 17(4): 479–492, August 2015. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-015-9629-9.

Maralit:2015:MAI

[1266] Benedict Arias Maralit, Mami Komatsu, Sheryll Grospe Hipolito, Ikuo Hirono, and Hidehiro Kondo. Microarray analysis of immunity against WSSV in response to injection of non-specific long dsRNA in kuruma shrimp, Marsupenaeus japonicus. Marine Biotechnology, 17(4): 493–501, August 2015. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-015-9637-9.

Koyama:2015:ISL

[1267] Takashi Koyama, Akiyuki Ozaki, Kazunori Yoshida, Junpei Suzuki, Kanako Fuji, Jun ya Aoki, Wataru Kai, Yumi Kawabata, Tatsuo Tsuzaki, Kazuo Araki, and Takashi Sakamoto. Identification of sex-linked SNPs and sex-determining regions in the yellowtail genome. Marine Biotechnology, 17(4):502–510, August 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-015-9636-x.

Ma:2015:GMM

[1268] Zai-Chao Ma, Nan-Nan Liu, Zhe Chi, Guang-Lei Liu, and Zhen-Ming Chi. Genetic modification of the marine-isolated yeast Aureobasidium melanogenum P16 for efficient pullulan production from inulin. Marine Biotechnology, 17(4):511–522, August 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-015-9638-8.

Miao:2015:ITP

[1269] Yan Miao, Lingling Zhang, Yan Sun, Wenqian Jiao, Yangping Li, Jin Sun, Yangfan Wang, Shi Wang, Zhenmin Bao, and Weizhi Liu. Integration of transcriptomic and proteomic approaches provides a core set of genes for understanding of scallop attachment. Marine Biotechnology, 17 (5):523–532, October 2015. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-015-9635-y.

Bolanos:2015:PDS

[1270] Jessica Bolaños, Luis Fernando De León, Edgardo Ochoa, José Darias, Huzefa A. Raja, Carol A. Shearer, Andrew N. Miller, Patrick Vanderheyden, Andrea Porras-Alfaro, and Catherina Caballero-George. Phylogenetic diversity of sponge-associated fungi from the Caribbean and the Pacific of Panama and their In Vitro effect on angiotensin and endothelin receptors. Marine Biotechnology, 17(5):533–564, October 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-015-9634-z.

Dong:2015:PEM

[1271] Chuanju Dong, Jian Xu, Baosen Wang, Jianxin Feng, Zsigmond Jeney, Xiaowen Sun, and Peng Xu. Phylogeny and evolution of multiple common carp (Cyprinus carpio L.) populations clarified by phylogenetic analysis based on complete mitochondrial genomes. Marine Biotechnology, 17 (5):565–575, October 2015. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-015-9639-7.

Verleih:2015:ITS

[1272] Marieke Verleih, Andreas Borchel, Aleksei Krasnov, Alexander Rebl, Tomáš Korytář, Carsten Kühn, and Tom Goldammer. Impact of thermal stress on kidney-specific gene expression in farmed regional and imported rainbow trout. Marine Biotechnology, 17(5):576–592, October 2015. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-015-9640-1.

Pang:2015:TAD

[1273] Shao-Chen Pang, Hou-Peng Wang, Zuo-Yan Zhu, and Yong-Hua Sun. Transcriptional activity and DNA methylation dynamics of the Gal4/ UAS system in zebrafish. Marine Biotechnology, 17(5):593–603, October 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-015-9641-0.

Qin:2015:IAF

[1274] Qinbo Qin, Juan Wang, Jing Dai, YuDe Wang, Yun Liu, and Shaojun Liu. Induced all-female autotriploidy in the allotetraploids of Carassius auratus red var. ([female sign]) × Megalobrama amblycephala ([male sign]). Marine Biotechnology, 17(5):604–612, October 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-015-9647-7.

Lu:2015:TAR

[1275] Jianguo Lu, Min Zheng, Jiajia Zheng, Jian Liu, Yongzhuang Liu, Lina Peng, Pingping Wang, Xiaofeng Zhang, Qiushi Wang, Peixian Luan, Shahid Mahbooband, and Xiaowen Sun. Transcriptomic analyses reveal novel genes with sexually dimorphic expression in yellow catfish (Pelteobagrus fulvidraco) brain. Marine Biotechnology, 17(5): 613–623, October 2015. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-015-9650-z.

Wang:2015:LTD

[1276] Kai-Ling Wang, Ying Xu, Liang Lu, Yongxin Li, Zhuang Han, Jun Zhang, Chang-Lun Shao, Chang-Yun Wang, and Pei-Yuan Qian. Lowtoxicity diindol-3-ylmethanes as potent antifouling compounds. Marine Biotechnology, 17(5):624–632, October 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-015-9656-6.

Tang:2015:ICF

[1277] Chuan-Ho Tang, Ping-Chang Ku, Ching-Yu Lin, Te-Hao Chen, Kuo-Hsin Lee, Shu-Hui Lee, and Wei-Hsien Wang. Intra-colonial functional differentiation-related modulation of the cellular membrane in a pocilloporid coral Seriatopora caliendrum. Marine Biotechnology, 17(5): 633–643, October 2015. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-015-9645-9.

Morita:2015:FSY

[1278] Tetsuro Morita, Kagayaki Morishima, Misako Miwa, Naoki Kumakura, Satomi Kudo, Kensuke Ichida, Toru Mitsuboshi, Yutaka Takeuchi, and Goro Yoshizaki. Functional sperm of the yellowtail (Seriola quinqueradiata) were produced in the small-bodied surrogate, jack mackerel (Trachurus japonicus). Marine Biotechnology, 17(5):644–654, October 2015. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-015-9657-5.

Ge:2015:ISL

[1279] Jianlong Ge, Qi Li, Hong Yu, and Lingfeng Kong. Identification of single-locus PCR-based markers linked to shell background color in the Pacific oyster (Crassostrea gigas). Marine Biotechnology, 17(5): 655–662, October 2015. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-015-9652-x.

Kaitetzidou:2015:IPF

[1280] Elisavet Kaitetzidou, Stavros Chatzifotis, Efthimia Antonopoulou, and Elena Sarropoulou. Identification, phylogeny, and function of fabp2 paralogs in two non-model teleost fish species. Marine Biotechnology, 17(5): 663–677, October 2015. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-015-9648-6.

Koyama:2015:ERL

[1281] Sumihiro Koyama, Shinro Nishi, Maki Tokuda, Moeka Uemura, Yoichi Ishikawa, Takeshi Seya, Seinen Chow, Yuji Ise, Yuji Hatada, Yoshihiro Fujiwara, and Taishi Tsubouchi. Electrical retrieval of living microorganisms from cryopreserved marine sponges using a potential-controlled electrode. Marine Biotechnology, 17(5):678–692, October 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-015-9651-y.

Cui:2015:BAG

[1282] Yan Cui, Hongwei Wang, Xuemei Qiu, Haijin Liu, and Runqing Yang. Bayesian analysis for genetic architectures of body weights and morphological traits using distorted markers in Japanese flounder Paralichthys olivaceus. Marine Biotechnology, 17(6):693–702, December 2015. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-015-9646-8.

Maryoung:2015:DGE

[1283] Lindley A. Maryoung, Ramon Lavado, Theo K. Bammler, Evan P. Gallagher, Patricia L. Stapleton, Richard P. Beyer, Federico M. Farin, Gary Hardiman, and Daniel Schlenk. Differential gene expression in liver, gill, and olfactory rosettes of Coho salmon (Oncorhynchus kisutch) after acclimation to salinity. Marine Biotechnology, 17(6):703–717, December 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-015-9649-5.

Synytsya:2015:SFA

[1284] Andriy Synytsya, Doo Jin Choi, Radek Pohl, Ye Seul Na, Peter Capek, Erika Lattová, Tomáš Taubner, Ji Won Choi, Chang Won Lee, Jae Kweon Park, Woo Jung Kim, Sung Min Kim, Jisun Lee, and Yong Il Park. Structural features and anti-coagulant activity of the sulphated polysaccharide SPS–CF from a green alga Capsosiphon fulvescens. Marine Biotechnology, 17(6):718–735, December 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-015-9643-y.

Rocha:2015:SBA

[1285] Lenilson C. Rocha, Mirna H. R. Seleghim, João V. Comasseto, Lara D. Sette, and André L. M. Porto. Stereoselective bioreduction of α-azido ketones by whole cells of marine-derived fungi. Marine Biotechnology, 17(6):736–742, December 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.com/

article/10.1007/s10126-015-9644-x.

Dolashki:2015:SCE

[1286] A. Dolashki, M. Radkova, E. Todorovska, M. Ivanov, S. Stevanovic, L. Molin, P. Traldi, W. Voelter, and P. Dolashka. Structure and characterization of Eriphia verrucosa hemocyanin. Marine Biotechnology, 17(6):743–752, December 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.com/

article/10.1007/s10126-015-9653-9.

Abernathy:2015:FSC

[1287] Jason Abernathy, Stéphane Panserat, Thomas Welker, Elisabeth Plagne-Juan, Dionne Sakhrani, David A. Higgs, Florence Audouin, Robert H. Devlin, and Ken Overturf. Food shortage causes differential effects on body composition and tissue-specific gene expression in salmon modified for increased growth hormone production. Marine Biotechnology, 17(6): 753–767, December 2015. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-015-9654-8.

Giannetto:2015:HIF

[1288] Alessia Giannetto, Maria Maisano, Tiziana Cappello, Sabrina Oliva, Vincenzo Parrino, Antonino Natalotto, Giuseppe De Marco, Chiara Barberi, Orazio Romeo, Angela Mauceri, and Salvatore Fasulo. Hypoxia-inducible factor α and Hif-prolyl hydroxylase characterization and gene expression in short-time air-exposed Mytilus galloprovincialis. Marine Biotechnology, 17(6):768–781, December 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-015-9655-7.

Kim:2015:SDM

[1289] Hae Sol Kim, Yu Jeong Chu, Chang-Ho Park, Eun Yeol Lee, and Hee Sook Kim. Site-directed mutagenesis-based functional analysis and characterization of endolytic lyase activity of N- and C-terminal domains of a novel oligoalginate lyase from Sphingomonas sp. MJ-3 possessing exolytic lyase activity in the intact enzyme. Marine Biotechnology, 17(6): 782–792, December 2015. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-015-9658-4.

Chavez-Mardones:2015:NGT

[1290] Jacqueline Chávez-Mardones and Cristian Gallardo-Escárate. Nextgeneration transcriptome profiling of the salmon louse Caligus rogercresseyi exposed to deltamethrin (AlphaMax^{T M}): Discovery of relevant genes and sex-related differences. Marine Biotechnology, 17(6): 793–810, December 2015. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-015-9661-9.

Cheung:2015:RCM

[1291] Man Kit Cheung, Ho Yin Yip, Wenyan Nong, Patrick Tik Wan Law, Ka Hou Chu, Hoi Shan Kwan, and Jerome Ho Lam Hui. Rapid change of microbiota diversity in the gut but not the hepatopancreas during gonadal development of the new shrimp model Neocaridina denticulata. Marine Biotechnology, 17(6):811–819, December 2015. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-015-9662-8.

Araujo:2015:NZP

[1292] Carlos Araújo, Estefanía Muñoz-Atienza, Tania Pérez-Sánchez, Patrícia Poeta, Gilberto Igrejas, Pablo E. Hernández, Carmen Herranz, Imanol Ruiz-Zarzuela, and Luis M. Cintas. Nisin Z production by Lactococcus lactis subsp. cremoris WA2-67 of aquatic origin as a defense mechanism to protect rainbow trout (Oncorhynchus mykiss, Walbaum) against Lactococcus garvieae. Marine Biotechnology, 17(6):820–830, December 2015. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-015-9660-x.

Xu:2015:GGT

[1293] Hongyan Xu, Caixia Li, Yan Li, Grace Hwee Boon Ng, Chunsheng Liu, Xiaoyan Zhang, and Zhiyuan Gong. Generation of Tg(cyp1a:gfp) transgenic zebrafish for development of a convenient and sensitive In Vivo assay for aryl hydrocarbon receptor activity. Marine Biotechnology, 17(6): 831–840, December 2015. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-015-9669-1.

Ghosh:2015:AMO

[1294] Bikramjit Ghosh, Barbara F. Nowak, and Andrew R. Bridle. Alginate microencapsulation for oral immunisation of finfish: Release characteristics, Ex Vivo intestinal uptake and In Vivo administration in Atlantic salmon, Salmo salar L. Marine Biotechnology, 17(6):841– 853, December 2015. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-015-9663-7.

Fu:2016:CCP

[1295] Ge-Yi Fu, Yi Lu, Zhe Chi, Guang-Lei Liu, Shou-Feng Zhao, Hong Jiang, and Zhen-Ming Chi. Cloning and characterization of a pyruvate carboxylase gene from Penicillium rubens and overexpression of the genein the yeast Yarrowia lipolytica for enhanced citric acid production. Marine Biotechnology, 18(1):1–14, February 2016. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-015-9665-5.

Takagi:2016:PAA

[1296] Toshiyuki Takagi, Hironobu Morisaka, Shunsuke Aburaya, Yohei Tatsukami, Kouichi Kuroda, and Mitsuyoshi Ueda. Putative alginate assimilation process of the marine bacterium Saccharophagus degradans 2-40 based on quantitative proteomic analysis. Marine Biotechnology, 18 (1):15–23, February 2016. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-015-9667-3.

Jerez:2016:ENS

[1297] Celia G. Jerez, José R. Malapascua, Magda Sergejevová, Félix L. Figueroa, and Jiří Masojídek. Effect of nutrient starvation under high irradiance on lipid and starch accumulation in Chlorella fusca (Chlorophyta). Marine Biotechnology, 18(1):24–36, February 2016. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-015-9664-6.

Yan:2016:EPA

[1298] Hui Yan, Yadong Chen, Shun Zhou, Chao Li, Guangye Gong, Xuejie Chen, Tianzi Wang, Songlin Chen, and Zhenxia Sha. Expression profile analysis of miR-221 and miR-222 in different tissues and head kidney cells of Cynoglossus semilaevis, following pathogen infection. Marine Biotechnology, 18(1):37–48, February 2016. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-015-9668-2.

Jiang:2016:CBL

[1299] Likun Jiang, Weiwei You, Xiaojun Zhang, Jian Xu, Yanliang Jiang, Kai Wang, Zixia Zhao, Baohua Chen, Yunfeng Zhao, Shahid Mahboob, Khalid A. Al-Ghanim, Caihuan Ke, and Peng Xu. Construction of the BAC library of small abalone (Haliotis diversicolor) for gene screening and genome characterization. Marine Biotechnology, 18(1): 49–56, February 2016. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-015-9666-4.

Kessuwan:2016:DGR

[1300] Kanonkporn Kessuwan, Satoshi Kubota, Qi Liu, Motohiko Sano, Nobuaki Okamoto, Takashi Sakamoto, Hirofumi Yamashita, Yoji Nakamura, and Akiyuki Ozaki. Detection of growth-related quantitative trait loci and high-resolution genetic linkage maps using simple sequence repeat markers in the kelp grouper (Epinephelus bruneus). Marine Biotechnology, 18(1):57–84, February 2016. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-015-9673-5.

Gonzalez:2016:CMS

[1301] Juan D. González, Jonás I. Silva-Marrero, Isidoro Metón, Albert Caballero-Solares, Ivan Viegas, Felipe Fernández, Montserrat Miñarro, Anna Fàbregas, Josep R. Ticó, John G. Jones, and Isabel V. Baanante. Chitosan-mediated shRNA knockdown of cytosolic alanine aminotransferase improves hepatic carbohydrate metabolism. Marine Biotechnology, 18(1):85–97, February 2016. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-015-9670-8.

Negrin-Baez:2016:IQT

[1302] Davinia Negrín-Báez, Ana Navarro, Silvia T. Rodríguez-Ramilo, Juan M. Afonso, and María J. Zamorano. Identification of quantitative trait loci associated with the skeletal deformity LSK complex in gilthead seabream (Sparus aurata L.). Marine Biotechnology, 18(1):98–106, February 2016. CODEN MABIFW. ISSN 1436-2228 (print), 1436-

2236 (electronic). URL https://link.springer.com/article/10.
1007/s10126-015-9671-7.

Liu:2016:MQR

[1303] Peng Liu, Le Wang, Zi Yi Wan, Bao Qing Ye, Shuqing Huang, Sek-Man Wong, and Gen Hua Yue. Mapping QTL for resistance against viral nervous necrosis disease in Asian seabass. Marine Biotechnology, 18(1): 107–116, February 2016. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-015-9672-6.

Feijo:2016:SGI

[1304] Rubens G. Feijó, André L. Braga, Carlos F. C. Lanes, Márcio A. Figueiredo, Luis A. Romano, Marta C. Klosterhoff, Luis E. M. Nery, Rodrigo Maggioni, Wilson Wasielesky, and Luis F. Marins. Silencing of gonad-inhibiting hormone transcripts in Litopenaeus vannamei females by use of the RNA interference technology. Marine Biotechnology, 18(1): 117–123, February 2016. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-015-9676-2.

Jiang:2016:IVG

[1305] Qun Jiang, Qi Li, Hong Yu, and Lingfeng Kong. Inheritance and variation of genomic DNA methylation in diploid and triploid Pacific oyster (Crassostrea gigas). Marine Biotechnology, 18(1):124–132, February 2016. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-015-9674-4.

Prechoux:2016:EAP

[1306] Aurélie Préchoux, Sabine Genicot, Hélène Rogniaux, and William Helbert. Enzyme-assisted preparation of furcellaran-like κ-/β-carrageenan. Marine Biotechnology, 18(1):133–143, February 2016. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-015-9675-3.

Hirayama:2016:HMSa

[1307] Makoto Hirayama, Hiromi Shibata, Koji Imamura, Takemasa Sakaguchi, and Kanji Hori. High-mannose specific lectin and its recombinants from a carrageenophyta Kappaphycus alvarezii represent a potent anti-HIV activity through high-affinity binding to the viral envelope glycoprotein gp120. Marine Biotechnology, 18(1):144–160, February 2016. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-015-9677-1.

Konning:2016:IPS

[1308] Doreen Könning, Stefan Zielonka, Carolin Sellmann, Christian Schröter, Julius Grzeschik, Stefan Becker, and Harald Kolmar. Isolation of a pHsensitive IgNAR variable domain from a yeast-displayed, histidine-doped master library. Marine Biotechnology, 18(2):161–167, April 2016. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-016-9690-z.

Niu:2016:DTA

[1309] Donghong Niu, Fei Wang, Shumei Xie, Fanyue Sun, Ze Wang, Maoxiao Peng, and Jiale Li. Developmental transcriptome analysis and identification of genes involved in larval metamorphosis of the razor clam, Sinonovacula constricta. Marine Biotechnology, 18(2):168–175, April 2016. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-016-9691-y.

Zhu:2016:ECS

[1310] Qihui Zhu, Linlin Zhang, Li Li, Huayong Que, and Guofan Zhang. Expression characterization of stress genes under high and low temperature stresses in the Pacific oyster, Crassostrea gigas. Marine Biotechnology, 18(2):176–188, April 2016. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-015-9678-0.

Lim:2016:GTA

[1311] Ee-Leen Lim, Rouh-San Siow, Raha Abdul Rahim, and Chai-Ling Ho. Global transcriptome analysis of Gracilaria changii (Rhodophyta) in response to agarolytic enzyme and bacterium. Marine Biotechnology, 18 (2):189–200, April 2016. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-015-9680-6.

Bibo-Verdugo:2016:CPB

[1312] Betsaida Bibo-Verdugo, Anthony J. O’Donoghue, Liliana Rojo-Arreola, Charles S. Craik, and Fernando García-Carreño. Complementary proteomic and biochemical analysis of peptidases in lobster gastric juice uncovers the functional role of individual enzymes in food digestion. Marine Biotechnology, 18(2):201–214, April 2016. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-015-9681-5.

Hirayama:2016:HMSb

[1313] Makoto Hirayama, Hiromi Shibata, Koji Imamura, Takemasa Sakaguchi, and Kanji Hori. High-mannose specific lectin and its recombinants from a carrageenophyta Kappaphycus alvarezii represent a potent anti-HIV activity through high-affinity binding to the viral envelope glycoprotein gp120. Marine Biotechnology, 18(2):215–231, April 2016. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-015-9684-2.

He:2016:CBC

[1314] Tianliang He and Xiaobo Zhang. Characterization of bacterial communities in deep-sea hydrothermal vents from three oceanic regions. Marine Biotechnology, 18(2):232–241, April 2016. CODEN MABIFW.

ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-015-9683-3.

Ren:2016:IVM

[1315] Jianfeng Ren, Zhanhui Hou, Haiyan Wang, Ming an Sun, Xiao Liu, Bin Liu, and Ximing Guo. Intraspecific variation in mitogenomes of five Crassostrea species provides insight into oyster diversification and speciation. Marine Biotechnology, 18(2):242–254, April 2016. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-016-9686-8.

Qin:2016:ELH

[1316] Zhenkui Qin, Yun Li, Baofeng Su, Qi Cheng, Zhi Ye, Dayan A. Perera, Michael Fobes, Mei Shang, and Rex A. Dunham. Editing of the luteinizing hormone gene to sterilize channel catfish, Ictalurus punctatus, using a modified zinc finger nuclease technology with electroporation. Marine Biotechnology, 18(2):255–263, April 2016. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-016-9687-7.

Bai:2016:ZET

[1317] Hong Bai, Wen-Wen Kong, Chang-Lun Shao, Yun Li, Yun-Zhang Liu, Min Liu, Fei-Fei Guan, and Chang-Yun Wang. Zebrafish embryo toxicity microscale model for ichthyotoxicity evaluation of marine natural products. Marine Biotechnology, 18(2):264–270, April 2016. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-016-9688-6.

Wang:2016:IFP

[1318] Jian Wang, Liyuan Fu, Prasanthi P. Koganti, Lei Wang, Jacqelyn M. Hand, Hao Ma, and Jianbo Yao. Identification and functional prediction of large intergenic noncoding RNAs (lincRNAs) in rainbow trout (Oncorhynchus mykiss). Marine Biotechnology, 18(2):271–282, April 2016. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-016-9689-5.

Peng:2016:IML

[1319] Lung-Hsiang Peng, Binesh Unnikrishnan, Chi-Yu Shih, Tung-Ming Hsiung, Jeng Chang, Pang-Hung Hsu, Tai-Chia Chiu, and Chih-Ching Huang. Identification of microalgae by laser desorption/ionization mass spectrometry coupled with multiple nanomatrices. Marine Biotechnology, 18(2):283–292, April 2016. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.com/

article/10.1007/s10126-016-9685-9.

Eythorsdottir:2016:AAM

[1320] Arnheidur Eythorsdottir, Sesselja Omarsdottir, and Hjorleifur Einarsson. Antimicrobial activity of marine bacterial symbionts retrieved from shallow water hydrothermal vents. Marine Biotechnology, 18(3): 293–300, June 2016. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-016-9695-7.

Chen:2016:PCC

[1321] Yin Chen, Wen-Jun Mao, Meng-Xia Yan, Xue Liu, Shu-Yao Wang, Zheng Xia, Bo Xiao, Su-Jian Cao, Bao-Qin Yang, and Jie Li. Purification, chemical characterization, and bioactivity of an extracellular polysaccharide produced by the marine sponge endogenous fungus Alternaria sp. SP-32. Marine Biotechnology, 18(3):301–313, June 2016. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-016-9696-6.

Martins:2016:BMC

[1322] Aline P. Martins, Nair S. Yokoya, and Pio Colepicolo. Biochemical modulation by carbon and nitrogen addition in cultures of Dictyota menstrualis (Dictyotales, Phaeophyceae) to generate oil-based bioproducts. Marine Biotechnology, 18(3):314–326, June 2016. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-016-9693-9.

Zou:2016:PAO

[1323] Shanmei Zou and Qi Li. Pay attention to the overlooked cryptic diversity in existing barcoding data: the case of mollusca with character-based DNA barcoding. Marine Biotechnology, 18(3):327–335, June 2016. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-016-9692-x.

Shi:2016:CMA

[1324] Yaohua Shi, Xing Zheng, Xin Zhan, Aimin Wang, and Zhifeng Gu. cDNA microarray analysis revealing candidate biomineralization genes of the pearl oyster, Pinctada fucata martensii. Marine Biotechnology, 18(3):336–348, June 2016. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-016-9699-3.

Shi:2016:GPW

[1325] Yingli Shi, Jianhai Xiang, Guangzhou Zhou, Tetsuzan Benny Ron, Hsin-I Tong, Wen Kang, Si Sun, and Yuanan Lu. The Pacific white shrimp βactin promoter: Functional properties and the potential application for transduction system using recombinant baculovirus. Marine Biotechnology, 18(3):349–358, June 2016. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-016-9700-1.

Romaidi:2016:BVV

[1326] Romaidi and Tatsuya Ueki. Bioaccumulation of vanadium by vanadiumresistant bacteria isolated from the intestine of Ascidia sydneiensis samea. Marine Biotechnology, 18(3):359–371, June 2016. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-016-9697-5.

Toubarro:2016:CCE

[1327] Duarte Toubarro, Analuce Gouveia, Raquel Mesquita Ribeiro, Nélson Simões, Gonçalo da Costa, Carlos Cordeiro, and Romana Santos. Cloning, characterization, and expression levels of the Nectin gene from the tube feet of the sea urchin Paracentrotus lividus. Marine Biotechnology, 18(3):372–383, June 2016. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-016-9698-4.

Karakostis:2016:CAT

[1328] Konstantinos Karakostis, Caterina Costa, Francesca Zito, Franz Brümmer, and Valeria Matranga. Characterization of an alpha type carbonic anhydrase from Paracentrotus lividus sea urchin embryos. Marine Biotechnology, 18(3):384–395, June 2016. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-016-9701-0.

Cristobal:2016:QGE

[1329] Héctor Antonio Cristóbal, Hugo Ramiro Poma, Carlos Mauricio Abate, and Verónica Beatriz Rajal. Quantification of the genetic expression of bgl-A, bgl, and CspA and enzymatic characterization of βglucosidases from Shewanella sp. G5 . Marine Biotechnology, 18(3): 396–408, June 2016. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-016-9702-z.

Chen:2016:DMI

[1330] Min Chen, Wei Zhang, Chang-Lun Shao, Zhen-Ming Chi, and Chang-Yun Wang. DNA methyltransferase inhibitor induced fungal biosynthetic products: Diethylene glycol phthalate ester oligomers from the marine-derived fungus Cochliobolus lunatus. Marine Biotechnology, 18 (3):409–417, June 2016. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-016-9703-y.

Brown:2016:FGA

[1331] Tyler D. Brown, Tiago S. Hori, Xi Xue, Chang Lin Ye, Derek M. Anderson, and Matthew L. Rise. Functional genomic analysis of the impact of camelina (Camelina sativa) meal on Atlantic salmon (Salmo salar) distal intestine gene expression and physiology. Marine Biotechnology, 18 (3):418–435, June 2016. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-016-9704-x.

Siddique:2016:EDT

[1332] Bhuiyan Sharmin Siddique, Shigeharu Kinoshita, Chaninya Wongkarangkana, Shuichi Asakawa, and Shugo Watabe. Evolution and distribution of teleost myomiRNAs: Functionally diversified myomiRs in teleosts. Marine Biotechnology, 18(3):436–447, June 2016. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-016-9705-9.

Hirayama:2016:EHM

[1333] Makoto Hirayama, Hiromi Shibata, Koji Imamura, Takemasa Sakaguchi, and Kanji Hori. Erratum to: High-mannose specific lectin and its recombinants from a carrageenophyta Kappaphycus alvarezii represent a potent anti-HIV activity through high-affinity binding to the viral envelope glycoprotein gp120. Marine Biotechnology, 18(3):448, June 2016. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-016-9694-8.

Dehler:2016:DEG

[1334] Carola E. Dehler, Pierre Boudinot, Samuel A. M. Martin, and Bertrand Collet. Development of an efficient genome editing method by CRISPR/ Cas9 in a fish cell line. Marine Biotechnology, 18(4):449–452, August 2016. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-016-9708-6.

Li:2016:TSC

[1335] Yangping Li, Lingling Zhang, Yan Sun, Xiaoli Ma, Jing Wang, Ruojiao Li, Meiwei Zhang, Shi Wang, Xiaoli Hu, and Zhenmin Bao. Transcriptome sequencing and comparative analysis of ovary and testis identifies potential key sex-related genes and pathways in scallop Patinopecten yessoensis. Marine Biotechnology, 18(4):453–465, August 2016. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-016-9706-8.

Kramer:2016:PRS

[1336] Annemarie Kramer, Antje Labes, and Johannes F. Imhoff. Phylogenetic relationship and secondary metabolite production of marine fungi producing the cyclodepsipeptides scopularide A and B. Marine Biotechnology, 18(4):466–474, August 2016. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-016-9707-7.

Ha:2016:SEB

[1337] Byung Geun Ha, Jung-Eun Park, and Yun Hee Shon. Stimulatory effect of balanced deep-sea water containing chitosan oligosaccharides on glucose uptake in C_{2}C_{12} myotubes. Marine Biotechnology, 18(4): 475–484, August 2016. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-016-9709-5.

Mu:2016:LNC

[1338] Chuang Mu, Ruijia Wang, Tianqi Li, Yuqiang Li, Meilin Tian, Wenqian Jiao, Xiaoting Huang, Lingling Zhang, Xiaoli Hu, Shi Wang, and Zhenmin Bao. Long non-coding RNAs (lncRNAs) of sea cucumber: Large-scale prediction, expression profiling, non-coding network construction, and lncRNA-microRNA-Gene interaction analysis of lncRNAs in Apostichopus japonicus and Holothuria glaberrima during LPS challenge and radial organ complex regeneration. Marine Biotechnology, 18 (4):485–499, August 2016. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-016-9711-y.

Jiang:2016:PPG

[1339] Hong Jiang, Yan Ma, Zhe Chi, Guang-Lei Liu, and Zhen-Ming Chi. Production, purification, and gene cloning of a β-fructofuranosidase with a high inulin-hydrolyzing activity produced by a novel yeast Aureobasidium sp. P6 isolated from a mangrove ecosystem. Marine Biotechnology, 18 (4):500–510, August 2016. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-016-9712-x.

Birolli:2016:BPP

[1340] Willian G. Birolli, Natália Alvarenga, Mirna H. R. Seleghim, and André L. M. Porto. Biodegradation of the pyrethroid pesticide esfenvalerate by marine-derived fungi. Marine Biotechnology, 18(4):511–520, August 2016. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-016-9710-z.

Huang:2016:TAC

[1341] Songqian Huang, Xiaojuan Cao, and Xianchang Tian. Transcriptomic analysis of compromise between air-breathing and nutrient uptake of posterior intestine in loach (Misgurnus anguillicaudatus), an air-breathing fish. Marine Biotechnology, 18(4):521–533, August 2016. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-016-9713-9.

Zhao:2016:GCE

[1342] Xuelin Zhao, Hong Yu, Lingfeng Kong, and Qi Li. Gene co-expression network analysis reveals the correlation patterns among genes in euryhaline adaptation of Crassostrea gigas. Marine Biotechnology, 18(5): 535–544, October 2016. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-016-9715-7.

Fan:2016:CET

[1343] Zhaofei Fan, Zhihao Wu, Lijuan Wang, Yuxia Zou, Peijun Zhang, and Feng You. Characterization of embryo transcriptome of gynogenetic olive flounder Paralichthys olivaceus. Marine Biotechnology, 18(5): 545–553, October 2016. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-016-9716-6.

Levy:2016:SIH

[1344] Tom Levy, Ohad Rosen, Brit Eilam, Dudu Azulay, Eliahu D. Aflalo, Rivka Manor, Assaf Shechter, and Amir Sagi. A single injection of hypertrophied androgenic gland cells produces all-female aquaculture. Marine Biotechnology, 18(5):554–563, October 2016. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-016-9717-5.

Jung:2016:CSC

[1345] Woongsic Jung, Eun Jae Kim, Se Jong Han, Han-Gu Choi, and Sanghee Kim. Characterization of Stearoyl–CoA desaturases from a psychrophilic Antarctic copepod, Tigriopus kingsejongensis. Marine Biotechnology, 18 (5):564–574, October 2016. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-016-9714-8.

Dong:2016:GSU

[1346] Linsong Dong, Shijun Xiao, Junwei Chen, Liang Wan, and Zhiyong Wang. Genomic selection using extreme phenotypes and pre-selection of SNPs in large yellow croaker (Larimichthys crocea). Marine Biotechnology, 18(5):575–583, October 2016. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-016-9718-4.

Ma:2016:IMG

[1347] Hao Ma, Gregory M. Weber, Hairong Wei, and Jianbo Yao. Identification of mitochondrial genome-encoded small RNAs related to egg deterioration caused by postovulatory aging in rainbow trout. Marine Biotechnology, 18(5):584–597, October 2016. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-016-9719-3.

Huang:2016:GWA

[1348] Baoyu Huang, Linlin Zhang, Xueying Tang, Guofan Zhang, and Li Li. Genome-wide analysis of alternative splicing provides insights into stress adaptation of the Pacific oyster. Marine Biotechnology, 18(5):598– 609, October 2016. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-016-9720-x.

Wu:2016:IPP

[1349] Cuiling Wu, Dan Liu, Xinghao Yang, Ribang Wu, Jiang Zhang, Jiafeng Huang, and Hailun He. Improving production of protease from Pseudoalteromonas sp. CSN423 by random mutagenesis. Marine Biotechnology, 18(5):610–618, October 2016. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-016-9721-9.

daSilva:2016:PES

[1350] Ana Cláudia Rodrigues da Silva, Luciana Garcia Ferreira, Maria Eugênia Rabello Duarte, Mutue Toyota Fujii, Eladio Flores Sanchez, Miguel Daniel Noseda, and André Lopes Fuly. Protective effect of the sulfated agaran isolated from the red seaweed Laurencia aldingensis against toxic effects of the venom of the snake, Lachesis muta. Marine Biotechnology, 18(6):619–629, December 2016. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-016-9722-8.

Hieke:2016:WDD

[1351] Anne-Sophie Charlotte Hieke, Robin Brinkmeyer, Kevin M. Yeager, Kimberly Schindler, Saijin Zhang, Chen Xu, Patrick Louchouarn, and Peter H. Santschi. Widespread distribution of Dehalococcoides mccartyi in the Houston Ship Channel and Galveston Bay, Texas, sediments and the potential for reductive dechlorination of PCDD/F in an estuarine environment. Marine Biotechnology, 18(6):630–644, December 2016. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-016-9723-7.

Liang:2016:ICL

[1352] Jian Liang, Jun Xie, Jing Gao, Chao-Qun Xu, Yi Yan, Gan-Chu Jia, Liang Xiang, Li-Ping Xie, and Rong-Qing Zhang. Identification and characterization of the Lysine-rich matrix protein family in Pinctada fucata: Indicative of roles in shell formation. Marine Biotechnology, 18(6): 645–658, December 2016. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-016-9724-6.

Li:2016:MAG

[1353] Zhiyong Li, Yuezhu Wang, Jinlong Li, Fang Liu, Liming He, Ying He, and Shenyue Wang. Metagenomic analysis of genes encoding nutrient cycling pathways in the microbiota of deep-sea and shallow-water sponges. Marine Biotechnology, 18(6):659–671, December 2016. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-016-9725-5.

Dang:2017:CLL

[1354] Hoang Tran Dang, Shinya Komatsu, Hideyuki Masuda, and Keiichi Enomoto. Characterization of LuxI and LuxR protein homologs of N-acylhomoserine lactone-dependent quorum sensing system in Pseudoalteromonas sp. 520P1. Marine Biotechnology, 19(1):1–10, February 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-016-9726-4.

Li:2017:CNV

[1355] Bi Jun Li, Hong Lian Li, Zining Meng, Yong Zhang, Haoran Lin, Gen Hua Yue, and Jun Hong Xia. Copy number variations in tilapia genomes. Marine Biotechnology, 19(1):11–21, February 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-017-9733-0.

Wang:2017:CAS

[1356] Xiudan Wang, Mengqiang Wang, Zhihao Jia, Limei Qiu, Lingling Wang, Anguo Zhang, and Linsheng Song. A carbonic anhydrase serves as an important acid–base regulator in Pacific oyster Crassostrea gigas exposed to elevated CO_{2}: Implication for physiological responses of mollusk to ocean acidification. Marine Biotechnology, 19(1):22–35, February 2017. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-017-9734-z.

Chatchaiphan:2017:NTC

[1357] Satid Chatchaiphan, Prapansak Srisapoome, Jin-Hyoung Kim, Robert H. Devlin, and Uthairat Na-Nakorn. De Novo transcriptome characterization and growth-related gene expression profiling of diploid and triploid bighead catfish (Clarias macrocephalus Günther, 1864). Marine Biotechnology, 19(1):36–48, February 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-017-9730-3.

Carneiro:2017:PBC

[1358] Rômulo Farias Carneiro, Renato Cézar Farias Torres, Renata Pinheiro Chaves, Mayron Alves de Vasconcelos, Bruno Lopes de Sousa, André Castelo Rodrigues Goveia, Francisco Vassiliepe Arruda, Maria Nágila Carneiro Matos, Helena Matthews-Cascon, Valder Nogueira Freire, Edson Holanda Teixeira, Celso Shiniti Nagano, and Alexandre Holanda Sampaio. Purification, biochemical characterization, and amino acid sequence of a novel type of lectin from Aplysia dactylomela eggs with antibacterial/antibiofilm potential. Marine Biotechnology, 19 (1):49–64, February 2017. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-017-9728-x.

Park:2017:IEE

[1359] Sang-Yong Park, Eunson Hwang, Yu-Kyong Shin, Don-Gil Lee, Jung-Eun Yang, Jae-Hee Park, and Tae-Hoo Yi. Immunostimulatory effect of enzyme-modified Hizikia fusiforme in a mouse model In Vitro and ex vivo. Marine Biotechnology, 19(1):65–75, February 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-017-9727-y.

Yuan:2017:CEO

[1360] Jianbo Yuan, Xiaojun Zhang, Chengzhang Liu, Hu Duan, Fuhua Li, and Jianhai Xiang. Convergent evolution of the osmoregulation system in decapod shrimps. Marine Biotechnology, 19(1):76–88, February 2017. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-017-9729-9.

Li:2017:RSA

[1361] Yuquan Li, Shoumin Lai, Renjie Wang, Yuchao Zhao, Hao Qin, Lingxu Jiang, Na Li, Qiang Fu, and Chao Li. RNA-Seq analysis of the antioxidant status and immune response of Portunus trituberculatus following aerial exposure. Marine Biotechnology, 19(1):89–101, February 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-017-9731-2.

Zarski:2017:TPE

[1362] Daniel Żarski, Thaovi Nguyen, Aurélie Le Cam, Jérôme Montfort, Gilbert Dutto, Marie Odile Vidal, Christian Fauvel, and Julien Bobe. Transcriptomic profiling of egg quality in sea bass (Dicentrarchus labrax) sheds light on genes involved in ubiquitination and translation. Marine Biotechnology, 19(1):102–115, February 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-017-9732-1.

Imran:2017:GSM

[1363] Md. Imran, Poonam Pant, Yogini P. Shanbhag, Samir V. Sawant, and Sanjeev C. Ghadi. Genome sequence of Microbulbifer mangrovi DD-13^{T} reveals its versatility to degrade multiple polysaccharides. Marine Biotechnology, 19(1):116–124, February 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-017-9737-9.

Leelatanawit:2017:TAM

[1364] Rungnapa Leelatanawit, Umaporn Uawisetwathana, Amornpan Klanchui, Jutatip Khudet, Suwanchai Phomklad, Somjai Wongtriphop, Pikul Jiravanichpaisal, and Nitsara Karoonuthaisiri. Transcriptomic analysis of male black tiger shrimp (Penaeus monodon) after polychaete feeding to enhance testicular maturation. Marine Biotechnology, 19(2):

125–135, April 2017. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-017-9738-8.

Park:2017:PAP

[1365] Geon-Tae Park, Ryeo-Eun Go, Hae-Miru Lee, Geum-A Lee, Cho-Won Kim, Jeong-Woo Seo, Won-Kyung Hong, Kyung-Chul Choi, and Kyung-A Hwang. Potential anti-proliferative and immunomodulatory effects of marine microalgal exopolysaccharide on various human cancer cells and lymphocytes In Vitro. Marine Biotechnology, 19(2): 136–146, April 2017. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-017-9735-y.

Kong:2017:CPE

[1366] Fanna Kong, Hailong Zhao, Weixun Liu, Na Li, and Yunxiang Mao. Construction of plastid expression vector and development of genetic transformation system for the seaweed Pyropia yezoensis. Marine Biotechnology, 19(2):147–156, April 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-017-9736-x.

Hikima:2017:SDD

[1367] Jun ichi Hikima, Masahiro Ando, Hiro o Hamaguchi, Masahiro Sakai, Masashi Maita, Kazunaga Yazawa, Haruko Takeyama, and Takashi Aoki. On-site direct detection of astaxanthin from salmon fillet using Raman spectroscopy. Marine Biotechnology, 19(2):157–163, April 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-017-9739-7.

Avelelas:2017:EEN

[1368] Francisco Avelelas, Roberto Martins, Tânia Oliveira, Frederico Maia, Eliana Malheiro, Amadeu M. V. M. Soares, Susana Loureiro, and João Tedim. Efficacy and ecotoxicity of novel anti-fouling nanomaterials in target and non-target marine species. Marine Biotechnology, 19(2): 164–174, April 2017. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-017-9740-1.

Reyes:2017:NMG

[1369] Carolina Reyes, Dominik Schneider, Marko Lipka, Andrea Thürmer, Michael E. Böttcher, and Michael W. Friedrich. Nitrogen metabolism genes from temperate marine sediments. Marine Biotechnology, 19(2):

175–190, April 2017. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-017-9741-0.

Shiel:2017:ETG

[1370] Brett P. Shiel, Nathan E. Hall, Ira R. Cooke, Nicholas A. Robinson, and Jan M. Strugnell. Epipodial tentacle gene expression and predetermined resilience to summer mortality in the commercially important greenlip abalone, Haliotis laevigata. Marine Biotechnology, 19(2): 191–205, April 2017. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-017-9742-z.

Paidi:2017:CDM

[1371] Murali Krishna Paidi, Parinita Agarwal, Prashant More, and Pradeep K. Agarwal. Chemical derivatization of metabolite mass profiling of the recretohalophyte Aeluropus lagopoides revealing salt stress tolerance mechanism. Marine Biotechnology, 19(3):207–218, June 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-017-9745-9.

Fawzy:2017:FAC

[1372] Mustafa A. Fawzy. Fatty acid characterization and biodiesel production by the marine microalga Asteromonas gracilis: Statistical optimization of medium for biomass and lipid enhancement. Marine Biotechnology, 19(3):219–231, June 2017. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-017-9743-y.

Im:2017:TBI

[1373] Sungoh Im, Ha-Nul Lee, Hyun Shin Jung, Sunghwan Yang, Eun-Jeong Park, Mi Sook Hwang, Won-Joong Jeong, and Dong-Woog Choi. Transcriptome-based identification of the desiccation response genes in marine red algae Pyropia tenera (Rhodophyta) and enhancement of abiotic stress tolerance by PtDRG2 in Chlamydomonas. Marine Biotechnology, 19(3):232–245, June 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-017-9744-x.

Barcelo-Villalobos:2017:PML

[1374] Marta Barceló-Villalobos, Félix L. Figueroa, Nathalie Korbee, Félix Álvarez-Gómez, and Maria H. Abreu. Production of mycosporine-like amino acids from Gracilaria vermiculophylla (Rhodophyta) cultured through one year in an integrated multi-trophic aquaculture (IMTA) system. Marine Biotechnology, 19(3):246–254, June 2017. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-017-9746-8.

Wang:2017:GWA

[1375] Le Wang, Peng Liu, Shuqing Huang, Baoqing Ye, Elaine Chua, Zi Yi Wan, and Gen Hua Yue. Genome-wide association study identifies loci associated with resistance to viral nervous necrosis disease in Asian seabass. Marine Biotechnology, 19(3):255–265, June 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-017-9747-7.

Li:2017:IDE

[1376] Haimei Li, Bo Zhang, Sigang Fan, Baosuo Liu, Jiaqi Su, and Dahui Yu. Identification and differential expression of biomineralization genes in the mantle of pearl oyster Pinctada fucata. Marine Biotechnology, 19 (3):266–276, June 2017. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-017-9748-6.

Yu:2017:ISD

[1377] Yang Yu, Xiaojun Zhang, Jianbo Yuan, Quanchao Wang, Shihao Li, Hao Huang, Fuhua Li, and Jianhai Xiang. Identification of sex-determining loci in Pacific white shrimp Litopeneaus vannamei using linkage and association analysis. Marine Biotechnology, 19(3):277–286, June 2017. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-017-9749-5.

Nunez-Acuna:2017:FDM

[1378] Gustavo Núñez-Acuña, Camille Détrée, Cristian Gallardo-Escárate, and Ana Teresa Gonçalves. Functional diets modulate lncRNA-coding RNAs and gene interactions in the intestine of rainbow trout Oncorhynchus mykiss. Marine Biotechnology, 19(3):287–300, June 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-017-9750-z.

Zhao:2017:GWA

[1379] Liang Zhao, Yangping Li, Yajuan Li, Jiachen Yu, Huan Liao, Shuyue Wang, Jia Lv, Jun Liang, Xiaoting Huang, and Zhenmin Bao. A genomewide association study identifies the genomic region associated with shell color in Yesso scallop, Patinopecten yessoensis. Marine Biotechnology, 19(3):301–309, June 2017. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-017-9751-y.

Huang:2017:ISF

[1380] Chang-Jen Huang, Chih-Ming Chou, Huang-Wei Lien, Cheng-Ying Chu, Jhih-Yun Ho, Yimin Wu, and Chia-Hsiung Cheng. IRF9-Stat2 fusion protein as an innate immune inducer to activate Mx and Interferonstimulated gene expression in zebrafish larvae. Marine Biotechnology, 19 (3):310–319, June 2017. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-017-9752-x.

Koganti:2017:EIN

[1381] Prasanthi P. Koganti, Jian Wang, Beth Cleveland, and Jianbo Yao. 17 β-estradiol increases non–CpG methylation in exon 1 of the rainbow trout (Oncorhynchus mykiss) MyoD gene. Marine Biotechnology, 19(4): 321–327, August 2017. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-017-9756-6.

Sabu:2017:MIC

[1382] Sanyo Sabu, I. S. Bright Singh, and Valsamma Joseph. Molecular identification and comparative evaluation of tropical marine microalgae for biodiesel production. Marine Biotechnology, 19(4):328–344, August 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-017-9754-8.

Liu:2017:NDD

[1383] Helu Liu, Hui Wang, Shanya Cai, and Haibin Zhang. A novel ω3desaturase in the deep sea giant tubeworm Riftia pachyptila. Marine Biotechnology, 19(4):345–350, August 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-017-9753-9.

Chen:2017:OST

[1384] Bi-Shuang Chen, Hui Liu, Fayene Zeferino Ribeiro de Souza, and Lan Liu. Organic solvent-tolerant marine microorganisms as catalysts for kinetic resolution of cyclic β-hydroxy ketones. Marine Biotechnology, 19 (4):351–360, August 2017. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-017-9755-7.

Payton:2017:THS

[1385] Laura Payton, Mickael Perrigault, Jean-Paul Bourdineaud, Anjara Marcel, Jean-Charles Massabuau, and Damien Tran. Trojan horse strategy for non-invasive interference of Clock gene in the oyster Crassostrea gigas. Marine Biotechnology, 19(4):361–371, August 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-017-9761-9.

Wei:2017:PSP

[1386] Mei-Yan Wei, Cui-Fang Wang, Kai-Ling Wang, Pei-Yuan Qian, Chang-Yun Wang, and Chang-Lun Shao. Preparation, structure, and potent antifouling activity of sclerotioramine derivatives. Marine Biotechnology, 19 (4):372–378, August 2017. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-017-9760-x.

Zhong:2017:GWA

[1387] Xiaoxiao Zhong, Xiaozhu Wang, Tao Zhou, Yulin Jin, Suxu Tan, Chen Jiang, Xin Geng, Ning Li, Huitong Shi, Qifan Zeng, Yujia Yang, Zihao Yuan, Lisui Bao, Shikai Liu, Changxu Tian, Eric Peatman, Qi Li, and Zhanjiang Liu. Genome-wide association study reveals multiple novel QTL associated with low oxygen tolerance in hybrid catfish. Marine Biotechnology, 19(4):379–390, August 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-017-9757-5.

Yaacob:2017:PCA

[1388] Eamy Nursaliza Yaacob, Jens Goethals, Aline Bajek, Kristof Dierckens, Peter Bossier, Bruno G. De Geest, and Daisy Vanrompay. Preparation and characterization of alginate microparticles containing a model protein for oral administration in gnotobiotic European sea bass (Dicentrarchus labrax) larvae. Marine Biotechnology, 19(4):391–400, August 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-017-9758-4.

Ye:2017:CLT

[1389] Lihai Ye, Ni Jiao, Xiaojun Tang, Yiyi Chen, Xiaolan Ye, Li Ren, Fangzhou Hu, Shi Wang, Ming Wen, Chun Zhang, Min Tao, and Shaojun Liu. Chimeras linked to tandem repeats and transposable elements in tetraploid hybrid fish. Marine Biotechnology, 19(4):401–409, August 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-017-9764-6.

Skriptsova:2017:NEW

[1390] Anna V. Skriptsova. Nitrogen effect on water-soluble polysaccharide accumulation in Streblonema sp. (Ectocarpales, Phaeophyceae). Marine Biotechnology, 19(4):410–419, August 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-017-9759-3.

Xu:2017:EHG

[1391] Lan Xu, Qi Li, Hong Yu, and Lingfeng Kong. Estimates of heritability for growth and shell color traits and their genetic correlations in the black shell strain of Pacific oyster Crassostrea gigas. Marine Biotechnology, 19 (5):421–429, October 2017. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-017-9772-6.

Matsumoto:2017:CAI

[1392] Hiroko Matsumoto, Shoko Fujiwara, Hisako Miyagi, Nobuhiro Nakamura, Yasuhiro Shiga, Toshihiro Ohta, and Mikio Tsuzuki. Carbonic anhydrase inhibitors induce developmental toxicity during zebrafish embryogenesis, especially in the inner ear. Marine Biotechnology, 19(5): 430–440, October 2017. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-017-9763-7.

Li:2017:GWQ

[1393] Hong Lian Li, Xiao Hui Gu, Bi Jun Li, Chao Hao Chen, Hao Ran Lin, and Jun Hong Xia. Genome-wide QTL analysis identified significant associations between hypoxia tolerance and mutations in the GPR132 and ABCG4 genes in Nile tilapia. Marine Biotechnology, 19(5): 441–453, October 2017. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-017-9762-8.

Versluis:2017:RPU

[1394] Dennis Versluis, Kyle McPherson, Mark W. J. van Passel, Hauke Smidt, and Detmer Sipkema. Recovery of previously uncultured bacterial genera from three Mediterranean sponges. Marine Biotechnology, 19(5): 454–468, October 2017. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-017-9766-4. See correction [1455].

Pan:2017:IAS

[1395] Chengqian Pan, Yutong Shi, Bibi Nazia Auckloo, Syed Shams ul Hassan, Najeeb Akhter, Kuiwu Wang, Ying Ye, Chen-Tung Arthur Chen, Xinyi Tao, and Bin Wu. Isolation and antibiotic screening of fungi from a hydrothermal vent site and characterization of secondary metabolites from a Penicillium isolate. Marine Biotechnology, 19(5):469–479, October 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-017-9765-5.

deFreitas:2017:DGS

[1396] Robert Cardoso de Freitas, Estácio Jussie Odisi, Chiaki Kato, Marcus Adonai Castro da Silva, and André Oliveira de Souza Lima. Draft genome sequence of the deep-sea bacterium Moritella sp. JT01 and identification of biotechnologically relevant genes. Marine Biotechnology, 19(5): 480–487, October 2017. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-017-9767-3.

Niu:2017:CFH

[1397] Donghong Niu, Yunchao Du, Ze Wang, Shumei Xie, Haideng Nguyen, Zhiguo Dong, Heding Shen, and Jiale Li. Construction of the first highdensity genetic linkage map and analysis of quantitative trait loci for growth-related traits in Sinonovacula constricta. Marine Biotechnology, 19(5):488–496, October 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-017-9768-2.

Liu:2017:LMM

[1398] Yang Liu, Min Wei, Hua Guo, Changwei Shao, Liang Meng, Wenteng Xu, Na Wang, Lei Wang, Deborah M. Power, Jilun Hou, Shahid Mahboob, Zhongkai Cui, Yingming Yang, Yangzhen Li, Fazhen Zhao, and Songlin Chen. Locus mapping, molecular cloning, and expression analysis of rps6kb2 , a novel metamorphosis-related gene in Chinese tongue sole (Cynoglossus semilaevis). Marine Biotechnology, 19(5): 497–516, October 2017. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-017-9769-1.

Wang:2017:QMR

[1399] Le Wang, Bin Bai, Shuqing Huang, Peng Liu, Zi Yi Wan, Baoqing Ye, Jinlu Wu, and Gen Hua Yue. QTL mapping for resistance to iridovirus in Asian seabass using genotyping-by-sequencing. Marine Biotechnology, 19 (5):517–527, October 2017. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-017-9770-8.

Zhu:2017:EIT

[1400] Xiangping Zhu, Zhengmei Lin, Zhihao Wu, Jiandong Li, and Feng You. Effect of initiation time of hydrostatic pressure shock on chromosome set doubling of tetraploidization in turbot Scophthalmus maximus. Marine Biotechnology, 19(5):528–540, October 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-017-9771-7.

Song:2017:BCR

[1401] Kai Song, Li Li, and Guofan Zhang. Bias and correction in RNAseq data for marine species. Marine Biotechnology, 19(5):541–550, October 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-017-9773-5.

Li:2017:DGE

[1402] Hong Lian Li, Hao Ran Lin, and Jun Hong Xia. Differential gene expression profiles and alternative isoform regulations in gill of Nile tilapia in response to acute hypoxia. Marine Biotechnology, 19(6):551– 562, December 2017. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-017-9774-4.

Noble:2017:ETZ

[1403] Sandra Noble, Vishal Saxena, Marc Ekker, and Robert Devlin. Expression of Thiaminase in zebrafish (Danio rerio) is lethal and has implications for use as a biocontainment strategy in aquaculture and invasive species. Marine Biotechnology, 19(6):563–569, December 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-017-9776-2.

Geng:2017:GWA

[1404] Xin Geng, Shikai Liu, Zihao Yuan, Yanliang Jiang, Degui Zhi, and Zhanjiang Liu. A genome-wide association study reveals that genes with functions for bone development are associated with body conformation in catfish. Marine Biotechnology, 19(6):570–578, December 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-017-9775-3.

Hamasaki:2017:PTP

[1405] Masaomi Hamasaki, Yutaka Takeuchi, Ryosuke Yazawa, Souta Yoshikawa, Kazushi Kadomura, Toshiyuki Yamada, Kadoo Miyaki, Kiyoshi Kikuchi, and Goro Yoshizaki. Production of tiger puffer Takifugu rubripes offspring from triploid grass puffer Takifugu niphobles parents. Marine Biotechnology, 19(6):579–591, December 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-017-9777-1.

Gomaa:2017:SOC

[1406] Mohamed Gomaa, Awatief F. Hifney, Mustafa A. Fawzy, and Khayria M. Abdel-Gawad. Statistical optimization of culture variables for enhancing agarase production by Dendryphiella arenaria utilizing Palisada perforata (Rhodophyta) and enzymatic saccharification of the macroalgal biomass. Marine Biotechnology, 19(6):592–600, December 2017. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-017-9778-0.

Sawayama:2017:IQT

[1407] Eitaro Sawayama, Shiho Tanizawa, Shin-Ichi Kitamura, Kei Nakayama, Kohei Ohta, Akiyuki Ozaki, and Motohiro Takagi. Identification of quantitative trait loci for resistance to RSIVD in red sea bream (Pagrus major). Marine Biotechnology, 19(6):601–613, December 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-017-9779-z.

Giannetto:2017:EOA

[1408] Alessia Giannetto, Maria Maisano, Tiziana Cappello, Sabrina Oliva, Vincenzo Parrino, Antonino Natalotto, Giuseppe De Marco, and Salvatore Fasulo. Effects of oxygen availability on oxidative stress biomarkers in the Mediterranean mussel Mytilus galloprovincialis. Marine Biotechnology, 19(6):614–626, December 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-017-9780-6.

Liu:2017:PGD

[1409] Feng Liu, Zhe Jin, Yu Wang, Yuping Bi, and James T. Melton. Plastid genome of Dictyopteris divaricata (Dictyotales, Phaeophyceae): Understanding the evolution of plastid genomes in brown algae. Marine Biotechnology, 19(6):627–637, December 2017. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-017-9781-5.

Jeong:2018:IBE

[1410] Da Woon Jeong, Jeong Eun Hyeon, Young-Chul Joo, Sang Kyu Shin, and Sung Ok Han. Integration of bacterial expansin on agarolytic complexes to enhance the degrading activity of red algae by control of gelling properties. Marine Biotechnology, 20(1):1–9, February 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-017-9782-4.

Lin:2018:MQO

[1411] Grace Lin, Le Wang, Si Te Ngoh, Lianghui Ji, Laszlo Orbán, and Gen Hua Yue. Mapping QTL for omega-3 content in hybrid saline tilapia. Marine Biotechnology, 20(1):10–19, February 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-017-9783-3.

Yang:2018:CTA

[1412] Xiaolong Yang, Mhd Ikhwanuddin, Xincang Li, Fan Lin, Qingyang Wu, Yueling Zhang, Cuihong You, Wenhua Liu, Yinwei Cheng, Xi Shi, Shuqi Wang, and Hongyu Ma. Comparative transcriptome analysis provides insights into differentially expressed genes and long non-coding RNAs between ovary and testis of the mud crab (Scylla paramamosain). Marine Biotechnology, 20(1):20–34, February 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-017-9784-2.

Jia:2018:IIC

[1413] Kuntong Jia, Yongming Yuan, Wei Liu, Lan Liu, Qiwei Qin, and Meisheng Yi. Identification of inhibitory compounds against Singapore grouper iridovirus infection by cell viability-based screening assay and droplet digital PCR. Marine Biotechnology, 20(1):35–44, February 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-017-9785-1.

Zhao:2018:PGR

[1414] Yunfeng Zhao, Wenzhu Peng, Huayang Guo, Baohua Chen, Zhixiong Zhou, Jian Xu, Dianchang Zhang, and Peng Xu. Population genomics reveals genetic divergence and adaptive differentiation of Chinese sea bass (Lateolabrax maculatus). Marine Biotechnology, 20(1): 45–59, February 2018. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-017-9786-0.

Abdelhamid:2018:PAE

[1415] Amal Abdelhamid, Meriem Jouini, Haifa Bel Haj Amor, Zeineb Mzoughi, Mehdi Dridi, Rafik Ben Said, and Abderrahman Bouraoui. Phytochemical analysis and evaluation of the antioxidant, anti-inflammatory, and antinociceptive potential of phlorotannin-rich fractions from three Mediterranean brown seaweeds. Marine Biotechnology, 20(1):60–74, February 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-

2236 (electronic). URL https://link.springer.com/article/10.
1007/s10126-017-9787-z.

Li:2018:CEB

[1416] Shangyong Li, Linna Wang, Xuehong Chen, Wenwen Zhao, Mi Sun, and Yantao Han. Cloning, expression, and biochemical characterization of two new oligoalginate lyases with synergistic degradation capability. Marine Biotechnology, 20(1):75–86, February 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-017-9788-y.

Delisle:2018:VDA

[1417] Lizenn Delisle, Marine Fuhrmann, Claudie Quéré, Marianna Pauletto, Vianney Pichereau, Fabrice Pernet, and Charlotte Corporeau. The voltage-dependent anion channel (VDAC) of Pacific oysters Crassostrea gigas is upaccumulated during infection by the ostreid herpesvirus-1 (OsHV-1): an indicator of the Warburg Effect. Marine Biotechnology, 20 (1):87–97, February 2018. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-017-9789-x.

Gu:2018:IMQ

[1418] Xiao Hui Gu, Dan Li Jiang, Yan Huang, Bi Jun Li, Chao Hao Chen, Hao Ran Lin, and Jun Hong Xia. Identifying a major QTL associated with salinity tolerance in Nile tilapia using QTL-Seq. Marine Biotechnology, 20(1):98–107, February 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-017-9790-4.

Badary:2018:GPM

[1419] Amr Badary, Shouhei Takamatsu, Mitsuharu Nakajima, Stefano Ferri, Peter Lindblad, and Koji Sode. Glycogen production in marine cyanobacterial strain Synechococcus sp. NKBG 15041c. Marine Biotechnology, 20 (2):109–117, April 2018. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-017-9792-2.

Zamora-Sillero:2018:PFP

[1420] Juan Zamora-Sillero, Adem Gharsallaoui, and Carlos Prentice. Peptides from fish by-product protein hydrolysates and its functional properties: an overview. Marine Biotechnology, 20(2):118–130, April 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-018-9799-3.

Feng:2018:FTI

[1421] Guofang Feng, Wei Sun, Fengli Zhang, Sandi Orlić, and Zhiyong Li. Functional transcripts indicate phylogenetically diverse active ammoniascavenging microbiota in sympatric sponges. Marine Biotechnology, 20 (2):131–143, April 2018. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-018-9797-5.

Ventura:2018:CPC

[1422] P. Ventura, G. Toullec, C. Fricano, L. Chapron, V. Meunier, E. Röttinger, P. Furla, and S. Barnay-Verdier. Cnidarian primary cell culture as a tool to investigate the effect of thermal stress at cellular level. Marine Biotechnology, 20(2):144–154, April 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-017-9791-3.

Ohmori:2018:NIN

[1423] Fumito Ohmori, Shigeharu Kinoshita, Daisuke Funabara, Hiroki Koyama, Kiyohito Nagai, Kaoru Maeyama, Kikuhiko Okamoto, Shuichi Asakawa, and Shugo Watabe. Novel isoforms of N16 and N19 families implicated for the nacreous layer formation in the pearl oyster Pinctada fucata. Marine Biotechnology, 20(2):155–167, April 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-017-9793-1.

Cai:2018:ZES

[1424] Mengxin Cai, Yufeng Si, Jianshe Zhang, Zhenjun Tian, and Shaojun Du. Zebrafish embryonic slow muscle is a rapid system for genetic analysis of sarcomere organization by CRISPR/Cas9, but not NgAgo. Marine Biotechnology, 20(2):168–181, April 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-018-9794-8.

Okino:2018:VER

[1425] Nozomu Okino, Hiroyoshi Wakisaka, Yohei Ishibashi, and Makoto Ito. Visualization of endoplasmic reticulum and mitochondria in Aurantiochytrium limacinum by the expression of EGFP with cell organellespecific targeting/retaining signals. Marine Biotechnology, 20(2):182– 192, April 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-

2236 (electronic). URL https://link.springer.com/article/10.
1007/s10126-018-9795-7.

Wang:2018:CTA

[1426] Zhicheng Wang, Jun Cui, Jian Song, Haoze Wang, Kailun Gao, Xuemei Qiu, Meng Gou, Xin Li, Ziwen Hu, Xiuli Wang, and Yaqing Chang. Comparative transcriptome analysis reveals growth-related genes in juvenile Chinese sea cucumber, Russian sea cucumber, and their hybrids. Marine Biotechnology, 20(2):193–205, April 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-018-9796-6.

Yue:2018:GTA

[1427] Chenyang Yue, Qi Li, and Hong Yu. Gonad transcriptome analysis of the Pacific oyster Crassostrea gigas identifies potential genes regulating the sex determination and differentiation process. Marine Biotechnology, 20(2):206–219, April 2018. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-018-9798-4.

Shi:2018:PTA

[1428] Yu Shi, Wenguang Liu, and Maoxian He. Proteome and transcriptome analysis of ovary, intersex gonads, and testis reveals potential key sex reversal/differentiation genes and mechanism in scallop Chlamys nobilis. Marine Biotechnology, 20(2):220–245, April 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-018-9800-1.

Ye:2018:CST

[1429] Hua Ye, Shijun Xiao, Xiaoqing Wang, Zhiyong Wang, Zhengshi Zhang, Chengke Zhu, Bingjie Hu, Changhuan Lv, Shuming Zheng, and Hui Luo. Characterization of spleen transcriptome of Schizothorax prenanti during Aeromonas hydrophila infection. Marine Biotechnology, 20(2): 246–256, April 2018. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-018-9801-0.

Moodie:2018:DBE

[1430] Lindon W. K. Moodie, Gunnar Cervin, Rozenn Trepos, Christophe Labriere, Claire Hellio, Henrik Pavia, and Johan Svenson. Design and biological evaluation of antifouling dihydrostilbene oxime hybrids. Marine Biotechnology, 20(2):257–267, April 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-018-9802-z.

Kumar:2018:BCB

[1431] Manish Kumar, Amandeep Brar, V. Vivekanand, and Nidhi Pareek. Bioconversion of chitin to bioactive chitooligosaccharides: Amelioration and coastal pollution reduction by microbial resources. Marine Biotechnology, 20(3):269–281, June 2018. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-018-9812-x.

Punitha:2018:ECV

[1432] Thillai Punitha, Siew-Moi Phang, Joon Ching Juan, and John Beardall. Environmental control of vanadium haloperoxidases and halocarbon emissions in macroalgae. Marine Biotechnology, 20(3):282–303, June 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-018-9820-x.

Sakurada:2018:IAP

[1433] Shunto Sakurada, Shoko Fujiwara, Michio Suzuki, Toshihiro Kogure, Tatsuya Uchida, Tomonari Umemura, and Mikio Tsuzuki. Involvement of acidic polysaccharide Ph–PS-2 and protein in initiation of coccolith mineralization, as demonstrated by In Vitro calcification on the base plate. Marine Biotechnology, 20(3):304–312, June 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-018-9818-4.

Lee:2018:ABP

[1434] Don-Gil Lee, Yu-Kyong Shin, Jae-Hee Park, Sang-Yong Park, Eunson Hwang, Jung-Eun Yang, Hae Jo, Ki-Young Kim, Gafurjon T. Mavlonov, and Tae-Hoo Yi. Alveolar bone protective effect of hiziki extracts on the progression of periodontitis. Marine Biotechnology, 20 (3):313–323, June 2018. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-018-9814-8.

Li:2018:RTS

[1435] Hanbo Li, Baofeng Su, Guyu Qin, Zhi Ye, Ahmed Elaswad, Ahmed Alsaqufi, Dayan A. Perera, Zhenkui Qin, Ramji Odin, Khoi Vo, David Drescher, Dalton Robinson, Sheng Dong, Dan Zhang, Mei Shang, Nermeen Abass, Sanjay K. Das, Max Bangs, and Rex A. Dunham. Repressible transgenic sterilization in channel catfish, Ictalurus punctatus, by knockdown of primordial germ cell genes with copper-sensitive constructs. Marine Biotechnology, 20(3):324–342, June 2018. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-018-9819-3.

Sansone:2018:MDA

[1436] Clementina Sansone, Genoveffa Nuzzo, Christian Galasso, Raffaella Casotti, Angelo Fontana, Giovanna Romano, and Adrianna Ianora. The marine dinoflagellate Alexandrium andersoni induces cell death in lung and colorectal tumor cell lines. Marine Biotechnology, 20(3):343–352, June 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-018-9817-5.

Loh:2018:BCC

[1437] Jiun Yan Loh, Gemma L. Kay, and Adeline Su Yien Ting. Bioencapsulation and colonization characteristics of Lactococcus lactis subsp. lactis CF4MRS in Artemia franciscana: a biological approach for the control of edwardsiellosis in larviculture. Marine Biotechnology, 20(3): 353–362, June 2018. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-018-9813-9.

Palaveniene:2018:CMD

[1438] Alisa Palaveniene, Volodymyr Harkavenko, Vitalina Kharchenko, Povilas Daugela, Mindaugas Pranskunas, Gintaras Juodzbalys, Nataliya Babenko, and Jolanta Liesiene. Cuttlebone as a marine-derived material for preparing bone grafts. Marine Biotechnology, 20(3):363–374, June 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-018-9816-6.

Murugan:2018:BLM

[1439] Vidhyapriya Murugan and Krishnan Sankaran. Bacterial lipid modification of ICP11 and a new ELISA system applicable for WSSV infection detection. Marine Biotechnology, 20(3):375–384, June 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-018-9815-7.

Thongda:2018:DSP

[1440] Wilawan Thongda, Honggang Zhao, Dongdong Zhang, Lauren N. Jescovitch, Ming Liu, Ximing Guo, Meagan Schrandt, Sean P. Powers, and Eric Peatman. Development of SNP panels as a new tool to assess the genetic diversity, population structure, and parentage analysis of the Eastern oyster (Crassostrea virginica). Marine Biotechnology, 20 (3):385–395, June 2018. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-018-9803-y.

Nguyen:2018:MSI

[1441] Thao V. Nguyen, Andrea C. Alfaro, Tim Young, Sridevi Ravi, and Fabrice Merien. Metabolomics study of immune responses of New Zealand Greenshell^{T M} mussels (Perna canaliculus) infected with pathogenic Vibrio sp. Marine Biotechnology, 20(3):396–409, June 2018. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-018-9804-x.

Lee:2018:IDE

[1442] Eun Jeong Lee, Ok Kyung Lee, and Eun Yeol Lee. Identification of 4-deoxy-L-etychro-hexoseulose uronic acid reductases in an alginolytic bacterium Vibrio splendidus and their uses for L-lactate production in an Escherichia coli cell-free system. Marine Biotechnology, 20(3): 410–423, June 2018. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-018-9805-9.

Wei:2018:MDA

[1443] Lei Wei, Fei Xu, Yuzhi Wang, Zhongqiang Cai, Wenchao Yu, Cheng He, Qiuyun Jiang, Xiqiang Xu, Wen Guo, and Xiaotong Wang. The molecular differentiation of anatomically paired left and right mantles of the Pacific oyster Crassostrea gigas. Marine Biotechnology, 20(4): 425–435, August 2018. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-018-9806-8.

Bouyoucef:2018:REM

[1444] Mouloud Bouyoucef, Rodolphe Rakic, Tangni Gómez-Leduc, Thomas Latire, Frédéric Marin, Sylvain Leclercq, Franck Carreiras, Antoine Serpentini, Jean-Marc Lebel, Philippe Galéra, and Florence Legendre. Regulation of extracellular matrix synthesis by shell extracts from the marine bivalve Pecten maximus in human articular chondrocytes — application for cartilage engineering. Marine Biotechnology, 20(4): 436–450, August 2018. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-018-9807-7.

Fu:2018:LMM

[1445] Jingqiang Fu, Minghui Shen, Yawei Shen, Wengang Lü, Miaoqin Huang, Xuan Luo, Jinjin Yu, Caihuan Ke, and Weiwei You. LC–MS/MS-based metabolome analysis of biochemical pathways altered by food limitation in larvae of ivory shell, Babylonia areolata. Marine Biotechnology, 20 (4):451–466, August 2018. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-018-9808-6.

Yu:2018:GTA

[1446] Lingyun Yu, Dongdong Xu, Huan Ye, Huamei Yue, Shioh Ooka, Hidehiro Kondo, Ryosuke Yazawa, and Yutaka Takeuchi. Gonadal transcriptome analysis of Pacific abalone Haliotis discus discus: Identification of genes involved in germ cell development. Marine Biotechnology, 20(4): 467–480, August 2018. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-018-9809-5.

Sawayama:2018:ICM

[1447] Eitaro Sawayama, Daiki Noguchi, Kei Nakayama, and Motohiro Takagi. Identification, characterization, and mapping of a novel SNP associated with body color transparency in juvenile red sea bream (Pagrus major). Marine Biotechnology, 20(4):481–489, August 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-018-9810-z.

Blay:2018:CPS

[1448] Carole Blay, Serge Planes, and Chin-Long Ky. Cultured pearl surface quality profiling by the shell matrix protein gene expression in the biomineralised pearl sac tissue of Pinctada margaritifera. Marine Biotechnology, 20(4):490–501, August 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-018-9811-y.

Corral:2018:ISP

[1449] Paulina Corral, Fortunato Palma Esposito, Pietro Tedesco, Angela Falco, Emiliana Tortorella, Luciana Tartaglione, Carmen Festa, Maria Valeria D’Auria, Giorgio Gnavi, Giovanna Cristina Varese, and Donatella de Pascale. Identification of a sorbicillinoid-producing Aspergillus strain with antimicrobial activity against Staphylococcus aureus: a new polyextremophilic marine fungus from Barents Sea. Marine Biotechnology, 20 (4):502–511, August 2018. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-018-9821-9.

Quiroz-Guzman:2018:API

[1450] Eduardo Quiroz-Guzmán, Ricardo Vázquez-Juárez, Antonio Luna-González, José L. Balcázar, Diana R. Barajas-Sandoval, and Sergio F. Martínez-Díaz. Administration of probiotics improves the brine shrimp production and prevents detrimental effects of pathogenic Vibrio species. Marine Biotechnology, 20(4):512–519, August 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-018-9822-8.

Sasuga:2018:DRS

[1451] Keiji Sasuga, Tomoya Yamanashi, Shigeru Nakayama, Syuetsu Ono, and Koji Mikami. Discolored red seaweed Pyropia yezoensis with low commercial value is a novel resource for production of agar polysaccharides. Marine Biotechnology, 20(4):520–530, August 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-018-9823-7.

Taylor:2018:SSI

[1452] A. Taylor, D. Mills, T. Wang, N. Ntalamagka, S. F. Cummins, and A. Elizur. A sperm spawn-inducing pheromone in the silver lip pearl oyster (Pinctada maxima). Marine Biotechnology, 20(4):531–541, August 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-018-9824-6.

Motone:2018:PCL

[1453] Keisuke Motone, Toshiyuki Takagi, Shunsuke Aburaya, Wataru Aoki, Natsuko Miura, Hiroyoshi Minakuchi, Haruko Takeyama, Yukio Nagasaki, Chuya Shinzato, and Mitsuyoshi Ueda. Protection of coral larvae from thermally induced oxidative stress by redox nanoparticles. Marine Biotechnology, 20(4):542–548, August 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-018-9825-5.

Honda:2018:LBT

[1454] Takashi Honda, Daichi Morimoto, Yoshihiko Sako, and Takashi Yoshida. LexA binds to transcription regulatory site of cell division gene ftsZ in toxic cyanobacterium Microcystis aeruginosa. Marine Biotechnology, 20 (4):549–556, August 2018. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-018-9826-4.

Versluis:2018:CRP

[1455] Dennis Versluis, Kyle McPherson, Mark W. J. van Passel, Hauke Smidt, and Detmer Sipkema. Correction to: Recovery of previously uncultured bacterial genera from three Mediterranean sponges. Marine Biotechnology, 20(4):557, August 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.com/

article/10.1007/s10126-018-9829-1. See [1394].

Liu:2018:GSU

[1456] Yang Liu, Sheng Lu, Feng Liu, Changwei Shao, Qian Zhou, Na Wang, Yangzhen Li, Yingming Yang, Yingping Zhang, Hejun Sun, Weiwei Zheng, and Songlin Chen. Genomic selection using BayesCπ and GBLUP for resistance against Edwardsiella tarda in Japanese flounder (Paralichthys olivaceus). Marine Biotechnology, 20(5):559–565, October 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-018-9839-z.

Li:2018:CCT

[1457] Da-Wei Li, Wei-Hong Xie, Ting-Bin Hao, Jia-Xi Cai, Tian-Bao Zhou, Srinivasan Balamurugan, Wei-Dong Yang, Jie-Sheng Liu, and Hong-Ye Li. Constitutive and chloroplast targeted expression of Acetyl–CoA carboxylase in oleaginous microalgae elevates fatty acid biosynthesis. Marine Biotechnology, 20(5):566–572, October 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-018-9841-5.

Zhou:2018:GSA

[1458] Zhixiong Zhou, Lin Chen, Chuanju Dong, Wenzhu Peng, Shengnan Kong, Jinsheng Sun, Fei Pu, Baohua Chen, Jianxin Feng, and Peng Xu. Genome-scale association study of abnormal scale pattern in Yellow River carp identified previously known causative gene in European mirror carp. Marine Biotechnology, 20(5):573–583, October 2018. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-018-9827-3.

Na:2018:GPD

[1459] Yeonju Na, Ha-Nul Lee, Jiwoong Wi, Won-Joong Jeong, and Dong-Woog Choi. PtDRG1, a desiccation response gene from Pyropia tenera (Rhodophyta), exhibits chaperone function and enhances abiotic stress tolerance. Marine Biotechnology, 20(5):584–593, October 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-018-9828-2.

Shinohara:2018:CTP

[1460] Mikihiro Shinohara, Shigeharu Kinoshita, Enkong Tang, Daisuke Funabara, Makoto Kakinuma, Kaoru Maeyama, Kiyohito Nagai, Masahiko Awaji, Shugo Watabe, and Shuichi Asakawa. Comparison of two pearl sacs formed in the same recipient oyster with different genetic background involved in yellow pigmentation in Pinctada fucata. Marine Biotechnology, 20(5):594–602, October 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-018-9830-8.

Sun:2018:GFG

[1461] Fei Sun, Rongjian Tu, Jun Hong Xia, Xiao Jun Liu, and Gen Hua Yue. The FTO gene is associated with growth and omega-3/-6 ratio in Asian seabass. Marine Biotechnology, 20(5):603–610, October 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-018-9831-7.

Tsalafouta:2018:CED

[1462] A. Tsalafouta, E. Sarropoulou, N. Papandroulakis, and M. Pavlidis. Characterization and expression dynamics of key genes involved in the gilthead sea bream (Sparus aurata) cortisol stress response during early ontogeny. Marine Biotechnology, 20(5):611–622, October 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-018-9833-5.

Feng:2018:PAC

[1463] Dan Qing Feng, Jian He, Si Yu Chen, Pei Su, Cai Huan Ke, and Wei Wang. The plant alkaloid camptothecin as a novel antifouling compound for marine paints: Laboratory bioassays and field trials. Marine Biotechnology, 20(5):623–638, October 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-018-9834-4.

Thiyagarajamoorthy:2018:MBC

[1464] Dhinesh Kumar Thiyagarajamoorthy, Charli Deepak Arulanandam, Hans-Uwe Dahms, Santhosh Gokul Murugaiah, Muthukumar Krishnan, and Arthur James Rathinam. Marine bacterial compounds evaluated by in silico studies as antipsychotic drugs against schizophrenia. Marine Biotechnology, 20(5):639–653, October 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-018-9835-3.

Gan:2018:TGI

[1465] Han Ming Gan, Christopher Austin, and Stuart Linton. Transcriptomeguided identification of carbohydrate active enzymes (CAZy) from the Christmas Island red crab, Gecarcoidea natalis and a vote for the inclusion of transcriptome-derived crustacean CAZys in comparative studies. Marine Biotechnology, 20(5):654–665, October 2018. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-018-9836-2.

Song:2018:MGL

[1466] Junlin Song, Qi Li, Yong Yu, Sai Wan, Lichen Han, and Shaojun Du. Mapping genetic loci for quantitative traits of golden shell color, mineral element contents, and growth-related traits in Pacific oyster (Crassostrea gigas). Marine Biotechnology, 20(5):666–675, October 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-018-9837-1.

Song:2018:RAI

[1467] Kai Song, Li Li, and Guofan Zhang. Relationship among intron length, gene expression, and nucleotide diversity in the Pacific oyster Crassostrea gigas. Marine Biotechnology, 20(5):676–684, October 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-018-9838-0.

Boncan:2018:CRG

[1468] Delbert Almerick T. Boncan, Anne Marjorie E. David, and Arturo O. Lluisma. A CAZyme-rich genome of a taxonomically novel rhodophyteassociated carrageenolytic marine bacterium. Marine Biotechnology, 20(6):685–705, December 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.com/

article/10.1007/s10126-018-9840-6.

Hughes:2018:CSC

[1469] Melanie Hebe Hughes, Héctor Juan Prado, María Cecilia Rodríguez, Karina Michetti, Patricia Inés Leonardi, and María Cristina Matulewicz. Carrageenans from Sarcothalia crispata and Gigartina skottsbergii: Structural analysis and interpolyelectrolyte complex formation for drug controlled release. Marine Biotechnology, 20(6):706–717, December 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-018-9842-4.

Mohammadi:2018:ICN

[1470] Mohsen Mohammadi, Behrouz Taheri, Niloofar Momenzadeh, Reza Salarinia, Iraj Nabipour, Zahra Farshadzadeh, and Afshar Bargahi. Identification and characterization of novel antimicrobial peptide from Hippocampus comes by in silico and experimental studies. Marine Biotechnology, 20(6):718–728, December 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-018-9843-3.

Tan:2018:IAS

[1471] Suxu Tan, Wenwen Wang, Xiaoxiao Zhong, Changxu Tian, Donghong Niu, Lisui Bao, Tao Zhou, Yulin Jin, Yujia Yang, Zihao Yuan, Dongya Gao, Rex Dunham, and Zhanjiang Liu. Increased alternative splicing as a host response to Edwardsiella ictaluri infection in catfish. Marine Biotechnology, 20(6):729–738, December 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-018-9844-2.

Tao:2018:MAT

[1472] Min Tao, Yi Zhou, Shengnan Li, Huan Zhong, Hong Hu, Liujiao Yuan, Mi Luo, Jie Chen, Li Ren, Jing Luo, Chun Zhang, and Shaojun Liu. MicroRNA alternations in the testes related to the sterility of triploid fish. Marine Biotechnology, 20(6):739–749, December 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-018-9845-1.

Wu:2018:IST

[1473] Ping Wu, Wuying Chu, Xuanming Liu, Xinhong Guo, and Jianshe Zhang. The influence of short-term fasting on muscle growth and fiber hypotrophy regulated by the rhythmic expression of clock genes and myogenic factors in Nile tilapia. Marine Biotechnology, 20(6):750– 768, December 2018. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-018-9846-0.

Wang:2018:PGT

[1474] Yangfan Wang, Guidong Sun, Qifan Zeng, Zhihui Chen, Xiaoli Hu, Hengde Li, Shi Wang, and Zhenmin Bao. Predicting growth traits with genomic selection methods in Zhikong scallop (Chlamys farreri). Marine Biotechnology, 20(6):769–779, December 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-018-9847-z.

Harish:2018:PAA

[1475] B. S. Harish and Kiran Babu Uppuluri. Potential anticoagulant activity of trypsin inhibitor purified from an isolated marine bacterium Oceanimonas Sp. BPMS22 and its kinetics. Marine Biotechnology, 20(6): 780–791, December 2018. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-018-9848-y.

Li:2018:FPG

[1476] Zhipeng Li, Xi Chen, Jun Li, Tong Meng, Lingwei Wang, Zhen Chen, Yanyan Shi, Xueping Ling, Weiang Luo, Dafeng Liang, Yinghua Lu, Qingbiao Li, and Ning He. Functions of PKS genes in lipid synthesis of Schizochytrium sp. by gene disruption and metabolomics analysis. Marine Biotechnology, 20(6):792–802, December 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-018-9849-x.

Chen:2018:IMC

[1477] Ze Chen, Zhou Zheng, Feng-Lian Wang, Yuan-Pu Niu, Jin-Lai Miao, and Hao Li. Intracellular metabolic changes of Rhodococcus sp. LH during the biodegradation of Diesel oil. Marine Biotechnology, 20(6): 803–812, December 2018. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-018-9850-4.

Lee:2018:LMW

[1478] Jisun Lee, Seul Lee, Andriy Synytsya, Peter Capek, Chang Won Lee, Ji Won Choi, Sarang Cho, Woo Jung Kim, and Yong Il Park. Low molecular weight mannogalactofucans derived from Undaria pinnatifida induce apoptotic death of human prostate cancer cells In Vitro and In Vivo. Marine Biotechnology, 20(6):813–828, December 2018. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-018-9851-3.

Pedrosa-Gerasmio:2018:EAA

[1479] Ivane R. Pedrosa-Gerasmio, Tohru Tanaka, Asuka Sumi, Hidehiro Kondo, and Ikuo Hirono. Effects of 5-aminolevulinic acid on gene expression, immunity, and ATP levels in Pacific white shrimp, Litopenaeus vannamei. Marine Biotechnology, 20(6):829–843, December 2018. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-018-9852-2.

Miura:2019:RSS

[1480] Natsuko Miura, Keisuke Motone, Toshiyuki Takagi, Shunsuke Aburaya, Sho Watanabe, Wataru Aoki, and Mitsuyoshi Ueda. Ruegeria sp. strains isolated from the reef-building coral Galaxea fascicularis inhibit growth of the temperature-dependent pathogen Vibrio coralliilyticus. Marine Biotechnology, 21(1):1–8, February 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-018-9853-1.

Chen:2019:IAL

[1481] Xuemei Chen, Yueru Chen, Xiaotong Shen, Jianwei Zuo, and Huarong Guo. The improvement and application of lentivirus-mediated gene transfer and expression system in penaeid shrimp cells. Marine Biotechnology, 21(1):9–18, February 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-018-9862-0.

Li:2019:PIT

[1482] Yuanyou Li, Ziyan Yin, Yewei Dong, Shuqi Wang, Óscar Monroig, Douglas R. Tocher, and Cuihong You. Pparγ is involved in the transcriptional regulation of liver LC–PUFA biosynthesis by targeting the Δ6Δ5 fatty acyl desaturase gene in the marine teleost Siganus canaliculatus. Marine Biotechnology, 21(1):19–29, February 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-018-9854-0.

Parisi:2019:ICM

[1483] J. R. Parisi, K. R. Fernandes, I. R. Avanzi, B. P. Dorileo, A. F. Santana, A. L. Andrade, P. R. Gabbai-Armelin, C. A. Fortulan, E. S. Trichês, R. N. Granito, and A. C. M. Renno. Incorporation of collagen from marine sponges (Spongin) into hydroxyapatite samples: Characterization and In Vitro biological evaluation. Marine Biotechnology, 21(1): 30–37, February 2019. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-018-9855-z.

Rocha:2019:CAA

[1484] Miguel Rocha, Paulo Antas, L. Filipe C. Castro, Alexandre Campos, Vítor Vasconcelos, Filipe Pereira, and Isabel Cunha. Comparative analysis of the adhesive proteins of the adult stalked goose barnacle Pollicipes pollicipes (Cirripedia: Pedunculata). Marine Biotechnology, 21 (1):38–51, February 2019. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-018-9856-y.

Lin:2019:COR

[1485] Cheng-Yung Lin, Po-Hsiang Zhang, You-Jei Chen, Chia-Lun Wu, and Huai-Jen Tsai. Conditional overexpression of rtn4al in muscle of adult zebrafish displays defects similar to human amyotrophic lateral sclerosis. Marine Biotechnology, 21(1):52–64, February 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-018-9857-x.

Gabbai-Armelin:2019:CCE

[1486] P. R. Gabbai-Armelin, H. W. Kido, M. A. Cruz, J. P. S. Prado, I. R. Avanzi, M. R. Custódio, A. C. M. Renno, and R. N. Granito. Characterization and cytotoxicity evaluation of a marine sponge biosilica. Marine Biotechnology, 21(1):65–75, February 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-018-9858-9.

Aung:2019:PGC

[1487] Thu Aung, Hong Jiang, Cheng-Cheng Chen, Guang-Lei Liu, Zhong Hu, Zhen-Ming Chi, and Zhe Chi. Production, gene cloning, and overexpression of a laccase in the marine-derived yeast Aureobasidium melanogenum strain 11-1 and characterization of the recombinant laccase. Marine Biotechnology, 21(1):76–87, February 2019. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-018-9860-2.

Yin:2019:BMD

[1488] Qi Yin, Jinyou Liang, Weipeng Zhang, Lv Zhang, Zhang-Li Hu, Yu Zhang, and Ying Xu. Butenolide, a marine-derived broadspectrum antibiofilm agent against both Gram-positive and Gramnegative pathogenic bacteria. Marine Biotechnology, 21(1):88–98, February 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-018-9861-1.

Shene:2019:PLP

[1489] Carolina Shene, Marcelo Garcés, Daniela Vergara, Jhonatan Peña, Stéphane Claverol, Mónica Rubilar, and Allison Leyton. Production of lipids and proteome variation in a Chilean Thraustochytrium striatum strain cultured under different growth conditions. Marine Biotechnology, 21(1):99–110, February 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-018-9863-z.

Si:2019:GMG

[1490] Yufeng Si, Haishen Wen, and Shaojun Du. Genetic mutations in jamb, jamc, and myomaker revealed different roles on myoblast fusion and muscle growth. Marine Biotechnology, 21(1):111–123, February 2019. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-018-9865-x.

Dhaneesha:2019:DBM

[1491] M. Dhaneesha, O. Hasin, K. C. Sivakumar, R. Ravinesh, C. Benjamin Naman, S. Carmeli, and T. P. Sajeevan. DNA binding and molecular dynamic studies of polycyclic tetramate macrolactams (PTM) with potential anticancer activity isolated from a sponge-associated Streptomyces zhaozhouensis subsp. mycale subsp. nov. Marine Biotechnology, 21(1): 124–137, February 2019. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-018-9866-9.

Qin:2019:RGG

[1492] Qinbo Qin, Liu Cao, Yude Wang, Li Ren, Qiwen Liu, Yuwei Zhou, Chongqing Wang, Huan Qin, Chun Zhao, and Shaojun Liu. Rapid genomic and genetic changes in the first generation of autotetraploid lineages derived from distant hybridization of Carassius auratus red var. ([female sign]) × Megalobrama amblycephala ([male sign]). Marine Biotechnology, 21(2):139–149, April 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-018-9859-8. See correction

[1493].

Qin:2019:CRG

[1493] Qinbo Qin, Liu Cao, Yude Wang, Li Ren, Qiwen Liu, Yuwei Zhou, Chongqing Wang, Huan Qin, Chun Zhao, and Shaojun Liu. Correction to: Rapid genomic and genetic changes in the first generation of autotetraploid lineages derived from distant hybridization of Carassius auratus red var. ([female sign]) × Megalobrama amblycephala ([male sign]). Marine Biotechnology, 21(2):150, April 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-018-09872-9. See [1492].

Yuan:2019:GEP

[1494] Xiangcheng Yuan, Hui Huang, Weihua Zhou, Yajuan Guo, Tao Yuan, and Sheng Liu. Gene expression profiles of two coral species with varied resistance to ocean acidification. Marine Biotechnology, 21(2): 151–160, April 2019. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-018-9864-y.

Rangamaran:2019:BPB

[1495] Vijaya Raghavan Rangamaran and Venkat Kumar Shanmugam. Biocalcification by piezotolerant Bacillus sp. NIOTVJ5 isolated from deep sea sediment and its influence on the strength of concrete specimens. Marine Biotechnology, 21(2):161–170, April 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-018-9867-8.

Zhang:2019:NDS

[1496] Xiaofan Zhang, Shaochen Pang, Chengjie Liu, Houpeng Wang, Ding Ye, Zuoyan Zhu, and Yonghua Sun. A novel dietary source of EPA and DHA: Metabolic engineering of an important freshwater species — common carp by fat1-transgenesis. Marine Biotechnology, 21(2): 171–185, April 2019. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-018-9868-7.

Zhang:2019:SAM

[1497] Wanwan Zhang, Peng Jia, Wei Liu, Kuntong Jia, and Meisheng Yi. Screening for antiviral medaka haploid embryonic stem cells by genome wide mutagenesis. Marine Biotechnology, 21(2):186–195, April 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-018-09870-x.

Li:2019:DDM

[1498] Yangping Li, Lingling Zhang, Yajuan Li, Wanru Li, Zhenyi Guo, Ruojiao Li, Xiaoli Hu, Zhenmin Bao, and Shi Wang. Dynamics of DNA methylation and DNMT expression during gametogenesis and early development of scallop Patinopecten yessoensis. Marine Biotechnology, 21 (2):196–205, April 2019. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-018-09871-w.

Ma:2019:CGB

[1499] Ke-Yi Ma, Shu-Hui Yu, Yu-Xin Du, Shi-Qing Feng, Liang-Jie Qiu, Dai-Yi Ke, Mei-Zhong Luo, and Gao-Feng Qiu. Construction of a genomic bacterial artificial chromosome (BAC) library for the prawn Macrobrachium rosenbergii and initial analysis of ZW chromosome-derived BAC inserts. Marine Biotechnology, 21(2):206–216, April 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-018-09873-8.

Ye:2019:AEE

[1500] Ding Ye, Lin Zhu, Qifeng Zhang, Feng Xiong, Houpeng Wang, Xiaosi Wang, Mudan He, Zuoyan Zhu, and Yonghua Sun. Abundance of early embryonic primordial germ cells promotes zebrafish female differentiation as revealed by lifetime labeling of germline. Marine Biotechnology, 21 (2):217–228, April 2019. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-019-09874-1.

Serra:2019:MMB

[1501] Immacolata Serra, Claudia Capusoni, Francesco Molinari, Loana Musso, Luisa Pellegrino, and Concetta Compagno. Marine microorganisms for biocatalysis: Selective hydrolysis of nitriles with a salt-resistant strain of Meyerozyma guilliermondii. Marine Biotechnology, 21(2): 229–239, April 2019. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-019-09875-0.

Jin:2019:ESM

[1502] Can Jin, Jing-Ying Zhao, Xiao-Jun Liu, and Jia-Le Li. Expressions of shell matrix protein genes in the pearl sac and its correlation with pearl weight in the first 6 months of pearl formation in Hyriopsis cumingii. Marine Biotechnology, 21(2):240–249, April 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-019-09876-z.

Jiang:2019:ILQ

[1503] Dan Li Jiang, Xiao Hui Gu, Bi Jun Li, Zong Xian Zhu, Hui Qin, Zi ning Meng, Hao Ran Lin, and Jun Hong Xia. Identifying a long QTL cluster across chrLG18 associated with salt tolerance in tilapia using GWAS and QTL-seq. Marine Biotechnology, 21(2):250–261, April 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-019-09877-y.

Kong:2019:CHD

[1504] Shengnan Kong, Qiaozhen Ke, Lin Chen, Zhixiong Zhou, Fei Pu, Ji Zhao, Huaqiang Bai, Wenzhu Peng, and Peng Xu. Constructing a high-density genetic linkage map for large yellow croaker (Larimichthys crocea) and mapping resistance trait against ciliate parasite Cryptocaryon irritans. Marine Biotechnology, 21(2):262–275, April 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-019-09878-x.

Reina:2019:SMA

[1505] José Carlos Reina, Marta Torres, and Inmaculada Llamas. Stenotrophomonas maltophilia AHL-degrading strains isolated from marine invertebrate microbiota attenuate the virulence of Pectobacterium carotovorum and Vibrio coralliilyticus. Marine Biotechnology, 21(2):276–290, April 2019. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-019-09879-
w.

Iijima:2019:PEH

[1506] Mariko Iijima, Ko Yasumoto, Jun Yasumoto, Mina Yasumoto-Hirose, Nami Kuniya, Ryota Takeuchi, Masashi Nozaki, Nobuyoshi Nanba, Takashi Nakamura, Mitsuru Jimbo, and Shugo Watabe. Phosphate enrichment hampers development of juvenile Acropora digitifera coral by inhibiting skeleton formation. Marine Biotechnology, 21(2):291–300, April 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-019-09880-3.

Yu:2019:TGD

[1507] Hong Yu, Huijuan Li, Qi Li, Rui Xu, Chenyang Yue, and Shaojun Du. Targeted gene disruption in Pacific oyster based on CRISPR/ Cas9 ribonucleoprotein complexes. Marine Biotechnology, 21(3):301– 309, June 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-

2236 (electronic). URL https://link.springer.com/article/10.
1007/s10126-019-09885-y.

Zhang:2019:SPE

[1508] Bo Zhang, Na Zhao, Lei Jia, Kang Peng, Jinyuan Che, Kunming Li, Xiaoxu He, Jinsheng Sun, and Baolong Bao. Seminal plasma exosomes: Promising biomarkers for identification of male and pseudomales in Cynoglossus semilaevis. Marine Biotechnology, 21(3):310–319, June 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-019-09881-2.

Waiho:2019:GME

[1509] Khor Waiho, Hanafiah Fazhan, Yin Zhang, Yueling Zhang, Shengkang Li, Huaiping Zheng, Wenhua Liu, Mhd Ikhwanuddin, and Hongyu Ma. Gonadal microRNA expression profiles and their potential role in sex differentiation and gonadal maturation of mud crab Scylla paramamosain. Marine Biotechnology, 21(3):320–334, June 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-019-09882-1.

Wang:2019:GAI

[1510] Wenwen Wang, Suxu Tan, Jian Luo, Huitong Shi, Tao Zhou, Yujia Yang, Yulin Jin, Xiaozhu Wang, Donghong Niu, Zihao Yuan, Dongya Gao, Rex Dunham, and Zhanjiang Liu. GWAS analysis indicated importance of NF-κ B signaling pathway in host resistance against motile Aeromonas septicemia disease in catfish. Marine Biotechnology, 21 (3):335–347, June 2019. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-019-09883-0.

Wang:2019:CHD

[1511] Le Wang, Nan Xie, Yubang Shen, Baoqing Ye, Gen Hua Yue, and Xiaoyu Feng. Constructing high-density genetic maps and developing sexing markers in Northern snakehead (Channa argus). Marine Biotechnology, 21(3):348–358, June 2019. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-019-09884-z.

Ribas:2019:CGE

[1512] L. Ribas, B. Crespo, N. Sánchez-Baizán, D. Xavier, H. Kuhl, J. M. Rodríguez, N. Díaz, S. Boltañá, S. MacKenzie, F. Morán, S. Zanuy, A. Gómez, and F. Piferrer. Characterization of the European sea bass (Dicentrarchus labrax) gonadal transcriptome during sexual development. Marine Biotechnology, 21(3):359–373, June 2019. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-019-09886-x.

Yu:2019:GSG

[1513] Yang Yu, Quanchao Wang, Qian Zhang, Zheng Luo, Yue Wang, Xiaojun Zhang, Hao Huang, Jianhai Xiang, and Fuhua Li. Genome scan for genomic regions and genes associated with growth trait in Pacific white shrimp Litopeneaus vannamei. Marine Biotechnology, 21 (3):374–383, June 2019. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-019-09887-w.

Li:2019:QMR

[1514] Bi Jun Li, Zong Xian Zhu, Xiao Hui Gu, Hao Ran Lin, and Jun Hong Xia. QTL mapping for red blotches in Malaysia red tilapia (Oreochromis spp.). Marine Biotechnology, 21(3):384–395, June 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-019-09888-9.

Yang:2019:RAE

[1515] Qiqi Yang, Lyana Salim, Chuan Yan, and Zhiyuan Gong. Rapid analysis of effects of environmental toxicants on tumorigenesis and inflammation using a transgenic zebrafish model for liver cancer. Marine Biotechnology, 21(3):396–405, June 2019. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-019-09889-8. See correction [1897].

Tsuzuki:2019:PPP

[1516] Mikio Tsuzuki, Katsuhiko Okada, Haruna Isoda, Masayuki Hirano, Tetsuo Odaka, Hirotaka Saijo, Risa Aruga, Hiroki Miyauchi, and Shoko Fujiwara. Physiological properties of photoautotrophic microalgae and cyanobacteria relevant to industrial biomass production. Marine Biotechnology, 21(3):406–415, June 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-019-09890-1.

Silva:2019:CCB

[1517] Tiago R. Silva, Renata S. N. Tavares, Ramon Canela-Garayoa, Jordi Eras, Marili V. N. Rodrigues, Iramaia A. Neri-Numa, Glaucia M. Pastore, Luiz H. Rosa, José A. A. Schultz, Hosana M. Debonsi, Lorena R. G. Cordeiro, and Valeria M. Oliveira. Chemical characterization and biotechnological applicability of pigments isolated from Antarctic bacteria. Marine Biotechnology, 21(3):416–429, June 2019. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-019-09892-z.

deMatos:2019:HHH

[1518] Iara Lisboa de Matos, Marcia Nitschke, and André Luiz Meleiro Porto. Hydrogenation of halogenated 2′-hydroxychalcones by mycelia of marinederived fungus Penicillium raistrickii. Marine Biotechnology, 21(3): 430–439, June 2019. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-019-09893-y.

Huang:2019:EMM

[1519] Jianfang Huang, Weiwei You, Zhiwei Xu, Qiuning Yan, Chenggang Shi, Bin Tang, Xuan Luo, Guang Li, and Caihuan Ke. An effective microinjection method and TALEN-mediated genome editing in Pacific abalone. Marine Biotechnology, 21(4):441–447, August 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-019-09901-1.

Ma:2019:SDG

[1520] Zengyu Ma, Baozhen Qu, Shenjie Zhong, Lan Yao, Zhan Gao, and Shicui Zhang. Subtle difference generates big dissimilarity: Comparison of enzymatic activity in KL1 and KL2 domains of lancelet klotho. Marine Biotechnology, 21(4):448–462, August 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-019-09891-0.

Wu:2019:CTA

[1521] Jun-Jie Wu, Yu-Lin Zhou, Zhong-Wei Wang, Guang-Hua Li, Fang-Peng Jin, Li-Li Cui, Hai-Tao Gao, Xin-Ping Li, Li Zhou, and Jian-Fang Gui. Comparative transcriptome analysis reveals differentially expressed genes and signaling pathways between male and female red-tail catfish (Mystus wyckioides). Marine Biotechnology, 21(4):463–474, August 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-019-09894-x.

Sun:2019:MIR

[1522] Jun Jun Sun, Li Guo Zheng, Cui Ying Chen, Jin Ying Zhang, Cui Hong You, Qing Hao Zhang, Hong Yu Ma, Óscar Monroig, Douglas R. Tocher, Shu Qi Wang, and Yuan You Li. MicroRNAs involved in the regulation of LC–PUFA biosynthesis in teleosts: miR-33 enhances LC– PUFA biosynthesis in Siganus canaliculatus by targeting insig1 which in turn upregulates srebp1 . Marine Biotechnology, 21(4):475–487, August 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-019-09895-w.

Zhu:2019:DTM

[1523] Zong Xian Zhu, Dan Li Jiang, Bi Jun Li, Hui Qin, Zi Ning Meng, Hao Ran Lin, and Jun Hong Xia. Differential transcriptomic and metabolomic responses in the liver of Nile tilapia (Oreochromis niloticus) exposed to acute ammonia. Marine Biotechnology, 21(4):488–502, August 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-019-09897-8.

Yudiati:2019:ASS

[1524] Ervia Yudiati, Alim Isnansetyo, Murwantoko, Triyanto, and Christina Retna Handayani. Alginate from Sargassum siliquosum simultaneously stimulates innate immunity, upregulates immune genes, and enhances resistance of Pacific white shrimp (Litopenaeus vannamei) against white spot syndrome virus (WSSV). Marine Biotechnology, 21(4):503–514, August 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-019-09898-7.

Hu:2019:ACN

[1525] Fangzhou Hu, Jingjing Fan, Chang Wu, Ming Zhu, Yunfan Zhou, Shi Wang, Chun Zhang, Min Tao, Rurong Zhao, Chenchen Tang, Kaikun Luo, Qinbo Qin, Ming Ma, Bo Chen, Jinpu Wang, Aiguo Zhou, Liangxiong Bai, and Shaojun Liu. Analysis of chromosomal numbers, mitochondrial genome, and full-length transcriptome of Onychostoma brevibarba. Marine Biotechnology, 21(4):515–525, August 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-019-09899-6. See correction

[1552].

Feng:2019:RII

[1526] Dandan Feng, Qi Li, and Hong Yu. RNA interference by ingested dsRNA-Expressing bacteria to study shell biosynthesis and pigmentation in Crassostrea gigas. Marine Biotechnology, 21(4):526–536, August 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-019-09900-2.

Tang:2019:TAR

[1527] Lin Tang, Jiaxing Chen, Zhifeng Ye, Mi Zhao, Zining Meng, Haoran Lin, Shuisheng Li, and Yong Zhang. Transcriptomic analysis revealed the regulatory mechanisms of oocyte maturation and hydration in orangespotted grouper (Epinephelus coioides). Marine Biotechnology, 21(4): 537–549, August 2019. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-019-09902-0.

Gao:2019:DMT

[1528] Chengbin Gao, Xin Cai, Qiang Fu, Ning Yang, Lin Song, Baofeng Su, Fenghua Tan, Baining Liu, and Chao Li. Dynamics of MiRNA transcriptome in turbot (Scophthalmus maximus L.) intestine following Vibrio anguillarum infection. Marine Biotechnology, 21(4):550–564, August 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-019-09903-z.

Messina:2019:BAS

[1529] Concetta Maria Messina, Simona Manuguerra, Giuseppe Renda, and Andrea Santulli. Biotechnological applications for the sustainable use of marine by-products: In Vitro antioxidant and Pro-apoptotic effects of astaxanthin extracted with supercritical CO_{2} from Parapeneus longirostris. Marine Biotechnology, 21(4):565–576, August 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-019-09904-y.

Diaz:2019:IAA

[1530] Roberto T. Abdala Díaz, V. Casas Arrojo, M. A. Arrojo Agudo, C. Cárdenas, S. Dobretsov, and F. L. Figueroa. Immunomodulatory and antioxidant activities of sulfated polysaccharides from Laminaria ochroleuca, Porphyra umbilicalis, and Gelidium corneum. Marine Biotechnology, 21(4):577–587, August 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-019-09905-x.

Rasal:2019:LSM

[1531] Kiran D. Rasal, Mir Asif Iquebal, Sarika Jaiswal, Sangita Dixit, Manohar Vasam, Samiran Nandi, Mustafa Raza, Lakshman Sahoo, U. B. Angadi, Anil Rai, Dinesh Kumar, and Jitendra Kumar Sundaray. Liver-specific microRNA identification in farmed carp, Labeo bata (Hamilton, 1822), fed with starch diet using high-throughput sequencing. Marine Biotechnology, 21(5):589–595, October 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-019-09912-y.

Liu:2019:TGC

[1532] Xingyu Liu, Li Li, Ao Li, Yingxiang Li, Wei Wang, and Guofan Zhang. Transcriptome and gene coexpression network analyses of two wild populations provides insight into the high-salinity adaptation mechanisms of Crassostrea ariakensis. Marine Biotechnology, 21(5):596–612, October 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/

s10126-019-09896-9. See correction [1533].

Liu:2019:CTG

[1533] Xingyu Liu, Li Li, Ao Li, Yingxiang Li, Wei Wang, and Guofan Zhang. Correction to: Transcriptome and gene coexpression network analyses of two wild populations provides insight into the high-salinity adaptation mechanisms of Crassostrea ariakensis. Marine Biotechnology, 21 (5):613, October 2019. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-019-09915-9. See [1532].

Song:2019:AEP

[1534] Kai Song, Shiyong Wen, and Guofan Zhang. Adaptive evolution patterns in the Pacific oyster Crassostrea gigas. Marine Biotechnology, 21(5): 614–622, October 2019. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-019-09906-w.

Zhao:2019:TPP

[1535] Liang Zhao, Yangping Li, Jiarun Lou, Zhihui Yang, Huan Liao, Qiang Fu, Zhenyi Guo, Shanshan Lian, Xiaoli Hu, and Zhemin Bao. Transcriptomic profiling provides insights into inbreeding depression in Yesso scallop Patinopecten yessoensis. Marine Biotechnology, 21(5):623–633, October 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/

s10126-019-09907-9. See correction [1577].

Chen:2019:MEM

[1536] Xiajun Chen, Zhiyi Bai, and Jiale Li. The mantle exosome and MicroRNAs of Hyriopsis cumingii involved in nacre color formation. Marine Biotechnology, 21(5):634–642, October 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-019-09908-8.

Wang:2019:MVQ

[1537] Le Wang, Elaine Chua, Fei Sun, Zi Yi Wan, Baoqing Ye, Hongyan Pang, Yanfei Wen, and Gen Hua Yue. Mapping and validating QTL for fatty acid compositions and growth traits in Asian seabass. Marine Biotechnology, 21(5):643–654, October 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-019-09909-7.

Zhou:2019:GWA

[1538] Zhixiong Zhou, Kunhuang Han, Yidi Wu, Huaqiang Bai, Qiaozhen Ke, Fei Pu, Yilei Wang, and Peng Xu. Genome-wide association study of growth and body-shape-related traits in large yellow croaker (Larimichthys crocea) using ddRAD sequencing. Marine Biotechnology, 21(5): 655–670, October 2019. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-019-09910-0.

Zhou:2019:VTS

[1539] Li Zhou, Xueying Wang, Qinghua Liu, Shihong Xu, Haixia Zhao, Mingming Han, Yunong Wang, Zongcheng Song, and Jun Li. Visualization of turbot (Scophthalmus maximus) primordial germ cells in vivo using fluorescent protein mediated by the 3′ untranslated region of nanos3 or vasa gene. Marine Biotechnology, 21(5):671–682, October 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-019-09911-z.

Mao:2019:HED

[1540] Junxia Mao, Wenjing Zhang, Xubo Wang, Jian Song, Donghong Yin, Ying Tian, Zhenlin Hao, Bing Han, and Yaqing Chang. Histological and expression differences among different mantle regions of the Yesso scallop (Patinopecten yessoensis) provide insights into the molecular mechanisms of biomineralization and pigmentation. Marine Biotechnology, 21 (5):683–696, October 2019. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-019-09913-x.

Palma:2019:IGD

[1541] Peter Palma, Josephine Nocillado, Joshua Superio, Evelyn Grace de Jesus-Ayson, Felix Ayson, Akihiro Takemura, Ming Wei Lu, and Abigail Elizur. Induction of gonadal development in protogynous grouper with orally delivered FSH DNA. Marine Biotechnology, 21(5):697– 706, October 2019. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-019-09914-w.

Wu:2019:FGW

[1542] Lina Wu, Yang Yang, Bijun Li, Wenhua Huang, Xi Wang, Xiaochun Liu, Zining Meng, and Junhong Xia. First genome-wide association analysis for growth traits in the largest coral reef-dwelling bony fishes, the giant grouper (Epinephelus lanceolatus). Marine Biotechnology, 21(5): 707–717, October 2019. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-019-09916-8.

Lian:2019:IBC

[1543] Shanshan Lian, Jing Wang, Lingling Zhang, Qiang Xing, Naina Hu, Sinuo Liu, Xiaoting Dai, Fengmei Zhang, Xiaoli Hu, Zhenmin Bao, and Shi Wang. Integration of biochemical, cellular, and genetic indicators for understanding the aging process in a bivalve mollusk Chlamys farreri. Marine Biotechnology, 21(5):718–730, October 2019. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-019-09917-7.

Yang:2019:MCM

[1544] Zujing Yang, Xuan Li, Huan Liao, Liping Hu, Cheng Peng, Shenhai Wang, Xiaoting Huang, and Zhenmin Bao. A molecular cytogenetic map of scallop (Patinopecten yessoensis). Marine Biotechnology, 21(6): 731–742, December 2019. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-019-09918-6.

Bovio:2019:SAF

[1545] Elena Bovio, Marilyne Fauchon, Yannick Toueix, Mohamed Mehiri, Giovanna Cristina Varese, and Claire Hellio. The sponge-associated fungus Eurotium chevalieri MUT 2316 and its bioactive molecules: Potential applications in the field of antifouling. Marine Biotechnology, 21(6): 743–752, December 2019. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-019-09920-y.

Qin:2019:AMR

[1546] Qinbo Qin, Yuwei Zhou, Chongqing Wang, Minghe Zhang, Huan Qin, Chun Zhao, and Shaojun Liu. Analysis on the meiosis-related gene (Dmc1, Ph1) expression in autotriploid Carassius auratus. Marine Biotechnology, 21(6):753–761, December 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-019-09921-x.

Munar:2019:DNG

[1547] Madison Pascual Munar, Hirokazu Takahashi, and Yoshiko Okamura. Discovery of a novel gene conferring tellurite tolerance through tellurite reduction to Escherichia coli transformant in marine sediment metagenomic library. Marine Biotechnology, 21(6):762–772, December 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/

s10126-019-09922-w. See correction [1639].

Gornati:2019:MSE

[1548] Rosalba Gornati, Maria Maisano, Cristina Pirrone, Tiziana Cappello, Federica Rossi, Marina Borgese, Alessia Giannetto, Simone Cappello, Giuseppe Mancini, Giovanni Bernardini, and Salvatore Fasulo. Mesocosm system to evaluate BF–MBR efficacy in mitigating oily wastewater discharges: an integrated study on Mytilus galloprovincialis. Marine Biotechnology, 21(6):773–790, December 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-019-09923-9.

Zhang:2019:TAC

[1549] Xiaolin Zhang, Weiye Li, Minhai Liu, Xia Zhang, Xiaolong Yin, Zhijing Xu, and Jianhua Liu. Transcriptomic analysis of the cold-pretreated Larimichthys crocea showing enhanced growth fitness in cold water. Marine Biotechnology, 21(6):791–805, December 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-019-09924-8.

Liu:2019:EGS

[1550] Guijia Liu, Linsong Dong, Linlin Gu, Zhaofang Han, Wenjing Zhang, Ming Fang, and Zhiyong Wang. Evaluation of genomic selection for seven economic traits in yellow drum (Nibea albiflora). Marine Biotechnology, 21(6):806–812, December 2019. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-019-09925-7.

Yoshitake:2019:HSN

[1551] Kazutoshi Yoshitake, Tatsuki Yoshinaga, Chikaya Tanaka, Nanami Mizusawa, Md. Shaheed Reza, Atsumi Tsujimoto, Takanori Kobayashi, and Shugo Watabe. HaCeD-Seq: a novel method for reliable and easy estimation about the fish population using haplotype count from eDNA. Marine Biotechnology, 21(6):813–820, December 2019. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-019-09926-6.

Hu:2019:CAC

[1552] Fangzhou Hu, Jingjing Fan, Chang Wu, Ming Zhu, Yunfan Zhou, Shi Wang, Chun Zhang, Min Tao, Rurong Zhao, Chenchen Tang, Kaikun Luo, Qinbo Qin, Ming Ma, Bo Chen, Jinpu Wang, Aiguo Zhou, Liangxiong Bai, and Shaojun Liu. Correction to: Analysis of chromosomal numbers, mitochondrial genome, and full-length transcriptome of Onychostoma brevibarba. Marine Biotechnology, 21(6): 821, December 2019. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-019-09919-5. See [1525].

Huang:2020:FMH

[1553] Wen Huang, Chuhang Cheng, Jinshang Liu, Xin Zhang, Chunhua Ren, Xiao Jiang, Ting Chen, Kaimin Cheng, Huo Li, and Chaoqun Hu. Fine mapping of the high-pH tolerance and growth trait-related quantitative trait loci (QTLs) and identification of the candidate genes in Pacific white shrimp (Litopenaeus vannamei). Marine Biotechnology, 22 (1):1–18, February 2020. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-019-09932-8.

Kakinuma:2020:TIR

[1554] Makoto Kakinuma, Ko Yasumoto, Michio Suzuki, Chiaki Kasugai, Mirai Koide, Kayo Mitani, Kaho Shidoji, Shigeharu Kinoshita, Fumihiro Hattori, Kaoru Maeyama, Masahiko Awaji, Kiyohito Nagai, and Shugo Watabe. Trivalent iron is responsible for the yellow color development in the nacre of akoya pearl oyster shells. Marine Biotechnology, 22(1): 19–30, February 2020. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-019-09927-5.

Gao:2020:CHD

[1555] Dong Gao, Min Zheng, Genmei Lin, Wenyu Fang, Jing Huang, Jianguo Lu, and Xiaowen Sun. Construction of high-density genetic map and mapping of sex-related loci in the yellow catfish (Pelteobagrus fulvidraco). Marine Biotechnology, 22(1):31–40, February 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-019-09928-4.

Zhou:2020:QFM

[1556] Ying Zhou, Haiyang Liu, Xinhua Wang, Beide Fu, Xiaomu Yu, and Jingou Tong. QTL fine mapping for sex determination region in bighead carp (Hypophthalmichthys nobilis) and comparison with silver carp (Hypophthalmichthys molitrix). Marine Biotechnology, 22(1): 41–53, February 2020. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-019-09929-3.

Chen:2020:SGI

[1557] Fan Chen, Xi-Yin Li, Li Zhou, Peng Yu, Zhong-Wei Wang, Zhi Li, Xiao-Juan Zhang, Yang Wang, and Jian-Fang Gui. Stable genome incorporation of sperm-derived DNA fragments in gynogenetic clone of gibel carp. Marine Biotechnology, 22(1):54–66, February 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-019-09930-w.

Bridges:2020:ECB

[1558] Mary C. Bridges, Cheryl M. Woodley, Esther C. Peters, Lisa A. May, and Sylvia B. Galloway. Expression and characterization of a bright farred fluorescent protein from the pink-pigmented tissues of Porites lobata. Marine Biotechnology, 22(1):67–80, February 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-019-09931-9.

Wu:2020:ESA

[1559] Ping Wu, Aimin Wang, Jia Cheng, Lin Chen, Yaxiong Pan, Honghui Li, Qi Zhang, Jiaqi Zhang, Wuying Chu, and Jianshe Zhang. Effects of starvation on antioxidant-related signaling molecules, oxidative stress, and autophagy in juvenile Chinese perch skeletal muscle. Marine Biotechnology, 22(1):81–93, February 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-019-09933-7.

Dettleff:2020:IEL

[1560] Phillip Dettleff, Elizabeth Hormazabal, Jorge Aedo, Marcia Fuentes, Claudio Meneses, Alfredo Molina, and Juan Antonio Valdes. Identification and evaluation of long noncoding RNAs in response to handling stress in red cusk-eel (Genypterus chilensis) via RNA-seq. Marine Biotechnology, 22(1):94–108, February 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-019-09934-6.

Yanez:2020:HTS

[1561] José M. Yáñez, Grazyella Yoshida, Agustín Barria, Ricardo Palma-Véjares, Dante Travisany, Diego Díaz, Giovanna Cáceres, María I. Cádiz, María E. López, Jean P. Lhorente, Ana Jedlicki, José Soto, Diego Salas, and Alejandro Maass. High-throughput single nucleotide polymorphism (SNP) discovery and validation through whole-genome resequencing in Nile tilapia (Oreochromis niloticus). Marine Biotechnology, 22(1):109– 117, February 2020. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-019-09935-5.

Coutinho:2020:ERP

[1562] Paula Coutinho, Martiña Ferreira, Isabel Freire, and Ana Otero. Enriching rotifers with “Premium” microalgae: Rhodomonas lens. Marine Biotechnology, 22(1):118–129, February 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-019-09936-4.

Shi:2020:CHD

[1563] Yue Shi, Zhixiong Zhou, Bo Liu, Shengnan Kong, Baohua Chen, Huaqiang Bai, Leibin Li, Fei Pu, and Peng Xu. Construction of a high-density genetic linkage map and QTL mapping for growthrelated traits in Takifugu bimaculatus. Marine Biotechnology, 22(1): 130–144, February 2020. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-019-09938-2.

Zhang:2020:GPC

[1564] Yaqun Zhang, Zhanjiang Liu, and Hengde Li. Genomic prediction of columnaris disease resistance in catfish. Marine Biotechnology, 22(1): 145–151, February 2020. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-019-09941-7.

Yang:2020:ICG

[1565] Yang Yang, Lina Wu, Xi Wu, Bijun Li, Wenhua Huang, Zhuoying Weng, Zixuan Lin, Leling Song, Yin Guo, Zining Meng, Xiaochun Liu, and Junhong Xia. Identification of candidate growth-related SNPs and genes using GWAS in brown-marbled grouper (Epinephelus fuscoguttatus). Marine Biotechnology, 22(2):153–166, April 2020. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-019-09940-8.

Liu:2020:IHT

[1566] Youli Liu, Qihui Zhu, Li Li, Wei Wang, and Guofan Zhang. Identification of HSF1 target genes involved in thermal stress in the Pacific oyster Crassostrea gigas by ChIP-seq. Marine Biotechnology, 22 (2):167–179, April 2020. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-019-09942-6.

Yi:2020:SBR

[1567] Shaokui Yi, Li-Fang Liu, Lai-Fang Zhou, Bo-Wen Zhao, Wei-Min Wang, and Ze-Xia Gao. Screening of biomarkers related to ovarian maturation and spawning in blunt snout bream (Megalobrama amblycephala) based on metabolomics and transcriptomics. Marine Biotechnology, 22 (2):180–193, April 2020. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-019-09943-5.

Bellan:2020:NCS

[1568] D. L. Bellan, E. Mazepa, S. M. P. Biscaia, J. P. Gonçalves, C. C. Oliveira, G. R. Rossi, L. G. Ferreira, M. D. Noseda, E. S. Trindade, M. E. R. Duarte, and C. R. C. Franco. Non-cytotoxic sulfated heterorhamnan from Gayralia brasiliensis green seaweed reduces driver features of melanoma metastatic progression. Marine Biotechnology, 22 (2):194–206, April 2020. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-020-09944-9.

Mitchell:2020:TTS

[1569] Michela L. Mitchell, Gerry Q. Tonkin-Hill, Rodrigo A. V. Morales, Anthony W. Purcell, Anthony T. Papenfuss, and Raymond S. Norton. Tentacle transcriptomes of the speckled anemone (Actiniaria: Actiniidae: Oulactis sp.): Venom-related components and their domain structure. Marine Biotechnology, 22(2):207–219, April 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-020-09945-8.

Dong:2020:MPA

[1570] Chuanju Dong, Xiaodi Duan, Laghari Muhammad Younis, Meng Zhang, Xiao Ma, Baohua Chen, Xuejun Li, and Peng Xu. Mitogenomic perspectives on the adaptation to extreme alkaline environment of Amur ide (Leuciscus waleckii). Marine Biotechnology, 22(2):220–232, April 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-020-09946-7.

Li:2020:AAR

[1571] Ao Li, Li Li, Wei Wang, and Guofan Zhang. Acetylome analysis reveals population differentiation of the Pacific oyster Crassostrea gigas in response to heat stress. Marine Biotechnology, 22(2):233–245, April 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-020-09947-6.

Shi:2020:PCI

[1572] Yu Shi, Mi Zhao, and Maoxian He. PfSMAD1/5 can interact with PfSMAD4 to inhibit PfMSX to regulate shell biomineralization in Pinctada fucata martensii. Marine Biotechnology, 22(2):246–262, April 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-020-09948-5.

Xue:2020:LTP

[1573] Xi Xue, Jennifer R. Hall, Albert Caballero-Solares, Khalil Eslamloo, Richard G. Taylor, Christopher C. Parrish, and Matthew L. Rise. Liver transcriptome profiling reveals that dietary DHA and EPA levels influence suites of genes involved in metabolism, redox homeostasis, and immune function in Atlantic salmon (Salmo salar). Marine Biotechnology, 22(2):263–284, April 2020. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-020-09950-x.

vanderBurg:2020:RRR

[1574] Chloé A. van der Burg, Ana Pavasovic, Edward K. Gilding, Elise S. Pelzer, Joachim M. Surm, Hayden L. Smith, Terence P. Walsh, and Peter J. Prentis. The rapid regenerative response of a model sea anemone species Exaiptasia pallida is characterised by tissue plasticity and highly coordinated cell communication. Marine Biotechnology, 22 (2):285–307, April 2020. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-020-09951-w.

deCastro:2020:AHW

[1575] Inês de Castro, Sónia Mendo, and Tânia Caetano. Antibiotics from haloarchaea: What can we learn from comparative genomics? Marine Biotechnology, 22(2):308–316, April 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-020-09952-9.

Birolli:2020:SKA

[1576] Willian Garcia Birolli, Lucas Lima Zanin, David Esteban Quintero Jimenez, and André Luiz Meleiro Porto. Synthesis of knoevenagel adducts under microwave irradiation and biocatalytic ene-reduction by the marine-derived fungus Cladosporium sp. CBMAI 1237 for the production of 2-cyano-3-phenylpropanamide derivatives. Marine Biotechnology, 22(2):317–330, April 2020. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-020-09953-8.

Zhao:2020:CTP

[1577] Liang Zhao, Yangping Li, Jiarun Lou, Zhihui Yang, Huan Liao, Qiang Fu, Zhenyi Guo, Shanshan Lian, Xiaoli Hu, and Zhenmin Bao. Correction to: Transcriptomic profiling provides insights into inbreeding depression in Yesso scallop Patinopecten yessoensis. Marine Biotechnology, 22(2):331, April 2020. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-020-09949-4. See [1535].

Vijayakumar:2020:CIR

[1578] Parameswaran Vijayakumar, João Cardeira, Vincent Laizé, Paulo J. Gavaia, and M. Leonor Cancela. Cells isolated from regenerating caudal fin of Sparus aurata can differentiate into distinct bone cell lineages. Marine Biotechnology, 22(3):333–347, June 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-019-09937-3.

Ferreira:2020:BMP

[1579] Irlon M. Ferreira, Anderson Fiamingo, Sergio P. Campana-Filho, and André L. M. Porto. Biotransformation of (E)-2-methyl-3-phenylacrylaldehyde using mycelia of Penicillium citrinum CBMAI 1186, both free and immobilized on chitosan. Marine Biotechnology, 22(3): 348–356, June 2020. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-020-09954-7.

Parisi:2020:EVB

[1580] Julia Risso Parisi, Kelly Rossetti Fernandes, Matheus de Almeida Cruz, Ingrid Regina Avanzi, Alan de França Santana, Giovanna Caroline Aparecida do Vale, Ana Laura Martins de Andrade, Cíntia Pereira de Góes, Carlos Alberto Fortulan, Eliandra de Sousa Trichês, Renata Neves Granito, and Ana Claudia Muniz Rennó. Evaluation of the In Vivo biological effects of marine collagen and hydroxyapatite composite in a tibial bone defect model in rats. Marine Biotechnology, 22 (3):357–366, June 2020. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-020-09955-6.

Teng:2020:TPA

[1581] Jian Teng, Yan Zhao, Hong Ju Chen, Hui Wang, and Xiang Shan Ji. Transcriptome profiling and analysis of genes associated with high temperature-induced masculinization in sex-undifferentiated Nile tilapia gonad. Marine Biotechnology, 22(3):367–379, June 2020. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-020-09956-5.

Zhang:2020:GWD

[1582] Jiabin Zhang, Shaojie Luo, Zefeng Gu, Yuewen Deng, and Yu Jiao. Genome-wide DNA methylation analysis of mantle edge and mantle central from pearl oyster Pinctada fucata martensii. Marine Biotechnology, 22(3):380–390, June 2020. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-020-09957-4.

Sumikawa:2020:SAA

[1583] Kana Sumikawa, Kentaro Takei, Yuya Kumagai, Takeshi Shimizu, Hajime Yasui, and Hideki Kishimura. In silico analysis of ACE inhibitory peptides from chloroplast proteins of red alga Grateloupia asiatica. Marine Biotechnology, 22(3):391–402, June 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-020-09959-2.

Shirodkar:2020:MTP

[1584] Priyanka V. Shirodkar, Usha Devi Muraleedharan, Samir Damare, and Seshagiri Raghukumar. A mesohaline thraustochytrid produces extremely halophilic alpha-amylases. Marine Biotechnology, 22(3):403– 410, June 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-

2236 (electronic). URL https://link.springer.com/article/10.
1007/s10126-020-09960-9.

Ancora:2020:AEC

[1585] Stefania Ancora, Federica Rossi, Marina Borgese, Cristina Pirrone, Ilaria Caliani, Simone Cappello, Giuseppe Mancini, Nicola Bianchi, Claudio Leonzio, Giovanni Bernardini, and Rosalba Gornati. Assessing the effect of contaminated and restored marine sediments in different experimental mesocosms using an integrated approach and Mytilus galloprovincialis as a model. Marine Biotechnology, 22(3):411–422, June 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-020-09961-8.

Wang:2020:SST

[1586] Yue Wang, Yang Yu, Shihao Li, Xiaojun Zhang, Jianhai Xiang, and Fuhua Li. Sex-specific transcriptome sequencing of zoea I larvae and identification of sex-linked genes using bulked segregant analysis in Pacific white shrimp Litopenaeus vannamei. Marine Biotechnology, 22 (3):423–432, June 2020. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-020-09962-7. See correction [1668].

Qin:2020:RGE

[1587] Qinbo Qin, Chongqing Wang, Yuwei Zhou, Huan Qin, Chun Zhao, Li Yang, Tingting Yu, and Shaojun Liu. Rapid genomic and epigenetic alterations in gynogenetic Carassius auratus red var. derived from distant hybridization. Marine Biotechnology, 22(3):433–442, June 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-020-09963-6.

Peng:2020:BFT

[1588] Liangyue Peng, Wen Fu, Xianlong Wu, Sheng He, Han Zhao, Jinhui Liu, Wenbin Liu, and Yamei Xiao. Bisexual fertile triploid zebrafish (Danio rerio): a rare case. Marine Biotechnology, 22(3):443– 455, June 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-

2236 (electronic). URL https://link.springer.com/article/10.
1007/s10126-020-09964-5.

Tan:2020:SMM

[1589] Kianann Tan, Miao Zhou, Huigong Jiang, Donghuo Jiang, Yanhe Li, and Weimin Wang. siRNA-Mediated MrIAG silencing induces sex reversal in Macrobrachium rosenbergii. Marine Biotechnology, 22(3): 456–466, June 2020. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-020-09965-4.

Li:2020:IMS

[1590] Ming Li, Hao Xu, Wenteng Xu, Qian Zhou, Xiwen Xu, Ying Zhu, Weiwei Zheng, Wensheng Li, Zunfang Pang, and Songlin Chen. Isolation of a male-specific molecular marker and development of a genetic sex identification technique in spotted knifejaw (Oplegnathus punctatus). Marine Biotechnology, 22(4):467–474, August 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-020-09966-3.

Sun:2020:MCI

[1591] Junjun Sun, Cuiying Chen, Cuihong You, Yang Liu, Hongyu Ma, Óscar Monroig, Douglas R. Tocher, Shuqi Wang, and Yuanyou Li. The miR-15/ 16 cluster is involved in the regulation of vertebrate LC–PUFA biosynthesis by targeting pparγ as demonostrated in rabbitfish Siganus canaliculatus. Marine Biotechnology, 22(4):475–487, August 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-020-09969-0.

Sakatoku:2020:MIC

[1592] Akihiro Sakatoku, Masahito Ishikawa, Kanna Yamazaki, Tomoya Nakamachi, Hiroyuki Kamachi, Daisuke Tanaka, and Shogo Nakamura. Molecular identification, characterization, and expression analysis of a metallothionein gene from Septifer virgatus. Marine Biotechnology, 22(4): 488–497, August 2020. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-020-09970-7.

Qiang:2020:FSD

[1593] Le Qiang, Yu Zhang, Chao Wu, Yingkuan Han, Suchun Wang, Yanyan Wang, Congcong Zhang, Guangzhou Liu, Qi Wu, Hong Liu, Ian R. Jenkinson, Jun Sun, and Lin Han. A facile and sensitive DNA sensing of harmful algal blooms based on graphene oxide nanosheets. Marine Biotechnology, 22(4):498–510, August 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-020-09971-6.

Caballero-Solares:2020:DIP

[1594] Albert Caballero-Solares, Xi Xue, Beth M. Cleveland, Maryam Beheshti Foroutani, Christopher C. Parrish, Richard G. Taylor, and Matthew L. Rise. Diet-induced physiological responses in the liver of Atlantic salmon (Salmo salar) inferred using multiplex PCR platforms. Marine Biotechnology, 22(4):511–525, August 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-020-09972-5. See [1681].

Liu:2020:FHD

[1595] Yang Liu, Haolong Wang, Haishen Wen, Yue Shi, Meizhao Zhang, Xin Qi, Kaiqiang Zhang, Qingli Gong, Jifang Li, Feng He, Yanbo Hu, and Yun Li. First high-density linkage map and QTL fine mapping for growth-related traits of spotted sea bass (Lateolabrax maculatus). Marine Biotechnology, 22(4):526–538, August 2020. CODEN MABIFW.

ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-020-09973-4.

Cleveland:2020:TRS

[1596] Beth M. Cleveland, Guangtu Gao, and Timothy D. Leeds. Transcriptomic response to selective breeding for fast growth in rainbow trout (Oncorhynchus mykiss). Marine Biotechnology, 22(4):539–550, August 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-020-09974-3.

Nemoto:2020:CGA

[1597] Michiko Nemoto, Sayako Iwaki, Hisao Moriya, Yuki Monden, Takashi Tamura, Kenji Inagaki, Shigeki Mayama, and Kiori Obuse. Comparative gene analysis focused on silica cell wall formation: Identification of diatom-specific SET domain protein methyltransferases. Marine Biotechnology, 22(4):551–563, August 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-020-09976-1. See correction [1640].

Wang:2020:PGL

[1598] Quanchao Wang, Yang Yu, Qian Zhang, Zheng Luo, Xiaojun Zhang, Jianhai Xiang, and Fuhua Li. The polymorphism of LvMMD2 and its association with growth traits in Litopenaeus vannamei. Marine Biotechnology, 22(4):564–571, August 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-020-09977-0.

Wang:2020:CGE

[1599] Kang Wang, Yulin Cui, Yinchu Wang, Zhengquan Gao, Tianzhong Liu, Chunxiao Meng, and Song Qin. Chloroplast genetic engineering of a unicellular green alga Haematococcus pluvialis with expression of an antimicrobial peptide. Marine Biotechnology, 22(4):572–580, August 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-020-09978-z.

Madaro:2020:RWG

[1600] Angelico Madaro, Ole Torrissen, Paul Whatmore, Santosh P. Lall, Jerome Schmeisser, Viviane Verlhac Trichet, and Rolf Erik Olsen. Red and white Chinook salmon (Oncorhynchus tshawytscha): Differences in the transcriptome profile of muscle, liver, and pylorus. Marine Biotechnology, 22(4):581–593, August 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-020-09980-5.

Zhou:2020:MMR

[1601] Mingcan Zhou, Xiwei Jia, Haifu Wan, Shuhong Wang, Xin Zhang, Ziping Zhang, and Yilei Wang. miR-9 and miR-263 regulate the key genes of the ERK pathway in the ovary of mud crab Scylla paramamosain. Marine Biotechnology, 22(4):594–606, August 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-020-09981-4.

Tsutsui:2020:DVC

[1602] Yuri Tsutsui, Teika Onoue, Jun ichi Hikima, Masahiro Sakai, and Tomoya Kono. Diel variation in CC chemokine gene expression in the Japanese pufferfish Takifugu rubripes. Marine Biotechnology, 22(5): 607–612, October 2020. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-020-09988-x.

Liu:2020:GEG

[1603] Chengjie Liu, Ding Ye, Houpeng Wang, Mudan He, and Yonghua Sun. Elovl2 but not Elovl5 is essential for the biosynthesis of docosahexaenoic acid (DHA) in zebrafish: Insight from a comparative gene knockout study. Marine Biotechnology, 22(5):613–619, October 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-020-09992-1.

Wu:2020:MMN

[1604] Ping Wu, Lin Chen, Jia Cheng, Yaxiong Pan, Xinhong Guo, Wuying Chu, Jianshe Zhang, and Xuanming Liu. MiRNAs-modulation of Nrf2 signaling networks in regulation oxidative stress of Chinese perch skeletal muscle after fasting treatment. Marine Biotechnology, 22(5): 620–630, October 2020. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-020-09982-3.

Kong:2020:GWA

[1605] Shengnan Kong, Zhixiong Zhou, Tao Zhou, Ji Zhao, Lin Chen, Huanling Lin, Fei Pu, Qiaozhen Ke, Huaqiang Bai, and Peng Xu. Genomewide association study of body shape-related traits in large yellow croaker (Larimichthys crocea). Marine Biotechnology, 22(5):631–643, October 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-020-09983-2.

Zhang:2020:TEG

[1606] Changqing Zhang, Ziheng Ren, and Zhiyuan Gong. Transgenic expression and genome editing by electroporation of zebrafish embryos. Marine Biotechnology, 22(5):644–650, October 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-020-09985-0.

Thuoc:2020:GMR

[1607] Doan Van Thuoc, Tran Thi Loan, Trieu Anh Trung, Nguyen Van Quyen, Quach Ngoc Tung, Phi Quyet Tien, and Kumar Sudesh. Genome mining reveals the biosynthetic pathways of polyhydroxyalkanoate and ectoines of the halophilic strain Salinivibrio proteolyticus M318 isolated from fermented shrimp paste. Marine Biotechnology, 22(5):651– 660, October 2020. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-020-09986-z.

Lin:2020:EGM

[1608] Shanmeng Lin, Tongtong Kong, Xin Ren, Shengkang Li, and Yi Gong. Elucidation of gut microbiota in mud crab Scylla paramamosain challenged to WSSV and Aeromonas hydrophila. Marine Biotechnology, 22 (5):661–672, October 2020. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-020-09987-y.

deQueiroz:2020:BEW

[1609] Thayane Melo de Queiroz, Javier Ellena, and André L. M. Porto. Biotransformation of ethinylestradiol by whole cells of Brazilian marinederived fungus Penicillium oxalicum CBMAI 1996. Marine Biotechnology, 22(5):673–682, October 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-020-09989-w.

Papadaki:2020:NCR

[1610] Maria Papadaki, Elisavet Kaitetzidou, Constantinos C. Mylonas, and Elena Sarropoulou. Non-coding RNA expression patterns of two different teleost gonad maturation stages. Marine Biotechnology, 22(5): 683–695, October 2020. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-020-09991-2.

Niu:2020:IMS

[1611] Donghong Niu, Beibei Li, Shumei Xie, Zhiguo Dong, and Jiale Li. Integrated mRNA and small RNA sequencing reveals regulatory expression of larval metamorphosis of the razor clam. Marine Biotechnology, 22(5): 696–705, October 2020. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-020-09993-0.

Zhang:2020:TAL

[1612] Jiahua Zhang, Yubang Shen, Xiaoyan Xu, Yafan Dai, and Jiale Li. Transcriptome analysis of the liver and muscle tissues of black carp (Mylopharyngodon piceus) of different growth rates. Marine Biotechnology, 22(5):706–716, October 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-020-09994-z.

Gratacap:2020:EGE

[1613] Remi L. Gratacap, Ye Hwa Jin, Marina Mantsopoulou, and Ross D. Houston. Efficient genome editing in multiple salmonid cell lines using ribonucleoprotein complexes. Marine Biotechnology, 22(5):717– 724, October 2020. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-020-09995-y.

Takeyama:2020:MMB

[1614] Haruko Takeyama, Hiroshi Saito, and Yohei Nishikawa. MBC2019: Marine biotechnology conference 2019. Marine Biotechnology, 22(6):725– 726, December 2020. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-020-10009-0.

Kerfahi:2020:RIB

[1615] Dorsaf Kerfahi, Ben P. Harvey, Sylvain Agostini, Koetsu Kon, Ruiping Huang, Jonathan M. Adams, and Jason M. Hall-Spencer. Responses of intertidal bacterial biofilm communities to increasing p CO_{2}. Marine Biotechnology, 22(6):727–738, December 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-020-09958-3.

Nishi:2020:IDI

[1616] Michika Nishi, Hiroki Kobayashi, Taro Amano, Yuto Sakate, Tomohiro Bito, Jiro Arima, and Katsuhiko Shimizu. Identification of the domains involved in promotion of silica formation in glassin, a protein occluded in hexactinellid sponge biosilica, for development of a tag for purification and immobilization of recombinant proteins. Marine Biotechnology, 22(6):739–747, December 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-020-09967-2.

Takagi:2020:TAI

[1617] Toshiyuki Takagi, Yuki Yoshioka, Yuna Zayasu, Noriyuki Satoh, and Chuya Shinzato. Transcriptome analyses of immune system behaviors in primary polyp of coral Acropora digitifera exposed to the bacterial pathogen Vibrio coralliilyticus under thermal loading. Marine Biotechnology, 22(6):748–759, December 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-020-09984-1.

Yoshida:2020:GTA

[1618] Masa aki Yoshida, Junichi Imoto, Yuri Kawai, Satomi Funahashi, Ryuhei Minei, Yuki Akizuki, Atsushi Ogura, Kazuhiko Nakabayashi, Kei Yura, and Kazuho Ikeo. Genomic and transcriptomic analyses of bioluminescence genes in the enope squid Watasenia scintillans. Marine Biotechnology, 22(6):760–771, December 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-020-10001-8.

Huang:2020:IFS

[1619] Chang-Wen Huang, Pei-Yun Chu, Yu-Fang Wu, Wei-Ren Chan, and Yeh-Hao Wang. Identification of functional SSR markers in freshwater ornamental shrimps Neocaridina denticulata using transcriptome sequencing. Marine Biotechnology, 22(6):772–785, December 2020. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-020-09979-y.

Nguyen:2020:GAS

[1620] Chan D. H. Nguyen, Gianluca Amoroso, Tomer Ventura, Jeremiah J. Minich, and Abigail Elizur. Atlantic salmon (Salmo salar l., 1758) gut microbiota profile correlates with flesh pigmentation: Cause or effect? Marine Biotechnology, 22(6):786–804, December 2020. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-019-09939-1.

Kashitani:2020:TDT

[1621] Maho Kashitani, Taiki Okabe, Hikaru Oyama, Kaede Noguchi, Haruka Yamazaki, Rei Suo, Tetsushi Mori, Haruo Sugita, and Shiro Itoi. Taxonomic distribution of tetrodotoxin in acotylean flatworms (Polycladida: Platyhelminthes). Marine Biotechnology, 22(6):805–811, December 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-020-09968-1.

Okabe:2020:GMC

[1622] Taiki Okabe, Naoyuki Suguro, Tomoko Koito, Kento Endo, Haruo Sugita, and Shiro Itoi. Genetic and morphological characteristics in the local population of the landlocked salmon Oncorhynchus masou originally distributed in Kanagawa Prefecture, Japan. Marine Biotechnology, 22(6):812–823, December 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-020-09975-2.

Nuryadi:2020:CMC

[1623] Handung Nuryadi, Shimpei Sumimoto, Toshiaki Teruya, Kiyotake Suenaga, and Shoichiro Suda. Characterization of macroscopic colonyforming filamentous cyanobacteria from Okinawan coasts as potential sources of bioactive compounds. Marine Biotechnology, 22(6):824– 835, December 2020. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-020-10010-7.

Saito:2020:CAP

[1624] Takeshi Saito, Takahiro Ichihara, Hidetoshi Inoue, Takafumi Uematsu, Saki Hamada, Takaaki Watanabe, Yasushi Takimura, and Jason Webb. Comparison of areal productivity of Nannochloropsis oceanica between lab-scale and industrial-scale raceway pond. Marine Biotechnology, 22 (6):836–841, December 2020. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.com/

article/10.1007/s10126-020-09990-3.

Cruz:2021:CMS

[1625] Matheus Almeida Cruz, Tiago Akira Araujo, Ingrid Regina Avanzi, Julia Risso Parisi, Ana Laura Martins de Andrade, and Ana Claudia Muniz Rennó. Collagen from marine sources and skin wound healing in animal experimental studies: a systematic review. Marine Biotechnology, 23 (1):1–11, February 2021. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-020-10011-6.

Zhou:2021:CSR

[1626] Qian Zhou, Haoyang Gao, Hao Xu, Haoran Lin, and Songlin Chen. A chromosomal-scale reference genome of the kelp grouper Epinephelus moara. Marine Biotechnology, 23(1):12–16, February 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-020-10003-6. See

correction [1627].

Zhou:2021:CCS

[1627] Qian Zhou, Haoyang Gao, Hao Xu, Haoran Lin, and Songlin Chen. Correction to: A chromosomal-scale reference genome of the kelp grouper Epinephelus moara. Marine Biotechnology, 23(1):17, February 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-021-10025-8. See [1626].

Lin:2021:TPR

[1628] Genmei Lin, Dong Gao, Jianguo Lu, and Xiaowen Sun. Transcriptome profiling reveals the sexual dimorphism of gene expression patterns during gonad differentiation in the half-smooth tongue sole (Cynoglossus semilaevis). Marine Biotechnology, 23(1):18–30, February 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-020-09996-x.

Shono:2021:GLR

[1629] Chika Shono, Dwi Ariyanti, Koichi Abe, Yuta Sakai, Ippei Sakamoto, Kaori Tsukakoshi, Koji Sode, and Kazunori Ikebukuro. A green lightregulated T7 RNA polymerase gene expression system for cyanobacteria. Marine Biotechnology, 23(1):31–38, February 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-020-09997-w.

Li:2021:FGW

[1630] Hui Li, Kehui Li, Zhongrui Guo, Han Xue, Jing Li, Sixue Ji, Jiqian Wang, and Hu Zhu. The function of β-1,4-glucuronosyltransferase WelK in the sphingan WL gum biosynthesis process in marine Sphingomonas sp. WG. Marine Biotechnology, 23(1):39–50, February 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-020-09998-9.

Levert:2021:AAM

[1631] Annabel Levert, Valentin Foulon, Marilyne Fauchon, Nathalie Tapissier-Bontemps, Bernard Banaigs, and Claire Hellio. Antifouling activity of meroterpenes isolated from the ascidian Aplidium aff. densum. Marine Biotechnology, 23(1):51–61, February 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-020-10000-9.

Tripp-Valdez:2021:GPT

[1632] M. A. Tripp-Valdez, F. Cicala, C. E. Galindo-Sánchez, K. D. Chacón-Ponce, E. López-Landavery, F. Díaz, D. Re-Araujo, and F. Lafarga-De la Cruz. Growth performance and transcriptomic response of warm-acclimated hybrid abalone Haliotis rufescens ([female sign]) × H. corrugata ([male sign]). Marine Biotechnology, 23(1):62–76, February 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-020-10002-7.

daSilvaBarbosa:2021:VAP

[1633] Jefferson da Silva Barbosa, Laís Cristina Gusmão Ferreira Palhares, Cynthia Haynara Ferreira Silva, Diego Araujo Sabry, Suely Ferreira Chavante, and Hugo Alexandre Oliveira Rocha. In vitro antitumor potential of sulfated polysaccharides from seaweed Caulerpa cupressoides var. flabellata. Marine Biotechnology, 23(1):77–89, February 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-020-10004-5.

Ma:2021:CGT

[1634] Xiaoli Ma, Baofeng Su, Max Bangs, Veronica Alston, Nathan J. C. Backenstose, Rhoda Mae Simora, Wenwen Wang, De Xing, Shangjia Li, Zhi Ye, Anthony G. Moss, Thuy-Yen Duong, Xu Wang, and Rex A. Dunham. Comparative genomic and transcriptomic analyses revealed twenty-six candidate genes involved in the air-breathing development and function of the bighead catfish Clarias macrocephalus. Marine Biotechnology, 23 (1):90–105, February 2021. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-020-10005-4.

Curiel-Maciel:2021:CEC

[1635] Nidya Fabiola Curiel-Maciel, Fernando Martínez-Morales, Alexei Fedorovish Licea-Navarro, Brandt Bertrand, A. Berenice Aguilar-Guadarrama, Nashbly Sarela Rosas-Galván, Daniel Morales-Guzmán, Nancy Rivera-Gómez, Rosa Maria Gutiérrez-Ríos, and María R. Trejo-Hernández. Characterization of Enterobacter cloacae BAGM01 producing a thermostable and alkaline-tolerant rhamnolipid biosurfactant from the Gulf of Mexico. Marine Biotechnology, 23(1):106–126, February 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-020-10006-3.

Lin:2021:GWA

[1636] Huanling Lin, Zhixiong Zhou, Ji Zhao, Tao Zhou, Huaqiang Bai, Qiaozhen Ke, Fei Pu, Weiqiang Zheng, and Peng Xu. Genome-wide association study identifies genomic loci of sex determination and gonadosomatic index traits in large yellow croaker (Larimichthys crocea). Marine Biotechnology, 23(1):127–139, February 2021. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-020-10007-2.

Chan:2021:SDG

[1637] Michelle T. T. Chan, Annette Muttray, Dionne Sakhrani, Krista Woodward, Jin-Hyoung Kim, Kris A. Christensen, Ben F. Koop, and Robert H. Devlin. Sexually dimorphic growth stimulation in a strain of growth hormone transgenic Coho salmon (Oncorhynchus kisutch). Marine Biotechnology, 23(1):140–148, February 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-020-10012-5.

Tzachor:2021:PCS

[1638] Asaf Tzachor, Or Rozen, Soliman Khatib, Sophie Jensen, and Dorit Avni. Photosynthetically controlled spirulina, but not solar spirulina, inhibits TNF-α secretion: Potential implications for COVID-19-related cytokine storm therapy. Marine Biotechnology, 23(1):149–155, February 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-021-10020-z.

Munar:2021:CDN

[1639] Madison Pascual Munar, Hirokazu Takahashi, and Yoshiko Okamura. Correction to: Discovery of a novel gene conferring tellurite tolerance through tellurite reduction to Escherichia coli transformant in marine sediment metagenomic library. Marine Biotechnology, 23(1): 156, February 2021. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-020-10008-1. See [1547].

Nemoto:2021:CCG

[1640] Michiko Nemoto, Sayako Iwaki, Hisao Moriya, Yuki Monden, Takashi Tamura, Kenji Inagaki, Shigeki Mayama, and Kiori Obuse. Correction to: Comparative gene analysis focused on silica cell wall formation:

Identification of diatom-specific SET domain protein methyltransferases. Marine Biotechnology, 23(1):157, February 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-020-09999-8. See [1597].

Jagtap:2021:OME

[1641] Ashok S. Jagtap and Cathrine S. Manohar. Overview on microbial enzymatic production of algal oligosaccharides for nutraceutical applications. Marine Biotechnology, 23(2):159–176, April 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-021-10027-6.

Yoshikawa:2021:GDP

[1642] Sota Yoshikawa, Masaomi Hamasaki, Kazushi Kadomura, Toshiyuki Yamada, Hisashi Chuda, Kiyoshi Kikuchi, and Sho Hosoya. Genetic dissection of a precocious phenotype in male tiger pufferfish (Takifugu rubripes) using genotyping by random amplicon sequencing, direct (GRAS-Di). Marine Biotechnology, 23(2):177–188, April 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-020-10013-4.

Ji:2021:PNM

[1643] Yinghui Ji, Xue Yang, Dong Yang, and Rongqing Zhang. PU14, a novel matrix protein, participates in pearl oyster, Pinctada fucata, shell formation. Marine Biotechnology, 23(2):189–200, April 2021. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-020-10014-3.

Wan:2021:PPI

[1644] Z. Y. Wan, V. C. L. Lin, and Yue Gen Hua. Pomc plays an important role in sexual size dimorphism in tilapia. Marine Biotechnology, 23 (2):201–214, April 2021. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-020-10015-2.

Li:2021:CCM

[1645] Huijuan Li, Hong Yu, Shaojun Du, and Qi Li. CRISPR/Cas9 mediated high efficiency knockout of myosin essential light chain gene in the Pacific oyster (Crassostrea gigas). Marine Biotechnology, 23(2): 215–224, April 2021. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-020-10016-1.

Jiang:2021:GWI

[1646] Dan Jiang, Wanbo Li, Zhiyong Wang, and Ming Fang. Genome-wide identification of cis-acting expression QTLs in large yellow croaker. Marine Biotechnology, 23(2):225–232, April 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-020-10017-0.

Wang:2021:WGR

[1647] Chongqing Wang, Huan Qin, Chun Zhao, Li Yang, Tingting Yu, Yuxin Zhang, Xiang Luo, Qinbo Qin, and Shaojun Liu. Whole-genome re-sequencing and transcriptome reveal oogenesis-related genes in autotetraploid Carassius auratus. Marine Biotechnology, 23(2):233–241, April 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-021-10018-7.

Zhao:2021:GWA

[1648] Ji Zhao, Tao Zhou, Huaqiang Bai, Qiaozhen Ke, Bijun Li, Mindong Bai, Zhixiong Zhou, Fei Pu, Weiqiang Zheng, and Peng Xu. Genome-wide association analysis reveals the genetic architecture of parasite (Cryptocaryon irritans) resistance in large yellow croaker (Larimichthys crocea). Marine Biotechnology, 23(2):242–254, April 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-021-10019-6.

Zhang:2021:ITM

[1649] Yufei Zhang, Zipeng Gu, Yudong Ren, Lu Wang, Jian Zhang, Chengwei Liang, Shanying Tong, Yitao Wang, Dong Xu, Xiaowen Zhang, and Naihao Ye. Integrating transcriptomics and metabolomics to characterize metabolic regulation to elevated CO_{2} in Chlamydomonas reinhardtii. Marine Biotechnology, 23(2):255–275, April 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-021-10021-y.

Santos:2021:ICF

[1650] Rafaela A. Santos, Aires Oliva-Teles, Pedro Pousão-Ferreira, Russell Jerusik, Maria J. Saavedra, Paula Enes, and Cláudia R. Serra. Isolation and characterization of fish-gut Bacillus spp. as source of natural antimicrobial compounds to fight aquaculture bacterial diseases. Marine Biotechnology, 23(2):276–293, April 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-021-10022-x.

Sun:2021:REC

[1651] Hong-Yan Sun, Yu-Ling Su, Pin-Hong Li, Jia-Yang He, He-Jia Chen, Gang Wang, Shao-Wen Wang, Xiao-Hong Huang, You-Hua Huang, and Qi-Wei Qin. The roles of Epinephelus coioides miR-122 in SGIV infection and replication. Marine Biotechnology, 23(2):294–307, April 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-021-10023-w.

Valenzuela-Munoz:2021:TPL

[1652] Valentina Valenzuela-Muñoz, Juan Antonio Váldes, and Cristian Gallardo-Escárate. Transcriptome profiling of long non-coding RNAs during the Atlantic salmon smoltification process. Marine Biotechnology, 23(2):308–320, April 2021. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-021-10024-9.

Camargo:2021:PCM

[1653] Tavani Rocha Camargo, Simon Khelissa, Nour Eddine Chihib, Emilie Dumas, Jian Wang, Wagner C. Valenti, and Adem Gharsallaoui. Preparation and characterization of microcapsules containing antioxidant fish protein hydrolysates: a new use of bycatch in Brazil. Marine Biotechnology, 23(2):321–330, April 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-021-10026-7.

Kanamoto:2021:DMF

[1654] Akihiko Kanamoto, Yuichi Kato, Erina Yoshida, Tomohisa Hasunuma, and Akihiko Kondo. Development of a method for fucoxanthin production using the haptophyte marine microalga Pavlova sp. OPMS 30543. Marine Biotechnology, 23(2):331–341, April 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-021-10028-5.

Yao:2021:IMM

[1655] Qiucheng Yao, Mengdi Zhang, Shaopo Zu, Hong Yang, Weitian Xie, Jinjun Chen, Zhibao Chen, Ye Ge, Weiwei Zeng, and Zhihui Zhao. Integrated mRNA and microRNA transcriptome sequencing characterizes sequence variants and mRNA-microRNA regulatory networks in grass carp fibroblasts infected with virulent and attenuated GCRV. Marine Biotechnology, 23(2):342–355, April 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-021-10029-4.

Teixeira:2021:VEP

[1656] Thaiz Rodrigues Teixeira, Karen Cristina Rangel, Renata Spagolla Napoleão Tavares, Camila Martins Kawakami, Gustavo Souza dos Santos, Silvya Stuchi Maria-Engler, Pio Colepicolo, Lorena Rigo Gaspar, and Hosana Maria Debonsi. In vitro evaluation of the photoprotective potential of quinolinic alkaloids isolated from the Antarctic marine fungus Penicillium echinulatum for topical use. Marine Biotechnology, 23 (3):357–372, June 2021. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-021-10030-x.

Kawamura:2021:ESC

[1657] Kaz Kawamura, Koki Nishitsuji, Eiichi Shoguchi, Shigeki Fujiwara, and Noriyuki Satoh. Establishing sustainable cell lines of a coral, Acropora tenuis. Marine Biotechnology, 23(3):373–388, June 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-021-10031-w.

Dai:2021:RST

[1658] Ya-Fan Dai, Yu bang Shen, Shen-Tong Wang, Jia-Hua Zhang, Yu-Hong Su, Sheng-Chen Bao, Xiao-Yan Xu, and Jia-Le Li. RNA-Seq transcriptome analysis of the liver and brain of the black carp (Mylopharyngodon piceus) during fasting. Marine Biotechnology, 23(3): 389–401, June 2021. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-021-10032-9.

Rosenblad:2021:GCB

[1659] Magnus Alm Rosenblad, Anna Abramova, Ulrika Lind, Páll Ólason, Stefania Giacomello, Björn Nystedt, and Anders Blomberg. Genomic characterization of the barnacle Balanus improvisus reveals extreme nucleotide diversity in coding regions. Marine Biotechnology, 23(3): 402–416, June 2021. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-021-10033-8.

Li:2021:IAC

[1660] Zhuanzhuan Li, Qi Li, Shikai Liu, Ziqiang Han, Lingfeng Kong, and Hong Yu. Integrated analysis of coding genes and non-coding RNAs associated with shell color in the Pacific oyster (Crassostrea gigas). Marine Biotechnology, 23(3):417–429, June 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-021-10034-7.

Huang:2021:ASO

[1661] Yung-Sen Huang, Wen-Chih Cheng, and Chung-Yen Lin. Androgenic sensitivities and ovarian gene expression profiles prior to treatment in Japanese eel (Anguilla japonica). Marine Biotechnology, 23 (3):430–444, June 2021. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-021-10035-6.

Wang:2021:EGO

[1662] Le Wang, Fei Sun, Yanfei Wen, and Gen Hua Yue. Effects of ocean acidification on transcriptomes in Asian seabass juveniles. Marine Biotechnology, 23(3):445–455, June 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-021-10036-5.

Tsuji:2021:CCC

[1663] Yoshinori Tsuji, George Kusi-Appiah, Noriko Kozai, Yuri Fukuda, Takashi Yamano, and Hideya Fukuzawa. Characterization of a CO_{2}- concentrating mechanism with low sodium dependency in the centric diatom Chaetoceros gracilis. Marine Biotechnology, 23(3):456–462, June 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-021-10037-4.

Shchapova:2021:CIR

[1664] Ekaterina Shchapova, Anna Nazarova, Uliana Vasilyeva, Anton Gurkov, Alexander Ostyak, Andrei Mutin, Renat Adelshin, Natalia Belkova, and Maxim Timofeyev. Cellular immune response of an endemic Lake Baikal amphipod to indigenous Pseudomonas sp. Marine Biotechnology, 23 (3):463–471, June 2021. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-021-10039-2.

Nyunoya:2021:LDA

[1665] Hayato Nyunoya, Tatsuki Noda, You Kawamoto, Yasuhiro Hayashi, Yohei Ishibashi, Makoto Ito, and Nozomu Okino. Lack of Δ5 desaturase activity impairs EPA and DHA synthesis in fish cells from red sea bream and Japanese flounder. Marine Biotechnology, 23(3): 472–481, June 2021. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-021-10040-9.

Roy:2021:TDT

[1666] Uttam K. Roy, Birthe V. Nielsen, and John J. Milledge. Tuning Dunaliella tertiolecta for enhanced antioxidant production by modification of culture conditions. Marine Biotechnology, 23(3):482–500, June 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-021-10041-8.

Klein:2021:DIT

[1667] A. H. Klein, C. A. Motti, A. K. Hillberg, T. Ventura, P. Thomas-Hall, T. Armstrong, T. Barker, P. Whatmore, and S. F. Cummins. Development and interrogation of a transcriptomic resource for the giant triton snail (Charonia tritonis). Marine Biotechnology, 23(3): 501–515, June 2021. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-021-10042-7.

Wang:2021:CSS

[1668] Yue Wang, Yang Yu, Shihao Li, Xiaojun Zhang, Jianhai Xiang, and Fuhua Li. Correction to: Sex-specific transcriptome sequencing of zoea I larvae and identification of sex-linked genes using bulked segregant analysis in Pacific white shrimp Litopenaeus vannamei. Marine Biotechnology, 23(3):516, June 2021. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-021-10038-3. See [1586].

S:2021:NAT

[1669] Anoop B. S., Jayesh Puthumana, Vrinda Sukumaran, Cijo George Vazhappilly, Salini Kombiyil, Rosamma Philip, and Isaac Sarojini Bright Singh. A novel approach of transducing recombinant baculovirus into primary lymphoid cells of Penaeus monodon for developing continuous cell line. Marine Biotechnology, 23(4):517–528, August 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-021-10043-6.

Monteiro:2021:EPM

[1670] M. Monteiro, A. S. Lavrador, R. Santos, F. Rangel, P. Iglesias, M. Tárraga, A. Couto, C. R. Serra, C. Tafalla, E. Da Costa, M. R. Domingues, A. Oliva-Teles, A. P. Carvalho, P. Enes, and P. Díaz-Rosales. Evaluation of the potential of marine algae extracts as a source of functional ingredients using zebrafish as animal model for aquaculture. Marine Biotechnology, 23(4):529–545, August 2021. CODEN MABIFW.

ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-021-10044-5.

Kwak:2021:GSI

[1671] Jun Soung Kwak and Ki Hong Kim. Generation of self-inhibitory recombinant viral hemorrhagic septicemia virus (VHSV) by insertion of viral P gene-targeting artificial MicroRNA into viral genome and effect of dicer gene knockout on the recombinant VHSV replication. Marine Biotechnology, 23(4):546–559, August 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-021-10045-4.

Jackson:2021:CSE

[1672] T. Jackson, E. Ishengoma, and C. Rhode. Cross-species exon capture and whole exome sequencing: Application, utility and challenges for genomic resource development in non-model species. Marine Biotechnology, 23 (4):560–575, August 2021. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-021-10046-3.

Kitamura:2021:SDC

[1673] Ruriko Kitamura, Natsuko Miura, Michihiro Ito, Toshiyuki Takagi, Hideyuki Yamashiro, Yumi Nishikawa, Yuna Nishimura, Keita Kobayashi, and Michihiko Kataoka. Specific detection of coral-associated Ruegeria, a potential probiotic bacterium, in corals and subtropical seawater. Marine Biotechnology, 23(4):576–589, August 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-021-10047-2.

Rakbanjong:2021:CGC

[1674] Natthida Rakbanjong, Tomoyuki Okutsu, Wilaiwan Chotigeat, Anida Songnui, and Monwadee Wonglapsuwan. Cryopreservation of germ cells of banana shrimp (Fenneropenaeus merguiensis) and black tiger shrimp (Penaeus monodon). Marine Biotechnology, 23(4):590–601, August 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-021-10048-1.

Leiva:2021:TGE

[1675] Francisco Leiva, Scarlet Bravo, Killen Ko Garcia, Javier Moya, Osiel Guzman, and Rodrigo Vidal. Temporal gene expression signature of plasma extracellular vesicles–MicroRNAs from post-smolt Coho salmon challenged with Piscirickettsia salmonis. Marine Biotechnology, 23(4):

602–614, August 2021. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-021-10049-0.

Chen:2021:GTA

[1676] Xiuxia Chen, Hui Gong, Hongshu Chi, Binfu Xu, Zaiyu Zheng, and Yulin Bai. Gill transcriptome analysis revealed the difference in gene expression between freshwater and seawater acclimated guppy (Poecilia reticulata). Marine Biotechnology, 23(4):615–627, August 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-021-10053-4.

Wang:2021:TNT

[1677] Yude Wang, Jiajun Yao, Yaxin Luo, Huifang Tan, Xu Huang, Shi Wang, Qinbo Qin, Chun Zhang, Min Tao, Konrad Dabrowski, and Shaojun Liu. Two new types of homodiploid fish and polyploid hybrids derived from the distant hybridization of female koi carp and male bighead carp. Marine Biotechnology, 23(4):628–640, August 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-021-10050-7.

Habuddha:2021:CTT

[1678] Valainipha Habuddha, Chanyatip Suwannasing, Aticha Buddawong, Kanokwan Seenprachawong, Thitinat Duangchan, Chanyarak Sombutkayasith, Aungkura Supokawej, Wattana Weerachatyanukul, and Somluk Asuvapongpatana. Characterization of thrombospondin type 1 repeat in Haliotis diversicolor and its possible role in osteoinduction. Marine Biotechnology, 23(4):641–652, August 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-021-10054-3.

Schmeisser:2021:MMI

[1679] Jerome Schmeisser, Viviane Verlhac-Trichet, Angelico Madaro, Santosh P. Lall, Ole Torrissen, and Rolf Erik Olsen. Molecular mechanism involved in carotenoid metabolism in post-smolt Atlantic salmon: Astaxanthin metabolism during flesh pigmentation and its antioxidant properties. Marine Biotechnology, 23(4):653–670, August 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-021-10055-2.

Burut-Archanai:2021:DMC

[1680] Surachet Burut-Archanai, Déborah Ubertino, Parichat Chumtong, Wuttichai Mhuantong, Sorawit Powtongsook, and Sanit Piyapattanakorn.

Dynamics of microbial community during nitrification biofilter acclimation with low and high ammonia. Marine Biotechnology, 23(4): 671–681, August 2021. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-021-10056-1.

Caballero-Solares:2021:CDI

[1681] Albert Caballero-Solares, Xi Xue, Beth M. Cleveland, Maryam Beheshti Foroutani, Christopher C. Parrish, Richard G. Taylor, and Matthew L. Rise. Correction to: Diet-induced physiological responses in the liver of Atlantic salmon (Salmo salar) inferred using multiplex PCR platforms. Marine Biotechnology, 23(4):682, August 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-021-10051-6. See [1594].

Schwartz:2021:NDI

[1682] Julie Schwartz, Emilie Réalis-Doyelle, Lorane Le Franc, and Pascal Favrel. A novel Dop2/invertebrate-type dopamine signaling system potentially mediates stress, female reproduction, and early development in the Pacific oyster (Crassostrea gigas). Marine Biotechnology, 23(5): 683–694, October 2021. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-021-10052-5. See correction [1683].

Schwartz:2021:CND

[1683] Julie Schwartz, Emilie Réalis-Doyelle, Lorane Le Franc, and Pascal Favrel. Correction to: A novel Dop2/invertebrate-type dopamine signaling system potentially mediates stress, female reproduction, and early development in the Pacific oyster (Crassostrea gigas). Marine Biotechnology, 23(5):695, October 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.com/

article/10.1007/s10126-021-10070-3. See [1682].

Zhang:2021:GWC

[1684] Xiaoyan Zhang, Peng Yu, Haishen Wen, Xin Qi, Yuan Tian, Kaiqiang Zhang, Qiang Fu, Yun Li, and Chao Li. Genome-wide characterization of Aquaporins (aqps) in Lateolabrax maculatus: Evolution and expression patterns during freshwater acclimation. Marine Biotechnology, 23(5): 696–709, October 2021. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-021-10057-0.

Nunez-Acuna:2021:CTA

[1685] Gustavo Núñez-Acuña, Valentina Valenzuela-Muñoz, Diego Valenzuela-Miranda, and Cristian Gallardo-Escárate. Comprehensive transcriptome analyses in sea louse reveal novel delousing drug responses through MicroRNA regulation. Marine Biotechnology, 23(5):710–723, October 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-021-10058-z.

Lefevre:2021:DRF

[1686] Mathilde Lefevre, Thomas Ederth, Thibault Masai, Ruddy Wattiez, Philippe Leclère, Patrick Flammang, and Elise Hennebert. Disentangling the roles of functional domains in the aggregation and adsorption of the multimodular sea star adhesive protein Sfp1. Marine Biotechnology, 23 (5):724–735, October 2021. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-021-10059-y.

Jiao:2021:GSV

[1687] Zexin Jiao, Yuan Tian, Boyang Hu, Qi Li, and Shikai Liu. Genome structural variation landscape and its selection signatures in the fast-growing strains of the Pacific oyster, Crassostrea gigas. Marine Biotechnology, 23 (5):736–748, October 2021. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-021-10060-5.

Guerrero-Tortolero:2021:TID

[1688] Danitzia A. Guerrero-Tortolero, Grecia Vázquez-Islas, and Rafael Campos-Ramos. A transcriptome insight during early fish larval development followed by starvation in Seriola rivoliana. Marine Biotechnology, 23(5):749–765, October 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-021-10061-4.

Zhao:2021:CMM

[1689] Na Zhao, Lei Jia, Guangli Li, Xiaoxu He, Chunhua Zhu, and Bo Zhang. Comparative mucous miRomics in Cynoglossus semilaevis related to Vibrio harveyi caused infection. Marine Biotechnology, 23(5):766–776, October 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-021-10062-3.

Zhu:2021:SBP

[1690] Yijing Zhu, Qi Li, Hong Yu, Shikai Liu, and Lingfeng Kong. Shell biosynthesis and pigmentation as revealed by the expression of tyrosinase and tyrosinase-like protein genes in Pacific oyster (Crassostrea gigas) with different shell colors. Marine Biotechnology, 23(5):777– 789, October 2021. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-021-10063-2.

Chakraborty:2021:LFT

[1691] Kajal Chakraborty, Vinaya Kizhakkepatt Kizhakkekalam, Minju Joy, and Rekha Devi Chakraborty. A leap forward towards unraveling newer anti-infective agents from an unconventional source: a draft genome sequence illuminating the future promise of marine heterotrophic Bacillus sp. against drug-resistant pathogens. Marine Biotechnology, 23(5): 790–808, October 2021. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-021-10064-1.

Anggraeni:2021:CMT

[1692] Santi Rukminita Anggraeni and Marion B. Ansorge-Schumacher. Characterization and modeling of thermostable GH50 agarases from Microbulbifer elongatus PORT2 . Marine Biotechnology, 23(5):809–820, October 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-021-10065-0.

Hasan:2021:REM

[1693] Sharmin Hasan, Shuichi Asakawa, Shugo Watabe, and Shigeharu Kinoshita. Regulation of the expression of the myosin heavy chain (MYH) gene myh14 in zebrafish development. Marine Biotechnology, 23(5): 821–835, October 2021. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-021-10066-z.

Suwannasing:2021:BMP

[1694] Chanyatip Suwannasing, Aticha Buddawong, Sarawut Khumpune, Valainipha Habuddha, Wattana Weerachatyanukul, and Somluk Asuvapongpatana. Bone morphogenetic protein 2/4 in mollusk, Haliotis diversicolor: Its expression and osteoinductive function In Vitro. Marine Biotechnology, 23(5):836–846, October 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-021-10071-2.

Tohfuku:2021:DIS

[1695] Takuma Tohfuku, Hitoshi Ando, Naho Morishita, Michiaki Yamashita, and Masakazu Kondo. Dietary intake of selenoneine enhances antioxidant activity in the muscles of the amberjack Seriola dumerili grown in aquaculture. Marine Biotechnology, 23(6):847–853, December 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-021-10067-y.

Yang:2021:ERN

[1696] Zituo Yang, Sek Man Wong, and Gen Hua Yue. Effects of rrm1 on NNV resistance revealed by RNA-seq and gene editing. Marine Biotechnology, 23(6):854–869, December 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-021-10068-x.

Abass:2021:GDG

[1697] Nermeen Y. Abass, Baofeng Su, Ahmed Alsaqufi, Ahmed Elaswad, Zhenkui Qin, Hanbo Li, Ramjie Odin, Zhi Ye, and Rex A. Dunham. Growth differences of growth hormone transgenic female and male channel catfish, Ictalurus punctatus, grown in earthen ponds to sexual maturation. Marine Biotechnology, 23(6):870–880, December 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-021-10069-w.

Camargo:2021:BPH

[1698] Tavani R. Camargo, Paulo Mantoan, Patrícia Ramos, José M. Monserrat, Carlos Prentice, Célio C. Fernandes, Willian F. Zambuzzi, and Wagner C. Valenti. Bioactivity of the protein hydrolysates obtained from the most abundant crustacean bycatch. Marine Biotechnology, 23(6): 881–891, December 2021. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-021-10072-1.

Tortorella:2021:EAE

[1699] Emiliana Tortorella, Rosa Giugliano, Marleen De Troch, Bruno Vlaeminck, Gercende Courtois de Viçose, and Donatella de Pascale. The ethyl acetate extract of the marine edible gastropod Haliotis tuberculata coccinea: a potential source of bioactive compounds. Marine Biotechnology, 23(6):892–903, December 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-021-10073-0.

Labriere:2021:SAR

[1700] Christophe Labriere, Gunnar Cervin, Henrik Pavia, Jørn H. Hansen, and Johan Svenson. Structure–activity relationship probing of the natural marine antifoulant barettin. Marine Biotechnology, 23(6):904– 916, December 2021. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-021-10074-z.

Cheun-Arom:2021:EDE

[1701] Thaniwan Cheun-Arom and Taksina Chuanasa. An efficient DNA extraction for a blue Xestospongia sp. sponge and its associated microorganisms containing cytotoxic substances. Marine Biotechnology, 23(6): 917–927, December 2021. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-021-10075-y.

Cleverley:2021:VOC

[1702] Robert Cleverley, David Webb, Stuart Middlemiss, Phillip Duke, Anthony Clare, Keiju Okano, Colin Harwood, and Nick Aldred. In vitro oxidative crosslinking of recombinant barnacle cyprid cement gland proteins. Marine Biotechnology, 23(6):928–942, December 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-021-10076-x. See

correction [1741].

Huang:2021:ITE

[1703] Yung-Sen Huang, Chung-Yen Lin, and Wen-Chih Cheng. Investigating the transcriptomic and expression presence–absence variation exist in Japanese eel (Anguilla japonica), a primitive teleost. Marine Biotechnology, 23(6):943–954, December 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-021-10077-w.

Ambili:2021:ESC

[1704] Ambili M and Denoj Sebastian. Evaluation of sensitivity and costeffectiveness of molecular methods for the co-detection of waterborne pathogens in India. Marine Biotechnology, 23(6):955–963, December 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-021-10078-9.

Shen:2021:SAC

[1705] Huiyu Shen, Tingting Song, Jiaqi Lu, Qiongfen Qiu, Jiong Chen, and Jinbo Xiong. Shrimp AHPND causing Vibrio anguillarum infection: Quantitative diagnosis and identifying antagonistic bacteria. Marine Biotechnology, 23(6):964–975, December 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-021-10079-8.

Xu:2021:MMI

[1706] Qiaoyue Xu, Hongtao Nie, Zhihui Yin, Yanming Zhang, Zhongming Huo, and Xiwu Yan. MiRNA-mRNA integration analysis reveals the regulatory roles of MiRNAs in shell pigmentation of the Manila clam (Ruditapes philippinarum). Marine Biotechnology, 23(6):976–993, December 2021. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-021-10080-1.

Anonymous:2022:TYN

[1707] Anonymous. Thank you note to all our 2021 reviewers. Marine Biotechnology, 24(1):1–4, February 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-022-10108-0.

Sun:2022:RSA

[1708] Jun Sun, Zhe Liu, Jinqiang Quan, Lanlan Li, Guiyan Zhao, and Junhao Lu. RNA-seq analysis reveals alternative splicing under heat stress in rainbow trout (Oncorhynchus mykiss). Marine Biotechnology, 24(1): 5–17, February 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-021-10082-z.

Sun:2022:GWD

[1709] Dongfang Sun, Hong Yu, and Qi Li. Genome-wide differential DNA methylomes provide insights into the infertility of triploid oysters. Marine Biotechnology, 24(1):18–31, February 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-021-10083-y.

Wang:2022:GRN

[1710] Quanchao Wang, Ying Liu, Yanxia Wang, Shaoyu Jiang, Chuanxin Zhang, and Baoquan Li. GWAS reveal novel sex-related markers and candidate genes in sea urchin Mesocentrotus nudus. Marine Biotechnology, 24(1):32–39, February 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-021-10084-x.

Reboleira:2022:UBA

[1711] João Reboleira, Rafael Félix, Tânia F. L. Vicente, Adriana P. Januário, Carina Félix, Marcelo M. R. de Melo, Carlos M. Silva, Ana C. Ribeiro, Jorge A. Saraiva, Narcisa M. Bandarra, Maria Sapatinha, Maria C. Paulo, Joana Coutinho, and Marco F. L. Lemos. Uncovering the bioactivity of Aurantiochytrium sp.: a comparison of extraction methodologies. Marine Biotechnology, 24(1):40–54, February 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-021-10085-w.

Xu:2022:FLT

[1712] Yu Xu, Hai Lin, Weihui Yan, Jiajia Li, Mengling Sun, Jiaping Chen, and Zhiqiang Xu. Full-length transcriptome of red swamp crayfish hepatopancreas reveals candidate genes in Hif-1 and antioxidant pathways in response to hypoxia–reoxygenation. Marine Biotechnology, 24(1): 55–67, February 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-021-10086-9.

He:2022:DSW

[1713] Shan He, Wei-Bing Peng, Xian-Jun Fu, Hong-Lei Zhou, and Zhen-Guo Wang. Deep sea water alleviates Tau phosphorylation and cognitive impairment via PI3K/Akt/GSK-3β pathway. Marine Biotechnology, 24 (1):68–81, February 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-021-10087-8.

Zeng:2022:MTP

[1714] Dan Zeng and Ximing Guo. Mantle transcriptome provides insights into biomineralization and growth regulation in the Eastern oyster (Crassostrea virginica). Marine Biotechnology, 24(1):82–96, February 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-021-10088-7.

Fu:2022:CTA

[1715] Qiang Fu, Yuqing Li, Hao Zhang, Min Cao, Lu Zhang, Chengbin Gao, Xin Cai, Defeng Chen, Ziying Yang, Jie Li, Ning Yang, and Chao Li. Comparative transcriptome analysis of spleen reveals potential regulation of genes and immune pathways following administration of Aeromonas salmonicida subsp . masoucida vaccine in Atlantic salmon (Salmo salar). Marine Biotechnology, 24(1):97–115, February 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-021-10089-6.

Wright-LaGreca:2022:GOA

[1716] Marissa Wright-LaGreca, Clara Mackenzie, and Timothy J. Green. Ocean acidification alters developmental timing and gene expression of ion transport proteins during larval development in resilient and susceptible lineages of the Pacific oyster (Crassostrea gigas). Marine Biotechnology, 24(1):116–124, February 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-022-10090-7.

Wan:2022:CTRa

[1717] Haifu Wan, Jinying Zhong, Ziping Zhang, Pengfei Zou, and Yilei Wang. Comparative transcriptome reveals the potential modulation mechanisms of Spfoxl-2 affecting ovarian development of Scylla paramamosain. Marine Biotechnology, 24(1):125–135, February 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-022-10091-6.

Ye:2022:SCT

[1718] Huan Ye, Yutaka Takeuchi, Hao Du, Huamei Yue, Rui Ruan, Chuangju Li, and Qiwei Wei. Spermatogonia from cryopreserved testes of critically endangered Chinese sturgeon efficiently colonized and preferentially proliferated in the recipient gonads of Yangtze sturgeon. Marine Biotechnology, 24(1):136–150, February 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-022-10092-5.

Xie:2022:TMA

[1719] Jiahui Xie, Yi Sun, Yue Cao, Lingshu Han, Yuanxin Li, Beichen Ding, Chuang Gao, Pengfei Hao, Xin Jin, Yaqing Chang, Jian Song, Donghong Yin, and Jun Ding. Transcriptomic and metabolomic analyses provide insights into the growth and development advantages of triploid Apostichopus japonicus. Marine Biotechnology, 24(1):151–162, February 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-022-10093-4.

Gong:2022:SSG

[1720] Jie Gong, Bijun Li, Ji Zhao, Zhixiong Zhou, Qiaozhen Ke, Qihui Zhu, Dongdong Xu, Tao Zhou, and Peng Xu. Sex-specific genomic region identification and molecular sex marker development of rock bream (Oplegnathus fasciatus). Marine Biotechnology, 24(1):163–173, February 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-022-10095-2.

Jin:2022:ADE

[1721] Yulin Jin, Tao Zhou, Wansheng Jiang, Ning Li, Xiaoyan Xu, Suxu Tan, Huitong Shi, Yujia Yang, Zihao Yuan, Wenwen Wang, Guyu Qin, Shikai Liu, Dongya Gao, Rex Dunham, and Zhanjiang Liu. Allelically and differentially expressed genes after infection of Edwardsiella ictaluri in channel catfish as determined by bulk segregant RNA-Seq. Marine Biotechnology, 24(1):174–189, February 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-022-10094-3.

Handayani:2022:NRG

[1722] Desy Putri Handayani, Alim Isnansetyo, Indah Istiqomah, and Jumina Jumina. New report: Genome mining untaps the antibiotics biosynthetic gene cluster of Pseudoalteromonas xiamenensis STKMTI.2 from a mangrove soil sediment. Marine Biotechnology, 24(1):190–202, February 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-022-10096-1.

Cao:2022:SCI

[1723] Xueqin Cao, Qian Zhang, Yanglu Zhu, Siju Li, Ying Cai, Pei Li, Deliang Liu, Yun Leng, Simin Ye, Zengmei Xu, Hao Li, Baochun Shen, Qiongfeng Liao, Lan Liu, and Zhiyong Xie. Structural characterization and immunoenhancing effects of a polysaccharide from the soft coral Lobophytum sarcophytoides. Marine Biotechnology, 24(1): 203–215, February 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10099-y.

Crandall:2022:CGR

[1724] Grace Crandall, Pamela C. Jensen, Samuel J. White, and Steven Roberts. Characterization of the gene repertoire and environmentally driven expression patterns in Tanner crab (Chionoecetes bairdi). Marine Biotechnology, 24(1):216–225, February 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-022-10100-8.

Kim:2022:GMR

[1725] Duck-Hyun Kim, Eunjin Byeon, Min-Sub Kim, Young Hwan Lee, Jun Chul Park, Atsushi Hagiwara, and Jae-Seong Lee. The genome of the marine rotifer Brachionus manjavacas: Genome-wide identification of 310 G protein-coupled receptor (GPCR) genes. Marine Biotechnology, 24(1):226–242, February 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-022-10102-6.

Man:2022:IEP

[1726] Chun Hung Man, Yohei Shimura, and Iwane Suzuki. Identification of extracellular proteases induced by nitrogen-limited conditions in the thraustochytrids Schizochytrium aggregatum ATCC 28209. Marine Biotechnology, 24(1):243–254, February 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-022-10103-5.

Barbosa:2022:TPM

[1727] David Aciole Barbosa, Bruno C. Araújo, Giovana Souza Branco, Alexandre S. Simeone, Alexandre W. S. Hilsdorf, Daniela L. Jabes, Luiz R. Nunes, Renata G. Moreira, and Fabiano B. Menegidio. Transcriptomic profiling and microsatellite identification in cobia (Rachycentron canadum), using high-throughput RNA sequencing. Marine Biotechnology, 24(1):255–262, February 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-021-10081-0.

Min:2022:HPM

[1728] Yue Min, Qi Li, and Hong Yu. Heme-peroxidase 2 modulated by POU2F1 and SOX5 is involved in pigmentation in Pacific oyster (Crassostrea gigas). Marine Biotechnology, 24(2):263–275, April 2022. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-022-10098-z.

Song:2022:CTG

[1729] Jingwei Song, James D. Austin, and Huiping Yang. Comparative transcriptomics of the Northern quahog Mercenaria mercenaria and southern quahog Mercenaria campechiensis in response to chronic heat stress. Marine Biotechnology, 24(2):276–292, April 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-022-10101-7.

Liu:2022:PDI

[1730] Qi Liu, Yushan Yao, Mahmoud A. A. Ibrahim, Ali Mahmoud El Halawany, Li Yang, and Xuewu Zhang. Production of dual inhibitory hydrolysate by enzymatic hydrolysis of squid processing by-product. Marine Biotechnology, 24(2):293–302, April 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-022-10104-4.

Chen:2022:ICM

[1731] Jian Chen, Ziqin Zhai, Lili Lu, Suping Li, Dan Guo, Lirong Bai, and Dahui Yu. Identification and characterization of miRNAs and their predicted mRNAs in the larval development of pearl oyster Pinctada fucata. Marine Biotechnology, 24(2):303–319, April 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-022-10105-3.

Hou:2022:CDP

[1732] Mingxi Hou, Ke Feng, Hongrui Luo, Yinjun Jiang, Wen Xu, Yongming Li, Yanlong Song, Ji Chen, Binbin Tao, Zuoyan Zhu, and Wei Hu. Complete depletion of primordial germ cells results in masculinization of Monopterus albus, a protogynous hermaphroditic fish. Marine Biotechnology, 24(2):320–334, April 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-022-10106-2.

Zhao:2022:CPP

[1733] Na Zhao, Lei Jia, Qiuxia Deng, Chunhua Zhu, and Bo Zhang. Comparative piRNAs profiles give a clue to transgenerational inheritance of sex-biased piRNAs in Cynoglossus semilaevis. Marine Biotechnology, 24 (2):335–344, April 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10109-z.

Duan:2022:LEG

[1734] Xuzhuo Duan, Xianze Jia, Kaishan Liang, Fengqi Huang, Jinhong Shan, Huitao Chen, Xinhe Ruan, Lihua Li, Huihong Zhao, and Qing Wang. Liposome-encapsulated Rec8 and Dmrt1 plasmids induce red-spotted grouper (Epinephelus akaara) testis maturation. Marine Biotechnology, 24(2):345–353, April 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10111-5.

Yu:2022:HSP

[1735] Fan Yu, Jian-Lin Li, Wen-Rong Feng, Yong-Kai Tang, Sheng-Yan Su, Pao Xu, and Huan Zhong. Heat shock procedure affects cell divisionassociated genes in gynogenetic manipulation. Marine Biotechnology, 24 (2):354–365, April 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10112-4.

Zhao:2022:HTS

[1736] Na Zhao, Jiamin Guo, Bo Zhang, Kai Liu, Yuting Liu, Yubang Shen, and Jiale Li. Heterogeneity of the tissue-specific mucosal microbiome of normal grass carp (Ctenopharyngodon idella). Marine Biotechnology, 24 (2):366–379, April 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10113-3.

Liu:2022:OSU

[1737] Chuang Liu, Haipeng Liu, Jingliang Huang, and Xin Ji. Optimized sensory units integrated in the chiton shell. Marine Biotechnology, 24 (2):380–392, April 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10114-2.

Kominami:2022:TPM

[1738] Sayaka Kominami, Hiroyuki Mizuta, and Toshiki Uji. Transcriptome profiling in the marine red alga Neopyropia yezoensis under light/dark cycle. Marine Biotechnology, 24(2):393–407, April 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-022-10121-3.

Liu:2022:DGN

[1739] Liangjie Liu, Tian Liu, Shaoxuan Wu, Yajuan Li, Huilan Wei, Lijing Zhang, Ya Shu, Yaxin Yang, Qiang Xing, Shi Wang, and Lingling Zhang. Discovery of Nanos1 and Nanos2/3 as germ cell markers during scallop gonadal development. Marine Biotechnology, 24(2): 408–416, April 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10124-0.

Fujihara:2022:PGC

[1740] Ryo Fujihara, Naoto Katayama, Sakiko Sadaie, Misako Miwa, Gabriela Angelica Sanchez Matias, Kensuke Ichida, Wataru Fujii, Kunihiko Naito, Makoto Hayashi, and Goro Yoshizaki. Production of germ cell-less rainbow trout by dead end gene knockout and their use as recipients for germ cell transplantation. Marine Biotechnology, 24(2):417–429, April 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-022-10128-w.

Cleverley:2022:CVO

[1741] Robert Cleverley, David Webb, Stuart Middlemiss, Phillip Duke, Anthony Clare, Keiju Okano, Colin Harwood, and Nick Aldred. Correction to: In vitro oxidative crosslinking of recombinant barnacle cyprid cement gland proteins. Marine Biotechnology, 24(2):430, April 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-022-10107-1. See

[1702].

Alviz-Gazitua:2022:MRB

[1742] P. Alviz-Gazitua, A. González, M. R. Lee, and C. P. Aranda. Molecular relationships in biofilm formation and the biosynthesis of exoproducts in Pseudoalteromonas spp. Marine Biotechnology, 24(3):431–447, June 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-022-10097-0.

Kaari:2022:ENB

[1743] Manigundan Kaari, Radhakrishnan Manikkam, and Abirami Baskaran. Exploring newer biosynthetic gene clusters in marine microbial prospecting. Marine Biotechnology, 24(3):448–467, June 2022. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-022-10118-y.

Maeda:2022:MIL

[1744] Yoshiaki Maeda and Tsuyoshi Tanaka. Molecular insights into lipoxygenases in diatoms based on structure prediction: a pioneering study on lipoxygenases found in Pseudo-nitzschia arenysensis and Fragilariopsis cylindrus. Marine Biotechnology, 24(3):468–479, June 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-022-10120-4.

Muhlia-Almazan:2022:UDP

[1745] Adriana Teresita Muhlia-Almazán and Analía Verónica Fernández-Gimenez. Understanding the digestive peptidases from crustaceans: from their biochemical basis and classical perspective to the biotechnological approach. Marine Biotechnology, 24(3):480–491, June 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-022-10122-2.

Amelia:2022:RAM

[1746] Tan Suet May Amelia, Ferr Angelus C. Suaberon, Johanne Vad, Afiq Durrani Mohd Fahmi, Jonel P. Saludes, and Kesaven Bhubalan. Recent advances of marine sponge-associated microorganisms as a source of commercially viable natural products. Marine Biotechnology, 24(3): 492–512, June 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10130-2.

Xing:2022:CCM

[1747] De Xing, Baofeng Su, Shangjia Li, Max Bangs, David Creamer, Michael Coogan, Jinhai Wang, Rhoda Simora, Xiaoli Ma, Darshika Hettiarachchi, Veronica Alston, Wenwen Wang, Andrew Johnson, Cuiyu Lu, Tasnuba Hasin, Zhenkui Qin, and Rex Dunham. CRISPR/Cas9-mediated transgenesis of the masu salmon (Oncorhynchus masou) elovl2 gene improves n-3 fatty acid content in channel catfish (Ictalurus punctatus). Marine Biotechnology, 24(3):513–523, June 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-022-10110-6.

Iijima:2022:VPS

[1748] Mariko Iijima, Jun Yasumoto, Kanami Mori-Yasumoto, Mina Yasumoto-Hirose, Akira Iguchi, Atsushi Suzuki, Nanami Mizusawa, Mitsuru Jimbo, Shugo Watabe, and Ko Yasumoto. Visualisation of phosphate in subcalicoblastic extracellular calcifying medium and on a skeleton of coral by using a novel probe, fluorescein-4-isothiocyanate-labelled alendronic acid. Marine Biotechnology, 24(3):524–530, June 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-022-10115-1.

Zhang:2022:SMP

[1749] Xin Zhang, Zehui Yin, Zhuojun Ma, Jian Liang, Zhen Zhang, Liping Yao, Xia Chen, Xiaojun Liu, and Rongqing Zhang. Shell matrix protein N38 of Pinctada fucata, inducing vaterite formation, extends the DING protein to the mollusca world. Marine Biotechnology, 24 (3):531–541, June 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10116-0.

Jahan:2022:GCE

[1750] Kifat Jahan, Zhihui Yin, Yanming Zhang, Xiwu Yan, and Hongtao Nie. Gene co-expression network analysis reveals the correlation patterns among genes in different temperature stress adaptation of Manila clam. Marine Biotechnology, 24(3):542–554, June 2022. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-022-10117-z.

Ji:2022:OPP

[1751] Zhehui Ji, Guangxiao Yao, Liang Jiang, and Shizhen Wang. One-pot purification and immobilization of phenylalanine dehydrogenase from Bacillus nanhaiensi by functional reduced graphene oxide. Marine Biotechnology, 24(3):555–565, June 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-022-10123-1.

Liu:2022:IUP

[1752] Feng Liu, Fei Sun, Gang Qiao Kuang, Le Wang, and Gen Hua Yue. The insertion in the 3′ UTR of Pmel17 is the causal variant for golden skin color in tilapia. Marine Biotechnology, 24(3):566–573, June 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-022-10125-z.

Gaudin-Zatylny:2022:INS

[1753] Céline Gaudin-Zatylny, Erwan Corre, Bruno Zanuttini, Maxime Endress, Benoît Bernay, Julien Pontin, Alexandre Leduc, and Joël Henry. Identification of a new set of polypeptidic sex pheromones from cuttlefish (Sepia officinalis). Marine Biotechnology, 24(3):574–587, June 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-022-10126-y.

Guan:2022:CTA

[1754] Wen-Zhi Guan, Kai Jiang, Xing-Lin Lai, Yao-Ting Dong, and Gao-Feng Qiu. Comprehensive transcriptome analysis of gonadal and somatic tissues for identification of sex-related genes in the largemouth bass Micropterus salmoides. Marine Biotechnology, 24(3):588–598, June 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-022-10127-x.

Qeshmi:2022:PNM

[1755] Fatemeh Izadpanah Qeshmi, Ahmad Homaei, Khosro Khajeh, Ehsan Kamrani, and Pedro Fernandes. Production of a novel marine Pseudomonas aeruginosa recombinant L-asparaginase: Insight on the structure and biochemical characterization. Marine Biotechnology, 24(3): 599–613, June 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10129-9.

Li:2022:AAN

[1756] Zhi Li, Fangwang Shen, Lili Song, and Shicui Zhang. Antifungal activity of NP20 derived from amphioxus midkine/pleiotrophin homolog against Aspergillus niger and Aspergillus fumigatus. Marine Biotechnology, 24 (3):614–625, June 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10131-1.

Breedon:2022:RAA

[1757] Sarah A. Breedon, Aakriti Gupta, and Kenneth B. Storey. Regulation of apoptosis and autophagy during anoxia in the freshwater crayfish, Faxonius virilis. Marine Biotechnology, 24(3):626–639, June 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-022-10132-0.

Qu:2022:TSA

[1758] Ang Qu, Yulin Bai, Xinyi Zhang, Junjia Zeng, Fei Pu, Linni Wu, Peng Xu, and Tao Zhou. Tissue-specific analysis of alternative splicing events and differential isoform expression in large yellow croaker (Larimichthys crocea) after Cryptocaryon irritans infection. Marine Biotechnology, 24 (3):640–654, June 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10133-z.

Lu:2022:CLG

[1759] Jianguo Lu, Dong Gao, Ying Sims, Wenyu Fang, Joanna Collins, James Torrance, Genmei Lin, Jingui Xie, Jian Liu, and Kerstin Howe. Chromosome-level genome assembly of Acanthopagrus latus provides insights into salinity stress adaptation of Sparidae. Marine Biotechnology, 24(3):655–660, June 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10119-x.

Datsomor:2022:MRB

[1760] Alex K. Datsomor, Gareth Gillard, Yang Jin, Rolf E. Olsen, and Simen R. Sandve. Molecular regulation of biosynthesis of long chain polyunsaturated fatty acids in Atlantic salmon. Marine Biotechnology, 24(4): 661–670, August 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10144-w.

Zheng:2022:HGC

[1761] Jian Zheng, Linlin Zhao, Xiang Zhao, Tianxiang Gao, and Na Song. High genetic connectivity inferred from whole-genome resequencing provides insight into the phylogeographic pattern of Larimichthys polyactis. Marine Biotechnology, 24(4):671–680, August 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-022-10134-y.

Yan:2022:DGM

[1762] Mengzhen Yan, Bijun Li, Jiaying Wang, Yulin Bai, Qiaozhen Ke, Tao Zhou, and Peng Xu. Disruption of mstn gene by CRISPR/Cas9 in large yellow croaker (Larimichthys crocea). Marine Biotechnology, 24 (4):681–689, August 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10135-x.

Kim:2022:SGC

[1763] Seong Cheol Kim, Hyeon Jeong Kim, Gi Eun Park, Chang Won Lee, Andriy Synytsya, Peter Capek, and Yong Il Park. Sulfated glucuronorhamnoxylan from Capsosiphon fulvescens ameliorates osteoporotic bone resorption via inhibition of osteoclastic cell differentiation and function In Vitro and In Vivo. Marine Biotechnology, 24(4):690–705, August 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-022-10136-w.

Goecke:2022:NDC

[1764] Franz Goecke, Amelia Gómez Garreta, Rafael Martín-Martín, Jordi Rull Lluch, Jorunn Skjermo, and Åshild Ergon. Nuclear DNA content variation in different life cycle stages of sugar kelp, Saccharina latissima. Marine Biotechnology, 24(4):706–721, August 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-022-10137-9.

Xiang:2022:RRD

[1765] Xueyan Xiang, Davide Poli, Bernard M. Degnan, and Sandie M. Degnan. Ribosomal RNA-depletion provides an efficient method for successful dual RNA-Seq expression profiling of a marine sponge holobiont. Marine Biotechnology, 24(4):722–732, August 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-022-10138-8.

Yang:2022:ESP

[1766] Tianjing Yang, Darryl Joy Juntila, Naomichi Fujihara, Takashi Inada, Kohei Yoneda, and Iwane Suzuki. Enhancement of squalene production by constitutive expression of the 3-Hydroxy-3-Methylglutaryl-CoA reductase in Aurantiochytrium sp. 18W-13a. Marine Biotechnology, 24 (4):733–743, August 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10139-7.

Song:2022:GWI

[1767] Kai Song. Genome-wide identification of long non-coding RNAs in Crassostrea gigas and their association with heat stress. Marine Biotechnology, 24(4):744–752, August 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-022-10140-0.

Cao:2022:IAM

[1768] Jia-Yi Cao, Si-Min Xu, Ying-Ying Wang, Xie-Dan Long, Shuo-Nan Ma, Cheng-Xu Zhou, Ji-Lin Xu, and Xiao-Jun Yan. Integrated analyses of miRNome and transcriptome reveal the critical role of miRNAs toward heat stress response in Isochrysis galbana. Marine Biotechnology, 24(4): 753–762, August 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10141-z.

Song:2022:LLC

[1769] Chaowei Song, Wanbo Li, and Zhiyong Wang. The landscape of liver chromatin accessibility and conserved non-coding elements in Larimichthys crocea, Nibea albiflora, and Lateolabrax maculatus. Marine Biotechnology, 24(4):763–775, August 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-022-10142-y.

Wu:2022:MFR

[1770] Shenji Wu, Jinqiang Huang, Yongjuan Li, Zhe Liu, and Lu Zhao. MiR-382 functions on the regulation of melanogenesis via targeting dct in rainbow trout (Oncorhynchus mykiss). Marine Biotechnology, 24(4): 776–787, August 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10143-x.

Maeda:2022:CSG

[1771] Yoshiaki Maeda, Ryosuke Kobayashi, Kahori Watanabe, Tomoko Yoshino, Chris Bowler, Mitsufumi Matsumoto, and Tsuyoshi Tanaka. Chromosome-scale genome assembly of the marine oleaginous diatom Fistulifera solaris. Marine Biotechnology, 24(4):788–800, August 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-022-10147-7.

Buscaglia:2022:MES

[1772] Manon Buscaglia, Fabienne Guérard, Philippe Roquefort, Thierry Aubry, Marilyne Fauchon, Yannick Toueix, Valérie Stiger-Pouvreau, Claire Hellio, and Gwenaëlle Le Blay. Mechanically enhanced Salmo salar gelatin by enzymatic cross-linking: Premise of a bioinspired material for food packaging, cosmetics, and biomedical applications. Marine Biotechnology, 24(4):801–819, August 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-022-10150-y.

Fan:2022:RSP

[1773] Yaoran Fan, Jianbin Feng, Nan Xie, Feiyue Ling, Zefei Wang, Keyi Ma, Xueming Hua, and Jiale Li. RNA-seq provides novel insights into response to acute salinity stress in oriental river prawn Macrobrachium nipponense. Marine Biotechnology, 24(4):820–829, August 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-022-10151-x.

deMuizon:2022:OMS

[1774] Capucine Jourdain de Muizon, Donata Iandolo, Dung Kim Nguyen, Ali Al-Mourabit, and Marthe Rousseau. Organic matrix and secondary metabolites in nacre. Marine Biotechnology, 24(5):831–842, October 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-022-10145-9.

Coogan:2022:IGH

[1775] Michael Coogan, Veronica Alston, Baofeng Su, Karim Khalil, Ahmed Elaswad, Mohd Khan, Andrew Johnson, De Xing, Shangjia Li, Jinhai Wang, Rhoda M. C. Simora, Cuiyu Lu, Patrick Page-McCaw, Wenbiao Chen, Max Michel, Wenwen Wang, Darshika Hettiarachchi, Tasnuba Hasin, Ian A. E. Butts, Roger D. Cone, and Rex A. Dunham. Improved growth and high inheritance of melanocortin-4 receptor (mc4r) mutation in CRISPR/Cas-9 gene-edited channel catfish, Ictalurus punctatus. Marine Biotechnology, 24(5):843–855, October 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-022-10146-8.

Zhang:2022:TAL

[1776] Wei Zhang, Xiaoyan Xu, Jiale Li, and Yubang Shen. Transcriptomic analysis of the liver and brain in grass carp (Ctenopharyngodon idella) under heat stress. Marine Biotechnology, 24(5):856–870, October 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-022-10148-6.

Ide:2022:ECC

[1777] Keigo Ide, Yoshikatsu Nakano, Michihiro Ito, Yohei Nishikawa, Hiroyuki Fujimura, and Haruko Takeyama. The effect of co-culture of two coral species on their bacterial composition under captive environments. Marine Biotechnology, 24(5):871–881, October 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-022-10149-5. See correction

[1802].

Windarto:2022:FRS

[1778] Seto Windarto, Meng-Chou Lee, Happy Nursyam, and Jue-Liang Hsu. First report of screening of novel angiotensin-i converting enzyme inhibitory peptides derived from the red alga Acrochaetium sp. Marine Biotechnology, 24(5):882–894, October 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-022-10152-w.

Kijima:2022:AAD

[1779] Yusuke Kijima, Wang Wantong, Yoji Igarashi, Kazutoshi Yoshitake, Shuichi Asakawa, Yutaka Suzuki, Shugo Watabe, and Shigeharu Kinoshita. Age-associated different transcriptome profiling in zebrafish and rats: an insight into the diversity of vertebrate aging. Marine Biotechnology, 24(5):895–910, October 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-022-10153-9.

Tan:2022:CMA

[1780] Chao Tan, Chenyu Shi, Yin Li, Wen Teng, Yongjing Li, Huiru Fu, Liting Ren, Hong Yu, Qi Li, and Shikai Liu. Comparative methylome analysis reveals epigenetic signatures associated with growth and shell color in the Pacific oyster, Crassostrea gigas. Marine Biotechnology, 24(5): 911–926, October 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10154-8.

Suo:2022:GWA

[1781] Ning Suo, Yidi Wu, Zhixiong Zhou, Qian He, Huaqiang Bai, Huanling Lin, Qiaozhen Ke, and Peng Xu. Genome-wide association and expression analysis revealed the candidate variants and molecular underpinnings of cold-stress response in large yellow croaker. Marine Biotechnology, 24(5):927–941, October 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-022-10155-7.

Zheng:2022:IRS

[1782] Yun Zheng, Shanjie Zha, Weifeng Zhang, Yinghui Dong, Jing He, Zhihua Lin, and Yongbo Bao. Integrated RNA-seq and RNAi analysis of the roles of the Hsp70 and SP genes in red-shell Meretrix meretrix tolerance to the pathogen Vibrio parahaemolyticus. Marine Biotechnology, 24(5): 942–955, October 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10156-6.

Wu:2022:CRA

[1783] Liqun Wu, Zhijie Yin, Zhijin Zheng, Yijun Tang, and Songlin Guo. Comprehensive relationship analysis of the long noncoding RNAs (lncRNAs) and the target mRNAs in response to the infection of Edwardsiella anguillarum in European eel (Anguilla anguilla) inoculated with Freund’s adjuvant. Marine Biotechnology, 24(5):956–968, October 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-022-10157-5.

Ma:2022:IMS

[1784] Yuting Ma, Yongshuang Xiao, Zhizhong Xiao, Yanduo Wu, Haixia Zhao, and Jun Li. Identification of male-specific molecular marker and development of PCR-based genetic sex identification technique in spotted knifejaw (Oplegnathus punctatus). Marine Biotechnology, 24(5): 969–978, October 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10160-w.

Seo:2022:NSS

[1785] Haneul Seo, Andre Ditya Maulana Lubis, and Sukchan Lee. A novel specific single-chain variable fragment diagnostic system for viral hemorrhagic septicemia virus. Marine Biotechnology, 24(5):979–990, October 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-022-10161-9.

Tzachor:2022:EIL

[1786] Asaf Tzachor, Asger Smidt-Jensen, Alfons Ramel, and Margrét Geirsdóttir. Environmental impacts of large-scale spirulina (Arthrospira platensis) production in Hellisheidi Geothermal Park Iceland: Life cycle assessment. Marine Biotechnology, 24(5):991–1001, October 2022. CO-DEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL

https://link.springer.com/article/10.1007/s10126-022-10162-

8. See correction [1803].

Zhang:2022:SMR

[1787] Yuan Zhang, Fangrui Lou, Jianwei Chen, Zhiqiang Han, Tianyan Yang, Tianxiang Gao, and Na Song. Single-molecule real-time (SMRT) sequencing facilitates transcriptome research and genome annotation of the fish Sillago sinica. Marine Biotechnology, 24(5):1002–1013, October 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-022-10163-7.

Endres:2022:MSM

[1788] Carla Teresinha Endres, Graziela Vargas Rigo, Luciana Angelo Loges, Melissa Fontes Landell, Denise Brentan Silva, Alexandre José Macedo, and Tiana Tasca. Mass spectrometry metabolomics approach reveals anti-Trichomonas vaginalis scaffolds from marine fungi. Marine Biotechnology, 24(5):1014–1022, October 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-022-10164-6.

Wu:2022:LMR

[1789] Ping Wu, Pengzheng Yong, Zhanxiong Zhang, Rui Xu, Renjie Shang, Jun Shi, Jianshe Zhang, Pengpeng Bi, Elizabeth Chen, and Shaojun Du.

Loss of myomixer results in defective myoblast fusion, impaired muscle growth, and severe myopathy in zebrafish. Marine Biotechnology, 24(5): 1023–1038, October 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10159-3.

Zhang:2022:CTA

[1790] Yibo Zhang, Weiliang Shen, Jie Ding, Xinming Gao, Xiongfei Wu, and Junquan Zhu. Comparative transcriptome analysis of head kidney of Aeromonas hydrophila-infected hypoxia-tolerant and normal large yellow croaker. Marine Biotechnology, 24(6):1039–1054, December 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-022-10158-4.

Wan:2022:CTRb

[1791] Haifu Wan, Yinzhen Sheng, Ziping Zhang, Xiwei Jia, and Yilei Wang. Comparative transcriptome reveals the potential modulation mechanisms of Spdsx affecting ovarian development of Scylla paramamosain. Marine Biotechnology, 24(6):1055–1065, December 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-022-10165-5.

Yang:2022:DSI

[1792] Jeong In Yang and Ki Hong Kim. Display of Streptococcus iniae αenolase on the surface of virus-like particles (VLPs) of nervous necrosis virus (NNV) using SpyTag/SpyCatcher. Marine Biotechnology, 24(6): 1066–1072, December 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10166-4.

Feng:2022:PIF

[1793] Xiaoqing Feng, Haihong Chen, Baoheng Xiao, Qiong Wu, Jingyu Zhang, Ni Zhang, Pingping Li, Lu Wang, Jingru Yin, and Zhenghong Sui. Ploidy identification by flow cytometry and application of the method to characterize seasonal ploidy variation of wild populations of the red alga Gracilariopsis lemaneiformis. Marine Biotechnology, 24(6):1073– 1083, December 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10167-3.

Yang:2022:SAS

[1794] Zituo Yang, Yepin Yu, Le Wang, Sek-Man Wong, and Gen Hua Yue. Silencing Asian seabass gab3 inhibits nervous necrosis virus replication.

Marine Biotechnology, 24(6):1084–1093, December 2022. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-022-10169-1.

Zhou:2022:DTR

[1795] Shun Zhou, Xujia Zheng, Zongrui Yang, Qing Huang, Jingyuan Yi, Lin Su, Baoshan Guo, and Yunji Xiu. Development of two recombinase polymerase amplification EXO (RPA-EXO) and lateral flow dipstick (RPA-LFD) techniques for the rapid visual detection of Aeromonas salmonicida. Marine Biotechnology, 24(6):1094–1109, December 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-022-10170-8.

Kumar:2022:IFD

[1796] S. Santhosh Kumar, Snehal Jamalpure, A. Nafeez Ahmed, G. Taju, S. Vimal, S. Abdul Majeed, S. Suryakodi, Syed Rahamathulla, Kishore M. Paknikar, Jyutika M. Rajwade, and A. S. Sahul Hameed. An indigenous, field-deployable, lateral flow immunochromatographic assay rapidly detects infectious myonecrosis in shrimp, Litopenaeus vannamei. Marine Biotechnology, 24(6):1110–1124, December 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-022-10172-6.

Ma:2022:EAA

[1797] Fang Ma, Ruilin Ma, Yali Zou, and Lei Zhao. Effect of astaxanthin on the antioxidant capacity and intestinal microbiota of Tsinling Lenok trout (Brachymystax lenok tsinlingensis). Marine Biotechnology, 24(6): 1125–1137, December 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10175-3.

Chen:2022:GWA

[1798] Geng Chen, Ying Zhou, Xiaomu Yu, Junru Wang, Weiwei Luo, Meixia Pang, and Jingou Tong. Genome-wide association study reveals SNPs and candidate genes related to growth and body shape in bighead carp (Hypophthalmichthys nobilis). Marine Biotechnology, 24(6):1138– 1147, December 2022. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10176-2.

Jia:2022:SAB

[1799] Li Jia, Yabiao Yu, Jinyang Zheng, Hao Zhou, Qiang Liu, Wei Wang, Xinxin Liu, Xiuming Zhang, Dongtao Ge, Wei Shi, and Yanan Sun.

Self-assembling bioadhesive inspired by the fourth repetitive sequence of Balanus albicostatus cement protein 20 kDa (Bal cp-20 k). Marine Biotechnology, 24(6):1148–1157, December 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-022-10177-1.

Oyama:2022:CTD

[1800] Hikaru Oyama, Masaaki Ito, Rei Suo, Naoko Goto-Inoue, Mizuki Morisasa, Tsukasa Mori, Haruo Sugita, Tetsushi Mori, Ryota Nakahigashi, Masaatsu Adachi, Toshio Nishikawa, and Shiro Itoi. Changes in tissue distribution of tetrodotoxin and its analogues in association with maturation in the toxic flatworm, Planocera multitentaculata. Marine Biotechnology, 24(6):1158–1167, December 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-022-10179-z.

Goel:2022:ISS

[1801] Nikky Goel, Rajendra Singh, Seema Sood, and Sunil Kumar Khare. Investigation of Streptomyces sp. strain EMB24 secondary metabolite profile has unraveled its extraordinary antibacterial potency against drug–resistant bacteria. Marine Biotechnology, 24(6):1168–1175, December 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-022-10168-2.

Ide:2022:CEC

[1802] Keigo Ide, Yoshikatsu Nakano, Michihiro Ito, Yohei Nishikawa, Hiroyuki Fujimura, and Haruko Takeyama. Correction: The effect of co-culture of two coral species on their bacterial composition under captive environments. Marine Biotechnology, 24(6):1176, December 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-022-10171-7. See

[1777].

Tzachor:2022:CEI

[1803] Asaf Tzachor, Asger Smidt-Jensen, Alfons Ramel, and Margrét Geirsdóttir. Correction to: Environmental impacts of large-scale spirulina (Arthrospira platensis) production in Hellisheidi Geothermal Park Iceland: Life cycle assessment. Marine Biotechnology, 24(6):1177, December 2022. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-022-10173-5. See [1786].

Bomkamp:2023:DMM

[1804] Claire Bomkamp, Lisa Musgrove, Diana M. C. Marques, Gonçalo F. Fernando, Frederico C. Ferreira, and Elizabeth A. Specht. Differentiation and maturation of muscle and fat cells in cultivated seafood: Lessons from developmental biology. Marine Biotechnology, 25(1):1– 29, February 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10174-4.

Dai:2023:CBC

[1805] Wenfang Dai, Jing Ye, Qinggang Xue, Sheng Liu, Hongqiang Xu, Minhai Liu, and Zhihua Lin. Changes in bacterial communities of kumamoto oyster larvae during their early development and following Vibrio infection resulting in a mass mortality event. Marine Biotechnology, 25(1): 30–44, February 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10178-0.

Guo:2023:ITD

[1806] Ziqi Guo, Jiabao Zuo, Jianbin Feng, Jiale Li, Shupeng Zhang, and Keyi Ma. Impact of titanium dioxide–graphene oxide (TiO_{2}-GO) composite nanoparticle on the juveniles of the giant river prawn, Macrobrachium rosenbergii: Physio–biochemistry and transcriptional response. Marine Biotechnology, 25(1):45–56, February 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-022-10180-6.

Bai:2023:IEA

[1807] Yulin Bai, Xintong Chen, Ang Qu, Yue Liu, Ji Zhao, Qiaozhen Ke, Fei Pu, Linni Wu, Hongshu Chi, Hui Gong, Tao Zhou, and Peng Xu. Identification and expression analysis of LncRNAs reveal the immune mechanism of visceral white-nodules disease resistance in large yellow croaker. Marine Biotechnology, 25(1):57–69, February 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-022-10181-5.

Wei:2023:GWE

[1808] Xin Wei, Miao Zhang, Zhe Chi, Guang-Lei Liu, and Zhen-Ming Chi. Genome-wide editing provides insights into role of unsaturated fatty acids in low temperature growth of the psychrotrophic yeast Metschnikowia bicuspidata var. australis W7-5. Marine Biotechnology, 25(1):70–82, February 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-

2236 (electronic). URL https://link.springer.com/article/10.
1007/s10126-022-10182-4.

Schwaner:2023:MFA

[1809] Caroline Schwaner, Sarah Farhat, Michelle Barbosa, Isabelle Boutet, Arnaud Tanguy, Emmanuelle Pales Espinosa, and Bassem Allam. Molecular features associated with resilience to ocean acidification in the northern quahog, Mercenaria mercenaria. Marine Biotechnology, 25(1): 83–99, February 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10183-3.

Xiao:2023:BNS

[1810] Baoheng Xiao, Yiyi Hu, Xiaoqing Feng, and Zhenghong Sui. Breeding of new strains of Gracilariopsis lemaneiformis with high agar content by ARTP mutagenesis and high osmotic pressure screening. Marine Biotechnology, 25(1):100–108, February 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-022-10184-2.

Fang:2023:NRC

[1811] Jianhao Fang, Luqi Yan, Minghui Tan, Ganghui Li, Yingyin Liang, and Kuntai Li. Nitrogen removal characteristics of a marine denitrifying Pseudomonas stutzeri BBW831 and a simplified strategy for improving the denitrification performance under stressful conditions. Marine Biotechnology, 25(1):109–122, February 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-022-10185-1.

Ding:2023:TTD

[1812] Jie Ding, Daojun Tang, Yibo Zhang, Xinming Gao, Chen Du, Weiliang Shen, Shan Jin, and Junquan Zhu. Transcriptomes of testes at different developmental stages in the Opsariichthys bidens predict key genes for testis development and spermatogenesis. Marine Biotechnology, 25(1): 123–139, February 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10186-0.

Xu:2023:DFF

[1813] Hao Xu, Wenbo Wang, Zhentao Nie, Xiaomin Miao, and Yun Li. Delayed first feeding chronically impairs larval fish growth performance, hepatic lipid metabolism, and visceral lipid deposition at the mouth–opening stage. Marine Biotechnology, 25(1):140–149, February 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-022-10187-z.

Lin:2023:TDM

[1814] Yi Long Lin, Zong Xian Zhu, Chun Hui Ai, Ying Ying Xiong, Tong De Liu, Hao Ran Lin, and Jun Hong Xia. Transcriptome and DNA methylation responses in the liver of yellowfin seabream under starvation stress. Marine Biotechnology, 25(1):150–160, February 2023. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-022-10188-y.

Mohindra:2023:TAR

[1815] Vindhya Mohindra, Labrechai Mog Chowdhury, Nishita Chauhan, Alisha Paul, Rajeev Kumar Singh, Basdeo Kushwaha, Rajesh Kumar Maurya, Kuldeep K. Lal, and J. K. Jena. Transcriptome analysis revealed osmoregulation related regulatory networks and hub genes in the gills of hilsa shad, Tenualosa ilisha, during the migratory osmotic stress. Marine Biotechnology, 25(1):161–173, February 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-022-10190-4.

Guo:2023:DEH

[1816] Ximing Guo, Jonathan B. Puritz, Zhenwei Wang, Dina Proestou, Standish Allen, Jessica Small, Klara Verbyla, Honggang Zhao, Jaime Haggard, Noah Chriss, Dan Zeng, Kathryn Lundgren, Bassem Allam, David Bushek, Marta Gomez-Chiarri, Matthew Hare, Christopher Hollenbeck, Jerome La Peyre, Ming Liu, Katie E. Lotterhos, Louis Plough, Paul Rawson, Scott Rikard, Eric Saillant, Robin Varney, Gary Wikfors, and Ami Wilbur. Development and evaluation of high-density SNP arrays for the eastern oyster Crassostrea virginica. Marine Biotechnology, 25(1): 174–191, February 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10191-3.

Liu:2023:CEP

[1817] Jianlin Liu, Hui Li, Xuanyu Zhang, Lin Yue, Wei Lu, Shaohua Ma, Ziyu Zhu, Dong Wang, Hu Zhu, and Jiqian Wang. Cost-efficient production of the sphingan WL gum by Sphingomonas sp. WG using molasses and sucrose as the carbon sources. Marine Biotechnology, 25(1): 192–203, February 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10193-1.

Wang:2023:HSR

[1818] Yiran Wang, Xinrui Huang, Qiao Zhou, Yuxin Tian, Jinmei Zuo, Zengzhi Yuan, Yichen Liu, Ju Li, and Jinsheng Sun. Hippo signaling regulates blastema formation during limb regeneration in Chinese mitten crab (Eriocheir sinensis). Marine Biotechnology, 25(1):204– 213, February 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-022-10194-0.

Sourani:2023:MIR

[1819] Zahra Sourani, Sadegh Shirian, Shafigh Shafiei, Nadia Mosayebi, and Amin Nematollahi. Modulation of immune-related gene expressions in zebrafish (Danio rerio) by dietary purslane (Portulaca oleracea) extract. Marine Biotechnology, 25(1):214–221, February 2023. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-022-10195-z.

Bravo:2023:URD

[1820] Scarleth Bravo, Francisco Leiva, Javier Moya, Osiel Guzman, and Rodrigo Vidal. Unveiling the role of dynamic alternative splicing modulation after infestation with sea lice (Caligus rogercresseyi) in Atlantic salmon. Marine Biotechnology, 25(2):223–234, April 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-023-10196-6.

Jiang:2023:TAR

[1821] Gaowei Jiang, Yin Li, Geng Cheng, Kunyin Jiang, Jianmin Zhou, Chengxun Xu, Lingfeng Kong, Hong Yu, Shikai Liu, and Qi Li. Transcriptome analysis of reciprocal hybrids between Crassostrea gigas and C. angulata reveals the potential mechanisms underlying thermo-resistant heterosis. Marine Biotechnology, 25(2):235–246, April 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-023-10197-5.

Chen:2023:LTA

[1822] Chunna Chen, Bo Zhou, Jue Lin, Quan Gong, Fei Xu, Zhengyi Li, and Yingying Huang. Liver transcriptome analysis reveals energy regulation and functional impairment of Onychostoma sima during starvation. Marine Biotechnology, 25(2):247–258, April 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-023-10201-y.

Sousa:2023:PSM

[1823] Karolyne dos Santos Jorge Sousa, Júlia Risso Parisi, Amanda de Souza, Matheus de Almeida Cruz, Rogério Erbereli, Jonas de Araújo Silva, Giovanna do Espirito Santo, Gustavo Oliva do Amaral, Cintia Cristina Santi Martignago, Carlos Alberto Fortulan, Renata Neves Granito, and Ana Claudia Muniz Renno. 3D printed scaffolds manufactured with biosilica from marine sponges for bone healing in a cranial defect in rats. Marine Biotechnology, 25(2):259–271, April 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-023-10202-x.

Nakayasu:2023:DRP

[1824] Mana Nakayasu, Momoka Amano, Tsuyoshi Tanaka, Ginga Shimakawa, and Yusuke Matsuda. Different responses of photosynthesis to nitrogen starvation between highly oil-accumulative diatoms, Fistulifera solaris and Mayamaea sp. JPCC CTDA0820. Marine Biotechnology, 25 (2):272–280, April 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10203-w.

Zhang:2023:GAP

[1825] Changqing Zhang, Ziheng Ren, and Zhiyuan Gong. Generation of albino phenotype in ornamental fish by CRISPR/Cas9-mediated genome editing of slc45a2 gene. Marine Biotechnology, 25(2):281–290, April 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-023-10204-9.

Cheng:2023:PPA

[1826] Jianxin Cheng, Yuqing Xia, Cheng Zhou, Xiaohao Li, and Pengfei Liu. Proteomics and phosphoproteomic analysis to identify spleen of Takifugu rubripes infected Cryptocaryon irritans. Marine Biotechnology, 25 (2):291–313, April 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10205-8.

Kang:2023:APN

[1827] Ji Young Kang, Ha-Yeon Song, and Jung-Mi Kim. Agarolytic pathway in the newly isolated Aquimarina sp. bacterial strain ERC-38 and characterization of a putative β-agarase. Marine Biotechnology, 25(2): 314–327, April 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10206-7.

Zhao:2023:FST

[1828] Zhenfang Zhao, Guanyu Hu, Zhou Fang, Jinhui Li, Bilin Liu, and Xinjun Chen. Feeding strategies and trophic niche divergence of three groups of Dosidicus gigas off Peru: Based on stable carbon and nitrogen isotopes and morphology of feeding apparatuses. Marine Biotechnology, 25 (2):328–339, April 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10207-6.

Wang:2023:EDD

[1829] Xinxing Wang, Rihao Cong, Ao Li, Wei Wang, Guofan Zhang, and Li Li. Experimental DNA demethylation reduces expression plasticity and thermal tolerance in Pacific oysters. Marine Biotechnology, 25 (3):341–346, June 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10208-5.

Gokcek:2023:CSN

[1830] Emel Özcan Gökçek, Raziye Işık, Bilge Karahan, and Kutsal Gamsız. Characterisation of single nucleotide polymorphisms and haplotypes of MSTN associated with growth traits in European sea bass (Dicentrarchus labrax). Marine Biotechnology, 25(3):347–357, June 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-023-10211-w.

Araujo:2023:PFM

[1831] Bruno C. Araújo, David Aciole Barbosa, Renato M. Honji, Giovana S. Branco, Fabiano B. Menegidio, Victor H. Marques, Renata G. Moreira, Marcelo V. Kitahara, Artur N. Rombenso, Paulo H. de Mello, and Alexandre W. S. Hilsdorf. Post-feeding molecular responses of cobia (Rachycentron canadum): RNA-sequencing as a tool to evaluate postprandial effects in hepatic lipid metabolism. Marine Biotechnology, 25 (3):358–371, June 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10209-4.

Guo:2023:AAS

[1832] Songlin Guo, Minxia Chen, Wanbo Li, Qijuan Wan, and Ming Xu. Analysis of alternative splicing and long noncoding RNAs after the Edwardsiella anguillarum infected the immunized European eels (Anguilla anguilla) revealed the role of outer membrane protein a in OmpA subunit vaccine. Marine Biotechnology, 25(3):372–387, June 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-023-10210-x.

Yan:2023:MCU

[1833] Fengying Yan, Xinxun Xiao, Chen Long, Lin Tang, Chongwei Wang, Mingqing Zhang, Jin Zhang, Haoran Lin, Hai Huang, Yong Zhang, and Shuisheng Li. Molecular characterization of U6 promoters from orange-spotted grouper (Epinephelus coioides) and its application in DNA vector-based RNAi technology. Marine Biotechnology, 25(3): 388–402, June 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10212-9.

Liu:2023:CEG

[1834] Mingtao Liu, Qihui Zhu, Huan Li, Ruiyi Chen, Weihua Hu, Simiao Liu, and Dongdong Xu. Characterization of early gonadal differentiation and induction of sex reversal in the rock bream Oplegnathus fasciatus. Marine Biotechnology, 25(3):403–414, June 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-023-10213-8.

Gunathilaka:2023:UMS

[1835] M. D. T. L. Gunathilaka. Utilization of marine seaweeds as a promising defense against COVID-19: a mini-review. Marine Biotechnology, 25 (3):415–427, June 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10214-7.

Yang:2023:MPN

[1836] Min Yang, Xin lei Li, Yu ting Zhang, Yue wen Deng, and Yu Jiao. miR-10a-3p participates in nacre formation in the pearl oyster Pinctada fucata martensii by targeting NPY. Marine Biotechnology, 25 (3):428–437, June 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10216-5.

Moghadam:2023:IGL

[1837] Hooman K. Moghadam, Brede Fannemel, Ingunn Thorland, Carlos Lozano, and Borghild Hillestad. Identification and genomic localization of autosomal sdy locus in a population of Atlantic salmon (Salmo salar). Marine Biotechnology, 25(3):438–446, June 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-023-10217-4.

Jiang:2023:ESR

[1838] Qun Jiang, Peng Ji, Shiqi Ao, Xiaojian Gao, and Xiaojun Zhang. Effects of starvation and refeeding on glucose metabolism and immune responses in Macrobrachium rosenbergii. Marine Biotechnology, 25(3): 447–462, June 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10218-3.

Xu:2023:ETI

[1839] Simin Xu, Jiayi Cao, Minnan Wu, Yijun Xu, Yuanyuan Wu, Kaixi Shang, Bin Ma, Lin Zhang, Deshui Chen, Xinyu Liu, Xiaojun Yan, and Jilin Xu. Enhancing the thermotolerance of Isochrysis zhangjiangensis through co-culturing with Algoriphagus marincola. Marine Biotechnology, 25 (3):463–472, June 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10219-2.

Razak:2023:NTL

[1840] Muhammad Raznisyafiq Razak, Ahmad Zaharin Aris, Fatimah Md Yusoff, Zetty Norhana Balia Yusof, Sang Don Kim, and Kyoung Woong Kim. De Novo transcriptomic and life-history responses of Moina micrura under stress environment conditions. Marine Biotechnology, 25 (3):473–487, June 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10220-9.

Koiwai:2023:SSA

[1841] Keiichiro Koiwai, Hidehiro Kondo, and Ikuo Hirono. scRNA-seq analysis of hemocytes of penaeid shrimp under virus infection. Marine Biotechnology, 25(3):488–502, June 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-023-10221-8.

Yamamoto:2023:NEP

[1842] Hana Yamamoto, Nanami Shimomura, Kazuma Oura, and Yasushi Hasegawa. Nacre extract from pearl oyster shell prevents D-galactose-induced brain and skin aging. Marine Biotechnology, 25(4):503–518, August 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-022-10192-2.

Cochereau:2023:HEB

[1843] Bastien Cochereau, Yoran Le Strat, Qiaolin Ji, Audrey Pawtowski, Ludovic Delage, Amélie Weill, Lisa Mazéas, Cécile Hervé, Gaëtan Burgaud, Nina Gunde-Cimerman, Yves François Pouchus, Nathalie Demont-Caulet, Catherine Roullier, and Laurence Meslet-Cladiere. Heterologous expression and biochemical characterization of a new chloroperoxidase isolated from the deep-sea hydrothermal vent black yeast Hortaea werneckii UBOCC-A-208029. Marine Biotechnology, 25(4):519– 536, August 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10222-7.

Li:2023:EDC

[1844] Zhuanzhuan Li, Chengxun Xu, Hong Yu, Lingfeng Kong, Shikai Liu, and Qi Li. Effects of dietary cystine and tyrosine supplementation on melanin synthesis in the Pacific oyster (Crassostrea gigas). Marine Biotechnology, 25(4):537–547, August 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-023-10223-6.

Scaranto:2023:GSD

[1845] Bianca Maria Soares Scaranto, Josiane Ribolli, Graziela Cleuza Vieira, João Paulo Ramos Ferreira, Carlos Henrique Araujo de Miranda Gomes, and Claudio Manoel Rodrigues de Melo. Genetic structure and diversity in wild and cultivated populations of the mangrove oyster Crassostrea gasar from Southern Brazil. Marine Biotechnology, 25(4): 548–556, August 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10224-5.

Yu:2023:TAS

[1846] Chengchen Yu, Minglin Wu, Yuchen Jiang, Xiaoyan Xu, Jiale Li, and Yubang Shen. Transcriptome analysis of the spleen provides insight into the immune regulation of GCRV resistance in grass carp (Ctenopharyngodon idella). Marine Biotechnology, 25(4):557–566, August 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-023-10225-4.

Yue:2023:WGM

[1847] Yanfeng Yue, Yabing Wang, Bianbian Zhang, Jiao Zeng, Qian Wang, Cuihua Wang, and Shiming Peng. Whole-genome methylation sequencing of large yellow croaker (Larimichthys crocea) liver under hypoxia and acidification stress. Marine Biotechnology, 25(4):567–579, August 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-023-10226-3.

Weise:2023:SPG

[1848] Ellen M. Weise, Mallory Van Wyngaarden, Cornelia Den Heyer, Joanna Mills Flemming, Tony Kess, Anthony L. Einfeldt, Jonathan A. D. Fisher, Reina Ditta, Guillaume Pare, and Daniel E. Ruzzante. SNP panel and genomic sex identification in Atlantic halibut (Hippoglossus hippoglossus). Marine Biotechnology, 25(4):580–587, August 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-023-10227-2.

Zhang:2023:HPT

[1849] Ruiqi Zhang, Xiang Shi, Zhe Liu, Jun Sun, Tongzhen Sun, and Mingquan Lei. Histological, physiological and transcriptomic analysis reveal the acute alkalinity stress of the gill and hepatopancreas of Litopenaeus vannamei. Marine Biotechnology, 25(4):588–602, August 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-023-10228-1.

Wang:2023:FGP

[1850] Jiaying Wang, Yulin Bai, Xiaoqing Zou, Chengyu Li, Junyi Yang, QiaoZhen Ke, Ji Zhao, Tao Zhou, and Peng Xu. First genomic prediction of single-step models in large yellow croaker. Marine Biotechnology, 25 (4):603–611, August 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10229-0.

Bustos:2023:SVG

[1851] Paulina Bustos, Paulina Schmitt, Donald I. Brown, and Rodolfo Farlora. Silencing of the Vasa gene by RNA interference affects embryonic development and reproductive output in the sea louse Caligus rogercresseyi. Marine Biotechnology, 25(4):612–623, August 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-023-10232-5.

Ibrahim:2023:WTA

[1852] Salifu Ibrahim, Chuangye Yang, Chenyang Yue, Xinyu Song, Yuewen Deng, Qi Li, and Wengang Lü. Whole transcriptome analysis reveals the global molecular responses of mRNAs, lncRNAs, miRNAs, circR-NAs, and their ceRNA networks to salinity stress in Hong Kong oysters, Crassostrea hongkongensis. Marine Biotechnology, 25(4):624–641, August 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-023-10234-3.

Yamashita:2023:SIJ

[1853] Michiaki Yamashita, Hiroyuki Yoshikawa, and Yoshihiro Shiraiwa. Special issue — JSMB Marine Biotechnology 2022 Conference. Marine Biotechnology, 25(5):643, October 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-023-10266-9.

Maeda:2023:DTS

[1854] Hiroto Maeda, Yuto Hirata, Hirokazu Takahashi, Kenshi Watanabe, Tsunehiro Aki, and Yoshiko Okamura. Development of a transformation system for Nitratireductor sp. Marine Biotechnology, 25(5): 644–651, October 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10198-4.

Hashimoto:2023:CMC

[1855] Kanako Hashimoto and Michiaki Yamashita. Comparison of muscle color and total selenium concentrations between spotted mackerel Scomber australasicus and Pacific mackerel S. japonicus. Marine Biotechnology, 25(5):652–656, October 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-023-10230-7.

Naser:2023:HEG

[1856] Insaf Naser, Yusuke Yabu, Yoshiaki Maeda, and Tsuyoshi Tanaka. Highly efficient genetic transformation methods for the marine oleaginous diatom Fistulifera solaris. Marine Biotechnology, 25(5):657–665, October 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/

s10126-022-10189-x. See correction [1930].

Yasukawa:2023:TDJ

[1857] Shino Yasukawa, Kyoko Shirai, Kaho Namigata, Masaaki Ito, Mei Tsubaki, Hikaru Oyama, Yukino Fujita, Taiki Okabe, Rei Suo, Shouzo Ogiso, Yukina Watabe, Hajime Matsubara, Nobuo Suzuki, Makoto Hirayama, Haruo Sugita, and Shiro Itoi. Tetrodotoxin detection in Japanese bivalves: Toxification status of scallops. Marine Biotechnology, 25(5): 666–676, October 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10199-3.

Matsumoto:2023:IMA

[1858] Eri Matsumoto, Shihomi Ito, and Tsutomu Nishimura. Identification of monomethylmonothioarsonic acid as the major thioarsenical generated during extraction processes for arsenic species analysis. Marine Biotechnology, 25(5):677–682, October 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-023-10200-z.

Shimokawa:2023:NSE

[1859] Yutaro Shimokawa, Kanako Abe, Mami Ohura, Manae Yamamoto, Hitoshi Ando, Takuma Tohfuku, Michiaki Yamashita, and Masakazu Kondo. Nutritional supplementation and enhanced antioxidant function by dietary intake of selenoneine and other selenium compounds in red seabream Pagrus major. Marine Biotechnology, 25(5):683–690, October 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-023-10215-6.

Fu:2023:CHH

[1860] Huiru Fu, Yongjing Li, Jing Tian, Ben Yang, Yin Li, Qi Li, and Shikai Liu. Contribution of HIF-1α to heat shock response by transcriptional regulation of HSF1/HSP70 signaling pathway in Pacific oyster, Crassostrea gigas. Marine Biotechnology, 25(5):691–700, October 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-023-10231-6.

Yu:2023:PID

[1861] Tieying Yu, Junhao Ning, Min Chen, Fukai Wang, Guilong Liu, Quanchao Wang, Xin Xu, Chunde Wang, and Xia Lu. Potential involvement of DNA methylation in hybrid sterility in hermaphroditic Argopecten scallops. Marine Biotechnology, 25(5):701–717, October 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-023-10233-4.

Maekawa:2023:SAS

[1862] Mari Maekawa, Emiri Yoshii, Yuri Akase, He Huang, Sota Yoshikawa, Masahiko Matsuda, Yosuke Kuruma, and Eitaro Sawayama. Sexassociated SNP confirmation of sex-reversed male farmed Japanese flounder Paralichthys olivaceus. Marine Biotechnology, 25(5):718–728, October 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-023-10235-2.

Mao:2023:GWA

[1863] Jinxia Mao, Jiali Lu, Sheng Liu, Youli Liu, Zhihua Lin, and Qinggang Xue. Genome-wide analysis of family I84 protease inhibitor genes in three bivalves reveals important information about the protein family’s evolution. Marine Biotechnology, 25(5):729–748, October 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-023-10236-1.

Houdelet:2023:CMI

[1864] Camille Houdelet, Eva Blondeau-Bidet, Mathilde Estevez-Villar, Xavier Mialhe, Sophie Hermet, François Ruelle, Gilbert Dutto, Aline Bajek, Julien Bobe, and Benjamin Geffroy. Circulating MicroRNAs indicative of sex and stress in the European seabass (Dicentrarchus labrax): Toward the identification of new biomarkers. Marine Biotechnology, 25(5): 749–762, October 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10237-0.

Zhang:2023:EMM

[1865] Jiahua Zhang, Jie Wang, Xiaodong Wang, Shikun Liu, Liang Zhou, and Xingguo Liu. Evaluation of microplastics and Microcystin–LR effect for Asian clams (Corbicula fluminea) by a metabolomics approach. Marine Biotechnology, 25(5):763–777, October 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-023-10238-z.

E:2023:AED

[1866] Zixuan E, Peng Luo, Chunhua Ren, Chuhang Cheng, Wenjie Pan, Xiao Jiang, Fajun Jiang, Bo Ma, Suzhong Yu, Xin Zhang, Ting Chen, and Chaoqun Hu. Applications of environmental DNA (eDNA) in monitoring the endangered status and evaluating the stock enhancement effect of tropical sea cucumber Holothuria scabra. Marine Biotechnology, 25(5): 778–789, October 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10239-y.

Wang:2023:SRM

[1867] Yayu Wang, Xiaoyue Zhu, Yang Gu, Zongyu Liu, Yingrui Mao, Xiaojun Liu, Zhiyi Bai, Guiling Wang, and Jiale Li. Study on the role of mitophagy receptor PHB2 in doubly uniparental inheritance of Hyriopsis cumingii. Marine Biotechnology, 25(5):790–799, October 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-023-10240-5.

Han:2023:GSM

[1868] Fei Han, Yuan Zhang, Anle Xu, Na Song, Geng Qin, Xiaoyan Wang, Siqing Chen, Li Bian, and Tianxiang Gao. Genomic structure and molecular characterization of toll-like receptors in black scraper Thamnaconus modestus and their expression response to two types of pathogens. Marine Biotechnology, 25(5):800–814, October 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-023-10241-4.

Song:2023:GCO

[1869] Na Song, Siyu Ma, Xiang Zhao, Jiabao Zhao, and Linlin Zhao. Genomic characteristics of Okamejei kenojei and the implications to its evolutionary biology study. Marine Biotechnology, 25(5):815–823, October 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-023-10242-3.

Chen:2023:IGR

[1870] Feng Chen, Wei Zhang, Xiaoyan Xu, Lang Gui, Yanfeng Lin, Minglin Wu, Jiale Li, and Yubang Shen. Identification of genes related to resistance to Ichthyophthirius multifiliis based on co-expression network analysis in grass carp. Marine Biotechnology, 25(5):824–836, October 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-023-10243-2.

Imaizumi:2023:ICS

[1871] Kentaro Imaizumi, Motohiko Sano, Hidehiro Kondo, and Ikuo Hirono. Insights into a chitin synthase of kuruma shrimp Penaeus japonicus and its role in peritrophic membrane and cuticle formation. Marine Biotechnology, 25(6):837–845, December 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-023-10244-1.

Hu:2023:CSI

[1872] Jiabao Hu, Man Zhang, Kaiheng Yan, Youyi Zhang, Yaya Li, Jiajie Zhu, Guanlin Wang, Xiangbing Wang, Yuanbo Li, Xiang Huang, Jie Tang, Rongyue Zheng, Shanliang Xu, Danli Wang, Yajun Wang, and Xiaojun Yan. Cold stress induces apoptosis in silver pomfret via DUSP–JNK pathway. Marine Biotechnology, 25(6):846–857, December 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-023-10245-0.

A:2023:TVA

[1873] Jose Priya T. A., Charutha Karunakaran, Aishwarya Nath, and Sudha Kappalli. Transcriptomic variation of Amphiprion Percula (Lacepède, 1802) in response to infection with Cryptocaryon irritans Brown, 1951. Marine Biotechnology, 25(6):858–890, December 2023. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-023-10246-z.

Yu:2023:WGR

[1874] Tieying Yu, Junhao Ning, Fukai Wang, Guilong Liu, Quanchao Wang, Xin Xu, Chunde Wang, and Xia Lu. Whole-genome re-sequencing and transcriptome reveal candidate genes and pathways associated with hybrid sterility in hermaphroditic Argopecten scallops. Marine Biotechnology, 25(6):891–906, December 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-023-10247-y.

Liu:2023:URG

[1875] Kai Liu, Nan Xie, Yuxi Wang, and Xinyi Liu. The utilization of referenceguided assembly and in silico libraries improves the draft genome of Clarias batrachus and Culter alburnus. Marine Biotechnology, 25(6): 907–917, December 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10248-x.

Malykin:2023:TLD

[1876] Grigorii V. Malykin, Peter V. Velansky, Daria I. Melnikova, and Timur Yu. Magarlamov. Tetrodotoxins in larval development of ribbon worm Cephalothrix cf. simula (Palaeonemertea, Nemertea). Marine Biotechnology, 25(6):918–934, December 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-023-10249-w.

Duan:2023:SOA

[1877] Yu Duan, Fuguo Liu, Chunyun Zhang, Yuanyuan Wang, and Guofu Chen. Screen and optimization of an aptamer for Alexandrium tamarense — a common toxin-producing harmful alga. Marine Biotechnology, 25(6): 935–950, December 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10251-2.

Wang:2023:TKM

[1878] Qingchun Wang, Yue Yan, Yifan Tao, Siqi Lu, Pao Xu, and Jun Qiang. Transcriptional knock-down of mstn encoding myostatin improves muscle quality of Nile tilapia (Oreochromis niloticus). Marine Biotechnology, 25(6):951–965, December 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-023-10252-1.

Zhu:2023:CCG

[1879] Xinhai Zhu, Lijie Qin, Yujie Zhu, Qieqi Qian, Xiaojian Gao, Qun Jiang, Jun Wang, Guoxing Liu, and Xiaojun Zhang. Characteristics and complete genome analysis of a pathogenic Aeromonas veronii SJ4 from diseased Siniperca chuatsi. Marine Biotechnology, 25(6):966– 982, December 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10253-0.

Sun:2023:AMM

[1880] Rong Sun, Shaokui Yi, Linlin Shi, Kianann Tan, Ruixue Shi, Siqi Yang, and Yanhe Li. Analysis of mRNA and MicroRNA expression profiles of nervous tissues and reproductive tissues in male Procambarus clarkii after silencing IAG. Marine Biotechnology, 25(6):983– 996, December 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10254-z.

Schwaner:2023:CRR

[1881] Caroline Schwaner, Sarah Farhat, Isabelle Boutet, Arnaud Tanguy, Michelle Barbosa, Denis Grouzdev, Emmanuelle Pales Espinosa, and Bassem Allam. Combination of RNAseq and RADseq to identify physiological and adaptive responses to acidification in the Eastern oyster (Crassostrea virginica). Marine Biotechnology, 25(6):997–1019, December 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-023-10255-y.

Seko:2023:DES

[1882] Takuya Seko, Yoko Sato, Michiko Kuniyoshi, Yuko Murata, Kenji Ishihara, Yumiko Yamashita, Sanjuro Fujiwara, Tomohiro Ueda, and Michiaki Yamashita. Distribution and effects of selenoneine by ingestion of extract from mackerel processing residue in mice. Marine Biotechnology, 25(6):1020–1030, December 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-023-10256-x.

Gong:2023:CTW

[1883] Jianwen Gong and Qi Li. Comparative transcriptome and WGCNA analysis reveal molecular responses to salinity change in larvae of the Iwagaki oyster Crassostrea nippona. Marine Biotechnology, 25(6):1031– 1042, December 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10257-w.

Li:2023:MHT

[1884] Jiahua Li and Xiaodong Zheng. Morphology, histology, and transcriptome analysis of gonadal development in Octopus minor (Sasaki, 1920). Marine Biotechnology, 25(6):1043–1056, December 2023. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-023-10258-9.

Kumar:2023:MTS

[1885] Pokanti Vinay Kumar, Kiran D. Rasal, Arpit Acharya, Diganta Dey, Arvind A. Sonwane, Dhalongsaih Reang, R. Rajeshkannan, Sachin S. Pawar, Nitin P. Kurade, Mukesh P. Bhendarkar, Kishore K. Krishnani, Naresh S. Nagpure, and Manoj P. Brahmane. Muscle transcriptome sequencing revealed thermal stress-responsive regulatory genes in farmed rohu, Labeo rohita (Hamilton, 1822). Marine Biotechnology, 25(6):1057– 1075, December 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10259-8.

Bessaid:2023:GRS

[1886] Mariem Bessaid, Jun Soung Kwak, and Ki Hong Kim. Generation of recombinant snakehead rhabdovirus (SHRV) expressing artificial MicroRNA targeting spring viremia of carp virus (SVCV) P gene and In Vivo therapeutic use against SVCV infection. Marine Biotechnology, 25(6):1076–1084, December 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-023-10260-1.

Hu:2023:CRN

[1887] Jiabao Hu, Youyi Zhang, Kaiheng Yan, Man Zhang, Yaya Li, Yuanbo Li, Weiwei Gu, Shanliang Xu, Yajun Wang, and Xiaojun Yan. Change and regulation of nutritional metabolism in silver pomfret during compensatory growth. Marine Biotechnology, 25(6):1085–1098, December 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-023-10261-0.

Ossa-Hernandez:2023:RFP

[1888] Natalia Ossa-Hernández, Luis Fernando Marins, Rodrigo Volcan Almeida, and Daniela Volcan Almeida. Red fluorescent protein variant with a dual-peak emission of fluorescence. Marine Biotechnology, 25(6): 1099–1109, December 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10262-z.

Huang:2023:MTA

[1889] Jiasheng Huang, Zixuan E, Wenjie Pan, Zhi Li, Tiehao Lin, Chunhua Ren, Peng Luo, Bo Ma, Yang Liu, Xugan Wu, Chaoqun Hu, Xiao Jiang, and Ting Chen. Metabolome and transcriptome association analysis reveals the link between pigmentation and nutrition utilization in the juveniles of sea cucumber Holothuria leucospilota. Marine Biotechnology, 25(6):1110–1122, December 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-023-10263-y.

Rasal:2023:DMP

[1890] Kiran D. Rasal, Sujata Mohapatra, Pokanti Vinay Kumar, Shasti Risha K, Prachi Asgolkar, Arpit Acharya, Diganta Dey, Siba Shinde, Manohar Vasam, Rajesh Kumar, and Jitendra Kumar Sundaray. DNA methylation profiling of ovarian tissue of climbing perch (Anabas testudienus) in response to monocrotophos exposure. Marine Biotechnology, 25(6): 1123–1135, December 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10264-x.

Xu:2023:CAT

[1891] Yuanfeng Xu, Yongkai Tang, Wenrong Feng, Yanan Yang, and Zhaoxia Cui. Comparative analysis of transposable elements reveals the diversity of transposable elements in decapoda and their effects on genomic evolution. Marine Biotechnology, 25(6):1136–1146, December 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-023-10265-w.

Dai:2023:ASS

[1892] Wenfang Dai, Zijuan Zhang, Yinghui Dong, Lin He, Qinggang Xue, and Zhihua Lin. Acute salinity stress disrupts gut microbiota homeostasis and reduces network connectivity and cooperation in razor clam Sinonovacula constricta. Marine Biotechnology, 25(6):1147–1157, December 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-023-10267-8.

Cahyani:2023:PPC

[1893] Ni Kadek Dita Cahyani, Noer Kasanah, Dewi Sri Kurnia, and Mark T. Hamann. Profiling prokaryotic communities and aaptamines of sponge Aaptos suberitoides from Tulamben, Bali. Marine Biotechnology, 25(6): 1158–1175, December 2023. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10268-7.

Tan:2023:SSF

[1894] Kianann Tan, Yaxin Dong, Karsoon Tan, Leong-Seng Lim, Khor Waiho, Jing Chen, Peng Xu, and Kit Yue Kwan. siRNA silencing of FpVtg induces ovarian cell apoptosis in redtail prawn, Fenneropenaeus penicillatus. Marine Biotechnology, 25(6):1176–1190, December 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-023-10269-6.

Zuo:2023:PTA

[1895] Anli Zuo, Yonghua Zhou, Yuxian Chen, Sipu Liu, Yuyao Lu, Yingjie Li, Shenping Cao, and Zhen Liu. Physiological and transcriptome analysis reveal the regulation mechanism underlying the muscle quality effect of dietary Schisandra chinensis in triploid crucian carp (Carassius auratus). Marine Biotechnology, 25(6):1191–1207, December 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-023-10270-z.

Okada:2023:NTM

[1896] Koki Okada, Yu Morimoto, Yukine Shiraishi, Takashi Tamura, Shigeki Mayama, Takashi Kadono, Masao Adachi, Kentaro Ifuku, and Michiko Nemoto. Nuclear transformation of the marine pennate diatom Nitzschia sp. strain NIES-4635 by multi-pulse electroporation. Marine Biotechnology, 25(6):1208–1219, December 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-023-10273-w.

Yang:2023:CRA

[1897] Qiqi Yang, Lyana Salim, Chuan Yan, and Zhiyuan Gong. Correction: Rapid analysis of effects of environmental toxicants on tumorigenesis and inflammation using a transgenic zebrafish model for liver cancer. Marine Biotechnology, 25(6):1220–1221, December 2023. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-023-10250-3. See [1515].

Dominguez-Maqueda:2024:EDP

[1898] Marta Domínguez-Maqueda, Jorge García-Márquez, Silvana T. Tapia-Paniagua, Carmen González-Fernández, Alberto Cuesta, Cristóbal Espinosa-Ruíz, María Ángeles Esteban, Francisco Javier Alarcón, María Carmen Balebona, and Miguel Ángel Moriñigo. Evaluation of the differential postbiotic potential of Shewanella putrefaciens Pdp11 cultured in several growing conditions. Marine Biotechnology, 26(1): 1–18, February 2024. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10271-y.

Ponce:2024:ABI

[1899] Marian Ponce, Eugenia Zuasti, Victoria Anguís, and Catalina Fernández-Díaz. Anti-bacterial and immunostimulatory properties of ulvan-loaded chitosan nanoparticles for use in aquaculture. Marine Biotechnology, 26 (1):19–27, February 2024. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10272-x.

Zhong:2024:EBF

[1900] Guo-Fang Zhong, Liang-Fa Zhang, Yi Zhuang, Qi Li, He Huang, Cong Cao, Zhan-Ying Zhu, Zhong-Yuan Huang, Nu-An Wang, and Kun Yuan. Effects of brown fishmeal on growth performance, digestibility, and lipid metabolism of the Chinese soft-shelled turtle (Pelodiscus sinensis). Marine Biotechnology, 26(1):28–36, February 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-023-10274-9.

Jin:2024:NKT

[1901] Can Jin, Kang Cheng, Rui Jiang, Yihang Zhang, and Wen Luo. A novel Kunitz-type serine protease inhibitor (HcKuSPI) is involved in antibacterial defense in innate immunity and participates in shell formation of Hyriopsis cumingii. Marine Biotechnology, 26(1):37–49, February 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-023-10275-8.

Lee:2024:CMA

[1902] Euihyeon Lee, Kyun-Woo Lee, Yeun Park, Ayeon Choi, Kae Kyoung Kwon, and Hye-Min Kang. Comparative microbiome analysis of Artemia spp. and potential role of microbiota in cyst hatching. Marine Biotechnology, 26(1):50–59, February 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-023-10276-7.

Wang:2024:SSL

[1903] Wenhao Wang, Junrou Huang, Yan Hu, Jianxiang Feng, Dong Gao, Wenyu Fang, Meng Xu, Chunlei Ma, Zhenqiang Fu, Qinglong Chen, Xuanguang Liang, and Jianguo Lu. Seascapes shaped the local adaptation and population structure of South China Coast yellowfin seabream (Acanthopagrus latus). Marine Biotechnology, 26(1):60–73, February 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-023-10277-6.

Cai:2024:IDD

[1904] Yuyang Cai, Li He, Shenping Cao, Peng Zeng, Linhan Xu, Yanan Luo, Xiang Tang, Qixiang Wang, Zhen Liu, Zhimin He, and Suchun Liu. Insights into dietary different co-forms of lysine and glutamate on growth performance, muscle development, antioxidation and related gene expressions in juvenile grass carp (Ctenopharyngodon idellus). Marine Biotechnology, 26(1):74–91, February 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-023-10278-5.

Zhang:2024:SAI

[1905] Saisai Zhang, Shuang Liu, Hongwei Liu, Hui Li, Jun Luo, Aili Zhang, Yinpeng Ding, Tongjun Ren, and Wenbo Chen. Stochastic assembly increases the complexity and stability of shrimp gut microbiota during aquaculture progression. Marine Biotechnology, 26(1):92–102, February 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-023-10279-4.

Shi:2024:ASR

[1906] Xiang Shi, Ruiqi Zhang, Zhe Liu, Guiyan Zhao, Jintao Guo, Xue Mao, and Baoyi Fan. Alternative splicing reveals acute stress response of Litopenaeus vannamei at high alkalinity. Marine Biotechnology, 26(1): 103–115, February 2024. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10281-w.

Zhou:2024:RVD

[1907] Shun Zhou, Zongrui Yang, Baoshan Guo, Jingyuan Yi, Yunfei Pang, Ruixin Feng, Jiaxue Song, and Yunji Xiu. Rapid visual detection of Spiroplasma eriocheiris by loop-mediated isothermal amplification with hydroxynaphthol blue dye. Marine Biotechnology, 26(1):116– 124, February 2024. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-023-10282-9.

Yang:2024:CTA

[1908] Qiong Yang, Hong Yu, and Qi Li. Comparative transcriptome analysis reveals the role of ribosome reduction in impeding oogenesis in female triploid Crassostrea gigas. Marine Biotechnology, 26(1):125– 135, February 2024. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-024-10283-2.

Reyes-Becerril:2024:BIR

[1909] Martha Reyes-Becerril and Tania Zenteno-Savin. Bisphenol A induces reactive oxygen species production and apoptosis-related gene expression in Pacific red snapper Lutjanus peru leukocytes. Marine Biotechnology, 26 (1):136–148, February 2024. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-024-10284-1.

Acquafredda:2024:TRA

[1910] Michael Acquafredda, Ximing Guo, and Daphne Munroe. Transcriptomic response of the Atlantic surfclam (Spisula solidissima) to acute heat stress. Marine Biotechnology, 26(1):149–168, February 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-024-10285-0.

Yu:2024:CLA

[1911] Hua Yu, Zhao xia Zou, Wei Wei, and Ying Li. Conjugated linoleic acid reduces lipid accumulation via down-regulation expression of lipogenic genes and up-regulation of apoptotic genes in grass carp (Ctenopharyngodon idella) adipocyte in vitro. Marine Biotechnology, 26(1):169– 180, February 2024. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-024-10286-z.

eReis:2024:FCS

[1912] Maíra Barbosa e Reis, Açucena Imparato Maximo, Jessica Maria Magno, Daniel de Lima Bellan, João Luiz Aldinucci Buzzo, Fernanda Fogagnoli Simas, Hugo Alexandre Oliveira Rocha, Edvaldo da Silva Trindade, and Carolina Camargo de Oliveira. A fucose-containing sulfated polysaccharide from Spatoglossum schröederi potentially targets tumor growth rather than cytotoxicity: Distinguishing action on human melanoma cell lines. Marine Biotechnology, 26(1):181–198, February 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-024-10287-y.

Colihueque:2024:AES

[1913] Nelson Colihueque and Margarita Parraguez. Assessing the effectiveness of sex-linked molecular markers to identify neomale breeders for the production of all-female progenies of rainbow trout. Marine Biotechnology, 26(1):199–204, February 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-024-10288-x.

Li:2024:AGM

[1914] Ju Li, Xiaohong Li, Simiao Fu, Yuxuan Meng, Xiaoyan Lv, Xin Zhang, Guozheng Liu, and Jinsheng Sun. Adaptation of glucose metabolism to limb autotomy and regeneration in the Chinese mitten crab. Marine Biotechnology, 26(1):205–213, February 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-024-10290-3.

Yang:2024:DTS

[1915] Chenhu Yang, Yanzhen Du, Xiaoqi Zeng, and Gang Ni. Development and testing of species-specific primers for detecting the presence of the Northern Pacific sea star (Asterias amurensis) from environmental DNA. Marine Biotechnology, 26(2):215–222, April 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-024-10292-1.

Sasaki:2024:AML

[1916] Shudai Sasaki, Tsukasa Mori, Hirofumi Enomoto, Sakiko Nakamura, Hideo Yokota, Hiroshi Yamashita, and Naoko Goto-Inoue. Assessing molecular localization of symbiont microalgae in coral branches through mass spectrometry imaging. Marine Biotechnology, 26(2): 223–229, April 2024. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-024-10294-z.

Yao:2024:OAA

[1917] Ting Yao, Fengjie Sun, Bingkui Zhu, Subing Han, Hao Zhang, Chunxiao Meng, Zhengquan Gao, and Yulin Cui. Oral administration of antimicrobial peptide NZ2114 through the microalgal bait Tetraselmis subcordiformis (Wille) butcher for improving the immunity and gut health in turbot (Scophthalmus maximus L.). Marine Biotechnology, 26(2): 230–242, April 2024. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-024-10289-w.

Casuso:2024:SBT

[1918] Antonio Casuso, Bárbara P. Benavente, Yeny Leal, Crisleri Carrera-Naipil, Valentina Valenzuela-Muñoz, and Cristian Gallardo-Escárate. Sex-biased transcription expression of vitellogenins reveals fusion gene and MicroRNA regulation in the sea louse Caligus rogercresseyi. Marine Biotechnology, 26(2):243–260, April 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-024-10291-2.

Campos-Sanchez:2024:PHI

[1919] Jose Carlos Campos-Sánchez, Jhon A. Serna-Duque, Carmen Alburquerque, Francisco A. Guardiola, and María Ángeles Esteban. Participation of hepcidins in the inflammatory response triggered by λcarrageenin in gilthead seabream (Sparus aurata). Marine Biotechnology, 26(2):261–275, April 2024. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-024-10293-0.

Sales:2024:ASD

[1920] Abdias Fernando Simon Sales, Karolyne dos Santos Jorge Sousa, Amanda de Souza, Mirian Bonifacio, Tiago Akira Tashiro Araújo, Matheus de Almeida Cruz, Márcia Busanello Costa, Daniel Araki Ribeiro, Lívia Assis, Cintia Cristina Santi Martignago, and Ana Cláudia Rennó. Association of a skin dressing made with the organic part of marine sponges and photobiomodulation on the wound healing in an animal model. Marine Biotechnology, 26(2):276–287, April 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-024-10295-y.

Xia:2024:DMA

[1921] Yu qing Xia, Yi Yang, Yan yun Liu, Jian xin Cheng, Ying Liu, Cheng hua Li, and Peng fei Liu. DNA methylation analysis reveals potential mechanism in Takifugu rubripes against Cryptocaryon irritans infection. Marine Biotechnology, 26(2):288–305, April 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-024-10296-x.

Chen:2024:TAH

[1922] Minxia Chen, Qijuan Wan, Ming Xu, Zihao Chen, and Songlin Guo. Transcriptome analysis of host anti- Vibrio harveyi infection revealed the pathogenicity of V. harveyi to American eel (Anguilla rostrata). Marine Biotechnology, 26(2):306–323, April 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-024-10298-9.

Massocato:2024:CBA

[1923] Thaís Fávero Massocato, Víctor Robles-Carnero, Bruna Rodrigues Moreira, Pablo Castro-Varela, José Bonomi-Barufi, Roberto Abdala-Díaz, Leonardo Rubi Rörig, and Félix Diego López Figueroa. Characterization and biological activities of the ulvan polysaccharide-rich fraction obtained from Ulva rigida and Ulva pseudorotundata and their potential for pharmaceutical application. Marine Biotechnology, 26(2): 324–337, April 2024. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-024-10299-8.

Yuliani:2024:VAR

[1924] Dewi Yuliani, Fumihiro Morishita, Takuya Imamura, and Tatsuya Ueki. Vanadium accumulation and reduction by vanadium-accumulating bacteria isolated from the intestinal contents of Ciona robusta. Marine Biotechnology, 26(2):338–350, April 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-024-10300-4.

Chen:2024:CTA

[1925] Shujian Chen, Ce Shi, Yangfang Ye, Ronghua Li, Weiwei Song, Changbin Song, Changkao Mu, Zhiming Ren, and Chunlin Wang. Comparative transcriptome analysis reveals the light spectra affect the growth and molting of Scylla paramamosain by changing the chitin metabolism. Marine Biotechnology, 26(2):351–363, April 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-024-10301-3.

Jiang:2024:TPA

[1926] Kunyin Jiang, Chengxun Xu, Hong Yu, Lingfeng Kong, Shikai Liu, and Qi Li. Transcriptomic and physiological analysis reveal melanin synthesis-related genes and pathways in Pacific oysters (Crassostrea gigas). Marine Biotechnology, 26(2):364–379, April 2024. CODEN MAB-IFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://

link.springer.com/article/10.1007/s10126-024-10302-2.

Zeng:2024:SPL

[1927] Junjia Zeng, Wei Liu, Yacheng Deng, Pengxin Jiang, Zhijun Wang, Yanhong Ou, Hongtao Lu, Yuanjingxi Hui, Hongli Xu, and Peng Xu. Swimming performance in large yellow croaker: Effects of group size, test protocol, and recovery time on critical swimming speed. Marine Biotechnology, 26(2):380–388, April 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-024-10303-1.

Tian:2024:CMO

[1928] Jing Tian, Hongxia Wang, Pin Huan, Xin Yue, and Baozhong Liu. Comprehensive multi-omics approaches provide insights to summer mortality in the clam Meretrix petechialis. Marine Biotechnology, 26(2): 389–403, April 2024. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-024-10304-0.

Kanwate:2024:PAA

[1929] Balaji Wamanrao Kanwate, Kalpana Patel, Sandesh Suresh Karkal, Deependra Rajoriya, Kunal Sharan, and Tanaji G. Kudre. Production of antioxidant, angiotensin-converting enzyme inhibitory and osteogenic gelatin hydrolysate from Labeo rohita swim bladder. Marine Biotechnology, 26(2):404–420, April 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-024-10305-z.

Naser:2024:CHE

[1930] Insaf Naser, Yusuke Yabu, Yoshiaki Maeda, and Tsuyoshi Tanaka. Correction: Highly efficient genetic transformation methods for the marine oleaginous diatom Fistulifera solaris. Marine Biotechnology, 26 (2):421–422, April 2024. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-024-10306-y.

El-Zaeem:2024:NIE

[1931] Samy Y. El-Zaeem, Amr El-Hanafy, Alaa A. El-Dahhar, Ayaat M. Elmaghraby, Sara F. Ghanem, and Amany M. Hendy. A novel investigation for early sex determination in alive adult European seabass (Dicentrarchus labrax) using cyp19a1a, dmrt1a, and dmrt1b genes expression in tail fin tissues. Marine Biotechnology, 26(3):423–431, June 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-024-10313-z.

Xu:2024:LMP

[1932] Qiaoyue Xu, Hongtao Nie, Qianying Ma, Jiadi Wang, Zhongming Huo, and Xiwu Yan. The lgi-miR-2d is potentially involved in shell melanin synthesis by targeting mitf in Manila clam Ruditapes philippinarum. Marine Biotechnology, 26(3):432–446, June 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-024-10307-x.

Rodriguez-Mejia:2024:NTA

[1933] José Luis Rodríguez-Mejía, Itzel Anahí Hidalgo-Manzano, Luis Felipe Muriel-Millán, Nancy Rivera-Gomez, Diana X. Sahonero-Canavesi, Edmundo Castillo, and Liliana Pardo-López. A novel thermo-alkaline stable GDSL/SGNH esterase with broad substrate specificity from a deep-sea Pseudomonas sp. Marine Biotechnology, 26(3):447–459, June 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-024-10308-w.

Jiang:2024:CMC

[1934] Kunyin Jiang, Hong Yu, Lingfeng Kong, Shikai Liu, and Qi Li. cAMPmediated CREM-MITF-TYR axis regulates melanin synthesis in Pacific oysters. Marine Biotechnology, 26(3):460–474, June 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-024-10309-9.

Yang:2024:DRD

[1935] Chao Yang, Yu Zhen, Jialin Hou, and Tiezhu Mi. Development of a rapid detection method to Prorocentrum lima by loop-mediated isothermal amplification with hydroxy naphthol blue. Marine Biotechnology, 26(3):475–487, June 2024. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-024-10310-2.

Li:2024:GAM

[1936] Xue Li, Miao Yang, Kunlian Mo, Yonghua Hu, Hanjie Gu, Dongmei Sun, Shixiang Bao, and Huiqin Huang. Genome analysis of multiple polysaccharide-degrading bacterium Microbulbifer thermotolerans HB226069: Determination of alginate lyase activity. Marine Biotechnology, 26(3):488–499, June 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-024-10311-1.

Ueda:2024:JPF

[1937] Hiroyuki Ueda, Masaaki Ito, Ryo Yonezawa, Kentaro Hayashi, Taiga Tomonou, Maho Kashitani, Hikaru Oyama, Kyoko Shirai, Rei Suo, Kazutoshi Yoshitake, Shigeharu Kinoshita, Shuichi Asakawa, and Shiro Itoi. Japanese planocerid flatworms: Difference in composition of tetrodotoxin and its analogs and the effects of ingestion by toxin-bearing fishes in the Ryukyu Islands, Japan. Marine Biotechnology, 26(3): 500–510, June 2024. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-024-10312-0.

Chen:2024:TTP

[1938] Zheyan Chen, Shouhao Zhu, Bingbing Feng, Min Zhang, Jinhua Gong, Huangen Chen, Brian Pelekelo Munganga, Xianji Tao, and Jianbin Feng. Temporal transcriptomic profiling reveals dynamic changes in gene expression of giant freshwater prawn upon acute saline-alkaline stresses. Marine Biotechnology, 26(3):511–525, June 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-024-10314-y.

Xie:2024:HTI

[1939] Xiaobin Xie, Yibo Wang, Fang Ma, Ruilin Ma, Leqiang Du, and Xin Chen. High-temperature-induced differential expression of miRNA mediates liver inflammatory response in Tsinling Lenok trout (Brachymystax lenok tsinlingensis). Marine Biotechnology, 26(3):526–538, June 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-024-10315-x.

Nahle:2024:OMC

[1940] Sarah Nahle, Camille Lutet-Toti, Yuto Namikawa, Marie-Hélène Piet, Alice Brion, Sylvie Peyroche, Michio Suzuki, Frédéric Marin, and Marthe Rousseau. Organic matrices of calcium carbonate biominerals improve osteoblastic mineralization. Marine Biotechnology, 26(3):

539–549, June 2024. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-024-10316-w.

Kim:2024:DPB

[1941] Suhyeok Kim, Jaebeom Shin, Nalin Medagoda, Sera Choi, So Yun Park, Jeung-Yil Park, and Kyeong-Jun Lee. Dietary poly-β-hydroxybutyrate improved the growth, non-specific immunity, digestive enzyme activity, intestinal morphology, phagocytic activity, and disease resistance against Vibrio parahaemolyticus of Pacific white shrimp, Penaeus vannamei. Marine Biotechnology, 26(3):550–561, June 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-024-10317-9.

Tiong:2024:HKB

[1942] Irene K. R. Tiong, Cher Chien Lau, Patrick Sorgeloos, Mimi Iryani Mat Taib, Tengku Sifzizul Tengku Muhammad, Muhd Danish-Daniel, Min Pau Tan, Liying Sui, Min Wang, and Yeong Yik Sung. Hsp70 knockdown in the brine shrimp Artemia franciscana: Implication on reproduction, immune response and embryonic cuticular structure. Marine Biotechnology, 26(3):562–574, June 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-024-10318-8.

Tan:2024:TTS

[1943] Kianann Tan, Xiaowan Ma, Boyu Su, Chen Zhan, Xin Yang, Khor Waiho, Leong-Seng Lim, and Kit Yue Kwan. Targeting ttvgr via siRNA knockdown elicits ovarian cell death in the tri-spine horseshoe crab. Marine Biotechnology, 26(3):575–587, June 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-024-10319-7.

Ye:2024:EVB

[1944] Xiaokang Ye, Jiali Lin, Qiuji Chen, Jiehuan Lv, Chunsheng Liu, Yuping Wang, Shuqi Wang, Xiaobo Wen, and Fan Lin. An efficient vector-based CRISPR/Cas9 system in zebrafish cell line. Marine Biotechnology, 26 (3):588–598, June 2024. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-024-10320-0.

Kim:2024:ABM

[1945] Ju-Won Kim, Julan Kim, Ja Young Cho, Younhee Shin, Hyojung Son, Subramaniyam Sathiyamoorthy, Bo-Seong Kim, Young-Ok Kim, Byeong chul Kang, and Hee Jeong Kong. Association between muscle growth and transcription of a mutant MSTN gene in olive flounder (Paralichthys olivaceus). Marine Biotechnology, 26(3):599–608, June 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-024-10322-y.

Zhang:2024:TAD

[1946] Jiahua Zhang, Jie Wang, Zhaojun Gu, and Xingguo Liu. Transcriptome analysis of different aquaculture substrates on the immune response of Babylonia areolata. Marine Biotechnology, 26(3):609–622, June 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-024-10324-w.

Ye:2024:EDM

[1947] Yucong Ye, Jiarong Huang, Siwen Li, Yiming Li, and Yunlong Zhao. Effects of dietary melatonin on antioxidant capacity, immune defense, and intestinal microbiota in red swamp crayfish (Procambarus clarkii). Marine Biotechnology, 26(4):623–638, August 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-024-10326-8.

Zhang:2024:ARS

[1948] Yunsheng Zhang, Hu Xia, Wei Peng, Lanhai Liu, Liangguo Liu, and Pinhong Yang. Application of repetitive sequences in fish cell depletion as a target for the CRISPR/Cas9 system. Marine Biotechnology, 26(4): 639–648, August 2024. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-024-10328-6.

Yonezawa:2024:TLT

[1949] Ryo Yonezawa, Kentaro Hayashi, Hikaru Oyama, Kazutoshi Yoshitake, Soshi Sato, Jayan Duminda M. Senevirathna, Ashley R. Smith, Taiki Okabe, Rei Suo, Shigeharu Kinoshita, Tomohiro Takatani, Osamu Arakawa, Shuichi Asakawa, and Shiro Itoi. Tissue localization of tetrodotoxin in the flatworm Planocera multitentaculata (Platyhelminthes: Polycladida). Marine Biotechnology, 26(4):649–657, August 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-024-10332-w.

Yamamoto:2024:APF

[1950] Juna Yamamoto, Hana Deguchi, Takechiyo Sumiyoshi, Kentaro Nakagami, Akatsuki Saito, Hiroshi Miyanishi, Masakazu Kondo, Tomoya Kono, Masahiro Sakai, Masato Kinoshita, and Jun ichi Hikima. Accumulation and phagocytosis of fluorescently visualized macrophages against Edwardsiella piscicida infection in established mpeg1.1-transgenic Japanese medaka Oryzias latipes. Marine Biotechnology, 26(4):658–671, August 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-024-10333-9.

Gao:2024:HTI

[1951] Xinming Gao, Yaoping Lv, Qingmin Dai, Ling Zhu, Siqi Liu, Zehui Hu, Junkai Lu, Haidong Zhou, and Zufei Mei. Histological and transcriptomic insights into the ovary development of Hemibarbus labeo injected with spawn-inducing hormones. Marine Biotechnology, 26(4): 672–686, August 2024. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-024-10335-7.

Ma:2024:DDI

[1952] Yuting Ma, Yongshuang Xiao, Zhizhong Xiao, and Jun Li. Development of DNA insertion-specific markers based on the intergenic region of Oplegnathus punctatus Cdkn1 /srsf3 for sex identification. Marine Biotechnology, 26(4):687–695, August 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-024-10336-6.

P:2024:NBD

[1953] Athira P. P., Anju M. V., Anooja V. V., Archana K., Neelima S., Muhammed Musthafa S., I. S. Bright Singh, and Rosamma Philip. A novel beta-defensin isoform from Malabar trevally, Carangoides malabaricus (Bloch & Schneider, 1801), an arsenal against fish bacterial pathogens: Molecular characterization, recombinant production, and mechanism of action. Marine Biotechnology, 26(4):696–715, August 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-024-10338-4.

Jin:2024:HNU

[1954] Can Jin, Fangmengjie Wei, Jiayi Zhang, Xiaoyang Tan, Taixia Fan, Wen Luo, and Jiale Li. HcN57, a novel unusual acidic silk-like matrix protein from Hyriopsis cumingii, participates in framework construction and nacre nucleation during nacreous layer formation. Marine Biotechnology, 26(4):716–731, August 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.springer.

com/article/10.1007/s10126-024-10339-3.

Ke:2024:GSL

[1955] Qiaozhen Ke, Jiaxing Liu, Ji Zhao, Jiaying Wang, Pengxin Jiang, Yacheng Deng, Xiaoying Zhou, Junjia Zeng, Tao Zhou, and Peng Xu. Genomic selection of large yellow croaker (Larimichthys crocea) with a high plant protein diet enhances the growth performance of offspring. Marine Biotechnology, 26(4):732–740, August 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-024-10341-9.

Li:2024:EGA

[1956] Xin Li, Qiaozhen Ke, Ang Qu, Jiaying Wang, Ji Zhao, Peng Xu, and Tao Zhou. Effects of gene alternative splicing events on resistance to cryptocaryonosis of large yellow croaker (Larimichthys crocea). Marine Biotechnology, 26(4):741–753, August 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-024-10342-8.

Ulloa:2024:ISN

[1957] Pilar E. Ulloa, Felipe Jilberto, Natalia Lam, Gonzalo Rincón, Luis Valenzuela, Valentina Cordova-Alarcón, Adrián J. Hernández, Patricio Dantagnan, Maria Cristina Ravanal, Sebastian Elgueta, and Cristian Araneda. Identification of single-nucleotide polymorphisms in differentially expressed genes favoring soybean meal tolerance in highergrowth zebrafish (Danio rerio). Marine Biotechnology, 26(4):754–765, August 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-024-10343-7.

Richman:2024:ASP

[1958] Jasmine Richman and Michael Phelps. Activin signaling pathway specialization during embryonic and skeletal muscle development in rainbow trout (Oncorhynchus mykiss). Marine Biotechnology, 26(4):766– 775, August 2024. CODEN MABIFW. ISSN 1436-2228 (print),

1436-2236 (electronic). URL https://link.springer.com/article/
10.1007/s10126-024-10345-5.

Risha:2024:DMP

[1959] K. Shasti Risha, Kiran D. Rasal, Dhalongsaih Reang, Mir Asif Iquebal, Arvind Sonwane, Manoj Brahmane, Aparna Chaudhari, and Naresh Nagpure. DNA methylation profiling in genetically selected Clarias magur (Hamilton, 1822) provides insights into the epigenetic regulation of growth and development. Marine Biotechnology, 26(4):776–789, August 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-024-10346-4.

Wang:2024:MAE

[1960] Fei Wang, Zi-Rou Zhong, Qing Xie, Jie Ou, Ning-Xia Xiong, Ming-Zhu Huang, Shi-Yun Li, Gang Hu, Zi-Le Qin, and Sheng-Wei Luo. Multiomics analyses explore the immunometabolic interplay in the liver of white crucian carp (Carassius cuvieri) after Aeromonas veronii challenge. Marine Biotechnology, 26(4):790–809, August 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https://link.

springer.com/article/10.1007/s10126-024-10347-3.

Gao:2024:EHR

[1961] Changsheng Gao and Hongtao Nie. Exploring the heat-responsive miRNAs and their target gene regulation in Ruditapes philippinarum under acute heat stress. Marine Biotechnology, 26(4):810–826, August 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-024-10348-2.

Li:2024:OGD

[1962] Xinchun Li, Yitian Bai, Chengxun Xu, Shikai Liu, Hong Yu, Lingfeng Kong, Shaojun Du, and Qi Li. OysterDB: A genome database for Ostreidae. Marine Biotechnology, 26(4):827–834, August 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236 (electronic). URL https:

//link.springer.com/article/10.1007/s10126-024-10327-7.

Ortega-Kindica:2024:SMR

[1963] Rose Chinly Mae H. Ortega-Kindica, Chinee S. Padasas-Adalla, Sharon Rose M. Tabugo, Joey Genevieve T. Martinez, Olive A. Amparado, Carlo Stephen O. Moneva, Rodelyn Dalayap, Carlos O. Lomeli-Ortega, and Jose Luis Balcazar. Shotgun metagenomics reveals taxonomic and functional patterns of the microbiome associated with Barbour’s seahorse (Hippocampus barbouri). Marine Biotechnology, 26(4):835–841, August 2024. CODEN MABIFW. ISSN 1436-2228 (print), 1436-2236

(electronic). URL https://link.springer.com/article/10.1007/
s10126-024-10330-y.